A Complete Bibliography of Publications in the Journal of the American Society for Information Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.04

Title word cross-reference

$1 [Har77a]. $1.50 [Har78a]. $10.00 [Har70b]. $10.50 [New73]. $12.00 [Sel71, Woo77b, Lan78, Wei79]. $12.50 [Ger79, Har78b]. $12.95 [Tsu79]. $13.50 [Shu78, Str78]. $13.95 [Die81]. $14.00 [Wel80b]. $15.00 [Sel71, Lan73]. $15.00/year [Lil78]. $16.95 [Ere81b]. $17.50 [Ste77]. $17.95 [Die82b]. $18.95 [Nov82]. $19.25 [Lun80]. $19.75 [Her82]. $19.95 [Fis81b]. $20.00 [Ere80]. $21.50 [Kra79]. $21.95 [Lan82a, Vag80]. $24.95 [Shi85]. $25.00 [Kun82]. $26.75 [Kaz79]. $27.50 [Mat85]. $29.95 [Pat85]. $3.00 [Ger79]. $32.50 [Bla79]. $33 [Rei85]. $34.50 [Mat85]. $35 [Ano74a]. $36.00/yr [Var78]. $39.95 [Mar85]. $4.00 [Lil78]. $4.95 [Fis79]. $43.50 [Mig79]. $5.00 [Rus85]. $5.95 [Boy85b, Mim85, Har70c]. $50 [Usd85]. $60 [Rei85]. $65 [Hei78]. $70 [Dro78]. $8.00 [Shu80]. 80/20 [Egg93]. $9.50 [Ols74]. $9.95 [Mar80]. ^{1} [TEC99]. ^{3} [CT87]. K [Jr88]. M [Jr88]. μ [Egg93]. n [Cle76]. θ [Jr88].

-dimensional [Cle76]. -Legal [Ere80]. -rule [Egg93].

0 [Fis79, Mig79, Boy85b, Fis81a, Shu78]. 0-12-118660-1 [Wel80b]. 0-12-785006-6 [Mar77b]. 0-13-527713-2 [Fis81a]. 0-13-585331-1 [Fis81b]. 0-208-01546-9 [Lan78]. 0-262-03056-X [Lan77a]. 0-387-06335-8 [Lev76a]. 0-440-54412-2 [Boy85b]. 0-442-80036-3 [Vag80]. 0-471-03205-0 [Fis79]. 0-8108-0899-4 [Gro77]. 0-8108-0949-4 [Har77a]. 0-8135-0807-X [Gro75b]. 0-837-1-8587-4 [Har77c]. 0-8371-9485-7 [Shu78]. 0-85142-106-7 [Mig79]. 0-87815-010-2 [Lar74]. 0-87845-0424 [Har76a]. 0-87845-043-2 [Swa77b]. 0-87845-044-0 [Ric77]. 0-88405-284-2 [Dav77b]. 0-89391-010-4 [Mig80a]. 0-910972-43-5 [Har76c]. 0.00 [Har77a]. 0024-2394 [Lil78]. 0163 [Mig80a]. 0163-5689 [Mig80a]. 0470-27501-4 [Mar85].

1 [OG70, Wel71, Wri71a]. 1.0 [Sie89a]. 1/71 [Wel71, Wri71a]. 1/71-1R [Wel71]. 1/71-2R [Wri71a]. 1/2 [Rei85]. 1/2- [Rei85]. 1/Number [McM76]. 105MM [Kuc76a]. 15th [Bro95a]. 17th [Kil93b]. 1917 [How89c]. 1940s [Shi86a]. 1945 [Tay95]. 1967 [Hin70]. 1969- [Har71b]. 1970 [Ste77]. 1970s [Bor99]. 1971 [Swa76c]. 1972 [JJ75, Ric77]. 1973 [Ano74a, Bor74, Har76b, Kol74, Art74, Har75c]. 1974 [Har77b]. 1975 [Kuc76a]. 1976 [Gro77]. 1977 [Wei79, Koc79, Tes79]. 1978 [Die82a]. 1979 [Lan80, Mea82b]. 1980’s [Ric89]. 1981 [Mea82a]. 1983 [Ber84, Dan84, Joh86]. 1983-Monthly [Bol86b]. 1985 [Woo90]. 1986 [Ble86, Les89]. 1988 [Old91]. 1989 [Buc91b]. 1994 [ATJ95, Hen96]. 1995 [Cra96, CML^{+}96, Deb96, Hay96a]. 1996 [Wei97a]. 1997 [BI98]. 1R [Wel70, Wel71, WH70].

2 [Fis81a, SBW93, Boy85b, Fis81a]. 2-3 [Wes00]. 2-7177-1206-2 [Gro76a]. 2001 [Wes00]. 2003 [Bro97]. 20036 [Rus85]. 20th [Boy91, Ano84j]. 21st [Fro94, Lan94, Nan95b]. 24-X [Har77b]. 245 [Ano08]. 2d [Wei76]. 2nd [Bea97, Bla79, Cob96, Col98, Dom78, Gro87, Har83, Her82, How87, Jen95a, Pen93, Pet96a, Sel71, Tom98b, Wei95b, Zwe95]. 2R [Hin70, Wri71a]. 2R. [Har70a].

3-7940-44 [Har77b]. 3-7940-4422-3 [Har76b]. 37th [Swa76c]. 38 [Aiy88]. 3R [Til70]. 3R. [Sie70]. 3rd [Lan78, O’C95, Wei94b].

4 [Don70, Hin70, Pre70, Til70, WH70, Nov82]. 4/70 [Don70, Hin70, Pre70, Til70, WH70]. 4/70-1R [WH70]. 4/70-2R [Hin70]. 4/70-3R [Til70]. 4/70-4R [Pre70]. 4/70-5R [Don70]. 4/1973 [McM76]. 40th [Har77b]. 4R [Pre70]. 4th [Swa76a, Ten94].

5 [Har70a, Sie70, Wel70]. 5/70 [Har70a, Sie70, Wel70]. 5/70-1R [Wel70].

5/70-2R. [Har70a]. 5/70-3R. [Sie70]. 50th [Bat99a]. 51 [Ano08]. 54-62638 [Lil78]. 5689 [Mig80a]. 5R [Don70]. 5th [Nel97a].

6th [Kri87].

7 [Dav76a, Mig79, Shu78]. 70-1R [Wel70, WH70]. 70-2R [Hin70]. 70-2R. [Har70a]. 70-3R [Til70]. 70-3R. [Sie70]. 70-4R [Pre70]. 70-5R [Don70]. 71-1R [Wel71]. 71-2R [Wri71a]. 73 [Lar74]. 73-77342 [Lar74]. 74 [Wel78]. 75-5345 [Har77c]. 76-1790 [Swa77b]. 76-236331 [Ric77]. 76-25422 [Har77a]. 77342 [Lar74]. 78-8166 [Vag80].

80/20 [Egg87]. 86 [Kri87].

90-247-2860-6 [Wil85b]. 92-3-001232-7 [Gro76b]. 99/4A [Mim85].

= [Gro76b].

A. [Deb92, Koe94]. AACR [Gro90]. AACR2 [HO89]. Abbate [Mal00]. Abbreviation [Low71]. abbreviations [Bai76, BB77]. ABCs [Adl97]. Abiteboul [Rap00]. ABLEX [Mig80a]. ABN [Gul85]. Aboutness [BSW00]. absence [Bai74, San74, Swa89c]. Abstract [DK75, Win95, Har78b]. Abstracting [Lan80, MUD99, Boy92, Coh95, Exn97, Flo72, FG70, LN82, LWZ99, Mai99, Mol95, Pri83, Ror93, RSZ71, Bea97, Bus95, Wei91]. abstracting/indexing [FG70]. abstraction [Wei96a]. Abstracts [Bol86b, Cra00, Byr75, Kin76, LBKO87, MRY73, RSZ71, SM91, Smi74b, Tho73a, Wei70, Flo99a, SA91, Sma80, Wei97c, Bol86a]. academe [Whi87]. Academic [ABB^{+}91a, ABB^{+}91b, BD88, Bas74, BS91b, BDBR97, BF91a, BD97, BDSA97, BFM88, CS91, DRCG91, D’E97, FW88, Goo97, GLS97, Har75d, KMB99, LS97, LD88, LH97, LB88, LD91, Mar77b, McC78, MT96, Mcd82, Pee96a, PW88, Pre70, RRB91, RH97, SA97, SB91, Swa77c, War88, Wel80b, Wes88, WMME88, ZvBR88, Afs98, BCG79, BM72, Bra77, Bud91, Che72, CB83a, CO94, Dep75, Eva94, Fra99, Gol94, Hei94, KRD95, Lap00, Lev99, McG73, McG96b, Mil90, Pou89, Ric87, Und98a, Voo99, Whi94, Fu96a, Fu96b]. academics [Sch94]. Accelerating [KT90]. acceptable [Kra94a]. acceptance [HJ89]. Access [EN85, JB85, Lap99, Mar94, Sve94, AM77, Ath72, Bat86, Bat98, BB96a, BC00, BB95, Boy86g, Boy86f, Bus95, Cha99a, CWBL00, Coo70b, DS91, Dra89, DW96, EH94, EQH00, Fre82, GGJ91, Gir91, Gor79, Har74, Haw90, Her96b, Kuc75, Mac82, MCBC89, OOA99, SW92, Sou96, SM90, Tur98, War81, WSQ94, Wil77a, Wol71, Yee91, PW92, Ber94, Lan77b, Hei94, Car95b]. accessibility [Cul85]. accessing [Lar90]. accommodate [SD97]. Accommodation [RG70]. Accounting [CMRWW93, Bri91, Mos70, Rep89, Pec85]. accreditation [Hay88, MF88]. accrediting [ERV00]. accuracy [MV95]. Aceto [McM76]. Achievements [WT93]. acknowledgement [Van97a]. Acknowledgments [DC93, CMRWW93]. Ackoff [Woo77b]. acoustics [Caw78]. acquired [BB95, Whi76]. acquisition [Coo91c, Erc92, HK80, OG93, Ozo79b, RM97, Rou74b]. Acquisitions [MBFK76, War79, Dep75, DCD81, WSS79]. acronyms [Bai76, BB77]. across [BFM88, Los97b]. act [Lev77, Gro96]. Acta [Lif93]. acted [Pap98]. Action [BHCC^{+}74, AdKK87, Kra92a, RCJS94, Kuc76b]. activation [CN95]. Active [Flo99b]. Activities [Pre71b, Swa77b, Ker77, Ric88, SP73, Wol87]. Activity [Col99, BW85, CN81, Che98a, Ere85, Spa99]. activity-theoretical [Che98a]. acts [Bro85, WT95]. Actuality [Avr79]. Ad [And78]. Adams [Sta75a]. Adaptation [Die82a, Mea83c, PM00]. adaptations [SFB^{+}00]. Adaptive [Str82, Ber94, MBVL99, WYSB91, Aba86, HC94]. Added [Tay82]. Addison [Har79]. Addison-Wesley [Har79]. additional [Kra89a]. addressing [BYN00, CNMS97]. adequacy [CN81]. Adjusted [Asa81a, Rou92a]. Administration [War79, Hal85, Mea79a, Mea83a, Pra83, SS93, WWSR00, Whi89a]. Administrative [Rus70a]. Admissibility [Dav84, Mea83d, Mea84c]. adolescent [Jul99, Tod99]. adopting [BSR93]. adoption [Klo89, Lon90, Sch96b, Wil92b]. Advanced [DC98, Hei78, Mac79a, RH99, SFV85, Swa78b, DH97a, Fuh00, Lar92b, LD88, And91, Bea91, Bes91, Buc91c, LL91, Lyn91a, Nug91, Rom91, Ban98]. advancement [Sch99]. Advances [Fau96, Hol96, Pre70, Div87, Exn98, Mar77b, Swa77c]. advantage [Hub98, Kan78, de 76, de 78]. advertisements [DH97b]. advisors [Gre92]. Advisory [Har70c]. aerial [MM98]. Aeronautical [Joh84]. affairs [Gar86]. affect [PJ90]. affecting [Bat77, FS83]. Affective [NT96, Bil00]. Affects [Wal72]. affirmative [Tal74]. Africa [Aiy88, Old87, Pou89, Tia89, MT96, Pee96a]. African [SR95]. after [Bla96, Mea94]. aftermath [Lev77]. afterword [Don77]. again [Spr99, Kra91b]. Against [Pom74, GD90]. Age [Art77, Alg95, Asa81a, BW85, Bro96, Bro99c, BMFK76, Egg97, GM86, Gup90, Kin98b, Mar94, Par82b, San99, Whi00, Bor93, Gro88, Mal99, Mat97b]. agencies [Kan94]. Agency [Ger79]. agenda [Fro94]. agent [Hur88, LR79, MBVL99]. agents [HG98, LH76, Mac96, Whi98b]. aggregate [Ren80]. Aging [ER00a, ER00c, Hal85, Xia92]. Agre [Lam99]. agree [Los00]. agreement [Col95]. agricultural [LM81, Lon90]. AGRIS [LM78]. Aid [HM87, Wel71, Bau72, Cra80a, LDS^{+}00, Zai77]. aided [Luc93, SS76]. AIDS [BB95, JJ75, RL76, Sew82]. Air [HK79]. Aires [Jac71]. Aiyepeku [Old88]. al. [Wal00a]. ALA [Bel81, Wel83, Usd85]. Alan [Dan84]. Alayamma [Bai75]. Albany [War79]. Albert [Afi99, Gro76a]. alchemy [Bac93]. alcoholism [Cha76]. alert [Kra91a]. alerting [Bot70]. Alexandria [Har77c, HCL^{+}00]. Alfred [Don74]. algebra [Dio83, Mat00]. Algorithm [FS91a, AYS96, Cod73, Gor91, Har75b, Kje74, MBVL99, Str74, Wri72, Yu74]. algorithmic [CN95, GS93a]. Algorithms [Wil00c, BS92, Bur95, CO84, CO85, Che95a, CSSI98, CYW99, Hul96, Lov71, Pai96, Ro88a, Ro88b]. aliasing [Bro93b]. Aligning [Pal99a]. All-Union [Sch73b]. Allan [Kaz77]. Allen [Coc98, Hir90]. Allerton [Har76a, Ric77]. alleviate [CA80, vB81]. Alleviating [CMK^{+}98a]. allied [Tay95]. allied-German [Tay95]. allocate [Smi81]. allocating [WA79]. Allocation [Boo74a, Mcg79, SR79, Buc79, Rou75]. almanac [How89a]. along [Cas98a]. alpha [Was74]. alpha-numeric [Was74]. alteration [He98, He00]. alternative [McL79, San98b, SCLL73]. alternatives [Boy86g, Ler79, Boy86e]. Alumni [Ano84m]. Alvin [Shu80]. ambiguity [Boo90b, ME98, Boo97, Bon89]. AMC [Rou89b]. America [Hin70, ASR85, Blu71, Ere81b, WWB97, Zur84, Pic86]. American [Ano84k, Ano84l, Ano85g, Bat99a, Boy85a, Boy91, Buc99, Gro96, Har70c, How89c, Kra87c, Kra99c, Kuc76a, Mea82b, Ols74, Pec85, Sel71, Smi99, Usd85, And94, Ano72a, Bar96, Cox92, DS79, Ell70, HHC70, Hum70, Joh84, KS70, Kra87a, Oro77, Sta77b, Eli87, Ste77]. Amey [Koc79]. Among [Pre71a, WFK73, And99, Cam96b, Cra71b, Cra72, Fen81, LBBR95, LS99, Mar94, Par82a, Rob80a, Wil98a, Yer83, YG90]. Amorphous [Caw74]. amounts [RSW91]. Amsterdam [Kaz79]. analogies [Col96]. analogues [Ham00]. analogy [LD85]. analyses [HF00, IC97, SCD98, WH91, Mar80]. Analysis [Ano70a, BSM90, COC90, DF82, LR72, LW90, McC90, McL90b, Pai90, SG89, Sta77a, Wel80a, Wri71a, AKS96, Ami70, BI00, Beg95, Ber72a, BR92, Boy89b, Boy90a, BMvR91a, BMvR91b, CO85, Cha81, Che86b, CMK^{+}98a, CPR84, Coo91b, CMK98b, CBM^{+}94, Cox83, Cri81, DDF^{+}90, Del90, DW98, Die79, Don72, DOO^{+}96, DW96, EK75, EP91, Fay73, Gar97, Goe87, Gol74a, GWDS86, HG00, Hay95, He00, HMNW86, HA95, Hub98, Jah74, JF87, KB98a, KB98b, Kan76b, KDS89, Kie98, KW97, Koe99, Koe79, Koh74, Kuc82, LJ70a, Lan71, Leh72, Lif93, Liu97c, LG75, MM89, MM97, MB86, MB81a, MCT73, Mar83c, MRY73, MB81b, McC91, McG73, McG96b, MPW73, MN83b, Nee81, Nis99, Nom84, NML99, Pal99a, Pao78, Per92, RW86, Ras90, Rob91, RW00, SKL97, SYY75, Sal79a, San83]. analysis [SP73, Sch00a, SKvB92, SvB94, Sch90, Sha92, Sha83b, She91, Sie90, Sny98, Sol93, Ste74a, Ste74b, Sum84, Sut94, Tan99, Tho90, Tod99, VG93, War93, Wei96a, Wes87, WM98, WY73, Woo70, Yar97, Yer83, YFR81a, YFR81b, YFR81c, YF85, BIFW71, Che97, Per82, Cal89a, Sch87]. analyst [Sie89a]. analysts [BR97]. analytic [Boy90b, Los96c]. Analytical [LC71, Mac72, Nan70, AP96b]. analyzed [AP97]. analyzer [CT73]. analyzing [Tan96]. anaphora [LBKO87]. and/of [Gli72]. Anderson [Ere81a, Kuc76b]. Andrew [Lan77b, Whi00]. Andrews [Mim85]. Angeles [Ano74b, Gol74b, Lea75]. angle [Coo82, ZK99a, ZK99b]. animation [CD97, LBBR95, LBBR96]. Ann [Pre71b]. Anne [Rei77]. annealing [BMSM90, CSSI98]. Anniversary [Bat99a, Kra87a]. annotated [FK99]. annotator [Wis00]. Announcement [Ano82a, Wes00]. Annual [Ano84j, Bor74, Boy91, McM76, Rei85, Bol86b, Har71b, Kra97, Ano74f, Hei88]. Anomalous [BN76]. anonymous [Kol75]. ANS [Cha77]. ANSI [Bra79]. Answer [DO81, O’C70, O’C80, O’C73]. Answer-passage [DO81, O’C80]. answer-passages [O’C73]. Answer-providing [O’C70]. answer-sentences [O’C73]. answering [Roy90a]. Antecedents [Don77, AL84]. Anthony [Die81, Har76b, Hei78]. anti [JD78]. anti-verb [JD78]. antibiotic [Qin00b]. anticipating [Deb86]. Antiquarian [Shi86a]. Anzlowar [Sel71]. Apendices [Har70a]. appendices [New73]. Appendix [Art75a, Vag80]. Appilcation [Ama84]. Apple [Mim85]. applicability [Ro88a, Ro88b, Cha91]. applicable [AN73]. Application [BMSM90, Blo70, DS84, Fed82a, Fed82b, GSKL00, Joh84, KW97, Lin70, MB91, Miy98, Rei84, SC76, SR79, Sum84, Asa81a, HS78, KH80, LMSC82, Mos70, RSZ71, Sch97b, Swa77a, Swa78a, Wan92, Wel83, Wis00, SP73]. Applications [And91, Bea91, Bes91, Buc91c, Egg90b, LL91, Lyn91a, McG84, Nug91, Rom91, Swa77b, Wil95a, BS92, Bro86a, Fuh00, HJD^{+}99, Lee96b, LS84b, Lyn91b, Rou98, Swa75b, Swa76b, Und98b, Bar96, HW71, Lyn98, Ere81a, Cal00, Lan77a, Rou76a, Bar98b, Bir97, SS95]. Applied [Hei78, Rob91, Swa78b, Coo89, EQH00, JB71a, Kra78, LJ70b, LL85, MN83b, SC85a, SC85b, Tur77]. Applying [CV80, Sta75b, DW98, RT81, AKS96]. appraisal [CA80]. Approach [Dav76a, Kra79, AF98, Asa81b, BS91b, Boo80, BP93, CN95, CNMS97, Che98a, CSSI98, CWBL00, CPR84, CT87, Dun96b, DS84, Erc92, Ere81c, Fag89, FBC79, FBK97, Goe87, GV87, GFHR97, GR74, Har75a, Har75b, HPH96, Jär89, KdK90, Kra73, KB78, Lan78, Lee96b, LK97, LS71, LFHG99, MR87, MBVL99, Mat82, McC95, Mer71, OCN97, Rad82a, Reg88, Sac76, Sav97b, SWB79, TO84, Tzv90, Voi97, Wis99, Yoo94, Zha98, Ano75a, HK79, Sta77a, Sta82, Par97]. Approaches [KR80, LC84, Bro87, Cha99a, Che99, SCLL73, SS95, YLYY00, HF00, Rep89]. appropriate [Sch73a]. approximate [BYN00]. April [Hay96a, Les89]. Arabic [ASAOE99, AKE94, Ama84, HAA87, HKE97, Mus86]. Arabic/ [Mus86]. Arbeiten [Wel71]. archeological [Tyn75]. Archetypes [Smi97]. Architecture [RNBV98, Wei95b, Zoe88, Jon86]. architectures [Bir97]. archival [AD79, Cox92, Cox94]. archives [Con93, Cox93, DKP00, Rou89b, Rou89a, Swa76a, Cox93]. archivists [Cox88]. Archon [Wel70]. area [PR98, Ren80, SP73, Sma80, Tom98b]. areas [RBR80, Smi74b, Blu89b]. arena [Bra89]. argument [BS91b]. Arise [Kra91a]. arising [HG00]. ARIST [Hei88, Hen00a]. armes [Kra91a]. Army [Pre71b]. arrangement [FS84]. Array [CPW88, MR87]. Arrive [Bra79]. Art [Pre71b, Wei95b, Bea97, Cim95, Cob96, Kos94, Mad96a, Mad96b, Mil00, Pen70, Wit81, Wei95b]. Artandi [Bas74]. Arthur [Che99]. Article [Ano78c, Art75a, Lan89b, Can95, Coo91a, Coo76, Cou98, Dic96, Die84b, He98, Her78, Lev98, Seg94, Wei90, WH91]. Articles [Jor89, WFK73, And78, Ano78c, Cam96a, EK75, Gup90, HST96, HNW93, Kim00, Kuc78, LWZ99, May78, Now81, Sch94, Wyl87]. Artificial [CM84, LS84b, Smi84b, BF91b, Cob96, Jon91, Kas96a, Kas96b, Mac96, Wal81, vdBL96]. arts [Bes91, SvB94]. Ascertaining [SP73]. Ashley [Fis79]. Asian [Joh86]. ASIS [Kra87a, Ano72c, Blo70, Bor73b, Eli87, HM70, Lei73, Lin70, Mat70, Mos72, OG70, Var99, Wel80a]. ASIS/ [Ano72c]. ASLIB [Lan81b, Mig79]. Aspects [Lip99, Rus70a, BMvR91a, BMvR91b, ER98, FB88, Kin98a, Mat82, MZ79, Pra83, Rob00, Tal79, Tor89, Tra98, Kat00]. assess [HN97, Kat72]. assessed [Die79]. Assessing [Bam79, Deb86, LM78, RR77, Koe83, LB83, SWT^{+}79]. assessment [Bom75, BCG79, HJD00, Kos94, LB83, LS97, MAN80, MMB99, Whi96]. assessments [DRCG91, Har96, War91, PLM00]. Assets [Han75, Hei94, Han75]. assist [SG89]. Assistance [Ste77, Cra00, Dos85, SBW90]. assistant [Wis00]. assisted [BS88, BBMS87, BBSM87, HHC70, HSP75, Mer71, SHS72, Jer70]. associated [Hoc89]. Association [Boy85a, How89c, Inh76, Kuc76a, Lev76b, Pec85, Pre71c, Sel71, Usd85, Coo71, PN98, Ano84j, Hor78, SD96]. association-based [PN98]. Associations [Ble86, Swa76c, CMK^{+}98a, FS97b, MB86]. Associative [Mar82a, Cag70c, FWR95, Sac76]. assuming [Har71d]. assumptions [CM84, Col81, Der94, Eve93, EP97, Gle93]. Assured [Wri71b]. Astronomers [Bro99c]. astronomy [MP74, PLM00]. ATA [Sel71]. Atherton [Hir89, Vea90]. Atomic [Art74]. Attending [CCN93]. Attitudes [Kuc75, AM77, Chr72, Kat72, KM73, MCT73, MWY95, Sch94]. attorney [Sut94]. attribute [SLC98]. attributes [SMC83]. attribution [BR95b, Rou92a]. Attunement [CD97]. Atypical [Hei86a]. audiences [Pap88]. audio [Wue97]. audiographic [Sla73]. auditing [CO94, Whi94]. Audrey [ATJ95]. augmentation [Bam79]. augmented [MS76]. August [CML^{+}96, Har76c, Wei97a]. aural [Sve94]. Australia [Gro75a, Gul85, Var75a, Tod99]. Australian [Gul85, Ten88, Gro75a]. Author [BSM90, COC90, Din87, Hin70, Law82, LW90, MM88, McC90, Pai90, Rei77, WG81, WC85, Bla71, BM81, Car96, Coo70b, DG91, Egg94, McC84, McC86, McC88, MZ80, Mea84c, Ozo80, SJ72, SR77, Whi81, Whi91, WM98, Lan81a]. authoring [JM97]. Authoritative [Sch98a]. authority [Boy86a, FPS00, Ros76, Ber84, Boy86a]. Authors [Dan00, BK79, Cam96b, ERV00, HG00, Hai97, Kuc78, McC90, Pap93, Van97b, WM00b, Ols74]. authorship [BR95b, Lip99, SR95, Wal97]. Authorships [Dav92b, Aji91, Aji92]. autobiographical [Gar76]. Automated [Bur80, Rod95, WBX98, Cor94a, DH70, MQP98, WRE^{+}70, WP94, Aus93, Boy85a, Liu98, Wei93a]. Automatic [Bel71, CYFS95, Cra82, Dil82, DPP78, Fed82b, Hum99, Lar92a, May97, NKM77, Pao78, RL76, RSZ71, Sal70, Sal79a, SBF83, Sch90, SM90, Zam80, BS74, BCM^{+}70, BM75, Boy90a, CT73, CN95, Coh95, DC80, DG83, Fag89, FS91a, GS93b, HST96, Haa97, HZ80, Har75a, Har75b, Har74, HKE97, JB71b, JGR90, Lan98a, Lan98b, Lar92c, LK97, LH95a, MRY73, MRW73, PN98, RC95, RW82, Sal72c, SYY75, SWY81, Sho71, Ste74a, VS87, WEWP97, Wei96a, WT95, CMK^{+}98a, GS93c]. automatically [Ror93]. Automating [Sta98]. Automation [Bel81, Don70, Lev76b, Rus77, Col83, Coo77a, Cor94b, CB83b, Fis70, Her97, LC83, Mar83b, Mic83, O’C95, Tor89, Wal86c, Zvi87]. autopoietic [Bra94]. Autor [Har70a]. Aux [Kra91a]. Availability [Don76, Gar91, Kan76b, Pit75, Ras90, Rou92b]. Available [Don70, Hoc89, LWG74, Rus85]. Avenue [Rus85]. average [YK93]. Award [Coc98, Mel72, Smi74b, Str74]. Awareness [ASS73, Ami70, BWV72, DH70, FRGL75, Lev76b, MS76, MP74, PS79]. axiom [Fug85]. Aylesbury [Mea82b].

B [Har76c, Rod93, Sta75a]. B. [TS92]. B.C. [Kil93b]. Back [DG91, Dio94, Spr99]. Background [SKCT88]. backgrounds [JJ75]. backlog [Kra89b, Kra90]. Backpropagation [Bir97]. Baden [GSK98]. Baden-Wuerttemberg [GSK98]. balance [Shi88]. balances [TvR90]. ball [Jac91]. Baltimore [BR94]. bandwagon [Key96]. Bane [Mim85]. bank [Mor78]. Banking [Nov82]. Barbara [Kaz75b, Lev76a]. barefoot [Wil94]. bargaining [Har76a]. barricades [Kra91a]. barrier [Sch74]. Barriers [Jul99]. Barter [Teb70]. Barwise [Lan80]. Base [Kra79, Wil75, Zam80, BW84, Haw74, HO89, JJ77, PW88, Per94, Rum90, VA73, WL80, War73, Wil77a]. Based [Swa76c, Ami70, Ang75, Ano00a, Arm94, BWV72, Bie95, Bil00, BR80a, Car75, CYW99, Coo98, CC92, Cra86, CO94, DKD71, DG83, DW95, Dog98, Eft00, Egg94, EW70, Fid84, For91, Gal91, Gor91, GL96, GD98, HF00, Hei74, Her99, HOS^{+}93, HM87, JN93, KSDR90, KPZ^{+}99, Kwo85, LUH^{+}96, Len83, LG99, Los92, MUD99, Nan95a, OOA99, Pai96, Pai71, Pao78, Par94, PN98, Rob80b, RNBV98, Sal84, Sav93, ST89, Sha91b, SD96, SFB^{+}00, SGS70, Tyn75, Und98b, Whi94, Yao95, Yu74, VC70, VTS97, Art75b, Boy86b, Smi00a, Wil95a]. bases [BM72, Bru78, Fre80, Mar72a, RF97, SWS71, Sch76, SF78, Wes89, Wil77b, Pet96b, Pet96c]. bases-history [Bru78]. Basic [Bro00d, CM84, FB78, LMM^{+}73, Leh72, Mon81, YFR81a, Fis99]. Basics [Par83a, Mea83b]. Basing [Kri89]. Basis [MA76, Ano08, JGR90, NK00, RR71, Wil00a]. batch [AM77, Fay73]. Be [Ano74c, Ano75b, Fag95, Hal81, HW99, Her78, Kra97, OMY^{+}89, OR78, Swa89c]. Bearing [DF82, BKR98]. Beaudiquez [Wel78]. Becker [Ano74b, Art74, Hay96a, New73]. becoming [Toc70]. been [Dos88, OV93]. begins [Mas97]. behavior [BR97, BHWG95, Bro99c, Bur93, Can89, CH00, Cha81, Cha91, Che95b, CCT95, CC98, COC90, Dal90, DC98, Dre99, Fid84, FDD^{+}99, Kan87, Kay98, KRC81, Mig80b, Moo72, NT96, Nah98, OV92, OV95, Pra83, Ras80, SC85a, ST86, SBW93, SJ00, Sol93, Sol97a, Sol97b, Sol97c, Spi96, Str82, SMC83, Yua97, Ros73]. Behavioral [McC89, GWDS86, NZ91]. behaviors [Bil00, Par00, Sch00a, Kri89]. behind [Rol87]. Beilby [War79]. Belgium [GBDH76]. Belief [dSM93, KdK90]. Belver [WM00a]. Belzer [Deb96]. Benefit [MBFK76, Hum00, Qin00a, Sny98]. benefits [Bor73b, BM75, LM78].

Benejam [Die84a]. Bequai [Lan82a]. Bernadine [Lar74]. Bernal [SS70]. Bernard [Ano74a, Tes79, Dav94]. Bernier [Wel80b]. Bertie [TS92]. besieged [Smi98b]. Best [Wei79, BK98, Bro00f]. Beta [Har75c, GS87]. beta-binomial [GS87]. Bethesda [Cer83]. Better [Mec99, Die84a, SF78, Die84a]. Between [Bon82, BR80b, BG94, BLPG99, CL86, DC98, Egg93, FS97b, GL96, HNW93, He00, Her78, Hub78, Hur80, Laz00, LBW00, Lin87b, MB81a, Mar92b, Mea00, NL83, ODVG87, Pic96, Qin00b, Raf75, Rou90, Sal72c, Sch74, Seg94, Sma73, SR77, SSD99, Tag72, TvR90, Wal86d, YAH90, YR77]. Beverly [Ste77]. Beyond [Key96, San98b, Shi85]. biases [Bla80]. biblio [Lun76]. biblio-profile [Lun76]. Bibliocryptography [Boo96]. Bibliographic [Ath72, CPW88, Dod79, Elc73, Fed82a, Fed82b, Gul85, Haw74, HD72, LBP76, Mor76, Rus77, AL84, BE80, BM72, Boy89a, Bro83, BC99, Coo77b, DW95, Elc78, FS83, GBDH76, GWDS86, KRC81, Kuc76b, MCT73, MK72, Mou77, OMY^{+}89, Ras80, Sch76, Sch81, Sho71, SL99, Smi96a, Smi96b, TF89, Wal86a, Wel80c, Wel75, Wol96, WD97b, YG90, Wel78]. Bibliographical [Hin70, Bro88]. bibliographically [Swa87]. bibliographies [Bro90b, Cor82, Egg90b]. bibliographique [Gro76a]. Bibliography [Lan81b, Nov82, Ray97, Rei77, WWB97, FM97, Fre77, Har75c, New73, RM99, SJ72, Smi84b, FF80, Mar75, Wei96a]. Bibliography/International [Ray97]. bibliogrpahic [Yer83]. Bibliometric [Ano75a, Can99, DCFQ00, KB98b, Koe83, LL85, Nic89, Par82a, Par82b, Par82c, VHdB^{+}98, BR92, DC93, DC98, Don72, Egg87, Gri88, Haw77, IC97, MP98, NML99, OV97, Pie92, Rou90, SP73, Sic85, Sic92, Tag81, de 76]. Bibliometrics [Bro90a, Bro87, KET99, Lan97, McC97, Pra77, Rob91, Sha92, McL90a, McC96]. Bibliothekar [Wel71]. Bibliothekar-Lehrinstitut [Wel71]. Bibliotheken [Wel71]. Bibliotheque [Gro76a]. Bigelow [New76]. bilingual [Mus86]. Bill [Mim85, Pat85, Wei79, Woo87a]. binary [Gre90, Har71d, Jan91, WY90a]. binational [War93]. binding [TO84]. Bing [Ere80]. Bingley [Tes79, Wel70]. binomial [GS87]. bio [Har75c]. bio-bibliography [Har75c]. biochemistry [Mar83c]. Biographic [Ols74]. Biographical [Van99, WS71]. Biological [Ano72a, Flo99a]. Biology [Gan74, Fei73, Hai97, Mar92b, Ano72a]. biomedical [Fay72a, Fay72b, Fay73, Ste90, SR95, TB93, NPG76]. BIOSIS [Ols74, LBP76, WFK73, WFK72]. Biotechnology [McC95]. BIRS [DCFQ00]. birth [Adl92]. Bit [EQH00, CYW99]. bits [Whi99b]. bitsy [CCRY98]. bivariate [BS91a]. Black [Pom74]. BLEND [Sha83a]. Blending [Glu95b]. blind [Mea84b]. Block [BYN00]. Blois [Bal89]. Blomquist [Kuc75]. Blum [Hin70]. Board [Lan80]. Bob [Kat00, Kra99a]. Bohnert [Sch91a]. Bommer [Rei75]. bones [Gro96]. Book [Aba86, Abe96b, Afi98, Afi99, Afs98, Afs99, Agh86, Aik97, Alg95, All98, And99, And88, And90, Ano74a, Ano74b, Ano75a, Ano77a, Art74, Art75b, Art76a, Art77, Aul87, Aus93, Bai75, Bak89, Ban98, BI98, Bar98b, Bas74, Bat97, Bat93, Bea97, Bea93, Beh96a, Beh96b, Bel81, Bel76, Ben86, Ber86a, Ber84, Bie95, Bir97, Bix96a, Bix96b, Bla79, Bla93, Ble86, Blu89a, Blu89b, Bol86a, Bol86b, Bol92, Bon89, Boo00, Bor86a, Bor94, Bor00, Bor90, Bor93, Bos87, Boy85a, Boy85b, Boy86g, Boy86d, Boy86e, Boy86b, Boy86a, Boy86f, Boy86c, Boy87b, Boy87a, Boy88, Boy89a, Boy89b, Boy90b, Boy90a, Boy91, Boy92, Bro86a, Bro93a, Bro94, Bro95a, Bro96, Bro98a, Bro98b, Bro00a, Bro00d, Bro00b, Bro00c, Bro97, Buc91b, Buc98, Bud91, Bue86a, Bue86b, Bus95, Cal89b, Cal89a]. Book [Cal00, Cam97, Car95a, Car95b, Car88, Cas92, Cas95, Cas96, Cas97, Cas98a, Cas98b, Cer83, Che86b, Che98a, Che99, Che76, Chi93, Cla75, Cob96, Col96, Col98, Con93, Cor94a, Cor94b, Cox93, Cox94, Cox98, Cox00, Cre95b, Cre95a, Cri81, Cul88, Cur97, Dan84, Dav76a, Dav77b, Dav79b, Dav92a, Dav96b, Dav98, Day99b, Deb86, DK90, DC76, Die81, Die82a, Die82b, Die84a, Die87, Dil97a, Dil98, Div87, Dog98, Dol98, Dom78, Don70, Don74, Don76, Don77, Dra89, Dro77, Dro78, Dro98, Dun96b, Dur76, Erc95, Ere80, Ere81a, Ere81b, Eva94, Exn97, Exn98, Fis99, Fis79, Fis81a, Fis81b, Fra86, Fro96, Fu96a, Fu96b, Gan74, Gar86, Gel96, Ger79, Glu95a, Gol74b, Goo99, Gre96, Gro75a, Gro75b, Gro76a, Gro76b, Gro76c, Gro77, Gro87, Gro88, Gro90, Gul85, Haa96, Han96, Har70a]. Book [Har87, Har88, Har93a, Har83, Har70d, Har70c, Har70b, Har71b, Har78a, Har79, Har75d, Har75c, Har76c, Har76b, Har76a, Har77a, Har77b, Har78b, Har77c, Has98, Has95a, Has95b, Haw90, Hay95, Haz98, Hei78, Hei86b, Hei86a, Hei87, Hei90, Hei95d, Hei94, Hen95, Her82, Her97, Hil96a, Hil96b, HC94, Hin70, Hjø99, Hjø00a, Hoc89, Hol99, How87, How89c, How89a, How89b, Hun96a, Hun96b, Inh76, Jac96, Jaf98, Jas97, Jen95a, Jen95b, Joh86, Jon86, Kan94, Kas96a, Kas96b, Kat00, Kaz75b, Kaz75a, Kaz77, Kaz79, Key96, Key97, Kie98, Kin98b, Koc79, Koe75, Kol75, Kra79, Kri86a, Kri86b, Kri87, Kri86c, Kri89, Kuc75, Kuc76b, Kuc76a, Kun82, Lam00a, Lam99, Lam00b, Lan77a, Lan78, Lan80, Lan82a, Lan87a, Lan87b, Lan89a, Lan90, Lan93, Lan73, Lan76, Lan77b]. Book [Lan81b, Lap99, Lap00, Lar74, Lar90, Lea98, Lea75, Lea76, Lee96a, Lee96b, Les89, Les96, Lev76a, Lev76b, Lew00, Lid96, Lie96a, Lie96b, Lil78, Lit96, Liu97a, Liu98, Los96a, Los96b, Lun80, Lyn90, Mac96, Mad96a, Mad96b, Mai99, Mai00, Mai90, Mal99, Mal00, Mar80, Mar95, Mar85, Mar75, Mar77b, Mar96, Mat85, Mat97b, Mat97a, McC96, MT96, McG95, McG96a, McK96a, McK96b, McL90a, McL90b, McM76, Mea82b, Mea82a, Mea83b, Mig79, Mig80a, Mil93, Mil90, Mim85, Miz98, Moo86, Mur93, Myb98a, Myb98b, Myb99, Nan95b, Nan95a, Nel97a, New73, New76, Nor96, Nor98, Nov82, O’C93, O’C95, Old91, Ols74, Ols75, Ols76, Ols89, Ora00, Par97, Pas90, Pat85, Pat90, Pec85, Pee96a, Pel93, Pen93, Pet96b, Pet96c, Pet96a, Pet98a, Pet98b, Pet99, PJ87, Pic86]. Book [Pra90, Pre71b, Pre71a, Pre71c, Pre70, Rag96, Rap00, Ras90, Rei85, Rei75, Rei77, Ric87, Ric89, Ric77, Rob96a, Rol87, Rou89b, Rou89a, Rou76a, Roy90a, Roy89, Roy90b, Rum90, Rus77, Rus85, Rus87, Rya80, Rya93, Sal72a, Sal86, Sam95, San97, San98a, San98b, San99, Sav97a, Sch96b, Sch97b, Sch00b, Sch87, Sch98a, Sch73b, Sel71, Sem98a, Sem98b, Sem00a, Sem00b, Sha97b, Sha98, Sha86b, Sha86a, Sha90, Sha96, She91, Shi85, Shi86b, Shi86a, Shi88, Shu78, Shu80, Sie70, Sie86, Sie89a, Sie89b, Sko98, Sle94, Smi96a, Smi96b, Smi97, Smi98c, Smi98b, Smi98a, Smi00b, Smi00a, Sny98, Sop96, Sou96, Spi94, Sta75a, Sta77a, Sta77b, Sta98, Ste99a, Ste00b, Ste00a, Ste77, Sti00, Str76, Str78, SS95, Swa75a, Swa75b, Swa76c, Swa76b, Swa76a, Swa77b]. Book [Swa77c, Swa78b, Tau74, Tay94, Tay95, Ten88, Ten90, Ten93, Ten94, Tes79, Tes86, Tho00, Til70, Tom98a, Tom98b, Tom96, Tom99, Tor89, Tsu79, Tur98, Tzv90, Und96, Und98a, Und98b, Usd85, Vag80, Van90, Van97a, Var75a, Var75b, Var78, Var97, Vei88, VG93, Wal00a, Wal86c, Wal86b, War79, War96a, War96b, War98a, War98b, Wei79, Wei88, Wei91, Wei93b, Wei93a, Wei94b, Wei94a, Wei95a, Wei95b, Wei96a, Wei76, Wei96b, Wel70, Wel71, Wel78, Wel80b, Wel94, Wel75, Wes89, Whi89a, Whi89b, Whi98a, Whi98b, Whi98c, Whi99b, Whi00, Whi98d, Wil98b, Wil00b, Wil00c, Wil85b, Wil94, Wil95a, Wil98c, Wit96a, Wit96b, Woo88a, Woo77b, Woo90, WH70, Wri71a, Wue97, Yin95, You75, Zvi86, Zvi87, Zwe95, BT84, Ben80, BF94, Coa86, Day97, DG91, Dio94]. book [Dun96a, FB81, GS87, Jai72, Men95b, SRS98, Wal87, Don76]. Books [Ano74c, Ano75b, Bel76, Gro76c, Har76c, Lan78, Lar99, Nov82, San98a, Str76, Wel70, GR74, Haa93, RV78, Tur77, Xia92, Kol75, Lap00, Wil98b, Cox98, Wei95a]. bookselling [Shi86a]. Boole [Kug71]. boolean [Ang75, Boo78, Boo80, Boo81, CCW84, Cro79, NKM77, Rad82a, Rad82b, SMB72, BP93, Cro86, DG87, FW90, FS91a, FSTV99, FBK97, HH95b, HH95c, Los97b, Sav97b, YS93]. Boon [Ber94]. Booth [CL90]. Borders [War98a]. Boris [Sel71]. Borko [Koc79, Wel80b]. Boston [Mea82a]. botany [MM97, Del90]. both [Win70a]. bothsides [DMZL82]. Boulder [Cri81, Kaz75b, Lun80]. bound [CN95]. boundaries [Cas92, Gir91, Shp92]. Boundary [Pie99, Ozo79b, SSD99, Wee92, YLYY00]. boundary-spanning [Wee92]. Bowker [Bro00c]. boycotts [MT96, Pee96a]. bracket [Cra86]. bracket-based [Cra86]. Bradford [Asa81b, Bas92, Bas94, Bro79, CL86, Che95b, CCT95, DG78a, Egg86, Egg90b, Egg90a, Has76, Hub76, Hub78, MK79, Mor81, OV93, OV95, OV97, Pop75, PK78]. brain [Gul87, Ray99]. branch [CN95]. branch-and-bound [CN95]. Brand [Ora00]. Brazil [Ros82a]. breadth [Eas89, Sub84]. Breakdown [Rou92a]. breakeven [EH94]. Breaking [Sch74]. breeds [ER95, Hub98, Tag81]. Brenni [Mar75]. Breve [Har70a]. Brewer [Mim85]. Brewster [Ste77]. bricks [Wil98b]. bridge [Jr88]. bridges [Lap00]. Brief [AP97, Ano70a, Ano70b, Ano71a, Ano72a, Ano72b, Art71, Bar78, Bau72, BK96, Bea80, Bel71, Ber71, Boo78, Bor73a, BP70, BM81, Bra70, Bra79, Bro70b, Bro80b, Buc95, BMFK76, BT99, Bur92b, Bur80, Cag70b, Cag70a, Car79, CC75, Cha88, CL90, Cod73, Coo71, Coo79, Cor82, Cra80a, Cra86, Cro94, CO95, CO96, CC99, Cro99a, DKD71, Dav76b, DDC78, DC93, DM99, DS91, DS96, DG00, Eas89, Eas91, Egg88, Ell76, Eve93, EB98, Fal98, Flo72, Fre77, GS93a, Gel79, Gle93, Gli72, Goy87, Gue78, HST96, Hai97, HY98, Hal85, Han91, HW80, Har71d, HN97, HAA87, Her84, HH70, Hoc74, HW71, Hol71, Hub78, Hur80, Jac71, Jac72, Jan91, JD78, Kan78, Kan87, KM73, KW77, KZC86, KF92, Kil93b, Kil93a, Kil95a, Kir73, Kje74, Koc72, Koc83]. Brief [Koe93, KH95, KK96, KG72, Kuc78, Kug71, LJ70a, LJ70b, Law76, Ler79, Lev73a, Lev74, Lou92a, MS76, MM87, MM88, MM97, MBH94, Mar72a, MZ78, MK79, Mau71, May78, McC88, McC78, Mea83c, Mon73, Mos70, Mur73, Nei80, NL83, Nic87, O’H70, OMY^{+}89, OE81, Onu82, Ozo79a, PKK74, Per83, Pic96, Pra83, RM97, Rob91, Ros79, SM91, Sal72d, Sal85, San75b, San77, Sch76, Sch75, Sha95, Sha86c, SC78, Smi84a, Ste85, SR95, Sub84, Swa72, Swa76d, Swa89b, Swa89a, SK98, Tag72, Tho73b, Tho72, Tyn75, Van97b, Voo74, Voo80, WV97, Wel80a, Whi91, Whi76, Wil84, Wil75, WL81, Wil79, WD97b, Wri71b, Wri72, Yag87, Yan97, Yar97, Yer71a, Yer71b, de 78, Has85, Kra99a]. Briet [Buc95]. Brigham [Don77]. Brigitte [Mat85]. Brillouin [BM81]. Bringing [San97]. British [Bro95a, Lap99, Sta77b, Bax85]. Brookes [TS92]. Brown [Fis79, Mea82a, Swa76a]. browsing [BHWG95, CHSS98, Los92, MM98, RCZS99]. Bruce [Har78b]. Brunswick [Gro75b]. Brush [Han92]. Brussels [Gro76a]. Bryant [Lan73]. Buck [CC79, SC80]. Buckingham [Kaz79]. Buckinghamshire [Mea82b]. Buckland [Don76, Wei90]. bucks [San98a]. Budapest [Bor73a]. budget [Smi74a]. Budgeting [KL79, HK80, OG93]. Buenos [Jac71]. Building [Cal00, KL79, Mac96, SF78, Mar00, Har76c, Han96]. buildings [Har76b]. bulletin [Ano78j]. Buneman [Rap00]. Bureau [Gro75b]. Bush [Buc92b, Bur92a, Koe92a, NK89]. Business [Ano74a, Don70, Key97, Pra83, Sch97b, Bol86a, Bol86b]. businesses [VTS97]. bust [Ber94]. Butterworths [Bla79]. bytes [Lap00, Lar99, San98a, Wil98b].

C [Afs99, Ano75a, Bla79, Bro00c, Har75c, Lan78, Lan73, Lun80, Nov82, Swa78b, TS92, Wel75, Woo77b, WM00a]. C-7 [Dav76a]. c1969 [Har70c]. cable [Mat85, Mat85]. Caesar [Hal95]. CAI [Woo76a]. calculating [Egg88]. Calculation [Lon92, Egg90b, Lou92a]. calculus [Bro79, vRL96]. Calibrating [FM86]. California [Lea75, Sut96]. Call [Ano83h, Ano85h, Ano86a, Ano86b, Ano86c, Ano00a, Ano00c, Ano00b, Bro99a, Mea78a, Kra92a, Par94]. CALP [Ste74b]. Calvin [Hen96]. Cambridge [Ano75a, Dav76a, Die82a, Har70c, Har70b, Lan77a, Rei75, Rou76a]. Cameron [Hei78]. Can [Her78, CH00, OMY^{+}89, Swa89c]. Canada [Inh76, Shu80]. Canadian [Ere80, Inh76, AI76, AC93, DS79, DCD81, Inh75a, Inh75b]. Canberra [Gro75a, Gul85]. Cancer [MH87]. Cane [Fis81a]. Cannon [Rei77]. cannot [Ley97]. can’t [Wat98]. capital [HK80]. CAPRI [Blo70]. captions [LBBR95, OOA99]. card [Erc98]. Cardiovascular [FG70, MG72]. cards [Har71d, LS70]. Care [Cim95, DG95, Gor95, HH95a, Her95, JS95, LH95b, MWH^{+}95, Rod95, SMJ95]. career [Bur92a, Jul99, Kuh99]. careers [Hei86a, Aul87]. Carfield [Dro77]. Carlo [Bur99]. Carmel [Var75a]. Carnot [Pre71a]. Carolina [Dou96]. Carter [Lan76]. CAS [Ols74, LBP76, WFK72, WFK73]. Case [Coz82, Lan77a, Nov82, Ros82a, Rou76a, Beg95, BCM^{+}70, BM75, Caw78, CB94, DW98, Dou96, Egg96, EH94, Erc92, Fid84, FS84, Har90, Hul96, Ker77, KB92, LL85, LBIP97, LK91, MS00, Mud98, PvR94, RM97, Rob80b, Rou92a, SH89, Sol93, Spa99, Sta75a, SS00, SR95, SS89, Sut94, TEC99, Und98b, WD97a, Wal75, WWSR00, Woo76b, Zen99, Zha00, EK97, Per82]. case-based [Und98b]. Case-Control [Per82]. case-study-based [Fid84]. cases [MUD99, McG95]. cash [War99b]. Cashless [Lan82a]. Catalog [MKB78, Til70, Bat77, Bel81, Car96, Coo70b, FS84, Her96b, Kil95a, Lar91, Lar92b, LMO^{+}96, SR77, Tag72, Bel81, BHWG95, Don70, Bor94, Kri86c]. cataloger [Ros76]. Cataloging [Can95, Dic96, HO89, Kil95a, Wei95b, Erc98, GO93, Gro90, KS70, Leh72, SM90, Tan96, Tan99, Tom99, Yee91, Roy90b]. catalogs [Bat86, Bor86b, BHWG95, Bor96, Boy86f, Con95, Dal90, DW96, Erc98, Har70b, Yee91, Dra89, VG93]. catalogue [Ell70, Jon89]. cataloguing [SS89]. catastrophe [Gil78, McG96a]. Catching [FW99]. categories [Fed77, Sav93, TSV99, Bro98b]. categorization [YW96]. categorized [Moo73]. category [CHSS98]. Cater [ER97]. Catholic [Hin70, Blu71]. CATLINE [TH85]. Causal [Seg94]. Cause [Lei73]. Causes [Mar83c]. Caveat [Pri72]. CBE [Gan74]. CCM [Har71b]. CD [CZ88, Dav88, EH94, Jas94, LS88, Rie88, Sch88a, Sch88b, Zoe88]. CD-ROM [CZ88, Dav88, EH94, Jas94, LS88, Rie88, Sch88a, Sch88b, Zoe88]. centenary [Buc95]. Centennial [Ric77]. Center [BDBR97, BD97, BDSA97, D’E97, Goo97, GLS97, LS97, LH97, Ric77, RH97, SA97, WH70, Wri71b, DS73, EW70, FW88, LR79, Mos70, Woo70, ABB^{+}91a, ABB^{+}91b, BDBR97, BF91a, BD97, LS97, RRB91, SA97, SB91]. centered [All00, Bal00, Bru94, Fid94, GQSdP98, Mor94, RFK97]. Centers [ABB^{+}91a, ABB^{+}91b, BS91b, BF91a, CS91, DRCG91, LD91, RRB91, SB91, DDC78, Dou72, Lap00, Mai90, Tal79, Wei73, Wes87, Wri73, New73, Swa78b, How87]. Central [Pic86, War79]. century [Ell70, Fro94, Kil93b, Lan94, Mil90, Nan95b, Nis99, Sut96, Wil98b]. CEOs [AC93]. Cerny [Rob79]. certain [Don72, Fal98]. Challenges [BDBR97, Bes96, HG98, Gar88, Wei94a]. Champaign [Har76a, Ric77, Swa76b, Swa77b]. Chandler [Ols75]. Change [Cha77, Kaz77, Hur88, Pee94, PR98, Rao80, Sem98a, Whi76, WMME88, vdBL96, Kra74b, Bea93, Lie96b]. changed [Bar78]. Changes [Mea81, BB95, MKDL83, Kir81]. changing [BNS^{+}00, Bur85, CS91, CS90, LD85, McC84, Wal98]. channels [BR97, Wee92]. chaos [Kir93, Nic00, McG96a]. Chaplin [Swa76c]. Chapman [Wri71a]. Chapter [Har75c]. character [Eas91]. characteristic [Her78, TT79]. Characteristics [Bon82, Her82, And78, Boy86e, DG91, JJ77, KDS89, Los96c, NK00, Wal97, Wri73, FG70]. Characterization [LM81, MN83a, Jon76, PZ83]. Characterizing [JJ80, CPR84]. characters [Tau84]. Charges [Lan80, Bra77]. charging [Pen70]. Charles [Die82b, Koe75, Wel80b, Whi99b, Lou92b]. chart [Wel72]. Charter [Che76, Dav77b, DC76]. Check [Tho73b]. checking [LNL99]. checksum [CC75]. CHEMCONF [O’H95]. Chemical [BR95a, Ols74, Wei97c, Bot70, Bro80b, Bun80, Keg70, MS76, Mes95, PvR94, PBvR95, Ros79, Ros82b, Rot79, SWK78, Toc70, Wig95]. Chemist [BR95a, Chr95, Fag95, Men95a, Mes95, O’H95, Wig95]. chemistry [AP96b, Koe75]. chemists [BWV72, Bro99c]. CHEMLINE [SWK78]. Chen [Lan77a, Rou76a]. Chernyi [Sch73b]. Cheshire [LMO^{+}96]. Chicago [Boy85a, Die82b, Kuc76a, Pec85, Usd85, You75]. Chichester [Kra79, Mar85, Mea83b]. Chief [Dro78, Pre71b]. chih [Lan77a, Rou76a]. child [BW85]. child-selected [BW85]. children [BW85, Mar94, MBS98, Bil00, BHWG95, Hir99, SCD98, Sol93]. chile [Fre77, FF80, Ano72b]. Chinese [CYW99, CWBL00, He98, He00, Kwo99, LNL99, Tau84, WEWP97, WT93, WT95, YLYY00]. Chinese-to-English [He98]. Ching [Lan77a, Rou76a]. Ching-chih [Lan77a, Rou76a]. CHMINF [Wig95]. CHMINF-L [Wig95]. choice [BW85, Glu95a]. choices [Pet96a, Boy85a]. chrestomathy [Woo70]. Chris [Fis81a]. Ciel [Lan76]. Circulation [Mcg79, Rou74a, BS91a, Boy85a, GS87, GO93, McG73, McG84, McG96b, Smi84a, Spe80, TT79, Tur77, Sha75]. circulation-use [TT79]. circulations [Bur90, BF94]. Citation [BCM^{+}70, Can89, CN81, Caw78, CPR84, CO94, Del90, Gar83, GS97, Gor82, HY98, Lif93, Lip83, Liu97c, MM97, MSPPM98, MG72, MS00, Poy82, TL74, Whi94, WW73, BMvR91a, BMvR91b, CH00, Caw74, Col92, Cro94, Die79, EP91, Gar97, GdCD81, Her78, Kra88a, LB83, Leh72, MM89, Mar83c, MAN80, MB81b, McC89, MN83b, MM78, NC75, Nee81, NL83, Nom84, NML99, PW89, Per92, PvR94, PBvR95, Pic96, Pra83, Raf75, Rob91, San75b, Sha92, Sha85, Sha91a, Sha91b, SH89, Sma73, Sma86, Sma99, Sum84, TvR90, Tsa98, Van98b, Wal97, Wei97b, WM98, Win70a, Yar97, Coz82, Mar83c, Per82]. Citation-based [CO94, Whi94]. Citations [Bor85a, KH71, Law82, Wel80a, AP97, Bat00, BCM^{+}70, Bro83, CH73, Cor82, Gup90, Swa89c, YAH90, Rei77]. Cited [Bon82, Cam96a, Gar79, HNW93, Her78, Hoo91, Hur80, Kwo85, Leh72, OMY^{+}89, OR78, OG80, Yoo94]. citedness [Seg94, SK98]. citer [Bro85, Bro86b]. Citing [Bon82, CH00, HNW93, Her78, Kwo85, MM87, MM88, WW99]. citoyens [Kra91a]. city [Aga99, TK96]. civil [EK97, MN83b]. civilization [Myl92]. Civilizing [Cas97]. clarification [Kra91c, Lip83]. CLARIT [Pai93a]. Class [ODVG87, Car79, New94, Pra77]. Classes [Fau96, Hol96]. classical [Dom00, Myl92, OV97]. classics [CPR84]. Classification [Bro00c, Dil82, FK73b, FK73a, Lar92a, Lar92c, Law82, Ric76, Ric77, Aul71, BG98, BCSZ79, Bie95, Dav76b, ER91, FBK97, Gli74, Gli78, Gre96, HZ80, JB71b, Jon73c, Jon74, LDS^{+}00, LH95a, May97, Moo73, MQP98, RF97, Soe72, Soe99, Yu76, Can89, Los92, McI97, ODVG87, Lea76, Bat93]. classifications [JN93, Wil84]. classifiers [DCRS98]. classify [Jon74]. classifying [OCN97]. classroom [LB00]. classrooms [GSKL00]. clearinghouses [Wes87]. Clelland [Rei75]. Cleverdon [Swa78b, Ken87b, Rob98]. Cliffs [Art76a, Dav79b, Dom78, Fis81a, Fis81b, Har83, Mar80, New76, Str78]. Clinic [Swa77b, Swa75b, Swa76b]. Clinical [War88, WRE^{+}70]. Clinician [Fri73]. Cliqueing [Ger78, Gli78]. Clive [Tes79, Wel70]. clock [Ora00]. close [Mat97b, Wat98]. Closed [CK96]. closer [Hei95b]. Closing [Die82a]. cloth [Woo77b, Fis81b]. cloudy [Jac91]. clue [Swa89c]. clues [Bar98a]. Clumping [BKR98]. cluster [Mat00, MPW73, Sha91b, Yer83]. cluster-based [Sha91b]. clustered [Cro79]. Clustering [Bro90b, CN73, Cro77, Goe87, RF97, BMSM90, Bur95, CO84, CO85, Cod73, DC80, FPS00, GV87, Gor91, Har93b, Koc83, Lov71, MN83b, MUD99, Sha91a, TZ92, Yu74, Swa73]. clusters [AM70, Ger78, Sma86]. CO [Har76c, Lun80, Boy85b, Cri81, Ere81a, Nom84, Sma73, Var78, BMvR91a, BMvR91b, CMK^{+}98a, CMK98b, Ley97, Ley98, NJWC97, PW91, Sha85, Sma86, WM98]. Co-citation [Nom84, Sma73, BMvR91a, BMvR91b, Sha85, Sma86, WM98]. co-occurrence [CMK^{+}98a, PW91]. co-operative [NJWC97]. co-word [CMK98b]. co-words [Ley97, Ley98]. Coal [De 80, FBC79]. coauthor [LS87]. COBOL [Fis81a, Dav92a]. Cochrane [Boy86f]. Cocitation [BSM90, COC90, LW90, McC90, Pai90, BE80, McC84, McC91, WG81, Whi83]. cocitations [BSM90]. Cocited [McC86, Whi81, McC88]. CODASYL [Kra79]. code [LY91, Los92]. CODEN [AB76]. codes [Gre90]. Coding [CC79, Boo74b, CM00, Cra82, Cra86, Goy84, SC80]. coefficient [CO87]. Cognition [Lin96, Mon81, Nor96, Sim81, Ano81j, Che99]. Cognitive [For00, PK00, PBvR95, Ber84, Bru94, CCN93, Dal90, Dav77a, ENN98, LBBR94, Mic94, NT96, Sch00a, Tod99, WS98, WW99, Bil00]. coherence [RSZ71]. coherent [GS93c]. coin [Fau96]. Coinage [Sha95]. Cold [Law74, MT96, Pee96a]. Colin [Bro00b]. Colinked [HC96]. collaborating [CMB^{+}95]. Collaboration [CH98, Fau96, Hol96, Pao82, IT89, MP98, QLA97, Nan95a]. Collaborative [Hol96, Kar98, Ano08, Che97, Mar99, McC88, Rou98, Wil00a]. colleagues [Ler79]. Collected [Gro90]. Collection [DS79, HJD^{+}99, PZ83, RTM99, CH73, Col00b, EW70, EP97, Koe79, Lee00a, LT74, Los90, RCZS99, San79a, SC76, SC78, YR98, Zen99, Sha91a, Car95a]. collection-control [SC76, SC78]. Collections [Kuc75, BC00, DKP00, RCZS99, RT81, Sal72b, SFB^{+}00]. Collective [Bot73, SK98, SK99, Har76a, Nan95a]. collectors [SB80]. College [Ano77a, Woo88b, Cox93, Nom84, Zal74, de 71]. colleges [CH99]. collocation [Car96]. colloquium [Bro95a, Har76b]. Colonicity [Per85]. Color [Pap88, CDV98]. Colorado [Kaz75b]. Columbia [Ere80, Har75c, ST89]. columnists [Dur76]. Combination [YLYY00]. combinations [FS84, Har90]. combined [BE80, BMvR91a, BMvR91b, EP91, Eve93, Gle93]. Combining [NML99, RC95]. coming [Kra99c]. Comment [Aiy88, Tia89, McG84]. Commentary [Woo87a, Kra99a]. Comments [Ano72c, Bor74, Cou98, Dun96a, Old88]. commerce [Nor98]. commercial [BM80]. Commission [Art74, BL85, Har70c, Kaz75b, Has85]. commitment [Bar96]. Committee [Ano71a, Har70c, Gan74, Hol71, Joh84, Var75b]. commodities [Mow92a, Mow92b, Mow92c]. Commodore [Mim85]. Common [Lei73, BSR93, Gra82, Yu76]. commonsense [BSW00, Sou96]. Communicating [BPK^{+}71, Pra90]. Communication [Ano72a, Bea80, Bro00f, Bro80b, BT99, Bur80, Cag70a, Cra80a, CC99, Cro99a, DDC78, DM99, DG00, Gel79, Gue78, HW80, Hur80, Koc83, Law76, Mea83c, Mig80a, Nei80, NL83, Per83, Pra83, Sub84, Voo80, Wel80a, Whi76, Wil84, Wil93b, Woo77b, Bai75, BA74, BR92, CH00, Cit98, Cov00, Cra71b, Cra72, CB83b, Dro95, Ehi90, Hap99, Har88, Har98, HJ89, Hur00b, KC99, KM99, KM00b, Koc81, Kor74, LS80a, Lyn77, Pai93b, Pep74, PR98, RB83, RCJS94, Rub92, SC85b, Sch94, Sch74, Sha83b, SPWH99, SRS98, SH94, Van97a, Var97, WKMG00, Wee92, Whi98d, Wil85b, Wol72, Bra94, Pre71a, Wal72, Boy85b, Die84a, Har70d, Swa75a, Afi98, Boy88, Has95b, Sha86a, Sie89b]. communication-graph [Sha83b]. communication-information [Rub92]. Communications [AP97, Ano70a, Ano70b, Ano71a, Ano72a, Ano72b, Art71, Bar78, Bau72, BK96, Bel71, Ber71, Boo78, Bor73a, BP70, BM81, Bra70, Bra79, Bro70b, Buc95, BMFK76, Bur92b, Cag70b, Car79, CC75, Cha88, CL90, Cod73, Coo71, Coo79, Cor82, Cra86, Cro94, CO95, CO96, DKD71, Dav76b, DC93, DS91, DS96, Eas89, Eas91, Egg88, Ell76, Eve93, EB98, Fal98, Flo72, Fre77, GS93a, Gle93, Gli72, Goy87, HST96, Hai97, HY98, Hal85, Han91, Har71d, HN97, HAA87, HH70, Hoc74, HW71, Hol71, Hub78, Jac71, Jac72, Jan91, JD78, Kan78, Kan87, KM73, KW77, KZC86, KF92, Kil93b, Kil93a, Kil95a, Kir73, Kje74, Koc72, Koe93, KH95, KK96, KG72, Kuc78, Kug71, Kun82, LJ70a, LJ70b, Ler79, Lev73a, Lev74, Lev76b, Lou92a, MS76, MM87]. Communications [MM88, MM97, MBH94, Mar72a, MZ78, MK79, Mau71, May78, McC88, McC78, Mon73, Mos70, Mur73, Nic87, O’H70, OMY^{+}89, OE81, Onu82, Osb98, Ozo79a, PKK74, Pic96, RM97, Rob91, Ros79, SM91, Sal72d, Sal85, San75b, San77, Sch76, Sch75, Sha95, Sha86c, SC78, Smi84a, Ste85, SR95, Swa72, Swa76d, Swa89b, Swa89a, SK98, Tag72, Tho73b, Tho72, Tyn75, Van97b, Voo74, WV97, Whi91, Wil75, WL81, Wil79, WD97b, Wri71b, Wri72, Yag87, Yan97, Yar97, Yer71a, Yer71b, de 78, Ano81k, Che86b, Mea83b, Sem98a, TB93, Wue97, Cer83, Sha86b]. communicators [BPK^{+}71]. Communities [KDJ94, DC98, GCS97]. Community [Ano77a, CNMS97, Bea92, CYFS95, Fra99, Haw74, Hei80, Hei91, Kra91a, Lev99, VS99, Wee99, ZK72, Wil75]. compact [NC86]. companies [Key97]. companion [Smi98a]. Company [Gol74b, Lea75, Mea82b, Nov82]. Comparative [Ros71a, Wel80a, Koh74, LBIP97, Ric70]. compare [GdCD81]. Comparing [AKE94, Cha99a, Los97b, Pai93a, Sma80, TC00a]. Comparison [CC75, DS79, DCD81, GH96, LBP76, MAN80, SWS71, BBDH00, Boo81, BM80, BNP99, Bur80, Coa86, Elc78, Jan94, Kin76, KPZ^{+}99, KG72, Mar83a, May78, NN92, OWF99, Ros79, Sal72c, Sch82, Sch76, SBW90, Wil92a, Wil98a, Zar82, Fay73]. comparisons [Aul71, NC75, Sal86]. COMPENDEX [Byr75]. compendium [Jac96, Rou89a]. competence [Glu95b]. competencies [Hei86b]. competing [Bod99]. competition [Cam96b, Rob80a]. competitive [Wil85b]. Compilation [Gol73, Goy87]. Compiled [Mar75]. compiler [Swa76a]. compilers [Gro76b]. compiling [GS93c]. Complementarity [Raf75]. complete [Gro90, Smi98a]. completion [Egg90b]. complex [Bro86b, RFK97, SCD98]. complexity [BF91b, BLPG99, Kuh99, WH91]. Components [Mom82, DG95, Mon81]. Composite [Dav77b, Sha91a, Sha91b]. composition [Law74]. Compound [Jon71, Bla71]. comprehension [CM00, LBBR94, LBBR95]. comprehensive [Bro90b, Hei94, Wil98c]. compressed [BM95, MB95]. Compression [CELY80, BMNM^{+}93, CYW99, CL82a, Jak88, LY91]. computacion [Har70a]. Computation [CO87, HW80, Bue86a]. Computational [BG98, Pao82, Wal81, Lid96, Wei96a]. Computer [Bea91, BZ85, Bro95a, Bry70, Col96, Don70, FC92, Gau92, Ger79, Har92a, KPZ^{+}99, Lan76, LH92, Mar80, Mar92a, MHA82a, Myl92, Nan95b, Nov82, RG70, Rou98, Sew82, Sha97a, Sha75, SS76, Sta77a, Vei88, Wit81, Ami70, Bam79, BWV72, BS88, BW85, Bel76, Ber94, BCM^{+}70, BM75, BR80a, Boy85b, Car88, Cas85, Che86b, Cli70, CCN93, Cra00, DKD71, Div87, Gli72, Haa96, HJ89, HHC70, KWR98, KK96, LMM^{+}73, Lap00, MR81a, MR81b, McL90b, MHA82b, Mim85, Nan95a, New76, NJWC97, NK83, Nor96, RB83, Rob80b, RR77, Sha98, SD96, SHS72, Ste74b, Str76, Sub84, SH96, WD97a, Wei96b, WC84, Zmu75, Ano81k, Car76, Jer70, RK79, Ros73, VC70, WKMG00, FW88, Art75b, Che76, DC76, Dil98, How89a, Kri86b, Par97, PJ87, Smi00a, Whi98d, Boy90a]. Computer-aided [SS76]. computer-assisted [BS88, HHC70, Jer70]. computer-based [BWV72, BR80a, DKD71, Nan95a, Rob80b, SD96, VC70, Art75b, Smi00a]. computer-mediated [HJ89, WKMG00, Whi98d]. computer-monitored [RB83]. Computer-readable [Kri86b]. computer-supported [KWR98, SH96]. Computerization [Pet96a]. Computerized [JBH77, Ama84, Bat97, CH98, GO93, Hoc89, SS89, Tzv90, Wil94, FRGL75, Har70b]. Computers [Art82, Has95b, Pre71c, Ren70, Rus70a, Swa77b, AN73, ASR85, Cas96, Mar83a, War96a, War96b, Kol75, Hei95d]. computing [Bat97, CMK^{+}98a, Mea83b, WY93, Wri72, Mar95]. comrades [Kra91a]. Comserve [SH94]. Con [Har70a, Sch99]. Concentration [Rou92b, BE91, Bur92b, Car79, Egg88, ER91, Pra77]. Concept [He98, VS87, Art71, CO87, CN95, CNMS97, CMK^{+}98a, CHSS98, DS84, FNNW74, Jon73c, Sch96a, SB91, WNAR99]. Conceptions [Smi98c]. Concepts [Dil97a, Wel80b, Art73, Bel75, Boy87a, Bro98b, Col96, Col95, Fis81b, Lan89a, Ozo79b, RT81, SNG00, LS84b, Ren70, Boy90a]. Conceptual [Aik97, Cas91, MA76, Sha97b, Sut99, BF91b, Cha86, FS83, He98, He00, JGR90, MR87, Roy90b, Wei96a, WY73, YFR81a]. Concerning [Lou91, Bac72, McI97, Mea82d, RV98]. concerns [Can95, Dic96, Win95]. Concise [Bel76]. conclusions [Ber70]. condition [Gre90]. conditions [Ere85, Hei80, Hei91]. Condorcet [vdVM99]. conduct [Sha96]. conductors [Caw74]. conduit [Day00a]. CONF [Ger79]. Conference [BI98, Gul85, RH99, Smi98c, Wes00, Dro95, Inh76, Joh84, Les89, OMY^{+}89, Old91, JJ75, Ten88]. conferences [KRD95, Wee99]. conferencing [O’H95]. confidence [LT74, OWF99]. confidentiality [Cre95b]. conflict [CH98]. Conflicts [Ric89, Pet96a, Pic86]. Confronting [Doc91]. CONFU [But99, Cre99a, Sun99]. CONFU-sed [But99]. Congress [Joh86, Lar92a, Til70, Don79, KS70, Mar71, Lar92c, ODVG87]. conjecture [DC93]. Conjunct [EAW99]. conjunctions [DG87]. Conn [Wel70]. connected [Ger78, Swa87]. Connecticut [Koe75]. Connecting [WKMG00]. connection [GL96]. connectionist [CN95, Che99]. connections [ABLS88, Swa89a, Wes00, Woo88b, Sou96]. Conroy [Kaz75b]. conscious [Gar71]. consciousness [Cha76]. consensus [Beg95]. consent [Cox93]. consequences [Day99a, HC84, KC99, Sun99, ERV00, OV98, Van97b, Bro00c]. conservation [MN81]. consideration [HP00]. considerations [CZ88, Coo98, HH70, Kim73, Mes95]. considering [Raf75]. consistency [Sch91b, SA91]. Consortia [WC92]. constancy [Koe99]. constant [Hen99]. Constituent [MHCW89, WA79]. constraints [MBS98, Wes91]. construct [Col94, SWT^{+}79]. Construction [FK73b, Lea75, Beh96a, Beh96b, Caw74, FB88, FS91a, KH80, Kim73, Kuh99, McK00, RL76, McD99, Lan81b]. constructs [How94]. Consultants [Rus85]. consultation [Anv91, CN95]. Consumer [JS95, Lar74, Klo89]. consumers [JS95, Tag75]. contact [McC78]. contagion [Wor73]. contained [Tau74]. contemporary [CMRWW93]. Content [Deb72, DF82, YCO00, BI00, Bel73, BKR98, DF84, Gro96, HG00, HAA87, MM97, Sch00a, Wei96a, WA96]. Content-Bearing [DF82, BKR98]. Contents [Kol74, Gue78]. Contents(R) [Mau71]. Context [Bar95, Kan76a, MSPPM98, McC95, OO97, Yer73, YF85, Bro98a, Chi93]. Contexts [Kay98, Erc98]. contextual [RSZ71, Gel96]. continuation [War81]. continuation/cancellation [War81]. continue [OR78]. continuing [Kaz75b, Var99, BB95, Hei87]. continuity [Lie96b, NK89]. continuous [CELY80, Jan91, Rou90]. continuum [KM99]. contours [Bro70c]. contracting [Sle94]. contractor [Wol87]. contractor-supported [Wol87]. Contribution [Mcg79, BM81, Koe92a]. Contributions [Ano77a, Shu78, Swa76c, EF96, Wal81]. Contributors [Ano85a, Ano80h, Ano82d]. Control [Fis79, FS91b, Aba86, Boy87b, Cro99b, DK90, Ere81b, Fay72a, Fay72b, Fay73, Har74, Kuc76b, LJ70b, Lan93, LCL85, MDS70, Rob96a, Sal72a, San79a, SC76, SC78, TF89, Wel80c, Wel75, Per82, Yin95]. Controlled [Bal88, BM89, LK97, Mud98, PN98, Qin00b, Sve86, Fid91b]. controversy [GD90, Pet96a]. conundrums [All00]. conventional [Lev73a, Sal72c]. Convergence [Koe92b, BR92, Koe92c]. converging [Whi99b]. Conversation [Wee99, Glu95a]. Conversion [Til70, Bor86a]. convertibility [Ang75]. Convey [Tes79]. convolution [Egg94]. Conway [Dav76a]. Cook [Boy85b, Har79, Lou91]. cookbook [Sou96]. Cooper [Coo76]. cooperating [EW70]. Cooperation [Rob80a, SH96, GH96, NT85a, Swa75b, Nor96]. cooperative [DW98, GQSdP98, KWR98, Nan95b, SH96]. coordinate [Gue78, Gul87, Jon71, Kil97, KW97, MDS70]. coordinated [vdVM99]. Coordinating [KWR98, Whi98a]. coordination [BK98, KB98a, Los87]. cope [Ker74, Tor89]. Copenhagen [Buc91b, Var75b]. Coping [Gil78]. copy [Ger79]. copying [Day99a]. Copyright [But99, Cre99b, Cre99a, CH99, Day99a, Fra99, GL93, Gas98, Gas99, KW81, Lev99, LCB99, RM99, Ste99b, Sun99, Wei96b, Wit80, DG78b, Mat97a, Tal74, War93, War99b, HC84]. Core [Fei73, Hir78, McC95, Oro77, GSKL00]. corners [O’H70]. Corp [Mig80a]. corpora [Sav99]. Corporation [Har71b, SL83]. corpus [AM70, YW96]. Correcting [SJ72]. Correction [PZ84, Zam80, BSY00, Van80]. correlates [KS77]. Correlation [Elc73, Cag70b, Cag70a, NL83, Wri72]. correlations [KPZ^{+}99]. Correspondence [DOO^{+}96, KL98c, OV98]. correspondent [FS97b]. Cosmochimica [Lif93]. Cosmology [LD85]. cosmopolitan [MP98]. Cost [AKS96, ASS73, BCB96, Elc78, Hay96b, MBFK76, BB83a, Coo72, CBM^{+}94, Fay73, Har70b, Jär89, Koe88, Lan71, McM73, Mos70, Sny98, Sou96, Spe71]. Cost-Benefit [MBFK76, Sny98]. Cost-effectiveness [Elc78, Lan71]. Costigan [Vag80]. costing [Hay96b, Kin98b]. Costs [AL96, Abe96a, KS70, Spe71, WA79, WC75]. Council [Gan74, Har70c, Har77b, Lan80, Lev73b, Rus85, Swa76c]. Counseling [Bai75]. Counter [EQH00]. Countering [Smi98b, Xia92]. counterpoint [Tia89]. counting [Egg96, Opp98, Pai96, Van97b]. countrie [ZPB98]. Countries [Ros82a, Ali90, DOO^{+}96, ERV00, Ere81c, EN85, Ere85, Gat92, Gor79, Ker77, SBA85, Rob96a]. country [EH94]. counts [BR95b, MSPPM98, Rou92a, Van98b]. Countway [Lan77a, Rou76a]. coupling [BE80, HD72, WD97a]. course [AdKK87, SHS72, SK99]. courseware [WA96]. court [BM75]. courtesy [Van97a]. courts [EK97]. cousins [BS92, Hor93]. cover [CO87, Mat85, Mat85, Shi85]. Coverage [Mau71, Poy82, WFK73, BR80b, FG70, Gar86, Glu90, Inh75a, Pic86, Sma80, Tho72]. Cowan [Woo77b]. Craziano [Lev76b]. Crazyna [Swa76a]. creating [Bro00d, FPS00, Mes95, Wee99]. Creation [Rie88, Boy86a, EFW96, Gat92]. creative [Col96, Haa96]. creativity [Haa96]. credits [Egg96]. criminal [EK97, Ere81b]. crises [TEC99]. crisis [Pra90]. criteria [Avr71, Bar94, GS78, HS78, Hir99, LB83, Lev74, RSZ71, Kaz75b]. Critical [Sha85, Bom75, Cas98a, MM89, MBH94, RW86, Ray99, Bor93, SS95, Und96]. critique [Wel83]. Crook [Lev76a]. Cross [AN73, BC00, FB78, FS97b, MDS70, Nee81, Pai90, Whi98d]. cross-citation [Nee81]. cross-cultural [Whi98d]. cross-fertilization [FS97b]. Cross-language [BC00]. Cross-national [FB78]. cross-referencing [MDS70]. Crossing [DPD99, Pie99]. crosslanguage [He00]. crossroads [Dav96b, Hei80, Hei91, Lan82a]. Crouch [Har79]. Crowers [Wel75]. CRT [FS84]. Crunch [Car76]. crystal [Jac91]. CT [Ano77a, Ere81b, Gro76c, Har77c, Kun82, Lan78, Mea82b, Ols76, Shu78]. cues [All89, BTSKC96, KL98c]. cult [Cob96]. cultural [CA80, Whi98d]. culture [Die81, Pet98a]. Cumulation [Hin70, McM76]. cumulations [Har71b, Bol86b]. Cumulative [Hub98, de 78, Kan78, Kuc76a, de 76]. Current [ASS73, Bix96a, Bix96b, MS76, Wan81, Ami70, BWV72, BB96a, Bro95a, DH70, Dun96a, FRGL75, MP74, PS79, FSTV99, Wel00, Wil82, Dav79b, Gue78, Har71b, Mau71, Tsu79]. current-awareness [Ami70, BWV72, DH70]. Curricula [BIFW71, BWKG75, Hay88]. Curricular [WA96]. Curriculum [BZ85, GM86, Kaz77, Coo85, WC84, Whi89a]. Curtis [Don77]. curves [TT79]. customization [BBFM99]. Cutting [DM99, Smi74a]. Cyberbrowsing [BBFM99]. cybernetic [Kha91]. cybernetics [Wel80c]. cyberspace [BLPG99, Cas97, KL98b, KL98c, Nug91, Rom91, Smi98b, War98a, Cre95a, Sha97b]. Cybics [Koc80, Wel81]. cycle [Lev82, Tay82]. Cyril [Ken87b, Rob98]. Cystic [Sha91a].

D [Gro77, Har77a, Lun80, Nov82]. D.C [Don74, Har77b, Hin70, Lan73, Pre71b, Til70]. Dagobert [Lea75]. Dahl [Mea82b]. Dan [Cal00, Rap00, Kra92c]. Daniel [Har77c, Vag80]. DARE [ZK99b]. Darsee [KB92]. Data [Ano74a, BMNM^{+}93, Bru78, Elc73, Fre80, Ger79, Gol73, IC97, Kra79, Lev73b, LY98, LY91, Swa77b, Wil77b, Zam80, Agh86, Aik97, BM95, Bon79, BM72, Boy88, CVK92, CD94, Cor99, DGS86, Dod79, DPP78, Eas91, Ell76, Exn98, FCJ^{+}92, FS97a, GO93, GBDH76, GFHR97, Haw74, HAA87, Her97, KG72, Kri86b, LT74, LCL85, Los90, LH76, Mar72a, MK79, Mea83b, MCD73, NK00, PN96, PW91, Rap00, Ren80, RB83, RFK97, Sal80, SWS71, Sch76, Sol74, SF78, Swa75b, Swa76b, TBCE94, TT79, VA73, Vei80, WL80, War73, WSQ94, Wes89, Whi77, Whi76, WC75, Wil77a, WL81, Win70b, WBX98, PW92, FCJ^{+}92, Wil75, Dom78, Rap00, Rum90, Sch97b, Str78, Wes89, Cox94]. data-base [VA73, War73]. Database [Fed82a, MHA82a, NC86, Zar82, Anv91, Ban98, Bat98, Boy86d, Coo85, EH94, Goo99, HT91, Hoc89, KET99, LCL85, Los90, Los96c, LFHG99, Lyn91b, McC95, MW91, MHA82b, NT96, OWF99, PK00, Qin00b, Sch88a, Sch81, SD97, SL99, TH85, VD98, WdBFF95, WBX98, ZC92, Zvi86, And91, Roy89]. database-filtering [McC95]. Databases [Ali90, Afi99, BMSM90, Boy89a, Boy90b, Bro88, Cha88, Col81, CBM^{+}94, Ens95, FM86, GWDS86, Jär89, KSDR90, KZC86, Kri86b, LM81, Lyn88, Mea87, Neu95, Par00, Pit82, PT87, Rie88, SW92, She91, SMJ95, SS93, Tes86, TF89, Wal86a, WL81, WLC00, YAH90, Yer83, YG90, Hoc89]. datalog [Fuh00]. date [BR80a, Din87, San75b]. Datenverarbeitung [Wel71]. David [Dav76a, Wil00c]. Davis [Koe75, Woo77b]. days [Kra92a, Kra91b]. DC [Gan74, Ols74, Rus85, Aul71]. DDC [Hjø99]. Dean [Buc92a]. Deans [CC99]. Dear [Byr76]. death [Sem98a]. debate [Tal74]. Debons [Hei78]. Debunking [Cov00]. decade [Bea93, Mas97, Raw93]. decades [Lip99]. December [Cra96, Hen96, War79]. decentralized [BS91b]. d’échange [Gro76a]. Decimal [McI97, ODVG87]. Decision [BF88, RM97, Arm94, BS75, CB83a, Cit98, Dep75, DS73, EF96, FGS81, GO93, Hig99, Jul99, KdK90, Kra73, KB78, Kra78, OG93, Pul74, Rei75, RSK74, Sch97b, Tay94, TB93, War81, YF85, YdKKM87, YK93, Kra74a]. decision-making [FGS81, Rei75, YF85]. decision-support [Arm94]. Decisions [Ere80, Sta77a, BM75, TO84, WW99, Fu96a, Fu96b]. decline [Lar91]. decoding [BM95, Col00a]. decrease [Gup90]. Dedication [Bal89]. Deductive [JN93, MW91]. deep [Nei80, Ste99a]. Defence [RC74]. Defense [Ano84m]. Deficiencies [Gor79]. define [LMSC82]. defined [Aba86, Bar94]. Definition [LMM^{+}73, Bel77, Dom00, Glu90, Los97a]. definitional [Lyn94]. Definitions [Buc91b, Bur71, ER00a, NKB72]. degraded [PN96]. Degree [Col95, Flo00]. Dekker [Bai75, Dan84, Swa75a]. delays [ER00c]. delegates [Dav84, Mea83d, Mea84c]. Delivery [Vag80, CS91, Cul85, Mat85, RMMS96, Rus85, SC85a, SC85b, WSB98, WA96]. Dell [Boy85b]. Delphi [Mil95, Neu95]. delusiveness [BSR93]. Demand [Mcd82, Ber72a, Cas78, War73, Mow92c]. demands [KR80]. Democracies [Lie94]. democracy [Der94, Lie94]. democratic [Bra94, Mar94]. democratization [SB94]. Densities [Bai74, San74, San75b]. density [Gup90, KL98b, MB81a]. density-of-use [MB81a]. Department [BS91b, Bax85, Mea79c, NN92]. departmental [MB81b]. Departmento [Har70a]. dependence [BSR98]. dependencies [BDP98, Cro86, WBX98]. dependency [Cra91]. dependent [ER95]. deployment [DL00, Sut99]. Depository [Rya93]. Depth [Mar79, Wei95b, BM89]. Derek [Koc84c, Ste84]. Derivation [Pit75, GK89, SM90]. Derivative [SL99, GS93a]. derive [Dil00]. derived [Ras80, Kir81]. Deriving [AM70, Bri91, Coo70a]. Describing [Mom82, WV97]. description [GSKL00, Tho71, Zen99, Fay72a]. Descriptions [LBP76, Cra86, Gor91, O’C70, TFTB94]. Descriptive [MMB99, Cra82, KDS89, LBBR96, Roy90b]. Descriptor [McC89, Avr71]. descriptors [Bro93b, HC96, Hum99, Swa77a, Swa78a]. deselection [RM97, TB93]. Design [CZ88, Coo98, Hei74, HKE97, Kaz77, Mes95, PZ84, SWK78, All00, Avr71, Bac72, Bat86, BCM89, Boy86e, Bro00d, BT99, CMB^{+}95, Col81, CCN93, Coo70a, CNO^{+}00, CT87, DW98, DW96, Fu96a, Fu96b, Gil78, Her99, HCL^{+}00, KBN90, LUH^{+}96, LS71, Mar00, MB71, MCBC89, RNBV98, Roy89, Rum90, San97, Sch73a, Sch97a, SA97, Sha96, Sut99, Sve86, SWB79, Tho71, VS99, Was80, WBX98, Yee91, ZvD00, Dav77b, Dav79b, DL00, GS78, MR81a, MHA82a, SFB^{+}00, VA73, Che76, Fis81b, Bro00b, Dil97a, Her97, Agh86, Boy86g, Kie98, McL90b, Und98b]. designed [CC98]. Designer [McD99]. Designing [Mil00, Mor78, Shn86, WT94, Woo77b, Smi00a, GMPS00, LMO^{+}96, Bro96, Car88, Lit96]. designs [Kat00, Kra93, Tra98]. desire [Mat97b]. desk [Law76, Rie88]. desk-top [Law76]. desktop [BR95a, Lea98, Spr99]. desktops [Fag95]. Desmond [SS70]. detailed [Hul96]. Detection [PZ84, Zam80, YLYY00]. determinants [PvR94]. Determination [MN81, Sol74, Gor82, Los96c]. determinations [Sut94]. Determining [KN00, Hir78, YR77]. Detroit [Bol86b, Var78]. developed [Ere81c, EN85, Ere85, Gor79]. Developer [Cal00]. Developing [CB83a, Ros82a, Ali90, EH94, Gat92, Rob96a, SBA85, Par82c, Cur97]. Development [AM77, Art77, Har78b, Kil97, McI97, MK72, Nie76, Tho82, Wei97b, Wei97c, WA96, Yov70, Aik97, Arm94, AL98, BIFW71, Ber86b, Bie95, Bur00, BNP99, Car95a, CH98, DG95, Dos85, Dow77, EF96, Fre80, Gol74a, GSK98, Jah74, Kat72, Kaz75b, KC99, Ley97, Moo72, Osb98, RNBV98, Sal73, Sch97b, SS93, SGS70, Swa77a, Swa78a, VTS97, Wal75, WWSR00, WC92, WY73, WK75, Wil93b, Wil95b, YFR81a, vdBL96, HW71, Gul85, War79, Sam95, Wei88]. developmental [For99, PR98]. Developments [Ber86b, Dav86, Dic86, Gib86, LL86, Mea86, She86, Bru78, FBC79, NvR98, Phi92, Wil77b, Jer70, Vin74]. Devil [Onu82]. Deward [Fis79]. Dewey [Ric77, ODVG87]. diabetes [AP96a]. Diagnosing [RFK97]. diagnosis [GQSdP98, Mar72b]. dialectic [NM74]. Dialectical [MWR72]. Dialog [Bra79, SBW90]. dialogism [Sch97a]. dialogue [Lam00a, Spi94]. dialtone [Abe96b]. Dianne [Gro77]. diatribe [Cer79, Rob78]. Dictionaries [War96a, War96b, LH95a]. Dictionary [Joh84, Lan97, McC97, Jak88, SWK78, Tin70, McC96, Mea83b]. did [Hal85]. Didsbury [Cer83]. dies [Coc98]. Dietrich [Wel71]. Differences [LBW00, All00, CCM00, Che00a, Dav77a, Hur80, KM00b, MB81a, Laz00, Mea00]. different [Egg90a, Fen81, Goy84, Hal81, IS98, Kje74, PM00, Bel88]. differential [AM77, Kat72, Par82a]. difficult [Shi88]. difficulty [LP99]. Diffusion [Cha86, AP97, BS88, Gro76a, Oro77, Har78b]. Digital [Ano08, BNS^{+}00, FHN^{+}93, GL93, HOS^{+}93, KMG^{+}93, Lap99, LF93, Raw93, Wil00a, Alg95, Bal00, Bat98, Bro86a, CK96, DL00, DG00, EK97, Erc98, Fau96, GSKL00, GMPS00, HJD^{+}99, HCL^{+}00, KN00, Kin98b, KC99, LaF00, LDS^{+}00, Lyn94, Mea83b, Mil00, Par00, Ros83, San98a, San99, Sha86a, Sun99, TC00a, WH98, YCO00, Tur98, Wue97]. digital-images [Sun99]. digitized [McC00, Zen99]. Digress [MW91]. dilemma [CH99, Ell96]. Dilemmas [Cas92]. Dillion [Per85]. dimension [Mar99]. dimensional [Cle76]. dimensionality [KM73]. Dimensions [ABLS88, BDL88, Bel88, Cox88, Dos88, Gar88, Hay88, Hur88, LC88, Lyt88, MF88, OC88, Pre88, SM88, Woo88b, Buc91b, CK96, Cul85, HBJR99, Moo72, Sha98]. DINFOS [Ano84m]. dinosaurs [NC86]. directed [DKP00]. Direction [Hur88, Cro99b]. Directions [ABLS88, AM80, BDL88, Bel88, Cox88, Dos88, Gar88, Hay88, LC88, Lyt88, MF88, OC88, Pre88, SM88, Tho82, Woo88b, Lan87b, Mas97]. Directive [Lev73b]. directory [Boy89a, Hoc89, Kri86b, LR72, Sel71, Var78, Boy86d]. Dirty [Kug71]. disambiguation [BD92]. disaster [Smi74a]. Disciplinary [Haa97, Bri91]. Discipline [Hir78, Yov70, Los97a, SH94, WM98, Dav98]. disciplines [Bon90, EAW99, LBIP97, Los97b, RBR80, And99]. disclosing [Cox93]. discount [Ols74]. discourse [CM00, Day00b, LBKO87]. discourses [IS98]. Discovering [Sol97a, Sol97b, Sol97c, WLC00]. discovers [Bea92]. discovery [BG98, Des97, Exn98, GCH99, GL96, GD98, LG99, PB99, SNG00]. Discriminant [DF82, Car74]. discriminate [Fen81]. discrimination [CO87]. discs [Ali90, NC86]. Discussion [Bes96, HG00, New73, Wig95]. disease [BB83a, EM94, Koe88]. Disinformation [Pap96]. disjunct [EAW99]. Disk [Par83a, Bro86a, GP83, Ros83]. disorder [Sha83b]. disparate [Rod95]. dispatch [CH98]. dispersion [ODVG87]. display [Bur81a, Cra91, EM94, KL98b, Wol71]. displayed [FS84]. displaying [Ell76]. displays [Car96, Cra80a, Lin97]. disposition [RSK74]. dispositions [BM75]. dissatisfiers [ZvD00]. disseminating [McL79]. Dissemination [Lun80, Mar72c, Bac72, Coh99, FRGL75, Lan71, SMB72, ZK72]. disseminators [Bra87]. Dissertations [Hin70]. Dissipative [Liu96]. Distance [Bar96, BB96b, BD96, Bur96, Dav96a, Dou96, EQH00, Fau96, Lin96, Sut96, Bes96, Don96, Gas99, Rob96b, Sem98a, WA96, ZK99a, ZK99b]. Distilling [SMHC93]. Distinguished [Lei73]. Distortion [Die82b]. distortions [LCL85]. Distributed [CPW88, Cit98, LBP76, DL00, DG95, DRM98, HG98, KMG^{+}93, Mea82b, Mea86, MSHB86, Wee99, Zha98]. Distribution [Bel71, Dav92b, Fau96, Har75a, Hol96, Aji91, Aji92, Asa81a, Asa81b, Bur93, Coi77, ER00c, EN98, Gel79, GS87, MK79, McC89, Mor81, NT85b, OV92, OV97, PK78, Rao80, Rie88, ST89, She91, Sic85, Sic92]. distributions [Bas92, Ben00, Boo90a, Boo90b, Boo97, BG78, Bur94, Egg87, ER95, Gri88, Kan78, KG72]. Distributor [Ere80]. District [KL79, ST89]. diversity [MF88, Rou92b]. Division [BR94, Bel81, Har94, Lan94, LB94, Pet94, Sap94, WT94]. Do [Chr95, CC99]. doctoral [Cov00, Lee00b, Old91]. Document [Art71, Bar98a, Cro81, Koc83, MPW73, Oh98, Sha91a, BI97, BE80, BBMS87, BBSM87, Bla80, Bod99, Buc92b, Buc97, Cag70c, Cod73, CT87, Cro94, DG87, Dav77a, DW98, EB88, EC92, Fag89, Gir91, Gor91, GLW86, HZ80, Jan94, Les89, LDS^{+}00, Los88, Los00, LG75, MR87, MSHB86, Pai93a, Par83b, PW91, PZSD96, Ror99b, RF97, SKL97, Sch96a, SWS71, WS98, WW99, Wat99, Wel80c, Wil93a, Wil98a, Wil84, Mac79b, WS98, Bar98b, Rus85]. document-boundaries [Gir91]. document-retrieval [Los88]. documentacion [Gro76b]. Documentary [Mol95, Wal97, FS91b, KC99]. Documentation [Ano71a, Buc95, Ray97, Ano72b, FS97b, Fre77, FF80, GS78, GGJ91, Gir91, Gro76b, Hal80, HHC70, Koe75, Lit96, Ric88, RL97, Swa76a, Var75b, Vin74, Wei76, Wil97, YdKKM87, Eli87]. Documents [Vag80, BTSKC96, BCB95, BSR93, Cla75, Coo78, Cro77, EFW96, FK73a, HKE97, Hoo91, Hum99, Kin76, Kwo85, Lar99, Los92, MQP98, O’C70, PJ90, RB89, Ror90, Ror93, Ros71b, Sal70, Sma73, Tho73a, TC00b, WW00, WRE^{+}70, Win70a, WP94, WNAR99, YR98, Yao95]. Does [Jon73a, PJ90, BF93, Erc98]. dogma [Fro96]. Dokumentation [Har76b, Har77b, Kuc76b, Gro76b, Wel80a]. Dolby [Har70b]. dollars [Hen99]. Domain [Beg95, Bat98, Coh95, Pal99a, Sca89, HA95, Wal97]. Domain-analysis [HA95]. domain-independent [Coh95]. domain-specific [Sca89]. dominant [AdKK87]. dominates [Die81]. Don [Jor89]. Donald [Art77, Lan73, Mar85, Pre71a]. Donohue [Ano75a]. don’t [Wat98]. doomed [Gas99, Gas99]. Dorothy [Kuc76b]. Double [Boo72, Mea84b]. Dougherty [Kuc75]. Dr. [Her96a]. drawing [BR95a]. dream [JB85]. Dreams [Hen95, Cas98a, Smi97]. Drexel [Wel75, Flo00, Woo88b]. drill [BMFS84]. driver [Has98]. drowning [EB98]. Drug [Ano84j]. DSSSL [Adl97]. dual [Egg86, Tin70]. Dublin [GSKL00]. Dugan [Ste00a]. Dumain [Hol87]. Duplication [Cor82, NL83]. during [BDBR97, Col99, IS98, Mic94, Rob00, Spi97b, WS98, WW99]. Dutch [Voo99]. duties [Cre95a]. dyads [Mar92b]. Dynamic [CCT95, Bur93, Che95b, HBJR99, KH71, NvR98, OV92, OV95, PN96, Art76a]. dynamical [Kir93, BMvR91b]. dynamics [Die84b, Lin78, Rou74a]. dynamism [Bru94]. dyslexia [PR98].

E-conferences [KRD95]. E-journal [HF00]. E-journals [CO95]. E-mail [HBG96, May97]. E. [Bra92]. each [Mim85]. earliest [Wei97b]. Early [Bro87, JBS^{+}77, Kol75, Kuh99, Hei88]. easy [Mim85]. EATPUT [Wan92]. ecological [Lee96b, Per85, Wis99]. ecology [Lee96a]. Economic [Wes91, Zai77, Ber72a, Bra77, DK90, Gre92, Kin98a, Ren80, Sny98, Wei88]. economical [SSM93]. Economics [Bra89, Coz82, Bag81, Bon79, Bou80, Bro70b, Day99a, Gan74, Koc72, McC91, Rep89, Sle94, Sny98, Var97, Gan74, Tom98a]. economist [Mac79a]. Economy [Rei77]. Ed [Ano74a, Art77, Cer83, Dan84, Her82, Kaz77, Kun82, Mat85, Mim85, Ric77, Shi85, Swa76b, Swa77b, Swa78b, Wei79, Wil85b, Bea97, Bla79, Col98, Dav79b, Dom78, Har83, Har76b, Har76a, Har77c, Jen95a, Kri87, Kuc76a, Lan78, Nel97a, O’C95, Pen93, Pet96a, Sch73b, Swa76c, Swa76a, Wei94b, Wei95b, Wei76, Wel78, Zwe95]. ed. [Gro87, How87, Tom98b, Cob96]. edge [War98b]. edges [Goe87, GV87, O’H70]. Edison [Bez82]. Edited [Gro75b, Har75d, Afi99, Che99, Don74, Lam99, Lap00, Mar75, Pet99]. Edition [Sel71, Har77a, Kuc76a, Ten94]. Editions [Don74, Har76c, Mau71]. Editor [Deb72, Dro78, Hin70, Lil78, Mea78b, Pre70, Sel71, WH70, Mea84a, SH89, Wel75, Aiy88, Aji92, AI76, Ano74d, Ano77b, Ano78a, Ano78b, Ano81c, Ano82b, Ath72, Bag81, Bal84, Bal88, Bas94, Bat99a, Bat00, Bat83, Bel77, Ber72b, BB77, Boh90, BSR94, Bor97, Bor87, Bra92, BL72, Bri72, Bro79, Bro90a, Bro00e, Bru78, Buc91a, Buc92a, Bun80, Bur93, Bur94, Byr76, Can95, Cha77, Che95b, Coo91a, Coo76, Cou98, Cox83, Cra72, CRC70, Cur93, Dan00, Dav92b, Dav79c, De 80, Deb92, DO81, Dic96, DH78, Din87, DG78b, Dum87, Dun80, Dun96a, ER97, Eve98, Fei73, Fid87, Flo99b, Fre80, FB81, Gar79, Gar83, Gar91, GS97, Gar97, Gar98, GD90, GT70, Gli78, Gol87, Gor92, Gri80, Gro78, Hal95, Han92, Har93b, HN98a, HN98b]. Editor [Haw87, Hen97, Hen00b, Her96a, HD72, Hir90, Hjø96, Hjø00b, Hol87, Hor78, Hor93, Hub77, Hum96, Hum00, Hun72, Jac81, Jor89, KW81, Kno77, KM78, Koc80, Koc84b, KBN90, Koe81, Koe88, Koe92a, Koe92b, Koe94, Kra85, Kra86a, Kra91c, Kra92c, Kri77, Kuc80, Kuv73, Lan81a, Lan89b, Lan97, Lan98a, Lan98b, Lar92a, Lev73b, LP83, Lev98, Ley98, Lin77a, Lin77b, LS80b, Lin87b, Lip83, Lon92, Lou91, Lou92b, Luc93, Luk96, LD97, Mac72, Mac97, Mar78, Mar72b, Mar72c, Mar88, McC97, McM73, Mea78a, Mea79b, Mea79d, Mea79c, Mea79a, Mea80, MZ80, Mea81, Mea82d, Mea82c, MKDL83, Mea83a, Mea83d, Mea84b, Mea84a, Mea87, Mea97, Mea98, Men95b, Mig80b, Mil87, Mil95, Mos72, MOV83, Now81, Old88, OV93, OV95, OG80, Opp98, Ozo80, Pap88, Pap93]. Editor [Pap96, Pap98, Pat81, PTH83, PLM^{+}79, Pri72, Qin00a, Ric97, RHH^{+}70, Rot81, RV98, SC80, Sal87a, Sal96, San79b, San83, SM78, Sch91b, SLW70, Sha84, Shc79, Shp92, SB80, SG89, Sor96, Spe76, Ste84, Stu78, Swa89c, Swa89d, Tau84, Tem94, TZ93, UAW82, Van98a, Van80, Vea90, Vin99, Wal72, Wei90, WM78, Wel81, Wel87, Whi94, Whi96, WM00b, Wil70, Win95, Wit80, Wit87, WA97, Woo76b, Wyl87, Yer91, Zhi97, vB81]. Editor-in-Chief [Dro78]. Editorial [Bat99a, Bor87, Boy99a, Boy99b, Boy99c, Boy99d, Boy99e, Boy99f, Boy99g, Boy99h, Boy99i, Boy99j, Boy00a, Boy00b, Boy00c, Boy00d, Boy00e, Cas95, Che00b, Che00c, Eli70a, Eli70b, Eli70c, Hur00a, Kra86b, Kra87a, Kra87b, Kra87c, Kra88b, Kra88a, Kra89a, Kra89b, Kra90, Kra91a, Kra91b, Kra91c, Kra92a, Kra92b, Kra93, Kra94a, Kra94b, Kra97, Kra99b, Mea94, Ric95, She72, Shi88, Sim83, Wal87, Wel00, Har71c, Mea81, MKDL83, Sta85]. Editorials [Ano00c, Ano00b, Boy99k, Boy99l, Boy00f, Boy00g, CCM00, Kra99c]. Editors [Gan74, Ols74, Cam96b, Pre71a, Pre71c, Ain85, Ano74f, Ano75f, Ano75g, Ano86e, Aul79, BHCC^{+}74, Bar74, BA74, Bez82, BZ85, Bor85a, Bor74, BK79, BPK^{+}71, Bur71, Con84, CHPC71, Dav79a, Dav74, Dav84, Die79, Gol74a, Gol79, Hen00a, Jon73b, KBR^{+}71, Ker84, Kol74, Kra74a, Kra85, Kra74b, Laz00, LHM^{+}86, Mea84c, Mea85, Mea00, Mur74, Nad85, NCH^{+}71, Opp74, Pap84, Pit82, Pom74, Pom75, Pul74, RC74, Ros82b, Rot79, Row84, RTH71, Sal79a, Sal79b, Sal84, San75a, San79a, SB73, Sew82, Str84, TL74, Tho79, WC85, WS75, Zar82]. EDP [Wel71]. EDS [SMHC93, Bel81, Boy85a, Gro77, Har77b, Hei78, Mar85, Pat85, Rei77, Swa77c, Gan74, Mar77b, McM76]. Educating [Cox88]. Education [ABLS88, BDL88, Bal89, BDZL89, Bel88, Bel70, BB96b, BD96, BR80a, Bur70, But99, Cox88, Cre99b, Cre99a, CH99, Day99a, Dos88, Dou96, Dun80, Fra99, Gar88, Gas99, Hay88, Hur88, Kaz77, Lev99, LC88, LB89, LCB99, Lyt88, MF88, OC88, Pre88, RM99, Rob80b, Rus70a, Rus70b, SM88, Shu80, Ste99b, Sun99, Swa78c, Woo88b, AL98, Bar96, Bor84, Cro99b, Don96, Hei87, Kaz75b, Kaz75a, KBN90, Mil00, Old91, Rob96b, RH97, SBA85, Var99, Whi89a, WA96, ZvBR88, Hay88, Kaz77, Lev76a, Har71b, Kaz79, Ols89, Pat90, Whi89b, Lan87b]. education/exploration [Mil00]. educational [MF88, Sut99, Wri73, LAT79, Bix96a, Bix96b]. educators [Mat85, SMC83, Sun99]. Edward [Fis81a, Lan73, Lev76b, Wri71a]. Effect [LBBR96, Lev74, Pal87, Bam79, Cod73, Dav77a, FS84, Han86, RR78, RW82, Sal72b, SMC83]. Effecting [Kaz77]. Effective [PM00, Cal00, Car88, CK96, DCRS98, Har91, JB71b, Sha96, Mai90]. Effectiveness [ASS73, BB96b, Dou96, Kil95b, Low71, RSW91, BSY00, BCM^{+}70, BM75, BP70, Bur95, Byr75, CK99, Coo73a, Coo73b, Coo83c, Elc78, EFW96, EK75, Fag89, Gor90, Har96, KMB99, Lan71, Mar83a, NK00, PW89, Par00, Ras90, Ro88a, Ro88b, RT81, RB72, SK88a, Sch00a, SS97, TK96, Woo70, Yao95, Yer73, PW92]. Effects [Bre79, HY93, KS70, BW85, EB88, KL98b, KL98c, Kir93, Nug91, TBCE94, Xia92, Stu72]. efficacy [New94]. efficiency [LG96, LT74, Mil71, RSW91, Wil93b]. Efficient [BM95, Goy87]. effort [Chr95]. Efforts [vB81, BR80a, BB83b, CA80]. EFTs [Lan82a]. Egyptian [Kil93b]. eight [Eli87]. Einsatz [Wel71]. elderly [MY85]. Electric [War96a, War96b]. Electronic [Ber94, CB94, Cre99a, DB94, Gol94, Jac94, Jas94, Kra94a, LP94, Lyn94, Pee94, Pee96b, Sch94, SD96, Sta85, SH94, TH82, WSB98, Abe96a, ALD96, Afi98, Bro99c, Buc92b, Car95a, Cas85, CYFS95, Cov00, Gro88, HDRD88, Har98, HP00, Hay96b, Her95, Hil96a, Hil96b, Hir99, HOS^{+}93, Jas97, Kim00, KM99, KM00b, Koe92a, KH95, Lan89a, LG96, Lie96a, LAD97, LLM91, Lyn91a, Mar89, MR94, OCN97, Pel93, PTH83, Pri83, RMMS96, Sha83a, Sha97b, SPWH99, SRS98, TK96, Wee99, Wei94a, Wel00, Woo88a, GL93, Afi99, Bro97, Lan87a, Vag80]. Electronics [BB83a, Koe88, Mea83b]. elektronischer [Wel71]. element [WL81]. elementary [GSKL00, SH89]. Elements [Elc73, MK79, Sho71, Zen99, Sha86a]. eleven [ST86, Mic94]. elimination [BDP98, TEC99]. elitism [Ami73, de 71]. Elizabeth [Sta75a]. Ellis [Mar85, Ols76]. Ellsworth [Gro77]. Elmer [Har77a]. Elmsford [Ano74a]. else [Gor92]. Elsevier [Koc79]. Elton [Woo77b]. elusive [Bas94, OV93]. ELVYN [RMMS96]. Emanuel [Buc92b, Buc92c]. embodied [Mac96]. Emergence [AdKK87, DRM98, Gra82]. emergency [Mon94, WWSR00]. Emerging [Pet96b, Pet96c, Dav98, DC98, How89c, Mar99, Pel93, PKK74, Smi96a, Smi96b, Wel00]. Emergy [Woo77b]. Emil [Ere81a]. emphasizes [Ano70b]. Empirical [SEW00, DG78a, HST96, LUH^{+}96, Lev76b, Nic89, Oh98, OV98, Pao86a, Pao86b, Par82c, Per85]. empirically [Wil92b]. employers [Pre88]. Employing [Kwo99, FK73a]. employment [Mar98]. enabling [Col99, Col00a, CM00]. enacted [Spa97]. encoding [BI97, Col00a, FCJ^{+}92, LY91]. encodings [CL82a]. encountered [BDBR97]. Encounters [Slo00]. encryption [Pet98b]. Encyclopedia [Dro78, Mar89, Dro78, Ble86, Dan84, Kri87]. End [RW00, SBW90, Yua97, BCM89, CK99, Goo97, GLS97, Han86, HC96, LS97, Mar92a, Mar95, MR81a, MR81b, NT96, Oja86, ST86, Sha90, SBW93, Spi96, SEW00]. End-user [Yua97, CK99, Han86, HC96, Mar92a, MR81a, MR81b, Oja86, ST86, SBW93, Spi96, SEW00]. End-users [SBW90, Goo97, GLS97, LS97, NT96]. ends [Har92a, Wil86]. enduring [NO85]. Energy [Art74, Ger79, Nie76, Van80, Ano70b, BT00, BCM89, FBC79]. enfants [Kra91a]. enforcement [Ere81a]. engine [Bil00, FFKS99, LP99, RW00, vdVM99]. Engineering [Ano74a, WA96, ALD96, CMK98b, LAD97, May78, MN83b, Ols76, PvR94, PBvR95, Shi86b, Yer71a, Byr76]. Engineers [Pre71b, Pre71a]. engines [Her99, LS99, Sch98b, Sch98a]. England [Kra79, Mea82b, Mea83b]. Englewood [Art76a, Dav79b, Dom78, Fis81a, Fis81b, Har83, Har76c, Mar80, New76, Str78]. English [He98, He00, Mea82b, Mus86, SM91, Str74]. English-to-Chinese [He98]. enhancing [LBBR96]. Ensuring [Cur97]. Entering [Koe92b, Koe92c]. enterprise [Bar98b]. Entertainment [Myb98a]. entitled [Aiy88]. Entries [RG70, SJ72, SR77]. Entropy [SD83, Bal84, Coo83c, KL98a]. Entry [BM89, DC98]. enumerative [Dav76b]. Environment [BDBR97, BD97, BDSA97, D’E97, Fed82a, Goo97, GLS97, LS97, LH97, RH97, SA97, Aga99, Bes96, BM72, Bur85, Che00a, CS91, Dav86, Eft00, Gib86, HH97, HJD00, Jon73c, Jon74, LaF00, LBBR96, LWG74, MR94, Old91, Pas90, PN96, RF97, SH96, ZK99b, Tho73b, Mad96a, Mad96b, Sta98]. Environmental [AC93, DC76, LMSC82, MN83b, Mon94, Ste99a, SS00]. Environments [ABB^{+}91a, ABB^{+}91b, BS91b, BF91a, CS91, DRCG91, LD91, RRB91, SB91, CCM00, For00, MBH94, Mar99, McD99, Mil00, MSHB86, Sca89, Wee99, Hil96a, Hil96b, HC94]. Envision [FHN^{+}93]. Envisioning [Mil93]. epidemic [Wor73]. Epidemiology [BN76]. Epilogue [NK89]. Epistemological [Hjø00b, Kha91]. equality [RCJS94]. equations [Coo70a, Hei74]. equipment [Agh86]. equity [Doc91]. equivalence [Ang75, DF85, Dor85, Dor88]. era [BB96a, Lan89a, Afs99, Alg95, Car95a, Mur93]. Ergonomic [Agh86]. Ergonomics [Has98]. ERIC [AM77, Kue72, Wri73, Pai71, Bur00, Mar72b]. ERIC-like [Kue72, Pai71]. ERICS [Bur00]. Errata [Ano70c, Ano71b, Ano72d, Ano73a, Ano74e, Ano76a, Ano81a, Ano87, Coo91a, Kra88c, Swa78a]. Erratum [Ano79a, Ano81b, Ano82c, Ano83a, Ano84a, Ano84b, Ano84c, Ano86d, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano96a, Ano96b, Ano97, Ano99a, Ano99b, Ano99c, Ano00d, Ano08, Koe92b, Kra91d, Lar92a, Wel85, Wil85a, Wit87]. Error [Lon92, Lov71, Lou92a, Pai96, RFK97, Swa89b]. Errors [PZ84, Zam80, Bro70c, PZ83]. ESPRIT [Sme87, Sme92]. essay [DB94]. Essays [Cor94b, Ben86, Dro77, Mar75]. Essential [Mal99]. Essentials [Yin95]. Essex [Mea82b]. Establishes [Lev73b]. Establishing [ABLS88, SK97a]. Estimating [Cas78, SSD99, War73, Van98b]. Estimation [Nic87, ST89, Bon79, Bru94, EM94, Los88]. ETC [Kaz75a]. Ethics [Ber85, Koc87, Cox93, Day00b, Bat97, Pet99, Dil98]. ethnography [CNO^{+}00]. Ethnomethodologically [CNO^{+}00]. Eugene [Dro77, Lev76b]. Euronet [Bre79]. Europe [Ing97, Jon89, Phi92, Sme92, Wil85b]. European [Mea82c, AL98, Bax85, DRM98, Ell86, Ens95, For91, Ing97, Jon89, KDJ94, MP98, Phi92, Sme87, Sme92, Wil82, Woo84]. evaluate [Hei90, Zwe95]. Evaluating [BRR96, Caw78, Gor90, HMNW86, JM97, Los96c, McC88, NJWC97, Pou89, SBPP89, Abe90, Coo72, DC80, GWDS86, Har79]. Evaluation [Abe90, BDBR97, BD97, BDSA97, Bro70a, CS90, D’E97, Gar90, Gol73, Goo97, GLS97, Kei70, KB78, LUH^{+}96, Lan73, Lar92b, LS97, LH97, LC90, MP74, RH97, SA97, Vir70, Voi97, WYSB91, AL94, BSY00, BWV72, Bla96, Bon79, BP93, BCM89, Boy86a, Bur99, Car96, Car74, Car75, Dun00, Eft00, ERV00, Gor82, HS78, Her94, HH95b, HH95c, HPH96, HCL^{+}00, Hul96, KB98a, Kir93, LS80b, Lip83, Liu97b, Lov71, MS00, Pai96, Pen93, Ro88a, Ro88b, RF00, Ros71a, Sal72b, Sha86c, SCLL73, Su94, SM75, TS96, Tho72, WSQ94, WC77, WSS79, YS93, ZvD00, Fay72b, Her82, LAT79, MHA82b, Oh98, Pat85, Swa78b, She91]. evaluations [CHSS98, MR81b]. evaluative [NML99]. Evangelist [SS70]. Evans [War79]. even [Kra90]. Everett [ER97]. Everybody [Has95a]. everyday [BJ79]. Evidence [Bro86b, DDC78, LB83, Liu97c, Sko98, Soe94]. evidential [Kas96a, Kas96b]. Evolution [EN98, SGS70, DHS00, Hal80, Har71a, Kan76a, Sch81, Wol87]. Evolutionary [Sta82]. evolved [Cas94]. Evolving [Woo88b, Bec84, Lyn98]. ex [LFS72]. Exact [Egg93, Gri88, Mor81]. examination [Boy85a, CK99, Der94, DG78a, GL91, GL96, Hur80, Pao86a, Pao86b]. examined [Ano74a]. Examining [Col81, FBC79]. example [Caw74, EM94, KDJ94, Kuc82, MMB99, Swa89a]. Examples [Fau96, Hol96, CSSI98, YFR81b, FB78]. exchange [Ano70b, HW98, Zal74]. Excite [RW00]. exclusivity [Jon73c]. exercise [Blo70, Cli70, HM70, Lin70, Mat70, OG70]. exercises [New73]. exhaustivity [Bur95, HW80, Jon73a]. exhibit [Mar00]. existence [Koe93]. Expanding [Fed77]. expansion [AL94, Dav76b, Eft00, PW91]. Expectations [Bra87, Ken87a, Lin87a, LC87, MH87, Pal87, Spr87, WCL87, Wes87, Wol87]. expected [FGS81]. experience [GS93c, Han86, HH95a, HY93, IT89, Kuh99, PK00, WMME88, Ker74, Lev76b, Cor94b]. experienced [Con95, HY93, Laz00, LBW00, Mea00]. Experiences [Wes88, Fen81, SFB^{+}00]. Experiment [Dil82, CNMS97, CSSI98, CCN93, DMZL82, Jas94, KE99, McC91, NM74, O’H95, SMB72, Sve72, TB93, Whi81, Woo76a]. experimental [BWV72, BP70, Col92, Lar92b, Mar83a, Reg88, Woo76a, Bot73, CB73]. experimental-psychology [Col92]. Experimentation [HHC70, Coo78]. Experiments [AYS96, Lar92a, Lar92c, WEWP97, YF85, All00, Lan98a, Lan98b, MPW73, RF00, YFR81a, YFR81b, YFR81c, PN96]. Expert [YM84, ZC92, Erc92, Fid86, FW90, GS93b, Har93a, HO89, Koe83, LK91, Tes86, TF89, Wel87, Wil94, Bor90, Dog98]. Experties [Bar74]. Expertise [BF91b, DS96, Col00b, Nor96]. experts [Wie86, Wil98a]. explained [de 78]. explanation [Egg94]. explanations [Sko98]. Explicit [Rob75, BCB95]. Exploiting [Coo83c, GCH99, Wol71]. exploration [Bal00, CN95, OOA99, Ror99b, Ror99a, WV97, Whi91, Yan97]. Explorations [Exn97]. exploratory [Aga99, Bar94, DG91, GB95, Spi96, Sut94]. Exploring [Cas92, Sch00a, Tra98]. Exponential [Gup90]. expressed [Moo73]. expressions [Der84]. Extended [GGJ91, Gir91, Lan80, BFM88, FW90]. Extending [MBH94, RFK97, SD97]. extensible [BI97]. extension [Fly97, Glu90, GL96, Wu98]. extensions [LY98]. extent [JJ77, Tau74]. External [YdKKM87]. extraction [BM75, Wei96a]. extracts [Hel71]. eye [Mur93].

F [Ano74b, Art75b, Cer83, Die84a, Her82, Kaz77, Lil78, Mai99, Pre71b, Swa76b, Swa77b, Swa78b]. F. [Sal72a, Woo87a]. face [Bar78]. facilities [Sol74]. facility [CH98, WT94]. fact [Bil00]. fact-based [Bil00]. facto [LFS72]. factor [Asa81a, Bar95, DOO^{+}96, DK75, HN97, ZvD00]. Factors [ALD96, Bat77, FS83, LAD97, Bat98, Cam96b, CBM^{+}94, EK75, GCS97, Gar98, HN98b, Hir78, MB71, McI97, Pet96b, Pet96c, RV98, Tau74, Bur00, Hig99, MV95, RV96]. facts [Oh98, OV98]. factual [OWF99]. Faculty [SPWH99, ALD96, BC97, LBIP97, LS97, LAD97, MS00, SKvB92, SvB94]. fails [Gre95]. Failure [Coo91b, DW96, KB98a, Wri71b]. Fair [But99, Cre99b, Cre99a, CH99, Day99a, Fra99, Gas99, HC84, Lev99, LCB99, RM99, Ste99b, Sun99]. Fair-Use [But99, Cre99b, Cre99a, CH99, Day99a, Fra99, Gas99, Lev99, LCB99, RM99, Ste99b, Sun99]. fairest [She72]. Fairthorne [WS71]. Falcolm [Har75c]. Fall [Ano74a, Lil78, Rei85]. Fall/ Winter [Rei85]. family [Bla71, Lea76]. FAMULUS [Bur81b]. far [OV93]. Farber [Har75d]. Farewell [Har77c]. fashion [Zen99]. FASIT [BD92, DG83]. fate [Spa97]. Fayen [Ano74b]. feasibility [Sal72d, Tan96]. Feature [BM80, DCRS98, Mea82c]. features [BLPG99, Dio94, Egg94, Fay72b]. February [Cra96]. federal [Ain84, Ber82, Bra87, Pal87, WWSR00, Bau72, Bra87, Bur70, FW99, Ken87a, Kos94, Lin87a, LC87, MH87, Nad85, Pal87, Spr87, WCL87, Wes87, Wol87, Ric89]. Federation [Lea76, Ray97, Swa76c, Swa76a, Var75b]. Fee [Boy86b]. Fee-based [Boy86b]. feedback [AdKK87, BCB98, CSSI98, Eft00, Fed77, FS91b, FSV93, LFHG99, SFV85, SB90, Spi97a, Spi97b]. Felix [Tau74]. Felleman [Ste99a]. fellow [Lyt88]. FEMA [WWSR00]. Fernandes [Fis79]. fertilization [FS97b]. few [Kra88b]. Fibrosis [Sha91a]. fiction [Beg95, Bie95, McK96a, McK96b]. FID [Ano71a, Hol71, Ray97, Sch73b, Swa76a, Var75b, Ano71a, Bor73a, FS97b, Har71c, Hol71, Jac71]. FID/II [Hol71, Var75b]. FID/LD [Ano71a]. Field [Vir70, AP96a, HBJR99, Pai71, Wei88, WH91, Wit96a, Wit96b, KM00b]. field-based [Pai71]. fields [Aul87, Lee00b, Len83, PS79]. Fifty [Eli87, Var99]. File [Elc73, Her97, Hum84, Blo70, Cli70, Cro79, Har71d, HM70, Hol70, KSDR90, Lin70, LS70, MS76, Mat70, OG70, Ras80, Sch74, Wei94b]. files [BR95a, Cro77, Dod79, FPS00, Kje74, MB95, WSQ94, Whi77, Gol74b, Boy86a]. filing [Gro76c, Wel83]. Film [McM76]. Filmdex [McM76]. filter [HST96, YS93]. filter/flow [YS93]. Filtered [PZSD96]. Filtering [MQP98, McC95]. Final [War79, Ric70]. Financial [Rus70a, Tal79]. Financing [Bos87]. find [BBDH00, DCD81, Ols75, Wat98]. finding [Aul87, Haz98, Tin70, GLL90, Hun96a, Hun96b]. finding-keys [Tin70]. findings [Tyn75, WD97b]. Fine [Bes91]. finite [KW97]. fire [Pic86]. firms [Hoc89]. First [Gul85, LS99, Sal86, Wil98b, Art74, Bue86b, Bur00, DG00, Dog98, Gol97, HW99, Mil90, Ten88, Dat79, Sut96, Bla93]. fish [GL96]. fisheries [Mar92b]. fit [KG72]. fits [Gri88]. Fitting [CMB^{+}95]. five [Bur95, Fug85, Lip99]. five-axiom [Fug85]. fixed [CL82a]. fixed-length [CL82a]. Flaming [Kay98]. Flanders [VHdB^{+}98]. Flexible [BI97, Dat79]. Flores [Dom78]. flow [And99, Col99, Die79, KET99, Liu97c, Vei80, Wel72, WY73, YFR81a, YFR81b, YFR81c, SWT^{+}79]. Flowchart [Mor76, Cha77]. Flowcharts [Wel83]. flows [FB78]. Fluctuations [San79b]. Fluid [Col96]. focus [Rob00]. Force [Ano72a, Til70]. Forecast [Man82a]. Forecasting [Klo89, KR80]. foreign [Ker74, Sal70]. foreign-language [Ker74]. forestry [SS00]. Foreword [Don77, Rus70c]. Forget [Lea76]. Forlaget [Ere80]. form [FS91a, Til70]. Formal [BLPG99, Cra72, LG75, Stu72, Wee92]. formalism [FCJ^{+}92]. Format [Gro76a, Mar71, GS97, GBDH76, Kra88a, Kuc76a, Win70a, Bor85a, Rou89b]. formats [Blu89b]. Formatted [Eas91, DGS86]. Forms [Swa75b, Ang75, Hub78, Tau74]. formula [Has76, War79]. formulation [Asa81b, Hom83, NE90, PK78, WY90b]. formulations [FS91a, Rad82a, Rad82b, SBF83]. Forsyth [Har70b]. Forthcoming [Ano78c]. Foskett [Lan78]. foster [FW90]. found [Fag95, OMY^{+}89]. Foundation [Fed82a, BS75, Sha86c]. Foundational [Con93]. Foundations [RH99, Afs99, DPD99, DH97a, Kim73, Roy90b, SS95, Hol99, Aik97]. four [Bon90, HG00, Leh72, BWKG75]. fourteen [Kra97]. fractal [Ott94]. Fractional [BR95b, Opp98, Van97b, Egg96]. Fraley [Pat85]. Framework [CB73, Alg95, BF91b, CP98, DHS00, FS83, Liu97c, PKK74, Sar75, SK97a, Spi97a, Sut99, Wat89]. France [Ger79, Wei76]. Francis [Lan77a, Rou76a]. Frank [Mea83b]. fraud [KB92, LaF00]. Fred [Hin70, McM76]. Frederick [Har76a]. free [FK73a, Fid91b, Sch90, TK96]. Free-Nets [TK96]. free-text [Fid91b]. Freedman [Bel81]. freedom [Bue86b, Cre76, Die87]. freestyle [LP99]. French [FS97b, Sav93, Sav99]. frequencies [Mea82b]. Frequency [Bel71, BG78, Hoo91, Bur94, CL90, Coi77, LJ70a, MSPPM98, PZSD96, ST89, Sic92, SSD99, Sve72, LCL85]. Frequency-imposed [LCL85]. Frequency-rank [BG78]. frequency-sorted [PZSD96]. Frieder [Wil00c]. friendliness [Kri86c]. friendly [Har92a]. friendship [HW98]. Fril [Kas96a, Kas96b]. Fritz [Ano77a, Kin83]. front [BCM89, Har92a, Mec99, SBW90, Wil86]. front-end [BCM89, SBW90]. front-ends [Har92a]. frontier [Sha97b, Jas97]. fronts [DS91, Per94]. Frustration [Mel72]. Fry [Ano74a, Dav94]. Full [GD90, SMJ95, Sie96, BMNM^{+}93, BW84, CCW84, EFW96, Mar89, NT96, Par00, Ro88a, Ro88b, RW82, Wei96a, WNAR99, WT95, Lan98a, Lan98b]. Full-text [GD90, SMJ95, Sie96, BMNM^{+}93, CCW84, EFW96, Mar89, NT96, Par00, Ro88a, Ro88b, Wei96a, WNAR99]. fully [DG83]. Function [DS79, Bur95, Car74, EM94, Gal91, GLL90, KdK90, Par82b, Swa77a, Swa78a, dSM93]. functional [AKS96, BDP98, Hoo91]. Functions [Blu89a, Wil71, Mar83b, SD97, Tho73a, Yu76]. fundamental [Bro98c, Col96, IC97]. funding [Hai97, Mec99]. Furth [JK91]. further [Kra90, Whi91]. fusion [NK00, YR98]. Fussler [You75, Rou88]. Future [Afs99, Coo77a, Fox86, Kun82, Tho82, Ano78j, Bel81, BCSZ79, Cer83, Cox88, Dav96a, Dic86, Fro96, Gar88, GSK98, Koc81, LN82, Les96, Pat90, San98b, Shi85, Smi99, Swa88, War73, War98b, Wes88, Hen95, Lan87a]. futures [San98b, Wes88]. fuzz [Cer79, Rob78]. Fuzzy [Boo80, BDP98, KS90, PT87, BP93, Bue85, CVK92, CP98, Jon76, Kas96a, Kas96b, KZC86, MBVL99, RT81, Sac76, TO84, Yin95, Rob79, Tes86].

G [Ano74b, Die84a, Fis81b, Koc79, Ols75, Pec85, Ray99]. g007702305 [War79]. gain [AL84]. gained [Ker74]. Gale [Bol86b, Mea82b, Var78]. Galland [Mea83b]. game [CV80]. games [de 78]. Gamma [Bur90]. Gamma-Poisson [Bur90]. gap [SJ00, Tyn75]. Garfield [Bra92]. Garvin [Har70a]. Gary [Fis79, Rus77]. gasification [De 80, FBC79]. gatekeepers [Aga99, Glo88, KM95]. gateways [Fu96a, Fu96b, Wil86, Afs98]. gedanken [Coo78]. Gellatly [Shi85]. gender [BW85, DH97b, MBS98]. General [Har77b, Swa76c, AdKK87, Asa81b, CVK92, DF84, HM70, LT74, Mer71, OV97, Sav99, War73, Wil95a, de 76, Kei70]. generality [Sal72b]. generalizations [Fal98]. Generalized [ER95, Elc73, Ger79, Mor76, Egg93, WY73, WY90a, YK93, Ger79]. generalizing [BP93, Sri90]. generated [Cra80a, FK73a, Tag72]. generates [BT00]. generating [OV98, Str74]. generation [CYFS95, CELY80, Cor94a, Cov00, Cra82, Dog98, FWR95, Fox86, Hom83, LMO^{+}96, Lyn77, MB95, WC75]. generator [Soe72]. generic [SNG00]. genetic [Che95a, CSSI98, Gor91, MBVL99, WLC00]. genre [DG00, McK96a, McK96b]. Genreflecting [McK96a, McK96b]. Genres [DG00]. gentle [GBL89]. geochemistry [Ste85]. Geochimica [Lif93]. Geoffrey [Bro00c]. geographic [CD94, RCZS99, WP94]. geographical [Dor88, ZC00]. Geological [MK72]. geometric [JF87]. Geopolitics [Die81]. Geoscience [Koh74]. Gerald [Sie70, CML^{+}96]. Geraldine [Lam00a]. Gerard [Art76a]. Gerhard [Mig80a]. German [Lev76a, Tay95, Hap99]. Gernot [Gro76b]. gestalt [Lev76b]. Get [Ber72b]. Getting [Fag95, Ste00a]. Getty [SBW93, BS92, Hor93]. gIBIS [CB89]. Gifford [War79]. gigabyte [HC90]. GIGP [BF93]. GILS [MMB99]. Gini [Car79]. GIPSY [WP94, MK72]. girls [Tod99]. GIS [Arm94]. GIS-based [Arm94]. given [Ere81a, Los96c, War79, Ere81a]. Giving [BR94]. Glenford [Dav77b]. Global [Jac94, Gul87, SLC98, SL99, War98a, Wue97, Has95b, Lee96a]. globally [DL00]. Gloria [Pre71c]. glossary [Usd85, PJ87]. Glyn [War79]. go [Aul79, Wri79]. goal [AdKK87]. goals [Her96b]. going [Pee94]. Goldberg [Buc92b, Buc92c]. Goldmann [Sal87a]. good [Bro00f, Kra86b]. goodness [KG72]. Goods [Lev82]. Gordian [DM99]. Gordon [Ere81b, JJ75]. Gore [Har77c]. governance [Lil78]. Government [Ken87a, Spr99, Ste00a, Wel75, Bol92, DPD99, Haw90, Hun96a, Hun96b, MMB99, Mec99, Nel97a, SS93, Bra87, Bur70, KJR99, Ken87a, Lin87a, LC87, MH87, Pal87, Spr87, WCL87, Wes87, Wol87, Ano74a]. GPO [Bra87]. Graduate [Gro75b, Har76a, Ric77, Swa76b, Whi96]. Grainger [Che99]. grammatical [Sav93]. grammers [Sag75]. Grant [War79, Wes88]. graph [Asa81b, AM70, LS87, Sha83b]. graph-oriented [Asa81b]. Graphic [Cra91, Cra80a]. Graphical [LLM91, BC99, McD99, OCN97, YS93]. graphics [PJ87, Wil98c, Vag80]. graphs [Ell76, GH96, Goe87, GV87, Sha85]. gratification [Cha91]. Graves [Gro96, McM76]. Gray [Los92]. Greek [Don77, Myl92]. Greeley [DC76]. Greenwood [Ano77a, Ber84, Ere81b, Har77c, Koe75, Kun82, Ols76, Shu78]. Greven [Wel71]. grid [LS92a]. Gries [Dav76a]. Griffith [WM00a]. Grosch [ATJ95]. Grossman [Wil00c]. grounded [Wil92b]. grounds [Kha91]. Group [Gro76a, Arm94, Bat83, Wil98a, Zha98, RNBV98, Bro95a]. grouping [MDS70]. groups [FK73a, Gra82, Sha98]. groupware [Sha98, Nan95b]. groves [Whi87]. growing [Mar72a, Yer71b]. grows [Pee96b]. Growth [Har77c, Lin78, Bau72, Bro97, Eft90, ER00a, For99, Har70b, KH71, Men92, Ste85]. Guest [Har71c, She72, Bat99a]. Guidance [Bai75]. Guide [Fis79, Har79, Lar74, Ols74, Wei95b, Bor94, Bro00d, Cal00, Cor94b, Fu96a, Fu96b, Har71d, Has95a, Hei94, Hoc89, Kin98b, LS70, Mat85, McK96a, McK96b, Sch74, Sch98a, Sie86, Sny98, Sop96, Sou96, Wal86c, Wei76, Wit96a, Wit96b, Mim85, Haz98, Lan76, Boy90b, Cre95b, Tom99, Wei93b]. guidebook [Mat97a]. Guided [SFB^{+}00]. Guidelines [And94, But99, Cre99b, Cre99a, CH99, Day99a, Fra99, Gas99, Kaz75b, Lev99, LCB99, Mai90, RM99, Ste99b, Sun99, Dod79, Wri71a]. guides [Jac96]. guild [Old91]. Guinet [Die84a]. Gupta [Buc75]. Guy [Har75d].

H [Bas74, Don77, Ere81a, Gro76c, Har70b, Har78a, Har79, Har77a, Koc79, Koe75, Mar77b, New76, Nov82, Rei77, Swa76c, Swa77c, War79, Wil85b, WH70, You75, Fly97, Fid87]. H. [Ray99]. habits [Chr72]. Hague [Swa76c, Swa76a, Wil85b]. Hal [Har76c]. half [Bro80b, Nis99, Tsa98]. half-century [Nis99]. half-life [Bro80b, Tsa98]. Hall [Art76a, Dav79b, Die82b, Dom78, Fis81a, Fis81b, Har83, Mar80, Mig79, New76, Str78]. Hamburg [Rei75]. Hamden [Gro76c, Lan78, Wel70]. Hammer [Art77]. Hamming [EQH00]. hand [Ber84]. handbook [How87, Jaf98, LH76, Whi89a, Wil00b, Wal00a, Sha96, Aus93, Sav97a]. handicaps [Kir81]. Handling [Ger79, Buc89, GWDS86, NE90]. hands [Gre92, Lar90]. hands-on [Lar90]. Hannah [Afs99, Tau74]. Hanneman [Mig80a]. Hans [Fid87]. Hansen [San99]. Hanson [Kra92c]. happens [Gro76c]. Happy [Kra87a, Kra91b, Kra92a]. harbors [But99]. hard [Bor86b, Bor96, Ger79, Mat85, Pat85, Shi85]. Hardesty [Lap00]. Hardware [Har83, Mea83b]. Harold [Wel80b]. Harris [Afs99, Har77a, Mar77b, Swa77c, Dil82, Afs99]. Harry [Str76]. Harter [RV98, TZ93]. Hartley [Whi00]. Hartvig [Mea82b]. Harvard [Lan77a, Rou76a]. Harvold [Ere80]. Hash [Boo74b]. hashing [Boo72]. Hauptman [Pet99]. haven [War99b]. Haworth [Pat85, Rei85, Shi85].

Hayes [Mar85, New73, Ano74b]. HCI [Sch97a]. headings [Byr75, Hum84, Jen95b]. Health [Bal89, BDZL89, BR94, Bra87, Cim95, DG95, Gor95, Har94, HH95a, Her95, JS95, Ken87a, Lan94, Lin87a, LC87, LB89, LB94, LH95b, MH87, MWH^{+}95, Nov82, Pal87, Pet94, Rod95, Sap94, SMJ95, Spr87, WT94, WCL87, Wes87, Wol87, BDL88, Ber86a, Bix96a, Bix96b, Jac96, Ras90, SJ00, JS95, Lin87a, Wes87]. Heath [Pre71a]. Hebrew [Wei97b]. Heft [Wel71]. Hegarty [Boy85a]. Heidelberg [Lev76a]. Heilprin [Rod93]. held [Har76b]. Helen [Cri81]. Helping [Tor89]. Henderson [Ric77, Afi99]. Herbert [Gro76c]. Here [Kra91b]. Herman [New73, Str78, You75]. Hernon [Ste00a]. heroin [Tod99]. Hershfield [Kaz77]. Heterogeneity [Law82]. heterogeneous [Bal00, MR81a, MR81b, SNG00]. heuristics [Wil00c]. Hewitt [Tau74]. Hierarchical [Bas92, WNAR99, Dum87, Hol86]. hierarchies [Gor82]. High [Neu95, Ano70b, Cob96, DH97b, FDD^{+}99, Fu96a, Fu96b, Key97, Mar83c, SSD99, Sha97b]. high- [SSD99]. high-energy [Ano70b]. high-tech [Cob96, Fu96a, Fu96b]. Higher [Kaz75b, AL98, Coo71]. Highlights [Coh95]. Highly [Gar79, OR78, OG80, Cag70c, Hur80]. highway [FW99, Rob96a, Cas98a]. highwaymen [Cas98b]. Hills [Kun82]. hint [Gol97]. Hirtle [Vea90]. historians [Cas91]. Historic [And91, SS89]. Historical [Bar96, Eli87, Flo99a, Gul87, Hei88, Hen88, Hir89, Jac91, Ken87b, Ken87c, Kil87, Kra87b, Ric88, Rob96b, Sal87b, Swa88, Whi87, Woo87b, Wyl87, Ben00, Bor87, Bus95, KW97, Ste80, War93, Zen99]. histories [MSPPM98]. History [FM97, Hap99, Mea87, BCSZ79, Bru78, Col00b, Day00b, DW98, Fre80, Her84, Mea98, Miz97, Myb98b, NC86, RL97, Shi86a, TEC99, Wal75, War98b, War99a, Wil77b, Woo90, Bag81, Bou80, Rus77, Wei97c, WWB97, Har77a]. hoax [Wha83]. hobgoblin [Soe76]. hoc [And78]. Hodina [Don74]. Hoffman [Gro76c]. Holland [Kaz79, Koc79]. Holography [JB71a, Voo80]. Homage [Mea79d]. home [DG00, Klo89, Mim85]. homes [FW90]. Homewood [Kaz75a]. Honor [Har75d]. Honoring [Kra88b, Kra89a, Kra92b]. Hopfield [CN95]. horizon [HA95, Hjø96, Sal96]. horse [Hal95]. Horton [Woo87a, Bro00d]. Horwood [Mar85]. Hostrop [Kaz75a]. hosts [BBDH00]. hot [San99, Hai97]. Hotel [MP98]. Houghton [Tes79]. House [Dro78, BM80]. households [DHS00]. Houser [Shu80]. Houten [Wil85b]. Howard [Cer83]. Hsu [Boy95a]. Hsu-Yuang [Boy95a]. HTML [Fly97, RNBV98]. HTML/ [RNBV98]. Huffman [LY91]. Human [Bat98, CB83b, FC92, Gau92, Har92a, LH92, Mar92a, MB71, Myl92, NK83, Osb98, Sha97a, Sha98, Bea91, Car88, Che99, Div87, GCS97, Hig99, Hoe72, Hum00, Lan98a, Lan98b, Lee96a, Lee96b, MQP98, Nor96, Pep74, Qin00a, Ano81j, Pet96b, Pet96c, Tor89]. Human-Computer [FC92, Gau92, Har92a, LH92, Mar92a, Myl92, Sha97a, Car88, Div87, Nor96]. human-machine [Lee96a, Lee96b]. humanities [And99, Bus95, Hub77, Mac82, Mur73, NZ91, NN92, SBW93, Tib94]. humans [Mar83a]. Humphrey [DRCG91, Her96a]. Hunter [Tau74]. hunting [Ten93]. hybrid [Dog98, LWZ99, PR98, Sav97b]. hypergeometric [Wil93a]. hyperlinking [Kim00]. hyperlinks [KW97]. Hypermedia [Lit96, Yan97, Beh96a, Beh96b, CK99, IT89, KE99]. Hypersonic [KET99]. Hypertext [CB89, IT89, LR89, Rad89, Sca89, SBPP89, BTSKC96, Che97, DW95, EFW96, FC92, GS93c, GGJ91, Gir91, KL98b, KL98c, PN96, Qiu93, Ray94, SW92, Wes91, Wol96, WD97b, Chi93, Glu95a, Lar90]. hypertext-based [DW95]. Hypotheses [Koe92b, Dor88, Koe92c]. Hypothesis [Per85, Ott94]. HyTime [KW97].

IAIMS [BD88, BFM88, FW88, LD88, LB88, PW88, War88, Wes88, WMME88, ZvBR88, BFM88, LD88, PW88, War88, Wes88, WMME88, WA97]. Ian [Cra96]. IBM [Don70, Han96, Mim85, Lin70]. IC [Mai90]. Ichthyology [Mar92b]. ICL [CPW88]. ICR [SK99]. ICSU [BCSZ79]. ID3 [CSSI98]. IDC [FNNW74]. Idea [Bat79a, Ray99]. ideas [And99, Pra90, Sch96b]. Identification [KM95, BCM^{+}70, HST96, Haa97, JBS^{+}77, TSV99]. identifier [Tur77, BCM^{+}70]. identifying [McC95]. identity [CL89, McD99, OC88, Var99, Wal98, Coz82]. Ier [Gro76a]. if [Sel71, Wat98, Hei90, Zwe95]. IFIA [Swa76c]. IFLA [Har76b, Har77b]. ignorance [Boy91, Shu78]. ignore [Ler79]. II [Wes00, ABLS88, BD88, BBSM87, Boo90b, BMvR91b, BF91a, Coo73b, FS97b, Fay72b, Fid91b, GGJ91, GB95, Inh75b, Jon74, KB98a, KS97, LR70, LMO^{+}96, Lin87a, MR81b, MHA82b, Mow92b, Par82b, Ric88, Ro88b, RSZ71, SK88a, Sha91b, Sme92, Sol97b, YFR81b]. II. [Har75b, LAD97, SK97b, SvB94, WW99]. IIB [FS97b]. IIC [BDSA97, Goo97, GLS97, LH97, RH97]. IIe [Mim85]. III [Koe92b, Fay73, FW88, FNNW74, Gir91, KS97, Lyt88, Mow92c]. III. [Boo97, Fid91c, Koe92c, Par82c, SK88b, Sol97c, Wes87, YFR81c]. ill [Aba86]. ill-defined [Aba86]. Illinois [Har76a, Kaz75a, Lil78, Ric77, Swa76b, Swa77b]. illus [Har77c]. illusion [Swa88]. illustrated [Vag80, Wil98c]. illustrating [Wel83]. illustration [Bri91]. illustrations [Hei78]. Image [CDV98, Ens95, RTM99, Bea91, HOS^{+}93, Jen91, Liu97a, Tur98, Wue97, ZC00, Col98]. image-based [HOS^{+}93]. Images [Ror99b, Sko98, Lay94, OOA99, RCZS99, Sun99, Wei94a]. Imaging [And91, Bea91, Bes91, Buc91c, LL91, Lyn91a, Nug91, Rom91, GQSdP98]. immediacy [Asa81a]. immunity [CH99]. immunodeficiency [BB95]. Impact [Abe90, BB96b, Bra87, BT99, CS90, Gar90, HP00, Ken87a, Kun82, Lan80, Lin87a, LC83, LC87, LC90, MH87, MV95, Pal87, RV96, Spr87, WCL87, Wes87, Wol87, AP96b, Asa81a, Bro97, Cam96b, Cit98, CB83b, ER00a, Gar98, HN97, HN98b, Har98, HF00, Hir78, KB98b, KN00, Kos94, Law74, Mar00, Men92, Mil90, MBS98, Myb98b, Pap93, Per83, Pri83, RR77, RV98, Sam95, Seg94, Smi96a, Smi96b, SS00, Van97b, VTS97, Vei80, War81, War91, Art77, DG78b]. Impacts [LS97, Goo97, GLS97, Van98b, Wei97c]. impediment [Bro98c]. Implementation [BDBR97, BD97, BDSA97, D’E97, Goo97, GLS97, LS97, LH97, OG93, RH97, SA97, ABB^{+}91b, DS80, HKE97, SWK78, YS93, Coo73b, FW88, Liu97b, Und98a, Dro98, Wal86b]. implementing [BD88, Wil84, Fuh00]. Implications [Goy87, Kun82, Mor81, RH97, Voo80, Art76b, Liu96, Lyn77, Par00, Sch94, Sho71, VA73, Whi98c, Zin00, Bri84, Cul85, Dic86, Gil78, Mar94, McG76, SRS98, Myb99]. implicit [Rob75, Swa89a]. imply [Bac72]. importance [LJ70a, Mar78, PS79, SYY75, SWY81, SK99]. imposed [LCL85]. impoverished [Cha96]. Imprecision [VD98]. improbable [Pai90]. improve [GCH99, Ro88a, Ro88b, SMB72]. improved [Boy86f]. Improvement [MRY73, Wel83]. Improving [BD92, MV95, OE81, Pai71, RT81, SB90, HC96, Kuc75]. In-Depth [Wei95b]. in-house [BM80]. in-library [Jai72]. inaccessibility [Spr99]. incidence [EM94]. including [New73]. inclusion [MDS70, Zmu75]. Incomplete [Haa93, CDV98]. inconservation [Yan90]. inconsistent [LK91]. incorporating [BF94]. increase [Coo83c]. increased [Cha76]. Increasing [CK99, WH91, LT74]. Incremental [Hum00, Qin00a]. indebtedness [DC93]. independence [WY90a]. Independent [BD96, Bes96, BB96b, Coh95, Los97a, Wil92a, WA96]. Index [Bol86b, CN81, Har71b, Hin70, Lan73, Lev73b, McM76, Nov82, Rei77, AKE94, Asa81a, Ben80, Bol86a, BP70, Bro70a, Car74, Car79, Cra82, Egg97, Ell70, FB81, Gar76, Gue78, Hei78, Lar91, LK97, LH76, MDS70, NT85b, New73, RC95, Ros71b, RR71, SJ72, Sve72, Tho72, Vag80, WM78, Yer73, vdVM99, JGR90, Byr76, Poy82, Sma80, Vir70]. index-handbook [LH76]. indexed [SK98]. Indexer [Sch91b]. Indexers [Boy91]. indexes [Aul71, Bla71, Bot70, Boy86f, DG91, Dio94, Gar83, Gar91, Gou74, Her78, HH70, Kil93a, Law76, PZSD96, Smi74b, Wei97b, Yer73, Bar74]. Indexing [And94, Bat98, Bel71, Bre91, Coo78, DDF^{+}90, Fid94, HW80, Lay94, LF94, Mig80b, Mil94, SA91, Soe94, Ste74a, Sve94, Tho79, Tib94, Wei95b, Wel94, Ano70a, Art76b, Blu71, BS74, BS75, Bur95, Bur80, CMK^{+}98a, CCW84, Coh95, Coo79, Cra77, Cra80a, Cra86, DKD71, DG83, DG91, DPP78, DS84, Exn97, Fag89, Flo72, Fug85, GS93a, GD90, GLL90, GD98, Gul87, Har75a, Har75b, Har74, HJD00, HHC70, HKE97, HM87, Hum99, Hum00, HSP75, Jon73a, JGR90, KSDR90, Kwo85, LN82, Lan98a, Lan98b, LR70, LK97, Lev74, Luc93, Mar77a, Mar79, OE81, PN98, Pri83, Qin00a, Ros71a, RW82, RSZ71, Sal72c, SWY81, Sal84, Sch82, Sha92, Sha91a, Sha91b, WEWP97, Wei91, Wei93b, Wei94a, Wei96a, Wel72, Wil79, WP94, WNAR99, YLYY00, CB73, Boy91, Boy92, Lea75]. Indexing [Mai99, Wei95a, Wel80b, And88, Wei93a]. India [Lan90]. indication [Hom83]. indicative [RSZ71]. indicator [CN81]. indicators [BB83b, Cro99a, HY98, Hen99, Koe83, LAD97, MKB78, NZ91, Pie92, VHdB^{+}98, Whi81, Whi83]. indices [BYN00, GH96]. indigenous [Wil94]. Indirect [SK99, SK98]. Indirect-collective [SK99, SK98]. Individual [All00, CCM00, Che00a, Lil78, Sie70, ASAOE99, Dav77a, Par00, Whi76, Wil98a, BR97]. individuals [Ber94, Rei85, Sha98]. inducing [Col00b]. induction [Wu98]. inductive [CSSI98, Sch00a]. Industrial [Win99, Afs99, Bon79, Hoc74, JJ77, JJ80, LWG74, Pra90, Sta75b, Tal79]. industries [AC93]. Industry [ASS73, Mat85, Tho82, Han96, HDRD88, Hol71, How89a, Sch94, Var75b, Wil85b, Var75b]. inequity [CA80, vB81]. Inference [LCL85, O’C70, RSZ71]. Influence [KRC81, Kuc75, ALD96, Cam96b, Cit98, CMRWW93, ER00c, EB88, Koc84c, LB83, LAD97, MAN80, McI97, SH89, TvR90, PR98]. influences [BR97, MM88]. influencing [Bur00]. Informactics [Gor95]. Informal [Cra71b, Ehi90, Kor74, Wee92, Cra72, HH97, Dre99]. Informatics [ASR85, BDL88, Bal89, BDZL89, Cim95, DG95, HH95a, Her95, JS95, LB89, LH95b, MWH^{+}95, Rod95, SMJ95, Vei80, Wel81, BT00, Kar98, KRH98, Mar99, Sor96, ZvBR88]. informatiia [Rya80]. Information [AL96, AKS96, Abe96a, Abe90, ASAOE99, ABB^{+}91a, ABB^{+}91b, Adl92, AKE94, ABLS88, Ano72a, Ano72b, Ano77a, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano85g, Art73, Art77, Art79, Bac93, Bai75, BR97, BDL88, BD88, Bal89, BDZL89, BR94, BHCC^{+}74, Bat79b, Bat99a, BGM85, BL85, Bea92, BS91b, BDBR97, BR76, Bel77, Bel81, Bel88, Bel73, Bla93, Boo83, BP93, Bra87, BCB96, BF91a, BD97, BDSA97, Bre79, BFM88, Bro95a, Bro99a, BC99, Bro99c, Buc89, Buc91c, Buc91d, Bur70, Bur81a, CZ88, Cim95, CS91, Cox88, Cox92, CS90, DRCG91, D’E97, Dan83, Dat79, Dav88, Deb72, DG95, Der84, Der94, Die82a, DH97b, Dil82, Doc91, Dos85, Dos88, DB94, Dro77, Dum87, EAW99, Ere80, FMN82, FW88, FK99, Fre77, FF80, Fri97, GP83, Gar92, Gar71, Gar90]. Information [Gar88, Ger79, Gol87, Gol74b, Goo97, GLS97, Gor95, Gre92, Gro75b, Gul87, HK79, Har71b, Har94, Har78a, Har79, HH92, Hay88, Hay96b, HAA87, Hei88, Hei95b, Hen99, HH95a, Her95, Hol86, Hum70, Hum96, Hur88, JS95, Kat00, Kaz77, Kaz79, Kei70, Ken87a, Kno77, Koc84a, Kra87c, Kra99c, Lan80, Lan73, Lan76, Lan77b, Lan94, Lar74, LAT79, Leh92, Lei73, LS97, Lev82, Lie94, Lin87a, LD88, Lon90, Los90, Luk96, LH97, LC87, LS88, LC88, LB88, LB89, LC90, LD91, LS92b, LB94, LH95b, Lyt88, Mac79a, Mar89, Mar92b, Mar98, MH87, MWH^{+}95, McM73, Mea78b, Mig79, Mol92, MF88, Mom82, Nah98, NKB72, Nit79, OC88, Ozo79a, Pal87, Pap84, PW88, Pet94, Pre71b, Pre88, RCZS99, RRB91, Ric88, Rie88, Rod95, RH97, RL99, RTM99]. Information [RH99, Ros82a, Rus70b, RTH71, Rus85, SC85a, Sap94, Sar99, Sch88a, Sch73b, Sch88b, SM88, SA97, SMJ95, Sim81, Sla85, Sma99, Sme87, Smi98c, Smi99, Spi97a, Spr87, SB91, Stu78, SWT^{+}79, SOM^{+}99, TC73, Ten88, TL74, Tes79, Tho82, Tom92, Tra98, Tsu79, VA73, Var75a, Var78, Wal72, War88, War98a, War99b, Was80, Wat99, WT94, WC92, WCL87, Wes88, Wes87, Wil82, WMME88, Wol87, Woo87a, Woo88b, WH70, Yer91, Yov70, YFR81a, YFR81b, YFR81c, ZvBR88, Zoe88, dR90, vRL96, Afs99, Aga99, AL94, Ain84, Alg95, AM77, AP96a, Ama84, Ami70, Ami73, And90, Ano81j, Ano00c, Art74, Art75b, Art76a, AL84, Aus93, Bac72, Bak89, Bal88, Bal00, BI00, Bar95, Bar96, Bas74, Bat99b, Bat99c, Bax85, BK96, Bea91, Bec71, Bec84, Bel75, BJ79]. information [BIFW71, BA74, BWKG75, Ben80, Ben86, Ben00, Ber82, BC00, BNS^{+}00, Bol92, Boo74a, Bor86b, BR92, BHH96, Bor84, Bor90, Bor93, Bor85b, Bos87, Boy86e, Boy87b, Boy87a, Boy90a, BM80, Bri72, Bri84, Bro70b, Bro80a, Bro88, Bro96, Bro98c, Bro99b, Bru98, Buc98, Buc99, BC97, Bue85, BT99, Bur81b, Bur85, Bus95, Car74, Car95b, Cas92, Cas94, Cas97, Cas98a, Cas98b, Cha99a, Cha81, Cha85, Cha86, Cha91, Che95a, CNMS97, Che98a, CWBL00, CMB^{+}95, Cob96, CA80, Coh99, Col81, Col94, Col99, Col00a, CM00, Col95, CC96, Coo70a, Coo72, Coo82, Coo91c, Coo98, Cox83, Cox93, CNO^{+}00, Cra71a, Cra81a, Cri81, Cro81, Cul85, Cul88, Cur93, Dan84, Dan00, DDC78, Dav96b, Day00a, Day00b, DK90, Die79, Die81, DS96, Dil98, Dil00, DW95, DCFQ00, Dio83, Dom00, Don72, Don79]. information [Dos86, Dou72, Dow77, Dre99, DK75, Dun00, Dun96b, DS73, ER98, EW70, EQH00, Ell86, Ell96, Eps80, Ere81b, Ere81c, Ere85, EN98, EK75, FDD^{+}99, FW99, For91, For99, Fox86, FFKS99, Fra86, FB78, FBC79, FB88, FS91b, FSV93, FRGL75, FNNW74, Fug85, Fuh00, Gar83, Gat92, Gau92, Gil78, Glu94, Glu95b, GQSdP98, Gor79, Gor90, GL91, Gra82, GM86, GMPS00, Gri80, GLW86, Gro88, GSD74, Hal80, Hal81, Hal85, Hap99, Har71a, Har92b, HNW93, Has85, HG98, Haw77, Haw86, Haw90, Hay95, HW98, Haz98, He98, He00, Hei78, Hei86a, Hei87, Hel71, Hen88, HH97, HMNW86, Her78, HS78, Her84, HPH96, Hig99, Hir99, HA95, Hjø96, HKE97, Hoc89, Hoe72, Hol71, Hol99, Hom83, Hoo91, Hun96a, Hun96b, Ing97, Inh76, Jac72, Jaf98, JR96, JN93, JB71a]. information [Joh86, Jon91, JB85, Jul99, KMG^{+}93, Kan76a, Kan94, KL98a, Kar98, Keg70, Ker77, Ker84, KZC86, Key96, Key97, KL98b, KL98c, Kil93b, Kil93a, Kin82, Kin98b, Kir73, KW81, KdK90, KRH98, KM95, Klo89, Koc72, Koc84c, Koc87, KBN90, Koe75, Koe92a, Kor74, KRD95, Kra73, KB78, Kra91a, Kri87, Kuc82, Kue72, Kuh99, Kwo99, LJ70b, Lan71, LM78, Lan87b, LMSC82, LMM^{+}73, LG96, Lan82b, Law74, Laz00, LBW00, LK91, Lev76a, Lev77, Lie96b, Lin97, Lin78, LS71, Liu96, Liu97a, Liu97c, Liu98, LR79, Los96a, Los96b, Los97b, Los97a, Los00, LFHG99, Lun76, LH76, LE85, Lyn77, Lyn91b, Lyn94, Mai90, Mal99, Man82b, Mar92a, MKB78, MBVL99, MCT73, Mar94, Mar99, Mat85, McC00, McC78, MMB99, McG76, McK00, McL79, ME98, Mea72, Mea83c, Mea86, Mea95]. information [MWY95, Mea00, MB91, Men92, Men85, Mes95, MHCW89, Mic94, Mil87, MWR72, Miy98, Miz98, Mog75, Mon94, Mon72, Moo72, Mor94, Mos70, Mow92a, Mow92b, Mow92c, Mud98, Myb98a, Myb99, Nad85, NE90, NT85a, NM74, Nel97a, Nis99, O’C93, Old91, OO97, OD70, Ozo79b, Ozo80, Pac82, PS79, Pai71, Pal99a, Pap98, Par82a, Par82b, Par82c, Pat90, Pen70, Pet96b, Pet96c, Pet99, Pie99, PLM00, Pop75, PB99, Pul74, Qiu93, Rad82a, RW86, Rag96, Ray97, Reg88, Rei85, Ren80, Rep89, RB83, Ric87, Ric89, RL97, Rob96a, RFK97, Rob75, Rob00, RF00, Ros74, Rou92b, Roy89, Rub92, Sag75, Sal72a, Sal73, Sal80, SBF83, SFV85, Sal85, Sal87b, Sal96, Sam95, Sar71, Sar75, SBA85, SKCT88, SK88a, SK88b, SK97a, SK97b, SKB99]. information [SMB72, Sch73a, Sch87, Sch99, Sha95, SD83, Sha90, She71, Shp92, SMHC93, Shu78, Sie86, Sie96, Sla75, Sle94, SJ00, Sme92, Smi74a, Smi00a, Sny98, Sol93, Sol97a, Sol97b, Sol97c, Spa99, SS97, Spi97b, Spr99, SNG00, SGS70, Sta75a, Ste99a, Ste00a, SCLL73, Str82, Str84, SS95, Stu72, SMC83, SS89, SB83, SEW00, SM75, Swa77a, Swa78b, Swa78a, Swa78c, SWB79, Tag75, TS96, Tal79, Tau74, Tem94, Ten90, Ten93, Ten94, Tho81, Tho71, Tho90, TEC99, Tod99, Tom96, Usd85, Var75b, VTS97, VS99, Voo99, Voo74, Voo80, Wal81, Wal75, Wal85, WEWP97, WVE85, Wan92, War73, War91, War98b, War99a, Was74, Wat89, Wei88, Wei94b, Wei96a, Wei76, Wei73, Whi89a, Whi77, Whi89b, WM98, WM00b, WY73, Wig95, Wil71, WK75, Wil86, Wil85b]. information [Wil95b, WW73, Win99, Win70b, Wis00, Wit80, Wol86, Wol71, WY92, Woo84, Woo77a, Woo70, Woo87b, Woo90, Wri73, Wue97, Xie00, Yag87, YR98, Yan90, Yan97, Yer71b, YG90, Yoo94, YF85, YdKKM87, YK93, Yua97, YM99, Zai77, ZK72, Zal74, Zhi97, dSM93, de 71, vB81, ABLS88, Art82, Bel70, BBFM99, Bot73, Buc91a, Buc99, Bur92a, Can99, Col93, DH97a, Die82b, Dun96a, FM97, Flo00, Fri73, Ger79, Gro96, Kra74a, Kra87a, Kuh91, Lun80, Lyt88, MV95, Rya80, Sie70, SA91, Sta82, Swa73, Swa88, TC73, Tay82, Tho73b, WWB97, Wri71b, WNAR99, Lan78, Don74, Lar74, Aik97, Ano74b, Bla79, Bro93a, Bro94, Bro95a, Bro98a, Bro00b, Buc91b, Cal89a, Che98a, Col98, Cox94, Dav98, Deb86, Dol98, Dro98, Ere81a, Fra86, Har87, Her82, Hil96a, Hil96b, Key97, Koc79]. Information [Kri89, Mad96a, Mad96b, Mal99, New73, Smi98c, Sta77b, Ten88, Van90, Var75a, Whi00, Wil00c, Wil98c, Lar90, Mil93, Tay94, Var97]. information-conscious [Gar71]. Information-processing [Sim81]. information-related [MF88]. information-retrieval [Bor86b]. Information-seeking [BR97, Mar89, SC85a, Cha81, Cha91, Dre99, HH97, OO97, Xie00, Yan97]. information-theoretic [WY92]. Information-value [McM73]. Informational [Die84b]. Informationless [TC73]. informative [CM00, Lev77, Toc70]. Informatology [OD70]. informed [CNO^{+}00]. Informetric [Boo90a, Boo90b, Boo97, ER95]. Informetrics [BI98, BI00]. Informing [Spa99, Fra86]. infostructure [Zur84]. INFOTRENDS [Cul88]. infrastructure [BNS^{+}00, DG95, GSK98, Lyn98, PW88, WA96, DB94, War98a, Wue97]. infringements [CH99]. initial [Tho73a]. initiative [BI97]. Initiatives [Lap99, LD88, MMB99, AL98]. Ink [Whi99b]. Inner [Aga99, Fal98]. Inner-city [Aga99]. Innovation [NK89, SM88, Und98a, Sch96b]. innovations [Lon90]. Innovative [Kuc75, Hei86a]. Input [Coo77b]. Input-output [Coo77b]. inquiring [MWR72]. inquiry [Ber84, DHS00, Har84, Par94, Sar71]. INS [Man82a]. Inside-out [Bur96]. insider [Flo00, Cor94b]. insourcing [Key96]. installations [SBPP89]. installing [How89b]. Instances [Bur96, Dav96a, Lin96, Sut96]. instantaneous [Lev74]. instead [Har92a]. Institute [Ano72a, Har76a, Hei78, Ray97, Ric77, Sch73b, DRCG91, Mac79a, MH87, MV95]. Institutes [Bel81, Swa78b, DH97a]. Institution [BR94, Har94, Lan94, LB94, Pet94, Sap94, WT94]. Institutional [Ano75e, Ano75c, Ano75d, Ano70b, BR97, Con93, LD88]. institutions [ALD96, BFM88, Hai97, LAD97, Pou89, Rei85, Lap00]. instruction [Hol96, Sla73, SHS72, Jer70, Smi96a, Smi96b]. Instructional [Kea83]. Instructions [Ano80h, Ano82d, Ano85a]. instrumenting [Coo98]. integer [AF98, Smi81]. integral [CYW99, Hoe72]. integral-bit [CYW99]. Integrated [ABB^{+}91a, ABB^{+}91b, BD88, BS91b, BDBR97, BF91a, BD97, BDSA97, BFM88, CS91, DRCG91, D’E97, FW88, Goo97, GLS97, LS97, LD88, LH97, LB88, LD91, Nie76, PW88, RRB91, RH97, SA97, SB91, ZvBR88, Par00, Van90, ABB^{+}91a, ABB^{+}91b, BD88, BDBR97, BF91a, BD97, LS97, RRB91, SA97, SB91, War88, Wes88, WMME88, Wal86b]. Integrating [GFHR97, Zur84, BR95a]. Integration [BFM88, Smi98c, Pac82, BD88, Lan94]. integrity [LaF00, Lyn94]. Intellectual [COC90, BSM90, Bra89, DC93, GD90, McC86, McC90, Per94, Wee99, Wei96b, WG81, Wil92b]. intelligence [BF91b, CM84, Cob96, Jon91, LS84b, Mac96, Nan95a, Smi84b, Wal81, vdBL96, JGR^{+}95, Kas96a, Kas96b]. Intelligent [Gau92, HG98, SLZC92, Has98, Mad96a, Mad96b, MSHB86, Mar85]. intent [Wes88]. intentions [Xie00]. Inter [HDRD88, Wol96, Hol96]. Inter-industry [HDRD88]. Inter-record [Wol96]. inter-university [Hol96]. Interaction [Art79, Bry70, Col00a, SKB99, SS97, Bea91, Col99, Inh75b, Jon76, Kar98, KH71, Sha97a, Bro93a, Car88, Div87]. interactions [Mar92b, Myb99]. Interactive [BM72, BC99, Eft00, JGR^{+}95, MHA82a, Mor76, AB76, BRR96, Bla80, HH95b, HH95c, Kat72, LUH^{+}96, MCT73, MB71, MHA82b, Mea83c, PM00, Rob00, Sch81, Sch74, Spi97b, SB83, Tho71, VA73, Wil71, Xie00, Dun00, RCJS94]. interchange [Ama84]. interchangeability [Zoe88]. intercorrelated [McG84, Smi84a]. intercounty [KS77]. Interdependence [MB86]. Interdisciplinarity [AI76, Her99]. interdisciplinary [ABLS88, CPR84, Nee81, Pal99b, Pie99, QLA97, Smi74b, SS00, Stu72]. interdocument [GLW86]. interest [Eas89, Gra82, Sub84, Yer71b]. interests [McK96a, McK96b]. Interface [FC92, Gau92, Ham00, Har92a, LH92, Mar92a, Myl92, Tho71, Bal00, BBMS87, BCM89, Car88, CD97, DCFQ00, Kie98, MR81a, MR81b, MCT73, MW91, Mel72, Wil94, Ros73]. Interfaces [BC99, KMG^{+}93, Mar92a, RCZS99, RL99, RTM99, RH99, Sma99, CK99, Cur97, FHN^{+}93, MWY95, Wil86, Yee91, Has98]. Interlibrary [DG78b, Gas99, Rei84, RR77, Spe71, War81]. Interlinking [HBG96]. INTERMARC [Gro76a, Gro76a]. intermediaries [Mar83a, Wil86]. Intermediary [MHA82a, MHA82b]. intermittent [SW91]. internal [GBDH76, YdKKM87]. International [Ano70b, BI98, Don70, Gar86, Har77b, Hei78, Kri87, Lan80, Lea76, NN92, NT85a, Nov82, Sal86, Smi98c, Swa76c, Swa76a, Swa78b, Ano74a, Bla71, Bor85b, Bra89, CN81, Cla75, Dav84, Dav86, Dos88, Hal80, Han75, Has95b, LM78, Les89, Liu97c, LE85, Mea83d, O’H95, Old91, Ols99, RL97, SB94, Var75b, Whi98a, ZK72, Zal74, BCSZ79, Ray97]. Internationality [Bri84]. Internationalizing [Whi89a]. Internet [Bor94, Jac96, Jaf98, Smi98a, Sou96, Sut99, Tom96, Wal00a, Bat98, BC00, Bru98, CHSS98, CB94, Dre99, Fag95, GCS97, Gas98, Her99, KB98b, Kin98a, LBIP97, Lyn98, Mal00, Nah98, Osb98, PPP98, RW00, SCD98, Sch98b, TL98, Voo99, Wat98, Zha00, Has95a, Haz98, Hun96a, Hun96b, Myb98b, Pet98a, Smi97, Smi98b, Tom98a, Wil00b]. interpersonal [Wol72, Bai75]. interpretation [Bon89, DG87, Gre92, Hub76, Gel96]. interpreter [SSM93]. Interrelationships [NCB72, Moo73]. Interscience [Kra79, Lan77b, Sie70, Wri71a]. interspace [Luk96]. Interstate [Kaz75b]. intervention [SS93, Gat92]. interview [AL84]. interviews [Sch00a, Abe96a]. introducción [Har70a]. introduce [Cov00]. Introduction [AL96, Ano70d, Ano80a, BT00, BD96, BP98, BH97, CT99, Che98b, Che00b, Che00c, CL82b, Cra80b, Cra81b, Cre99b, D’E97, Dil97b, Dog98, Erc99, EL85, FB99, Gri81, Hur00a, LP97, Lun82, Lun83, LS84a, LH84, LH86, LL86, LC87, LS88, LC88, LB88, LR89, LB89, LC90, LW90, LD91, LL91, LH92, LS92b, LF93, LS93, LP94, LF94, LB94, LH95b, McN98, Men95a, RDS98, RL99, TL98, vR98, Bat99a, Che76, Dav76a, Gau92, Gre96, KRH98, Kri86a, Lar90, McC80, Pat81, Rus87, Wel70, BB96a, CCM00, Fro94, Glu94, Pre71c, Sie96, Bas74, Ber86a, Mar80, Nel97a, O’C95, Str76, Wei91]. Introductory [Tes79, Ten94, Wei93b]. Introductory-Manual [Tes79]. inventing [Smi97, Mal00]. invention [Bre91]. inverted [MB95]. investigate [CH00]. investigation [Bel85, Bro83, Bro85, Bro95b, Chr95, EQH00, Fra86, GH96, Goy84, HT91, Hum99, Kim82, LS87, Qin00b, Bar74]. investments [Hen99]. invigorate [Pap88]. Invisible [Gre92, Bat99b, Nom84, Zal74, de 71]. Invited [Sha83a, Wha83]. Invoked [CSR^{+}98]. Involving [HW71]. IPCT [BC96, CB94]. IR [Rob75, BSR93, Kar98, Pai93a]. IR-systems [Pai93a]. Ira [Har75d]. Irish [Fis81a]. ISBN [Boy85b, Dav77b, Dav79b, Fis79, Fis81a, Fis81b, Gro75b, Gro76a, Gro76b, Gro77, Har76c, Har76b, Har76a, Har77a, Har77b, Har77c, Lan77a, Lan78, Lan82a, Lar74, Lev76a, Mar77b, Mig79, Mig80a, Nov82, Ols76, Ric77, Shu78, Swa77b, Vag80, Wei79, Wel80b, Wil85b, Mar85]. ISBN-0-8371-7955-6 [Ols76]. ISI [Dro77, DS91, HN97]. isolated [Swa89a]. isolation [IC97, MT96, Pee96a]. ISORID [Hal80]. Isotopes [WH70, WH70]. Israel [Ker77]. ISSN [Lil78, Mig80a]. Issue [Ano74a, Boy98j, Boy00h, Boy00i, Boy00j, Boy00k, Boy00l, Lie94, Lil78, Ano00a, Ano00c, Ano00b, BB96a, Boy95b, Boy95c, Boy95d, Boy95e, Boy95f, Boy95g, Boy95h, Boy96a, Boy96b, Boy96c, Boy96d, Boy96e, Boy96f, Boy96g, Boy96h, Boy97a, Boy97b, Boy97c, Boy97d, Boy97e, Boy97f, Boy97g, Boy97h, Boy97i, Boy97j, Boy98a, Boy98b, Boy98c, Boy98d, Boy98e, Boy98f, Boy98g, Boy98h, Boy99k, Boy99a, Boy99b, Boy99c, Boy99d, Boy99e, Boy99f, Boy99l, Boy99g, Boy99h, Boy99i, Boy99j, Boy00a, Boy00b, Boy00c, Boy00d, Boy00e, Boy00f, Boy00g, Bro99a, Cam97, CT99, Che00b, Che00c, Coo76, Fro94, Glu94, Wel75, Boy98i]. Issues [ABB^{+}91b, Bes96, Car95a, Dra89, Gas98, Haw90, Kin98a, Lyn98, Mea72, Nah98, Osb98, PPP98, SBA85, Sch98b, TL98, Wat98, Whi98c, Ano78j, Bec71, Bor93, CT99, Gro96, HF00, IC97, KWR98, Kra94b, Kra97, LL85, Lay94, Lil78, LE85, Lyn88, Lyn94, Pit75, PT87, RRB91, Sha97b, Und96, Var78, Wan81, Wei96b, Wel00, IT89, JR96, Und98b, Pas90, Ric89]. Issues/Human [Osb98]. Issues/Information [Nah98]. Issues/Information [Wat98]. Issues/Technological [Lyn98]. Issues/Publishing [Gas98, Kin98a, PPP98]. Issues/Technological [Sch98b, TL98]. item [All89, Egg96, Spe80]. items [Egg96, KMB99, KM00a, Par82a]. iteration [KS98]. ither [Fly97]. Ithiel [Koc84b]. itsy [CCRY98]. IV [Bel88, Bra87]. Ivan [Dom78, Har75d].

J [Bel81, Bla79, Boy85b, Bro00d, Dav77b, DC76, Die84a, Don77, Gro77, Har83, Har70b, Hei78, Lun80, Mar75, Mar77b, McM76, Mea83b, Mig79, Mig80a, Pre70, Ste77, Swa77c, Whi00, Wil85b, Woo77b]. Jack [Deb96]. Jacobs [Che99]. Jacqueline [Lea76]. Jahodn [Sie70]. James [Dro78, Koe75, Tau74, Woo77b, Ano72e]. Jane [McM76]. Janes [Wal00a, Lam00a]. Janet [Mal00]. January [ATJ95, Deb96, Wei97a]. Janzing [Swa76a]. Japanese [Cor82, HY98]. JASIS [Aiy88, Kra87c, Kra99c, Smi99, Ano00a, Ano00c, Ano00b, Bat99c, Bro99a, Bro00f, Eli87, Glu94, HH92, Kol74, Kra88b, Kra88a, Kra89a, Kra89b, Kra90, Kra92b, Kra94a, Kra94b, Kra97, Kra99b, Kuc80, Lip99, Mea79d, Mea79c, Mea79a, MZ79, Mea80, MZ80, MKDL83, Mea83a, Mea84a, Mea94, Nis99, Per94, Row84, SM78, WM78]. JASIST [Ano08]. Java [Cal00, Lea98]. JavaScript [RNBV98]. Jean [Mar85, Wel70, Nel97b]. Jeffrey [Tsu79]. Jeffry [Har79]. Jerusalem [BI98]. Jesse [Don77, And90]. Job [Fis79, Aul87]. jobless [Fro96]. jobs [Sha78]. jobtype [McC78]. John [Don74, Fis79, Mea83b, New73, Sta77a, Sta77b, Ste99a, Ste00a, Tes79, Wri71a, SS70]. Johnson [Har77a]. Jon [Ere80]. Jonathan [Che99, Mim85]. Jordan [Rei77]. Joseph [Ano75a, Art74, Boy85a, Lam00a, Wal00a, Hay96a]. Journal [Bur93, CH73, FK73b, Han75, LS80b, Mau71, OV92, RV96, RSK74, TH82, WC77, WFK73, And78, Bro99b, Cam96b, Cas85, CB94, Die84b, DF85, Dor85, Egg88, Glu90, Gor82, HN97, HF00, Har78b, Hir78, Hum99, Kuc78, MM97, MB81a, Mar92b, MB81b, McC91, Nis99, PPP98, Pit75, Pit82, PTH83, Qua96, Rob91, RMMS96, SKvB92, SvB94, Seg94, SG80, Sma80, Sum84, TB93, Tsa98, Wal86d, WA79, WL81, Caw78, MV95, RV96, Ano74a, Cam97, Ano84k, Ano84l, Ano85g, Bat99a, BC96, CB94, Kra87c, Kra99c, SH89, Smi99, Wal97, Wel80a]. Journal-Oriented [FK73b]. Journals [Har71b, Mcd82, AP96b, AP97, Bai74, BK96, Ber72a, Bra77, BR80b, CN73, Caw78, Che72, CO95, Dor88, EP91, FG70, Glo88, Haa93, HY98, HNW93, Har98, HP00, Inh75a, Inh75b, Mar83c, MAN80, McC95, Moo73, MM78, NCB72, NZ91, Pap93, San74, Sha83a, SPWH99, TvR90, Wel00, WFK72, Pie92]. Joy [Dro78, Boy85b]. Jr [Bas74, Cer83, Str76, WH70, Lou92b]. judges [Ros71a]. judgment [PJ90]. judgments [Dav77a, EB88, Jan91, Jan94, KS90, LK91, NZ91, OWF99, Ror99a]. Judi [Fis79]. July [ATJ95, Gul85, Hay96a, Tau74, WH70]. Jumpstarting [VS99]. June [BI98, Deb96, Joh86]. Jung [Jor89, Lan89b]. just [KM00b]. Justification [Bel71].

Kaimann [Gol74b]. Kantor [de 78]. Karl [Ere81b]. Kashyap [Smi00a]. Kathleen [San99]. Kathryn [Ric77]. Katz [Pat85, Wei79]. Katzan [Str76]. Katzer [Har79]. keep [Ber71]. Keigh [Har78a]. Kench [Var75a]. Kenneth [Har79]. Kenney [Mat85]. Kent [Dan84]. Kevin [Boy85a, Lou91]. Key [Bec71, Cam97, Har77b, Boo74b, Tes86, Low71]. keys [Fid86, Fid91a, Fid91b, Fid91c, Ras80, SLC98, SR77, Tin70]. Keyword [Hin70, WH70, BHWG95, Car75, DKD71, Har75a, Har75b, JB71b, OE81, Qin00b, Wei73, Yer73]. keyword-in-context [Yer73]. keyword-selection [Car75]. keywords [AP96a, GS93a]. Kidder [Mea82a]. Kilgour [Kje74]. Kim [Art75a]. kind [SK99]. King [Lan73]. Kingdom [Dro78, Woo84, Bax85, Wil82]. Kintsch [CM00]. Kiviat [Ell76]. Kiviat-graphs [Ell76]. Klapp [Die82a]. Klaus [Lev76a]. knot [DM99]. knowing [Ber85]. Knowledge [Bac93, BBMS87, BBSM87, Dro78, For91, HM87, Jen91, Kir93, LS93, Mat85, McG76, MA76, Pai93b, Afs98, Bat93, Ber84, BW84, Boy91, CN95, CDV98, Dog98, Erc92, Exn98, HO89, Hjø99, HY93, LUH^{+}96, LS92a, MCBC89, Nei80, RL99, Shu78, SJ00, Spa97, Swa89d, Wie86, WdBFF95, Wil94, WLC00, YCO00, Jen95a, Bro98b, Wil95a, Beh96a, Beh96b]. Knowledge-assisted [BBMS87, BBSM87]. Knowledge-based [For91, HM87, LUH^{+}96, Wil95a]. known [All89, KMB99, KM00a, Spe80]. known-item [All89, Spe80]. Kochen [GBL89]. Koenig [Buc92a]. Kolmogorov [Lon92, Bur94, Lou92a]. Köln [Wel71]. Korfhage [Bor99, Kra99a, Ras99]. Kroeber [Koe81]. Kudos [Mil95]. Kurdish [MS00]. KWIC [BP70, Bro70a]. kwoc [Bur80, Aul71, WM78].

L [Har70a, Har70b, Kuc75, Mat85, Mig79, Pre71c, Rei75, Shu80, Wel80b, Woo77b, Wri71a, Wig95]. Labeled [BTSKC96]. Laboratory [WH70, Col92, Ker74, Ric70, VC70]. Lack [Kil93a, WL81]. Lafferty [WH70]. Lancaster [Ano74b, Her82, Sal72a, Swa76b, Swa77b, Swa78b, Mai99, McC97]. land [KW77]. Landes [Wel71]. Landmarks [Gro77]. landscape [Buc99, Lam99]. Language [SD96, WNAR99, BJ79, BW84, BC00, BBMS87, Dos86, ER00b, FK73a, Gri84, HH95b, HH95c, Jr88, Ker74, Kim82, KPZ^{+}99, KG72, ME98, Mud98, Pep74, RSW91, Sal70, SSM93, PW92, Coh95, Joh84, Mac79b, Fis79, All98, Lev76b, O’C93, She91]. Language- [Coh95]. Language-operational-gestalt [Lev76b]. languages [Avr71, GS78, Lea75, McL90b, Sch82, Sve94, Wei96a, Zvi87]. Large [Kaz79, Whi00, Zam80, Bla80, Cag70b, Cag70a, Che72, CN95, CBM^{+}94, Cro77, Fay72a, Fay72b, Fay73, Gri88, Haw74, KSDR90, LJ70b, Lyn88, MM98, RCZS99, Ras80, RT81, RSW91, RF97, Sal72b, SR77, SF78, WC77, WSS79]. large-information [LJ70b]. larger [Cra91]. Larne [Cal00]. Larry [Lap00]. Laskowski [Har83]. last [HW99, Kra88b, Kra92b, Nis99]. latent [DDF^{+}90, GD98]. Latin [ASR85]. Laura [Kuc75]. Laurence [Rod93]. Law [Ere81a, Lon92, Cre95a, DG78b, Egg99, Mat97a, Mil90, OV97, Sut94, Yan90, Bas92, Bas94, Bro79, CL86, CL90, Che95b, CCT95, DG78a, Egg86, Egg90b, Egg94, Fed82a, Has76, Hub76, Hub77, Hub78, KH95, Kri77, Lou92a, Mur73, Nic89, OV93, OV95, Pao86a, Pao86b, Pop75, RK79, Rou92a, Sch75, New76, Mar96]. Lawrence [VC70]. Laws [Coo89, Egg96, KK96, Rou90]. layout [Smi81]. LC [Gro77, Har77a, Har77c, Lar74, Lil78, Ric77, Swa77b, Vag80, Aul71]. LCM [Hay96b]. Lea [Sch91a]. leading [ER95]. Learned [Har70c, Bor86b, WMME88]. learner [AL98, Yan97]. Learning [Bur96, Dav96a, Lin96, Nah98, Sut96, Afs98, Boo83, Che95a, CSSI98, Fau96, Gas99, HC94, LC84, LBBR96, LK97, NJWC97, OO97, Par83b, RFK97, Sew82, SBW90, Var99, Wee99, Wit96a, Wit96b, Fis81a, Wit96a, Wit96b]. leasing [Sle94]. least [Chr95]. Lebanon [Pic86]. Lecture [Lei73]. Legal [RRB91, Sor96, BS88, BCM^{+}70, EK97, Gas98, Haz98, MUD99, Sha92, Wei96b, Ere80]. legislation [HC84]. legitimacy [SPWH99]. legitimize [PPP98]. Lehrinstitut [Wel71]. Leigh [Bro00c]. Leimkuhler [Egg90b]. length [BK79, CL82a, Han91, Kuc78]. Leny [Bas74]. Leone [Har77b]. Less [CCW84, Ere81c, EN85, Ere85, Gor79]. Lessons [WMME88, Bor86b, Klo89]. Lester [Har83]. let [Nug91]. Letter [Aul79, Bar74, Bez82, BK79, Bro00e, Dav79a, De 80, Die79, Flo99b, Gol79, Gro78, Hen97, Hen00a, Hen00b, Hjø00b, Hum00, Koc84b, Laz00, Mea00, Pit82, Pom74, Pom75, Qin00a, Ros82b, Rot79, Sal79a, Sal79b, San79a, Sew82, Spe76, Tho79, Vin99, Wel81, Wil82, Zar82, Bra92, Buc92a, Deb92, Fid87, FS84, Her96a, Hir90, Hol87, Kra92c, Lou92b, Sal87a, TZ93, Woo84, Mea82c]. Letters [Ain85, Aiy88, Aji92, AI76, Ano74f, Ano75f, Ano75g, Ano81c, Ano86e, AL98, Ath72, Bag81, Bal84, Bal88, BHCC^{+}74, Bas94, Bat00, Bat83, Bax85, Bel77, BA74, Ber72b, BB77, Boh90, BZ85, BSR94, Bor85a, Bor97, Bor74, Bor87, Bra92, BL72, Bri72, BPK^{+}71, Bro79, Bro90a, Bru78, Buc91a, Buc92a, Bun80, Bur71, Bur93, Bur94, Byr76, Can95, Cha77, Che95b, Con84, Coo91a, Coo76, Cou98, Cox83, Cra72, CHPC71, CRC70, Cur93, Dan00, Dav74, Dav84, Dav92b, Dav79c, Deb72, Deb92, DRM98, DO81, Dic96, DH78, Din87, DG78b, Dum87, Dun80, Dun96a, ER97, Ell86, Ens95, Eve98, Fei73, Fid87, For91, Fre80, FB81, Gar79, Gar83, Gar91, GS97, Gar97, Gar98, GD90, GT70, Gli78, Gol87, Gol74a, Gor92, Gri80, Hal95, Han92]. Letters [Har93b, HN98a, HN98b, Haw87, Her96a, HD72, Hir90, Hjø96, Hol87, Hor78, Hor93, Hub77, Hum96, Hun72, Ing97, Jac81, Jon73b, Jon89, Jor89, KBR^{+}71, Ker84, KW81, Kno77, KM78, Koc80, KBN90, Koe81, Koe88, Koe92a, Koe92b, Koe94, Kol74, Kra74a, Kra85, Kra92c, Kra74b, Kri77, Kuc80, Kuv73, Lan81a, Lan89b, Lan97, Lan98a, Lan98b, Lar92a, Lev73b, LP83, LHM^{+}86, Lev98, Ley98, Lin77a, Lin77b, LS80b, Lin87b, Lip83, Lon92, Lou91, Lou92b, Luc93, Luk96, LD97, Mac72, Mac97, Mar78, Mar72b, Mar72c, Mar88, McC97, McM73, MZ80, Mea83d, Mea84c, Mea85, Mea87, Mea97, Mea98, Men95b, Mig80b, Mil87, Mil95, Mos72, MOV83, Mur74, Nad85, NCH^{+}71, Now81, Old88, OV93, OV95, Opp74, OG80, Opp98, Ozo80, Pap84, Pap88, Pap93, Pap96]. Letters [Pap98, Pat81, Phi92, PTH83, PLM^{+}79, Pri72, Pul74, Ric97, RHH^{+}70, RC74, Rot81, RV98, Row84, RTH71, SC80, Sal84, Sal87a, Sal96, San75a, San79b, San83, SM78, SB73, Sch91b, SLW70, Sha84, Shc79, Shp92, SB80, SG89, Sme87, Sme92, Sor96, Ste84, Str84, Stu78, Swa89c, Swa89d, Tau84, Tem94, TL74, TZ93, UAW82, Van98a, Van80, Vea90, Wal72, Wei90, WM78, Wel87, WC85, Whi94, Whi96, WM00b, WS75, Wil70, Win95, Wit80, Wit87, WA97, Woo76b, Wyl87, Yer91, Zhi97, vB81, Ano78b, Kra91c, Yer71a]. level [BS91b, Coo71, Erc98, LS70, Los87, SW92]. Levels [KL79, Pic96, PM00, QLA97, War73]. Levine [Ere81a]. lexical [LG99]. lexicon [Sta98]. Lexington [Nov82, Pre71a]. LEXtractor [BM75]. Leyden [Hei78, Swa78b]. Leydesdorff [Cou98]. LHNCBC [Woo76a]. liable [CH99]. librarian [Erc95, McC78, Sou96, Bud91, Hei86b]. Librarians [Dav96b, Joh86, Aul87, Hay95, Mat85, O’C95, Pec85, San77, Fis99, Hei94, Ols89, Rus87, Smi98a]. Librarianship [Ano77a, Buc91b, Shu78, Tau74, And90, Gro75b, SS95, Don77, Mar77b, Pre70, Sch75, Swa77c, Bix96a, Bix96b]. Libraries [BL85, But99, Cre99b, Cre99a, CH99, DS79, Day99a, FHN^{+}93, Fra99, GL93, Gas99, Har77a, Hin70, HOS^{+}93, KMG^{+}93, Lan77a, Lap00, Lev99, LF93, LCB99, Pre71c, Raw93, RM99, SR79, Ste99b, Sun99, Wel71, Wil98b, You75, Afs98, Aus93, Bat98, Bat83, BB83a, Bea80, BNS^{+}00, Bix96a, Bix96b, Blu89a, Blu89b, Bom75, BCG79, Bro97, DH70, DDC78, Dep75, Des97, DCD81, EK97, Erc98, Fau96, Fu96a, Fu96b, GO93, GMPS00, Har70c, Has85, Hen95, Hoc89, How87, Kan94, KS70, Koe88, Lin87a, LD85, McG73, Pal87, Par00, Qua96, Ric87, Rou76a, RMMS96, San98a, Sav97a, SC76, Slo00, Swa78b, Tom92, Tom99, Wil97, Xia92, Har77c, Swa77b, Har76a, Lil78, Ols76, Rei85, Wil00b, Gro88, Har93a, Boy86b, Bro86a, Dav92a, Eva94, Lyn90, Mil90, Sch96b, Shi86b, Und98a]. Library [AL96, AKS96, Abe96a, Ano80i, Bai75, BHWG95, Boy85a, BCB96, DS79, Don76, GS87, GBDH76, GWDS86, Gro75a, Gro75b, Gro77, Gul85, Har75d, Har75c, Har76a, Har77b, Hay96b, How89c, Kaz77, KMB99, Lan77a, Lap99, Lar92a, Lin87a, Mom82, MBFK76, MY85, Pec85, Pre71b, Rei77, Ric77, Rou76a, Rus85, Shu80, Smi98c, Ste77, Ste80, Swa76c, Swa77b, Til70, Tsa98, War79, Wel75, Wes88, Afs99, AF98, Ano08, Art76a, AM80, Bak89, Bal00, BS91a, Bar96, Bas74, Bax85, Ben00, Bor84, Bro70a, Bro70c, Buc79, Buc98, BMFK76, Bur90, BF94, Cas78, Che72, CB83a, Chr95, CV80, Coo77a, Coo85, Cor94a, Cor94b, CK96, Dal90, Dan84, DL00, Day99a, DRM98, Don79, Dun96b, Eps80, Erc95, EP97, Fay72a, Fay72b, Fay73, Fis70, FHN^{+}93, Glo88, Gro76c, Har70b, Har78a]. library [HNW93, Har76c, Har76b, Hay88, Hei80, Hei91, Hei86a, Hei87, Her97, HJD^{+}99, HCL^{+}00, Hir89, Hoc74, HOS^{+}93, Hol99, JJ77, JJ80, Jai72, Joh86, KR80, Kan76a, KN00, Kaz75b, KS77, KS70, KRD95, Kra91a, Kri86c, Kuc75, Kuc76a, Lan87b, Lan90, Law74, LC71, Lev76b, Lil78, Lin77a, Lin77b, LDS^{+}00, Los90, Mac72, MB81a, Mar71, Mat97a, MB71, MB81b, McG84, McG96b, Mic94, MN81, Nan70, Nis99, Old91, OG93, Pas90, Pat90, Pen93, Ras90, Rie88, Rob80a, RT81, Rou75, Rou76b, RR78, Rou79, Rya93, RB72, San98b, SP73, SK97a, SK97b, Sch96b, Sha78, Shi85, Shu78, Smi81, Smi84a, Smi00a, Sop96, SS76, Spe71, Sta75b, SHS72, Ste74b, Swa75b, Swa76b, TC00a, Tor89, Tur77, Usd85, Wal00a, Wal85, Wal86c, Wal86b, WA79, WC77, WSS79, Whi89a]. library [Whi89b, Wil00a, WH98, Woo84, YCO00, YG90, Ano77a, Can99, Cas79, FS97b, Hor78, Lar92c, ODVG87, Pre71c, Sha75, Wes00, Wri71a, Zwe95, Don70, FW88, Tom98b, Bea93, Cre95b, Mar96, Rei75, Swa76a, Wei79, Hei90, Hjø99, Lie96a, San98b]. Library/ [FW88]. Library/information [MY85]. Liebesny [Tau74]. Life [Cha91, Tay82, BR94, Bot70, Bro80b, Cha96, Cha99b, Lev82, Mac96, Mea95, SK99, Tsa98, VS87]. life-world [Cha96]. Lifetime [GdCD81]. Lifetime-citation [GdCD81]. like [Kue72, Pai71]. likeness [Fri97]. Lima [Har70a]. limitations [Boy86e, PW91]. limits [Cre99a, Sun99, Sve94]. line [AB76, Bac72, BR95a, Bou80, Coo77b, CR80, Cua71, DH78, Elc78, Ell76, Fay72a, Fay72b, Fay73, Haw77, JJ77, Kat72, KM73, Koe79, LS80a, LWG74, Mar00, McC80, Mec99, MRW73, Mog75, Mor76, Nie76, PK00, Pet98b, Ras80, RBR80, Ros79, Sch00a, SWK78, Slo00, SR77, WL80, Wil77a, Wol71, Ano74b, Mig79, Tes79]. linear [Mat00, WY90b, WYSB91]. Linguistic [KK96, Whi98d, BP93, War91, Art76b]. Lingüistica [Har70a]. Linguistics [Ano71a, Bri72, Mon72, Gli72, Wal81, Wei96a]. link [CK99, Cas85, Cro77, Her78, KL98b, KL98c, San99]. linkage [Wil84, Wol96]. linkages [War88]. linked [Lyn90]. linking [LR79]. links [BTSKC96, EFW96]. Linnet [Gro76c, Lan78]. LIS [Whi96]. LISREL [VTS97]. LISREL-based [VTS97]. List [Flo99a, RG70, Haw86, JD78, MDS70, Mil87, Moo73, Ros76, Sav99, Wig95, Tho73b]. Lister [Fis81a]. Listing [Hin70]. lists [Hir78, WFK72]. Lit [Wei79]. literacy [Mal99]. literary [Beg95, War93]. Literature [Avr79, Bon82, FM97, Gro77, Har75a, LG99, McM76, Poy82, Swa77b, TL74, Wal97, WWB97, Bai76, Bau72, Ber71, BB95, BCM^{+}70, BM81, Bro70a, Bro80b, Bun80, CH00, Cha76, CMK98b, Del90, DS93, Don72, DG78a, Eft90, ER00c, EN85, FCJ^{+}92, FG70, FRGL75, GL96, GD98, Gup90, Hap99, HM87, Hun72, JJ75, KDS89, KB92, Koh74, KH71, Kuc82, Law76, MP74, McC89, McC91, MWH^{+}95, MG72, PN79, Pop75, Qin00b, Ros79, Ros82b, Rot79, SP73, Sch74, SG80, Sha92, Sie90, Sma73, SG89, Sma80, Ste85, Ste90, Sum84, Tay95, Tho71, WG81, Whi81, WW73, NPG76, RK79, Tau74, Bus95]. Literature-based [LG99, GL96]. literatures [Ano75a, Ber70, Bro90b, GWDS86, Nee81, Pie99, Swa87, Swa89b, Swa89a, Swa89d]. Little [Mea82a, KL79]. Liverpool [Swa76c]. lives [Sem98a]. Living [Bat97]. loan [AF98, BF94, Gas99, RR77]. loans [DG78b, Spe71]. local [Don70, Ren80, Lee96b, Tom98b]. Localist [Che99]. Locating [Jaf98, Kil93b]. location [Can89]. locator [MMB99]. logging [AB76]. Logic [Wat89, HT91, SMB72, WD97a]. logical [Fuh00, Ham00, Soe94]. logically [Swa87, Swa89b, Swa89d]. logically-related [Swa89b]. logics [Ozo79b]. loglinear [EP91, Eve93, Gle93]. loglinear/MDS [EP91, Eve93, Gle93]. London [Bla79, Kol75, Lan81b, Lea76, Mig79, Pre70, Tau74, Tes79]. long [KPZ^{+}99, Kuc76b, Ora00, Vin74, Lar91]. long-range [KPZ^{+}99]. Long-term [Lar91]. Longfellow [Lou92b]. Longitudinal [McC84, Yua97]. look [Bea80, MBH94, Wat98]. looks [Har71c]. loop [AdKK87]. loss [Pap84].

Lotka [Lon92, CL86, Coi77, Egg94, Hub77, KH95, Kri77, Lou92a, Mur73, Nic89, Pao86a, Pao86b, RK79, Rou92a, Sch75, Voo74]. Loughborough [SOM^{+}99]. low [Mar83c, Sou96, SSD99]. low-cost [Sou96]. low-frequency [SSD99]. lower [FS84]. Ltd [Art75b, Mar85]. Lubans [Wri71a]. Lucy [Whi00]. Luddite [Cal89a, Cob96]. Luther [Ric77]. Lyle [Har75d]. Lynch [Art75b, Bro00d].

M [Ano74a, Art75b, Che99, Cri81, Dav94, Gro76a, Har78a, Har78b, Kuc75, Lar74, Lev76a, Lun80, New73, Rei75, Rus77, Sch91a, Shu80, Vag80, Wel70, Woo77b, San74]. M.I.T [Har70c, Har70b]. M.I.T. [Bai74]. M.S. [Flo00]. MA [Har79, Lan77a, Nov82, Rei75, Rou76a]. Maccafferty [Lan81b]. Machine [Art76b, Che95a, Haw81, HSP75, Luc93, Mer71, Til70, Blu71, Byr75, CSSI98, Gol74a, Jah74, LC84, Lee96a, Lee96b, Mel72, SGS70, Wil71, Bur92a, Mea82a]. Machine-assisted [HSP75, Mer71]. machine-based [SGS70]. Machine-Readable [Til70, Haw81]. Machines [Don70]. Machlup [Kin83]. Macleod [Cra96]. MacMillan [Mim85, Lan76, Mim85]. Macro [Kir81, Koc83, Wal98]. Macro- [Kir81]. macro-theory [Wal98]. macroeconomics [McC84]. Macrothesaurus [Wei88]. Madame [Buc95]. Madan [Smi00a]. Made [RG70]. Madison [Han75]. madness [Hen95]. Maguire [Var75a]. mail [HBG96, May97]. maintenance [Boy86a, GGJ91, Hum84, MN81, Ros76, Lea75]. Major [BR94, Har94, Lan94, LB94, Pet94, Sap94, WT94, Ric77]. maker [YdKKM87, YK93]. makes [JB71b]. Making [LWG74, TC73, BF88, FGS81, GO93, Hig99, Joh84, Jul99, KdK90, Rei75, Sol97a, Sol97b, Sol97c, Tay94, War81, YF85, Bur96]. Malay [AYS96, BSY00]. Malinconico [Bel81]. mall [FDD^{+}99]. Man [Art79, Gol74a, GBL89, Jah74, Mel72, Wil71]. man-machine [Mel72, Wil71]. manage [Hen99]. Management [BD88, BFM88, Elc73, FW88, GL93, Ger79, LD88, LB88, Pas90, PW88, War88, Wes88, WMME88, ZvBR88, ABB^{+}91b, Alg95, Aus93, Bar95, Bar98b, BS91b, Blo70, Buc98, Car95a, Cli70, Coo85, Cox93, Cox94, EK75, Eva94, GO93, HM70, LL85, Lin70, Liu96, LFHG99, Lyt88, Mad96a, Mad96b, Mat70, Mon94, Nee81, OG70, Par82a, Par82b, Par82c, Rep89, RT81, Sav97a, Spr87, Sta75b, Swa78b, WWSR00, Wei94b, YCO00, FMN82, Kra79, Dom78, Sta75a, Blu89a, Bor90, Rum90]. management-by-objectives [Sta75b]. manager [Sop96, Wal86c, Erc95]. Managerial [Sta77a, Wol86, Mos70]. managers [Wei94b, Pec85, Mai90, Miz98]. Managing [LFS72, SNG00, Con93, Kaz75a, Sle94, Tom96]. Mandelbrot [Egg99]. Manfred [GBL89]. manipulation [Les89]. Manpower [Man82a, GSD74]. Mansell [Kol75]. Manual [Tes79, Elc78, LT74, McG73, RC95, Ten94, Wei94b, Boy86c]. manuals [Gro90]. manuel [Gro76a]. manufacturer [Var75a]. Manufacturers [Var75a]. manuscript [HP00]. manuscripts [Rou89b, Rou89a]. many [LS70, Pai90]. Manzer [Har78b]. Map [Lin97, CHSS98, Ley97, NvR98, Whi83, KDJ94, Sch75, WNAR99]. Mapping [BSM90, BZ85, BMvR91a, BMvR91b, HBJR99, McC90, McC91, WC84, vdBL96, EP91, Eve93, Gle93, KS98, McC84, McC86, NML99, Sma99]. Mappings [Len83]. Maps [PLM00, Erc98]. MARC [Gro75a, Rou89b, Rou89a]. Marcel [Bai75, Dan84, Swa75a]. Marcelle [Wel78]. March [CML^{+}96]. Marchant [Har78a]. Marcos [Har70a]. Margaret [Gro75b]. Marion [Har75c]. Mark [Mim85, Wil00b]. marker [CK99]. Market [Bre79, SS93, Mow92a, Mow92b, Mow92c]. Marketing [Kin82, Kue72, Tho82, Lun76, Ken87a, Bak89]. Marketing/planning [Bak89]. markets [Sch88b]. Markov [BT84, Qiu93]. Marsden [Bal89]. Martin [San99, Woo77b]. Martinus [Wil85b]. Mary [War79]. Maryland [And91, BR94, Wes00, BD88]. Mass [Dav76a, Har70c, Har70b, Pre71a, Sie89b]. Massachusetts [Ano75a, Rus85]. Massive [Coh99]. Master [CH73]. Masthead [Ano70e, Ano70f, Ano70g, Ano70h, Ano70i, Ano70j, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano73g, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano75m, Ano76b, Ano76c, Ano76d, Ano76e, Ano76f, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano79b, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano80g, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano81i, Ano82e, Ano82f, Ano82g, Ano82h, Ano82i, Ano82j, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano83g, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano85b, Ano85c, Ano85d, Ano85e]. Masthead [Ano85f, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k]. Matched [Per82]. Matching [YAH90, BSY00, BS92, CPW88, FW90, Gal91, Gre95, Los87, Yu76]. material [Hir89, Par82b]. Materials [Gol73, WA96, BMFK76, Mer71, OMY^{+}89, OG93, Sve94, Sop96]. Mathai [Bai75]. mathematical [Dom00, Egg93, GdCD81, Lyn77, ML90, RR71, Sch76, WD97a, Rou79]. mathematicians [Bro99c]. matrices [Cag70b, Cag70a, HW80, Wri72, Wu98]. matter [Jon73a, KM00b, Mar95, RW82]. Matthews [Boy85a]. Maureen [Tau74]. Maurer [Str78]. Maurice [Bel81, Har78a, Har75c]. maximally [Ger78]. Maximum [Bal84, Coo83c, KL98a]. Maxine [Lan81b]. May [Joh86]. May-3 [Joh86]. Maybeth [Woo77b]. Mayor [Har70a]. Mayra [Gro75b]. MD [Cer83]. Meadow [Whi99b]. mean [Egg93, Egg97, Erc98]. meaning [Dun96b, Lan82b, Mar77a, Pie92]. meanings [Gar98, HN98b, War96a, War96b]. means [Ell76, FK73a, Pie99]. measure [Bel73, BM81, Car79, Coo73a, Coo73b, Egg87, HN97, Kwo85, Mat00, Pra77, Ras90, Sma73, TvR90, WG81, WY92, Yu76, ZK99a]. Measurement [Ber71, Bro80a, Mcd82, CH73, Ell96, Goo97, GLS97, Har96, Kan76a, Lev76a, Pen93, SWY81, BR80b, EFW96, Rag96]. Measures [Keg70, RB72, BE91, Bur92b, CC98, Coo71, Cro94, ER91, ER00b, Fal98, Fri97, JF87, Los00, Men85, MM78, PT87, Rad82b, Reg88, Ros71b, ZK99b, ZPB98, Fen81]. Measuring [HJ89, Kid90, LP99, VTS97, Yao95, LG96, Tur77, AP96b, AP96a, Los96a, Los96b, Sal86, Sam95]. meat [Gol87]. mechanical [Mat97b]. mechanics [Bar78, Mud98, Lyn94]. mechanism [FSV93, LCL85, OOA99, OV98]. Mechanisms [Abe90, Col96, Mon81]. mechanized [SJ72]. Media [Bai75, Fau96, Hol96, KL79, Lan87a, New94, BL72, Cit98, DHS00, Gel96, HW98, How87, KM00b, Pic86, Whi99b, Cal89b, Day99b]. median [Egg97]. mediated [HJ89, McD99, Spi97b, SBW90, WKMG00, Whi98d]. Medical [Cim95, DG95, Gor95, HH95a, Her95, JS95, LH95b, MWH^{+}95, OWF99, Rod95, SM91, SMJ95, WdBFF95, WNAR99, ZvBR88, Anv91, GQSdP98, GWDS86, He98, He00, HH95b, HH95c, HM87, LH76, MN83a, MQP98, Spe71, Swa87, Swa89b, Swa89a, TC00a, Tsa98, McC89]. medicine [Her94, HH95a, Hum96, SMJ95, GWDS86, Lan77a, Lin87a, Rou76a, Wes88, Shi88]. medicus [Tho72, Sma80, Vir70]. Meditz [Woo77b]. medium [Dav88]. MEDLARS [Sal72c, Mil71, Vir70]. MEDLINE [Fed82b, HT91, Hum84, McC80, MSJM91, Pat81, Tag75]. MEDLINE/full [MSJM91]. MEDLINE/full-text [MSJM91]. meet [Yee91]. Meeting [Ano84j, Bor73b, Boy91, Mac79a, Mos72, OCN97, Ano74f, Bor74, GM86]. meetings [Dav84, Mea83d]. meets [BT00]. Melbourne [Gul85, Rei77]. Melvil [Gro75b]. Melville [Gol74b, Lea75]. Melvin [Mar77b, Mig80a, Pre70, Swa77c]. Members [Ano75e, Ano75c, Ano75d, Flo99b, LBIP97, Mar92b]. membership [Flo99b, Hor78]. memex [NK89, Buc92b, Bur92a]. Memoriam [Ano72e, ATJ95, Bor99, Boy95a, Coc98, Cra96, CML^{+}96, Dav94, Deb96, GBL89, Hay96a, Hen96, JK91, Kra99a, Nel97b, Ols99, Ras99, Rob98, Rod93, Sch91a, TS92, Wei97a, Ano73h, Ano74m, Che86a, Koc84b]. Memoriams [Ano73h]. memory [Bal89, Cas91, Mon81, NK83, OWF99, WM00a]. mental [Sut94, TFTB94]. mention [Her78]. menu [Shn86]. menus [CCN93, TK96]. merger [Ano72c]. merging [Boo78]. Merit [Coc98]. mes [Kra91a]. MeSH [Hum84, Kir81]. MeSH-derived [Kir81]. message [CM00, May97, NKB72]. messages [May97]. Meta [Hig99]. Meta-information [Hig99]. Metadata [Zen99, BNP99, Cor99, GSKL00, HJD^{+}99, Sut99]. Metaknowledge [Con95]. metalanguage [Pep74]. metaphor [Cas91, Day00a, TK96, Wal00b, Bro96]. metaphors [Ham00, Smi97]. metascience [OD70]. metatheory [Dow77]. Metcalfe [Sta77b]. Method [Pai96, Bro96, Bur99, CC92, Cro77, Hir78, LWZ99, McG84, Per92, PN98, Ror93, RF00, SG80, Smi84a, Ste85, Tur77, Wal86a, WY93, YR77, ZK99a]. methodological [IC97, Kim73, MM88, Mol95, SD97, WV97]. methodologies [ASAOE99]. Methodology [Dav79b, Mcd82, Har90, MWR72, SKL97, SKCT88, Sum84, Was80, SK97b]. Methods [Bes96, ERV00, BSY00, BG98, BHH96, Bri91, DC98, EQH00, Fis99, Goy84, GWDS86, McL79, MN83b, MQP98, Ros71a, SFV85, Wal85, Wil92a, Wel80b, Bel76, Lan89a, Liu98]. metrical [Bro80a]. metrics [Ror99b]. Metuchen [Art77, Gro77, Har75d, Har77a, Har78b, Kuc75, Mar75, Rei77, Rus77, Shu80, Sta77b, Ste77, Wei79]. Meyer [Pre71c]. MI [Var78]. Michael [Afs99, Bel81, Don76, Har77a, Mar77b, Mim85, Swa77c, Wei90]. Michie [Mar85]. micro [How87, Wal98]. Micro-studies [Wal98]. Microcard [Don74, Har76c]. Microcomputer [Cra80a, KDS89, Pas90, Sie86]. Microcomputer-generated [Cra80a]. microcomputers [Har83]. Microfiche [Chr72, Tsu79, Ger79, Kuc76a]. Microfiches [Dav79a]. Microfilm [AD79, Tin70]. Micrographics [Gro87]. Micronesian [Wan92]. Microprocessors [Har83]. microthesauri [Kir81]. Mikhailov [Koe93, Koe94]. Mill [Ano74a]. millennium [Kra99c, Bro99a, Nan95b]. minds [WKMG00]. mined [WBX98]. mini [AN73]. mini-computers [AN73]. miniature [Kil95a]. minicomputer [HC90, KH80]. minicomputers [BJ79, Swa76b]. Minimal [Erc98]. minimum [CSI87]. mining [Ano00a, LY98, Sch97b, Exn98]. Minnesota [Gro90, BDBR97, BD97]. Mira [Dun00]. Miranda [Boy95a]. mirrored [BBDH00]. MIS [Mai90, Lev73a, Fis81b]. Miscellaneous [Ano78j, Ano80h, Ano80i, Ano81k, Ano81j, Ano83h, Ano84j, Ano84k, Ano84l, Ano84m, Ano85g, Ano85h, BL85, CL82b, CR80, Lun82, LE85]. misconduct [Whi98c]. misinformation [Fra86]. misinforming [Fra86]. mismatching [Pap88]. Misnio [Har70a]. misnomer [HN97]. Missing [Hjø00b, Cas85]. Mistakes [Mur74]. Misunderstanding [Cal89b]. Mitchell [Dro78]. MITRE [SL83]. Mitroff [DMZL82]. Mixed [MCD73, Smi81]. MLIS [Dou96]. mode [AM77]. Model [Dav92b, Fed82b, MBFK76, ABB^{+}91a, AL94, Aji91, Aji92, AP96b, AP96a, Anv91, BS91a, BCB95, Bat86, BT84, BSR94, BCG79, BP93, BF91a, Buc75, BK81, Bur90, BF94, Cha86, CL89, CVK92, Cle76, CCW84, CM00, Coo72, Cra81a, Dal90, DC98, Dep75, DGS86, DS73, EP91, Eve93, FS97b, FWR95, Fid84, FGS81, For99, GdCD81, Gle93, Gol94, Goo97, Haa96, Har84, Har93b, Hay96b, Hei74, Hur00b, IS98, Jr88, Kan87, KDJ94, KS98, Kra78, KH71, LMSC82, LT74, Lin77a, Lin77b, LY98, Mac91, ML90, Mol95, Nan70, Par82c, Qua96, RW86, Rou74a, Rou76b, RR78, Rou92b, San79a, Sch76, SC76, SC78, Spi96, SSD99, TZ92, TB93, VD98, Wan92, WS98, WW99, War73, Wil92b, WY90a, WYSB91, Wor73, YK93, ZvD00, ZPB98, Kei70, Liu97b, MR87, Sri90]. Model [dSM93]. Modeling [CP98, GCS97, SW91, Aik97, BSR93, Che97, CD94, CC92, HMNW86, LS92a, MR87, Nic89, Rou79, Tur77, Yar97, Yin95, Zai77, Col99, Col00a, Che86b]. modelling [Jär89]. Models [FC92, Lan77a, Mcg79, San77, SR79, Yer73, Bod99, BS74, Bri91, Col83, Col96, CK96, Cro81, Cro86, Kha91, KS90, LC71, Los88, Mac72, NKB72, NT85b, OV97, Par82a, Par82b, Par82c, Qiu93, Rou76a, Sim81, Sut94, TFTB94, Wel00, Wil93a, WY90b, Los90, Mar82a, Nei80, Wes89]. Modern [Dio83, Hjø99, Rei84, SD96, Miz98]. modernization [Lan90]. Modernizing [Men95b]. modes [SHS72]. modified [Moo73]. Modular [WA96, Har74]. Mohan [Smi00a]. molecular [SS93]. moment [Cro99a]. Monica [RCJS94]. monitored [RB83, WFK72]. Monitoring [NvR98, BHH96, Coo98, Ros73]. Monograph [Pre71c, WSS79]. monographies [Gro76a]. Monographs [Lan77a, Rou76a, KF92]. Monte [Bur99]. month [She72]. Monthly [Bol86b, Har71b]. Mooers [Hen96]. Moore [Mar80]. Morse [BT84, Che86a]. Moscow [Rya80, Sch73b]. most [Cam96a]. most-cited [Cam96a]. motivation [MM88]. Motivations [Kim00, Bro85, Bro86b, MBS98]. Mott [Bas74]. Mount [Ols76]. movement [Man82b, Whi83]. movements [Wil97]. Moving [SH94]. Mu [Har75c]. much [OMY^{+}89, MR94]. Muchen [Kuc76b]. multi [BCB98, Egg99, MCD73]. multi-purpose [MCD73]. multi-query [BCB98]. multi-word [Egg99]. multiauthored [Opp98, Van97b]. multiauthorship [Rou92a]. Multidatabase [SLC98, NL83]. multidimensional [KS98, McG84]. multidisciplinary [Ker74, McC95, SD83, SF78, Stu72]. multidiscipline [Pai90]. multifaceted [LH76]. multilingual [BC00, Mor78, PN96, Sta98]. Multimedia [Beh96a, Beh96b, Gel96, Gor92, LBBR94, LBBR95, DW98, Ens95, Goo99, JM97, LBBR96, Ste99b, Sun99]. Multinational [Vei88, Nov82]. Multiple [Spi96, SR95, Egg96, Erc92, FFKS99, Gar98, HN98b, HS78, KSDR90, Kwo99, Rod95, SJ72, YR98, Her78]. multiple-criteria [HS78]. multiple-observation [Erc92]. multiple-term [SJ72]. multipliers [Egg90a]. multiuser [McD99]. multivariate [Ami70]. Multiversion [FSV93]. multiweighted [Goe87]. mumps [EM94]. Munchen [Har76b]. Mundaneum [RL97]. Munich [Har77b]. Murdick [Fis81b]. museology [SS89]. Museum [Mar99, BT00, Mar00, PBL^{+}00, PM00, SBPP89, Wal00b, YCO00]. music [Coo89, DS93, BL72]. musicology [Pao82]. mutual [Jon73c]. mutually [Swa89a]. my [FS97a]. Myers [Dav77b, Gro75b]. Mysteries [Nei80]. myth [Cov00, Lie96a]. myths [Smi97].

N [ATJ95, Fis79, Hen96, Pet99]. N.J [Art76a, Har75d, New76, Sel71]. N.W [Rus85]. N.Y [Don70]. NACA [Joh84]. Nachrichten [Wel80a]. Nacional [Har70a]. Nahum [Sal87a]. naive [Han86, NK00]. Name [Col00b, BS92, Fag95, SM90]. name-matching [BS92]. names [Bla71, Cor82, SR77]. naming [Col95]. NAPS [Kra74b]. Naranan [Hub76]. narrative [Sko98]. Narratives [DHS00, Sma86]. NASA [RTM99]. Nathan [Har78a]. nation [MY85]. National [And94, BL85, Bor85b, Bra87, Gro75a, Gul85, Har70c, Has85, Lev73b, NC75, Nit79, Wes88, Woo77b, WH70, Bur00, Don79, FB78, Hal85, Lan90, Mou77, Neu95, Old91, Whi77, ZPB98, DB94, GWDS86, Lin87a, MH87, Spr87, Woo87b, Har76b, Har77b]. nationale [Gro76a]. Native [Gro96]. NATO [Hei78, Swa78b, DH97a]. Natural [Dos86, Gri84, Mud98, BW84, BBMS87, CK96, HH95b, HH95c, Kim82, KPZ^{+}99, ME98, Pep74, RSW91, SSM93, War98b, PW92]. natural-language [BBMS87, PW92]. nature [Bel81, Cer79, Day00a, Jan91, Koc84a, Mow92a, MM78, Ott94, Rob78, Shp92, SK99]. Nauka [Rya80]. nauki [Rya80]. Naur [Bel76]. navigation [IS98, JGR^{+}95]. NCSTRL [DL00]. nearest [Wil84]. Necessary [Hei80, Hei91]. necessity [BSR98]. need [Cur93, FB88, Hom83, Tem94, Ste00a]. needed [Dre99]. Needs [Mil94, Var75a, BT99, CB83b, Gor95, GM86, GSD74, Har94, WA79, WA96, Man82a, Deb86]. negation [ME98, NE90]. neglected [Sha92, HH70]. negotiation [MB91]. neighbors [Wil84]. Nelson [Die82b, DMZL82, Pre71a]. Nelson-Hall [Die82b]. neo [Cob96]. neo-Luddite [Cob96]. neon [Wal00b]. Nephis [Cra77, Cra82]. nested [Cra77]. nested-phrase [Cra77]. Net [TvR90, Bra94, CN95, KS97, Smi97]. Netherlands [Hei78, Swa76c, Swa76a, Swa78b]. Netizens [Myb98b]. Nets [Nov82, Vei88, TK96]. Network [Gul85, Nit79, SR79, WWSR00, Ano70b, CN95, Coo77a, Don79, DF85, Dor85, Dor88, KR80, Lin77a, Lin77b, LR79, Lyn98, Nan70, NvR98, PKK74, PR98, Rou76b, RR78, SS76, SB94, VS99, Woo76a, Yar97, HK79, WCL87]. Networked [AL98, JR96, BC97, HW98, HJD00, MBH94]. Networking [Swa75b, AM80, DC98, Lyn90, Tom98b, Woo76a, Sha98]. Networks [Mea86, Wil77a, ALD96, Ano80i, Caw74, Che95a, CV80, Eps80, Gul85, Has95b, KRD95, LAD97, NKB72, Pai90, Par97, Rei84, Rob80a, RR77, Sch97b, Smi81, Ste80, Wit96a, Wit96b, Koc79, SWT^{+}79, Rei85, Boy88, Che86b, Gar86]. neural [NvR98, Sch97b, Che95a]. Neurological [WCL87]. Neveling [Gro76b]. Newkirk [Har77b]. news [HDRD88, WSB98, WW00]. newsgroups [Kay98, Osb98]. newspaper [LWZ99]. newspapers [DKD71, San99]. next [LMO^{+}96, Raw93, Ric87]. next-generation [LMO^{+}96]. nexus [War99b]. Nigerian [Ehi90]. NII [FW99, HW99]. NIICC [SK98]. Nijhoff [Wil85b]. nineteen [MCT73]. nineteenth [Ell70]. Nintendo [Cov00]. Nisonger [RV98]. NJ [Art77, Dav79b, Dom78, Fis81a, Fis81b, Gro75b, Gro77, Har83, Har77a, Har78b, Kuc75, Mar80, Mar75, Mig80a, Rei77, Rus77, Shu80, Sta77b, Ste77, Str78, Wei79]. NLM [Rod95, TH85]. NLM. [LD88]. NMA [Kuc76a]. No [Ano74a, Ano77a, Ere81a, Har78a, Har76a, Har77b, Hei78, Kaz77, Lan81b, Lil78, Pre71b, Pre71c, Ric77, Shu78, SBW93, Swa76a, Swa78b, War79, Gul85, ENN98]. node [KE99]. noisy [TBCE94, WLC00]. nominations [Len83]. Non [Hjø00b, SK98, Boo74b, CCW84, McC95, Tra98, YAH90]. non-boolean [CCW84]. non-core [McC95]. Non-indexed [SK98]. Non-Relevance [Hjø00b]. non-standardized [YAH90]. non-traditional [Tra98]. non-unique [Boo74b]. nonauthor [DG91]. Nonbibliographic [WL80]. nonbook [Sop96, Sve94]. noninteractive [Swa89b, Swa89d]. Nonmaterialized [Lyn91b]. nonspecific [Kir93]. Nonstandard [RG70]. nonsyntactic [Fag89]. nontraditional [Hen88]. noon [Sha97b]. Noordhoff [Hei78, Swa78b]. NORDINFO [Buc91b]. normalization [BE91]. Norman [Gro77]. normative [MM87]. North [Kaz79, Koc79, Tau74, Bar96, WWB97]. North-Holland [Kaz79, Koc79]. Norton [Tsu79]. Norway [Ere80]. Norwood [Mig80a]. Nos [Wel75, Rei85].

Nostrand [Vag80]. not-for-profit [Pec85]. Notation [Mom82]. notations [Bra70]. Note [Ain85, Ano76g, Ano77b, Ano77h, Ano78a, Ano78b, Ano84o, Eli87, Gul87, Kil87, Mea78a, Mea79a, Mea80, Mea81, Mea82c, MKDL83, Mea83a, Mea84b, Mea84a, Ric88, Sal87b, Sic92, Whi87, Woo87b, Bur92b, Coo79, Egg90a, Flo99a, Hir89, Jac91, Kan78, MM88, Per83, RR71, Sal85, Swa88, Van99, Vea90, Wil84, Yag87, Zal74]. Notes [Ano74d, Ano82b, Buc75, Ken87b, Ken87c, Kra86a, Kra91c, Mea78b, Mea79b, Mea79d, Mea79c, Mea82d, RH99, Bor87, Kra87b]. Notice [Mea82d]. notices [Mea82d]. notion [Col94, Sar75]. notions [Fra86]. Nottingham [Les89]. noun [TC00a]. novel [Sch73a, Wis00]. novelty [McC89]. November [Har70c, Ric77, Wes00]. novice [HY93, Laz00, LBW00, Mea00, NT96]. novices [Mar89, Wie86]. NSF [Smi74a]. NSF-OSIS [Smi74a]. NTIS [Bra87]. Nu [Har75c]. Nuclear [Ger79]. Number [CC79, Han75, BK79, Har71d, Kuc78, SC80, TSV99, Whi91]. numeric [Dod79, Jär89, Was74]. numerical [BR95b]. nutshell [Lea98]. NY [Bai75, Che76, Kaz77, Mat85, Sta75a]. Nycum [New76].

O.R. [KL79]. Oak [WH70]. oasis [Pai90]. Obituary [Ano74m, Kin83]. Object [CD94, CC92, MHCW89, Zen99]. object-based [CC92]. Object-oriented [CD94]. objective [Avr71, Wal86a, Bro80a]. objectives [CM84, Per92, Sta75b]. objects [BSR93, Bro85, Col95, FHN^{+}93, MUD99]. observation [ENN98, Erc92, Spe71]. Observations [LaF00, ST86, Sla75]. observed [ER00c]. Observing [CB73]. Obshchestvennye [Rya80]. Obsolescence [Avr79, Bro70c, DS93, Gup90, AEP00, Bai74, ER00a, Koh74, Kuc82, Lin87b, San74, San83, Wal86d]. Occasional [Har78a, Sta85]. occupation [MBS98]. Occurrence [Yer71a, CMK^{+}98a, PW91, SSD99]. occurrences [Pao78]. occurring [Kir81]. OCLC [Bea80, DDC78, DCD81, Kil87]. OCR [SLZC92]. October [Hen96, Kol74, She72]. OECD [Ger79]. off [GK89]. Office [Art74, LC83, Pre71b, War79, Col83, Coo91b, Coo91c, CB83b, Gli74, Mic83, MSHB86, Ken87a, Mar83b]. offices [Don70]. oil [GL96]. OK [JGR^{+}95]. Old [Col83, Gar79, Hur80, OR78, OG80, MBH94]. Olle [Kra79]. On-line [AB76, Bou80, CR80, Cua71, Fay72a, Fay72b, Fay73, Koe79, LS80a, Mar00, Mor76, Nie76, RBR80, Bac72, BR95a, Coo77b, DH78, Elc78, Ell76, Haw77, Kat72, KM73, LWG74, McC80, MRW73, Mog75, PK00, Ras80, Ros79, SWK78, SR77, WL80, Wil77a, Wol71, Haw77, Slo00, Ano74b, Mig79, Tes79]. oncology [Sch88a]. One [Jak88, Lun76, LH76, CH73]. One-pass [Jak88]. ongoing [Law76]. Online [Bag81, Dos86, Fen81, Fid84, Fox86, Haw86, Haw87, Jon89, Lan80, Lan81a, LH86, MCBC89, Mil87, NC86, Oja86, Rot81, SBW93, Swa89b, TH85, Wil86, ZC92, AL84, Bat86, BB96a, Bel85, BF88, Bor86b, BHWG95, BHH96, Bor96, BMFS84, Boy86f, BM89, Boy90b, Bro98c, Car96, Cha88, Cha81, Con84, Con95, Coo83a, Coo83b, Cre95a, DS91, Dra89, DW96, Egg88, EH94, Erc98, FS83, GGJ91, Gir91, Har84, Haw81, HH95b, HH95c, Her96b, HJD00, Hoc89, IS98, IC97, Jen95b, Kan87, Kay98, KRC81, Kri86a, Lan93, Lar91, Lar92b, LMO^{+}96, Lit96, McG95, O’H95, Pat81, Ren80, Sch81, Sew82, ST86, Sie90, Spe80, SH94, SH96, Ten88, Vig83, VG93, Wal86b, Whi81, Whi00, Woo84, Yee91, Sie90, Boy86d, Boy86e, Boy87a, Boy89a, Boy89b, Hoc89, Lam00a]. Online [Lan80, Sha90, Sie89a, Spi94, Ten90, Ten93, Ten94, Sie86, Wit96a, Wit96b]. Online-Expert [ZC92]. only [Rus85]. Ontario [LAT79]. ontologies [Soe99]. OPAC [BB96a, Eft90, Slo00, Sol93]. open [CK96, Inh76]. Opening [Die82a]. operating [Boy86e]. operation [BB83a, Koe88, MR81a, MR81b]. operational [BCM^{+}70, BM75, Lev76b]. Operationalizing [Col94]. Operations [Lan77a, Col00a, Fay73, GR74, Kra73, LJ70b, Rou76a, Bom75, Lei73, Wri71b, Kan94]. operative [NJWC97]. operator [MZ78, Mar88]. operators [MB86]. Ophir [Wil00c]. Ophthalmology [Vir70]. Opinion [Art73, Bom75, Bre79, Bur85, Cre76, GS78, Har71a, HC84, Koc87, LR70, LMM^{+}73, LD85, Mar82a, Mea72, Min78, Old87, OD70, Pai71, Per92, Pri83, Ric70, Ros74, Sla75, TC73, Tal74, Teb70, Vin74, WK75, Wil77b, Woo87a, Koe83]. opportunity [Mar85]. Optical [Ros83, Ali90, Bro86a, Par83a]. Optimal [CC96, GV87, Gre90, GR74, MOV83, OV82, Rou74b, Rou75, Buc75, CC98, MB91, MN81, RV78, Ror99b, Sha91b, TSV99]. optimally [Smi81]. Optimization [SR79, HT91]. Optimizing [AF98, BCB98, LS70]. Optimum [Har71d]. Options [Bel88]. Oral [Don77]. ORBIT(R) [Mor76]. Order [CR00, EB88, PJ90]. Order-theoretical [CR00]. ordered [Sel71]. Ordering [Car96, Cod73, Los92]. orderliness [Ror99a]. Orderly [Glu95a]. orders [ER00b, San79b]. ordinal [Gre90]. Organization [Dil82, Kuc75, MA76, Swa76c, Wie86, Cas91, CCN93, Cro79, HW98, JR96, KDJ94, Lid96, LC83, Mad96a, Mad96b, NK83, Tra98, VA73, Wal81, War73, Bat93]. Organizational [CK96, EK97, Buc89, BT99, CB83b, COC90, Dan83, EF96, Kat00, Tra98, Und98a, WWSR00, Tay94]. organizations [KSDR90, Stu72, Cla75]. organized [NvR98]. organizing [Lar90, OCN97, TK96, Jen95a]. orientation [GSK98]. Oriented [FK73b, Asa81b, CD94, Gli72, HPH96, WSQ94, McL90b]. orientedness [Tan99]. Origin [Har78b, Pie92]. original [May78]. Origins [Kil97, Sha92, DK90, McI97, Ray97, Shi86a]. ORNL [JBH77]. Orrin [Die82a]. orthodoxies [Bea93]. Orthography [Bro98c]. Oryx [Bel81]. OSIS [Smi74a]. Oslo [Ere80]. Other [Jan94, Mor76, Mur74, Pre71b, Swa75b, Aul87, Bur92a, Cam96b, CK99, Coo70b, Cro99a, Gas98, Hei95b, Hen99, Hoc89, O’C73, Pec85, Tau74, Wes87, de 76]. Otlet [Day97, RL97, Day97, Ray94, Ray97, RL97]. ototoxic [LH76]. Ottawa [Inh76]. our [Boy91, Sem98a]. out-of-specialty [JJ75]. outcome [Bel85]. outcomes [BR97]. outer [Cre99a]. output [Che76, Coo77b, Gor79, Haw81, KS77, NKM77, OCN97]. outputs [LMSC82]. outreach [BD88]. outsiders [Cha96]. outsourcing [Key96]. Ovellete [DC76]. Overbey [DC76]. Overlap [DS79, Mau71, WFK72, WFK73, BR80b, Glu90, Har90, SK88b]. overlaps [McC89]. Overseas [Ste77]. Overview [LH84, AL96, BD96, Boo90a, Can99, CCM00, Cre99b, D’E97, Dou72, EL85, Has85, Hur00a, Lan89a, Lun83, LS84a, LH86, LL86, LC87, LS88, LC88, LB88, LR89, LB89, LC90, LW90, LD91, LL91, LH92, LS92b, LF93, LS93, LP94, LF94, LB94, LH95b, Man82b, Mar83b, McC90, Men95a, Men85, Rad89, RL99, SD97, SD83, Ste80, TL98, Tho81, LUH^{+}96]. overviews [GMPS00]. Ovshinsky [Caw74]. Own [TL74, WW73]. Ownership [Lan77b, Rob96a, Gro96]. owning [Sle94]. owns [Car95b]. Oxford [Die81, Die84a, Har70d, Ols75].

P [Art77, DC76, Har83, Har78a, Lan80, Lea76, New76, RL97, Sta77a, Wel75, TZ93, Dom78, Wel78]. PA [Dro77, Woo77b]. PAC [Lin70]. package [Lun76]. Packages [Mar80, How89b]. page [Bra77, BLPG99, Coo76, DG00, HG00, Jen91, Koe99]. pages [Bat99c, Don74, HG00, Kuc76a, New73, Var78, Sal72a, Sch73b]. paintbrush [Wal00b]. Painter [Pre71b]. pair [RW00]. Palaces [Woo87b]. paleontology [Ste85]. Pamela [Afs99]. panel [Gar90]. Pao [Boy95a]. Paper [Bre79, Cre76, Har78a, LR70, Mar82a, OD70, Pri83, Ric70, Teb70, Wil77b, Aiy88, Die82b, GS78, Lou91, Mel72, Min78, Ros74, SD96, Smi74a, Smi74b, Str74, Tal74, Vin74, WK75, Bom75, Sla75, TC73]. Papere [Har77b]. Paperless [Aiy88, Lan81a, Old87, Rot81, Tia89]. Papers [Ano86a, Ano86b, Ano86c, Ano00a, Art73, Bro99a, Bur85, Gul85, Har71a, Koc87, LMM^{+}73, LD85, Mea72, Old87, Old91, Pai71, Per92, Ric77, Sha83a, Wha83, Woo87a, Ano83h, Ano85h, Ano00c, Ano00b, Boy86f, Bro00f, Dro95, Gar79, Hai97, Hur80, Kra86b, MN83a, Mea78a, Mea80, Mon73, OR78, OG80, Pic96, SR95, Toc70]. paradigm [Koc84c, Par94, WV97]. paradox [Lie94]. parallel [Bue86a, CMK^{+}98a, CBM^{+}94, LFHG99, RF97]. parallelism [GCH99, Car76]. Parameter [Los88, RR78]. parameters [Mar72a, Nic87]. Parenthood [Lea76]. Paris [Ger79, Gro76a, Gro76b, Lan80, Wei76, Wel78]. Park [Har76a, Ric77]. parsing [MHCW89]. part [Boo90a, Boo90b, Che00c, Coo73b, Fay72b, Hoe72, ALD96, ABLS88, BD88, Bel88, BF91a, Che00b, CS91, Fay73, FW88, Gar88, Gir91, Har75a, Har75b, Hay88, Inh75a, Inh75b, Jon74, LAD97, LD88, Lyt88, SK97a, SK97b, SKvB92, SvB94, Sha91a, Sha91b]. Partial [BK98, ER00b, KB98a]. Particular [Kil97, McI97, Wei97b, Wei97c, Bla71, Wri73]. partitioning [CO84, CO85, SG80]. Partners [Wes00]. partnership [FW88]. parts [Fis81a, GGJ91]. pass [Jak88, YR77]. passage [DO81, O’C80]. passages [O’C73]. Passman [Har70d]. past [Ano78j, Cox88, Sal87b, Wes88, Smi99]. Patent [Gli74, Hun72, Gli78, Tau74, Wei96b]. patents [NO85]. pathology [DPP78]. paths [SFB^{+}00]. Patrick [Bai75, Ber84, Bro00d, Shu78, Swa75a]. Patron [Spe80, Coa86, FS84, Dra89]. patrons [Mar96, Tor89]. pattern [CPW88, DKP00, Hub98, Str82]. pattern-directed [DKP00]. Patterns [Art79, Bra87, Hal81, Ken87a, Lin87a, LC87, MH87, MB81b, Pal87, San75b, Spr87, WCL87, Wes87, Wol87, BB95, Bon90, Che72, Coo83b, DOO^{+}96, Koh74, Lar91, LLM91, McG84, MG72, Qiu93, SH89, Smi84a, TH85, Yer83]. Paul [Har70a, Tau74, Wri71a, Day97, Ray94, Ray97, RL97]. Pauline [Boy86f]. paving [FW99]. PBS [SSM93]. PC [Mim85]. peace [Hum70]. Peer [Abe90, CS90, Gar90, LC90, ANM78, KM78, Kra87c, LB83, MAN80, MS00, Shi88, Wel00]. Pekowsky [Cal00]. PEN [RCJS94]. [Ano72c, Dav74, FW88, HM70, Mus86, RNBV98, Sch99, Wol87]. 1973 [McM76]. 2- [Rei85]. 20 [Egg87]. 4A [Mim85]. abstract [Mal74]. cancellation [War81]. Charter [Sta75a]. document [CO87]. exploration [Mil00]. flow [YS93]. full-text [MSJM91]. II [Hol71, Var75b]. indexing [FG70]. Information [Wat98, MY85]. International [Ray97]. LD [Ano71a]. MDS [EP91, Eve93, Gle93]. Number [McM76]. of [Gli72]. optical [GP83]. planning [Bak89]. Reengineering [BR94, Har94, Lan94, LB94, Pet94, Sap94, WT94]. service [SC85a, SC85b]. Technological [Lyn98]. Winter [Rei85]. Penland [Bai75, Swa75a]. Pennsylvania [Woo77b]. People [Bro95b, Ber71, Jaf98, Jan94]. Peoples [Mur74]. perceived [Cul85, WA79]. Perception [Bro00b, Whi96, DS96, TC00b]. perceptions [Bro95b, Jan91, SPWH99, Wat98, WD97b, Kuh99]. Peregrinus [Art75b]. Performance [Ben80, BCG79, EQH00, FB81, Har77c, Kje74, Reg88, Bur99, Bur80, CC98, CBM^{+}94, Cra71a, CO94, Ell76, Glu95b, Key97, Koe83, Los96c, McC88, ML90, MWY95, MS00, NZ91, NN92, PK00, Ras90, RR77, Rum90, SB90, Sha91b, Soe94, SBW90, VHdB^{+}98, Whi94, Wil98a, WD97b, Yoo94, ZPB98, SCD98]. Performing [SM75, Key97]. Pergamon [Die84a, Don76, Har70d, Ols75]. perils [Boo78]. period [DOO^{+}96]. periodical [DS73, Han75, KDS89, Pop75, TO84, Wel94, Yer71a]. periodicals [AB76, Bro70c, Hoc74, MN81, San79b, War81]. Periodicity [McG96b]. periods [BF94]. peripheral [EN85]. permanence [AD79, Koe99]. Permuted [HH70, Cra82]. permuterm [Gar76]. Perreault [Wel70]. Perry [Ano72e]. persecution [Ler79]. Perseus [Yan97]. persistence [KB92]. persists [Bar78]. Person [SC85b, Sol97c]. Person-to-person [SC85b]. personal [Bar95, BS92, Bor99, Bur81b, Bur85, Cas85, Cox93, DG00, How94, McK00, OWF99, SKvB92, SvB94, SMC83, WdBFF95]. personalized [Kil87]. personnel [Kaz75b]. Perspective [But99, Cre99b, Cre99a, CH99, Day99a, Fra99, Gas99, Hjø00b, Kin98a, Lev99, LCB99, Lyn98, PPP98, RM99, Sch98b, Ste99b, Sun99, TL98, BSW00, DRM98, Hay88, Hei88, Kuh91, Mac82, NvR98, Ols99, Pee94, RL97, Rob96b, Rub92, Smi98c, Wyl87, Los96a, Los96b, Rya93]. Perspectives [AL96, AKS96, Abe96a, Abe90, ABB^{+}91a, ABB^{+}91b, Adl92, ABLS88, And91, And94, Ano81k, Ano81j, Ano84n, Bac93, BDL88, BD88, Bal89, BDZL89, BR94, Bar96, BR95a, BSM90, Bea91, BL85, Bea92, BS91b, BDBR97, Bel88, Ber94, Ber86b, Bes91, BD96, Bra87, BCB96, BF91a, BD97, BDSA97, BFM88, BC99, Buc91c, Chr95, CZ88, Cim95, Col83, CS91, CB94, CB89, Cox88, Cox92, CS90, CB83b, COC90, DRCG91, D’E97, Dan83, Dav88, Dav86, DG95, Dic86, Dos88, Dos86, DB94, Fag95, FW88, Fid94, Fox86, FHN^{+}93, FC92, Gar92, Gar90, GL93, Gar88, Gas98, Gau92, Gib86, Gol94, Goo97, GLS97, Gor95, Gre92, Har92a, Har94, Haw86, Hay88, Hay96b, HH95a, Her95, HOS^{+}93, Hur88, IT89, Jac81, Jac94, Jas94, JS95, KMG^{+}93, Ken87a, Ken87c, Kir93, Lan94]. Perspectives [Lay94, Leh92, LS97, Lin87a, LD88, LH97, Lun83, LC83, LS84b, LE85, LH86, LL86, LC87, LS88, LC88, LB88, LR89, LB89, LC90, LW90, LD91, LL91, LH92, LS92b, LF93, LS93, LP94, LF94, LB94, LH95b, Lyn91a, Lyn94, Lyt88, Mar92a, Mar83b, MH87, McC90, MWH^{+}95, Mea86, Men95a, Mes95, Mic83, Mil94, Mol92, MF88, Myl92, Nah98, NC86, Nug91, O’H95, OC88, Oja86, Osb98, Pai90, Pai93b, Pal87, Pee94, PW88, Pet94, Pre88, Rad89, RCZS99, Raw93, RRB91, Rie88, Rob96b, Rod95, RH97, Rom91, RL99, RTM99, RH99, Sap94, Sca89, Sch88a, Sch88b, SM88, SA97, She86, SBPP89, SMJ95, Sma99, Soe94, Spr87, SH94, SB91, Sve94, Tib94, Tom92, War88, Wat98, WT94, WC92, WCL87, Wel94, Wes88, Wes87, Wig95, Wil86, WMME88]. Perspectives [Wol87, Woo88b, ZvBR88, Zoe88, Bol92, FSTV99, Kue72, Lee96a, Mar95, TS96, WA97, Her94, Hei78, Bol92, Whi98d]. Pertinence [How94, NL83]. Peru [Har70a]. Peter [Art75b, Bel76, Rap00, Shi85, Ste00a, Woo77b, Vea90]. Petrocelli [Bel76, Che76, Dav77b, DC76, Sta75a, Str76]. Petrocelli-Charter [Che76, Dav77b]. Petrocelli/Charter [Sta75a]. Ph.D. [Col00b]. pharmaceuticai [FRGL75]. Pharmaceutical [ASS73, WRE^{+}70]. phenomena [Pao78]. phenomenological [Bro95b]. phenomenology [Lev76b]. phenomenon [Bas94, BR76, OV93, Stu78, Tag81, de 71]. Phi [Har75c]. Philadelphia [Dro77, Woo77b]. Philip [Kun82, Lam99, Che86a]. philosophical [Aik97, Hei80, Hei91, Ste80]. philosophy [Coo73b, Sch73a]. Phoenix [Bel81]. photocopies [Spe71]. Photocopying [Pom74]. photoduplication [Hir89]. Photographic [SB80, Mur93]. photographs [MM98]. photography [Buc92c]. phrase [Cra77, Fag89, Gal91, JGR90]. phrase-based [Gal91]. phrase-indexing [JGR90]. phrases [Egg99]. phrasing [TC00a]. Phyllis [Coc98]. physical [Bil00, Raf75, San75b, Sop96, Gup90]. physicians [Gor95]. physicists [Bro99c]. Physics [BB77, Ano70b, Bai74, Bai76, BK96, BCSZ79, BB83b, Che72, FB78, Gup90, MP74, San74]. physiology [Mar83c]. pictorial [SB80]. Pictures [JF87]. Pierre [Wri71a]. pilot [Oh98, Pra83]. Pioneer [Ric88]. pitfalls [Cua71]. Pitkin [Rus77]. Pittsburgh [Old91]. PL [Bas74, Dav76a, Dav74, Dav76a, HM70]. PL/1 [Dav76a]. PL/C [Dav76a]. PL/C-7 [Dav76a]. PL/I [Bas74]. PL/ [Dav74, HM70]. place [Ere81a]. Placebo [Kir93]. Plains [Don70, Mat85]. plan [Kuc76b, Sha96]. Planned [Lea76]. planner [LC83]. Planning [Har77b, KL79, Mic83, Sut96, Bro97, Lea76, LC71, LD88, Mac72, Rei75, Rou74a, Rou98, Wal86b, BD88, Cor94a]. plant [Mar83c]. plastics [Ano70a]. player [Har87]. plus [KM00a, Lan87a]. Plus^{T M} [GS93a]. poets [Whi99a]. point [BM89, EH94]. Pointing [Exn97]. points [Coo70b, SM90]. Poisson [Bur90, Eve93, Gle93, Sri90]. Poland [FMN82, Nit79]. Policies [Ros82a, Bol92, HP00, HW99, Ric89, RSK74, Whi89a, Wel75]. Policy [Bra79, Cas97, Ain84, AF98, Ber82, Bor85b, FW99, Hen99, HMNW86, Ker77, Kra91c, Kuc76b, MMB99, Men92, Nad85, Sha84, Spr87, Ste99a, War98a, Wil00b, Cas79, Spr87, Shu78]. political [New94, RCJS94]. politicized [Dos88]. politics [Cha85, Day00a, How89c, Mea78b, Pet98b, Rob96a]. Pollach [Har76b]. Pollock [Pre71a]. Pollution [HK79]. Polytechnic [Tau74]. Pool [Koc84b]. popular [Coo89]. popularity [Chr95]. Population [Lea76]. por [Har70a]. Portrayals [DH97b]. Position [Smi74a]. positive [Rou92b]. Possibilities [Nie76, Ric76]. Possibility [Tzv90]. Possible [Pri83, And78, EB88]. post [Afs99, BK98, Gue78, Hjø99, Jon71, LFS72, Mur93, Pra90]. post-coordinate [Gue78, Jon71]. post-coordination [BK98]. post-industrial [Afs99, Pra90]. post-modern [Hjø99]. post-photographic [Mur93]. postcards [Ste74b]. Postcoordinate [Sch82]. posterity [Lar99]. Postindustrial [Art82]. Postmodern [Bro99b, Bro96]. Postscript [Ano89, CO96]. Potential [Mcd82, Ali90, Bea80, Bra94, Coo70b, CB83b, DDC78, MUD99, ST89, Swa89d, Wee99, Wol71, Sch88b, Smi74a, Jen95b]. potentials [Mar83c]. pour [Gro76a]. power [Boh90, Buc89, Cas97, Sca89, Exn97]. pp [Ano74b, Ano75a, Art74, Art76a, Art77, Bai75, Bas74, Bel76, Bla79, Boy85b, Che76, Dav76a, Dav77b, Dav79b, DC76, Die81, Die82b, Don70, Don76, Dur76, Ere80, Fis79, Fis81a, Fis81b, Gan74, Ger79, Gro75a, Gro75b, Gro76a, Gro76b, Gro76c, Gro77, Han75, Har70c, Har70b, Har78a, Har75d, Har75c, Har76c, Har76b, Har76a, Har77a, Har77b, Har78b, Har77c, Hei78, Inh76, Kaz75b, Kaz75a, Kaz77, Kaz79, Kol75, Kra79, Kuc75, Kuc76b, Kun82, Lan78, Lan82a, Lan73, Lan76, Lan77b, Lan81b, Lar74, Lea75, Lea76, Lev76a, Lev76b, Lil78, Lun80, Mar80, Mar75, Mar77b, Mat85, McM76, Mig79, Mig80a, Mim85, New76, Nov82, Ols74, Ols75, Ols76, Pat85, Pre71b, Rei75, Rei77, Ric77, Rus77, Rus85, Sel71, Shi85, Shu80, Sta77a, Ste77, Str76, Str78]. pp [Swa75a, Swa76c, Swa76b, Swa76a, Swa77b, Swa77c, Swa78b, Tau74, Tsu79, Usd85, Vag80, Var75a, Var75b, War79, Wei79, Wei76, Wel71, Wel80b, Wel75, Wil85b, Woo77b, You75]. pp. [Har70a, Hin70, Pre71a, Pre71c, Pre70, Sie70, Til70, WH70, Wri71a, Bel81, Boy85a, Cer83, Cri81, Die84a, Don77, Gul85, Har83, Har70d, Har79, Mea83b, Pec85, Rya80, Sta77b]. practical [Cor94b, Kin98b, Myb99, DPD99, San98a]. Practice [Row84, Bix96a, Bix96b, BMFS84, Bur96, Cas78, Jen95b, Lyt88, MWH^{+}95, Pen70, Sta98, Wes00, Tes79, Lam00a, Mai99, Ten94, Whi00, Boy92, Rou89a, Spi94, Und96]. practices [Cov00, Kid90, MR94, Raf75, Wel75, Con93]. practitioner [Ben00, Rob96b, Rya93]. practitioners [Mar72c]. Pragmatic [Dav76b, DS84, Sla75]. pragmatical [Lev76b]. pragmaticism [NK89]. praise [Boh90]. Pratt [Car79, Egg87]. pre [BK98, IS98]. pre-coordination [BK98]. pre-online [IS98]. preach [CC99]. PRECIS [MB86, Ric76]. precision [GK89, LS99, MOV83, O’H70, OV82, SMB72, Su94, BG94]. precision-recall [O’H70]. Preclinical [Poy82]. predicament [Sta85]. predict [BS91a, Bur90]. predictability [And78]. Predicting [Cra71a, Mcg79, NK00, Ros71b]. predictions [RR78]. predictive [Lev77]. Predictors [Bon82, PK00]. Preface [Hei95a]. preference [Gou74, Yao95]. preferences [CK99, Dio94, ER00b, Par00]. prefix [Gre90]. prefix-condition [Gre90]. prefixes [Bla71]. prehistory [Kil87]. Preliminary [WD97b, Del90, GB95, Hum99, KB98a, Moo72, Tho72, LAD97, YFR81c]. premises [Ros74]. Prentice [Art76a, Dav79b, Dom78, Fis81a, Fis81b, Har83, Mar80, New76, Str78]. Prentice-Hall [Art76a, Dav79b, Dom78, Fis81a, Fis81b, Har83, Mar80, New76, Str78]. prepaid [Rus85, Sel71]. Preparation [Law76, Wol72, Bla71, CT73, Lun76, SJ72]. Prepared [Hin70, Rus85]. preprint [Ano70b]. preschool [BW85]. prescientific [Hei95c]. presearch [Han86]. presence [MBS98]. Present [Pat90, AdKK87, Cox88, Hei95c, Smi99]. presentation [EB88, KW97, PJ90]. presentations [Pap88]. presented [Old91, Ric77]. presenting [FCJ^{+}92]. Presentor [Pri72]. Preserving [Lar99]. Press [Ano75a, Ano77a, Art77, Bas74, Bel81, Ber84, Cri81, Die81, Die82a, Die84a, Don76, Don77, Dro77, Ere80, Ere81b, Gro76b, Gro77, Har70c, Har70b, Har75d, Har75c, Har77a, Har78b, Har77c, Hin70, Kuc75, Kun82, Lan77a, Lan73, Lar74, Lun80, Mar75, Mar77b, Ols75, Ols76, Pat85, Pre70, Rei85, Rei75, Rei77, Rou76a, Rus77, Sal72a, Shi85, Shu78, Shu80, Sta77b, Ste77, Swa77c, Wei79, Woo77b, You75, Koe75]. pressure [SS93]. prestige [SPWH99]. prevalence [Sch97a]. Previews [GMPS00]. Price [Bla79, Boy85b, Die81, Die82b, Dro78, Egg97, Ere80, Ere81b, Fis79, Har78a, Har78b, Hei78, Her82, Kaz79, Kra79, Kun82, Lan78, Lan82a, Lun80, Mar80, Mar85, Mat85, Mig79, Mim85, Nov82, Shi85, Shu78, Shu80, Str78, Tsu79, Vag80, Var78, War79, Wei79, Wel80b, Ere81a, Fis81b, Lil78, Mow92b, Mow92c, Pap98, Koc84c, KH95, Ste84]. Pricing [AL96, AKS96, Abe96a, BCB96, Cas79, Hay96b, Kin98b, Zai77, Zmu75]. Primary [Jas94, BR80b, EN85, GSKL00, Van97a]. Prime [CC79, SC80]. principle [Bal84, Chr95, Coo83c, ER95, GL91, GL92, KL98a]. Principles [Bar98b, Dav86, Gib86, LL86, Mea86, Tes79, Wal86b, Whi00, Art75b, Boy87a, Bro00d, Con93, ER91, Ten94, Zin00, MR94, Goo99, Bie95]. print [Afi98]. Printed [Kun82, LBP76]. Printer [Onu82]. Printing [Ken87a, Pom74, Kra91c]. prior [HP00]. privacy [Car95b, Lam99, Sal80, Woo88a, Pet98b]. Private [Bro85, Gar92, Hal85, Shu78, Wol87]. Probabilistic [Aji92, Avr71, BS74, Dav92b, Fuh00, Los87, Aji91, Cro81, Cro86, Egg93, Har75a, Har75b, KS90, Kwo85, LK91, Los88, Los97b, Rad82a, RC95, Tho90, WY90a, WY93, Mar82a]. probabilities [Gre90]. Probability [Ben00, GL91, GL92]. Problem [Mcg79, Buc92c, Bue85, CSI87, CNMS97, Col99, Cor82, Gir91, Goe87, GV87, HK80, Jon71, Jon86, Ker74, PR98, SG89, SH96, WdBFF95, Xia92, Zha98].

Problems [Gol73, Har77c, MM89, Pac82, Bla71, GGJ91, HF00, Her95, MHCW89, New73, Rob00, SCD98, Sch97b, Wil95a, WT93, Per92, Wil75, Cer83, Swa76b, Sch73b]. procedural [LBBR96]. procedure [DKD71, Sav99, Whi89a, Kra74b]. procedures [CC75]. Proceedings [Bel81, Boy91, Bro95a, Gul85, Joh86, Les89, Smi98c, Ten88, Buc91b, Har76b, Pap93, BI98, Inh76, Swa75b, Swa76b, Swa77b]. process [BBSM87, Boo83, BHH96, Dun96b, Gro78, Haa96, IS98, Kra87c, Kuh91, Kuh99, LS80a, Leh92, Rob91, SCD98, Sta85, WC92, Wor73, CB73, Swa76d, Cur97, Har87, Van97a]. processed [Whi76]. processes [Dum87, Hol86, Lev76b, Mon81, Nic89, Tag81, Yan97, de 76, Tay82]. Processing [Swa77b, Ano81j, BSR94, Boy90a, Dos86, GBDH76, Gri84, Les89, MHCW89, Mic94, Mus86, RSW91, Sag75, Sal70, Sal72c, SW91, SLC98, Sim81, Sol74, Sop96, SLZC92, Swa75b, Swa76b, Tzv90, VS87, Wei88, Zvi86, Art76a]. processor [AB76, CPW88]. Produced [Don70, Cra00]. producing [Ger78]. product [BHH96, Cur97, DH97b, Fal98, LH76, Sch88a]. Production [Swa76d, DW98, Egg96, Gro78, Lev82, Lun76, RV96, TSK71, RSZ71]. productive [FG70, McC95]. Productivity [Mon73, Bam79, BGM85, Bur93, Coi77, Col92, CB83b, Eas89, KB98b, Lin87b, OV92, Rao80, Sal86, ST89, Sub84, Wal86d, Cas96]. products [Kin82, Lan73, Lit96]. Profession [Shu80, Swa76c, Cox92, How89c, KM95, Mar92b, Swa78c]. Professional [ABLS88, BDL88, Bel88, Cox88, Dos88, ENN98, Gar88, Hay88, Hur88, LC88, Lyt88, Mcd82, MF88, OC88, Pre88, Sel71, SM88, Woo88b, AL98, Ber71, Cro99b, Day00b, FS97b, Ols89, Sch94, SvB94, SMC83, Var99, Wal98, Hei86b]. professionals [Bel88, KBN90, KRD95, Pat90, RH97, Mat85, Hay95, Sny98, Wei96b, Whi89b]. professions [BDL88, BDZL89, Hei87, Mat97a]. Professor [Che86a, Kin83]. proficiency [WdBFF95]. profile [Lun76, RL76, SP86, SBW93, YdKKM87]. profile-query [SP86]. profiles [CT73, SF78]. profit [Pec85]. Profiting [Cul88]. program [BCM^{+}70, BM75, BMFS84, CO96, Fis81a, Flo00, Goy87, LM78, Wes88, Lap99]. Programmatic [Wes88]. Programme [Sha83a]. Programming [Dav79b, AF98, Bas74, Dav76a, Mar95, Smi81, Str78, Wie86, WLC00, Dav74]. Programs [Don96, KL79, Bor73b, Dav92a, Kaz75b, KK96, MH87, MF88, NK83, Pal87, SvB94, SBW90, Whi96, Woo88b]. Progress [MBFK76, TH82, Wil75, BWKG75, Sal80, Fre77, FF80, Har88, Mig80a, Wil95a]. Project [Pre71b, SBW93, BHWG95, BF91a, BDSA97, Hun72, Lyn90, Mar71, MSJM91, RCJS94, RH97, SA97, WS98, WW99, WSB98, Zen99, HM87, Rod95, Ere81b]. projection [Liu97a]. Projections [LS80a]. projects [Blu71, Ens95, Law76, Wes87]. promise [LM78, LS84b, Cua71]. promises [Cer83, Her95]. Promising [Jer70]. promote [OO97, Zoe88]. Promotion [SCLL73]. proofs [Egg93]. Proper [Bat00]. properties [BKR98, Goy87, Ror99b, YK93]. Property [Gol73, Bra89, Rou92a, Wei96b]. proportion [Flo99b]. Proposal [KDJ94, RT81, RV98, Teb70]. proposed [BP70, Don79, Har90, HN97, Yu76]. prose [CELY80]. prospect [Ano84n]. Prospecte [Cer83]. Prospects [Ber86b, Car76, Dav86, Dic86, Gib86, LL86, Mea86, She86, Pat90]. Protecting [Woo88a]. Protection [Han75, Pom74, Wei96b, Gro96, Han75]. Protein [Wis00]. protest [Ler79]. proto [Cro99a]. proto-indicators [Cro99a]. prototype [YS93]. providers [GP83, Wol87]. Providing [Bus95, Hoc89, O’C70]. Provo [Don77, Har78a]. proximity [Wee99]. psycholinguistics [Wan81]. Psychological [Die82b, Har92b, PN79, SMC83, Sma80]. Psychology [Gli72, BP70, BM81, Col92, Con84, Dor85, Lev76a, Sma80, Vig83]. Publ [Var75b]. Public [DS79, FS97b, BB96a, Bro85, Car95b, Cas94, Cox93, DH70, GO93, Hal85, Her96b, KS77, OG93, San98b, Sav97a, Sch96b, Slo00, Spa97, Wal00a, WWSR00, Yee91, Gar92, Wol87, Haw90, Shu78]. Public/ [Wol87]. Publication [ANM78, BB83b, Har77b, Lee00b, Swa76a, BB95, DOO^{+}96, Egg94, ER00c, Gan74, HP00, KN00, Kuc80, Lan81a, MZ80, NC75, Raf75, Rot81, San75b, ST89, Tau74]. Publications [Bol86b, Har76b, Kaz75a, Mat85, TL74, Ano74a, Bol86a, Col92, ERV00, IC97, LB83, MZ79, Opp98, Oro77, PBvR95, Van97b, WW73, Wel75, Gan74]. publicly [Hoc89]. publish [CC99]. Published [Dro78, Har75c, EK75, Gol97, Gou74, Ste85, Win70a]. Publisher [Ano77h, RMMS96, Ano84o]. Publishers [Dav76a, Hei78, Tsu79, Gol94, PPP98]. Publishing [Ber94, Boy85b, CB94, DB94, Ere81a, Gas98, Gol94, Gol74b, Han75, Jac94, Jas94, Kin98a, Lea75, LP94, Lyn94, Mig80a, Pee94, PPP98, Raw93, SH94, Swa78b, Bro97, Day99a, Jas97, Ken87a, KM99, KH95, Lan87a, Lan89a, MR94, Pee96b, Pra90, Pri83, Sch94, SRS98, Sta85, Afi99, Alg95, Cox98]. Pullach [Kuc76b]. purpose [HM70, MCD73, Bor73b]. purposes [NML99].

Qualitative [Yan97, Kim00, LS97, War91]. Quality [LJ70b, NZ91, AD79, Buc91b, Cha76, Cit98, KM78, Los90, LP99, Los00, Mat00, NJWC97, TF89, Van98b]. Quantitative [War91, MB86, YdKKM87, Bur93, OV92]. quantities [Sko98]. Quantity [Ols74]. Quarterly [Ano74a, Lil78, Wel75, Han75]. queries [BSR93, BC99, Cro79, Cro86, FWR95, Gor90, Los97b, ME98, Tag72, Yag87, YS93, Yu74, ALD96]. Query [NE90, WY90b, ASAOE99, AB76, Ang75, BCB98, BSR98, CSSI98, Coh99, Eft00, FS91a, FBK97, GS93b, Jär89, Los96c, LP99, MB91, PW91, SBF83, SSM93, SLC98, SP86, TSK71, Wei96a, vdVM99, Zvi86]. Question [Hom83, Gol74a, Ham00, Jah74, SPWH99, Vei80]. Questionable [Gle93]. questioning [MB91]. Questions [Dil82, HG00, JR96, Roy90a, SK88a, Sve86]. Queueing [SR79]. queuing [Smi81]. quick [Lea98]. Quotations [KF92, Haa93].

R [Bai75, Bor99, Boy85a, Caw74, DC76, Fis81a, Har75d, Kaz79, Ras99, Rei75, Sel71, Swa75a, Swa76a, Whi00, CT87]. R&D [Bax85, McC95]. race [New94]. RaD [KW77]. Radiation [VC70]. radically [Lev76b]. raised [Gro96]. Raising [Nic00]. Ralph [Hol87]. Ramahgam [Sta77a]. Ramist [Rei75]. Random [Dro78, Spe71, Whi87]. randomness [Boo97]. range [KPZ^{+}99]. rank [BG78, Bur94, NT85b]. rank-frequency [Bur94]. ranked [Gri88, NKM77, San79b]. Ranking [RB89, Sav97b, CR00, DM99, Gor82, GL91, GL92, HC90, MB81a, MS00, Ro88a, Ro88b, LP99]. rankings [CO96, Los00, Nis99]. Rapid [Bur92a, Coh99, JBS^{+}77]. Rasch [AP96b, AP96a]. rate [NL83]. rates [Coa86, DCD81, GdCD81, Rei85]. Rating [WW00]. ratings [ANM78]. Rational [BE91, CK96]. rationale [Bra77, Dil97a, Zin00]. Rationalization [BR80b]. Ravindran [Buc75]. raw [MB81a]. raw-use [MB81a]. Raymond [Dav79b]. Raynaud [GL96]. RCA [OG70]. re [Ano08, Bod99, GL96, Vea90]. re-examination [GL96]. re-unification [Bod99]. reaching [Var99]. reacting [Ler79]. reactions [OOA99, LB00]. Read [Ano78j]. readability [DK75, Kin76, May78]. Readable [Til70, Haw81, Kri86b]. Reader [Cla75, Don74, Har76c]. Readers [HG00]. readership [BC96]. Reading [Gro96, Har79, BTSKC96, Chr72, Haw86, McK96a, McK96b, Mil87, Rad89, Sum84, WW99]. Readings [Ano83i, Var75b, Ano83i, Pet99, Whi98b]. Realigning [Sap94]. reality [JB85, Nug91, Hen95]. really [BF93, Sic85]. reason [Gol97]. reasoning [Kas96a, Kas96b, Und98b, Liu97a]. reasons [CH00, CB89, OR78]. Reassessing [Har94]. reassessment [Ray99]. Recall [All89, GK89, Lev98, Mar78, MOV83, O’H70, OV82, Su94, BG94]. Recall-precision [GK89]. recallable [KM00a]. recognition [DF84, Mar71, Sho71, VS87, WH98]. recollections [Flo00]. recommendations [Har70c]. RECON [Til70]. reconfigured [Mur93]. reconsidered [Fro94]. reconstruction [vdBL96]. record [CS90, Her95, San99, Wol96]. Recordkeeping [HW99]. Records [Til70, Boy86a, Car96, Cre95b, FS84, HC90, JJ75, VC70]. recovery [RF00]. Red [KL79, Pom74]. Redeploying [Pet94]. redescribing [Gor91]. Redesign [BR94, Har94, Lan94, LB94, Pet94, Sap94, WT94, Boy86f]. Redesign/Reengineering [BR94, Har94, Lan94, LB94, Pet94, Sap94, WT94]. Redesigning [Sha78]. Reducing [Kra89b, Kra90, KJR99]. Reduction [WC75, PT87]. redundancy [BDP98, CVK92]. redundant [YW96]. redux [Bla96]. reeducating [Sap94]. Reexamining [Dro95]. refereeing [Mea80, Mea84b, SM78]. referees [Cam96b, Cam96a, Glo88, Kra88b, Kra89a, Kra92b, Mea79d]. Reference [Gol74a, Jah74, Lan76, Swa77b, Bac72, Ber86a, Bla71, BR80a, BCB96, Bud91, Egg97, EW70, Gro90, Hay96b, Her78, Hoc89, Kat72, Mer71, Pat85, Rie88, Rob80b, Roy90a, Smi96a, Smi96b, SHS72, Ten90, Wil95a, Wri73, Lea98, Wil98c]. referenced [BK96, KF92]. references [Dod79, KH71, Mou77, Yoo94]. referencing [Din87, Kid90, MDS70, SK99]. referral [Don79, Hoc89]. reflected [How94, Mar92b, Wal97]. reflection [Mea94]. Reflections [BCB96, Dun00, Ber82, LD85, Mac79a]. reflective [Bur96, DB94]. reformation [GS93b]. reformulation [Dal90, FBK97]. Regarding [Bra92, Dav92b, Hir90, Lon92, Wei90, Wit87]. Regional [KL79, GSK98, Spe71]. Registry [Wei97c, BR95a]. Regulatory [Ber86b, Haz98]. reinforcement [Rou92b]. Reinhold [Vag80]. reinterpretation [Lyn77]. reinvention [Soe99]. rejected [Cam96a]. Rejoinder [Bor97, Buc91a, Buc92a, Deb92, Eve98, Fid87, HN98a, Hjø96, Hol87, Kra92c, Lou92b, LD97, Mac97, Mea97, Old88, Ric97, Sal87a, TZ93, Van98a, WM00b, Laz00]. Rejoiner [HN98b]. Related [Swa77b, Bro90b, Ere85, Kir81, Mas97, McC00, MN83b, MF88, Moo72, Swa76b, Swa89b, Swa89a, Swa89d, ZPB98]. Relatedness [Bon82]. relating [ABB^{+}91b]. Relation [BK79, Kuc78, Pre71b, Egg87, Egg93, Egg97, Lip83]. Relational [WVE85, Bro88, CVK92, Cra81a, GFHR97, LFHG99, Lyn91b, Mac91, MW91, PT87, Tes86, WEWP97]. Relations [Rou90, AM70, DH97b, ER00a, FNNW74, HDRD88, Lyn91b, PBvR95, Pic96, Rap00, Tag72]. Relationship [Lin87b, Bel85, BG94, CL86, DC98, Egg94, Hub78, LBBR95, Rub92, Seg94, SP86, Sma73, SL99, SR77, Wal86d, YdKKM87]. relationships [CP98, Coo77b, Gre95, GB95, HNW93, KS98, She71, SL99, WY93, YR77, BLPG99]. Relative [Byr75, BB83b, Egg88, Los00, Van98b]. relativity [Bar78]. Relevance [Boo79, Fro94, Her94, Hjø00b, Mea85, Miz97, RJ76, Sar75, Bar94, Bar98a, BCB98, Bru94, CSSI98, Dav77a, Eft00, EB88, Glu95b, Gre95, GB95, Har92b, Har96, Hir99, Jan91, Jan94, JF87, KS90, Lev98, LFHG99, MKB78, Mea98, Ott94, Par94, Ror99a, SB90, Su94, Sut94, TSV99, TC00b, RC74]. relevant [Wil95b]. reliability [Lyn88, Nic00, SWT^{+}79]. Reliable [Dav77b]. remarks [Hum70, de 71]. Remediation [Day99b]. remembrance [Bor99]. Remembrances [GBL89]. remote [Wol71]. remove [YW96]. renewal [Coa86]. renewals [Coa86]. Repatriation [Gro96]. repertory [LS92a]. Repetitive [Bro88]. replacement [Fly97]. replication [DMZL82, GL96, KM73, Nic00]. Reply [Bor74, Ley98, Qin00a, Bro00e, Cer79, Coo76, Kra85, Lan81a, MZ80, Mea84c, Ozo80, de 78, Art75a, Rob79, WC85]. Report [BWV72, Ger79, MBFK76, Pre71b, SBW93, TH82, Var75a, War79, Woo77b, AB76, BWKG75, Fre77, FF80, Mar71, Pen70, Ric70, Rou98, Rus85, Sal80, Tho72, Tyn75, Zen99, DG91, Dou96]. Reporters [Pic86]. reports [Kuc76a, Caw78]. Representation [FNNW74, Whi76, Bat93, Che98a, Cro81, Jen91, McC86, O’C93, Oh98, TSK71, YS93, Zen99, Zha98]. Representational [PT87]. Representations [Mon81, Bar98a, Kwo99, LLM91, Pai93a, RC95, Sha91a, Sha91b]. representative [MUD99]. Representing [BCB95, KS98]. reprint [Onu82]. Reproduction [Wal97]. request [Don70, Rad82a, Rad82b]. requests [Boo80, Onu82]. Requirements [Arm94, CC92, DRCG91]. Resarch [Kuh99]. Research [ALD96, ASAOE99, Aga99, AYS96, AL94, Aji91, AF98, AKE94, Ali90, All89, AdKK87, AP96b, AP96a, Anv91, AL98, AL84, AC93, BYN00, Bal00, BR97, BS88, BTSKC96, BS91a, Bar95, Bar74, Bar94, BCB95, BCB98, BM95, Bas92, Bat86, Bat98, Bat99c, Bat83, Bax85, Beg95, BT84, BMSM90, BMNM^{+}93, BN76, BBFM99, BBDH00, BC00, BB95, Bil00, BBMS87, BBSM87, BF88, Bod99, BSR93, BSR98, BE91, Bon90, Boo90a, Boo90b, Boo96, Boo97, BKR98, BP93, Bor86b, BCM89, BR92, BS92, BHWG95, BM89, Boy95d, Boy95e, Boy95f, Boy95g, Boy95h, BMvR91a, BMvR91b, Bra89, Bre91, Bri91, Bro87, Bro86b, Bro88, Bro90b, Bro93b, Bro95b, Bro98c, Bro99b, Bro00e, Bro99c, Bru94, Bru98, BSW00, Buc89, BF91b, Buc91d, Buc92b, Buc92c, BG94, Buc99, BLPG99, BC97]. Research [Bur00, BD92, Bur95, Bur99, Bur92a, BNP99, Bur90, BF93, BF94, BR95b, Cam96b, Cam96a, CO84, CO87, Can89, Can99, CK99, CR00, CPW88, Cas91, Cas94, CH00, CSI87, Cha86, Cha91, Cha96, Cha99b, CT99, Che86a, CL86, CL89, CVK92, CN95, CYFS95, CCT95, Che95a, CNMS97, CSSI98, Che00a, CYW99, CPR84, CMB^{+}95, Coa86, Coh95, Coh99, Col93, Col94, Col99, Col92, CCN93, Col95, CC96, CC98, Con95, Coo89, Coo91b, Coo91c, Coo98, CC92, Cor99, CMK98b, CBM^{+}94, CNO^{+}00, Cra91, Cro86, CT87, CMRWW93, CO94, Dal90, DG87, DHS00, DL00, DPD99, Day00a, DDF^{+}90, Del90, DRM98, DGS86, Des97, DH97b, Dil00, DW95, DG91, DS93, Dio94, Doc91, Dom00, DOO^{+}96, Dor88, Dro95, Eas89, Eft90, Egg86, Egg87, Egg90b, Egg90a, ER91, Egg93, Egg94, ER95]. Research [Egg96, Egg97, ER98, Egg99, ER00c, ERV00, Ehi90, EB88, EH94, Ell86, ENN98, Ens95, EF96, Erc92, Erc98, Erc99, EN98, EP91, EC92, EP97, Fag89, FCJ^{+}92, FWR95, Fid84, Fid86, Fid91a, Fid91b, FDD^{+}99, FM86, FW99, FB99, Flo00, For91, For00, FW90, FFKS99, FK99, FB88, FS91a, FS91b, FSV93, FSTV99, FBK97, Fri97, Fro94, FS84, Fuh00, GCH99, Gal91, Gat92, GS93b, GH96, GS87, GSKL00, GS93c, GGJ91, Gir91, GO93, Glo88, Glu90, Glu95b, Goe87, GV87, GLL90, GK89, Gor90, Gor91, GL91, GL92, GL96, GD98, GM86, Gre95, GB95, Gre90, GWDS86, Gri88, GLW86, Gro75b, Gro96, GFHR97, Gup90, Haa93, Haa97, HG00, Ham00, Han86, Hap99, HC90, Har91, HDRD88, Har90, HH92, Har92b, HNW93, HC96, HP00, HG98]. Research [He98, HW99, Hei91, Hei95b, Hei95a, Hei95c, HT91, Hen99, Hen00b, HMNW86, Her99, Her94, HH95c, HBG96, Hig99, HJD^{+}99, HCL^{+}00, HJ89, Hir99, HO89, HA95, Hjø00b, HKE97, Hol86, Hoo91, How94, HY93, Hub98, HM87, Hum99, IS98, Ing97, IC97, JJ75, Jak88, Jan94, Jär89, JN93, Jen91, Jr88, JF87, Jon89, JGR90, Jon91, JGR^{+}95, Jul99, KB98b, KL98a, KJR99, KDS89, KSDR90, KL98b, Kha91, Kid90, Kil95b, KMB99, KM00a, KE99, Kim00, Kin92, KdK90, KM99, KM95, Klo89, KB92, Koe99, Koe92c, KS97, KPZ^{+}99, KS90, Kos94, KET99, KT90, KRD95, Kuh91, Kwo99, Lan77a, LG96, LBBR94, LBBR95, LBBR96, Lar99, Lar91, Lar92b, Lar92c, LS92a, LBIP97, LBW00, LK91, LNL99, Lee00b, LS99, Lei73, LK97, Lev76b, Ley97, LBKO87, LAD97, Lif93]. Research [LLM91, Lin97, LG99, Lip99, LY91, Liu96, Liu97c, Liu97b, LWZ99, LDS^{+}00, LS87, Lon90, Los87, Los88, Los90, Los92, LH95a, Los97b, Los97a, LP99, LFHG99, Lyn88, Lyn91b, MM89, MR87, Mac91, MB86, Mar89, MBVL99, Mar92b, Mat00, May97, McC84, McC86, McC89, McC91, McC95, ML90, MMB99, McD99, McG96b, MSJM91, MW91, ME98, Mea82b, Mea82c, MCBC89, Mea95, MWY95, Mec99, MB91, MS00, Men92, MHCW89, Mic94, MR94, MV95, MUD99, MB95, Mol95, Mor94, Mow92a, Mow92b, Mow92c, Mud98, MSHB86, Mus86, NE90, NZ91, NN92, Neu95, NJWC97, Nic89, Nis99, NML99, OOA99, OWF99, ODVG87, Oh98, OO97, OV92, OV98, OCN97, Ott94, OG93, Pai96, Pai93a, Pal99b, PK00, Pao86b, PW89, Par94, PJ90, PW91, PR98, PZSD96, Per94, PvR94]. Research [PBvR95, Phi92, Pie92, Pie99, PN98, Pou89, PT87, PB99, QLA97, Qin00b, Qiu93, Qua96, RB89, RW86, RC95, Ras90, Ray94, Ray99, Reg88, Rep89, Ro88a, Ro88b, RFK97, Rob00, Ror90, Ror93, Ror99b, Ror99a, RW00, Rou88, Rou90, Rou92a, Rou92b, RV96, Rub92, RSW91, RF97, SW92, SB90, SSM93, SW91, SKCT88, SK88a, SK88b, SK97a, SK97b, Sav93, Sav97b, Sav99, SCD98, Sch94, ST89, SKvB92, SvB94, Sch99, Sch90, Seg92, Seg94, ST86, Sha92, Sha91a, Sha91b, SD96, SP86, SFB^{+}00, Shn86, SMHC93, Sic92, SBW93, SH89, Sie90, SA91, SS93, Sma86, Sme87, Sme92, SL99, Smi99, Sol93, Sol97a, Sol97b, Sol97c, Spa97, SPWH99, Spi96, SS97, Spi97a, Spi97b, SRS98, Spr99, Sri90, Sta98, Ste90, Su94, Sub84, SBW90, SS89]. Research [SLZC92, SSD99, Sut94, Sut99, Sve86, SM90, Swa87, SH96, SK99, TBCE94, TSV99, Tan96, Tan99, TZ92, Tho82, Tho90, Tia89, TvR90, TEC99, Tod99, TF89, TK96, Tra98, TB93, Tsu79, Van98b, Var78, Var99, VTS97, VS99, VS87, Voi97, Voo99, WD97a, Wal86a, Wal86d, WEWP97, Wan92, WS98, WW99, WWSR00, War91, War93, War99b, Wat89, WSQ94, WSB98, Wat99, WW00, Wee92, Wee99, Wes91, WH91, WM98, Wie86, Wil92a, Wil93a, Wil92b, WdBFF95, Wil00a, Wil82, Wil93b, Wil95b, Wis99, Wis00, Wol86, Wol96, WY90a, WY90b, WYSB91, WY92, WY93, WNAR99, WT93, WT95, Xia92, YR98, Yan90, YW96, YAH90, Yao95, Yee91, YG90, Yoo94, YS93, YdKKM87, YK93, Yua97, YM99, ZC92, Zen99, ZK99a, ZK99b, Zha00, ZC00, dSM93, dR90, PW92, vRL96, vdBL96]. Research [vdVM99, Ano00b, Bat93, BB96a, Bom75, Bro95a, BB83b, Che72, Cov00, Cra71b, Cra72, CS90, COC90, Day99a, DS91, DCFQ00, Dre99, Dun96a, EK97, Ell96, FS83, Fis99, FB78, Gol87, GR74, Hal80, Har87, Har70c, Har78a, Hoc74, HSP75, Jon76, Kan94, KWR98, KS70, Koc84a, Koe83, Kra73, LWG74, Law76, Mat97a, MN83a, McC00, Mil90, Mil94, Nic00, NvR98, Pap96, Pep74, RB83, Rou76a, Row84, SC85a, SC85b, Sal85, Sha84, Sla75, SS00, Ste85, SL83, Stu72, Sum84, Tho71, Und98a, VHdB^{+}98, WRE^{+}70, WC77, WSS79, Wes00, Woo84, Can99, FMN82, Har79, Lyt88, Wil82, Whi98c, You75, Afs98]. research-related [McC00]. research/service [SC85a, SC85b]. researchers [BCM89, Keg70, Teb70, Sie70]. resemblance [PBvR95]. Resented [Har77b]. reserves [Cre99a]. Resilience [Boo90b]. resistance [Qin00b]. Resnikoff [Har70b]. resolution [Bon89, Coh99]. Resource [SR79, Ano80i, BS91b, Bra87, Gul85, Hei94, LB00, Lyt88, MWH^{+}95, Rei84, Rou75, Ste80, Rei85]. resource-sharing [Rei84]. Resources [Don96, Lan73, Lar74, Lie94, Rus85, Sal72a, Abe96a, Ano08, Boo74a, Buc79, Dan83, Hay96b, Hir99, Hoe72, Kuc75, Lev82, Pet94, Smi81, Sut99, Wil00a, Wri73, Zmu75, Spr87, Cul88, Jac96]. respect [Egg88, Pac82, Yu76]. Response [Coo83a, Gar90, Her96a, Swa89c, Che95b, DW95, FS84, Koe94, Kra92a, McC97, Key96]. responsibility [Cre76, Ora00]. responsibly [Bat97]. Responsive [War79]. rests [Woo76b]. result [Bur92b]. resulting [Whi76]. results [DCFQ00, O’C70, Sar71, Swa89d, Voi97, YFR81b, BC97, Neu95, Kaz75a]. retention [Cre95b]. Rethinking [BHH96, Sch96a]. Retrieval [ASAOE99, Ano74b, Bai75, BC99, Byr76, CZ88, Dal90, Dat79, Dav88, Her82, KMB99, Kim82, Kra74a, Lan98a, Lan98b, LS88, Mar72a, Mar82a, Mig79, PW89, RCZS99, Ras80, Rie88, RL99, RTM99, RH99, Sch88a, Sch88b, Sie70, Sma99, Sta77b, Tes79, Wel71, Wil92a, Wil93a, Wil82, Zoe88, AL94, AM77, All89, Ami70, Ano00a, Ano00c, Art71, AL84, Aus93, Bac72, BYN00, BSY00, Bal88, BJ79, BMNM^{+}93, Ben80, BW84, BE80, BBMS87, BBSM87, Bla80, Bla93, Bod99, BSR98, Boo78, Boo81, Boo83, Bor86b, BHH96, BMFS84, Boy87a, Boy89b, Boy90a, BM80, Bro93a, Bro94, Bro95a, Bro98c, Buc91c, Buc92b, Buc92c, BK81, Bue85, Bur95, Bur99, Bus95, Byr75, Cag70c, Cas91, Che95a, CNMS97, CWBL00, Cle76, Col81, Col99, Col00a, CM00, Col95, CC96, CC98, Coo70a]. retrieval [Coo72, Coo73a, Coo73b, Coo83c, Coo98, CDV98, Cor99, Cra81a, Cro81, Cro86, CT87, DG87, DO81, DW95, DH78, DCFQ00, Dom00, Dos86, Dro98, Dun96a, Dun00, ER97, ER98, EQH00, Ell76, EFW96, Ell96, EC92, Fag89, Fay72b, FS83, For91, FFKS99, FK99, FB88, FS91b, FSV93, FNNW74, Fuh00, Gar83, Gau92, GS78, Gor90, GL91, GLW86, Har90, Har96, Haw77, Haw86, Haw87, Hei74, Her78, Her94, HH95a, HPH96, HKE97, Hoc89, Hol70, Hum96, JR96, JN93, Jen95b, JB71a, Jon71, Jon91, KL98a, Kar98, Kat72, KM73, KZC86, KL98b, KL98c, Kin92, KC99, Koc83, Koe75, Kra73, KB78, Kra78, Kwo99, Lam00a, Lan71, LG96, Lar92b, LK91, Lev73a, Lev74, Lin97, Liu97a, Liu98, Los87, Los88, Los92, Los96c, Los97b, Los00, Low71, LG75, Luk96, LFHG99]. retrieval [Lyn91b, MR87, Mac91, MBH94, MR81a, MR81b, Mar77a, MBVL99, McC89, ML90, McG76, ME98, Mea83c, MWY95, MHCW89, Mic94, Mil71, Mil87, MPW73, MRW73, MCD73, Miy98, Mog75, Mud98, MSHB86, Myb99, NE90, O’C70, O’C73, O’C80, Oh98, Pap84, PW91, PZSD96, Pul74, Qiu93, Rad82a, RW86, RC95, Reg88, Ro88a, Ro88b, Rob75, Rob00, Rod95, RF00, Ros83, Sac76, Sal72b, SBF83, SFV85, Sal87b, SB90, Sar71, SKB99, Sch82, Sch73a, Sew82, Sha91b, Sie96, Sme87, Sme92, Soe94, Sol93, Spi94, SS97, Spi97b, SNG00, SS89, Swa88, TBCE94, TS96, Ten94, Tho71, Tho90, WEWP97, WVE85, War73, Was74, Wat89, Wat99, WRE^{+}70, Wei94b, Wei96a, Wel80c, Whi81, Wil98a, Wil71, Wil00c, Win70b, Wis00, Wol71, Wol96, WD97b, WY90b, Woo84, WT93]. retrieval [WT95, Xie00, Yag87, YR98, Yao95, Yer73, Yer83, Yoo94, Yua97, YM99, ZK99b, ZC00, dSM93, vRL96, AKE94, Bla79, BP93, Kei70, Mac79b, McC89, Cri81, Sal72a, Boy87b, O’C93, Roy89, Ten90, Tur98]. retrieved [SR77]. Retrieving [HC90, SB83, SKCT88, SK88a, SK88b]. Retrospect [Ano84n]. Retrospective [Til70, Elc78, FRGL75, Bor86a]. return [Coa86, HC84]. revealed [WH91]. reverse [FS84]. Reversibility [Bra79]. Review [Aba86, Abe90, Abe96b, Afi98, Afi99, Afs98, Afs99, Agh86, Aik97, Alg95, All98, And99, And88, And90, Ano74a, Ano74b, Ano75a, Ano77a, Art74, Art75b, Art76a, Art77, Aul87, Aus93, Bai75, Bak89, Ban98, BI98, Bar98b, Bas74, Bat97, Bat93, Bea97, Bea93, Beh96a, Beh96b, Bel81, Bel76, Ben86, Ber86a, Ber84, Bie95, Bir97, Bix96a, Bix96b, Bla79, Bla93, Ble86, Blu89a, Blu89b, Bol86a, Bol86b, Bol92, Bon89, Bor86a, Bor94, Bor00, Bor90, Bor93, Bos87, Boy85a, Boy85b, Boy86g, Boy86d, Boy86e, Boy86b, Boy86a, Boy86f, Boy86c, Boy87b, Boy87a, Boy88, Boy89a, Boy89b, Boy90b, Boy90a, Boy91, Boy92, Bro86a, Bro90a, Bro93a, Bro94, Bro95a, Bro96, Bro98a, Bro98b, Bro00d, Bro00b, Bro00c, Bro97, Buc91b, Buc98, Bud91, Bue86a, Bue86b, Bus95, Cal89b, Cal89a]. Review [Cal00, Cam97, Car95a, Car95b, Car88, Cas92, Cas96, Cas97, Cas98a, Cas98b, Cer83, Che86b, Che98a, Che99, Che76, Chi93, Cla75, Cob96, Col96, Col98, Con93, Cor94a, Cor94b, Cox93, Cox94, Cox98, Cox00, CS90, Cre95b, Cre95a, Cri81, Cul88, Cur97, Dan84, Dav76a, Dav77b, Dav79b, Dav92a, Dav96b, Dav98, Day99b, Deb86, DK90, DC76, Die81, Die82a, Die82b, Die84a, Die87, Dil97a, Dil98, Div87, Dog98, Dol98, Dom78, Don70, Don74, Don76, Don77, Dra89, Dro77, Dro78, Dro98, Dun96b, Dur76, Erc95, Ere80, Ere81a, Ere81b, Eva94, Exn97, Exn98, Fis99, Fis79, Fis81a, Fis81b, Fra86, Fro96, Fu96a, Fu96b, Gan74, Gar86, Gar90, Gel96, Ger79, Glu95a, Gol74b, Goo99, Gre96, Gro75a, Gro75b, Gro76a, Gro76b, Gro76c, Gro77, Gro87, Gro88, Gro90, Gul85, Haa96, Han96]. Review [Har70a, Har87, Har88, Har93a, Har83, Har70d, Har70c, Har70b, Har71b, Har78a, Har79, Har75d, Har75c, Har76c, Har76b, Har76a, Har77a, Har77b, Har78b, Har77c, Has98, Has95a, Has95b, Haw90, Hay95, Haz98, Hei78, Hei88, Hei86b, Hei86a, Hei87, Hei90, Hei95d, Hei94, Hen95, Her82, Her97, Hil96a, Hil96b, HC94, Hin70, Hjø99, Hoc89, Hol99, HC84, How87, How89c, How89a, How89b, Hun96a, Hun96b, Inh76, Jac96, Jaf98, Jas97, Jen95a, Jen95b, Joh86, Jon86, Kan94, Kas96a, Kas96b, Kat00, Kaz75b, Kaz75a, Kaz77, Kaz79, Key96, Key97, Kie98, Kin98b, Koc79, Koe75, Kol75, Kra79, Kri86a, Kri86b, Kri87, Kri86c, Kri89, Kuc75, Kuc76b, Kuc76a, Kun82, Lam00a, Lam99, Lam00b, Lan77a, Lan78, Lan80, Lan82a, Lan87a, Lan87b, Lan89a, Lan90, Lan93, Lan73]. Review [Lan76, Lan77b, Lan81b, Lap99, Lap00, Lar74, Lar90, Lea98, Lea75, Lea76, Lee96a, Lee96b, Les89, Les96, Lev76a, Lev76b, Lid96, Lie96a, Lie96b, Lil78, Lit96, Liu97a, Liu98, Los96a, Los96b, Lun80, LC90, Lyn90, Mac96, Mad96a, Mad96b, Mai99, Mai90, Mal99, Mal00, Mar80, Mar95, Mar85, Mar75, Mar77b, Mar96, Mat85, Mat97b, Mat97a, McC96, MT96, McG95, McG96a, McK96a, McK96b, McL90a, McL90b, McM76, Mea82b, Mea82a, Mea83b, Mig79, Mig80a, Mil93, Mil90, Mim85, Miz98, Moo86, Mur93, Myb98a, Myb98b, Myb99, Nan95b, Nan95a, Nel97a, New73, New76, Nor96, Nor98, Nov82, O’C93, O’C95, Old91, Ols74, Ols75, Ols76, Ols89, Ora00, Par97, Pas90, Pat85, Pat90, Pec85, Pee96a, Pel93, Pen93, Pet96b, Pet96c, Pet96a, Pet98a, Pet98b, Pet99, PJ87]. Review [Pic86, Pra90, Pre71b, Pre71a, Pre71c, Pre70, Rag96, Rap00, Rei85, Rei75, Rei77, Ric87, Ric89, Ric77, Rob96a, Rol87, Rou89b, Rou89a, Rou76a, Roy90a, Roy89, Roy90b, Rum90, Rus77, Rus85, Rus87, Rya80, Rya93, Sal72a, Sal86, Sam95, San97, San98a, San98b, San99, Sav97a, Sch96b, Sch97b, Sch87, Sch98a, Sch73b, Sel71, Sem98a, Sem98b, Sem00a, Sha97b, Sha98, Sha86b, Sha86a, Sha90, Sha96, She91, Shi86b, Shi86a, Shi88, Shu78, Shu80, Sie70, Sie86, Sie89a, Sie89b, Sko98, Sle94, Smi96a, Smi96b, Smi97, Smi98c, Smi98b, Smi98a, Smi00a, Sny98, Sop96, Sou96, Spi94, Sta75a, Sta77a, Sta77b, Sta98, Ste99a, Ste00a, Ste77, Str76, Str78, SS95, Swa75a, Swa75b, Swa76c, Swa76b, Swa76a, Swa77b, Swa77c, Swa78b, Tau74, Tay94, Tay95, Ten88]. Review [Ten90, Ten93, Ten94, Tes79, Tes86, Til70, Tom98a, Tom98b, Tom96, Tom99, Tor89, Tsu79, Tur98, Tzv90, Und96, Und98a, Und98b, Usd85, Vag80, Van90, Van97a, Var75a, Var75b, Var78, Var97, Vei88, VG93, Wal00a, Wal86c, Wal86b, War79, War96a, War96b, War98a, War98b, Wei79, Wei88, Wei91, Wei93b, Wei93a, Wei94b, Wei94a, Wei95a, Wei95b, Wei96a, Wei76, Wei96b, Wel70, Wel71, Wel78, Wel80b, Wel75, Wes89, Whi89a, Whi89b, Whi98a, Whi98b, Whi98c, Whi99b, Whi00, Whi98d, Wil98b, Wil00b, Wil00c, Wil85b, Wil94, Wil95a, Wil98c, Wit96a, Wit96b, Woo88a, Woo77b, Woo90, WH70, Wri71a, Wue97, Yin95, You75, Zvi86, Zvi87, Zwe95, Cha86, Dun96a, Gar76, Glu90, Gro78, Hap99, KWR98, KM78, Kra87c, MM89, Nis99, Ric70, San75b, Sar75]. review [Sum84, Wel00, BR80a, Gup90, HG98, Pre71b, Sie90, Swa76d, Koe75, Vei88]. Reviewed [Ano74c, Ano75b]. Reviewing [Glo88, And99, Cas95, Wal87]. Reviews [Boo00, Bro00a, Hjø00a, Lew00, Mac79a, Mai00, Sch00b, Sem00b, Shi85, Smi00b, Ste00b, Sti00, Tho00, Men95b, SRS98, Woo77a]. Revised [HK79]. Revising [And94]. revisited [BT84, Bur81b, Col93, Eas89, HAA87]. revolution [DK90, Doc91, Jac94, Joh86, Sem98a, War98b]. Rheinland [Wel71]. Rheinland-Westfalen [Wel71]. rheumatology [Tho72]. rib [Gro75b]. Richard [Dav76a, Dur76, Gol74b, Hol99, Kaz75a, Kuc75, Mar80, Nov82, Pet99]. Richmond [Coc98]. richness [GLL90]. ride [FW99]. Ridge [WH70]. right [Bro93b, Fag95, GLL90, Hal81, Lev99]. RightPages^{T M} [HOS^{+}93]. Rights [Woo87a, Bra89, Cre95a]. Rijsbergen [Bla79]. rise [Soe99]. risks [KJR99]. rivalry [Tay95]. Ro [Jor89, Lan89b]. Road [Ano74a]. roadmaps [KET99]. roam [Nug91]. Robert [Fis81b, Har77b, Kaz77, New76, Pet99, Pre71c, Ste00a, WH70, Bor99, Ras99, Van99, WS71]. Robertson [Cer79]. Robin [Var75a]. Robustness [CC98, Rou92a]. Rocchio [Cod73]. Roessler [Wel71]. Roger [Mim85, War79]. Role [Bur70, Pre71b, Abe90, Afs98, AL98, Bax85, Bud91, Cas91, Don79, Dro95, Erc95, Gra82, Har74, Jon91, KJR99, Kuh99, LR79, LD85, MB86, Men92, Mon94, Pai93b, Sch99, Ste99a, Sut94, Van97a, Var99]. roles [DH97b, Gar92, Woo77a]. ROM [CZ88, Dav88, EH94, Jas94, LS88, Rie88, Sch88a, Sch88b, Zoe88]. Roman [Jac72]. Romanization [Bra79]. Rome [Har76b]. Ronald [Har83]. root [LJ70a]. root-stem [LJ70a]. roots [AKE94, Win99]. Rotenberg [Lam99]. rough [BG98, LY98, Miy98]. round [Cha99b]. Rousseau [Bur92b]. route [Mac96]. routine [Fid91a]. Rowley [Lun80]. Royal [Buc91b, Ell70, GBDH76]. royale [Gro76a]. Rubin [Hol99]. Rubinsky [Wei97a]. Rule [Wu98, Egg87, Egg93, Kra78]. rules [JGR^{+}95, Tan96, Tan99, Wel83]. ruminations [Whi87]. Rush [Koe75]. Russell [Woo77b]. Russian [Rya80, GS93c, JBH77]. Rutgers [Gro75b, ABLS88]. Ruth [Fis79, Har75d, Pat85].

S [Bal89, Bel81, Caw74, DC76, Kaz77, Pre71c, Tau74, Van99, Hjø00b, Shp92]. Sacherschliessung [Wel71]. Sachs [Woo77b]. Sacks [Mim85]. safe [But99]. safety [Has98]. Saharan [Aiy88, Old87, Tia89]. salience [Cro99a]. Salton [Art76a, CML^{+}96]. same [SSD99]. same-frequency [SSD99]. sample [Pit75]. Sampling [Han91, Jai72, LT74, Rou88, Spe71, Bro70c]. Samuelson [Koc79]. San [Har70a]. Sands [Woo87b]. Santa [RCJS94]. SAPHIRE [HH95a]. Sarah [Bro00d, Lar74]. Sarson [Fis81a]. satellite [Sha86a]. satisfaction [AL84, Bru98, Soe76, SBW90]. satisfactory [Sol74]. Satisfiers [ZvD00]. Saw [Ano74a]. scalability [GCH99, Ror90]. scalable [Cal00]. scale [Sol74]. scaling [McG84, Ror99b]. scanning [AC93, Mad96a, Mad96b]. Scarborough [Gro75b]. Scarecrow [Art77, Gro77, Har75d, Har75c, Har77a, Har78b, Kuc75, Mar75, Rei77, Rus77, Shu80, Sta77b, Ste77, Wei79]. SCATT [Woo77b, DS80]. Scatter [YG90, PS79]. scattering [Coo89, DG78a]. Sceince [Hei78]. Sceince-No [Hei78]. schedules [Aul71]. Schell [Har76c]. schema [WBX98]. Scheme [CC79, Cra86, Dav76b, EW70, KSDR90, SC80]. Schemes [FK73b, Sav97b, Low71]. Schlichtmeier [Tyn75]. Schlipf [Har76a]. scholar [GBL89, Van97a]. scholarliness [Per83]. Scholarly [BK96, Har98, KM99, KRD95, PR98, PZ84, Cro99a, Dro95, GCS97, Hap99, HP00, KB98b, KF92, Kim00, PPP98, PZ83, Sch94, SPWH99, SRS98, And99, Jas97]. scholars [Kil95b, KM00a, SBW93]. scholarship [Kra86b, MS00]. School [Ano84m, Buc91b, Gro75b, Har78a, Har75c, Har76a, Kaz77, KL79, Lil78, SM88, SH89, Swa76b, Tau74, Coo85, FDD^{+}99, GSKL00, Hay88, Mar94, Neu95, Old91]. school-age [Mar94]. Schoolhouse [KL79]. schools [ABLS88, MN83a, Woo84, Dol98]. Schrader [Shu80]. SCI [CN81, Fro96, Boy86b]. Sci-tech [Fro96, Boy86b]. Science [Ano77a, Ano81c, Bal89, BDZL89, BR94, BHCC^{+}74, Bel81, BHWG95, Bot73, Bro99a, Can99, Deb72, Don74, Ere81a, Flo00, Gol79, Gro75b, GSK98, Har94, Har79, Har76a, Hei88, Hei95c, Kaz77, KS98, Kra87a, Kra87c, Kra99c, Lan94, Lei73, LB89, LB94, Pet94, Poy82, RK79, Ric77, Rus70b, Sal79b, Sap94, Shc79, Shu78, Shu80, SA91, Smi99, Swa76b, Swa78b, WD97a, WT94, Yov70, ALD96, Ami73, AI76, ABLS88, Ano84n, Art75b, Bar96, Bat99b, Bat99c, Bel75, BR76, Bel77, BIFW71, BWKG75, Ben00, Bez82, BR92, Bor84, BMvR91a, BMvR91b, Bri72, Bro80a, Bro99b, Buc99, Coo82, Cox83, CR80, Cur93, Dan00, Dav98, Day00b, Dio83, Dod79, Don72, Dou72, Dow77, Ell70, Ell86, EAW99, EF96, For99, FRGL75, Gar71, Gar88, Glu94, Gol87, Gro88, Gul87, GSD74, Har71a]. science [HH92, Har92b, HNW93, Has85, Hei95b, Hen99, Her84, HA95, Hjø96, Hoe72, Hoo91, Hur88, Ing97, Inh76, Jac96, Jas94, Ker77, Ker84, Kil93a, KDJ94, KRH98, Kno77, Koc84a, Koc84c, Koc87, KRD95, Kra91a, Kuc82, Lan87b, LMM^{+}73, Law74, Lev77, LAD97, Lyn77, Mar92b, Mar83c, Mar99, Mea78b, MWR72, MF88, Mon72, Mos70, Nee81, Nis99, Old91, Ols76, Pal99b, Pop75, Pou89, Ray97, RBR80, Rep89, RB83, Ric88, RL97, Ros74, RTH71, Sag75, Sal73, Sal85, Sal96, Sar75, SBA85, Sar99, SS70, Sch99, Seg92, Sha95, She71, Shi86b, Shp92, Sla75, Sma99, Smi74a, Smi98c, Spa99, Spi97a, Stu78, SOM^{+}99, Swa75a, SM75, Swa78c, Tem94, Tsa98, Van90, Voo74, Wal81, Wal85, War91, Wei96a, Whi89a, Whi77, WM98, WM00b, WK75, Win99, Woo90, Wri73]. science [Yer71b, YG90, Zhi97, de 71, BL85, Bel70, Buc99, CN81, DH97a, FM97, Hum70, Mac79a, Min78, Poy82, SKvB92, Sch73b, Sta82, Swa73, WWB97, Art74, Bas74, Dan84, DC76, Hei78, Hol99, Str76, Usd85, Ano84k, Ano84l, Ano85g, Bat99a, Smi98c, And90, Cas92, Che98a, Hei86a, Rag96]. Science-No [Swa78b]. science-profession [Mar92b]. Science-technology [WD97a]. Sciences [Ano74a, Har78a, Rya80, And99, Bix96a, Bix96b, Bot70, Bri84, Bro79, Fre82, Gor82, GWDS86, Hei95b, Kim82, Ley97, Lin87a, LM81, Moo73, NZ91, Pac82, Pal87, Per85, Pie92, QLA97, Raf75, Ras90, SS00, VS87, Ano72a, Lan76, McC89, Mig80a, Ber86a, Har88]. Scientific [Abe90, Ami73, Avr79, CS90, Cre76, Gar90, Ger79, Har84, Kir73, Lan80, LC90, MV95, PZ84, Sch73b, Shu80, Tho82, Tsu79, Woo77b, Yov70, Afi98, Ain84, AP96b, AP97, Ami70, And78, Ano75a, BSM90, Ber72a, Bol92, Bon90, CN73, CN81, Caw74, Caw78, CNMS97, Coi77, Col81, DG78a, ER00c, Ehi90, EN98, FCJ^{+}92, FK73a, GCH99, Gan74, Gat92, Gor79, Hoc74, HSP75, Hur00b, Inh75a, Inh75b, Jon76, KM00b, Kor74, KH71, LaF00, LB83, LBKO87, MAN80, Men92, MM78, NCB72, Now81, NvR98, PZ83, Rao80, Ros74, ST89, SC76, Sma73, SG89, Swa78b, SK99, Tag78, Tau74, TvR90, Vin74, Voo99, WKMG00, vdBL96, Blu89a, Blu89b, Har70d, Tay95]. Scientist [TC73, Whi99a, Bec84, GdCD81, Ler79, Dro77, Ben86]. Scientist-poets [Whi99a]. Scientists [Pre71a, TL74, Bro99c, Cra71b, Cra72, HH92, JJ75, Man82b, MCT73, Sub84, WW73, ZK72]. scientometric [vdBL96]. Scientometrics [BI98]. scofflaws [Smi98b]. Scope [Mea79a, Mea83a, Wei97c]. scores [PvR94]. Scottish [DHS00]. screening [Tho73a]. SDI [CT73, Col81, Mal74, PS79, SF78]. SEARCH [Ere81b, AL84, Caw74, Har90, Mor76, Shu80, Bat77, Bat79b, Bel85, Bil00, Boo74b, BHH96, BM72, Boy86d, BM89, BM80, BC99, Buc79, CN95, CMK^{+}98a, Coo83a, Coo83b, CBM^{+}94, DCFQ00, DKP00, FWR95, Fid86, Fid91a, Fid91b, Fid91c, FSTV99, GGJ91, Gol74a, HT91, Her99, Hoc89, HY93, IS98, Jah74, Kir81, Koc84a, Kra86b, Kuh91, Kuh99, Lan93, LWG74, LS99, LLM91, Lin78, LDS^{+}00, LP99, Mar83a, McC88, McG95, Mic94, Nic00, PK00, Qiu93, Rad82a, Rad82b, Ras80, RJ76, RW00, Sch98a, Sch98b, Sie89a, Spi96, SS97, Swa72, Swa89b, Tag72, TH85, Voi97, Zin00, Ano84m, Fri73, Mar82a, Sie89b]. search-engine [LP99]. search-support [GGJ91]. search-terms [Tag72]. Searcher [DW95, Fid91a, Fid91b, Fid91c, Bel85, Jaf98, Sch74, WD97b]. searchers [Con95, HY93, SK88b]. searches [Boo80, DW96, Han86, Haw81, Kil95b, KMB99, KM00a, NL83, NKM77, SK88b, SBW90]. Searching [Bla80, Bun80, Dos86, Fox86, Haw86, KL98b, KL98c, Lan80, LH86, MHA82a, NC86, Oja86, Ros79, Ros82b, Rot79, Swa77b, Voo99, Wil86, Ang75, BF88, BHWG95, CHSS98, CCW84, Con84, Con95, Coo77b, CBM^{+}94, Dal90, Elc78, Fen81, Fid84, Fid91b, FDD^{+}99, FRGL75, Har71d, Har84, HC96, HH95b, HH95c, Kil97, KT90, Lan81a, Lar91, Laz00, LBW00, LR72, Mal74, MCT73, McC80, MHA82b, Mea00, Mic94, NT96, O’C70, O’C73, O’C80, PW89, Par00, Pat81, RW00, Rot81, SCD98, ST86, Sha90, SBW93, Sie90, Slo00, Smi74b, Spe80, SEW00, SH96, Vig83, WdBFF95, Yer73, Yua97, Fid91c, Kri86a, Lan93]. SEBD [Ano74a]. Second [Smi98c, Cor94a, Dog98, Inh76, Mas97, Swa89a, Ber84]. Second-hand [Ber84]. Secondary [Coo82, Ber82, BR80b, Dol98, EN85, Fre82, Gra82, Kin82, Mac82]. Section [Bra87, Lin87a, Spr87, Wes87, Wol87]. sector [Gar92, Hal85, Mar98, WWSR00]. securities [BR97]. security [But99, Nor98, Sal80, Lev73b]. sed [But99]. Seeing [Wal00b]. seeking [BR97, Bro98a, Bro99c, Bru98, Cha81, Cha91, Che98a, Deb86, Der84, Dre99, GMPS00, HH97, Hil96a, Hil96b, Hir99, Jul99, Kri89, Kuh91, Mar89, Mar92a, OO97, Pal99a, SC85a, SKCT88, SK88a, SK88b, SMC83, Var99, Whi00, Xie00, Yan97, Dun96b]. seen [CMK98b]. segmentation [LNL99, WT93, WT95]. Seidell [Hir89, Vea90]. Seitz [Die84a]. Select [DF82]. Selected [Boy86f, Sar71, WFK73, BW85, FM97, Hap99, New73, RM99, Smi84b, Var75b]. Selecting [Zmu75, Coo73a, Coo73b, FW90, How89b]. Selection [Cha88, Hoc74, Car75, DCRS98, DS73, FWR95, Fid86, Fid91a, Fid91b, Fid91c, FSV93, HC96, IS98, JJ75, MDS70, MUD99, Rou74b, Shn86, SS76, WS98, ZC92, SS97]. selective [And78, Coh99, FRGL75, SMB72, TEC99]. Selector [Bur92a]. Self [Fis79, Law82, EP97, FK73a, Kuv73, Lit96, OCN97, Sch97a, Sla73, Spe71, NvR98]. Self-Citations [Law82]. self-generated [FK73a]. self-instruction [Sla73]. self-observation [Spe71]. Self-organized [NvR98]. self-organizing [OCN97]. self-sufficient [EP97]. self-supporting [Lit96]. Self-teaching [Fis79]. selves [McD99]. Semantic [HNW93, Liu97b, Qin00b, AM77, Che00a, DDF^{+}90, GD98, Kat72, PW89, YR77, Bon89]. semantics [CDV98, Dil00]. semi [Caw74, Har71b]. semi-annual [Har71b]. semi-conductors [Caw74]. seminar [Buc91b, Wes00]. semiotic [Art76b]. semistructured [Rap00]. senior [MS00]. sense [Bur96, Sol97a, Sol97b, Sol97c]. sensitivity [Tho90]. sentence [Cra91, Haa93]. sentences [O’C73]. SENTINEL [FFKS99]. September [Har76b, Old91, Jac71, Kol74]. SEQUEL [Mac79b]. sequential [Boo83]. Serge [Rap00]. Serial [HK80, RG70, FG70, MG72, VC70]. serials [Fay72a, Fay72b, Fay73, Fei73, Koe79, RM97, Swa76a, Rus77]. Series [Hei78, Kaz77, Swa78b, BL72, Mim85]. server [Cal00]. server-side [Cal00]. servers [KMG^{+}93]. Service [Har75c, Wei97c, BWV72, BM80, Cas78, Col81, DH70, Kan76a, Lyn98, MP74, MMB99, Mor94, Smi96a, Smi96b, Bra87, HOS^{+}93, Var78, Wal72, Pen93]. Services [AL96, AKS96, Abe96a, Ano72a, Art74, BR94, BCB96, Har94, Hay96b, Lan73, Lan94, LB94, New73, Pet94, Sap94, Sel71, Tho82, War79, WT94, Afs99, Art75b, BD88, Ber82, Bix96a, Bix96b, Bot70, BR80a, Boy86b, Bra87, BR80b, Byr75, Cas94, CS91, Coo82, Cul85, DPD99, DRM98, Dol98, Dou72, Egg88, FW88, Fre82, HDRD88, Hei86a, Hel71, Hoc89, KR80, Key97, Kin82, Klo89, KRC81, Kri87, Kri89, LN82, LS99, LD88, Lin78, Los96a, Los96b, LH97, Mac82, Mea86, MY85, NT85a, Pac82, Pen70, RFK97, Rob80b, SK97a, SK97b, Spr87, SCLL73, Str84, WL80, Wel78, Wol86, Zai77, Cas79, Kuc75, Lan80, Pat85, Shi85, Wei76, Bak89, Bos87, Buc98, Dun96b, Kin98b, Myb98a, Sle94, Tom96, Woo88a]. Session [Har77b, Swa76c]. sessions [Spi96]. set [Bra70, Gro90, IC97, Jon76, KZC86, LY98, RT81, TO84]. sets [Art71, Bue85, Kin92, Miy98, Rob79, Ror99b, Sac76]. setting [Bon79, BF91a, Gar92, SC85a, SC85b]. seven [Kra92b]. several [KG72]. severe [PS79]. SGML [Gol97, MS97, Mas97, SD97]. SGML. [FS97a]. Shakespeare [Han92]. shall [HW99]. Shanghai [VTS97]. Shannon [San99, Col93, Lyn77]. Shape [RF00, DS96, Dil00]. Shaping [Han96, Cas94, KM00b]. share [PBL^{+}00]. Sharing [Bry70, McC00, Ano80i, Gul85, Hei94, Rei85, Rei84, Ste80]. Sharpe [Mim85]. Shcharansky [Ano81c, Gol79, Min78, Sal79b, Shc79]. shelf [MN81]. shelves [Los92]. Shera [And90, Don77]. Sherrod [Don74]. Shifting [Woo87b]. Shifts [Rob00, Xie00]. Shining [Woo87b]. Short [Bro00f]. showcase [Jac72]. Shpackov [Deb92]. Shuler [Ste00a]. side [Cal00]. sides [Fau96]. Sidney [Har70d]. SIG [Sch99]. SIG/ [Sch99]. signature [KSDR90, KT90]. significance [EB98, Gel79, Ste74a]. Silva [McM76]. similar [Coa86, MHCW89]. similarities [BCB95, Qin00b]. Similarity [CO85, Car79, Rad82b, BSY00, BCB98, Cro94, Fal98, GLW86, HW80, He98, JF87, Kwo85, PT87, Ror99b, WW00, ZK99a]. Simon [CL89, OV98]. simple [Ror90]. simpler [Mol92]. simplified [Her78]. simulated [BMSM90, CSSI98, Gor90]. Simulating [TT79]. Simulation [Qua96, CC75, ENN98, YFR81a, YFR81b, YFR81c, Man82a, Sha75, YFR81b]. Singapore [Joh86]. Single [YR77, Coa86, Cro77, Fau96, Wil84]. single-link [Cro77]. Single-pass [YR77]. Sir [Byr76]. SIRE [NKM77]. site [BLPG99, Koe99]. sites [Bro00d]. situ [MB95]. situated [Mac96]. Situating [BNS^{+}00]. situation [Kra73, Pul74, Kra74a]. situational [Bru94, Glu95a]. Sixth [BI98, Joh86]. Size [DS79, GR74, KE99, Kin92, NT85b, RV78, Sic92, Whi91]. size-rank [NT85b]. sketch [WS71]. skewness [Seg92]. skill [Coo91c]. skills [Mal99]. SLA [Pre71c, Ano72c]. sleep [Cra71b, Cra72]. SLIC [RR71]. slide [Pap88]. slogan [EB98]. Slovene [PW92]. Slype [Die84a]. small [ALD96, ASR85, Cha91, How87, Ker77, LAD97, Mar95, Tom98b, Var75a, VTS97, Wei73]. SMART [Sal72c, Ber72b, CCRY98]. Smirnov [Lon92, Bur94, Lou92a]. Smith [Die81, Har78a, Pec85, Pre71c, Wil00b]. smoothly [EM94]. Snee [Cal00]. Social [Die82b, Ell86, Har79, KRH98, Lan77b, Rya80, Wri73, And99, Beh96a, Beh96b, Bri84, Bro79, Cas94, Day97, DH97b, Doc91, Dod79, Gor82, Kar98, Kay98, Kim82, Men85, NZ91, Nee81, Pet96a, Pie92, Raf75, Rao80, Sol97b, Vei80, Wei88, Whi77, Whi81, Whi83, Whi98d, Fre82]. Societies [Har70c]. Society [Aiy88, Ano84k, Ano84l, Ano85g, Art79, BI98, Bat99a, Boy91, Bro95a, Cer83, Hum70, Kra87c, Kra99c, Lan82a, Ols74, Smi99, Gar71, Gat92, Lie96b, Mar94, Old87, Tia89, War99b, Art82, Buc99, Die82a, Ell70, Kra87a, DK90, Pra90]. socio [Caw74, Mar99]. socio-scientific [Caw74]. socio-technological [Mar99]. Sociodemographic [KS77]. Socioeconomic [Ere85]. socioemotional [ZK72]. sociological [She71, SWB79]. sociology [CMRWW93, Oro77]. Sociomedia [Beh96a, Beh96b]. sociotechnological [Mar00]. Soergel [Lea75, Coo76]. soft [Mat85, War98b]. Software [CMK98b, Dav77b, Dav79b, Ger79, HW71, CZ88, How89b, Mea83b, Zoe88, Har83, Wei96b, Jaf98, Les96, San97]. soil [KDJ94, Mud98]. Sola [Koc84b]. SOLAR [MRW73]. Solla [Koc84c, Ste84]. solution [HM70, Kir81, OG70, TEC99]. Solutions [Har77c, HJD^{+}99]. Solving [Sch97b, SG89, SH96, WdBFF95, Zha98, Jon86]. Some [Bel75, Ber82, Kin98a, Lay94, Lyn88, MZ79, Pra83, Sho71, Tag72, Tag78, TS96, Wal87, Wil82, de 71, Cam96a, Egg87, FB88, Kim73, Kir81, LMM^{+}73, MP98, Sha83a, Bri91, McI97, O’C70, Ric76]. Sons [Fis79, Mea83b, New73, Sta77a]. Soon [Jor89, Lan89b]. sorted [PZSD96]. Sorting [Bro00c]. Soul [Mea82a]. Source [Egg96, BI00, Cra86, Kin76, Soe72]. Source-item [Egg96]. sourcebook [Kuc76a, Mig79, Boy89a, Kri86b, Sha90]. Sourcerer [Rod95]. Sources [ASS73, Lan76, Lar74, WNAR99, BR97, Ber86a, Egg96, GSKL00, Hai97, HG98, Hoc89, Kuh99, McC78, Mic94, Rod95, Rou74b, SMC83, Swa89d, Wig95, Nel97a]. South [Dou96, MT96, Pee96a, SR95, Var75a, Pou89]. Southeast [Joh86]. Soviet [Bec71, Bel75, Kir73]. Soviets [Ano81c, Gol79, Min78, Sal79b, Shc79]. Space [Har77c, Liu97b, MA76, BSR93, BSR98, Bro80a, Cam96b, CNMS97, CHSS98, Day97, Dil00, EM94, Fu96a, Fu96b, McC90, McK00, RW86, All98]. spaces [KW97, McG76, PLM00]. Spain [Can99, FM97]. spanning [Wee92]. Spatial [Dil00, Glu94, TFTB94, Che00a, CP98, FS84, KDJ94, Liu97a, Mon94]. spatial-semantic [Che00a]. Spatial-semantics [Dil00]. Speakers [JJ75]. Special [Blu89b, Bro99a, Egg94, Kra94b, Lie94, Pre71c, Bea80, BB96a, Bro70c, CT99, Che00b, Che00c, DDC78, DCD81, Fro94, Glu94, JJ77, JJ80, Nug91, SK99, Wes87, Wil97, Ano00a, Ano00c, Ano00b, Tom99]. Specialist [Bro95a, Ger79, EW70, Har87]. specialists [Cra71a, Bai75]. specialized [DS73, EW70, Dou72]. specialties [MM78, vdBL96]. Specialty [Har75a, Col92, JJ75, McC88, RBR80, Sma86]. specific [Kil95a, LMSC82, Sca89]. Specification [Dav79b, Gro75a]. specifications [Cam97, Tau74]. specificity [WY92]. spectral [McG96b]. spectrum [Sha86b]. Speculations [Cox88]. Speech [Phi92, WH98, CWBL00]. spell [LNL99]. Spelling [PZ84, Zam80, BSY00, PZ83]. spelling-correction [BSY00]. spider [CCRY98]. spinoff [Lev73a]. Split [Coz82, NT85b]. spoken [CWBL00, Mea82b, TC00b]. spoken-access [CWBL00]. sponsored [Old91]. Spread [Sha86b]. Spring [Han75, Cam97]. Springer [Lev76a]. Springer-Verlag [Lev76a]. SPSS [Har78a]. square [O’H70]. Squibb [FRGL75]. St [Wri71a]. stability [CO85, Yu76]. Staff [WH70, Bam79, Mea81, Sap94, Tor89, Kaz75b]. Stage [Koe92b, IS98, Koe92c, KS97, MN83b]. stages [WW99]. STAIRS [Bla96]. Stan [Afs99]. Standard [And94, Gar92, Adl92, Cha77, Mou77]. Standardization [Leh92, Pit82, Gre92, Whi98a, WL81]. standardized [Win70a, YAH90]. Standardizing [Van98b]. Standards [Adl92, And94, Bea92, Cox92, Gar92, Gre92, Leh92, LS92b, Mol92, Tom92, Wei70, WC92, Kra99b, Mas97, Rus87, Bra79]. Star [Bro00c]. Start [Flo99a]. State [Pre71b, Sut96, Bar78, Boy91, Bra94, Cha81, CH99, FSTV99, Fre82, Gat92, Koe92c, Pen70, Qiu93, Ren80, Whi76, YdKKM87, Cim95, Kos94, Wit81]. State-Of-The-Art [Pre71b, Kos94]. statement [Har70c, Kra86b, Tal74]. States [Art74, Bol92, Nel97a, Rus77, Shi86a, Wel75, Mor78, ST89, Ste80, Wil97, Gar92, Mea72, Rus85, War99b]. Statistical [DF84, Mar80, Sha83b, Ami70, DS73, HC90, JGR90, Kim82, LK97, LT74, LCL85, MZ79, RFK97, Ros71b, Wal85, War73]. statistically [Car75, Wil92a]. Statistics [LG96, Kuc80, LG99, MZ80, YW96]. status [Bat83, Coa86]. stem [LJ70a]. Stemming [ASAOE99, Hul96, Sav93, AYS96, Lov71, Pai96, Sav99, PW92]. stems [AKE94]. step [Fis70]. Stephen [JK91, TZ93]. Stevens [Gro77]. Stevenson [Pec85]. Stewart [Ora00]. STI [KW77]. Stichler [Pet99]. Still [Dav98, Bor96, CH99]. STINFO [Lev73a]. Stirling [Har78a]. stochastic [OV97]. Stop [Flo99a]. stopping [Kan87]. stopword [Sav99]. Storage [CZ88, Dav88, LS88, Rie88, Sch88a, Sch88b, Sie70, Zoe88, Buc75, Bur81a, Goy87, GR74, JB71a, MRW73, RV78, Ros83, Sch73a, Was74, WRE^{+}70, Kei70]. store [Par82a, Par82b, Par82c]. story [Rob96b, Rol87]. storytelling [Haa96]. Strategic [Eva94, Sav97a]. Strategies [Die82a, Aul87, Car75, Car88, FPS00, KWR98, Lan93, Lan94, Mar89, MBH94, Pal99b, Sie89b, Swa72, Xie00, Xia92, Haw90, Whi98c]. strategy [Bro93b, Caw74, Gol74a, Jah74, MB91, Swa89b, Zin00, Boy89b, McG95]. strength [Wil85b]. string [BSY00, Cra80a, Cra86, Goy84, And88]. string-similarity [BSY00]. strongly [Sic92]. Structural [DF85, Dor85, Yar97, Dor88, GCH99, MRY73, Str74, BMvR91a]. structural-equivalence [Dor88]. Structure [BCM^{+}70, BM75, Deb72, NPG76, PN79, BSM90, Buc89, COC90, Dan83, HG00, Her97, Kno77, Liu96, LH97, McC84, McC86, Nah98, PR98, SW92, Sha91a, Tau84, WG81, Win70b, Wol96]. Structured [Dav76a, Rum90, Fis81a, GFHR97, SM91, Zin00, Gol74b]. Structures [Pal99b, BR95a, Bri91, Cra91, Dom78, Her97, HJD00, MCD73, Str78]. structuring [NvR98]. struggle [Ere81b]. Struminger [Bas74]. Student [Bry70, Smi74b, KRC81, Mel72, Neu95, OWF99, Old91, Str74, Wat98, WdBFF95]. Student-to-Student [Bry70]. students [Col00b, Cov00, FDD^{+}99, Lee00b]. Studies [Und98a, Whi98a, Bac72, Beg95, Bor86b, Cur93, Day00a, ERV00, Haw77, HCL^{+}00, Hur80, MLC80, Pal99a, Spe71, SEW00, SM75, Tem94, Wal98, Whi96, YM99, Per82, SM88, Woo88b, Sta75a, Roy90a]. Study [Coz82, Ger79, Hei78, Lan77a, Mac79a, Rou76a, Spi97b, Swa78b, Bar94, Bon90, Bur93, BR95b, Cas92, CH00, Caw78, CYW99, Chr72, CCN93, CB94, CMK98b, DH97a, DG91, Dou96, EB88, EH94, EK97, Erc92, Fid84, GB95, Gro78, Gul85, Har90, Har98, Her99, Hul96, JJ77, Jan91, Key97, Kim00, LL85, LBBR94, LBIP97, LK91, LBKO87, Lie96b, MWY95, MS00, MP98, Mil95, Mud98, Neu95, NML99, OV92, PKK74, PvR94, Pra83, Raf75, RM97, Rob80b, Ros71b, Sal72d, SKCT88, SK88a, SK88b, SMB72, Sha83a, SH89, Spi96, SS00, SR95, Str82, SS89, Tan96, Var75b, VTS97, WWSR00, War99b, Xia92, Yua97, YM99, Zha00, Swa76d, WS98, WW99]. Studying [AKS96, SK97a, SK97b, McG84, Smi84a]. style [Bro00d, Dav77a, Din87, PK00]. styles [Fid84, Fid91c, For00]. stylistic [Tag78]. Sub [Aiy88, Old87, Tia89]. Sub-Saharan [Aiy88, Old87, Tia89]. subfield [BSM90]. subfields [BB83b, Ste85]. Subject [Bat86, Con95, Jac96, Mcg79, Nie76, Sha91a, Sha91b, Bat77, Bau72, BB95, Blu89b, Boy86f, Byr75, Che98a, Coa86, Coo70b, DS91, DW96, DS84, Egg88, Gar76, GGJ91, Gir91, Gor91, Gue78, HY93, Jen95b, Kuv73, Lan78, Lar91, RW82, SP73, Sma80, Sve94, Wil98a, Wel71, Boy86g, Sch87, VG93, Wei93b]. Subjective [FGS81, Bro80a, Col94, Gor82]. subjects [EAW99, Kje74]. Sublanguage [LH95a, Sag75, Haa97]. sublanguages [Bon90]. submission [Bra77, Rob91]. submissions [Kra94a]. Submitted [Har70c, RW00]. suboptimal [GL92]. subscription [Lil78, Qua96, Rei85, War81]. subscriptions [SKvB92, SvB94]. Subsidiary [New73]. subspecialties [EF96]. Substances [Gol73]. substrate [Bat99b]. Subsystem [Sha75]. suburbanization [New94]. subword [Jak88]. Success [Zin00, Bat77, ER95, Han86, Hub98, Tag81, Roy90a]. success-breeds-success [ER95, Hub98, Tag81]. Successful [Aul87]. Successors [Day97, Ray97, RL97]. Suciu [Rap00]. sufficient [EP97, Hei80, Hei91]. suffixing [Har91]. Suggested [Ano83i, Dod79]. Suggestions [Wol71, Wel83]. summaries [Aul71, BM75, HSP75]. summarizing [ENN98]. Summers [Lil78]. Suny [War79]. Supercomputers [Bue86a]. superhighway [Cas97, Cas98a, Cas98b, Dav96b]. superorganics [Koe81]. supersonic [KET99]. supply [Fug85, Mow92b]. Support [RNBV98, Arm94, AL98, CB83a, EF96, GGJ91, GL96, GMPS00, Lyn91b, OG93, RM97, Rou98, SW92, Wei93a, Wol86, Sch97b]. supported [GCS97, KWR98, Nan95b, NJWC97, SH96, Wol87]. Supporting [Des97, IT89, Bax85, KM00b, Lit96, McC95]. surfaces [Los96c]. surfing [EB98]. Surname [KM00a, Kil95b, KMB99]. surname-title-word [KMB99]. surname-title-words [Kil95b]. surrogates [GMPS00, Str74]. Surrounded [Mol92]. Survey [Dil82, Var75a, Aga99, Bel76, BC96, BS92, BC97, Del90, Glo88, GLS97, KZC86, Tag75, Tal79, Tho71, Voo99, Zha00, MK72]. surveys [SD96]. survival [Hei80, Hei91]. Susan [Bas74, Bro00c, Che76, New76]. sustainability [Ste99a]. Sutcliffe [Nel97b]. Suzanne [Buc95]. Swanson [GL96, Jor89]. Swets [Hei74]. Switching [Nie76, Cha88]. Symbiosis [Art82]. Symbiotic [Wal75]. symbolic [Che95a, CN95, Liu97a]. Symposium [McL79, Rus70b, Swa76c, Gro75b, Sal86]. Syndicated [Dur76]. syndrome [BB95]. Syndrome^{T M} [Hor93]. synergetics [Liu96]. Synoname^{T M} [BS92]. synopses [May78]. Syntactic [Bon90, RSZ71, Sch90]. Syntactically [Sal84, DG83]. synthesis [Sma86]. synthetic [Car75]. Syracuse [Kaz77, SM88]. System [Elc73, Fed82b, PZ84, Yee91, AKE94, AM77, Ami70, Ami73, Ben80, BW84, BCSZ79, Bie95, Boo74a, BM80, BK81, Bur00, Bur99, Cag70c, Car74, CYFS95, Col99, Col00a, CM00, CC96, Coo83a, Coo83b, Coo85, Coo91b, CNO^{+}00, Cra77, DW98, DG83, DW95, DCFQ00, Fay72a, Fay72b, Fay73, FW90, FFKS99, FS91b, FRGL75, FNNW74, GCH99, GS93b, GQSdP98, Hal80, Hal85, HMNW86, HM70, HCL^{+}00, Hol70, JGR90, JBS^{+}77, Kat72, KE99, LJ70b, Lev73a, LFHG99, LH76, Mat70, MB71, ML90, MRW73, MCD73, MK72, Moo73, MSHB86, Mus86, NK00, Oh98, OG93, Pai71, Qiu93, Ren80, Ros76, Rub92, Sha83a, SBPP89, Sla73, Spe80, SGS70, Str82, Swa77a, Swa78a, Tan96, Teb70, Tho71, Tho90, TF89, VS87, WEWP97, Wan92, Was74, Was80, WRE^{+}70, Wei73, Whi77, Wil71]. system [Wol96, WD97b, WT95, Yer71b, YK93, ZC92, ZC00, Zmu75, ASAOE99, CNMS97, Dat79, FCJ^{+}92, GL93, Kei70, LAT79, Lev73b, MCBC89, RNBV98, SMHC93, VC70, Wei97c, Woo77b, Yan97, Sta77a]. system-theoretic [Rub92]. Systematic [DM99, Hoe72, Pep74, Swa89b, Smi74b]. Systematization [Lan90]. Systemic [Del90]. systemics [Sch97a]. System(R) [WNAR99]. Systems [BD88, BFM88, Bry70, Bur70, Bur71, FW88, Ger79, Her82, Koc79, Lan82b, LD88, LB88, Mom82, Mor76, New73, PW88, Ric77, RH99, Rus85, Sie70, Sta82, Tes79, War88, Wes88, WMME88, Wri71a, ZvBR88, AL94, Aik97, Ano00a, Arm94, Bac72, Bag81, BS88, Bar95, Bea91, BRR96, BB96a, BJ79, BMNM^{+}93, Ber82, Bla80, Boo78, Boo81, BHH96, Bor90, Bou80, Boy86g, Boy86e, Bro95a, Bur81b, Bur85, CB83a, CMB^{+}95, Coo70a, Coo72, Coo91c, Coo98, CBM^{+}94, CK96, CR80, CT87, Cua71, Cul85, DG95, DH78, Dog98, DK75, Dro98, Dum87, Dun96a, Ell76, EAW99, EF96, Erc92, Erc95, Ere81c, EK75, EC92, Fid86, Fox86, FB88, FSV93, Gil78, GO93, Glu95b, Gor90, GL96, GLW86, Har92a, Haw77, Hei74, Hen88, HS78, Her97, HBG96, HM70, HJ89, HO89]. systems [Hol86, JB71a, Jon71, Jon86, KMG^{+}93, Kan87, KM73, Key96, Key97, Kir93, Kir73, KC99, Koe92a, KB78, Kri87, Kue72, Lan71, LS80a, LUH^{+}96, LK91, Lev76b, LT74, LS71, Liu96, Los97b, LG75, Lyn90, Lyn91b, Mal74, Man82b, MR81a, MR81b, Mar85, Mat82, McG73, McG76, McG84, ME98, MWR72, Miz98, Mog75, Mon94, Moo72, NE90, NT85a, NM74, Par97, PM00, PW91, Pet96b, Pet96c, Rad82a, RBR80, Reg88, Ric70, Rob75, Ros83, Rou75, Rou79, Rou92b, Sar71, Sav97b, Sch73a, Shn86, Smi84a, Smi00a, SNG00, Stu72, SWB79, TS96, Und96, VA73, Voo80, Wal75, Wal86b, Wei94b, Wel80c, Wel87, Wil86, Wil94, Wil95a, Wol71, Woo87b, Yag87, YR98, YM84, Yar97, dR90, Ere80, Kaz79, LAT79, Pai93a, Rei75, Cri81, Hay95, Aba86, Ban98]. systems [Cor94a, Dog98, Goo99, Har93a, Lee96a, Lee96b, Tes86, Wil94, Zvi86]. Szigethy [Har75c].

T [Dav79b, Kra79, Lan80, War79, Whi99b, San74]. table [Bra70]. tabulations [AN73]. tactics [Bat79b, HY93, Bat79a]. Tague [Nel97b]. Tague-Sutcliffe [Nel97b]. taken [FB78]. tale [Bur00, ZK99b]. tangled [Bur92a]. target [LP99]. Task [Til70, WSQ94, BF91b, DW98, HPH96, Kie98, Wil98a, ZK72, Ano72a]. Task-oriented [WSQ94, HPH96]. tasks [Bil00, OO97]. Tauber [Har75c]. taxonomies [Gre96]. taxonomy [SK97b]. Taylor [Kaz77, Van99]. Teacher [Dav96a]. Teaching [Coo85, Flo72, Afs98, Bur96, DMZL82, Fis79, Mat97a, Wit96a, Wit96b, Bot73, Woo84]. team [Har87]. teams [CMB^{+}95]. teamwork [Cit98]. tech [Boy86b, Cob96, DH97b, Fro96, Fu96a, Fu96b]. Technical [Bra87, Dou72, HK79, Har75a, MH87, Sch73b, Ste77, Tho82, Tsu79, Woo77b, Ain84, And78, Ber71, Blu89a, Blu89b, Bol92, Col81, DDC78, Dos85, EN98, Haw74, Keg70, McC90, Men92, Now81, Shi85, Tau74, Tho73a, TB93, Wei76, Rus87]. technicians [Wei94b]. technique [DC80, Ger78, MN83b, Rou88, Sta75b, Ros73]. Techniques [Kil97, McI97, Wei97b, Wei97c, AL94, And91, AL84, Bar98b, BBDH00, CHSS98, DC76, Dil97a, Koc83, LL85, Lar92b, LS84b, Rum90, TC00a, Wis00, Mar82a, Swa73, Art75b, Str78, Smi00a, Boy87a, Gre96]. Technological [DK90, Fau96, Hol96, Sch98b, TL98, DC98, Har70d, Hei80, Hei91, Kir73, KM95, Lan87a, LH97, Mar99, Men92, NK89, WV97, PW88]. technologically [McD99]. technologies [Bro86a, Doc91, HC94, Hol96, Lan89a, Lyn91a, Mar94, Mar99, Pai93b, Rei84, Smi96a, Smi96b, Wil92b, Die87]. Technology [Abe90, Adl92, Bea92, Cim95, Cox92, CS90, DG95, Gar92, Gar90, Gor95, Gre92, Hei88, HH95a, Her95, Hur88, JS95, Koc81, Leh92, LC90, LS92b, LH95b, MWH^{+}95, Men92, Mol92, Rod95, SMJ95, Tom92, WC92, Ano81k, Art75b, Bag81, Bar96, Bec84, Ber94, Bou80, Bro96, Cal89a, Coo82, CR80, DH97b, Dou72, Eps80, Ere81c, FBC79, GP83, Glu95a, GSK98, Gul87, Hei95c, Hei86b, JBS^{+}77, KM95, Klo89, KS98, Kuc76a, LMSC82, Lan94, Mar83b, Mar98, Mat97b, Mea72, Mil90, Moo73, Myb98a, Nor96, PKK74, Pet99, Phi92, RBR80, Ric87, RR77, Rya93, SW91, She86, Sla85, Smi74a, Sta75a, SS95, Tag75, Van80, WD97a, Wal00b, Wat98, Whi98a, Wit81, Kun82, Mar82b, Par83a, You75, Swa75a, Buc98, Lam99, Mat97a, Abe96b, Han96]. technology-principles [Art75b]. Tedd [Whi00]. Telecommunications [Ber86b, Dav86, Dic86, Gib86, LL86, Mea86, She86, Rob96a, Whi98a]. telecommuting [MBS98]. telephone [SB94]. teleprocessing [Gib86].

Television [Gar86]. TELLTALE [PN96]. Tempe [Lev76b]. TemplateFiller [SMHC93]. Ten [Sie90, Bla96, Mea94, Bur81b]. Tenn [WH70]. tenure [CO95]. Term [Cag70b, Cag70a, AM70, BSR98, CO87, Cro86, DC80, Haa97, Har90, IS98, Kuc76b, Lar91, PW91, RW00, SYY75, SWY81, Sha95, SJ72, Sve72, Vin74, WY92, WY93, Zhi97]. term-by-term [WY93]. Term-term [Cag70b, Cag70a]. term/document [CO87]. Terminal [Ros73]. terminologia [Gro76b]. terminologie [Gro76b]. terminology [Ham00, Mor78, RTH71, Swa72, Gro76b]. Terms [MBFK76, AKE94, Ano70a, BMFK76, Coh95, Col93, DF84, LMM^{+}73, LK97, LH95a, MDS70, NT85b, RJ76, Ros71b, SS97, TSK71, Tag72, YR77, vdVM99]. Terse [Ber70, Ber70]. Test [CC79, Lon92, BP70, Bro93b, Bur94, Cha86, DW96, EP97, Gel79, Lou92a, MM87, Moo72, Per85]. Testing [BW84, Dor88, Her82, KL98a, Sun99, Ben80, FB81, LUH^{+}96, OG93, Sar71, Sha96, Wil92a, Wil93a]. Tests [GWDS86, Pit75, Sha96]. Text [CL82a, LNL99, Mac91, O’C73, PZ84, BYN00, BI97, BMNM^{+}93, BW84, Boy90a, Buc91c, CT73, Car75, Cas91, CL89, CYW99, CWBL00, CCW84, CBM^{+}94, DF84, EFW96, Fid91b, GD90, GFHR97, HST96, HC90, HAA87, Jak88, KSDR90, Kil93b, KT90, Lan98a, Lan98b, LBBR95, Mal74, Mar89, MKB78, MSJM91, MHCW89, Mus86, NT96, O’C70, O’C80, Pao78, Par00, PZ83, Ro88a, Ro88b, RW82, RSW91, Sal72c, SYY75, Sal79a, Sch90, SMHC93, SMJ95, Sie96, Str74, TBCE94, VS87, Wei94a, Wei96a, Wis99, WP94, WNAR99, WT93, WT95, YW96, Lin70, Dat79, Dav74, Bla93, Les89]. TEXT-PAC [Lin70]. text-processing [VS87]. text-searching [O’C70]. textbook [HH95b, HH95c]. Textile [And91]. texts [BSY00, CELY80, FK99, GS93c, KPZ^{+}99, LBBR96, Leh72, SSD99]. textual [DGS86, HAA87, SW92, PW92]. textuality [Bro95b]. texture [MM98]. Their [Har70b, JJ75, TL74, Aga99, Ali90, Art73, BCB95, Bel85, Ber71, Bro86a, Kin76, KG72, Liu96, MB86, OV98, Sch00a, Sun99, Wel83, WW73, And78, MF88]. them [KRD95]. Thematic [Cox83, Kuc82, San83]. Theoretic [SWT^{+}79, BS75, Coo79, GL91, Rub92, Wil79, WY92]. Theoretical [Fed82a, Kim73, Myb99, AM70, BS91b, Bie95, CR00, Che98a, Cur93, Pet96b, Pet96c, SK97a, Sla75, Tem94, YFR81a]. theories [Liu97c, McG96a]. Theory [Bir97, Dro98, Jen95b, Kra74a, Liu98, Bar78, Bel73, BA74, Boy92, Bra70, Cha86, Cha91, Cha99b, CV80, CM00, Egg90b, Egg94, FNNW74, Fug85, HS78, KZC86, Kir93, Kra73, KB78, Kwo85, Lam00a, LR70, Lan82b, Liu96, Lyn77, MM87, Mai99, MB91, Nic00, Par94, Par83b, Per92, Pul74, RFK97, Sac76, SYY75, Spa99, Spi94, TO84, Tzv90, Und96, Wal98, War91, Wel80c, WK75, de 76, Cas78, YFR81a, YFR81b, YFR81c, Lid96]. there [Mol92, NO85]. thermodynamics [Hei95b]. thesauri [Kir81, Lan81b, Lea75, RCZS99, Wal75, WVE85, Boy90b, Lan81b]. Thesaurus [MDS70, Wei95b, CN95, CYFS95, JGR^{+}95, KH80, Kim73, MM98, RB89, Soe72, WEWP97, Wel72, RTM99, Wei95b, WH70]. thesaurus-construction [Kim73]. thesaurus-updating [Kim73]. Theses [CH73, Hin70]. thing [Buc91a, Buc91d, Yer91]. things [Bro00c]. thinking [Bur96, DMZL82, Sar75, Cob96]. Third [Har77a, Sla85, Spi97a]. Thirty [Eli87, Kuc76a, Sal87b]. Thirty-eight [Eli87]. Thomas [Bas74, Lar74, Woo77b, Bez82]. Thompson [Har76b]. Thorofare [Sel71]. those [Ste85]. thought [Col96]. Thoughts [Sal72d, Wal87, Bla96]. threaded [Hol70]. threaded-file [Hol70]. Three [Var78, Abe96a, DCD81, HM70, Lee00b, Ros79, SBPP89, ABLS88]. threshold [DPD99, Goe87, GV87, Kra78]. thresholds [Sha85]. throughout [Wel78]. TI [Mim85]. tie [Tau74]. Tiered [Cro94]. Time [Bry70, Ora00, Sch00a, Coo83a, Coo91b, EM94, ER95, Hig99, Hub98, KM00b, MN81, SKB99, Spe71, Sol97a]. time-dependent [ER95]. Time-line [Sch00a]. Time-Sharing [Bry70]. times [Cam96a, Hei95c, Sch96a]. timing [Bon79, Sol97a]. Timothy [Fis81a]. TIP [Mat70]. TISA [Pre71b]. Title [Bot70, BL72, Mal74, BK79, Coo70b, Dro78, HH70, Jen91, Kil95b, KMB99, KM00a, Kuc78, SJ72, Whi91, WH91, WL81]. title/abstract [Mal74]. titles [Byr75, Cra82, Die84b, HZ80, He98, MS76, Mil71, Pit82, Toc70, Yer71a]. Titular [Per85]. Tocci [Har83]. today [Bro94, Bue86b]. together [PBL^{+}00, Var99]. token [CL89]. Tom [Cal00]. tomography [KET99]. tomorrow [Myb98a]. too [Chr95]. tool [CB89, JM97, Lyn90, Nic00, SD96, Har78a]. Toolbox [Bro90a, McL90a]. tools [BR95a, OMY^{+}89, TC00a, Her97]. top [Law76]. Topic [Lie94, Osb98, Ano00a, Ano00c, Ano00b, BB96a, Bro99a, CT99, Che00b, Che00c, Fro94, Glu94, Gre95, Ror99b, Pai93a]. topic-document [Ror99b]. Topical [Gre95, GB95]. topicality [Jan94]. topics [Kra94b, RW00, TSK71, YG90]. Topological [ER98]. Topology [EC92, ER97, Swa77a]. TOSAR [FNNW74]. Total [Lev98, Flo99b, KS70]. tour [Bat99c]. Townley [Cri81]. toxicity [Mon73]. TOXLINE [Sch81]. Tracking [Hen88]. Tracy [Mea82a]. trade [GK89, Mea95]. trade-off [GK89]. tradition [SM88]. traditional [PPP98, Tra98]. traditions [BNP99]. traffic [BLPG99]. trails [Pai90]. trainers [Gro90]. training [Bel88, BR80a, Dun80, Jac72, Mog75, Rob80b]. traits [Bel85]. transborder [Vei80]. Transfer [ER91, Ere81c, Bri84, Dos85, JBS^{+}77, KW81, LMSC82, Lon90, Men92, NKB72, Pie99, Wit80, Woo77a, Tsu79]. Transferring [Hal85]. transformation [Afi98, Boy90a, Bur00, Hur00b]. Transforming [Bac93]. Transition [AL96, AKS96, Abe96a, BCB96, Hay96b, Cha81, Pao78, ZPB98]. transitions [WV97]. Translated [Lev76a]. translating [MR81a, MR81b, Old91]. translation [Die84a, He98, He00, IS98, JBH77, Liu97c]. Translations [Now81, And78]. Translators [Sel71, Mor78]. Translingual [He00]. transliteration [Bla71]. Transmission [Tho81, dR90]. Transnational [Die79, ZPB98]. Transparent [Wil86]. travelers [Lyt88]. Travica [Kat00]. treatise [Cob96, FNNW74]. treatment [CVK92]. TREC [BRR96, Ror99b, Ror99a]. trees [KT90]. Trends [Bor84, Dav79b, GSD74, Sal86, She86, Bel88, Bru78, Fre80, Haw90, Lar91, Lee00b, Lil78, Wil77b, Mar98, SKB99, Lil78]. trial [Swa89b]. trial-and-error [Swa89b]. tribute [Wei97a]. Tropes [Day00b]. trouble [Cas96]. True [Yov70, Cob96]. truncated [Ras80, SR77]. truncation [Kje74, LJ70a]. truth [Mur93]. Trygve [Ere80]. tuples [BMSM90]. Turner [Lun80]. Tutorial [Rou79, YM84]. Twenty [Wil98b, Mil90, Ric87, Sut96]. Twenty-First [Wil98b, Mil90, Sut96]. twist [ABLS88]. Two [Blu71, CO84, Mcg79, Roy90a, SHS72, Swa87, AC93, Bod99, Boo81, Bur00, BNP99, CO85, CCN93, Dal90, GH96, Hub78, Hur80, LS70, Lun76, MN83b, Pai93a, Ros71a, SW92, Sch82, Sma73, Woo88a, Yu76, ZK99b, ZvD00, All00, Fau96, Sri90]. two-experiment [CCN93]. two-factor [ZvD00]. two-in-one [Lun76]. two-level [LS70, SW92]. Two-Poisson [Sri90]. two-stage [MN83b]. two-way [Woo88a]. type [CO84, CO85, CL89, KG72, Law74, May97, San99]. type-token [CL89]. typed [BTSKC96]. Types [QLA97, Buc79, Eas91, Fen81, PM00]. Typologies [Gre96]. typology [GB95, Swa78a]. Tyranny [Art82].

U.S [Ere80, Gar86, Pic86, Pre71b, Shu80, Til70]. U.S. [Ber86b, Bre79, FS97b, Ken87a, Koh74, MN83a, MK72, ST89, Ste00a]. UDC [Wel70]. UFO [VD98]. UL [OG70]. UL/1 [OG70]. Ulrich [Gro76b]. Unanswered [Sve86]. uncertainties [RR78]. uncertainty [AdKK87, CMK^{+}98a, Col93, HMNW86, Kuh99, SLC98, VD98, Tzv90]. Uncitedness [Ste90]. uncontrolled [LFS72]. Unconventional [DH78, Haw77]. underlying [Der94, Goo97, CM84]. Understanding [Day99b, Glu95b, Col99, For99, Leh92, Lie94, Ano75a, Myb99]. undiscovered [Spa97]. uneasy [Sch96a]. UNESCO [Wei76, Gro76b, Wel78]. unformatted [Eas91]. unification [Bod99, Hei95b]. unified [Dom00, Goe87, GV87, Boo90a, WNAR99]. unimodal [Sic92]. uninhibited [Kay98]. Union [Sch73b, SR77, Kir73]. Unions [Lan80]. unique [Boo74b]. uniquely [DG00]. unisist [Sal72d, Yer71b]. unit [Col92, Spe71]. United [Art74, Bol92, Dro78, Nel97a, Rus77, Rus85, Shi86a, Wel75, Mor78, Ste80, Wil97, Woo84, Bax85, Gar92, Mea72, War99b, Wil82, Bol92]. Unitized [FCJ^{+}92]. universal [Soe72, McI97, Kuc76b]. universe [Hjø99, JJ77, JJ80, Nei80]. Universidad [Har70a]. Universitets [Ere80]. Universities [Sha84, ANM78, CH99, Ehi90, MP98, Ric87]. University [And91, Blu71, Die81, Die82a, Don77, Ere80, Gro75b, Har75c, Har76a, Hin70, Kaz77, Lan77a, Lil78, Old91, Ric77, Rou76a, SM88, Swa76b, Swa77b, Var75a, Wes00, Woo88b, Woo77b, You75, Cov00, Cox93, Hol96, Kan76a, Lan94, RB72, SKvB92, SvB94, Tur77, VHdB^{+}98, Xia92, BR94, BDBR97, BD97, BC97, Dou96, SOM^{+}99, Sut96, War88, Ols76, Shi86b, Pel93]. unlimited [Bro93b]. unnoticed [Swa89a]. unprecedented [Mar85]. unpublished [Win70a]. untrained [Wil98a]. Unused [Wil95b, MZ78]. updating [Kim73]. upon [Pap93, Pap98]. upper [FS84]. upper-lower [FS84]. Urbana [Har76a, Lil78, Ric77, Swa76b, Swa77b]. Urbana-Champaign [Har76a, Ric77, Swa76b, Swa77b]. urn [KH95, de 78]. USA [Mea82b]. Usability [Par00, Cas96, Cur97, EK97, HJD00, Sha96]. usable [Boy86c, MLC80]. Usage [Coo83b, Haw74, LH95a]. Use [Ama84, BMFK76, But99, CT73, Car75, Cor99, Cre99b, Cre99a, CH99, Day99a, DF82, Fra99, Gas99, Ger79, Lan77a, Lan80, Lan77b, Lev99, LCB99, Mar80, MBFK76, Mor76, Nah98, Pre71c, RM99, Rou76a, SRS98, Ste99b, Sun99, Swa77b, Wat98, ALD96, Aga99, Ami70, Bai74, Bai76, BK96, BT84, BE80, Bil00, BNS^{+}00, Bor86b, Bor96, Boy86e, Boy86a, CO87, Caw74, Che72, Col81, CC96, Coo85, Coo91b, CK96, DGS86, Egg88, Elc73, Hal81, HZ80, HW98, Hel71, Hen99, Hoo91, HC84, Jai72, KB98b, KZC86, KC99, KRD95, Lar91, LS92a, LBIP97, LAD97, MB81a, Mar92b, MRY73, MB81b, Neu95, New94, Pal99a, Pit75, Raf75, RB83, Rou92b, Rya93, San74, SMB72, SD96, SJ00, Smi81, Spe80, SMC83, TC00a, TT79, Tsa98, Tur77, VA73, Wal81, Wal85, WS98]. use [WW99, WdBFF95, Wil84, Wol86, YF85, YM99, Par82b, Dil97a]. used [Boy90b, Con95, Mic94, Ros71b, Ste85]. Usefulness [Cas96, Coo70b, Keg70]. Usenet [Myb98b, Osb98, Kay98]. User [Bar94, Bru98, Coo91c, Dio94, Fid94, Gor91, Gou74, Har92a, Her96b, Kuc75, Lan80, Mea83c, Mim85, Mog75, OOA99, Ros73, TC00b, AM77, All00, And91, AL84, Bal00, Bea92, BCM89, Bor94, Boy86c, Bru94, CK99, Car88, Chr72, Col00a, CM00, CC98, Cur97, Deb86, Eft00, EB88, FHN^{+}93, GQSdP98, Han86, Har94, HC96, HCL^{+}00, Jan91, Jan94, Kat72, KM73, KS90, Kri86c, Kri89, Kuh91, LC83, Mar92a, Mar95, MR81a, MR81b, MCT73, MWY95, MLC80, Mor94, Oja86, Par94, Par00, PJ90, RFK97, Rob00, RW00, SW91, ST86, Sie86, SBW93, Soe76, Spi96, Su94, SEW00, Tho73a, Voo99, Yao95, Yee91, Yu74, Yua97, YM99, CHSS98, Don76, HW71, Yee91, Kie98, McL90b]. User-based [Gor91, Eft00, Par94]. User-centered [Fid94, All00, Bal00, Bru94, GQSdP98, Mor94, RFK97]. User-defined [Bar94]. User-friendly [Har92a]. User-oriented [McL90b]. Users [FHN^{+}93, Har79, Der84, Dil00, Fen81, Goo97, GLS97, Kan87, KMB99, LMSC82, LB00, Laz00, LBW00, LS97, Mea00, NT96, PM00, Pen70, Sha90, SF78, Sta85, SBW90, Tag72, SK88a]. uses [DH78, Haw77, Lun76]. Using [Bur90, CM00, Cro79, Dav76a, Fau96, FW90, FPS00, GO93, GD98, GLW86, Hol96, Man82a, Mat70, OO97, San79a, SC78, YW96, Zha00, AB76, BSY00, BCB95, BCB98, Ber94, BM89, CPW88, Cha88, CSSI98, CCW84, CELY80, CL82a, CBM^{+}94, Cra00, Cro77, EH94, Fri97, FNNW74, Gor90, HC90, HM70, Hoc89, Koc83, KET99, Ley98, LY98, LY91, Mar89, MPW73, MK72, MQP98, OWF99, OMY^{+}89, Oh98, OG70, PT87, Ros71a, SR77, TT79, Wat98, WBX98, WLC00, Yan97, Rob91, Jen95b]. Usoe [War79]. UT [Don77, Har78a]. Utah [War88]. Utility [Coo79, Wil79, Art73, BA74, Bro70c, Can89, FGS81, GL91, HS78, Jan94, Kan76a, Par82a, Har70b]. Utility-theoretic [Coo79, Wil79]. Utilization [Bac93, Kir93, LS93, Pai93b, Tod99, ER00a, Hoe72, KdK90, Mar72c, Tag75, War73]. Utilizing [WA79, DCFQ00, YK93].

V [Har70b, Rya80, Dav79b, Gar88, Wol87]. valid [McC86]. Validating [ZC00]. Validation [Gar97, LS92a]. Validity [LB83, Bro83, Eve93, Gle93, Nic89]. valuable [Swa89c]. Value [Ozo79a, Pet96a, Tay82, AKS96, Gri80, Gro96, Her99, Hub98, McM73, Mow92a, Mow92b, Mow92c, Ozo80, Rep89, SK97a, SK97b, Str84, YF85]. Value-Added [Tay82]. values [CO87, MOV83, OV82, SLC98]. Vannevar [Buc92b, Bur92a, NK89]. variable [CL82a]. variable-to [CL82a]. variables [Rob75, YM99]. Variation [MM78, Haa97]. Variations [Har96, Coo83a, KS77, Tag78]. Variety [Lyn77, CELY80, Mat82]. various [LBIP97, Rob00]. varying [EM94]. Veaner [Hir90]. Vector [BSR94, FW90, RW86, Liu97b]. vehicles [FW99]. Veit [Ano77a]. Veith [Nov82]. vendors [Abe96a, Ros79]. verb [JD78]. verbal [Lev76a]. verbatim [Mea82b]. Verification [Swa89d]. Verlag [Har76b, Har77b, Kuc76b, Lev76a, Wel71]. versatile [SGS70]. Versig [Gro76b]. version [RMMS96, Sie89a]. versions [Rou90]. versus [ANM78, Bur80, EH94, Gor82, Hei94, HC84, Koe83, KS90, Lan98a, Lan98b, Mal74, MSPPM98, Mar82a, MS00, Mud98]. Vertebrate [Ste85]. Vertical [EQH00]. very [JBS^{+}77, KSDR90, Lyn88]. via [BJ79, He98, Pep74, Ros79, SH96]. viable [dR90]. Victoria [Gul85]. video [FS84, Klo89, WH98, Abe96b]. Videodisc [Mar82b, SL83, GP83, Kea83, Par83a, Bro86a]. videodisc/optical [GP83]. videodiscs [SB83]. videotex [Cas94]. view [Bec84, Bru94, Cas98a, Coo82, Har74, Kar98, Kra73, Lin87a, Mol92, Pul74, SOM^{+}99, War99a, Kra74a, Van90]. viewpoint [Ben00]. Views [Bol92, Dic86, Oja86, Mec99]. Vincent [McM76]. Vinogradov [Rya80]. Virtual [Nug91, CCM00, Che00a, Des97, For00, McD99, PBL^{+}00, Sem98b, Sha98, Wat99, Sha98]. virtue [Exn97]. visible [Gre92]. Vision [KBN90, BR94, Koe92a]. Visions [Ray94]. visit [Ain84, FDD^{+}99, PBL^{+}00]. Visiting [PBL^{+}00]. Visual [BC99, Mur93, RCZS99, RL99, RTM99, RH99, Sma99, GMPS00, Jen91, Ror93, Ror99b, Ror99a, RF00, Sve94, Ano00a, Sko98]. Visualization [RH99, Bro00b, FFKS99, Wis99, RL99]. Visualizing [LDS^{+}00, Sma99, WM98]. Vito [Mar75]. Vleeschauer [Don77]. vocabularies [Cha88, Cor99, Sve86]. Vocabulary [Cim95, Nie76, Aul71, Bal88, BM89, CSI87, CNMS97, Fid91b, GLL90, HC96, LFS72, Mud98, PN98, Qin00b, Sal72a, Boy87b, Lan93]. Voight [Pre70, Swa77c]. Voigt [Mar77b, Mig80a]. Vol [Ano74a, Dan84, Dro77, Han75, Lil78, Rei85]. Volume [Bix96a, Bix96b, Blu89a, Blu89b, Cam97, Dav79b, Mar77b, McM76, Mig80a, Pre70, Swa77c, HHC70, KS70, Kra97, Kri87]. Vosper [Har77b]. vs [CN95, Wel80a].

W [Ano74b, DC76, Fly97, Har79, Kaz75a, Lan73, Mai99, Mar80, Rob98, Sal72a, Swa78b, Tau74, Woo87a]. W.O. [Old88]. WAIS [MBH94]. Wales [Var75a, Var75a]. Walker [Lam00a]. Walling [Har75d]. want [GLL90, Hei90, Pre88, Zwe95]. wanted [Whi99a]. War [FS97b, Ric88, Mat97b]. Ware [Bro00b]. Warholian [Cro99a]. warning [JBS^{+}77]. warrant [Beg95]. Warriors [Cas98b]. wartime [Ric88, Tay95]. Was [BHCC^{+}74]. Washington [Don74, Gan74, Har77b, Hin70, Lan73, Lar74, Ols74, Pre71b, Rus85, Sal72a, Til70]. wasteland [Ain84]. Water [BN76]. watercraft [SS89]. way [Aul79, Lar90, Woo88a, Wri79]. Wayne [Har79]. ways [MBH94]. Web [Cal00, Cam97, Rap00, Sch98a, Ste00a, Bro00d, Coo98, DCFQ00, Fly97, HG00, Koe99, LB00, Nic00, Whi99b, DG00, Ano00a, Ano00b, BI00, BT00, BBFM99, Bil00, BLPG99, Cam97, CCRY98, CSR^{+}98, DPD99, FDD^{+}99, GCS97, HF00, KL98b, KL98c, Laz00, LBW00, LS99, Mar00, McC00, Mea00, PK00, PPP98, Rod95, RNBV98, Sch98b, SFB^{+}00, Bro00d, Nor98]. web-based [Coo98, HF00, RNBV98, SFB^{+}00]. website [ZvD00]. Weddington [Lar74]. weeding [Wal86a]. weekly [Gue78]. weighted [Ang75, BSR98, Boo78, Boo80, Boo81, BK81, GV87, Yag87]. weighting [DM99, RJ76]. Weiner [Dur76]. Weisman [New73]. well [Ber71, Chr95, CC98, SM91]. well-designed [CC98]. Wellisch [Fid87]. Wells [Ray99]. Weltner [Lev76a]. Wesley [Har79]. Wessel [Lan77b]. West [Die82b]. Western [Die81, Har77a, Kaz75b]. Westfalen [Wel71]. Westport [Ano77a, Ere81b, Har77c, Koe75, Kun82, Ols76, Shu78]. Westview [Cri81, Lun80]. whats [Gli74]. Whence [Sha97a]. Where [BHCC^{+}74, Pee94, Pai90, Gol87]. which [Sic85]. White [Don70, Mat85]. Whitfield [Rei75]. Whither [Rom91, Sha97a]. Whitney [Ano72e]. Whitworth [Gel79]. Who [Leh72, CD97, Car95b]. whole [Bor94, Mea98, Miz97]. whom [CD97]. whys [Gli74]. Wide [BT00, Cam97, DPD99, GCS97, KL98b, KL98c, Laz00, LBW00, LS99, Mar00, McC00, Mea00, Rod95, Coo82]. wide-angle [Coo82]. Wiley [Ano74b, Fis79, Her82, Kra79, Lan82a, Lan77b, Mea83b, New73, Sie70, Sta77a, Wri71a]. Wiley-Becker [Ano74b]. Wiley-Interscience [Kra79, Lan77b, Sie70, Wri71a]. Wilfrid [Her82, Swa76b, Swa77b]. will [Kra97, Sem98a]. William [Boy85b, Hei78, Kra79, Lil78]. Williams [Art75a]. Wilson [Ber84, Shu78, Coo79]. window [Gar86]. Windsor [Opp74]. winner [Coc98]. Winthrop [Dav76a]. wiretapping [Pet98b]. Wisconsin [Han75]. within [ABB^{+}91a, ABB^{+}91b, And99, BS91b, BF91a, CP98, CS91, DRCG91, LD91, McD99, RRB91, SR79, Smi81, SB91, War73, Zmu75]. without [ENN98, Cox93]. Wladimir [Woo77b]. Women [Gro75b]. Woods [Cal00]. Woolridge [Che76]. Word [Kun82, BMvR91a, BMvR91b, CL90, CYW99, CMK98b, Egg99, FWR95, Har93b, KMB99, KM00a, LJ70a, LY91, MDS70, Pao78, SW91, TZ92, Mea82b, Rol87]. word-based [CYW99]. word-frequency [LJ70a]. Words [DF82, Har75a, ASAOE99, AYS96, AKE94, BKR98, Bro95b, GLL90, Jon71, Kil95b, Ley97, Ley98, Sav93, SSD99, War96a, War96b, YW96]. Wordtree [Moo86]. Work [HW98, Ano08, Boy86a, BF93, Car96, Chr95, GdCD81, GL96, Gro78, HC94, Hoe72, Jac72, KWR98, SL99, WKMG00, War93, Wil00a, Wil95a, Bot73, Swa76d, Fro96]. workable [BF93]. worker [Kuh99]. workers [Wil98a]. Working [Til70, Nan95b]. workplace [Sem98b]. Works [Bon82, Kil93a, Sic85]. Workstation [BR95a, Chr95, Fag95, Men95a, Mes95, O’H95, Wig95]. workstations [Chr95]. World [Cam97, Cer83, TC73, Cha91, Cha96, Gar86, KS97, Mea78b, PBL^{+}00, Sla85, Wel78, BT00, DPD99, FS97b, GCS97, Har77a, KL98b, KL98c, Laz00, LBW00, LS99, Mar00, McC00, Mea00, Ray99, Ric88, Rod95, Die81, Cam97, Wei76, Bat97, Cre95a]. Worlds [Bel71]. worldwide [Pel93, Yer71b]. Worm [CNMS97]. Wright [Don77]. write [Boy86c]. Writing [Che97, KM78, Rad89, Ric95, War93, Day97, Hei95d, KW97, Lit96, Tag78, Wei70]. Written [TSK71, Keg70]. wrong [AEP00, Aul79, Fra99, Wri79]. wrongs [Ric95]. Wuerttemberg [GSK98]. WWW [BBDH00, CK99].

X [Gro75b, Har77b, Koc79, Lan77a]. X1iii [Lea75]. Xanadu [Ray94]. xii [Kuc75]. xiv [Fis81a, Fis81b]. XML [Ano00c, Rap00]. xvi [New73]. xvii [Nov82, Vag80]. xviii [Lev76b].

Yahooligans [Bil00]. Year [KJR99, BWKG75, Kri77, Var78]. Year-2000 [KJR99]. years [Bla96, Bur81b, Eli87, Kra88b, Kra92b, Lin87a, Mea94, Sal87b, ST86, Sie90, Var99, Ric87]. Yeh [Dav79b]. Yes [BF93]. yield [Has76]. York [Ano74a, Bai75, Bas74, Bel76, Boy85b, Che76, Dan84, Dav77b, DC76, Die81, Die82a, Don76, Dro78, Dur76, Fis79, Har70d, Har71b, Her82, Koc79, Lan82a, Lan76, Lan77b, Lev76a, Mar77b, McM76, Mim85, New73, Pat85, Pre71c, Pre70, Rei85, Shi85, Sie70, Sta75a, Sta77a, Str76, Swa75a, Swa77c, Tsu79, Vag80, War79, Wel80b, Wri71a]. Young [Don77]. Yourdon [Fis81a]. youth [CT99, Dre99]. Yuang [Boy95a]. Yule [CL89]. Yuri [Wei97a].

Z39.4 [And94]. Zealand [Jas94]. Zenk [Ere81b]. Zipf [CL86, Egg99, Fed82a, KG72, Nic87]. Zipf-type [KG72]. Zipfian [Fed82b]. zooming [CD97].

References

Allen:1976:LLP

[AB76] Albert H. Allen and Eugene F. Beirne. On-line logging in of periodicals by CODEN using interactive query report processor. Journal of the American Society for Information Science, 27(4): 230–234, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abarbanel:1986:BRB

[Aba86] Andrew Abarbanel. Book review: Adaptive control of ill-defined systems. Journal of the American Society for Information Science, 37(3):171, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Adams:1991:PLIa

[ABB^{+}91a] Carl Adams, Cynthia Mathis Beath, Richard Bolan, Joseph Branin, George D’Elia, Nancy Freeman Rohde, and Detmar Straub. Perspectives on … integrated information centers within academic environments: A model academic Integrated Information Center. Journal of the American Society for Information Science, 42(2):143–146, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Adams:1991:PLIb

[ABB^{+}91b] Carl Adams, Cynthia Mathis Beath, Richard Bolan, Joseph Branin, George D’Elia, Nancy Freeman Rohde, and Detmar Straub. Perspectives on … integrated information centers within academic environments: Issues relating to the implementation and management of an Integrated Information Center. Journal of the American Society for Information Science, 42(2):147–151, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abelson:1990:PLE

[Abe90] Philip Abelson. Perspectives on … evaluation of scientific information: Peer review and the impact of new information technology: Mechanisms for evaluating scientific information and the role of peer review. Journal of the American Society for Information Science, 41(3):216–222, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abels:1996:PLCc

[Abe96a] Eileen G. Abels. Perspectives on … costs and pricing of library and information services in transition: Pricing of electronic resources: Interviews with three vendors. Journal of the American Society for Information Science, 47(3):235–246, March 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abeysundara:1996:BRB

[Abe96b] Bandula Abeysundara. Book review: Video dialtone technology. Journal of the American Society for Information Science, 47(6): 477–478A, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anderson:1988:PLE

[ABLS88] James D. Anderson, Nicholas J. Belkin, Linda C. Lederman, and Tefko Saracevic. Perspectives on … education of the information professional: New dimensions, new directions: Part II. Three schools with a new twist. Information science at Rutgers: Establishing new interdisciplinary connections. Journal of the American Society for Information Science, 39(5):327–330, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Auster:1993:RES

[AC93] Ethel Auster and Chun Wei Choo. Research: Environmental scanning by CEOs in two Canadian industries. Journal of the American Society for Information Science, 44(4):194–203, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Avedon:1979:MPA

[AD79] Don M. Avedon and Ann M. Devilliers. Microfilm permanence and archival quality. Journal of the American Society for Information Science, 30(2):100–102, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Alo:1987:RED

[AdKK87] Richard A. Alò, André de Korvin, and Robert Kleyle. Research: Emergence of a dominant course of action in a general feedback loop when goal uncertainty is present. Journal of the American Society for Information Science, 38(2):111–117, March 1987. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Adler:1992:PLI

[Adl92] Sharon Adler. Perspectives on … information technology standards: The birth of a standard. Journal of the American Society for Information Science, 43(8):556–558, September 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Adler:1997:AD

[Adl97] Sharon C. Adler. The “ABCs” of DSSSL. Journal of the American Society for Information Science, 48(7):597–602, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Alvarez:2000:WWO

[AEP00] Pedro Alvarez, Isabel Escalona, and Antonio Pulgarín. What is wrong with obsolescence? Journal of the American Society for Information Science, 51(9):812–815, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Al-Fares:1998:ROL

[AF98] Hesham K. Al-Fares. Research: Optimizing a library’s loan policy: an integer programming approach. Journal of the American Society for Information Science, 49(13):1169–1176, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Afifi:1998:BRB

[Afi98] Marianne Afifi. Book review: From print to electronic: the transformation of scientific communication. Journal of the American Society for Information Science, 49(7):670, May 15, 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Afifi:1999:BRB

[Afi99] Marianne Afifi. Book review: Electronic databases and publishing, edited by Albert Henderson. Journal of the American Society for Information Science, 50(8):725, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Afshar:1998:BRB

[Afs98] Ebrahim Afshar. Book review: Gateways to knowledge: the role of academic libraries in teaching, learning, and research. Journal of the American Society for Information Science, 49(2):190–192, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Afshar:1999:BRB

[Afs99] Ebrahim Afshar. Book review: Into the Future: the foundations of library and information services in the post-industrial era, by Michael Harris, Stan A. Hannah, and Pamela C. Harris. Journal of the American Society for Information Science, 50(9):839–840, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Agada:1999:RIC

[Aga99] John Agada. Research: Inner-city gatekeepers: an exploratory survey of their information use environment. Journal of the American Society for Information Science, 50(1):74–85, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Aghazadeh:1986:BRB

[Agh86] F. Aghazadeh. Book review: Ergonomic data for equipment design. Journal of the American Society for Information Science, 37(3):172–173, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ammar:1976:LEI

[AI76] Suad S. Ammar and Herbert Inhaber. Letters to the Editor: Interdisciplinarity and Canadian science. Journal of the American Society for Information Science, 27(4):264–265, July 1976. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Aiken:1997:BRB

[Aik97] Peter Aiken. Book review: Information systems development and data modeling: Conceptual and philosophical foundations. Journal of the American Society for Information Science, 48(10):959– 960, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Aines:1984:VWF

[Ain84] Andrew A. Aines. A visit to the wasteland of federal scientific and technical information policy. Journal of the American Society for Information Science, 35(3):179–184, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Aines:1985:LEN

[Ain85] Andrew A. Aines. Letters to the Editors: Note. Journal of the American Society for Information Science, 36(5):352, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Aiyepeku:1988:LEC

[Aiy88] W. Olabode Aiyepeku. Letters to the Editor: Comment on the paper entitled “Sub-Saharan Africa and the Paperless Society” (JASIS, 38(4): 298–304; 1987). Journal of the American Society for Information Science, 39(3):219–220, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Old87] and rejoinder [Old88].

Ajiferuke:1991:RPM

[Aji91] Isola Ajiferuke. Research: A probabilistic model for the distribution of authorships. Journal of the American Society for Information Science, 42(4):279–289, May 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See comment [Dav92b].

Ajiferuke:1992:LEP

[Aji92] Isola Ajiferuke. Letters to the Editor: Probabilistic model for distribution of authorships. Journal of the American Society for Information Science, 43(1):96, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Al-Kharashi:1994:RCW

[AKE94] Ibrahim A. Al-Kharashi and Martha W. Evens. Research: Comparing words, stems, and roots as index terms in an Arabic Information Retrieval system. Journal of the American Society for Information Science, 45(8):548–560, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abels:1996:PLCb

[AKS96] Eileen G. Abels, Paul B. Kantor, and Tefko Saracevic. Perspectives on … cost and pricing of library and information services in transition: Studying the cost and value of library and information services: Applying functional cost analysis to the library in transition. Journal of the American Society for Information Science, 47(3):217–227, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Auster:1984:RSI

[AL84] Ethel Auster and Stephen B. Lawton. Research: Search interview techniques and information gain as antecedents of user satisfaction with online bibliographic retrieval. Journal of the American Society for Information Science, 35(2):90–103, March 1984. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Aigrain:1994:RME

[AL94] Philippe Aigrain and Véronique Longueville. Research: A model for the evaluation of expansion techniques in information retrieval systems. Journal of the American Society for Information Science, 45(4):225–234, May 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abels:1996:PLCa

[AL96] Eileen G. Abels and Lois F. Lunin. Perspectives on … costs and pricing of library and information services in transition: Introduction and overview. Journal of the American Society for Information Science, 47(3):208–209, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ashton:1998:ERL

[AL98] Sarah Ashton and Philippa Levy. European research letters: Networked learner support in higher education: Initiatives in professional development and research for a new role. Journal of the American Society for Information Science, 49(9):850–853, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abels:1996:RFI

[ALD96] Eileen G. Abels, Peter Liebscher, and Daniel W. Denman. Research: Factors that influence the use of electronic networks by science and engineering faculty at small institutions. Part I. Queries. Journal of the American Society for Information Science, 47(2):146–158, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Algon:1995:BRB

[Alg95] Jacqueline Algon. Book review: Publishing in the information age: a new management framework for the digital era. Journal of the American Society for Information Science, 46(5):392–393, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ali:1990:RDO

[Ali90] S. Nazim Ali. Research: Databases on optical discs and their potential in developing countries. Journal of the American Society for Information Science, 41(4):238–244, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Allen:1989:RRC

[All89] Bryce Allen. Research: Recall cues in known-item retrieval. Journal of the American Society for Information Science, 40(4):246– 252, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Allen:1998:BRB

[All98] Bryce Allen. Book review: Language and space. Journal of the American Society for Information Science, 49(4):381–382, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Allen:2000:IDC

[All00] Bryce Allen. Individual differences and the conundrums of usercentered design: Two experiments. Journal of the American Society for Information Science, 51(6):508–520, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Augustson:1970:DTR

[AM70] J. Gary Augustson and Jack Minker. Deriving term relations for a corpus by graph theoretical clusters. Journal of the American Society for Information Science, 21(2):101–111, March 1970. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Allen:1977:DSD

[AM77] Sheilah Allen and Joyce Matheson. Development of a semantic differential to access users’ attitudes towards a batch mode information retrieval system (ERIC). Journal of the American Society for Information Science, 28(5):268–272, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Avram:1980:DLN

[AM80] Henriette D. Avram and Sally H. McCallum. Directions in library networking. Journal of the American Society for Information Science, 31(6):438–444, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Aman:1984:AUA

[Ama84] Mohammed M. Aman. Appilcation: Use of Arabic in computerized information interchange. Journal of the American Society for Information Science, 35(4):204–210, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Amick:1970:MSA

[Ami70] Daniel James Amick. A multivariate statistical analysis of the use of a scientific computer based current-awareness information retrieval system. Journal of the American Society for Information Science, 21(3):171–178, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Amick:1973:SEI

[Ami73] Daniel James Amick. Scientific elitism and the information system of science. Journal of the American Society for Information Science, 24(5):317–327, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ackley:1973:CTA

[AN73] Paul R. Ackley and Robert A. Nardello. Cross tabulations applicable to mini-computers. Journal of the American Society for Information Science, 24(6):437–444, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anderson:1978:AHS

[And78] James D. Anderson. Ad hoc and selective translations of scientific and technical journal articles: Their characteristics and possible predictability. Journal of the American Society for Information Science, 29(3):130–135, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anderson:1988:BRB

[And88] James D. Anderson. Book review: String indexing. Journal of the American Society for Information Science, 39(6):435, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anderson:1990:BRB

[And90] Margaret Anderson. Book review: Jesse Shera, librarianship, and information science. Journal of the American Society for Information Science, 41(3):211–212, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anderson:1991:PLI

[And91] Clarita S. Anderson. Perspectives on … imaging. Advanced applications: A user’s applications of imaging techniques: The University of Maryland Historic Textile Database. Journal of the American Society for Information Science, 42(8):597–599, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anderson:1994:PLI

[And94] James D. Anderson. Perspectives on … indexing: Standards for indexing: Revising the American National Standard Guidelines Z39.4. Journal of the American Society for Information Science, 45(8):628–636, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Andersen:1999:BRB

[And99] Jack Andersen. Book review: Scholarly book reviewing in the social sciences and humanities. The flow of ideas within and among disciplines. Journal of the American Society for Information Science, 50(13):1259–1261, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Angione:1975:EBW

[Ang75] Pauline V. Angione. On the equivalence of boolean and weighted searching based on the convertibility of query forms. Journal of the American Society for Information Science, 26(2):112–124, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anderson:1978:PRV

[ANM78] Richard C. Anderson, Francis Narin, and Paul McAllister. Publication ratings versus peer ratings of universities. Journal of the American Society for Information Science, 29(2):91–103, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:BCA

[Ano70a] Anonymous. Brief communications: Analysis of indexing terms for plastics. Journal of the American Society for Information Science, 21(2):164–166, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:BCI

[Ano70b] Anonymous. Brief communications: International high-energy physics preprint network emphasizes institutional exchange. Journal of the American Society for Information Science, 21(1): 95–97, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:E

[Ano70c] Anonymous. Errata. Journal of the American Society for Information Science, 21(3):186, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:I

[Ano70d] Anonymous. Introduction. Journal of the American Society for Information Science, 21(4):267, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:Ma

[Ano70e] Anonymous. Masthead. Journal of the American Society for Information Science, 21(1):fmi, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:Mb

[Ano70f] Anonymous. Masthead. Journal of the American Society for Information Science, 21(2):fmi, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:Mc

[Ano70g] Anonymous. Masthead. Journal of the American Society for Information Science, 21(3):fmi, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:Md

[Ano70h] Anonymous. Masthead. Journal of the American Society for Information Science, 21(4):fmi, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:Me

[Ano70i] Anonymous. Masthead. Journal of the American Society for Information Science, 21(5):fmi, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1970:Mf

[Ano70j] Anonymous. Masthead. Journal of the American Society for Information Science, 21(6):fmi, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:BCF

[Ano71a] Anonymous. Brief communications: FID Committee on Linguistics in Documentation (FID/LD). Journal of the American Society for Information Science, 22(3):216–224, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:EE

[Ano71b] Anonymous. Errata. Journal of the American Society for Information Science, 22(4):297, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:Ma

[Ano71c] Anonymous. Masthead. Journal of the American Society for Information Science, 22(1):fmi, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:Mb

[Ano71d] Anonymous. Masthead. Journal of the American Society for Information Science, 22(2):fmi, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:Mc

[Ano71e] Anonymous. Masthead. Journal of the American Society for Information Science, 22(3):fmi, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:Md

[Ano71f] Anonymous. Masthead. Journal of the American Society for Information Science, 22(4):fmi, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:Me

[Ano71g] Anonymous. Masthead. Journal of the American Society for Information Science, 22(5):fmi, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1971:Mf

[Ano71h] Anonymous. Masthead. Journal of the American Society for Information Science, 22(6):fmi, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:BCA

[Ano72a] Anonymous. Brief communications: American Institute of Biological Sciences: Task Force on Communication and Information Services in Biology. Journal of the American Society for Information Science, 23(4):280–281, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:BCI

[Ano72b] Anonymous. Brief communications: Information and documentation in Chile ( 1 ). Journal of the American Society for Information Science, 23(4):283–285, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:CAS

[Ano72c] Anonymous. Comments: ASIS/ SLA merger comments. Journal of the American Society for Information Science, 23(1):1–2, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:E

[Ano72d] Anonymous. Errata. Journal of the American Society for Information Science, 23(5):336, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:MJW

[Ano72e] Anonymous. In memoriam: James Whitney Perry (1907–1971). Journal of the American Society for Information Science, 23(1): 2, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:Ma

[Ano72f] Anonymous. Masthead. Journal of the American Society for Information Science, 23(1):fmi, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:Mb

[Ano72g] Anonymous. Masthead. Journal of the American Society for Information Science, 23(2):fmi, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:Mc

[Ano72h] Anonymous. Masthead. Journal of the American Society for Information Science, 23(3):fmi, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:Md

[Ano72i] Anonymous. Masthead. Journal of the American Society for Information Science, 23(4):fmi, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:Me

[Ano72j] Anonymous. Masthead. Journal of the American Society for Information Science, 23(5):fmi, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1972:Mf

[Ano72k] Anonymous. Masthead. Journal of the American Society for Information Science, 23(6):fmi, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:E

[Ano73a] Anonymous. Errata. Journal of the American Society for Information Science, 24(6):465, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:Ma

[Ano73b] Anonymous. Masthead. Journal of the American Society for Information Science, 24(1):fmi, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:Mb

[Ano73c] Anonymous. Masthead. Journal of the American Society for Information Science, 24(2):fmi, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:Mc

[Ano73d] Anonymous. Masthead. Journal of the American Society for Information Science, 24(3):fmi, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:Md

[Ano73e] Anonymous. Masthead. Journal of the American Society for Information Science, 24(4):fmi, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:Me

[Ano73f] Anonymous. Masthead. Journal of the American Society for Information Science, 24(5):fmi, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:Mf

[Ano73g] Anonymous. Masthead. Journal of the American Society for Information Science, 24(6):fmi, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1973:MM

[Ano73h] Anonymous. Memoriams: In memoriam. Journal of the American Society for Information Science, 24(2):i, March 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:BRBa

[Ano74a] Anonymous. Book review: Government publications review; an international journal. 1973. Quarterly. $35. Ed: Bernard M. Fry. SEBD Publications, Inc., Sciences, Engineering and Business Data, 380 Saw Mill Road, Elmsford, New York 10523. Issue examined: Vol. 1, No. 1, Fall 1973. Journal of the American Society for Information Science, 25(4):273, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:BRBb

[Ano74b] Anonymous. Book review: Information Retrieval On-Line. F. W. Lancaster and E. G. Fayen. Los Angeles: Wiley-Becker & Hayes. 417 pp. (1974). Journal of the American Society for Information Science, 25(5):336–337, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:BR

[Ano74c] Anonymous. Books to be reviewed. Journal of the American Society for Information Science, 25(4):273, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:EN

[Ano74d] Anonymous. Editor notes. Journal of the American Society for Information Science, 25(4):275–276, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:EE

[Ano74e] Anonymous. Errata. Journal of the American Society for Information Science, 25(6):384, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:LEA

[Ano74f] Anonymous. Letters to the Editors: 1973 Annual Meeting. Journal of the American Society for Information Science, 25(2):134– 135, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See reply [Bor74].

Anonymous:1974:Ma

[Ano74g] Anonymous. Masthead. Journal of the American Society for Information Science, 25(1):fmi, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:Mb

[Ano74h] Anonymous. Masthead. Journal of the American Society for Information Science, 25(2):fmi, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:Mc

[Ano74i] Anonymous. Masthead. Journal of the American Society for Information Science, 25(3):fmi, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:Md

[Ano74j] Anonymous. Masthead. Journal of the American Society for Information Science, 25(4):fmi, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:Me

[Ano74k] Anonymous. Masthead. Journal of the American Society for Information Science, 25(5):fmi, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:Mf

[Ano74l] Anonymous. Masthead. Journal of the American Society for Information Science, 25(6):fmi, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1974:OM

[Ano74m] Anonymous. Obituary: In memoriam. Journal of the American Society for Information Science, 25(3):138, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:BRB

[Ano75a] Anonymous. Book review: Understanding scientific literatures: A Bibliometric Approach. Joseph C. Donohue, Cambridge, Massachusetts: The MIT Press, 101 pp. (1973). Journal of the American Society for Information Science, 26(2):129, March 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:BR

[Ano75b] Anonymous. Books to be reviewed. Journal of the American Society for Information Science, 26(1):68, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:IMa

[Ano75c] Anonymous. Institutional members. Journal of the American Society for Information Science, 26(1):69–70, January 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:IMb

[Ano75d] Anonymous. Institutional members. Journal of the American Society for Information Science, 26(3):i, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:IMI

[Ano75e] Anonymous. Institutional members: Institutional members. Journal of the American Society for Information Science, 26(4):i, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:LEa

[Ano75f] Anonymous. Letters to the Editors. Journal of the American Society for Information Science, 26(3):192–194, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:LEb

[Ano75g] Anonymous. Letters to the Editors. Journal of the American Society for Information Science, 26(4):265, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:Ma

[Ano75h] Anonymous. Masthead. Journal of the American Society for Information Science, 26(1):fmi, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:Mb

[Ano75i] Anonymous. Masthead. Journal of the American Society for Information Science, 26(2):fmi, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:Mc

[Ano75j] Anonymous. Masthead. Journal of the American Society for Information Science, 26(3):fmi, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:Md

[Ano75k] Anonymous. Masthead. Journal of the American Society for Information Science, 26(4):fmi, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:Me

[Ano75l] Anonymous. Masthead. Journal of the American Society for Information Science, 26(5):fmi, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1975:Mf

[Ano75m] Anonymous. Masthead. Journal of the American Society for Information Science, 26(6):fmi, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1976:E

[Ano76a] Anonymous. Errata. Journal of the American Society for Information Science, 27(3):191, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1976:Ma

[Ano76b] Anonymous. Masthead. Journal of the American Society for Information Science, 27(1):fmi, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1976:Mb

[Ano76c] Anonymous. Masthead. Journal of the American Society for Information Science, 27(2):fmi, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1976:Mc

[Ano76d] Anonymous. Masthead. Journal of the American Society for Information Science, 27(3):fmi, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1976:Md

[Ano76e] Anonymous. Masthead. Journal of the American Society for Information Science, 27(4):fmi, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1976:Me

[Ano76f] Anonymous. Masthead. Journal of the American Society for Information Science, 27(5):fmi, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1976:N

[Ano76g] Anonymous. Note. Journal of the American Society for Information Science, 27(2):126, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:BRB

[Ano77a] Anonymous. Book review: The Community College Library Fritz Veit. Westport, CT: Greenwood Press (1975). (Contributions in Librarianship and Information Science, No. 14). Journal of the American Society for Information Science, 28(2):123–124, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:EN

[Ano77b] Anonymous. Editor’s note. Journal of the American Society for Information Science, 28(6):311–312, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:Ma

[Ano77c] Anonymous. Masthead. Journal of the American Society for Information Science, 28(1):fmi, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:Mb

[Ano77d] Anonymous. Masthead. Journal of the American Society for Information Science, 28(2):fmi, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:Mc

[Ano77e] Anonymous. Masthead. Journal of the American Society for Information Science, 28(3):fmi, May 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:Md

[Ano77f] Anonymous. Masthead. Journal of the American Society for Information Science, 28(5):fmi, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:Me

[Ano77g] Anonymous. Masthead. Journal of the American Society for Information Science, 28(6):fmi, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1977:PN

[Ano77h] Anonymous. Publisher note. Journal of the American Society for Information Science, 28(2):70, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:EN

[Ano78a] Anonymous. Editor’s note. Journal of the American Society for Information Science, 29(2):50, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:ENL

[Ano78b] Anonymous. Editor’s note: On letters to the Editor. Journal of the American Society for Information Science, 29(1):1–3, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:FAF

[Ano78c] Anonymous. Forthcoming article: Forthcoming articles. Journal of the American Society for Information Science, 29(1):3, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:Ma

[Ano78d] Anonymous. Masthead. Journal of the American Society for Information Science, 29(1):fmi, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:Mb

[Ano78e] Anonymous. Masthead. Journal of the American Society for Information Science, 29(2):fmi, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:Mc

[Ano78f] Anonymous. Masthead. Journal of the American Society for Information Science, 29(3):fmi, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:Md

[Ano78g] Anonymous. Masthead. Journal of the American Society for Information Science, 29(4):fmi, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:Me

[Ano78h] Anonymous. Masthead. Journal of the American Society for Information Science, 29(5):fmi, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:Mf

[Ano78i] Anonymous. Masthead. Journal of the American Society for Information Science, 29(6):fmi, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1978:MRA

[Ano78j] Anonymous. Miscellaneous: Read about your future in past issues of the bulletin. Journal of the American Society for Information Science, 29(5):254, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1979:E

[Ano79a] Anonymous. Erratum. Journal of the American Society for Information Science, 30(3):172, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1979:Ma

[Ano79b] Anonymous. Masthead. Journal of the American Society for Information Science, 30(1):fmi, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1979:Mb

[Ano79c] Anonymous. Masthead. Journal of the American Society for Information Science, 30(2):fmi, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1979:Mc

[Ano79d] Anonymous. Masthead. Journal of the American Society for Information Science, 30(3):fmi, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1979:Md

[Ano79e] Anonymous. Masthead. Journal of the American Society for Information Science, 30(4):fmi, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1979:Me

[Ano79f] Anonymous. Masthead. Journal of the American Society for Information Science, 30(5):fmi, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1979:Mf

[Ano79g] Anonymous. Masthead. Journal of the American Society for Information Science, 30(6):fmi, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:II

[Ano80a] Anonymous. Introduction: Introduction. Journal of the American Society for Information Science, 31(3):154, May 1980. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:Ma

[Ano80b] Anonymous. Masthead. Journal of the American Society for Information Science, 31(1):fmi, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:Mb

[Ano80c] Anonymous. Masthead. Journal of the American Society for Information Science, 31(2):fmi, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:Mc

[Ano80d] Anonymous. Masthead. Journal of the American Society for Information Science, 31(3):fmi, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:Md

[Ano80e] Anonymous. Masthead. Journal of the American Society for Information Science, 31(4):fmi, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:Me

[Ano80f] Anonymous. Masthead. Journal of the American Society for Information Science, 31(5):fmi, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:Mf

[Ano80g] Anonymous. Masthead. Journal of the American Society for Information Science, 31(6):fmi, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:MIC

[Ano80h] Anonymous. Miscellaneous: Instructions for contributors. Journal of the American Society for Information Science, 31(6):452, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1980:MLN

[Ano80i] Anonymous. Miscellaneous: Library networks and resource sharing. Journal of the American Society for Information Science, 31(6):403, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Ea

[Ano81a] Anonymous. Errata. Journal of the American Society for Information Science, 32(2):158, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Eb

[Ano81b] Anonymous. Erratum. Journal of the American Society for Information Science, 32(6):420, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:LES

[Ano81c] Anonymous. Letters to the Editor: Science, Shcharansky, and the Soviets. Journal of the American Society for Information Science, 32(1):78, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Ma

[Ano81d] Anonymous. Masthead. Journal of the American Society for Information Science, 32(1):fmi, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Mb

[Ano81e] Anonymous. Masthead. Journal of the American Society for Information Science, 32(2):fmi, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Mc

[Ano81f] Anonymous. Masthead. Journal of the American Society for Information Science, 32(3):fmi, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Md

[Ano81g] Anonymous. Masthead. Journal of the American Society for Information Science, 32(4):fmi, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Me

[Ano81h] Anonymous. Masthead. Journal of the American Society for Information Science, 32(5):fmi, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:Mf

[Ano81i] Anonymous. Masthead. Journal of the American Society for Information Science, 32(6):fmi, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:MPOb

[Ano81j] Anonymous. Miscellaneous: Perspectives on …. Cognition: Human information processing. Journal of the American Society for Information Science, 32(5):343, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1981:MPOa

[Ano81k] Anonymous. Miscellaneous: Perspectives on …. Computer and communications technology. Journal of the American Society for Information Science, 32(2):121, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:A

[Ano82a] Anonymous. Announcement. Journal of the American Society for Information Science, 33(1):7, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:EN

[Ano82b] Anonymous. Editor’s notes. Journal of the American Society for Information Science, 33(3):i, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:E

[Ano82c] Anonymous. Erratum. Journal of the American Society for Information Science, 33(2):81, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:IC

[Ano82d] Anonymous. Instructions for contributors. Journal of the American Society for Information Science, 33(4):255, July 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:Ma

[Ano82e] Anonymous. Masthead. Journal of the American Society for Information Science, 33(1):fmi, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:Mb

[Ano82f] Anonymous. Masthead. Journal of the American Society for Information Science, 33(2):fmi, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:Mc

[Ano82g] Anonymous. Masthead. Journal of the American Society for Information Science, 33(3):fmi, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:Md

[Ano82h] Anonymous. Masthead. Journal of the American Society for Information Science, 33(4):fmi, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:Me

[Ano82i] Anonymous. Masthead. Journal of the American Society for Information Science, 33(5):fmi, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1982:Mf

[Ano82j] Anonymous. Masthead. Journal of the American Society for Information Science, 33(6):fmi, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:E

[Ano83a] Anonymous. Erratum. Journal of the American Society for Information Science, 34(4):302, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:Ma

[Ano83b] Anonymous. Masthead. Journal of the American Society for Information Science, 34(1):fmi, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:Mb

[Ano83c] Anonymous. Masthead. Journal of the American Society for Information Science, 34(2):fmi, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:Mc

[Ano83d] Anonymous. Masthead. Journal of the American Society for Information Science, 34(3):fmi, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:Md

[Ano83e] Anonymous. Masthead. Journal of the American Society for Information Science, 34(4):fmi, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:Me

[Ano83f] Anonymous. Masthead. Journal of the American Society for Information Science, 34(5):fmi, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:Mf

[Ano83g] Anonymous. Masthead. Journal of the American Society for Information Science, 34(6):fmi, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:MCP

[Ano83h] Anonymous. Miscellaneous: Call of papers. Journal of the American Society for Information Science, 34(5):363, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1983:SRS

[Ano83i] Anonymous. Suggested readings: Suggested Readings. Journal of the American Society for Information Science, 34(6):439–440, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Ea

[Ano84a] Anonymous. Erratum. Journal of the American Society for Information Science, 35(1):57, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Eb

[Ano84b] Anonymous. Erratum. Journal of the American Society for Information Science, 35(2):134, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Ec

[Ano84c] Anonymous. Erratum. Journal of the American Society for Information Science, 35(2):138, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Ma

[Ano84d] Anonymous. Masthead. Journal of the American Society for Information Science, 35(1):fmi, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Mb

[Ano84e] Anonymous. Masthead. Journal of the American Society for Information Science, 35(2):fmi, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Mc

[Ano84f] Anonymous. Masthead. Journal of the American Society for Information Science, 35(3):fmi, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Md

[Ano84g] Anonymous. Masthead. Journal of the American Society for Information Science, 35(4):fmi, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Me

[Ano84h] Anonymous. Masthead. Journal of the American Society for Information Science, 35(5):fmi, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:Mf

[Ano84i] Anonymous. Masthead. Journal of the American Society for Information Science, 35(6):fmi, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:MAM

[Ano84j] Anonymous. Miscellaneous: 20th Annual Meeting Drug Information Association. Journal of the American Society for Information Science, 35(2):139–140, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:MBAa

[Ano84k] Anonymous. Miscellaneous: Journal of the American Society for Information Science. Journal of the American Society for Information Science, 35(1):58, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:MBAb

[Ano84l] Anonymous. Miscellaneous: Journal of the American Society for Information Science. Journal of the American Society for Information Science, 35(2):142, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:MDI

[Ano84m] Anonymous. Miscellaneous: Defense Information School (DIN-FOS) Alumni Search. Journal of the American Society for Information Science, 35(2):141, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:PLI

[Ano84n] Anonymous. Perspectives on …: Information science: Retrospect and prospect. Journal of the American Society for Information Science, 35(3):155, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1984:PN

[Ano84o] Anonymous. Publisher’s note. Journal of the American Society for Information Science, 35(6):324, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:ICI

[Ano85a] Anonymous. Instructions for contributors: Instructions for contributors. Journal of the American Society for Information Science, 36(1):58, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:Ma

[Ano85b] Anonymous. Masthead. Journal of the American Society for Information Science, 36(2):fmi, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:Mb

[Ano85c] Anonymous. Masthead. Journal of the American Society for Information Science, 36(3):fmi, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:Mc

[Ano85d] Anonymous. Masthead. Journal of the American Society for Information Science, 36(4):fmi, July 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:Md

[Ano85e] Anonymous. Masthead. Journal of the American Society for Information Science, 36(5):fmi, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:Me

[Ano85f] Anonymous. Masthead. Journal of the American Society for Information Science, 36(6):fmi, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:MBA

[Ano85g] Anonymous. Miscellaneous: Journal of the American Society for Information Science. Journal of the American Society for Information Science, 36(6):421, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1985:MCP

[Ano85h] Anonymous. Miscellaneous: Call for papers. Journal of the American Society for Information Science, 36(2):140, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:CPCa

[Ano86a] Anonymous. Call for papers: Call for papers. Journal of the American Society for Information Science, 37(1):i, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:CPCb

[Ano86b] Anonymous. Call for papers: Call for papers. Journal of the American Society for Information Science, 37(3):179, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:CPCc

[Ano86c] Anonymous. Call for papers: Call for papers. Journal of the American Society for Information Science, 37(6):443, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:E

[Ano86d] Anonymous. Erratum. Journal of the American Society for Information Science, 37(2):95, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:LE

[Ano86e] Anonymous. Letters to the Editors. Journal of the American Society for Information Science, 37(3):177–178, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:Ma

[Ano86f] Anonymous. Masthead. Journal of the American Society for Information Science, 37(1):fmi, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:Mb

[Ano86g] Anonymous. Masthead. Journal of the American Society for Information Science, 37(2):fmi, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:Mc

[Ano86h] Anonymous. Masthead. Journal of the American Society for Information Science, 37(3):fmi, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:Md

[Ano86i] Anonymous. Masthead. Journal of the American Society for Information Science, 37(4):fmi, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:Me

[Ano86j] Anonymous. Masthead. Journal of the American Society for Information Science, 37(5):fmi, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1986:Mf

[Ano86k] Anonymous. Masthead. Journal of the American Society for Information Science, 37(6):fmi, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1987:E

[Ano87] Anonymous. Errata. Journal of the American Society for Information Science, 38(6):448, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1989:PP

[Ano89] Anonymous. Postscript: Postscript. Journal of the American Society for Information Science, 40(3):221, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1994:Ea

[Ano94a] Anonymous. Erratum. Journal of the American Society for Information Science, 45(6):441, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1994:Eb

[Ano94b] Anonymous. Erratum. Journal of the American Society for Information Science, 45(8):568, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1994:Ec

[Ano94c] Anonymous. Erratum. Journal of the American Society for Information Science, 45(10):802, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1995:Ea

[Ano95a] Anonymous. Erratum. Journal of the American Society for Information Science, 46(6):473, July 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1995:Eb

[Ano95b] Anonymous. Erratum. Journal of the American Society for Information Science, 46(10):801, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1996:Ea

[Ano96a] Anonymous. Erratum. Journal of the American Society for Information Science, 47(3):260, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1996:Eb

[Ano96b] Anonymous. Erratum. Journal of the American Society for Information Science, 47(7):584, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1997:E

[Ano97] Anonymous. Erratum. Journal of the American Society for Information Science, 48(1):91–93, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1999:Ea

[Ano99a] Anonymous. Erratum. Journal of the American Society for Information Science, 50(6):553, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1999:Eb

[Ano99b] Anonymous. Erratum. Journal of the American Society for Information Science, 50(7):635, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:1999:Ec

[Ano99c] Anonymous. Erratum. Journal of the American Society for Information Science, 50(9):842, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:2000:CPC

[Ano00a] Anonymous. Call for papers: Call for papers: Special topic issue of JASIS: Visual based retrieval systems and Web mining. Journal of the American Society for Information Science, 51(7):688, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:2000:ECPb

[Ano00b] Anonymous. Editorials: Call for papers: Special topic issue of JASIS: Web research. Journal of the American Society for Information Science, 51(14):1338, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:2000:ECPa

[Ano00c] Anonymous. Editorials: Call for papers: Special topic issue of JASIS: XML and information retrieval. Journal of the American Society for Information Science, 51(13):1250, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:2000:E

[Ano00d] Anonymous. Erratum. Journal of the American Society for Information Science, 51(6):583, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Anonymous:2008:ERD

[Ano08] Anonymous. Erratum re: “Digital library resources as a basis for collaborative work” JASIST 51(3) 2000, 228–245. Journal of the American Society for Information Science and Technology: JASIST, 59(9):1535, July 2008. CODEN JASIEF. ISSN 1532-2882 (print), 1532-2890 (electronic). See [Wil00a].

Anvari:1991:RDM

[Anv91] Morteza Anvari. Research: A database model for medical consultation. Journal of the American Society for Information Science, 42(8):554–557, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Alvarez:1996:RRMb

[AP96a] Pedro Alvarez and Antonio Pulgarin. Research: The Rasch model. Measuring information from keywords: the diabetes field. Journal of the American Society for Information Science, 47(6): 468–476, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Alvarez:1996:RRMa

[AP96b] Pedro Alvarez and Antonio Pulgarín. Research: The Rasch model. Measuring the impact of scientific journals: Analytical chemistry. Journal of the American Society for Information Science, 47(6):458–467, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Alvarez:1997:BCD

[AP97] Pedro Alvarez and Antonio Pulgarín. Brief communications: The diffusion of scientific journals analyzed through citations. Journal of the American Society for Information Science, 48(10):953–958, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Armstrong:1994:RDG

[Arm94] Marc P. Armstrong. Requirements for the development of GISbased group decision-support systems. Journal of the American Society for Information Science, 45(9):669–677, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1971:BCD

[Art71] Susan Artandi. Brief communications: Document retrieval and the concept of sets. Journal of the American Society for Information Science, 22(4):289–290, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1973:OPI

[Art73] Susan Artandi. Opinion papers: Information concepts and their utility. Journal of the American Society for Information Science, 24(4):242–245, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1974:BRB

[Art74] Susan Artandi. Book review: The first book of information science, Joseph Becker. United States Atomic Energy Commission Office of Information Services, 91 pp., 1973. Journal of the American Society for Information Science, 25(6):382, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1975:ARW

[Art75a] Susan Artandi. Appendix: Reply to the Williams and Kim article. Journal of the American Society for Information Science, 26 (1):9, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [WK75].

Artandi:1975:BRB

[Art75b] Susan Artandi. Book review: Computer-based information services in science and technology-principles and techniques, M. F. Lynch. Peter Peregrinus Ltd. (1974). Journal of the American Society for Information Science, 26(5):307, September 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1976:BRB

[Art76a] Susan Artandi. Book review: Dynamic library and information processing. Gerard Salton. Englewood Cliffs, N.J.: Prentice-Hall, Inc., 523 pp. (1975). Journal of the American Society for Information Science, 27(2):124, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1976:MIL

[Art76b] Susan Artandi. Machine indexing: Linguistic and semiotic implications. Journal of the American Society for Information Science, 27(4):235–239, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1977:BRB

[Art77] Susan Artandi. Book review: The Information Age: Its Development, Its Impact. Donald P. Hammer, Ed. Metuchen, NJ: Scarecrow Press, 275 pp. (1976). Journal of the American Society for Information Science, 28(6):373–374, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1979:MIS

[Art79] Susan Artandi. Man, information, and society: New patterns of interaction. Journal of the American Society for Information Science, 30(1):15–18, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Artandi:1982:CPS

[Art82] Susan Artandi. Computers and the Postindustrial Society: Symbiosis or Information Tyranny? Journal of the American Society for Information Science, 33(5):302–307, September 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Asai:1981:AAD

[Asa81a] Isao Asai. Adjusted age distribution and its application to impact factor and immediacy index. Journal of the American Society for Information Science, 32(3):172–174, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Asai:1981:GFB

[Asa81b] Isao Asai. A general formulation of Bradford’s distribution: the graph-oriented approach. Journal of the American Society for Information Science, 32(2):113–119, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Abu-Salem:1999:RSM

[ASAOE99] Hani Abu-Salem, Mahmoud Al-Omari, and Martha W. Evens. Research: Stemming methodologies over individual query words for an Arabic Information Retrieval System. Journal of the American Society for Information Science, 50(6):524–529, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Alvarez:1985:ISC

[ASR85] Jose Alvarez, Steven M. Smiley, and Francisco Rohrmann. Informatics and small computers in Latin America. Journal of the American Society for Information Science, 36(4):259–267, July 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ashmole:1973:CEC

[ASS73] R. F. Ashmole, D. E. Smith, and B. T. Stern. Cost effectiveness of current awareness sources in the pharmaceutical industry. Journal of the American Society for Information Science, 24(1): 29–39, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Atherton:1972:LEB

[Ath72] Pauline Atherton. Letters to the Editor: Bibliographic access to IS. Journal of the American Society for Information Science, 23(4):286–287, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Arth:1995:MAG

[ATJ95] Janet M. Arth, Roy Tally, and Joan D. Jensen. In memoriam: Audrey N. Grosch, January 10, 1934–July 16, 1994. Journal of the American Society for Information Science, 46(2):81–82, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Auld:1971:KIV

[Aul71] Larry Auld. KWOC indexes and vocabulary comparisons of summaries of LC and DC classification schedules. Journal of the American Society for Information Science, 22(5):322–325, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Auld:1979:LEW

[Aul79] Larry Auld. Letter to the Editors: The wrong way to go. Journal of the American Society for Information Science, 30(6):372, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Auld:1987:BRB

[Aul87] Lawrence W. S. Auld. Book review: Careers in other fields for librarians: Successful strategies for finding the job. Journal of the American Society for Information Science, 38(4):312–313, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Auster:1993:BRB

[Aus93] Ethel Auster. Book review: Automated information retrieval in libraries: a management handbook. Journal of the American Society for Information Science, 44(7):428–429, August 1993. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Avramescu:1971:PCO

[Avr71] Aurel Avramescu. Probabilistic criteria for the objective design of descriptor languages. Journal of the American Society for Information Science, 22(2):85–95, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Avramescu:1979:AOS

[Avr79] Aurel Avramescu. Actuality and obsolescence of scientific literature. Journal of the American Society for Information Science, 30 (5):296–303, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ahmad:1996:RES

[AYS96] Fatimah Ahmad, Mohammed Yusoff, and Tengku M. T. Sembok. Research: Experiments with a stemming algorithm for Malay words. Journal of the American Society for Information Science, 47(12):909–918, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belzer:1974:LEU

[BA74] Jack Belzer and Susan Artandi. Letters to the Editors: The utility of information and communication theory. Journal of the American Society for Information Science, 25(3):208, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Back:1972:WID

[Bac72] Harry B. Back. What information dissemination studies imply concerning the design of on-line reference retrieval systems. Journal of the American Society for Information Science, 23(3):156– 163, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Backer:1993:POK

[Bac93] Thomas E. Backer. Perspectives on … knowledge utilization: Information alchemy: Transforming information through knowledge utilization. Journal of the American Society for Information Science, 44(4):217–221, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bagley:1981:LEO

[Bag81] Philip R. Bagley. Letters to the Editor: Online systems: History, technology, and economics. Journal of the American Society for Information Science, 32(1):78, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bailey:1974:DUA

[Bai74] Martha J. Bailey. Densities of use and absence of obsolescence in physics journals at M.I.T. Journal of the American Society for Information Science, 25(6):383, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bailey:1975:BRB

[Bai75] Martha J. Bailey. Book review: Interpersonal communication: Counseling, Guidance and Retrieval for Media, Library, and Information Specialists. Patrick R. Penland and Alayamma Mathai, New York, NY: Marcel Dekker, Inc., 171 pp. (1974). Journal of the American Society for Information Science, 26(5):307, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bailey:1976:UAA

[Bai76] Martha J. Bailey. The use of abbreviations and acronyms in the physics literature. Journal of the American Society for Information Science, 27(2):81–84, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baker:1989:BRB

[Bak89] Sharon L. Baker. Book review: Marketing/planning library and information services. Journal of the American Society for Information Science, 40(6):442, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Balasubramanian:1984:LEM

[Bal84] V. Balasubramanian. Letters to the Editor: Maximum entropy principle. Journal of the American Society for Information Science, 35(3):153, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Balasubramanian:1988:LEC

[Bal88] V. Balasubramanian. Letters to the Editor: Controlled vocabulary in information retrieval. Journal of the American Society for Information Science, 39(4):285–286, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ball:1989:PLIa

[Bal89] Marion J. Ball. Perspectives on … information science and health informatics education: Dedication: In memory of Marsden S. Blois, 1919–1988. Journal of the American Society for Information Science, 40(5):367, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baldonado:2000:RUC

[Bal00] Michelle Q. Wang Baldonado. Research: A user-centered interface for information exploration in a heterogeneous digital library. Journal of the American Society for Information Science, 51(3): 297–310, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bamford:1979:AEC

[Bam79] Harold E. Bamford, Jr. Assessing the effect of computer augmentation on staff productivity. Journal of the American Society for Information Science, 30(3):136–142, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Banerjee:1998:BRB

[Ban98] Kyle Banerjee. Book review: Advanced database systems. Journal of the American Society for Information Science, 49(4):388–389, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barry:1974:LER

[Bar74] S. G. Barry. Letter to the Editors: Research Investigation into Experties Indexes. Journal of the American Society for Information Science, 25(4):274–275, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baratz:1978:BCW

[Bar78] Stephen S. Baratz. Brief communications: When is a changed state not a changed state? When a theory of mechanics persists in the face of a theory of relativity. Journal of the American Society for Information Science, 29(3):163–164, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barry:1994:RUD

[Bar94] Carol L. Barry. Research: User-defined relevance criteria: an exploratory study. Journal of the American Society for Information Science, 45(3):149–159, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barreau:1995:RCF

[Bar95] Deborah K. Barreau. Research: Context as a factor in personal information management systems. Journal of the American Society for Information Science, 46(5):327–339, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barron:1996:HPD

[Bar96] Daniel D. Barron. Historical perspectives: Distance education in North American library and information science education: Applications of technology and commitment. Journal of the American Society for Information Science, 47(11):805–810, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barry:1998:DRC

[Bar98a] Carol L. Barry. Document representations and clues to document relevance. Journal of the American Society for Information Science, 49(14):1293–1303, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barry:1998:BRB

[Bar98b] Richard E. Barry. Book review: Document management for the enterprise: Principles, techniques and applications. Journal of the American Society for Information Science, 49(1):94–95, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bassler:1974:BRB

[Bas74] Richard A. Bassler. Book review: Introduction to PL/I programming for library and information science. Thomas H. Mott, Jr., Susan Artandi and Leny Struminger. New York: Academic Press, 231 pp. (1972). Journal of the American Society for Information Science, 25(2):133, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Basu:1992:RHD

[Bas92] Aparna Basu. Research: Hierarchical distributions and Bradford’s Law. Journal of the American Society for Information Science, 43(7):494–500, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Basu:1994:LEE

[Bas94] Aparna Basu. Letters to the Editor: The elusive phenomenon of Bradford’s Law. Journal of the American Society for Information Science, 45(1):59–60, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1977:FAS

[Bat77] Marcia J. Bates. Factors affecting subject catalog search success. Journal of the American Society for Information Science, 28(3): 161–169, May 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1979:IT

[Bat79a] Marcia J. Bates. Idea Tactics. Journal of the American Society for Information Science, 30(5):280–289, September 1979. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1979:IST

[Bat79b] Marcia J. Bates. Information search tactics. Journal of the American Society for Information Science, 30(4):205–214, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Battin:1983:LER

[Bat83] Patricia Battin. Letters to the Editor: Research libraries group status. Journal of the American Society for Information Science, 34(4):296–297, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1986:RSA

[Bat86] Marcia J. Bates. Research: Subject access in online catalogs: a design model. Journal of the American Society for Information Science, 37(6):357–376, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Batty:1993:BRB

[Bat93] David Batty. Book review: Classification research for knowledge representation and organization. Journal of the American Society for Information Science, 44(10):593–595, December 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Battenfeld:1997:BRB

[Bat97] Robert L. Battenfeld. Book review: Ethics and computing: Living responsibly in a computerized world. Journal of the American Society for Information Science, 48(7):671–672, July 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1998:RIA

[Bat98] Marcia J. Bates. Research: Indexing and access for digital libraries and the Internet: Human, database, and domain factors. Journal of the American Society for Information Science, 49(13): 1185–1205, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1999:EAB

[Bat99a] Marcia J. Bates. Editorial: The 50th anniversary of the Journal of the American Society for Information Science: Guest Editor introduction. Journal of the American Society for Information Science, 50(11):960–964, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1999:ISI

[Bat99b] Marcia J. Bates. The invisible substrate of information science. Journal of the American Society for Information Science, 50(12):

1043–1050, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:1999:RTI

[Bat99c] Marcia J. Bates. Research: A tour of information science through the pages of JASIS. Journal of the American Society for Information Science, 50(11):975–993, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bates:2000:LEP

[Bat00] Marcia J. Bates. Letters to the Editor: Proper citations. Journal of the American Society for Information Science, 51(9):882, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baughman:1972:BCF

[Bau72] James C. Baughman. Brief communications: Federal aid and the growth of a subject literature. Journal of the American Society for Information Science, 23(2):129–130, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baxter:1985:ERL

[Bax85] Paul Baxter. European research letters: The role of the British library R&D department in supporting library and information research in the United Kingdom. Journal of the American Society for Information Science, 36(4):275–277, July 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bingley:1977:LEP

[BB77] Clive Bingley and Martha J. Bailey. Letters to the Editor: Physics abbreviations and acronyms. Journal of the American Society for Information Science, 28(2):129, March 1977. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baumol:1983:ECD

[BB83a] William J. Baumol and Sue Anne Batey Blackman. Electronics, the cost disease, and the operation of libraries. Journal of the American Society for Information Science, 34(3):181–191, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bujdoso:1983:PIR

[BB83b] Ernö Bujdosó and Tibor Braun. Publication indicators of relative research efforts in physics subfields. Journal of the American Society for Information Science, 34(2):150–155, March 1983. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bierbaum:1995:RLA

[BB95] Esther Green Bierbaum and Terrence A. Brooks. Research: The literature of acquired immunodeficiency syndrome (AIDS): Continuing changes in publication patterns and subject access. Journal of the American Society for Information Science, 46(7):530– 536, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beaulieu:1996:NEO

[BB96a] Micheline Beaulieu and Christine L. Borgman. A new era for OPAC research: Introduction to special topic issue on current research in online public access systems. Journal of the American Society for Information Science, 47(7):491–492, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Besser:1996:EDE

[BB96b] Howard Besser and Maria Bonn. Effectiveness of distance education: Impact of distance independent education. Journal of the American Society for Information Science, 47(11):880–883, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bharat:2000:RCT

[BBDH00] Krishna Bharat, Andrei Broder, Jeffrey Dean, and Monika R. Henzinger. Research: A comparison of techniques to find mirrored hosts on the WWW. Journal of the American Society for Information Science, 51(12):1114–1122, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Berghel:1999:RCI

[BBFM99] Hal Berghel, Daniel Berleant, Thomas Foy, and Marcus McGuire. Research: Cyberbrowsing: Information customization on the Web. Journal of the American Society for Information Science, 50(6):505–513, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Biswas:1987:RKAa

[BBMS87] Gautam Biswas, James C. Bezdek, Marisol Marques, and Viswanath Subramanian. Research: Knowledge-assisted document retrieval: I. The natural-language interface. Journal of the American Society for Information Science, 38(2):83–96, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Biswas:1987:RKAb

[BBSM87] Gautam Biswas, James C. Bezdek, Viswanath Subramanian, and Marisol Marques. Research: Knowledge-assisted document retrieval: II. The retrieval process. Journal of the American Society for Information Science, 38(2):97–110, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Berge:1996:BJR

[BC96] Zane L. Berge and Mauri P. Collins. IPCT Journal readership survey. Journal of the American Society for Information Science, 47(9):701–710, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Budd:1997:RUF

[BC97] John M. Budd and Lynn Silipigni Connaway. Research: University faculty and networked information: Results of a survey. Journal of the American Society for Information Science, 48(9): 843–852, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1999:POI

[BC99] Martin Brooks and Jennifer Campbell. Perspectives on … visual information retrieval interfaces: Interactive graphical queries for bibliographic search. Journal of the American Society for Information Science, 50(9):814–825, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bian:2000:RCL

[BC00] Guo-Wei Bian and Hsin-Hsi Chen. Research: Cross-language information access to multilingual collections on the Internet. Journal of the American Society for Information Science, 51(3): 281–296, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bartell:1995:RRD

[BCB95] Brian T. Bartell, Garrison W. Cottrell, and Richard K. Belew. Research: Representing documents using an explicit model of their similarities. Journal of the American Society for Information Science, 46(4):254–271, May 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brandt:1996:PLC

[BCB96] Kerryn A. Brandt, Jayne M. Campbell, and Willard F. Bryant, Jr. Perspectives on … cost and pricing of library and information services in transition: Reflections on reference services. Journal of the American Society for Information Science, 47(3):210–216, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bartell:1998:ROS

[BCB98] Brian T. Bartell, Garrison W. Cottrell, and Richard K. Belew. Research: Optimizing similarity using multi-query relevance feedback. Journal of the American Society for Information Science, 49(8):742–761, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bommer:1979:PAM

[BCG79] Michael R. Bommer, Ronald W. Chorba, and Walter Grattidge. Performance assessment model for academic libraries. Journal of the American Society for Information Science, 30(2):93–99, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borkowski:1970:SEC

[BCM^{+}70] Casimir Borkowski, Louis Cepanec, J. Sperling Martin, Virginia Salko, and Siegfried Treu. Structure and effectiveness of The Citation Identifier, an operational computer program for automatic identification of case citations in legal literature. Journal of the American Society for Information Science, 21(1):8–15, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1989:RDE

[BCM89] Christine L. Borgman, Donald O. Case, and Charles T. Meadow. Research: The design and evaluation of a front-end user interface for energy researchers. Journal of the American Society for Information Science, 40(2):99–109, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Berthelot:1979:IIC

[BCSZ79] A. Berthelot, Peter Clague, Samuel Schiminovich, and Wolfgang Zwirner. The ICSU AB International classification system for physics: Its history and future. Journal of the American Society for Information Science, 30(6):343–352, November 1979. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ball:1988:PLI

[BD88] Marion J. Ball and Judith V. Douglas. Perspectives on … integrated academic information management systems (IAIMS): Part II. Planning and implementing integrated information services. Integration and outreach: Integrated Academic Information Management Systems (IAIMS) at Maryland. Journal of the American Society for Information Science, 39(2):107–112, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burgin:1992:RID

[BD92] Robert Burgin and Martin Dillon. Research: Improving disambiguation in FASIT. Journal of the American Society for Information Science, 43(2):101–114, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Besser:1996:PLD

[BD96] Howard Besser and Stacey Donahue. Perspectives on … distance independent education: Introduction and overview. Journal of the American Society for Information Science, 47(11):801–804, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Branin:1997:PLIa

[BD97] Joseph Branin and George D’Elia. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: The implementation of the Integrated Information Center at the University of Minnesota. Journal of the American Society for Information Science, 48(5):442–445, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beath:1997:PLI

[BDBR97] Cynthia Beath, George D’Elia, Joseph Branin, and Nancy Rohde. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: Challenges encountered during the implementation of the Integrated Information Center at the University of Minnesota. Journal of the American Society for Information Science, 48(5):438–441, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ball:1988:PLE

[BDL88] Marion J. Ball, Judith V. Douglas, and Lois F. Lunin. Perspectives on … education of the information professional: New dimensions, new directions: Informatics and education in the health professions. Journal of the American Society for Information Science, 39(5):344–347, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bosc:1998:FFD

[BDP98] Patrick Bosc, Didier Dubois, and Henri Prade. Fuzzy functional dependencies and redundancy elimination. Journal of the American Society for Information Science, 49(3):217–235, March 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Branin:1997:PLIb

[BDSA97] Joseph Branin, George D’Elia, Thomas Shaughnessy, and Carl Adams. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: The evaluation of the IIC project. Journal of the American Society for Information Science, 48(5):472–473, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ball:1989:PLIb

[BDZL89] Marion J. Ball, Judith V. Douglas, John L. Zimmerman, and Lois F. Lunin. Perspectives on … information science and health informatics education: Informatics education and the professions. Journal of the American Society for Information Science, 40(5): 368–377, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bichteler:1980:CUB

[BE80] Julie Bichteler and Edward A. Eaton III. The combined use of bibliographic coupling and cocitation for document retrieval. Journal of the American Society for Information Science, 31(4): 278–282, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bonckaert:1991:RRN

[BE91] P. Bonckaert and L. Egghe. Research: Rational normalization of concentration measures. Journal of the American Society for Information Science, 42(10):715–722, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beasley:1980:BCA

[Bea80] Ruth Ann Beasley. Brief communication: Another look at OCLC’s potential for special libraries. Journal of the American Society for Information Science, 31(4):300–301, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beard:1991:PLI

[Bea91] David Volk Beard. Perspectives on … imaging. Advanced applications: Computer human interaction for image information systems. Journal of the American Society for Information Science, 42(8):600–608, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bearman:1992:PLI

[Bea92] David Bearman. Perspectives on … information technology standards: A user community discovers IT standards. Journal of the American Society for Information Science, 43(8):576–578, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bearman:1993:BRB

[Bea93] Toni Carbo Bearman. Book review: Library orthodoxies: a decade of change. Journal of the American Society for Information Science, 44(6):365, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bean:1997:BRB

[Bea97] Carol A. Bean. Book review: The art of abstracting (2nd ed.). Journal of the American Society for Information Science, 48(7):

670–671, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beck:1971:KIS

[Bec71] Leonard N. Beck. Key issues in Soviet information. Journal of the American Society for Information Science, 22(2):81–84, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Becker:1984:ISV

[Bec84] Joseph Becker. An information scientist’s view on evolving information technology. Journal of the American Society for Information Science, 35(3):164–169, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beghtol:1995:RDA

[Beg95] Clare Beghtol. Research: Domain analysis, literary warrant, and consensus: the case of fiction studies. Journal of the American Society for Information Science, 46(1):30–44, January 1995. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beheshti:1996:BRBa

[Beh96a] Jamshid Beheshti. Book review: Sociomedia: Multimedia, hypermedia, and the social construction of knowledge. Journal of the American Society for Information Science, 47(5):402–403A, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beheshti:1996:BRBb

[Beh96b] Jamshid Beheshti. Book review: Sociomedia: Multimedia, hypermedia, and the social construction of knowledge. Journal of the American Society for Information Science, 47(5):402–403A, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belzer:1970:EIS

[Bel70] Jack Belzer. Education in Information Science. Journal of the American Society for Information Science, 21(4):269–273, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belzer:1971:BCJ

[Bel71] Jack Belzer. Brief communications: Justification for automatic indexing by frequency distribution of worlds. Journal of the American Society for Information Science, 22(3):226, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belzer:1973:ITM

[Bel73] Jack Belzer. Information theory as a measure of information content. Journal of the American Society for Information Science, 24(4):300–304, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belkin:1975:SSC

[Bel75] Nicholas J. Belkin. Some Soviet concepts of information for information science. Journal of the American Society for Information Science, 26(1):56–64, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belzer:1976:BRB

[Bel76] Jack Belzer. Book review: Concise survey of computer methods. Peter Naur. New York: Petrocelli Books, 397 pp. (1975). Journal of the American Society for Information Science, 27(2):125–126, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belkin:1977:LEI

[Bel77] Nicholas J. Belkin. Letters to the Editor: “information” in information science — a definition. Journal of the American Society for Information Science, 28(2):129, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bellardo:1981:BRB

[Bel81] Trudi Bellardo. Book review: The nature and future of the catalog: Proceedings of the ALA’s Information Science and Automation Division’s 1975 and 1977 Institutes on the Catalog. Maurice J. Freedman and S. Michael Malinconico, Eds. Phoenix, AZ: Oryx Press; 1979: 317 pp. Journal of the American Society for Information Science, 32(5):337, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bellardo:1985:IOS

[Bel85] Trudi Bellardo. An investigation of online searcher traits and their relationship to search outcome. Journal of the American Society for Information Science, 36(4):241–250, July 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bellardo:1988:PLE

[Bel88] Trudi Bellardo. Perspectives on … education of the information professional: New dimensions, new directions: Part IV. Different perspectives. Options and trends in the training of information professionals. Journal of the American Society for Information Science, 39(5):348–350, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bennion:1980:PTB

[Ben80] Bruce C. Bennion. Performance testing of a book and its index as an information retrieval system. Journal of the American Society for Information Science, 31(4):264–270, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bensman:1986:BRB

[Ben86] Stephen J. Bensman. Book review: Essays of an information scientist. Journal of the American Society for Information Science, 37(3):169, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bensman:2000:PDL

[Ben00] Stephen J. Bensman. Probability distributions in library and information science: a historical and practitioner viewpoint. Journal of the American Society for Information Science, 51(9):816– 833, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bernier:1970:TLT

[Ber70] Charles L. Bernier. Terse literatures: I. Terse conclusions. Journal of the American Society for Information Science, 21(5):316– 319, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bernier:1971:BCM

[Ber71] Charles L. Bernier. Brief communications: Measurement of how well professional people keep up with their technical literature.

Journal of the American Society for Information Science, 22(4): 292–293, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Berg:1972:EAD

[Ber72a] Sanford V. Berg. An economic analysis of the demand for scientific journals. Journal of the American Society for Information Science, 23(1):23–29, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bernier:1972:LEG

[Ber72b] Charles L. Bernier. Letters to the Editor: Get smart? Journal of the American Society for Information Science, 23(5):335, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Berninger:1982:SRF

[Ber82] Douglas E. Berninger. Some reflections on federal policy for secondary information systems and services, 1945–1981. Journal of the American Society for Information Science, 33(3):162–167, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bernier:1984:BRB

[Ber84] Charles L. Bernier. Book review: Second-hand knowledge. An inquiry into cognitive authority. Patrick Wilson. Greenwood Press, 1983. Journal of the American Society for Information Science, 35(4):254–255, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bernier:1985:EK

[Ber85] Charles L. Bernier. Ethics of knowing. Journal of the American Society for Information Science, 36(3):211–212, May 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Berk:1986:BRB

[Ber86a] Robert Berk. Book review: Introduction to reference sources in the health sciences. Journal of the American Society for Information Science, 37(3):174–176, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bernard:1986:PLT

[Ber86b] Keith E. Bernard. Perspectives on … telecommunications: Perspectives, developments, prospects: Regulatory development in the U.S. Journal of the American Society for Information Science, 37(6):409–413, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Berliss:1994:PLE

[Ber94] Jane Berliss. Perspectives on … electronic publishing: Boon or bust? Access to electronic publishing by individuals using adaptive computer technology. Journal of the American Society for Information Science, 45(10):753–759, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Besser:1991:PLI

[Bes91] Howard Besser. Perspectives on … imaging. Advanced applications: Imaging: Fine arts. Journal of the American Society for Information Science, 42(8):589–596, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Besser:1996:DMI

[Bes96] Howard Besser. Discussion of methods: Issues and challenges for the distance independent environment. Journal of the American Society for Information Science, 47(11):817–820, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bezilla:1982:LET

[Bez82] Robert Bezilla. Letter to the Editors: Thomas Edison and science. Journal of the American Society for Information Science, 33(1):60, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blackshaw:1988:RDM

[BF88] Lyn Blackshaw and Baruch Fischhoff. Research: Decision making in online searching. Journal of the American Society for Information Science, 39(6):369–389, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Branin:1991:PLI

[BF91a] Joseph Branin and Charles Finn. Perspectives on … integrated information centers within academic environments: Part II. A model academic Integrated Information Center. The setting for the project. Journal of the American Society for Information Science, 42(2):137–138, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1991:RET

[BF91b] Michael K. Buckland and Doris Florian. Research: Expertise, task complexity, and artificial intelligence: a conceptual framework. Journal of the American Society for Information Science, 42(9):635–643, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burrell:1993:RYG

[BF93] Quentin L. Burrell and Michael R. Fenton. Research: Yes, the GIGP really does work — and is workable! Journal of the American Society for Information Science, 44(2):61–69, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burrell:1994:RML

[BF94] Quentin L. Burrell and Michael R. Fenton. Research: A model for library book circulations incorporating loan periods. Journal of the American Society for Information Science, 45(2):101–116, March 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Broering:1988:PLI

[BFM88] Naomi C. Broering, Cyril Feng, and Nina W. Matheson. Perspectives on … integrated academic information management systems (IAIMS): Integration across institutions: IAIMS extended. Journal of the American Society for Information Science, 39(2): 131–134, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brookes:1978:FRD

[BG78] Bertram C. Brookes and Jose M. Griffiths. Frequency-rank distributions. Journal of the American Society for Information Science, 29(1):5–13, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1994:RRB

[BG94] Michael Buckland and Fredric Gey. Research: The relationship between Recall and Precision. Journal of the American Society for Information Science, 45(1):12–19, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bell:1998:CMR

[BG98] D. A. Bell and J. W. Guan. Computational methods for rough classification and discovery. Journal of the American Society for Information Science, 49(5):403–414, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bearman:1985:IP

[BGM85] Toni Carbo Bearman, Polly Guynup, and Sandra N. Milevski. Information and productivity. Journal of the American Society for Information Science, 36(6):369–375, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1997:I

[BH97] Michael Buckland and Trudi Bellardo Hahn. Introduction. Journal of the American Society for Information Science, 48(4):285– 288, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bar-Hillel:1974:LTE

[BHCC^{+}74] Y. Bar-Hillel, R. Carnap, E. C. Cherry, E. Garfield, D. W. King, F. W. Lancaster, J. C. R. Licklider, D. M. Mackay, J. W. Perry, D. J. De S. Price, G. Salton, C. Shannon, M. Taube, B. C. Vickery, and A. E. Cawkell. Letters to the Editors: Where the action is and was in information science. Journal of the American Society for Information Science, 25(5):340, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1996:ROM

[BHH96] Christine L. Borgman, Sandra G. Hirsh, and John Hiller. Rethinking online monitoring methods for information retrieval systems: From search product to search process. Journal of the American Society for Information Science, 47(7):568–583, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1995:RCS

[BHWG95] Christine L. Borgman, Sandra G. Hirsh, Virginia A. Walter, and Andrea L. Gallagher. Research: Children’s searching behavior on browsing and keyword online catalogs: The Science Library Catalog project. Journal of the American Society for Information Science, 46(9):663–684, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barnard:1997:TEI

[BI97] David T. Barnard and Nancy M. Ide. The text encoding initiative: Flexible and extensible document encoding. Journal of the American Society for Information Science, 48(7):622–628, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bar-Ilan:1998:BRB

[BI98] Judit Bar-Ilan. Book review: Proceedings of the Sixth Conference of the International Society for Scientometrics and Informetrics. Jerusalem, June 16–19, 1997. Journal of the American Society for Information Science, 49(6):568–569, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bar-Ilan:2000:WIS

[BI00] Judit Bar-Ilan. The Web as an information source on informetrics? A content analysis. Journal of the American Society for Information Science, 51(5):432–443, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bierbaum:1995:BRB

[Bie95] Esther Green Bierbaum. Book review: The classification of fiction: the development of a system based on theoretical principles. Journal of the American Society for Information Science, 46(5): 389–390, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belzer:1971:CIS

[BIFW71] Jack Belzer, Akkanad Isaac, Eugene Finkelstein, and James Williams. Curricula in information science: Analysis and development. Journal of the American Society for Information Science, 22(3):193–223, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bilal:2000:RCU

[Bil00] Dania Bilal. Research: Children’s use of the Yahooligans! Web search engine: I. Cognitive, physical, and affective behaviors on fact-based search tasks. Journal of the American Society for Information Science, 51(7):646–665, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Biron:1997:BRB

[Bir97] Paul V. Biron. Book review: Backpropagation: Theory, architectures, and applications. Journal of the American Society for Information Science, 48(1):88–89, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bixler:1996:BRBa

[Bix96a] Linda S. Bixler. Book review: Educational services in health sciences libraries. Volume 2: Current practice in health sciences librarianship. Journal of the American Society for Information Science, 47(12):955–956, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bixler:1996:BRBb

[Bix96b] Linda S. Bixler. Book review: Educational services in health sciences libraries. Volume 2: Current practice in health sciences librarianship. Journal of the American Society for Information Science, 47(12):955–956, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bell:1979:TEL

[BJ79] Colin Bell and Kevin P. Jones. Towards everyday language information retrieval systems via minicomputers. Journal of the American Society for Information Science, 30(6):334–339, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borko:1979:LER

[BK79] Harold Borko and Matthew Koll. Letter to the Editors: Relation of title length to number of authors. Journal of the American Society for Information Science, 30(3):175–176, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buell:1981:MWR

[BK81] Duncan A. Buell and Donald H. Kraft. A model for a weighted retrieval system. Journal of the American Society for Information Science, 32(3):211–216, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bayer:1996:BCS

[BK96] Bernard Bayer and Frederick G. Kilgour. Brief communications: Scholarly use of referenced information in physics journals. Journal of the American Society for Information Science, 47(2):170– 172, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bodoff:1998:PCB

[BK98] David Bodoff and Ajit Kambil. Partial coordination. I. The best of pre-coordination and post-coordination. Journal of the American Society for Information Science, 49(14):1254–1269, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1998:RCP

[BKR98] A. Bookstein, S. T. Klein, and T. Raita. Research: Clumping properties of content-bearing words. Journal of the American Society for Information Science, 49(2):102–114, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brazell:1972:LET

[BL72] Troy Brazell and F. F. Leimkuhler. Letters to the Editor: Title: Music media series. Journal of the American Society for Information Science, 23(4):286, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bearman:1985:MPL

[BL85] Toni Carbo Bearman and Lois F. Lunin. Miscellaneous: Perspectives on … The National Commission on Libraries and Information Science. Journal of the American Society for Information Science, 36(6):357, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blanken:1971:PIA

[Bla71] Robert R. Blanken. The preparation of international author indexes, with particular reference to the problems of transliteration, prefixes, and compound family names. Journal of the American Society for Information Science, 22(1):51–63, January 1971. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blair:1979:BRB

[Bla79] David C. Blair. Book review: Information Retrieval, 2nd ed. C. J. Van Rijsbergen. London: Butterworths; 1979: 208 pp. Price: $32.50. Journal of the American Society for Information Science, 30(6):374–375, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blair:1980:SBL

[Bla80] David C. Blair. Searching biases in large interactive document retrieval systems. Journal of the American Society for Information Science, 31(4):271–277, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blair:1993:BRB

[Bla93] David C. Blair. Book review: Text retrieval: Information first. Journal of the American Society for Information Science, 44(2): 113–115, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blair:1996:SRT

[Bla96] David C. Blair. STAIRS redux: Thoughts on the STAIRS evaluation, ten years after. Journal of the American Society for Information Science, 47(1):4–22, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blewett:1986:BRB

[Ble86] Daniel K. Blewett. Book review: Encyclopedia of Associations, 1986. Journal of the American Society for Information Science, 37(4):276–277, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bloom:1970:ACA

[Blo70] P. W. Bloom. Application of CAPRI to the ASIS file management exercise. Journal of the American Society for Information Science, 21(3):219–223, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bucy:1999:RFF

[BLPG99] Erik P. Bucy, Annie Lang, Robert F. Potter, and Maria Elizabeth Grabe. Research: Formal features of cyberspace: Relationships between Web page complexity and site traffic. Journal of the American Society for Information Science, 50(13):1246–1256, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blum:1971:TMI

[Blu71] Fred Blum. Two machine indexing projects at the Catholic University of America. Journal of the American Society for Information Science, 22(2):105–106, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blue:1989:BRBa

[Blu89a] Richard I. Blue. Book review: Scientific and technical libraries. Volume 1: Functions and management. Journal of the American Society for Information Science, 40(2):135–136, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Blue:1989:BRBb

[Blu89b] Richard I. Blue. Book review: Scientific and technical libraries. Volume 2: Special formats and subject areas. Journal of the American Society for Information Science, 40(2):135–137, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borman:1972:ISB

[BM72] Lorraine Borman and Benjamin Mittman. Interactive search of bibliographic data bases in an academic environment. Journal of the American Society for Information Science, 23(3):164–171, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borkowski:1975:SEB

[BM75] Casimir Borkowski and J. Sperling Martin. Structure, effectiveness and benefits of LEXtractor, an operational computer program for automatic extraction of case summaries and dispositions from court decisions. Journal of the American Society for Information Science, 26(2):94–102, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyle:1980:FCH

[BM80] Stephen O. Boyle and A. Patricia Miller. Feature comparison of an in-house information retrieval system with a commercial search service. Journal of the American Society for Information Science, 31(5):309–317, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1981:BCB

[BM81] Bert R. Boyce and David Martin. Brief communications: The Brillouin measure of an author’s contribution to a literature in psychology. Journal of the American Society for Information Science, 32(1):73–76, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1989:REP

[BM89] Bert R. Boyce and John P. McLain. Research: Entry point depth and online search using a controlled vocabulary. Journal of the American Society for Information Science, 40(4):273–276, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bassiouni:1995:RED

[BM95] Mostafa A. Bassiouni and Amar Mukherjee. Research: Efficient decoding of compressed data. Journal of the American Society for Information Science, 46(1):1–8, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bulick:1976:BCU

[BMFK76] Stephen Bulick, K. Leon Montgomery, John Fetterman, and Allen Kent. Brief communications: Use of library materials in terms of age. Journal of the American Society for Information Science, 27(3):175–178, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1984:DPP

[BMFS84] Bert R. Boyce, David Martin, Barbara Francis, and Mary Ellen Sievert. A drill and practice program for online retrieval. Journal of the American Society for Information Science, 35(2):129–134, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bell:1993:RDC

[BMNM^{+}93] Timothy C. Bell, Alistair Moffat, Craig G. Nevill-Manning, Ian H. Witten, and Justin Zobel. Research: Data compression in full-text retrieval systems. Journal of the American Society for Information Science, 44(9):508–531, October 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bell:1990:RAS

[BMSM90] D. A. Bell, F. J. McErlean, P. M. Stewart, and S. I. McClean. Research: Application of simulated annealing to clustering tuples in databases. Journal of the American Society for Information Science, 41(2):98–110, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Braam:1991:RMSa

[BMvR91a] Robert R. Braam, Henk F. Moed, and Anthony F. J. van Raan. Research: Mapping of science by combined co-citation and word analysis. I. Structural aspects. Journal of the American Society for Information Science, 42(4):233–251, May 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Braam:1991:RMSb

[BMvR91b] Robert R. Braam, Henk F. Moed, and Anthony F. J. van Raan. Research: Mapping of science by combined co-citation and word analysis. II: Dynamical aspects. Journal of the American Society for Information Science, 42(4):252–266, May 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bennion:1976:ERA

[BN76] Bruce C. Bennion and Laurence A. Nsuton. The epidemiology of research on “anomalous water”. Journal of the American Society for Information Science, 27(1):53–56, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burnett:1999:RCT

[BNP99] Kathleen Burnett, Kwong Bor Ng, and Soyeon Park. Research: A comparison of the two traditions of metadata development. Journal of the American Society for Information Science, 50(13): 1209–1217, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bishop:2000:DLS

[BNS^{+}00] Ann Peterson Bishop, Laura J. Neumann, Susan Leigh Star, Cecelia Merkel, Emily Ignacio, and Robert J. Sandusky. Digital libraries: Situating use in changing information infrastructure. Journal of the American Society for Information Science, 51(4): 394–413, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bodoff:1999:RRU

[Bod99] David Bodoff. Research: A re-unification of two competing models for document retrieval. Journal of the American Society for Information Science, 50(1):49–64, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bohnert:1990:LEP

[Boh90] Lea M. Bohnert. Letters to the Editor: In praise of power. Journal of the American Society for Information Science, 41(6):427, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bolner:1986:BRBa

[Bol86a] Smittie Bolner. Book review: Business publications index and abstracts. Journal of the American Society for Information Science, 37(4):278, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bolner:1986:BRBb

[Bol86b] Smittie Bolner. Book review: Business Publications Index and Abstracts. Detroit: Gale; 1983-Monthly with annual cumulations. Journal of the American Society for Information Science, 37(4): 278, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bolner:1992:BRB

[Bol92] Myrtle Smith Bolner. Book review: United States government information policies: Views and perspectives; and United States scientific and technical information policies: Views and perspectives. Journal of the American Society for Information Science, 43(6):455, July 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bommer:1975:OPO

[Bom75] Michael Bommer. Opinion Paper. Operations research in libraries: a critical assessment. Journal of the American Society for Information Science, 26(3):137–139, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bondi:1979:ETD

[Bon79] A. A. Bondi. On the economics and timing of data evaluation and estimation in an industrial setting. Journal of the American Society for Information Science, 30(6):360–363, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bonzi:1982:CLP

[Bon82] Susan Bonzi. Characteristics of a literature as predictors of relatedness between cited and citing works. Journal of the American Society for Information Science, 33(4):208–216, July 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bonzi:1989:BRB

[Bon89] Susan Bonzi. Book review: Semantic interpretation and the resolution of ambiguity. Journal of the American Society for Information Science, 40(6):438, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bonzi:1990:RSP

[Bon90] Susan Bonzi. Research: Syntactic patterns in scientific sublanguages: a study of four disciplines. Journal of the American Society for Information Science, 41(2):121–131, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1972:DH

[Boo72] Abraham Bookstein. Double hashing. Journal of the American Society for Information Science, 23(6):402–405, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1974:ARI

[Boo74a] Abraham Bookstein. Allocation of resources in an information system. Journal of the American Society for Information Science, 25(1):52–58, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1974:HCN

[Boo74b] Abraham Bookstein. Hash coding with a non-unique search key. Journal of the American Society for Information Science, 25(4): 232–236, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1978:BCP

[Boo78] A. Bookstein. Brief communications: On the perils of merging boolean and weighted retrieval systems. Journal of the American Society for Information Science, 29(3):156–158, May 1978. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1979:R

[Boo79] Abraham Bookstein. Relevance. Journal of the American Society for Information Science, 30(5):269–273, September 1979. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1980:FRA

[Boo80] Abraham Bookstein. Fuzzy requests: an approach to weighted boolean searches. Journal of the American Society for Information Science, 31(4):240–247, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1981:CTS

[Boo81] A. Bookstein. A comparison of two systems of weighted boolean retrieval. Journal of the American Society for Information Science, 32(4):275–279, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1983:IRS

[Boo83] Abraham Bookstein. Information retrieval: a sequential learning process. Journal of the American Society for Information Science, 34(5):331–342, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1990:RIDa

[Boo90a] A. Bookstein. Research: Informetric distributions, part I: Unified overview. Journal of the American Society for Information Science, 41(5):368–375, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1990:RIDb

[Boo90b] A. Bookstein. Research: Informetric distributions, part II: Resilience to ambiguity. Journal of the American Society for Information Science, 41(5):376–386, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1996:RB

[Boo96] A. Bookstein. Research: Bibliocryptography. Journal of the American Society for Information Science, 47(12):886–895, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1997:RID

[Boo97] A. Bookstein. Research: Informetric distributions. III. Ambiguity and randomness. Journal of the American Society for Information Science, 48(1):2–10, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Booker:2000:BR

[Boo00] Queen Esther Booker. Book reviews. Journal of the American Society for Information Science, 51(7):685–686, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borko:1973:BCF

[Bor73a] Harold Borko. Brief communications: FID Budapest 1972. Journal of the American Society for Information Science, 24(2):164– 165, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borko:1973:MAM

[Bor73b] Harold Borko. Meeting: The 1973 ASIS meeting: Purpose, programs and benefits. Journal of the American Society for Information Science, 24(4):239–241, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borko:1974:LTE

[Bor74] Harold Borko. Letters to the Editors: A reply to comments on the 1973 Annual Meeting. Journal of the American Society for Information Science, 25(5):340, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Ano74f].

Borko:1984:TLI

[Bor84] Harold Borko. Trends in library and information science education. Journal of the American Society for Information Science, 35(3):185–193, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1985:LEC

[Bor85a] Christine L. Borgman. Letters to the Editors: Citations Format. Journal of the American Society for Information Science, 36(6): 420, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bortnick:1985:NII

[Bor85b] Jane Bortnick. National and international information policy. Journal of the American Society for Information Science, 36(3): 164–168, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1986:BRB

[Bor86a] Christine L. Borgman. Book review: Retrospective conversion. Journal of the American Society for Information Science, 37(2): 91, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1986:RWO

[Bor86b] Christine L. Borgman. Research: Why are online catalogs hard to use? Lessons learned from information-retrieval studies. Journal of the American Society for Information Science, 37(6):387–400, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borko:1987:LEE

[Bor87] Harold Borko. Letters to the Editor: Editorial on historical notes. Journal of the American Society for Information Science, 38(5): 387, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borko:1990:BRB

[Bor90] Harold Borko. Book review: Expert systems for information management. Journal of the American Society for Information Science, 41(3):210, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borko:1993:BRB

[Bor93] Harold Borko. Book review: Critical issues in the information age. Journal of the American Society for Information Science, 44 (2):113, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1994:BRB

[Bor94] Christine L. Borgman. Book review: The whole Internet user’s guide & catalog. Journal of the American Society for Information Science, 45(2):117–118, March 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1996:WOC

[Bor96] Christine L. Borgman. Why are online catalogs still hard to use? Journal of the American Society for Information Science, 47(7): 493–503, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1997:LER

[Bor97] Christine L. Borgman. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 48(10):964– 964c, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1999:MRR

[Bor99] Christine L. Borgman. In memoriam: Robert R. Korfhage: a personal remembrance from the 1970s. Journal of the American Society for Information Science, 50(4):289–290, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:2000:BR

[Bor00] Christine L. Borgman. Book review. Journal of the American Society for Information Science, 51(1):90–91, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boss:1987:BRB

[Bos87] Richard W. Boss. Book review: Financing information services. Journal of the American Society for Information Science, 38(3): 222–223, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bottle:1970:TIA

[Bot70] Robert T. Bottle. Title indexes as alerting services in the chemical and life sciences. Journal of the American Society for Information Science, 21(1):16–21, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bottle:1973:CEW

[Bot73] R. T. Bottle. Collective Experimental Work in Information Science Teaching. Journal of the American Society for Information Science, 24(1):54–56, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bourne:1980:LSH

[Bou80] Charles P. Bourne. On-line systems: History, technology, and economics. Journal of the American Society for Information Science, 31(3):155–160, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1985:BRBa

[Boy85a] Bert Boyce. Book review: Automated circulation: an examination of choices. Joseph R. Matthews and Kevin E. Hegarty, Eds., Chicago: American Library Association; 1984: 126 pp. Journal of the American Society for Information Science, 36(2):137, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1985:BRBb

[Boy85b] Bert R. Boyce. Book review: The joy of computer communication. William J. Cook. New York: Dell Publishing Co.; 1984: 128 pp. Price: $5.95. ISBN 0-440-54412-2. Journal of the American Society for Information Science, 36(2):137, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1986:BRBe

[Boy86a] Bert R. Boyce. Book review: Authority work: the creation, use, maintenance, and evaluation of authority records and files. Journal of the American Society for Information Science, 37(4):272, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1986:BRBd

[Boy86b] Bert R. Boyce. Book review: Fee-based services in sci-tech libraries. Journal of the American Society for Information Science, 37(3):176, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1986:BRBg

[Boy86c] Bert R. Boyce. Book review: How to write a usable user manual. Journal of the American Society for Information Science, 37(4): 277–278, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1986:BRBb

[Boy86d] Bert R. Boyce. Book review: Online database search directory. Journal of the American Society for Information Science, 37(3): 170, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1986:BRBc

[Boy86e] Bert R. Boyce. Book review: Online information systems use and operating characteristics, limitations, and design alternatives. Journal of the American Society for Information Science, 37(3):171–172, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1986:BRBf

[Boy86f] Bert R. Boyce. Book review: Redesign of catalogs and indexes for improved online subject access: Selected papers of Pauline A. Cochrane. Journal of the American Society for Information Science, 37(4):275, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1986:BRBa

[Boy86g] Bert R. Boyce. Book review: Subject access systems, alternatives in design. Journal of the American Society for Information Science, 37(2):90–91, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1987:BRBb

[Boy87a] Bert R. Boyce. Book review: Online information retrieval concepts, principles, and techniques. Journal of the American Society for Information Science, 38(4):316, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1987:BRBa

[Boy87b] Bert R. Boyce. Book review: Vocabulary control for information retrieval. Journal of the American Society for Information Science, 38(4):315, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1988:BRB

[Boy88] Bert R. Boyce. Book review: Networks and data communication. Journal of the American Society for Information Science, 39(1): 24, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1989:BRBa

[Boy89a] Bert R. Boyce. Book review: Online bibliographic databases: a directory and sourcebook. Journal of the American Society for Information Science, 40(2):134, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1989:BRBb

[Boy89b] Bert R. Boyce. Book review: Online retrieval analysis and strategy. Journal of the American Society for Information Science, 40 (6):442–443, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1990:BRBb

[Boy90a] Bert R. Boyce. Book review: Concepts of information retrieval and automatic text processing: the transformation analysis, and retrieval of information by computer. Journal of the American Society for Information Science, 41(2):150–151, March 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1990:BRBa

[Boy90b] Bert R. Boyce. Book review: Thesauri used in online databases, an analytic guide. Journal of the American Society for Information Science, 41(1):71, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1991:BRB

[Boy91] Bert R. Boyce. Book review: Indexing: the state of our knowledge and the state of our ignorance: Proceedings of the 20th Annual Meeting of the American Society of Indexers. Journal of the American Society for Information Science, 42(10):757, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1992:BRB

[Boy92] Bert R. Boyce. Book review: Indexing and abstracting in theory and practice. Journal of the American Society for Information Science, 43(6):456, July 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:MMH

[Boy95a] Bert R. Boyce. In memoriam: Miranda Hsu-Yuang Pao. Journal of the American Society for Information Science, 46(10):723, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:Ia

[Boy95b] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 46(7):477, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:Ib

[Boy95c] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 46(8):561, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:RIa

[Boy95d] Bert R. Boyce. Research: In this issue. Journal of the American Society for Information Science, 46(3):161, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:RIb

[Boy95e] Bert R. Boyce. Research: In this issue. Journal of the American Society for Information Science, 46(4):243, May 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:RIc

[Boy95f] Bert R. Boyce. Research: In this issue. Journal of the American Society for Information Science, 46(5):321, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:RId

[Boy95g] Bert R. Boyce. Research: In this issue. Journal of the American Society for Information Science, 46(6):399, July 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1995:RIe

[Boy95h] Bert R. Boyce. Research: In this issue. Journal of the American Society for Information Science, 46(9):645, October 1995. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:Ia

[Boy96a] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(2):107, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:Ib

[Boy96b] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(3):183, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:Ic

[Boy96c] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(4):301, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:Id

[Boy96d] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(5):337, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:Ie

[Boy96e] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(6):411, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:If

[Boy96f] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(8):585, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:Ig

[Boy96g] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(10):729, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1996:Ih

[Boy96h] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 47(12):885, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ia

[Boy97a] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(1):1, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ib

[Boy97b] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(2):95, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ic

[Boy97c] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(3):189, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Id

[Boy97d] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(5):381, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ie

[Boy97e] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(6):483, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:If

[Boy97f] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(7):577, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ig

[Boy97g] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(8):673, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ih

[Boy97h] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(9):771, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ii

[Boy97i] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(10):865, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1997:Ij

[Boy97j] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 48(12):1085, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ia

[Boy98a] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(1):1, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ib

[Boy98b] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(2):101, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ic

[Boy98c] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(4):303, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Id

[Boy98d] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(5):393, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ie

[Boy98e] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(6):485, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:If

[Boy98f] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(8):673, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ig

[Boy98g] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(9):769, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ih

[Boy98h] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(10):865, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ii

[Boy98i] Bert R. Boyce. In This Issue. Journal of the American Society for Information Science, 49(13):1143, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1998:Ij

[Boy98j] Bert R. Boyce. In this issue. Journal of the American Society for Information Science, 49(14):1245, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIb

[Boy99a] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(2):97, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIc

[Boy99b] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(3):193, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EId

[Boy99c] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(4):287, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIe

[Boy99d] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(5):385, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIf

[Boy99e] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(6):481–482, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIg

[Boy99f] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(7):555, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIi

[Boy99g] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(9):735, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIj

[Boy99h] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(10):843–844, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIk

[Boy99i] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(13):1163, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIl

[Boy99j] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 50(14):1263, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIa

[Boy99k] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science, 50(1):4, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:1999:EIh

[Boy99l] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science, 50(8):637–638, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:EIa

[Boy00a] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 51(2):93–94, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:EIb

[Boy00b] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 51(5):415–416, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:EIc

[Boy00c] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 51(6):497, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:EId

[Boy00d] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 51(8):689–690, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:EIe

[Boy00e] Bert R. Boyce. Editorial: In this issue. Journal of the American Society for Information Science, 51(10):885–886, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:EIf

[Boy00f] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science, 51(13):1157–1158, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:EIg

[Boy00g] Bert R. Boyce. Editorials: In this issue. Journal of the American Society for Information Science, 51(14):1251, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:IIa

[Boy00h] Bert R. Boyce. In this issue: In this issue. Journal of the American Society for Information Science, 51(1):1, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:IIb

[Boy00i] Bert R. Boyce. In this issue: In this issue. Journal of the American Society for Information Science, 51(7):585–586, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:IIc

[Boy00j] Bert R. Boyce. In this issue: In this issue. Journal of the American Society for Information Science, 51(9):791–792, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:IId

[Boy00k] Bert R. Boyce. In this issue: In this issue. Journal of the American Society for Information Science, 51(11):969–970, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boyce:2000:IIe

[Boy00l] Bert R. Boyce. In this issue: In this issue. Journal of the American Society for Information Science, 51(12):1067–1068, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bottle:1970:BCP

[BP70] Robert T. Bottle and Cynthia I. Preibish. Brief communications: The proposed KWIC index for psychology: an experimental test of its effectiveness. Journal of the American Society for Information Science, 21(6):427–428, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bordogna:1993:RFL

[BP93] Gloria Bordogna and Gabriella Pasi. Research: A fuzzy linguistic approach generalizing Boolean Information Retrieval: a model and its evaluation. Journal of the American Society for Information Science, 44(2):70–82, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bordogna:1998:I

[BP98] Gloria Bordogna and Gabriella Pasi. Introduction. Journal of the American Society for Information Science, 49(3):193–194, March 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bromberg:1971:LEC

[BPK^{+}71] Erik Bromberg, Michael W. Page, Howard M. Kibbel, Hans Wellisch, and Harvey Marron. Letters to the Editors: Communicating to communicators. Journal of the American Society for Information Science, 22(3):225–227, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belkin:1976:ISP

[BR76] Nicholas J. Belkin and Stephen E. Robertson. Information science and the phenomenon of information. Journal of the American Society for Information Science, 27(4):197–204, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bourne:1980:ETC

[BR80a] Charles P. Bourne and Jo Robinson. Education and training for computer-based reference services: Review of training efforts to date. Journal of the American Society for Information Science, 31(1):25–35, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brittain:1980:RSS

[BR80b] J. Michael Brittain and Stephen A. Roberts. Rationalization of secondary services: Measurement of coverage of primary journals and overlap between services. Journal of the American Society for Information Science, 31(3):131–142, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1992:RCI

[BR92] Christine L. Borgman and Ronald E. Rice. Research: The convergence of information science and communication: a bibliometric analysis. Journal of the American Society for Information Science, 43(6):397–411, July 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ball:1994:PLR

[BR94] Marion J. Ball and Errol L. Reese. Perspectives on … the redesign/reengineering of an information services division in a major health science institution: Information services at the University of Maryland at Baltimore: Giving the vision life. Journal of the American Society for Information Science, 45(5):326–330, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baumgras:1995:POC

[BR95a] Jan L. Baumgras and Anne E. Rogers. Perspectives on … the Chemist’s workstation: Chemical structures at the desktop: Integrating drawing tools with on-line registry files. Journal of the American Society for Information Science, 46(8):623–631, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burrell:1995:RFC

[BR95b] Quentin Burrell and Ronald Rousseau. Research: Fractional counts for authorship attribution: a numerical study. Journal of the American Society for Information Science, 46(2):97–102, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baldwin:1997:RIS

[BR97] Nancy Sadler Baldwin and Ronald E. Rice. Research: Information-seeking behavior of securities analysts: Individual and institutional influences, information sources and channels, and outcomes. Journal of the American Society for Information Science, 48(8):674–693, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brandhorst:1970:BCT

[Bra70] W. T. Brandhorst. Brief communications: A table of set theory notations. Journal of the American Society for Information Science, 21(6):428, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Braunstein:1977:ERP

[Bra77] Yale M. Braunstein. An economic rationale for page and submission charges by academic journals. Journal of the American Society for Information Science, 28(6):355–358, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brandhorst:1979:BCA

[Bra79] Ted Brandhorst. Brief communications: ANSI 239 Romanization Standards and “reversibility”: A dialog to arrive at a policy. Journal of the American Society for Information Science, 30(1): 55–59, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bracken:1987:PLF

[Bra87] Darcia D. Bracken. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Section IV. NTIS — GPO: Federal information disseminators. The National Technical Information Service: a federal resource for health information and services. Journal of the American Society for Information Science, 38(1):65–67, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Braunstein:1989:REI

[Bra89] Yale M. Braunstein. Research: Economics of intellectual property rights in the international arena. Journal of the American Society for Information Science, 40(1):12–16, January 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Braunstein:1992:LER

[Bra92] Yale M. Braunstein. Letters to the Editor: Regarding E. Garfield’s letter. Journal of the American Society for Information Science, 43(1):96, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Braman:1994:STI

[Bra94] Sandra Braman. The autopoietic state: Communication and democratic potential in the net. Journal of the American Society for Information Science, 45(6):358–368, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brenner:1979:OPE

[Bre79] Everett H. Brenner. Opinion paper: Euronet and its effects on the U.S. information market. Journal of the American Society for Information Science, 30(1):5–8, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Breton:1991:RII

[Bre91] Ernest J. Breton. Research: Indexing for invention. Journal of the American Society for Information Science, 42(3):173–177, April 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Briner:1972:LEL

[Bri72] L. L. Briner. Letters to the Editor: Linguistics and information science. Journal of the American Society for Information Science, 23(6):408, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brittain:1984:ISS

[Bri84] J. Michael Brittain. Internationality of the social sciences: Implications for information transfer. Journal of the American Society for Information Science, 35(1):11–18, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bricker:1991:RDD

[Bri91] Robert Bricker. Research: Deriving disciplinary structures: Some new methods, models, and an illustration with accounting. Journal of the American Society for Information Science, 42(1):27–35, January 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brodie:1970:EKI

[Bro70a] Nancy E. Brodie. Evaluation of a KWIC index for library literature. Journal of the American Society for Information Science, 21(1):22–28, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bromberg:1970:BCE

[Bro70b] Erik Bromberg. Brief communications: The economics of information. Journal of the American Society for Information Science, 21(1):95, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brookes:1970:OSL

[Bro70c] B. C. Brookes. Obsolescence of special library periodicals: Sampling errors and utility contours. Journal of the American Society for Information Science, 21(5):320–329, September 1970. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brookes:1979:LEB

[Bro79] B. C. Brookes. Letters to the Editor: The Bradford Law: a new calculus for the social sciences? Journal of the American Society for Information Science, 30(4):233–234, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brookes:1980:MIS

[Bro80a] Bertram C. Brookes. Measurement in information science: Objective and subjective metrical space. Journal of the American Society for Information Science, 31(4):248–255, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brown:1980:BCH

[Bro80b] Pauline Brown. Brief communication: The half-life of the chemical literature. Journal of the American Society for Information Science, 31(1):61–63, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Broadus:1983:IVB

[Bro83] Robert N. Broadus. An investigation of the validity of bibliographic citations. Journal of the American Society for Information Science, 34(2):132–135, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1985:PAP

[Bro85] Terrence A. Brooks. Private acts and public objects: an investigation of citer motivations. Journal of the American Society for Information Science, 36(4):223–229, July 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1986:BRB

[Bro86a] Terrence A. Brooks. Book review: Videodisc and optical digital disk technologies and their applications in libraries. Journal of the American Society for Information Science, 37(2):89, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1986:REC

[Bro86b] Terrence A. Brooks. Research: Evidence of complex citer motivations. Journal of the American Society for Information Science, 37(1):34–36, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Broadus:1987:REA

[Bro87] Robert N. Broadus. Research: Early approaches to bibliometrics. Journal of the American Society for Information Science, 38(2): 127–129, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1988:RRB

[Bro88] Terrence A. Brooks. Research: Repetitive bibliographical information in relational databases. Journal of the American Society for Information Science, 39(1):2–7, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1990:LER

[Bro90a] Terrence A. Brooks. Letters to the Editor: Review of “The Bibliometrics Toolbox”. Journal of the American Society for Information Science, 41(6):427, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1990:RCC

[Bro90b] Terrence A. Brooks. Research: Clustering in comprehensive bibliographies and related literatures. Journal of the American Society for Information Science, 41(3):183–192, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1993:BRB

[Bro93a] Terrence A. Brooks. Book review: Information retrieval interaction. Journal of the American Society for Information Science, 44(5):303–304, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1993:RAR

[Bro93b] Terrence A. Brooks. Research: All the right descriptors: a test of the strategy of unlimited aliasing. Journal of the American Society for Information Science, 44(3):137–147, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1994:BRB

[Bro94] Terrence A. Brooks. Book review: Information retrieval today. Journal of the American Society for Information Science, 45(10): 800–801, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1995:BRB

[Bro95a] Terrence A. Brooks. Book review: Information retrieval: New systems and current research. Proceedings of the 15th research colloquium of the British Computer Society Information Retrieval Specialist Group. Journal of the American Society for Information Science, 46(9):713–714, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1995:RPW

[Bro95b] Terrence A. Brooks. Research: People, words, and perceptions: a phenomenological investigation of textuality. Journal of the American Society for Information Science, 46(2):103–115, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1996:BRB

[Bro96] Terrence A. Brooks. Book review: Designing information technology in the postmodern age: From method to metaphor. Journal of the American Society for Information Science, 47(9):719–720A, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brown:1997:BRB

[Bro97] Mary E. Brown. Book review: Electronic publishing and libraries, planning for the impact and growth to 2003. Journal of the American Society for Information Science, 48(3):276–278, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1998:BRBa

[Bro98a] Terrence A. Brooks. Book review: Information seeking in context. Journal of the American Society for Information Science, 49(6): 573–574, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1998:BRBb

[Bro98b] Terrence A. Brooks. Book review: Knowledge, concepts, and categories. Journal of the American Society for Information Science, 49(7):671–672, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1998:ROF

[Bro98c] Terrence A. Brooks. Research: Orthography as a fundamental impediment to online information retrieval. Journal of the American Society for Information Science, 49(8):731–741, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1999:CPS

[Bro99a] Terrence A. Brooks. Call for papers: Special topic issue of JASIS: Information science at the Millennium. Journal of the American Society for Information Science, 50(8):733, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:1999:RPI

[Bro99b] Terrence A. Brooks. Research: Postmodern information science and its “journal”. Journal of the American Society for Information Science, 50(11):1030–1031, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brown:1999:RIS

[Bro99c] Cecelia M. Brown. Research: Information seeking behavior of scientists in the electronic information age: Astronomers, chemists, mathematicians, and physicists. Journal of the American Society for Information Science, 50(10):929–943, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:2000:BR

[Bro00a] Terrance A. Brooks. Book reviews. Journal of the American Society for Information Science, 51(5):490, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:2000:BRBb

[Bro00b] Terrence A. Brooks. Book review: Information visualization: Perception for design, by Colin Ware. Journal of the American Society for Information Science, 51(11):1062–1063, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:2000:BRBc

[Bro00c] Terrence A. Brooks. Book review: Sorting things out: Classification and its consequences, by Geoffrey C. Bowker and Susan Leigh Star. Journal of the American Society for Information Science, 51(12):1149–1150, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:2000:BRBa

[Bro00d] Terrence A. Brooks. Book review: Web style guide: Basic design principles for creating web sites, by Patrick J. Lynch and Sarah Horton. Journal of the American Society for Information Science, 51(10):964–965, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:2000:RLL

[Bro00e] Terrence A. Brooks. Research letter: Letter to the Editor (reply). Journal of the American Society for Information Science, 51(5): 496, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brooks:2000:SCH

[Bro00f] Terrence A. Brooks. Short communication: How good are the best papers of JASIS? Journal of the American Society for Information Science, 51(5):485–486, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beaulieu:1996:EIS

[BRR96] Micheline Beaulieu, Stephen Robertson, and Edie Rasmussen. Evaluating interactive systems in TREC. Journal of the American Society for Information Science, 47(1):85–94, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Brunenkant:1978:LED

[Bru78] Edward J. Brunenkant. Letters to the Editor: Data bases-history of developments and trends. Journal of the American Society for Information Science, 29(3):165, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bruce:1994:RCV

[Bru94] Harry W. Bruce. Research: A cognitive view of the situational dynamism of user-centered relevance estimation. Journal of the American Society for Information Science, 45(3):142–148, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bruce:1998:RUS

[Bru98] Harry Bruce. Research: User satisfaction with information seeking on the Internet. Journal of the American Society for Information Science, 49(6):541–556, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bryan:1970:SSI

[Bry70] Glenn L. Bryan. Student-to-student interaction in computer time-sharing systems. Journal of the American Society for Information Science, 21(4):279–284, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1974:PMA

[BS74] Abraham Bookstein and Don R. Swanson. Probabilistic models for automatic indexing. Journal of the American Society for Information Science, 25(5):312–316, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bookstein:1975:DTF

[BS75] Abraham Bookstein and Don R. Swanson. A decision theoretic foundation for indexing. Journal of the American Society for Information Science, 26(1):45–50, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barnett:1988:RDC

[BS88] George A. Barnett and Gary Siegel. Research: The diffusion of computer-assisted legal research systems. Journal of the American Society for Information Science, 39(4):224–234, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barr:1991:RBM

[BS91a] Aiala Barr and Herbert S. Sichel. Research: A bivariate model to predict library circulation. Journal of the American Society for Information Science, 42(8):546–553, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beath:1991:PLI

[BS91b] Cynthia Mathis Beath and Detmar W. Straub. Perspectives on … integrated information centers within academic environments: Department level information resource management: a theoretical argument for a decentralized approach. Journal of the American Society for Information Science, 42(2):124–127, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Borgman:1992:RGS

[BS92] Christine L. Borgman and Susan L. Siegfried. Research: Getty’s Synoname^{T M} and its cousins: a survey of applications of personal name-matching algorithms. Journal of the American Society for Information Science, 43(7):459–476, August 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bayer:1990:POA

[BSM90] Alan E. Bayer, John C. Smart, and Gerald W. McLaughlin. Perspectives on … author cocitation analysis: Mapping intellectual structure of a scientific subfield through author cocitations. Journal of the American Society for Information Science, 41(6):444– 452, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bollmann-Sdorra:1993:RDA

[BSR93] Peter Bollmann-Sdorra and Vijay V. Raghavan. Research: On the delusiveness of adopting a common space for modeling IR objects: Are queries documents? Journal of the American Society for Information Science, 44(10):579–587, December 1993. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bollmann-Sdorra:1994:LEV

[BSR94] Peter Bollmann-Sdorra and Vijay V. Raghavan. Letters to the Editor: Vector processing model. Journal of the American Society for Information Science, 45(7):540–541, August 1994. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bollmann-Sdorra:1998:RNT

[BSR98] Peter Bollmann-Sdorra and Vijay V. Raghavan. Research: On the necessity of term dependence in a query space for weighted retrieval. Journal of the American Society for Information Science, 49(13):1161–1168, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bruza:2000:RAC

[BSW00] P. D. Bruza, D. W. Song, and K. F. Wong. Research: Aboutness from a commonsense perspective. Journal of the American Society for Information Science, 51(12):1090–1105, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bakar:2000:ERE

[BSY00] Zainab Abu Bakar, Tengku Mohd T. Sembok, and Mohammed Yusoff. An evaluation of retrieval effectiveness using spellingcorrection and string-similarity matching methods on Malay texts. Journal of the American Society for Information Science, 51(8):691–706, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beheshti:1984:RMM

[BT84] Jamshid Beheshti and Jean M. Tague. Research: Morse’s Markov model of book use revisited. Journal of the American Society for Information Science, 35(5):259–267, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burke:1999:BCI

[BT99] M. E. Burke and Karen Tulett. Brief communication: Impact of information needs an organizational design. Journal of the American Society for Information Science, 50(4):380–381, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bearman:2000:IWM

[BT00] David Bearman and Jennifer Trant. Introduction: When museum informatics meets the World Wide Web, it generates energy. Journal of the American Society for Information Science, 51(1):3–4, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baron:1996:RLT

[BTSKC96] Lisa Baron, Jean Tague-Sutcliffe, Mark T. Kinnucan, and Tom Carey. Research: Labeled, typed links as cues when reading hypertext documents. Journal of the American Society for Information Science, 47(12):896–908, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1975:NGR

[Buc75] Michael K. Buckland. Notes on the Gupta–Ravindran optimal storage model. Journal of the American Society for Information Science, 26(6):351–352, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [GR74].

Buckland:1979:TSA

[Buc79] Michael K. Buckland. On types of search and the allocation of library resources. Journal of the American Society for Information Science, 30(3):143–147, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1989:RIH

[Buc89] Michael K. Buckland. Research: Information handling, organizational structure, and power. Journal of the American Society for Information Science, 40(5):329–333, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1991:LER

[Buc91a] Michael Buckland. Letters to the Editor: Rejoinder to “Information as thing”. Journal of the American Society for Information Science, 42(10):758, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1991:BRB

[Buc91b] Michael K. Buckland. Book review: Information quality: Definitions and dimensions; proceedings of a NORDINFO seminar, Royal School of Librarianship, Copenhagen, 1989. Journal of the American Society for Information Science, 42(6):463, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1991:PLI

[Buc91c] Michael K. Buckland. Perspectives on … imaging. Advanced applications: Information retrieval of more than text. Journal of the American Society for Information Science, 42(8):586–588, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1991:RIT

[Buc91d] Michael K. Buckland. Research: Information as thing. Journal of the American Society for Information Science, 42(5):351–360, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1992:LER

[Buc92a] Michael K. Buckland. Letters to the Editor: Rejoinder to Dean Koenig’s letter. Journal of the American Society for Information Science, 43(10):678, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1992:REG

[Buc92b] Michael K. Buckland. Research: Emanuel Goldberg, electronic document retrieval, and Vannevar Bush’s Memex. Journal of the American Society for Information Science, 43(4):284–294, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1992:RRP

[Buc92c] Michael K. Buckland. Research: The retrieval problem in photography (1932). Emanuel Goldberg (1881–1970). Journal of the American Society for Information Science, 43(4):295–298, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1995:BCC

[Buc95] Michael K. Buckland. Brief communications: The centenary of “Madame Documentation”: Suzanne Briet, 1894–1989. Journal of the American Society for Information Science, 46(3):235–237, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1997:WD

[Buc97] Michael K. Buckland. What is a “document”? Journal of the American Society for Information Science, 48(9):804–809, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1998:BRB

[Buc98] Michael Buckland. Book review: Technology and management in library and information services. Journal of the American Society for Information Science, 49(4):385–386, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buckland:1999:RLI

[Buc99] Michael Buckland. Research: The landscape of information science: The American Society for Information Science at 62. Journal of the American Society for Information Science, 50(11):970– 974, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Budd:1991:BRB

[Bud91] John Budd. Book review: The role of the academic reference librarian. Journal of the American Society for Information Science, 42(2):112, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buell:1985:PIR

[Bue85] Duncan A. Buell. A problem in information retrieval with fuzzy sets. Journal of the American Society for Information Science, 36 (6):398–401, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buell:1986:BRBa

[Bue86a] Duncan A. Buell. Book review: Supercomputers and parallel computation. Journal of the American Society for Information Science, 37(2):91–92, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buell:1986:BRBb

[Bue86b] Duncan A. Buell. Book review: The first freedom today. Journal of the American Society for Information Science, 37(5):349–351, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buntrock:1980:LES

[Bun80] Robert E. Buntrock. Letters to the Editor: Searching the chemical literature. Journal of the American Society for Information Science, 31(5):379–380, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burchinal:1970:RFG

[Bur70] Lee G. Burchinal. The role of the Federal Government in information systems in education. Journal of the American Society for Information Science, 21(4):274–278, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burns:1971:LES

[Bur71] Robert W. Burns, Jr. Letters to the Editors: Systems definitions. Journal of the American Society for Information Science, 22(1): 65, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burress:1980:BCA

[Bur80] Elaine P. Burress. Brief communication: Automated indexing versus kwoc: a performance comparison. Journal of the American Society for Information Science, 31(1):60–61, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burns:1981:ISD

[Bur81a] Christopher Burns. Information storage and display. Journal of the American Society for Information Science, 32(2):141–147, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burton:1981:FRT

[Bur81b] Hilary D. Burton. FAMULUS revisited: Ten years of personal information systems. Journal of the American Society for Information Science, 32(6):440–443, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burton:1985:OPC

[Bur85] Hilary D. Burton. Opinion papers: The changing environment of personal information systems. Journal of the American Society for Information Science, 36(1):48–52, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burrell:1990:RUG

[Bur90] Quentin L. Burrell. Research: Using the Gamma-Poisson model to predict library circulations. Journal of the American Society for Information Science, 41(3):164–170, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burke:1992:ROM

[Bur92a] Colin Burke. Research: The other Memex: the tangled career of Vannevar Bush’s Information Machine, The Rapid Selector. Journal of the American Society for Information Science, 43(10): 648–657, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burrell:1992:BCN

[Bur92b] Quentin L. Burrell. Brief communications: A note on a result of Rousseau for concentration measures. Journal of the American Society for Information Science, 43(6):452–454, July 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burrell:1993:LEJ

[Bur93] Quentin Burrell. Letters to the Editor: Journal productivity distribution: Quantitative study of dynamic behavior. Journal of the American Society for Information Science, 44(3):181–182, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burrell:1994:LEK

[Bur94] Quentin L. Burrell. Letters to the Editor: The Kolmogorov– Smirnov test and rank-frequency distributions. Journal of the American Society for Information Science, 45(1):59, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burgin:1995:RRE

[Bur95] Robert Burgin. Research: The retrieval effectiveness of five clustering algorithms as a function of indexing exhaustivity. Journal of the American Society for Information Science, 46(8):562–572, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burge:1996:IDL

[Bur96] Elizabeth J. Burge. Instances of distance learning: Inside-out thinking about distance teaching: Making sense of reflective practice. Journal of the American Society for Information Science, 47 (11):843–848, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burgin:1999:RMC

[Bur99] Robert Burgin. Research: The Monte Carlo method and the evaluation of retrieval system performance. Journal of the American Society for Information Science, 50(2):181–191, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Burchinal:2000:RTT

[Bur00] Lee G. Burchinal. Research: The tale of two ERICS: Factors influencing the development of the first ERIC and its transformation into a national system. Journal of the American Society for Information Science, 51(6):567–575, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Busch:1995:BRB

[Bus95] Joseph A. Busch. Book review: Abstracting, information retrieval and the humanities: Providing access to historical literature. Journal of the American Society for Information Science, 46(3):238, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Buttler:1999:PLC

[But99] Dwayne K. Buttler. Perspective on … copyright and fair-use guidelines for education and libraries: CONFU-sed: security, safe harbors, and fair-use guidelines. Journal of the American Society for Information Science, 50(14):1308–1312, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Bernstein:1984:TNL

[BW84] Lionel M. Bernstein and Robert E. Williamson. Testing of a natural language retrieval system for a full text knowledge base. Journal of the American Society for Information Science, 35(4): 235–247, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Beeson:1985:EGA

[BW85] Betty Spillers Beeson and R. Ann Williams. The effects of gender and age on preschool children’s choice of the computer as a childselected activity. Journal of the American Society for Information Science, 36(5):339–341, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Belzer:1975:CIS

[BWKG75] Jack Belzer, James Williams, John Kronebusch, and A. B. Gupta. Curricula in information science: Four year progress report. Journal of the American Society for Information Science, 26(1):17–32, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Barker:1972:REE

[BWV72] Frances H. Barker, Barry K. Wyatt, and Douglas C. Veal. Report on the evaluation of an experimental computer-based currentawareness service for chemists. Journal of the American Society for Information Science, 23(2):85–99, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Baeza–Yates:2000:RBA

[BYN00] Ricardo Baeza-Yates and Gonzalo Navarro. Research: Block addressing indices for approximate text retrieval. Journal of the American Society for Information Science, 51(1):69–82, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Byrne:1975:RET

[Byr75] Jerry R. Byrne. Relative effectiveness of titles, abstracts, and subject headings for machine retrieval from the COMPENDEX services. Journal of the American Society for Information Science, 26(4):223–229, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Byrne:1976:LER

[Byr76] Jerry R. Byrne. Letters to the Editor: Retrieval from Engineering Index: Dear Sir. Journal of the American Society for Information Science, 27(1):78, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Boll:1985:LEM

[BZ85] John J. Boll and Douglas Zweizig. Letters to the Editors: Mapping a curriculum by computer. Journal of the American Society for Information Science, 36(5):352–353, September 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cochrane:1980:CAE

[CA80] Glynn Cochrane and Pauline Atherton. The cultural appraisal of efforts to alleviate information inequity. Journal of the American Society for Information Science, 31(4):283–292, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cagan:1970:BCTb

[Cag70a] Carl Cagan. Brief communication: Term-term correlation for large matrices. Journal of the American Society for Information Science, 21(3):228, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cagan:1970:BCTa

[Cag70b] Carl Cagan. Brief communications: Term-term correlation for large matrices. Journal of the American Society for Information Science, 21(2):163, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cagan:1970:HAD

[Cag70c] Carl Cagan. A highly associative document retrieval system. Journal of the American Society for Information Science, 21(5): 330–337, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Calabrese:1989:BRBb

[Cal89a] Andrew M. Calabrese. Book review: Information technology: a Luddite analysis. Journal of the American Society for Information Science, 40(2):141–142, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Calabrese:1989:BRBa

[Cal89b] Andrew M. Calabrese. Book review: Misunderstanding media. Journal of the American Society for Information Science, 40(2): 139–140, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Calarco:2000:BRB

[Cal00] Pascal V. Calarco. Book review: Developer’s guide to the Java Web server: Building effective and scalable server-side applications, by Dan Woods, Larne Pekowsky, and Tom Snee. Journal of the American Society for Information Science, 51(10):965–966, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Campanario:1996:RRR

[Cam96a] Juan Miguel Campanario. Research: Have referees rejected some of the most-cited articles of all times? Journal of the American Society for Information Science, 47(4):302–310, April 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Campanario:1996:RCJ

[Cam96b] Juan Miguel Campanario. Research: The competition for journal space among referees, editors, and other authors and its influence on journals’ impact factors. Journal of the American Society for Information Science, 47(3):184–192, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Campbell:1997:BRB

[Cam97] Karen Campbell. Book review: World Wide Web journal: Volume 1, issue 2, spring 1996. Key specifications of the World Wide Web. Journal of the American Society for Information Science, 48(5):476–477, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cano:1989:RCB

[Can89] V. Cano. Research: Citation behavior: Classification, utility, and location. Journal of the American Society for Information Science, 40(4):284–290, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

CannCasciato:1995:LEC

[Can95] Daniel CannCasciato. Letters to the Editor: Cataloging article concerns. Journal of the American Society for Information Science, 46(10):798, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cano:1999:RBO

[Can99] V. Cano. Research: Bibliometric overview of Library and Information Science Research in Spain. Journal of the American Society for Information Science, 50(8):675–680, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carmon:1974:DFI

[Car74] James L. Carmon. A discriminant function index for information system evaluation. Journal of the American Society for Information Science, 25(2):118–122, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carroll:1975:UST

[Car75] John M. Carroll. Use of synthetic text in the evaluation of statistically based keyword-selection strategies. Journal of the American Society for Information Science, 26(4):261–262, July 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carroll:1976:PPC

[Car76] John M. Carroll. Prospects for Parallelism and the Computer Crunch. Journal of the American Society for Information Science, 27(1):63–69, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carpenter:1979:BCS

[Car79] Mark P. Carpenter. Brief communications: Similarity of Pratt’s measure of class concentration to the Gini index. Journal of the American Society for Information Science, 30(2):108–110, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carver:1988:BRB

[Car88] D. L. Carver. Book review: Designing the user interface, strategies for effective human-computer interaction. Journal of the American Society for Information Science, 39(1):22, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carrigan:1995:BRBa

[Car95a] Dennis Carrigan. Book review: Collection management and development: Issues in an electronic era. Journal of the American Society for Information Science, 46(3):239–240, April 1995. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carrigan:1995:BRBb

[Car95b] Dennis Carrigan. Book review: Who owns information? From privacy to public access. Journal of the American Society for Information Science, 46(9):714–716, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carlyle:1996:OAW

[Car96] Allyson Carlyle. Ordering author and work records: an evaluation of collocation in online catalog displays. Journal of the American Society for Information Science, 47(7):538–554, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Casper:1978:EDL

[Cas78] Cheryl A. Casper. Estimating the demand for library service: Theory and practice. Journal of the American Society for Information Science, 29(5):232–237, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Casper:1979:PPL

[Cas79] Cheryl A. Casper. Pricing Policy for Library Services. Journal of the American Society for Information Science, 30(5):304–309, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1985:PCM

[Cas85] Donald Case. The personal computer: Missing link to the electronic journal? Journal of the American Society for Information Science, 36(5):309–313, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1991:RCO

[Cas91] Donald Owen Case. Research: Conceptual organization and retrieval of text by historians: the role of memory and metaphor. Journal of the American Society for Information Science, 42 (9):657–668, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1992:BRB

[Cas92] Donald O. Case. Book review: Dilemmas in the study of information: Exploring the boundaries of information science. Journal of the American Society for Information Science, 43(10):674–675, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1994:RSS

[Cas94] Donald O. Case. Research: The social shaping of videotex: How information services for the public have evolved. Journal of the American Society for Information Science, 45(7):483–497, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1995:EBR

[Cas95] Donald O. Case. Editorial: Book reviewing. Journal of the American Society for Information Science, 46(5):325–326, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1996:BRB

[Cas96] Donald O. Case. Book review: The trouble with computers: Usefulness, usability, and productivity. Journal of the American Society for Information Science, 47(3):255–256, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1997:BRB

[Cas97] Donald Case. Book review: Civilizing cyberspace: Policy, power, and the information superhighway. Journal of the American Society for Information Science, 48(3):278–279, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1998:BRBa

[Cas98a] Donald O. Case. Book review: Highway of dreams: a critical view along the information superhighway. Journal of the American Society for Information Science, 49(2):185–186, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:1998:BRBb

[Cas98b] Donald O. Case. Book review: The highwaymen: Warriors of the information superhighway. Journal of the American Society for Information Science, 49(4):389–390, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cawkell:1974:SSC

[Caw74] Anthony E. Cawkell. Search strategy, construction and use of citation networks, with a socio-scientific example: “Amorphous semi-conductors and S. R. Ovshinsky”. Journal of the American Society for Information Science, 25(2):123–130, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cawkell:1978:ESJ

[Caw78] A. E. Cawkell. Evaluating scientific journals with Journal Citation Reports — a case study in acoustics. Journal of the American Society for Information Science, 29(1):41–46, January 1978. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Clarke:1973:EFO

[CB73] Dan C. Clarke and John L. Bennett. An Experimental Framework For Observing the Indexing Process. Journal of the American Society for Information Science, 24(1):9–24, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chorba:1983:DAL

[CB83a] Ronald W. Chorba and Michael R. W. Bommer. Developing academic library decision support systems. Journal of the American Society for Information Science, 34(1):40–50, January 1983. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Culnan:1983:POH

[CB83b] Mary J. Culnan and James H. Bair. Perspectives On…: human communication needs and organizational productivity: the potential impact of office automation. Journal of the American Society for Information Science, 34(3):215–221, May 1983. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Conklin:1989:PLH

[CB89] Jeff Conklin and Michael L. Begeman. Perspectives on … hypertext: gIBIS: a tool for all reasons. Journal of the American Society for Information Science, 40(3):200–213, May 1989. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Collins:1994:PLE

[CB94] Mauri P. Collins and Zane L. Berge. Perspectives on … electronic publishing: IPCT Journal: a case study of an electronic journal on the Internet. Journal of the American Society for Information Science, 45(10):771–776, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Couvreur:1994:RAP

[CBM^{+}94] T. R. Couvreur, R. N. Benzel, S. F. Miller, D. N. Zeitler, D. L. Lee, M. Singhal, N. Shivaratri, and W. Y. P. Wong. Research: An analysis of performance and cost factors in searching large text databases using parallel search systems. Journal of the American Society for Information Science, 45(7):443–464, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carroll:1975:BCC

[CC75] John M. Carroll and John Cakarnis. Brief communications: Comparison of checksum procedures by simulation. Journal of the American Society for Information Science, 26(2):125–128, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cleveland:1979:TBP

[CC79] Ana Divino Cleveland and Donald B. Cleveland. A test of Buck’s prime number coding scheme. Journal of the American Society for Information Science, 30(5):290–295, September 1979. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cordes:1992:ROB

[CC92] David W. Cordes and Doris L. Carver. Research: An objectbased requirements modeling method. Journal of the American Society for Information Science, 43(1):62–71, January 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Conlon:1996:ROU

[CC96] John R. Conlon and Sumali J. Conlon. Research: Optimal use of an information retrieval system. Journal of the American Society for Information Science, 47(6):449–457, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Conlon:1998:RRW

[CC98] John R. Conlon and Sumali J. Conlon. Research: Robustness of well-designed retrieval performance measures under optimal user behavior. Journal of the American Society for Information Science, 49(4):356–363, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1999:BCD

[CC99] Blaise Cronin and Holly Crawford. Brief communication: Do Deans publish what they preach? Journal of the American Society for Information Science, 50(5):471–474, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:2000:EID

[CCM00] Chaomei Chen, Mary Czerwinski, and Robert Macredie. Editorials: Individual differences in virtual environments — Introduction and overview. Journal of the American Society for Information Science, 51(6):499–507, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Coll:1993:RAC

[CCN93] Joan H. Coll, Richard Coll, and Renuka Nandavar. Research: Attending to cognitive organization in the design of computer menus: a two-experiment study. Journal of the American Society for Information Science, 44(7):393–397, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1998:SIB

[CCRY98] Hsinchun Chen, Yi-Ming Chung, Marshall Ramsey, and Christopher C. Yang. A smart itsy bitsy spider for the Web. Journal of the American Society for Information Science, 49(7):604–618, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1995:RDB

[CCT95] Ye-Sho Chen, P. Pete Chong, and Morgan Y. Tong. Research: Dynamic behavior of Bradford’s Law. Journal of the American Society for Information Science, 46(5):370–383, June 1995. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cleveland:1984:LTF

[CCW84] Donald B. Cleveland, Ana D. Cleveland, and Olga B. Wise. Less than full-text indexing using a non-boolean searching model. Journal of the American Society for Information Science, 35(1): 19–28, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Clementini:1994:OOM

[CD94] Eliseo Clementini and Paolino Di Felice. Object-oriented modeling of geographic data. Journal of the American Society for Information Science, 45(9):694–670, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chui:1997:WZW

[CD97] Michael Chui and Andrew Dillon. Who’s zooming whom? Attunement to animation in the interface. Journal of the American Society for Information Science, 48(11):1067–1072, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Corridoni:1998:IRC

[CDV98] Jacopo M. Corridoni, Alberto Del Bimbo, and Enrico Vicario. Image retrieval by color semantics with incomplete knowledge. Journal of the American Society for Information Science, 49(3): 267–282, March 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1980:CCP

[CELY80] David Cooper, Michael A. Emly, Michael F. Lynch, and A. Robin Yeates. Compression of continuous prose texts using variety generation. Journal of the American Society for Information Science, 31(3):201–207, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cerny:1979:RRD

[Cer79] Barbara A. Cerny. A reply to Robertson’s diatribe on the nature of fuzz. Journal of the American Society for Information Science, 30(6):356–357, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Rob78, Rob79].

Cerny:1983:BRB

[Cer83] Barbara Cerny. Book review: Communications and the future: Prospecte, promises and problems. Howard F. Didsbury, Jr., Ed. Bethesda, MD: World Future Society; 1982: 357 pp. Journal of the American Society for Information Science, 34(4):301, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chambers:1973:JCM

[CH73] George R. Chambers and James S. Healey. Journal citations in Master’s Theses: One measurement of a journal collection. Journal of the American Society for Information Science, 24(5): 397–401, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Clement:1998:CCD

[CH98] Andrew Clement and Chris Halonen. Collaboration and conflict in the development of a computerized dispatch facility. Journal of the American Society for Information Science, 49(12):1090–1100, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crews:1999:PLCc

[CH99] Kenneth D. Crews and Georgia K. Harper. Perspective on … copyright and fair-use guidelines for education and libraries: The immunity dilemma: Are state colleges and universities still liable for copyright infringements? Journal of the American Society for Information Science, 50(14):1350–1352, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Case:2000:RHC

[CH00] Donald O. Case and Georgeann M. Higgins. Research: How can we investigate citation behavior? A study of reasons for citing literature in communication. Journal of the American Society for Information Science, 51(7):635–645, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chafetz:1976:TIQ

[Cha76] Morris E. Chafetz. Toward increased quality consciousness for alcoholism literature. Journal of the American Society for Information Science, 27(3):162–170, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chapin:1977:LEC

[Cha77] Ned Chapin. Letters to the Editor: Change the ANS flowchart standard? Journal of the American Society for Information Science, 28(6):377, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chapman:1981:STA

[Cha81] Janet L. Chapman. A state transition analysis of online information-seeking behavior. Journal of the American Society for Information Science, 32(5):325–333, September 1981. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chartrand:1985:PI

[Cha85] Robert Lee Chartrand. The politics of information. Journal of the American Society for Information Science, 36(6):376–382, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chatman:1986:RDT

[Cha86] Elfreda A. Chatman. Research: Diffusion theory: a review and test of a conceptual model in information diffusion. Journal of the American Society for Information Science, 37(6):377–386, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chamis:1988:BCS

[Cha88] Alice Y. Chamis. Brief communications: Selection of online databases using switching vocabularies. Journal of the American Society for Information Science, 39(3):217–218, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chatman:1991:RLS

[Cha91] Elfreda A. Chatman. Research: Life in a small world: Applicability of gratification theory to information-seeking behavior. Journal of the American Society for Information Science, 42(6): 438–449, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chatman:1996:RIL

[Cha96] Elfreda A. Chatman. Research: The impoverished life-world of outsiders. Journal of the American Society for Information Science, 47(3):193–206, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chalmers:1999:CIA

[Cha99a] Matthew Chalmers. Comparing information access approaches. Journal of the American Society for Information Science, 50(12):

1108–1118, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chatman:1999:RTL

[Cha99b] Elfreda A. Chatman. Research: A theory of life in the round. Journal of the American Society for Information Science, 50(3): 207–217, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1972:UPP

[Che72] Ching-Chih Chen. The use patterns of physics journals in a large academic research library. Journal of the American Society for Information Science, 23(4):254–270, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cheydleur:1976:BRB

[Che76] Benjamin F. Cheydleur. Book review: Computer output design. Susan Woolridge. New York, NY: Petrocelli-Charter, 468 pp. + 12 introduction. (1975). Journal of the American Society for Information Science, 27(5):352, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1986:RPP

[Che86a] Ching-Chih Chen. Research: Professor Philip Morse: In memoriam. Journal of the American Society for Information Science, 37(1):45–46, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1986:BRB

[Che86b] Tau Chen. Book review: Modeling and analysis of computer communications networks. Journal of the American Society for Information Science, 37(5):351–352, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1995:RML

[Che95a] Hsinchun Chen. Research: Machine learning for information retrieval: Neural networks, symbolic learning, and genetic algorithms. Journal of the American Society for Information Science, 46(3):194–216, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1995:LED

[Che95b] Ye-Sho Chen. Letters to the Editor: On the dynamic behavior of Bradford’s Law — response. Journal of the American Society for Information Science, 46(10):799–800, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1997:WCH

[Che97] Chaomei Chen. Writing with collaborative hypertext: Analysis and modeling. Journal of the American Society for Information Science, 48(11):1049–1066, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1998:BRB

[Che98a] Chaomei Chen. Book review: Information seeking and subject representation: an activity-theoretical approach to information science. Journal of the American Society for Information Science, 49(11):1042–1043, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1998:I

[Che98b] Hsinchun Chen. Introduction. Journal of the American Society for Information Science, 49(7):579–581, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1999:BRB

[Che99] Chaomei Chen. Book review: Localist connectionist approaches to human cognition, edited by Jonathan Grainger and Arthur M. Jacobs. Journal of the American Society for Information Science, 50(8):725–726, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:2000:RID

[Che00a] Chaomei Chen. Research: Individual differences in a spatialsemantic virtual environment. Journal of the American Society for Information Science, 51(6):529–542, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:2000:EISa

[Che00b] Hsinchun Chen. Editorial: Introduction to the special topic issue: Part 1. Journal of the American Society for Information Science, 51(3):213–215, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:2000:EISb

[Che00c] Hsinchun Chen. Editorial: Introduction to the special topic issue: part 2. Journal of the American Society for Information Science, 51(4):311–312, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chignell:1993:BRB

[Chi93] Mark Chignell. Book review: Hypertext in context. Journal of the American Society for Information Science, 44(2):115–116, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cuadra:1971:LE

[CHPC71] Carlos A. Cuadra, P. O’N. Honey, Mrs. Maria Patermann, and Elaine Caruso. Letters to the Editors. Journal of the American Society for Information Science, 22(4):296–297, July 1971. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Christ:1972:MSU

[Chr72] C. W. Christ, Jr. Microfiche: a study of user attitudes and reading habits. Journal of the American Society for Information Science, 23(1):30–35, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chrzastowski:1995:POC

[Chr95] Tina E. Chrzastowski. Perspectives on … the Chemist’s workstation: Do workstations work too well? An investigation into library workstation popularity and the “principle of least effort”. Journal of the American Society for Information Science, 46(8): 638–641, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1998:IBS

[CHSS98] Hsinchun Chen, Andrea L. Houston, Robin R. Sewell, and Bruce R. Schatz. Internet browsing and searching: User evaluations of category map and concept space techniques. Journal of the American Society for Information Science, 49(7):582–603, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cimino:1995:PLM

[Cim95] James J. Cimino. Perspectives on … medical informatics: Information technology in health care: Vocabulary and health care information technology: State of the art. Journal of the American Society for Information Science, 46(10):777–782, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Citera:1998:DTI

[Cit98] Maryalice Citera. Distributed teamwork: the impact of communication media on influence and decision quality. Journal of the American Society for Information Science, 49(9):792–800, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Covi:1996:ODE

[CK96] Lisa Covi and Rob Kling. Organizational dimensions of effective digital library use: Closed rational and open natural systems models. Journal of the American Society for Information Science, 47(9):672–689, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carlson:1999:RIL

[CK99] John R. Carlson and Charles J. Kacmar. Research: Increasing link marker effectiveness for WWW and other hypermedia interfaces: an examination of end-user preferences. Journal of the American Society for Information Science, 50(5):386–398, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1982:TCU

[CL82a] David Cooper and Michael F. Lynch. Text compression using variable-to fixed-length encodings. Journal of the American Society for Information Science, 33(1):18–31, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1982:MI

[CL82b] Marianne Cooper and Lois F. Lunin. Miscellaneous: Introduction. Journal of the American Society for Information Science, 33(3):144–145, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1986:RRB

[CL86] Ye-Sho Chen and Ferdinand F. Leimkuhler. Research: A relationship between Lotka’s Law, Bradford’s Law, and Zipf’s Law. Journal of the American Society for Information Science, 37(5): 307–314, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1989:RTT

[CL89] Ye-Sho Chen and Ferdinand F. Leimkuhler. Research: A typetoken identity in the Simon–Yule model of text. Journal of the American Society for Information Science, 40(1):45–53, January 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1990:BCB

[CL90] Ye-Sho Chen and Ferdinand F. Leimkuhler. Brief communications: Booth’s Law of word frequency. Journal of the American Society for Information Science, 41(5):387–388, July 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Clarke:1975:BRB

[Cla75] Isabel V. Clarke. Book review: Reader in documents of international organizations. Journal of the American Society for Information Science, 26(4):264, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cleveland:1976:DRM

[Cle76] Donald B. Cleveland. An n-dimensional retrieval model. Journal of the American Society for Information Science, 27(5):342–347, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Climenson:1970:ECF

[Cli70] W. Douglas Climenson. An exercise in computer file management. Journal of the American Society for Information Science, 21(3):201–203, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cercone:1984:AIU

[CM84] Nick Cercone and Gordon McCalla. Artificial intelligence: Underlying assumptions and basic objectives. Journal of the American Society for Information Science, 35(5):280–290, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:2000:UKD

[CM00] Charles Cole and Bertie Mandelblatt. Using Kintsch’s discourse comprehension theory to model the user’s coding of an informative message from an enabling information retrieval system. Journal of the American Society for Information Science, 51(11): 1033–1046, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Citera:1995:RFI

[CMB^{+}95] Maryalice Citera, Michael D. McNeese, Clifford E. Brown, Jonathan A. Selvaraj, Brian S. Zaff, and Randall D. Whitaker. Research: Fitting information systems to collaborating design teams. Journal of the American Society for Information Science, 46(7):551–559, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1998:ASU

[CMK^{+}98a] Hsinchun Chen, Joanne Martinez, Amy Kirchhoff, Tobun D. Ng, and Bruce R. Schatz. Alleviating search uncertainty through concept associations: Automatic indexing, co-occurrence analysis, and parallel computing. Journal of the American Society for Information Science, 49(3):206–216, March 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Coulter:1998:RSE

[CMK98b] Neal Coulter, Ira Monarch, and Suresh Konda. Research: Software engineering as seen through its research literature: a study in co-word analysis. Journal of the American Society for Information Science, 49(13):1206–1223, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crouch:1996:MGS

[CML^{+}96] Carolyn Crouch, Mike McGill, Michael Lesk, Karen Sparck Jones, Edward A. Fox, Donna Harman, and Donald H. Kraft. In memoriam: Gerald Salton, March 8, 1927–August 28, 1995. Journal of the American Society for Information Science, 47(2):108–115, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1993:RAI

[CMRWW93] Blaise Cronin, Gail McKenzie, Lourdes Rubio, and Sherrill Weaver-Wozniak. Research: Accounting for influence: Acknowledgments in contemporary sociology. Journal of the American Society for Information Science, 44(7):406–412, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carpenter:1973:CSJ

[CN73] Mark P. Carpenter and Francis Narin. Clustering of scientific journals. Journal of the American Society for Information Science, 24(6):425–436, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carpenter:1981:ASC

[CN81] Mark P. Carpenter and Francis Narin. The adequacy of the Science Citation Index (SCI) as an indicator of international scientific activity. Journal of the American Society for Information Science, 32(6):430–439, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1995:RAA

[CN95] H. Chen and T. Ng. Research: An algorithmic approach to concept exploration in a large knowledge network (automatic thesaurus consultation): Symbolic branch-and-bound search vs. connectionist Hopfield net activation. Journal of the American Society for Information Science, 46(5):348–369, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1997:RCS

[CNMS97] Hsinchun Chen, Tobun D. Ng, Joanne Martinez, and Bruce R. Schatz. Research: A concept space approach to addressing the vocabulary problem in scientific information retrieval: an experiment on the Worm Community System. Journal of the American Society for Information Science, 48(1):17–31, January 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crabtree:2000:REI

[CNO^{+}00] Andy Crabtree, David M. Nichols, Jon O’Brien, Mark Rouncefield, and Michael B. Twidale. Research: Ethnomethodologically informed ethnography and information system design. Journal of the American Society for Information Science, 51(7):666–682, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Can:1984:RTP

[CO84] Fazli Can and Esen A. Ozkarahan. Research: Two partitioning type clustering algorithms. Journal of the American Society for Information Science, 35(5):268–276, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Can:1985:SSA

[CO85] Fazli Can and Esen A. Ozkarahan. Similarity and stability analysis of the two partitioning type clustering algorithms. Journal of the American Society for Information Science, 36(1):3–14, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Can:1987:RCT

[CO87] Fazli Can and Esen A. Ozkarahan. Research: Computation of term/document discrimination values by use of the cover coefficient concept. Journal of the American Society for Information Science, 38(3):171–183, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1994:RCB

[CO94] Blaise Cronin and Kara Overfelt. Research: Citation-based auditing of academic performance. Journal of the American Society for Information Science, 45(2):61–72, March 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1995:BCJ

[CO95] Blaise Cronin and Kara Overfelt. Brief communications: Ejournals and tenure. Journal of the American Society for Information Science, 46(9):700–703, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1996:BCP

[CO96] Blaise Cronin and Kara Overfelt. Brief communications: Postscript on program rankings. Journal of the American Society for Information Science, 47(2):173–176, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Coady:1986:RCS

[Coa86] Reginald P. Coady. Research: A comparison of single book renewals by subject and patron status for similar rates of renewal and return. Journal of the American Society for Information Science, 37(2):78–85, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cobbs:1996:BRB

[Cob96] Pamela Cobbs. Book review: The cult of information: a neo-Luddite treatise on high-tech, artificial intelligence, and the true art of thinking, 2nd ed. Journal of the American Society for Information Science, 47(3):251–252, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Culnan:1990:POA

[COC90] Mary J. Culnan, Charles A. O’Reilly III, and Jennifer A. Chatman. Perspectives on … author cocitation analysis: Intellectual structure of research in organizational behavior, 1972–1984: a cocitation analysis. Journal of the American Society for Information Science, 41(6):453–458, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cochrane:1998:MPA

[Coc98] Pauline Atherton Cochrane. In memoriam: Phyllis Allen Richmond: Award of Merit winner dies at 76. Journal of the American Society for Information Science, 49(14):1246–1247, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cody:1973:BCE

[Cod73] Roger Cody. Brief communications: The effect of document ordering in Rocchio’s clustering algorithm. Journal of the American Society for Information Science, 24(3):232–233, May 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cohen:1995:RHL

[Coh95] Jonathan D. Cohen. Research: Highlights: Language- and domain-independent automatic indexing terms for abstracting. Journal of the American Society for Information Science, 46(3): 162–174, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cohen:1999:RMQ

[Coh99] Jonathan D. Cohen. Research: Massive query resolution for rapid selective dissemination of information. Journal of the American Society for Information Science, 50(3):195–206, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Coile:1977:LFD

[Coi77] Russell C. Coile. Lotka’s frequency distribution of scientific productivity. Journal of the American Society for Information Science, 28(6):366–370, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:1981:EDA

[Col81] Elliot Cole. Examining design assumptions for an information retrieval service: SDI use for scientific and technical databases. Journal of the American Society for Information Science, 32(6): 444–450, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:1983:POO

[Col83] Elliot Cole. Perspectives On…: old and new models for office automation. Journal of the American Society for Information Science, 34(3):234–239, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Coleman:1992:RLP

[Col92] S. R. Coleman. Research: The laboratory as a productivity and citation unit in the publications of an experimental-psychology specialty. Journal of the American Society for Information Science, 43(9):639–643, October 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:1993:RSR

[Col93] Charles Cole. Research: Shannon revisited: Information in terms of uncertainty. Journal of the American Society for Information Science, 44(4):204–211, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:1994:RON

[Col94] Charles Cole. Research: Operationalizing the notion of information as a subjective construct. Journal of the American Society for Information Science, 45(7):465–476, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Collantes:1995:RDA

[Col95] Lourdes Y. Collantes. Research: Degree of agreement in naming objects and concepts for information retrieval. Journal of the American Society for Information Science, 46(2):116–132, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:1996:BRB

[Col96] Charles Cole. Book review: Fluid concepts and creative analogies: Computer models of the fundamental mechanisms of thought. Journal of the American Society for Information Science, 47(5): 403–404A, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:1998:BRB

[Col98] Charles Cole. Book review: Information of the image (2nd ed.). Journal of the American Society for Information Science, 49(14): 1333–1334, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:1999:RAU

[Col99] Charles Cole. Research: Activity of understanding a problem during interaction with an “enabling” information retrieval system: Modeling information flow. Journal of the American Society for Information Science, 50(6):544–552, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:2000:IEI

[Col00a] Charles Cole. Interaction with an enabling information retrieval system: Modeling the user’s decoding and encoding operations. Journal of the American Society for Information Science, 51(5): 417–426, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cole:2000:NCP

[Col00b] Charles Cole. Name collection by Ph.D. history students: inducing expertise. Journal of the American Society for Information Science, 51(5):444–455, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Conger:1984:LEP

[Con84] Lucinda D. Conger. Letters to the Editors: The psychology of online searching. Journal of the American Society for Information Science, 35(1):57, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Conway:1993:BRB

[Con93] Paul Conway. Book review: Managing institutional archives: Foundational principles and practices. Journal of the American Society for Information Science, 44(7):428–429, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Connell:1995:RSS

[Con95] Tschera Harkness Connell. Research: Subject searching in online catalogs: Metaknowledge used by experienced searchers. Journal of the American Society for Information Science, 46(7):506–518, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1970:DDE

[Coo70a] William S. Cooper. On deriving design equations for information retrieval systems. Journal of the American Society for Information Science, 21(6):385–395, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1970:PUC

[Coo70b] William S. Cooper. The potential usefulness of catalog access points other than author, title, and subject. Journal of the American Society for Information Science, 21(2):112–127, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1971:BCH

[Coo71] W. S. Cooper. Brief communications: On higher level association measures. Journal of the American Society for Information Science, 22(5):354–355, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1972:CME

[Coo72] Michael D. Cooper. A cost model for evaluating information retrieval systems. Journal of the American Society for Information Science, 23(5):306–312, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1973:SMRa

[Coo73a] William S. Cooper. On selecting a measure of retrieval effectiveness. Journal of the American Society for Information Science, 24(2):87–100, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1973:SMRb

[Coo73b] William S. Cooper. On selecting a measure of retrieval effectiveness part II. Implementation of the philosophy. Journal of the American Society for Information Science, 24(6):413–424, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1976:LEC

[Coo76] William S. Cooper. Letters to the Editor: Cooper’s reply to Soergel’s article on page 256 of this issue. Journal of the American Society for Information Science, 27(4):263–264, July 1976. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Soe76].

Cooke:1977:FLN

[Coo77a] Michael Cooke. Future library network automation. Journal of the American Society for Information Science, 28(5):254–258, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1977:IOR

[Coo77b] Michael D. Cooper. Input-output relationships in on-line bibliographic searching. Journal of the American Society for Information Science, 28(3):153–160, May 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1978:IDG

[Coo78] William S. Cooper. Indexing documents by gedanken experimentation. Journal of the American Society for Information Science, 29(3):107–119, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1979:BCU

[Coo79] William S. Cooper. Brief communications: Utility-theoretic indexing: a note on Wilson’s note. Journal of the American Society for Information Science, 30(3):170–172, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1982:SIS

[Coo82] Marianne Cooper. Secondary information services in science and technology: a wide-angle view. Journal of the American Society for Information Science, 33(3):152–156, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1983:RTV

[Coo83a] Michael D. Cooper. Response time variations in an online search system. Journal of the American Society for Information Science, 34(6):374–380, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1983:UPO

[Coo83b] Michael D. Cooper. Usage patterns of an online search system. Journal of the American Society for Information Science, 34(5): 343–349, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1983:EME

[Coo83c] William S. Cooper. Exploiting the maximum entropy principle to increase retrieval effectiveness. Journal of the American Society for Information Science, 34(1):31–39, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1985:TDM

[Coo85] Michael D. Cooper. Teaching database management system use in a library school curriculum. Journal of the American Society for Information Science, 36(5):330–338, September 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cook:1989:RLS

[Coo89] Kevin L. Cook. Research: Laws of scattering applied to popular music. Journal of the American Society for Information Science, 40(4):277–283, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cook:1991:LEE

[Coo91a] Kevin L. Cook. Letters to the Editor: Errata for article. Journal of the American Society for Information Science, 42(2):113–114, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1991:RFT

[Coo91b] Michael D. Cooper. Research: Failure time analysis of office system use. Journal of the American Society for Information Science, 42(9):644–656, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1991:RUS

[Coo91c] Michael D. Cooper. Research: User skill acquisition in office information systems. Journal of the American Society for Information Science, 42(10):735–746, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cooper:1998:RDC

[Coo98] Michael D. Cooper. Research: Design considerations in instrumenting and monitoring web-based information retrieval systems. Journal of the American Society for Information Science, 49(10): 903–919, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cornell:1982:BCD

[Cor82] Laurel L. Cornell. Brief communications: Duplication of Japanese names: a problem in citations and bibliographies. Journal of the American Society for Information Science, 33(2):102– 104, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Corbin:1994:BRB

[Cor94a] John Corbin. Book review: Planning second generation automated library systems. Journal of the American Society for Information Science, 45(3):218–219, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cortez:1994:BRB

[Cor94b] Ed Cortez. Book review: Insider’s guide to library automation: Essays of practical experience. Journal of the American Society for Information Science, 45(5):315–316, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cortez:1999:RUM

[Cor99] Edwin M. Cortez. Research: Use of metadata vocabularies in data retrieval. Journal of the American Society for Information Science, 50(13):1218–1223, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Courtial:1998:LEC

[Cou98] J. P. Courtial. Letters to the Editor: Comments on Leydesdorff’s article. Journal of the American Society for Information Science, 49(1):98, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Covi:2000:DMN

[Cov00] Lisa M. Covi. Debunking the myth of the Nintendo generation: How doctoral students introduce new electronic communication practices into university research. Journal of the American Society for Information Science, 51(14):1284–1294, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cox:1983:LET

[Cox83] Bruce B. Cox. Letters to the Editor: Thematic analysis in information science. Journal of the American Society for Information Science, 34(4):299–300, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cox:1988:PLE

[Cox88] Richard J. Cox. Perspectives on … education of the information professional: New dimensions, new directions: Educating archivists: Speculations on the past, present, and future. Journal of the American Society for Information Science, 39(5):340– 343, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cox:1992:PLI

[Cox92] Richard J. Cox. Perspectives on … information technology standards: The American archival profession and information technology standards. Journal of the American Society for Information Science, 43(8):571–575, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cox:1993:BRB

[Cox93] Richard J. Cox. Book review: Without consent: the ethics of disclosing personal information in public archives; and The management of college and university archives. Journal of the American Society for Information Science, 44(3):179–180, April 1993. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cox:1994:BRB

[Cox94] Richard J. Cox. Book review: Information management and archival data. Journal of the American Society for Information Science, 45(4):284–285, May 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cox:1998:BRB

[Cox98] Richard J. Cox. Book review: Publishing books. Journal of the American Society for Information Science, 49(4):390–391, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cox:2000:BR

[Cox00] Richard J. Cox. Book review. Journal of the American Society for Information Science, 51(1):91–92, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cozzens:1982:SCI

[Coz82] Susan E. Cozzens. Split Citation Identity: A case study from economics. Journal of the American Society for Information Science, 33(4):233–236, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cobb:1998:MSR

[CP98] Maria A. Cobb and Frederick E. Petry. Modeling spatial relationships within a fuzzy framework. Journal of the American Society for Information Science, 49(3):253–266, March 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chubin:1984:RCC

[CPR84] Daryl E. Chubin, Alan L. Porter, and Frederick A. Rossini. Research: “citation classics” analysis: an approach to characterizing interdisciplinary research. Journal of the American Society for Information Science, 35(6):360–368, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carroll:1988:RBP

[CPW88] David M. Carroll, Christine A. Pogue, and Peter Willett. Research: Bibliographic pattern matching using the ICL Distributed Array Processor. Journal of the American Society for Information Science, 39(6):390–399, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1980:MLS

[CR80] Susan Crawford and Alan M. Rees. Miscellaneous: On-line systems in science and technology. Journal of the American Society for Information Science, 31(3):153, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carpineto:2000:ROT

[CR00] Claudio Carpineto and Giovanni Romano. Research: Ordertheoretical ranking. Journal of the American Society for Information Science, 51(7):587–601, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cravens:1971:PPI

[Cra71a] David W. Cravens. Predicting performance of information specialists. Journal of the American Society for Information Science, 22(1):5–11, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1971:ICA

[Cra71b] Susan Crawford. Informal communication among scientists in sleep research. Journal of the American Society for Information Science, 22(5):301–310, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1972:LEF

[Cra72] Susan Crawford. Letters to the Editor: Formal and informal communication among scientists in sleep research. Journal of the American Society for Information Science, 23(4):287–288, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Craven:1977:NNP

[Cra77] Timothy C. Craven. Nephis: a nested-phrase indexing system. Journal of the American Society for Information Science, 28(2):

107–114, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Craven:1980:BCM

[Cra80a] Timothy C. Craven. Brief communication: Microcomputergenerated graphic displays as an aid in string indexing. Journal of the American Society for Information Science, 31(2):123–124, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1980:I

[Cra80b] Susan Crawford. Introduction. Journal of the American Society for Information Science, 31(6):404, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1981:RMI

[Cra81a] Robert G. Crawford. The relational model in information retrieval. Journal of the American Society for Information Science, 32(1):51–64, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1981:II

[Cra81b] Susan Crawford. Introduction: Introduction. Journal of the American Society for Information Science, 32(2):122–123, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Craven:1982:ANC

[Cra82] Timothy C. Craven. Automatic Nephis coding of descriptive titles for permuted index generation. Journal of the American Society for Information Science, 33(2):97–101, March 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Craven:1986:BCB

[Cra86] Timothy C. Craven. Brief communications: A bracket-based coding scheme for source descriptions in string indexing. Journal of the American Society for Information Science, 37(3):166–168, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Craven:1991:RGD

[Cra91] Timothy C. Craven. Research: Graphic display of larger sentence dependency structures. Journal of the American Society for Information Science, 42(5):323–331, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1996:MIM

[Cra96] Robert G. Crawford. In memoriam: Ian A. Macleod, February 15, 1943–December 15, 1995. Journal of the American Society for Information Science, 47(7):487, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Craven:2000:APU

[Cra00] Timothy C. Craven. Abstracts produced using computer assistance. Journal of the American Society for Information Science, 51(8):745–756, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cummings:1970:LE

[CRC70] Martin M. Cummings, Frank B. Rogers, and A. E. Cawkell. Letters to the Editor. Journal of the American Society for Information Science, 21(1):98–100, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cretsos:1976:OPS

[Cre76] James M. Cretsos. Opinion paper: Scientific freedom and responsibility. Journal of the American Society for Information Science, 27(4):211–212, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crews:1995:BRBb

[Cre95a] Kenneth D. Crews. Book review: Cyberspace and the law: Your rights and duties in the online world. Journal of the American Society for Information Science, 46(10):795–796, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crews:1995:BRBa

[Cre95b] Kenneth D. Crews. Book review: Library records: a retention and confidentiality guide. Journal of the American Society for Information Science, 46(9):707, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crews:1999:PLCb

[Cre99a] Kenneth D. Crews. Perspective on … copyright and fair-use guidelines for education and libraries: Electronic reserves and fair use: the outer limits of CONFU. Journal of the American Society for Information Science, 50(14):1342–1345, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crews:1999:PLCa

[Cre99b] Kenneth D. Crews. Perspective on … copyright and fair-use guidelines for education and libraries: Introduction and overview. Journal of the American Society for Information Science, 50(14): 1304–1307, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crickman:1981:BRB

[Cri81] Robin D. Crickman. Book review: Systems analysis for information retrieval. Helen M. Townley. Boulder, Co: Westview Press; 1978: 121 pp. Journal of the American Society for Information Science, 32(5):339, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Croft:1977:CLF

[Cro77] W. Bruce Croft. Clustering large files of documents using the single-link method. Journal of the American Society for Information Science, 28(6):341–344, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Croft:1979:UBQ

[Cro79] W. Bruce Croft. Using boolean queries with a clustered file organization. Journal of the American Society for Information Science, 30(6):358–360, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Croft:1981:DRP

[Cro81] W. Bruce Croft. Document representation in probabilistic models of information retrieval. Journal of the American Society for Information Science, 32(6):451–457, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Croft:1986:RBQ

[Cro86] W. Bruce Croft. Research: Boolean queries and term dependencies in probabilistic retrieval models. Journal of the American Society for Information Science, 37(2):71–77, March 1986. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1994:BCT

[Cro94] Blaise Cronin. Brief communications: Tiered citation and measures of document similarity. Journal of the American Society for Information Science, 45(7):537–538, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1999:BCW

[Cro99a] Blaise Cronin. Brief communication: The Warholian moment and other proto-indicators of scholarly salience. Journal of the American Society for Information Science, 50(10):953–955, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crowley:1999:CDP

[Cro99b] Bill Crowley. The control and direction of professional education. Journal of the American Society for Information Science, 50(12): 1127–1135, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Crawford:1990:PLE

[CS90] Susan Crawford and Loretta Stucki. Perspectives on … evaluation of scientific information: Peer review and the impact of new information technology: Peer review and the changing research record. Journal of the American Society for Information Science, 41(3):223–228, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Collins:1991:PLI

[CS91] Rosann Collins and Detmar Straub. Perspectives on … integrated information centers within academic environments: Part I. The delivery of information services within a changing information environment. The changing information environment. Journal of the American Society for Information Science, 42(2):120– 123, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chandrasekharan:1987:RMV

[CSI87] N. Chandrasekharan, R. Sridhar, and S. S. Iyengar. Research: On the minimum vocabulary problem. Journal of the American Society for Information Science, 38(4):234–238, July 1987. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cronin:1998:IW

[CSR^{+}98] Blaise Cronin, Herbert W. Snyder, Howard Rosenbaum, Anna Martinson, and Ewa Callahan. Invoked on the Web. Journal of the American Society for Information Science, 49(14):1319–1328, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1998:RML

[CSSI98] Hsinchun Chen, Ganesan Shankaranarayanan, Linlin She, and Anand Iyer. Research: A machine learning approach to inductive query by examples: an experiment using relevance feedback, ID3, genetic algorithms, and simulated annealing. Journal of the American Society for Information Science, 49(8):693–705, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Carroll:1973:UAT

[CT73] John M. Carroll and Jean M. Tague. Use of an automatic text analyzer in preparation of SDI profiles. Journal of the American Society for Information Science, 24(4):277–281, July 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Croft:1987:RNA

[CT87] W. B. Croft and R. H. Thompson. Research: I^{3}R: a new approach to the design of document retrieval systems. Journal of the American Society for Information Science, 38(6):389–404, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chelton:1999:RIW

[CT99] Mary K. Chelton and Nancy P. Thomas. Research: Introduction: Why a special topic issue on youth issues? Journal of the American Society for Information Science, 50(1):7–9, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cuadra:1971:LSP

[Cua71] Carlos A. Cuadra. On-line systems: Promise and pitfalls. Journal of the American Society for Information Science, 22(2):107–114, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Culnan:1985:DPA

[Cul85] Mary J. Culnan. The dimensions of perceived accessibility to information: Implications for the delivery of information systems and services. Journal of the American Society for Information Science, 36(5):302–308, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Culnan:1988:BRB

[Cul88] Mary J. Culnan. Book review: INFOTRENDS: Profiting from your information resources. Journal of the American Society for Information Science, 39(6):435–436, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Curras:1993:LEN

[Cur93] Emilia Curras. Letters to the Editor: The need for theoretical studies in information science. Journal of the American Society for Information Science, 44(7):430, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Curthoys:1997:BRB

[Cur97] Patty Curthoys. Book review: Developing user interfaces: Ensuring usability through product and process. Journal of the American Society for Information Science, 48(6):575–576, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cohen:1980:AGT

[CV80] Jacob Cohen and Wim Vijverberg. Applying game theory to library networks. Journal of the American Society for Information Science, 31(5):369–374, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1992:RGT

[CVK92] Guoqing Chen, Jacques Vandenbulcke, and Etienne E. Kerre. Research: A general treatment of data redundancy in a fuzzy relational data model. Journal of the American Society for Information Science, 43(4):304–311, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chien:2000:SAA

[CWBL00] Lee-Feng Chien, Hsin-Min Wang, Bo-Ren Bai, and Sun-Chien Lin. A spoken-access approach for Chinese text and speech information retrieval. Journal of the American Society for Information Science, 51(4):313–323, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Chen:1995:RAT

[CYFS95] Hsinchun Chen, Tak Yim, David Fye, and Bruce Schatz. Research: Automatic thesaurus generation for an electronic community system. Journal of the American Society for Information Science, 46(3):175–193, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cheng:1999:RSW

[CYW99] Kwok-Shing Cheng, Gilbert H. Young, and Kam-Fai Wong. Research: A study on word-based and integral-bit Chinese text compression algorithms. Journal of the American Society for Information Science, 50(3):218–228, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Cichocki:1988:PLC

[CZ88] Edward M. Cichocki and Susan M. Ziemer. Perspectives on … CD-ROM for information storage and retrieval: Design considerations for CD-ROM retrieval software. Journal of the American Society for Information Science, 39(1):43–46, January 1988. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dalrymple:1990:RRR

[Dal90] Prudence W. Dalrymple. Research: Retrieval by reformulation in two library catalogs: Toward a cognitive model of searching behavior. Journal of the American Society for Information Science, 41(4):272–281, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Daniel:1983:POI

[Dan83] Evelyn Daniel. Perspectives On…: information resources and organizational structure. Journal of the American Society for Information Science, 34(3):222–228, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Daniel:1984:BRB

[Dan84] Evelyn H. Daniel. Book review: Encyclopedia of library and information science, Vol. 36. Alan Kent (Ed.). New York: Marcel Dekker, 1983. Journal of the American Society for Information Science, 35(6):373, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Danton:2000:LEA

[Dan00] J. Periam Danton. Letters to the Editor: Authors of information science. Journal of the American Society for Information Science, 51(9):882, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dattola:1979:FFI

[Dat79] Robert T. Dattola. FIRST: Flexible Information Retrieval System for Text. Journal of the American Society for Information Science, 30(1):9–14, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1974:LEP

[Dav74] Charles H. Davis. Letters to the Editors: PL/ i Programming Text. Journal of the American Society for Information Science, 25(6):383, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1976:BRB

[Dav76a] Charles H. Davis. Book review: An introduction to programming: A Structured Approach Using PL/1 and PL/C-7. Richard Conway and David Gries. Cambridge, Mass.: Winthrop Publishers, Inc., 509 pp. (1975). Journal of the American Society for Information Science, 27(2):126, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1976:BCP

[Dav76b] Charles H. Davis. Brief communications: Pragmatic expansion of an enumerative classification scheme. Journal of the American Society for Information Science, 27(3):174–175, May 1976. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davidson:1977:EID

[Dav77a] David Davidson. The effect of individual differences of cognitive style on judgments of document relevance. Journal of the American Society for Information Science, 28(5):273–284, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1977:BRB

[Dav77b] Charles H. Davis. Book review: Reliable Software through Composite Design. Glenford J. Myers. New York: Petrocelli-Charter, 159 pp. (1975) (ISBN: 0-88405-284-2). Journal of the American Society for Information Science, 28(5):303, September 1977. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davidson:1979:LEM

[Dav79a] Colin H. Davidson. Letter to the Editors: Microfiches. Journal of the American Society for Information Science, 30(6):372, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1979:BRB

[Dav79b] Charles H. Davis. Book review: Current Trends in Programming Methodology, Volume I: Software Specification and Design. Raymond T. Yeh, ed. Englewood Cliffs, NJ: Prentice-Hall; 1977: 275 pp. ISBN: 0–13–195701–5 (v. 1). Journal of the American Society for Information Science, 30(1):61–62, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davison:1979:LE

[Dav79c] David Davison. Letters to the Editor. Journal of the American Society for Information Science, 30(1):60, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1984:LEA

[Dav84] Charles H. Davis. Letters to the Editors: Admissibility of delegates to international meetings. Journal of the American Society for Information Science, 35(1):56, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Mea83d] and reply [Mea84c].

Davison:1986:PLT

[Dav86] Wayne E. Davison. Perspectives on … telecommunications: Principles, developments, prospects: The international telecommunications environment. Journal of the American Society for Information Science, 37(6):418–422, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davies:1988:PLC

[Dav88] David H. Davies. Perspectives on … CD-ROM for information storage and retrieval: The CD-ROM medium. Journal of the American Society for Information Science, 39(1):34–42, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1992:BRB

[Dav92a] Charles H. Davis. Book review: COBOL programs for libraries. Journal of the American Society for Information Science, 43(6): 456, July 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1992:LER

[Dav92b] Charles H. Davis. Letters to the Editor: Regarding “A Probabilistic Model for the Distribution of Authorships”. Journal of the American Society for Information Science, 43(1):96, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Aji91].

Davis:1994:MBM

[Dav94] Charles H. Davis. In memoriam: Bernard M. Fry, 1915–1994. Journal of the American Society for Information Science, 45(8): 547, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1996:IDL

[Dav96a] Ben H. Davis. Instances of distance learning: Teacher of the future. Journal of the American Society for Information Science, 47(11):849–853, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1996:BRB

[Dav96b] Charles H. Davis. Book review: At the crossroads: Librarians on the information superhighway. Journal of the American Society for Information Science, 47(10):789–790A, October 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1998:BRB

[Dav98] Charles H. Davis. Book review: Information science: Still an emerging discipline. Journal of the American Society for Information Science, 49(6):569, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Day:1997:POH

[Day97] Ron Day. Paul otlet and his successors: Paul Otlet’s book and the writing of social space. Journal of the American Society for Information Science, 48(4):310–317, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Day:1999:PLC

[Day99a] Colin Day. Perspective on … copyright and fair-use guidelines for education and libraries: The economics of publishing: the consequences of library and research copying. Journal of the American Society for Information Science, 50(14):1346–1349, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Day:1999:BRB

[Day99b] Ronald Day. Book review: Remediation: Understanding new media. Journal of the American Society for Information Science, 50(8):730–732, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Day:2000:RCM

[Day00a] Ronald E. Day. Research: The “conduit metaphor” and the nature and politics of information studies. Journal of the American Society for Information Science, 51(9):805–811, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Day:2000:THE

[Day00b] Ronald E. Day. Tropes, history, and ethics in professional discourse and information science. Journal of the American Society for Information Science, 51(5):469–475, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Doty:1994:PLE

[DB94] Philip Doty and Ann P. Bishop. Perspectives on … electronic publishing: The National Information Infrastructure and electronic publishing: a reflective essay. Journal of the American Society for Information Science, 45(10):785–799, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Deregt:1976:BRB

[DC76] Maurits P. Deregt and Joseph Caponio. Book review: Computer techniques in environmental science. R. P. Ovellete, R. S. Greeley and J. W. Overbey. New York: Petrocelli Charter. 248 pp. (1975). Journal of the American Society for Information Science, 27(3): 188, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1980:TEA

[DC80] Martin Dillon and P. Caplan. A technique for evaluating automatic term clustering. Journal of the American Society for Information Science, 31(2):89–96, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1993:BCA

[DC93] Charles H. Davis and Blaise Cronin. Brief communications: Acknowledgments and intellectual indebtedness: a bibliometric conjecture. Journal of the American Society for Information Science, 44(10):590–592, December 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Debackere:1998:ABM

[DC98] Koenraad Debackere and Bart Clarysse. Advanced bibliometric methods to model the relationship between entry behavior and networking in emerging technological communities. Journal of the American Society for Information Science, 49(1):49–58, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dingle-Cliff:1981:CRA

[DCD81] Susan Dingle-Cliff and Charles H. Davis. Comparison of recent acquisitions and OCLC find rates for three Canadian special libraries. Journal of the American Society for Information Science, 32(1):65–69, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ding:2000:BIR

[DCFQ00] Ying Ding, Gobinda G. Chowdhury, Schubert Foo, and Weizhong Qian. Bibliometric information retrieval system (BIRS): a web search interface utilizing bibliometric research results. Journal of the American Society for Information Science, 51(13):1190–1204, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Deogun:1998:FSE

[DCRS98] Jitender S. Deogun, Suresh K. Choubey, Vijay V. Raghavan, and Hayri Sever. Feature selection and effective classifiers. Journal of the American Society for Information Science, 49(5):423–434, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1978:BCE

[DDC78] Charles H. Davis and Susan Dingle-Cliff. Brief communication: Evidence of OCLC’s potential for special libraries and technical information centers. Journal of the American Society for Information Science, 29(5):255–256, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Deerwester:1990:RIL

[DDF^{+}90] Scott Deerwester, Susan T. Dumais, George W. Furnas, Thomas K. Landauer, and Richard Harshman. Research: Indexing by latent semantic analysis. Journal of the American Society for Information Science, 41(6):391–407, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

deSollaPrice:1971:SRE

[de 71] Derek J. de Solla Price. Some remarks on elitism in information and the invisible college phenomenon in science. Journal of the American Society for Information Science, 22(2):74–75, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

deSollaPrice:1976:GTB

[de 76] Derek de Solla Price. A general theory of bibliometric and other cumulative advantage processes. Journal of the American Society for Information Science, 27(5):292–306, September 1976. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

deSollaPrice:1978:BCC

[de 78] Derek de Solla Price. Brief communications: Cumulative advantage urn games explained: a reply to Kantor. Journal of the American Society for Information Science, 29(4):204–206, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Kan78].

DeHaen:1980:LEC

[De 80] Paul De Haen. Letter to the Editor: Coal gasification. Journal of the American Society for Information Science, 31(2):125, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

DElia:1997:PLI

[D’E97] George D’Elia. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: Introduction and overview. Journal of the American Society for Information Science, 48(5):430–433, May 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Debons:1972:LEI

[Deb72] Anthony Debons. Letters to the editor: Information science, content & structure. Journal of the American Society for Information Science, 23(2):139, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Debons:1986:BRB

[Deb86] Anthony Debons. Book review: Information seeking: Assessing and anticipating user needs. Journal of the American Society for Information Science, 37(3):173, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Debons:1992:LER

[Deb92] Anthony Debons. Letters to the Editor: Rejoinder to A. A. Shpackov’s letter. Journal of the American Society for Information Science, 43(10):681–682, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Shp92].

Debons:1996:MJB

[Deb96] Anthony Debons. In memoriam: Jack Belzer, January 16, 1910– June 22, 1995. Journal of the American Society for Information Science, 47(8):586–587, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Delendick:1990:RCA

[Del90] Thomas J. Delendick. Research: Citation analysis of the literature of Systemic Botany: a preliminary survey. Journal of the American Society for Information Science, 41(7):535–543, October 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Depew:1975:ADM

[Dep75] John N. Depew. An acquisitions decision model for academic libraries. Journal of the American Society for Information Science, 26(4):237–246, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Derr:1984:ISE

[Der84] Richard L. Derr. Information seeking expressions of users. Journal of the American Society for Information Science, 35(2):124– 128, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dervin:1994:IDE

[Der94] Brenda Dervin. Information democracy: an examination of underlying assumptions. Journal of the American Society for Information Science, 45(6):369–385, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Desai:1997:RSD

[Des97] Bipin C. Desai. Research: Supporting discovery in virtual libraries. Journal of the American Society for Information Science, 48(3):190–204, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1982:UDA

[DF82] Martin Dillon and Peggy Federhart. The use of discriminant analysis to select content-bearing words. Journal of the American Society for Information Science, 33(4):245–253, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1984:SRC

[DF84] Martin Dillon and Peggy Federhart. Statistical recognition of content terms in general text. Journal of the American Society for Information Science, 35(1):1–10, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Doreian:1985:SEJ

[DF85] Patrick Doreian and Thomas J. Fararo. Structural equivalence in a journal network. Journal of the American Society for Information Science, 36(1):28–37, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drott:1978:EEB

[DG78a] M. Carl Drott and Belver C. Griffith. An empirical examination of Bradford’s Law and the scattering of scientific literature.

Journal of the American Society for Information Science, 29(5): 238–246, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drott:1978:LEI

[DG78b] M. Carl Drott and Belver C. Griffith. Letters to the Editor: Interlibrary loans: Impact of the new copyright law. Journal of the American Society for Information Science, 29(5):259–260, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1983:FFA

[DG83] Martin Dillon and Ann S. Gray. FASIT: a fully automatic syntactically based indexing system. Journal of the American Society for Information Science, 34(2):99–108, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Das-Gupta:1987:RBI

[DG87] Padmini Das-Gupta. Research: Boolean interpretation of conjunctions for document retrieval. Journal of the American Society for Information Science, 38(4):245–254, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diodato:1991:RBB

[DG91] Virgil Diodato and Gretchen Gandt. Research: Back of book indexes and the characteristics of author and nonauthor indexing: Report of an exploratory study. Journal of the American Society for Information Science, 42(5):341–350, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Deibel:1995:PLM

[DG95] Stephan R. A. Deibel and Robert A. Greenes. Perspectives on … medical informatics: Information technology in health care: An infrastructure for the development of health care information systems from distributed components. Journal of the American Society for Information Science, 46(10):765–771, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:2000:BCG

[DG00] Andrew Dillon and Barbara A. Gushrowski. Brief communication: Genres and the WEB: Is the personal home page the first uniquely digital genre? Journal of the American Society for Information Science, 51(2):202–205, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Desai:1986:RDM

[DGS86] B. C. Desai, P. Goyal, and F. Sadri. Research: A data model for use with formatted and textual data. Journal of the American Society for Information Science, 37(3):158–165, May 1986. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1970:ACA

[DH70] Charles H. Davis and Peter Hiatt. An automated currentawareness service for public libraries. Journal of the American Society for Information Science, 21(1):29–33, January 1970. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dimsdale:1978:LEU

[DH78] Christine Dimsdale and Donald T. Hawkins. Letters to the Editor: Unconventional uses of on-line retrieval systems. Journal of the American Society for Information Science, 29(4):209, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Debons:1997:NAS

[DH97a] Anthony Debons and Esther E. Horne. NATO advanced study institutes of Information Science and foundations of Information Science. Journal of the American Society for Information Science, 48(9):794–803, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dilevko:1997:RIT

[DH97b] Juris Dilevko and Roma M. Harris. Research: Information technology and social relations: Portrayals of gender roles in high tech product advertisements. Journal of the American Society for Information Science, 48(8):718–727, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davenport:2000:RNN

[DHS00] Elisabeth Davenport, Martin Higgins, and Ian Somerville. Research: Narratives of new media in Scottish households: the evolution of a framework of inquiry. Journal of the American Society for Information Science, 51(10):900–912, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dick:1986:PLT

[Dic86] Richard Dick. Perspectives on … telecommunications: Perspectives, developments, prospects: Views on telecommunications: Implications for the future. Journal of the American Society for Information Science, 37(6):423–428, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dickson:1996:LEC

[Dic96] Jean Dickson. Letters to the Editor: Cataloging article concerns. Journal of the American Society for Information Science, 47(2): 182, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diener:1979:LET

[Die79] Richard A. V. Diener. Letter to the Editors: Transnational information flow as assessed by citation analysis. Journal of the American Society for Information Science, 30(2):114–115, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diener:1981:BRB

[Die81] Richard A. V. Diener. Book review: Geopolitics of information: How Western culture dominates the world. Anthony Smith. New York: Oxford University Press; 1980; 192 pp. Price: $13.95. Journal of the American Society for Information Science, 32(5):340– 341, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diener:1982:BRBa

[Die82a] Richard A. V. Diener. Book review: Opening and Closing: Strategies of Information Adaptation in Society. Orrin E. Klapp. New York: Cambridge University Press; 1978. Journal of the American Society for Information Science, 33(2):109, March 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diener:1982:BRBb

[Die82b] Richard A. V. Diener. Book review: The Social and Psychological Distortion of Information. Charles K. West. Chicago: Nelson-Hall; 1981: 147 pp. Price: $17.95 (8.95 paper). Journal of the American Society for Information Science, 33(6):371, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diener:1984:BRB

[Die84a] Richard A. V. Diener. Book review: Better translation for better communication. G. Van Slype, J. F. Guinet, F. Seitz, and E. Benejam. Oxford: Pergamon Press; 1983: 194 pp. Journal of the American Society for Information Science, 35(6):373, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diener:1984:IDJ

[Die84b] Richard A. V. Diener. Informational dynamics of journal article titles. Journal of the American Society for Information Science, 35(4):222–227, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diener:1987:BRB

[Die87] Richard A. V. Diener. Book review: Technologies of freedom. Journal of the American Society for Information Science, 38(4): 313, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1982:ACH

[Dil82] Martin Dillon. Automatic classification of Harris survey questions: An experiment in the organization of information. Journal of the American Society for Information Science, 33(5):294–301, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1997:BRB

[Dil97a] Andrew Dillon. Book review: Design rationale: Concepts, techniques, and use. Journal of the American Society for Information Science, 48(8):762–763, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1997:I

[Dil97b] Andrew Dillon. Introduction. Journal of the American Society for Information Science, 48(11):965–969, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1998:BRB

[Dil98] Andrew Dillon. Book review: Computer and information ethics. Journal of the American Society for Information Science, 49(9):

861, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:2000:RSS

[Dil00] Andrew Dillon. Research: Spatial-semantics: How users derive shape from information space. Journal of the American Society for Information Science, 51(6):521–528, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dingle:1987:LEA

[Din87] Susan Dingle. Letters to the Editor: Author–date style of referencing. Journal of the American Society for Information Science, 38(4):308–309, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diodato:1983:MAI

[Dio83] Virgil Diodato. Modern algebra and information science. Journal of the American Society for Information Science, 34(4):257–261, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diodato:1994:RUP

[Dio94] Virgil Diodato. Research: User preferences for features in back of book indexes. Journal of the American Society for Information Science, 45(7):529–536, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Divilbiss:1987:BRB

[Div87] James L. Divilbiss. Book review: Advances in human-computer interaction. Journal of the American Society for Information Science, 38(4):314, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dronberger:1975:ARF

[DK75] Gladys B. Dronberger and Gerald T. Kowitz. Abstract readability as a factor in information systems. Journal of the American Society for Information Science, 26(2):108–111, March 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Debons:1990:BRB

[DK90] Anthony Debons and Fred Kohun. Book review: The control revolution: Technological and economic origins of the information society. Journal of the American Society for Information Science, 41(5):389, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1971:BCC

[DKD71] Charles H. Davis, W. Robert Kearney, and Bonnie M. Davis. Brief communications: A computer-based procedure for keyword indexing of newspapers. Journal of the American Society for Information Science, 22(5):348–351, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dworman:2000:PDS

[DKP00] Garett O. Dworman, Steven O. Kimbrough, and Chuck Patch. On pattern-directed search of archives and collections. Journal of the American Society for Information Science, 51(1):14–23, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:2000:RND

[DL00] James R. Davis and Carl Lagoze. Research: NCSTRL: Design and deployment of a globally distributed digital library. Journal of the American Society for Information Science, 51(3):273–280, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1999:BCS

[DM99] Charles H. Davis and Geoffrey W. McKim. Brief communication: Systematic weighting and ranking: Cutting the Gordian knot. Journal of the American Society for Information Science, 50(7): 626–628, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

DeGeus:1982:RNM

[DMZL82] Jan De Geus, Frans Mulder, Bert Zuurke, and Marilyn M. Levine. A replication of the Nelson and Mitroff experiment in teaching “bothsides” thinking. Journal of the American Society for Information Science, 33(2):76–81, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Depew:1981:LEA

[DO81] C. Henry Depew and John O’Connor. Letters to the Editor: Answer-passage retrieval. Journal of the American Society for Information Science, 32(3):235, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Doctor:1991:RIT

[Doc91] Ronald D. Doctor. Research: Information technologies and social equity: Confronting the revolution. Journal of the American Society for Information Science, 42(3):216–228, April 1991. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dodd:1979:BRN

[Dod79] Sue A. Dodd. Bibliographic references for numeric social science data files: Suggested guidelines. Journal of the American Society for Information Science, 30(2):77–82, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dogru:1998:BRB

[Dog98] Sait Dogru. Book review: Expert systems: Introduction to first and second generation and hybrid knowledge based systems. Journal of the American Society for Information Science, 49(10):952– 953, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Doll:1998:BRB

[Dol98] Carol A. Doll. Book review: Information services for secondary schools. Journal of the American Society for Information Science, 49(6):575–576, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dominick:1978:BRB

[Dom78] Wayne D. Dominick. Book review: Data structures and management. 2nd ed. Ivan Flores. Englewood Cliffs, NJ: Prentice-Hall; 1977. 390 p. Journal of the American Society for Information Science, 29(6):317–318, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dominich:2000:RUM

[Dom00] Sándor Dominich. Research: A unified mathematical definition of classical information retrieval. Journal of the American Society for Information Science, 51(7):614–624, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Donley:1970:BRB

[Don70] Mary Donley. Book review: 4/70-5R Library Automation — Computer Produced Book Catalog. 1969. International Business Machines, White Plains, N.Y. 41 pp. Available on request at local IBM offices. Journal of the American Society for Information Science, 21(4):307, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Donohue:1972:BAC

[Don72] Joseph C. Donohue. A bibliometric analysis of certain information science literature. Journal of the American Society for Information Science, 23(5):313–317, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Donohue:1974:BRB

[Don74] Joseph C. Donohue. Book review: Reader in Science Information, edited by John Sherrod and Alfred Hodina, Washington, D.C.: Microcard Editions, 403 pages (1973). Journal of the American Society for Information Science, 25(5):337–339, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Donohue:1976:BRB

[Don76] Joseph C. Donohue. Book review: Book Availability and the Library User. Michael K. Buckland. New York: Pergamon Press, 196 pp. (1975). Journal of the American Society for Information Science, 27(1):75–76, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Donohue:1977:BRB

[Don77] Joseph C. Donohue. Book review: The Oral Antecedents of Greek Librarianship. H. Curtis Wright. Foreword by Jesse H. Shera, afterword by H. J. de Vleeschauer. Provo, UT: Brigham Young University Press (1977). 237 pp. Journal of the American Society for Information Science, 28(6):371–372, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Donohue:1979:LCP

[Don79] Joseph C. Donohue. The library of congress: a proposed role in a national information and referral network. Journal of the American Society for Information Science, 30(4):202–204, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Donahue:1996:PRP

[Don96] Stacey Donahue. Programs and resources: Programs and resources in distance education. Journal of the American Society for Information Science, 47(11):870–874, November 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dore:1996:RCF

[DOO^{+}96] Jean-Christophe Doré, Tiiu Ojasoo, Yoshiko Okubo, Thomas Durand, Gérard Dudognon, and Jean-François Miquel. Research: Correspondence factor analysis of the publication patterns of 48 countries over the period 1981–1992. Journal of the American Society for Information Science, 47(8):588–602, August 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Doreian:1985:SEP

[Dor85] Patrick Doreian. Structural equivalence in a psychology journal network. Journal of the American Society for Information Science, 36(6):411–417, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Doreian:1988:RTS

[Dor88] Patrick Doreian. Research: Testing structural-equivalence hypotheses in a network of geographical journals. Journal of the American Society for Information Science, 39(2):79–85, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dosa:1985:ITT

[Dos85] Marta L. Dosa. Information transfer as technical assistance for development. Journal of the American Society for Information Science, 36(3):146–152, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Doszkocs:1986:PLO

[Dos86] Tamas E. Doszkocs. Perspectives on … online searching: Natural language processing in information retrieval. Journal of the American Society for Information Science, 37(4):191–196, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dosa:1988:PLE

[Dos88] Marta Dosa. Perspectives on … education of the information professional: New dimensions, new directions: Has international education been politicized? Journal of the American Society for Information Science, 39(5):351–354, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Douville:1972:TIC

[Dou72] Judith A. Douville. Technical information centers: Specialized services to science and technology. An overview. Journal of the American Society for Information Science, 23(3):176–184, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Douglas:1996:EDE

[Dou96] Gayle Douglas. Effectiveness of distance education: MLIS distance education at the University of South Carolina: Report of a case study. Journal of the American Society for Information Science, 47(11):875–879, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dow:1977:MDS

[Dow77] John T. Dow. A metatheory for the development of a science of information. Journal of the American Society for Information Science, 28(6):323–332, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dawes:1999:RCT

[DPD99] Sharon S. Dawes, Theresa A. Pardo, and Ann DiCaterino. Research: Crossing the threshold: Practical foundations for government services on the World Wide Web. Journal of the American Society for Information Science, 50(4):346–353, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dunham:1978:AIP

[DPP78] George S. Dunham, Milos G. Pacak, and Arnold W. Pratt. Automatic indexing of pathology data. Journal of the American Society for Information Science, 29(2):81–90, March 1978. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

deRaadt:1990:RIT

[dR90] J. D. R. de Raadt. Research: Information transmission in viable systems. Journal of the American Society for Information Science, 41(2):111–120, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drabenstott:1989:BRB

[Dra89] Karen Markey Drabenstott. Book review: Patron access: Issues for online catalogs. Journal of the American Society for Information Science, 40(6):441–442, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

DElia:1991:PLI

[DRCG91] George D’Elia, Nancy Freeman Rohde, James Cogswell, and Kathy Gorman. Perspectives on … integrated information centers within academic environments: Information requirements assessments of the Humphrey Institute. Journal of the American Society for Information Science, 42(2):139–142, March 1991. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dresang:1999:MRN

[Dre99] Eliza T. Dresang. More research needed: Informal informationseeking behavior of youth on the Internet. Journal of the American Society for Information Science, 50(12):1123–1124, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dempsey:1998:ERL

[DRM98] Lorcan Dempsey, Rosemary Russell, and Robin Murray. European research letters: The emergence of distributed library services: A European perspective. Journal of the American Society for Information Science, 49(10):942–951, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drott:1977:BRB

[Dro77] M. Carl Drott. Book review: Essays of an Information Scientist, Vol. I and 11. Eugene Carfield. Philadelphia, PA: ISI Press (1977). Journal of the American Society for Information Science, 28(6): 374–375, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drott:1978:BRB

[Dro78] M. Carl Drott. Book review: The Random House Encyclopedia. Mitchell, James. Editor-in-Chief. New York: Random House; 1977. Price: $70. [Published in the United Kingdom under the title: The Joy of Knowledge Encyclopedia]. Journal of the American Society for Information Science, 29(3):167, May 1978. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drott:1995:RRR

[Dro95] M. Carl Drott. Research: Reexamining the role of conference papers in scholarly communication. Journal of the American Society for Information Science, 46(4):299–305, May 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drott:1998:BRB

[Dro98] M. Carl Drott. Book review: Information retrieval systems: Theory and implementation. Journal of the American Society for Information Science, 49(7):668–670, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dym:1973:SDM

[DS73] Eleanor D. Dym and Donald L. Shirey. A statistical decision model for periodical selection for a specialized information center. Journal of the American Society for Information Science, 24(2): 110–119, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1979:COF

[DS79] Charles H. Davis and Debora Shaw. Collection overlap as a function of library size: A comparison of American and Canadian public libraries. Journal of the American Society for Information Science, 30(1):19–24, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Davis:1980:IS

[DS80] Peter Davis and Wladimir Sachs. The implementation of SCATT. Journal of the American Society for Information Science, 31(2): 67–74, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dutta:1984:APA

[DS84] S. Dutta and P. K. Sinha. Application: Pragmatic approach to subject indexing: a new concept. Journal of the American Society for Information Science, 35(6):325–331, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

DeHart:1991:BCI

[DS91] Florence E. DeHart and Larry Scott. Brief communications: ISI research fronts and online subject access. Journal of the American Society for Information Science, 42(5):386–388, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Diodato:1993:ROM

[DS93] Virgil Diodato and Fran Smith. Research: Obsolescence of music literature. Journal of the American Society for Information Science, 44(2):101–112, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dillon:1996:BCE

[DS96] Andrew Dillon and Dille Schaap. Brief communications: Expertise and the perception of shape in information. Journal of the American Society for Information Science, 47(10):786–788, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

daSilva:1993:RBF

[dSM93] Wagner Teixeira da Silva and Ruy Luiz Milidiú. Research: Belief Function Model for information retrieval. Journal of the American Society for Information Science, 44(1):10–18, January 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dumain:1987:LEI

[Dum87] Ralph Dumain. Letters to the Editor: Information processes. I. Hierarchical information systems. Journal of the American Society for Information Science, 38(4):310, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Duncan:1980:LEE

[Dun80] E. E. Duncan. Letters to the Editor: Education and training. Journal of the American Society for Information Science, 31(5): 379, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dunlop:1996:LEC

[Dun96a] Mark D. Dunlop. Letters to the Editor: Comments on a book review: Information retrieval: New systems and current research. Journal of the American Society for Information Science, 47(6): 484–485, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dunman:1996:BRB

[Dun96b] Susan Dunman. Book review: Seeking meaning: a process approach to library and information services. Journal of the American Society for Information Science, 47(3):249–250, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dunlop:2000:RMI

[Dun00] Mark Dunlop. Reflections on Mira: Interactive evaluation in information retrieval. Journal of the American Society for Information Science, 51(14):1269–1274, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Durkin:1976:BRB

[Dur76] Kay Durkin. Book review: Syndicated columnists. Richard Weiner, New York: Richard Weiner, 148 pp. (1975). Journal of the American Society for Information Science, 27(3):191, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Dimitroff:1995:RSR

[DW95] Alexandra Dimitroff and Dietmar Wolfram. Research: Searcher response in a hypertext-based bibliographic information retrieval system. Journal of the American Society for Information Science, 46(1):22–29, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Drabenstott:1996:FAS

[DW96] Karen M. Drabenstott and Marjorie S. Weller. Failure analysis of subject searches in a test of a new design for subject access to online catalogs. Journal of the American Society for Information Science, 47(7):519–537, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Denley:1998:CHA

[DW98] Ian Denley and Andy Whitefield. A case history in applying task analysis in the design of a multimedia cooperative document production system. Journal of the American Society for Information Science, 49(9):817–831, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eastman:1989:BCR

[Eas89] Caroline M. Eastman. Brief communications: Research productivity and breadth of interest revisited. Journal of the American Society for Information Science, 40(5):359–360, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eastman:1991:BCF

[Eas91] Caroline M. Eastman. Brief communications: Formatted and unformatted character data types. Journal of the American Society for Information Science, 42(2):110–111, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ellis:1999:ISI

[EAW99] David Ellis, David Allen, and Tom Wilson. Information science and information systems: Conjunct subjects disjunct disciplines. Journal of the American Society for Information Science, 50(12): 1095–1107, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eisenberg:1988:ROE

[EB88] Michael Eisenberg and Carol Barry. Research: Order effects: a study of the possible influence of presentation order on user judgments of document relevance. Journal of the American Society for Information Science, 39(5):293–300, September 1988. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Exner:1998:BCD

[EB98] Frank Exner and Little Bear. Brief communications: From drowning to surfing: a slogan’s significance. Journal of the American Society for Information Science, 49(1):92–93, ???? 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Everett:1992:RTD

[EC92] Daniel M. Everett and Steven C. Cater. Research: Topology of document retrieval systems. Journal of the American Society for Information Science, 43(10):658–673, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eom:1996:RCO

[EF96] Sean B. Eom and Roy S. Farris. Research: The contributions of organizational science to the development of decision support systems research subspecialties. Journal of the American Society for Information Science, 47(12):941–952, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Efthimiadis:1990:RGO

[Eft90] Efthimios N. Efthimiadis. Research: The growth of the OPAC literature. Journal of the American Society for Information Science, 41(5):342–347, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Efthimiadis:2000:IQE

[Eft00] Efthimios N. Efthimiadis. Interactive query expansion: a userbased evaluation in a relevance feedback environment. Journal of the American Society for Information Science, 51(11):989–1003, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ellis:1996:CHL

[EFW96] David Ellis, Jonathan Furner, and Peter Willett. On the creation of hypertext links in full-text documents: Measurement of retrieval effectiveness. Journal of the American Society for Information Science, 47(4):287–300, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1986:RDB

[Egg86] L. Egghe. Research: The dual of Bradford’s Law. Journal of the American Society for Information Science, 37(4):246–255, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1987:RPM

[Egg87] L. Egghe. Research: Pratt’s measure for some bibliometric distributions and its relation with the 80/20 rule. Journal of the American Society for Information Science, 38(4):288–297, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1988:BCR

[Egg88] L. Egghe. Brief communications: The relative concentration of a journal with respect to a subject and the use of online services in calculating it. Journal of the American Society for Information Science, 39(4):281–284, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1990:RND

[Egg90a] L. Egghe. Research: A note on different Bradford multipliers. Journal of the American Society for Information Science, 41(3): 204–209, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1990:RAT

[Egg90b] L. Egghe. Research: Applications of the theory of Bradford’s Law to the calculation of Leimkuhler’s Law and to the completion of bibliographies. Journal of the American Society for Information Science, 41(7):469–492, October 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1993:REP

[Egg93] L. Egghe. Research: Exact probabilistic and mathematical proofs of the relation between the mean μ and the generalized 80/20rule. Journal of the American Society for Information Science, 44(7):369–375, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1994:RSF

[Egg94] L. Egghe. Research: Special features of the author–publication relationship and a new explanation of Lotka’s Law based on convolution theory. Journal of the American Society for Information Science, 45(6):422–427, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1996:RSI

[Egg96] L. Egghe. Research: Source-item production laws for the case that items have multiple sources with fractional counting of credits. Journal of the American Society for Information Science, 47 (10):730–748, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1997:RPI

[Egg97] L. Egghe. Research: Price index and its relation to the mean and median reference age. Journal of the American Society for Information Science, 48(6):564–573, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1999:RLZ

[Egg99] L. Egghe. Research: On the law of Zipf–Mandelbrot for multiword phrases. Journal of the American Society for Information Science, 50(3):233–241, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

El-Hadidy:1994:RBP

[EH94] Bahaa El-Hadidy. Research: The breakeven point for using CD-ROM versus online: a case study for database access in a developing country. Journal of the American Society for Information Science, 45(4):273–283, May 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ehikhamenor:1990:RIS

[Ehi90] Fabian A. Ehikhamenor. Research: Informal scientific communication in Nigerian universities. Journal of the American Society for Information Science, 41(6):419–426, September 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Etnyre:1975:AEF

[EK75] Vance Etnyre and William J. Kretlow. An analysis of effectiveness factors in published articles on management information systems. Journal of the American Society for Information Science, 26(4): 247–252, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elliott:1997:OUD

[EK97] Margaret Elliott and Rob Kling. Organizational usability of digital libraries: Case study of legal research in civil and criminal courts. Journal of the American Society for Information Science, 48(11):1023–1035, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eres:1985:IO

[EL85] Beth Krevitt Eres and Lois F. Lunin. Introduction and overview. Journal of the American Society for Information Science, 36(3):

144–145, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elchesen:1973:CBD

[Elc73] Dennis R. Elchesen. A correlation of bibliographic data elements for use in a generalized file management system. Journal of the American Society for Information Science, 24(1):45–53, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elchesen:1978:CEC

[Elc78] Dennis R. Elchesen. Cost-effectiveness comparison of manual and on-line retrospective bibliographic searching. Journal of the American Society for Information Science, 29(2):56–66, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elias:1970:Ea

[Eli70a] A. W. Elias. Editorial. Journal of the American Society for Information Science, 21(1):1–3, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elias:1970:Eb

[Eli70b] A. W. Elias. Editorial. Journal of the American Society for Information Science, 21(4):236, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elias:1970:Ec

[Eli70c] A. W. Elias. Editorial. Journal of the American Society for Information Science, 21(6):384, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elias:1987:HNF

[Eli87] A. W. Elias. Historical note: Fifty years of ASIS — thirty-eight years of JASIS (American Documentation). Journal of the American Society for Information Science, 38(5):385–386, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elliott:1970:RSC

[Ell70] Clark A. Elliott. The Royal Society catalogue as an index to nineteenth century American science. Journal of the American Society for Information Science, 21(6):396–401, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Elliott:1976:BCK

[Ell76] Roger W. Elliott. Brief communications: Kiviat-graphs as a means for displaying performance data for on-line retrieval systems. Journal of the American Society for Information Science, 27(3):178–182, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ellis:1986:ERL

[Ell86] David Ellis. European research letters: Social science information research. Journal of the American Society for Information Science, 37(2):86–88, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ellis:1996:DMI

[Ell96] David Ellis. The dilemma of measurement in information retrieval research. Journal of the American Society for Information Science, 47(1):23–36, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eddy:1994:EED

[EM94] William F. Eddy and Audris Mockus. An example of the estimation and display of a smoothly varying function of time and space — The incidence of the disease mumps. Journal of the American Society for Information Science, 45(9):686–693, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eres:1985:APS

[EN85] Beth Krevitt Eres and K. T. Bivins Noerr. Access to primary and secondary literature from peripheral or less developed countries. Journal of the American Society for Information Science, 36(3): 184–191, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Esler:1998:RES

[EN98] Sandra L. Esler and Michael L. Nelson. Research: Evolution of scientific and technical information distribution. Journal of the American Society for Information Science, 49(1):82–91, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Endres-Niggemeyer:1998:RPS

[ENN98] Brigitte Endres-Niggemeyer and Elisabeth Neugebauer. Research: Professional summarizing: No cognitive simulation without observation. Journal of the American Society for Information Science, 49(6):486–506, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Enser:1995:ERL

[Ens95] Peter Enser. European research letters: Image databases for multimedia projects. Journal of the American Society for Information Science, 46(1):60–64, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Everett:1991:RCL

[EP91] James E. Everett and Antony Pecotich. Research: A combined loglinear/MDS model for mapping journals by citation analysis. Journal of the American Society for Information Science, 42(6): 405–413, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Exon:1997:RSS

[EP97] F. C. A. Exon and Keith F. Punch. Research: The self-sufficient library collection: a test of assumptions. Journal of the American Society for Information Science, 48(1):11–16, January 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Epstein:1980:TLI

[Eps80] Hank Epstein. The technology of library and information networks. Journal of the American Society for Information Science, 31(6):425–437, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

El-Qawasmeh:2000:PIH

[EQH00] Eyas El-Qawasmeh and Ismail Hmeidi. Performance investigation of Hamming Distance Bit Vertical Counter applied to access methods in information retrieval. Journal of the American Society for Information Science, 51(5):427–431, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1991:RTP

[ER91] L. Egghe and R. Rousseau. Research: Transfer principles and a classification of concentration measures. Journal of the American Society for Information Science, 42(7):479–489, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1995:RGS

[ER95] Leo Egghe and Ronald Rousseau. Research: Generalized successbreeds-success principle leading to time-dependent informetric distributions. Journal of the American Society for Information Science, 46(6):426–445, July 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1997:LEE

[ER97] Leo Egghe and Ronald Rousseau. Letters to the Editor: Everett and Cater’s retrieval topology. Journal of the American Society for Information Science, 48(5):479–480, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:1998:RTA

[ER98] Leo Egghe and Ronald Rousseau. Research: Topological aspects of information retrieval. Journal of the American Society for Information Science, 49(13):1144–1160, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:2000:AOI

[ER00a] Leo Egghe and Ronald Rousseau. Aging, obsolescence, impact, growth, and utilization: Definitions and relations. Journal of the American Society for Information Science, 51(11):1004–1017, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:2000:POM

[ER00b] Leo Egghe and Ronald Rousseau. Partial orders and measures for language preferences. Journal of the American Society for Information Science, 51(12):1123–1130, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:2000:RIP

[ER00c] Leo Egghe and Ronald Rousseau. Research: The influence of publication delays on the observed aging distribution of scientific literature. Journal of the American Society for Information Science, 51(2):158–165, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ercegovac:1992:RMO

[Erc92] Zorana Ercegovac. Research: A multiple-observation approach in knowledge acquisition for expert systems: a case study. Journal of the American Society for Information Science, 43(7):506–517, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ercegovac:1995:BRB

[Erc95] Zorana Ercegovac. Book review: The systems librarian: the role of the library systems manager. Journal of the American Society for Information Science, 46(9):712–713, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ercegovac:1998:RML

[Erc98] Zorana Ercegovac. Research: Minimal level cataloging: What does it mean for maps in the contexts of card catalogs, online catalogs, and digital libraries? Journal of the American Society for Information Science, 49(8):706–719, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ercegovac:1999:RI

[Erc99] Zorana Ercegovac. Research: Introduction. Journal of the American Society for Information Science, 50(13):1165–1168, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eres:1980:BRB

[Ere80] Beth Krevitt Eres. Book review: -Legal Decisions and Information Systems. Jon Bing and Trygve Harvold. Oslo, Norway: Universitets Forlaget; 1977: 272 pp. Price: $20.00 (U.S. and Canadian Distributor: Columbia University Press). Journal of the American Society for Information Science, 31(5):383–384, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eres:1981:BRBa

[Ere81a] Beth Krevitt Eres. Book review: Information Science: Law enforcement applications. Emil H. Levine. No place given: Anderson Publishing Co.; 1979. No price given. Journal of the American Society for Information Science, 32(5):339–340, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eres:1981:BRBb

[Ere81b] Beth Krevitt Eres. Book review: Project SEARCH: the struggle for control of criminal information in America. Gordon Karl Zenk. Westport, CT: Greenwood Press; 1979. Price: $16.95. Journal of the American Society for Information Science, 32(5): 339–340, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eres:1981:TIT

[Ere81c] Beth Krevitt Eres. Transfer of information technology to less developed countries: a systems approach. Journal of the American Society for Information Science, 32(2):97–102, March 1981. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eres:1985:SCR

[Ere85] Beth Krevitt Eres. Socioeconomic conditions related to information activity in less developed countries. Journal of the American Society for Information Science, 36(3):213–219, May 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Egghe:2000:RMA

[ERV00] Leo Egghe, Ronald Rousseau, and Guido Van Hooydonk. Research: Methods for accrediting publications to authors or countries: Consequences for evaluation studies. Journal of the American Society for Information Science, 51(2):145–157, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Evans:1994:BRB

[Eva94] G. Edward Evans. Book review: Strategic management for academic libraries. Journal of the American Society for Information Science, 45(9):723–724, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Everett:1993:BCV

[Eve93] James E. Everett. Brief communications: The validity of Poisson assumptions in a combined loglinear/MDS mapping model. Journal of the American Society for Information Science, 44(3): 175–178, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Evens:1998:LER

[Eve98] Martha W. Evens. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 49(6):577, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Eggins:1970:RCS

[EW70] H. O. W. Eggins and M. J. D. Willsher. A reference collection scheme for a specialized information center based on the “cooperating specialist”. Journal of the American Society for Information Science, 21(2):128–132, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Exner:1997:BRB

[Exn97] Frank Exner. Book review: Explorations in indexing and abstracting: Pointing, virtue, and power. Journal of the American Society for Information Science, 48(3):282–283, March 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Exner:1998:BRB

[Exn98] Frank Exner (Little Bear). Book review: Advances in knowledge discovery and data mining. Journal of the American Society for Information Science, 49(4):386–387, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fagan:1989:REN

[Fag89] Joel L. Fagan. Research: The effectiveness of a nonsyntactic approach to automatic phrase indexing for document retrieval. Journal of the American Society for Information Science, 40(2): 115–132, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fagan:1995:POC

[Fag95] Phoebe Fagan. Perspectives on … the Chemist’s workstation: Getting your name on the right desktops, or how to be found on the Internet. Journal of the American Society for Information Science, 46(8):618–622, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Falkowski:1998:BCC

[Fal98] Bernd-Jürgen Falkowski. Brief communications: On certain generalizations of inner product similarity measures. Journal of the American Society for Information Science, 49(9):854–858, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Faulhaber:1996:ECU

[Fau96] Charles B. Faulhaber. Examples of classes using technological advances for media distribution and collaboration: Distance learning and digital libraries: Two sides of a single coin. Journal of the American Society for Information Science, 47(11):854–856, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fayollat:1972:LSCa

[Fay72a] James Fayollat. On-line serials control system in a large biomedical library: (1) Description of the system. Journal of the American Society for Information Science, 23(5):318–322, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fayollat:1972:LSCb

[Fay72b] James Fayollat. On-line serials control system in a large biomedical library part II. Evaluation of retrieval features. Journal of the American Society for Information Science, 23(6):353–358, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fayollat:1973:LSC

[Fay73] James Fayollat. On-line serials control system in a large biomedical library. Part III: Comparison of on-line and batch operations and cost analysis. Journal of the American Society for Information Science, 24(2):80–86, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frame:1978:CNI

[FB78] J. Davidson Frame and John J. Baum. Cross-national information flows in basic research: Examples taken from physics. Journal of the American Society for Information Science, 29(5):247– 252, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fugmann:1981:LEP

[FB81] Robert Fugmann and Bruce C. Bennion. Letters to the Editor: Performance testing of a book index. Journal of the American Society for Information Science, 32(5):334–335, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frants:1988:RNI

[FB88] Valery I. Frants and Craig B. Brush. Research: The need for information and some aspects of information retrieval systems construction. Journal of the American Society for Information Science, 39(2):86–91, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fletcher:1999:RI

[FB99] Patricia D. Fletcher and John Carlo Bertot. Research: Introduction. Journal of the American Society for Information Science, 50(4):295–298, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frame:1979:IAE

[FBC79] J. Davidson Frame, John J. Baum, and Michael Card. An information approach to examining developments in an energy technology: Coal gasification. Journal of the American Society for Information Science, 30(4):193–201, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

French:1997:RCA

[FBK97] James C. French, Donald E. Brown, and Nam-Ho Kim. Research: A classification approach to Boolean query reformulation. Journal of the American Society for Information Science, 48(8):694– 706, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frisse:1992:PLH

[FC92] Mark F. Frisse and Steve B. Cousins. Perspectives on … humancomputer interface: Models for hypertext. Journal of the American Society for Information Science, 43(2):183–191, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fan:1992:RUD

[FCJ^{+}92] David P. Fan, Julie M. Curtsinger, Lisa Johnson, Santosh Kalkhar, and Jeanette St. John. Research: Unitized Data System: a new formalism for encoding and presenting data from the scientific research literature. Journal of the American Society for Information Science, 43(5):371–383, June 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1999:RVI

[FDD^{+}99] Raya Fidel, Rachel K. Davies, Mary H. Douglass, Jenny K. Holder, Carla J. Hopkins, Elisabeth J. Kushner, Bryan K. Miyagishima, and Christina D. Toney. Research: A visit to the information mall: Web searching behavior of high school students. Journal of the American Society for Information Science, 50(1): 24–37, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fedanzo:1977:ECF

[Fed77] Anthony J. Fedanzo, Jr. Expanding the categories of feedback. Journal of the American Society for Information Science, 28(5): 239–246, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fedorowicz:1982:TFZ

[Fed82a] Jane Fedorowicz. The theoretical foundation of Zipf’s Law and its application to the bibliographic database environment. Journal of the American Society for Information Science, 33(5):285–293, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fedorowicz:1982:ZMA

[Fed82b] Jane Fedorowicz. A Zipfian model of an automatic bibliographic system: An application to MEDLINE. Journal of the American Society for Information Science, 33(4):223–232, July 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Feinberg:1973:LEC

[Fei73] Edwin H. Feinberg. Letters to the Editor: Core serials in biology. Journal of the American Society for Information Science, 24(6): 465, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fenichel:1981:OSM

[Fen81] Carol Hansen Fenichel. Online searching: Measures that discriminate among users with different types of experiences. Journal of the American Society for Information Science, 32(1):23–32, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Freudenthal:1980:IDC

[FF80] Juan R. Freudenthal and Héctor Gómez Fuentes. Information and documentation in chile: Progress report. Bibliography 1974– 1976. Journal of the American Society for Information Science, 31(6):445–448, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fox:1999:RSM

[FFKS99] Kevin L. Fox, Ophir Frieder, Margaret M. Knepper, and Eric J. Snowberg. Research: SENTINEL: a multiple engine information retrieval and visualization system. Journal of the American Society for Information Science, 50(7):616–625, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frick:1970:CSL

[FG70] Barbara F. Frick and John M. Ginski. Cardiovascular serial literature: Characteristics, productive journals, and abstracting/ indexing coverage. Journal of the American Society for Information Science, 21(5):338–344, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fischhoff:1981:SEU

[FGS81] Baruch Fischhoff, Bernard Goitein, and Zur Shapira. Subjective expected utility: a model of decision-making. Journal of the American Society for Information Science, 32(5):391–399, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fox:1993:POD

[FHN^{+}93] Edward A. Fox, Deborah Hix, Lucy T. Nowell, Dennis J. Brueni, William C. Wake, Lenwood S. Heath, and Durgesh Rao. Perspectives on … digital libraries: Users, user interfaces, and objects: Envision, a digital library. Journal of the American Society for Information Science, 44(8):480–491, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1984:ROS

[Fid84] Raya Fidel. Research: Online searching styles: a case-study-based model of searching behavior. Journal of the American Society for Information Science, 35(4):211–221, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1986:RTE

[Fid86] Raya Fidel. Research: Towards expert systems for the selection of search keys. Journal of the American Society for Information Science, 37(1):37–44, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1987:LER

[Fid87] Raya Fidel. Letters to the Editor: Rejoinder to Hans H. Wellisch’s letter. Journal of the American Society for Information Science, 38(3):217, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1991:RSSa

[Fid91a] Raya Fidel. Research: Searchers’ selection of search keys: I. The selection routine. Journal of the American Society for Information Science, 42(7):490–500, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1991:RSSb

[Fid91b] Raya Fidel. Research: Searchers’ selection of search keys: II. Controlled vocabulary or free-text searching. Journal of the American Society for Information Science, 42(7):501–514, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1991:SSS

[Fid91c] Raya Fidel. Searchers’ selection of search keys: III. Searching styles. Journal of the American Society for Information Science, 42(7):515–527, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1994:PLI

[Fid94] Raya Fidel. Perspectives on … indexing: User-centered indexing. Journal of the American Society for Information Science, 45(8): 572–576, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fischer:1970:STL

[Fis70] Capt. Stephen B. Fischer. A step towards library automation. Journal of the American Society for Information Science, 21(1): 34–39, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fishman:1979:BRB

[Fis79] Marilyn S. Fishman. Book review: Job Control Language (Wiley Self-teaching Guide). Ruth Ashley and Judi N. Fernandes, with Gary Deward Brown. New York: John Wiley & Sons, Inc.; 1978: 157 pp. Price: $4.95. ISBN: 0-471-03205-0. Journal of the American Society for Information Science, 30(1):62, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fishman:1981:BRBa

[Fis81a] Marilyn Fishman. Book review: Learning to program in structured COBOL, parts 1 and 2. Edward Yourdon, Chris Cane, Irish Sarson, and Timothy R. Lister. Englewood Cliffs, NJ: Prentice-Hall; 1979: xiv, 478 pp. ISBN 0-13-527713-2. Journal of the American Society for Information Science, 32(5):337–338, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fishman:1981:BRBb

[Fis81b] Marilyn Fishman. Book review: MIS concepts and design. Robert G. Murdick. Englewood Cliffs, NJ: Prentice-Hall; 1980: xiv, 610 pp. ISBN 0-13-585331-1. price $19.95 cloth. Journal of the American Society for Information Science, 32(5):338, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fischer:1999:BRB

[Fis99] William T. Fischer. Book review: Basic research methods for librarians. Journal of the American Society for Information Science, 50(5):477–478, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Feinman:1973:CSD

[FK73a] R. D. Feinman and K. L. Kwok. Classification of scientific documents by means of self-generated groups employing free language. Journal of the American Society for Information Science, 24(5): 382–396, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Feinman:1973:CJO

[FK73b] R. D. Feinman and K. L. Kwok. The construction of journaloriented classification schemes. Journal of the American Society for Information Science, 24(1):71–72, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fraenkel:1999:RIR

[FK99] Aviezri S. Fraenkel and Shmuel T. Klein. Research: Information retrieval from annotated texts. Journal of the American Society for Information Science, 50(10):845–854, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Flood:1972:BCT

[Flo72] Barbara Flood. Brief communications: Teaching abstracting and indexing. Journal of the American Society for Information Science, 23(5):332–333, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Flood:1999:HNS

[Flo99a] Barbara J. Flood. Historical note: The start of a stop list at Biological Abstracts. Journal of the American Society for Information Science, 50(12):1066, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Flood:1999:LEA

[Flo99b] Barbara J. Flood. Letter to the Editor: Active members as a proportion of total membership. Journal of the American Society for Information Science, 50(7):634, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Flood:2000:RDI

[Flo00] Barbara Flood. Research: Drexel’s Information Science M.S. degree program, 1963–1971: an insider’s recollections. Journal of the American Society for Information Science, 51(12):1137–1148, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Flynn:1997:WWE

[Fly97] Peter Flynn. W[h]ither the web? The extension or replacement of HTML. Journal of the American Society for Information Science, 48(7):614–621, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fischhoff:1986:RCD

[FM86] Baruch Fischhoff and Donald MacGregor. Research: Calibrating databases. Journal of the American Society for Information Science, 37(4):222–233, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fernandez:1997:LHI

[FM97] Félix Sagredo Fernández and Antonia García Moreno. Literature: History of Information Science in Spain: a selected bibliography. Journal of the American Society for Information Science, 48(4): 369–372, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Faber:1982:IRM

[FMN82] Roman F. Faber, Mieczyslaw M. Murastkiewicz, and Zbigniew M. Nowicki. Information Research Management in Poland. Journal of the American Society for Information Science, 33(5):338–340, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fugmann:1974:RCR

[FNNW74] Robert Fugmann, Herbert Nickelsen, Ingeborg Nickelsen, and Jakob H. Winter. Representation of concept relations using the TOSAR system of the IDC: Treatise III on information retrieval theory. Journal of the American Society for Information Science, 25(5):287–307, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ford:1991:ERL

[For91] Nigel Ford. European research letters: Knowledge-based information retrieval. Journal of the American Society for Information Science, 42(1):72–74, January 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ford:1999:GUI

[For99] Nigel Ford. The growth of understanding in information science: Towards a developmental model. Journal of the American Society for Information Science, 50(12):1141–1152, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ford:2000:RCS

[For00] Nigel Ford. Research: Cognitive styles and virtual environments. Journal of the American Society for Information Science, 51(6): 543–557, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fox:1986:PLO

[Fox86] Christopher Fox. Perspectives on … online searching: Future generation information systems. Journal of the American Society for Information Science, 37(4):215–219, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

French:2000:UCS

[FPS00] James C. French, Allison L. Powell, and Eric Schulman. Using clustering strategies for creating authority files. Journal of the American Society for Information Science, 51(8):774–786, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frakes:1986:BRB

[Fra86] William B. Frakes. Book review: Information and misinformation: an investigation of the notions of information, misinformation, informing, and misinforming. Journal of the American Society for Information Science, 37(1):48–49, January 1986. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frazier:1999:PLC

[Fra99] Kenneth Frazier. Perspective on … copyright and fair-use guidelines for education and libraries: What’s wrong with fair-use guidelines for the academic community? Journal of the American Society for Information Science, 50(14):1320–1323, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Freudenthal:1977:BCI

[Fre77] Juan R. Freudenthal. Brief communications: Information and documentation in chile: Progress report and bibliography 1970– 1975. Journal of the American Society for Information Science, 28(1):58–60, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Freeman:1980:LED

[Fre80] Robert R. Freeman. Letters to the Editor: Data bases — history of development and trends. Journal of the American Society for Information Science, 31(6):453, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Freides:1982:SSA

[Fre82] Thelma Freides. The state of secondary access services: Social sciences. Journal of the American Society for Information Science, 33(3):157–161, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frycki:1975:SSI

[FRGL75] S. J. Frycki, P. A. Roskos, C. F. Gerity, and I. S. Levett. Squibb science information system: Computerized selective dissemination, current awareness, and retrospective searching of pharmaceuticai literature. Journal of the American Society for Information Science, 26(3):174–183, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Friedlander:1973:CSI

[Fri73] Janet Friedlander. Clinician Search for Information. Journal of the American Society for Information Science, 24(1):65–69, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fricke:1997:RIU

[Fri97] Martin Frické. Research: Information using likeness measures. Journal of the American Society for Information Science, 48 (10):882–892, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Froehlich:1994:RRR

[Fro94] Thomas J. Froehlich. Research: Relevance reconsidered — Towards an agenda for the 21st century: Introduction to special topic issue on relevance research. Journal of the American Society for Information Science, 45(3):124–134, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frohmann:1996:BRB

[Fro96] Bernd Frohmann. Book review: The jobless future: Sci-tech and the dogma of work. Journal of the American Society for Information Science, 47(2):177–179, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fidel:1983:FAO

[FS83] Raya Fidel and Dagobert Soergel. Factors affecting online bibliographic retrieval: a conceptual framework for research. Journal of the American Society for Information Science, 34(3):163–180, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fryser:1984:RES

[FS84] Benjamin S. Fryser and Keith H. Stirling. Research: The effect of spatial arrangement, upper-lower case letter combinations, and reverse video on patron response to CRT displayed catalog records. Journal of the American Society for Information Science, 35(6):344–350, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frants:1991:RAA

[FS91a] Valery I. Frants and Jacob Shapiro. Research: Algorithm for automatic construction of query formulations in Boolean form. Journal of the American Society for Information Science, 42(1): 16–26, January 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frants:1991:RCF

[FS91b] Valery I. Frants and Jacob Shapiro. Research: Control and feedback in a documentary information retrieval system. Journal of the American Society for Information Science, 42(9):623–634, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fausey:1997:AMD

[FS97a] Jon Fausey and Keith Shafer. All my data is in SGML. Now what? Journal of the American Society for Information Science, 48(7):638–643, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fayet-Scribe:1997:CFU

[FS97b] Sylvie Fayet-Scribe. The cross-fertilization of the U.S. Public Library model and the French documentation model (IIB, French correspondent of FID) through the French professional associations between World War I and World War II. Journal of the American Society for Information Science, 48(9):782–793, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frants:1999:RBS

[FSTV99] Valery I. Frants, Jacob Shapiro, Isak Taksa, and Vladimir G. Voiskunskii. Research: Boolean search: Current state and perspectives. Journal of the American Society for Information Science, 50(1):86–95, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Frants:1993:RMI

[FSV93] Valery I. Frants, Jacob Shapiro, and Vladimir G. Voiskunskii. Research: Multiversion information retrieval systems and feedback with mechanism of selection. Journal of the American Society for Information Science, 44(1):19–27, January 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fu:1996:BRBa

[Fu96a] Tina C. Fu. Book review: Academic libraries as high-tech gateways: a guide to design and space decisions. Journal of the American Society for Information Science, 47(5):401–402A, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fu:1996:BRBb

[Fu96b] Tina C. Fu. Book review: Academic libraries as high-tech gateways: a guide to design and space decisions. Journal of the American Society for Information Science, 47(5):401–402A, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fugmann:1985:FAT

[Fug85] Robert Fugmann. The five-axiom theory of indexing and information supply. Journal of the American Society for Information Science, 36(2):116–129, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fuhr:2000:RPD

[Fuh00] Norbert Fuhr. Research: Probabilistic datalog: Implementing logical information retrieval for advanced applications. Journal of the American Society for Information Science, 51(2):95–110, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Feng:1988:PLI

[FW88] Cyril C. H. Feng and Frieda O. Weise. Perspectives on … integrated academic information management systems (IAIMS): Part III. Implementation of integrated information services. Library/ computer center partnership. Journal of the American Society for Information Science, 39(2):126–130, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fox:1990:RUV

[FW90] Edward A. Fox and Sheila G. Winett. Research: Using vector and extended Boolean matching in an expert system for selecting foster homes. Journal of the American Society for Information Science, 41(1):10–26, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Fletcher:1999:RCR

[FW99] Patricia D. Fletcher and Lisa K. Westerback. Research: Catching a ride on the NII: The Federal policy vehicles paving the information highway. Journal of the American Society for Information Science, 50(4):299–304, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ferber:1995:RAM

[FWR95] Reginald Ferber, Manfred Wettler, and Reinhard Rapp. Research: An associative model of word selection in the generation of search queries. Journal of the American Society for Information Science, 46(9):685–699, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Galbiati:1991:RPB

[Gal91] Giulia Galbiati. Research: A phrase-based matching function. Journal of the American Society for Information Science, 42(1): 36–48, January 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gannett:1974:BRB

[Gan74] E. K. Gannett. Book review: Economics of scientific publications, CBE committee on economics of publication (eds.), Washington, DC: Council of Biology Editors, 78 pp. (1973). Journal of the American Society for Information Science, 25(6):381–382, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1971:ISI

[Gar71] Eugene Garfield. Information science and the informationconscious society. Journal of the American Society for Information Science, 22(2):71–73, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1976:PSI

[Gar76] Eugene Garfield. The permuterm subject index: an autobiographical review. Journal of the American Society for Information Science, 27(5):288–291, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1979:LEH

[Gar79] Eugene Garfield. Letters to the Editor: Highly cited old papers. Journal of the American Society for Information Science, 30(4): 234–235, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1983:LEC

[Gar83] Eugene Garfield. Letters to the Editor: Citation indexes in information retrieval. Journal of the American Society for Information Science, 34(4):298–299, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garay:1986:BRB

[Gar86] Ronald Garay. Book review: Television’s window on the world: International affairs coverage on the U.S. networks. Journal of the American Society for Information Science, 37(3):173–174, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garrison:1988:PLE

[Gar88] Guy Garrison. Perspectives on … education of the information professional: New dimensions, new directions: Part V. The future. Challenges to information science education. Journal of the American Society for Information Science, 39(5):362–366, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1990:PLE

[Gar90] Eugene Garfield. Perspectives on … evaluation of scientific information: Peer review and the impact of new information technology: Response to the panel on evaluation of scientific information and the impact of new information technology. Journal of the American Society for Information Science, 41(3):229–230, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1991:LEA

[Gar91] Eugene Garfield. Letters to the Editor: Availability of indexes. Journal of the American Society for Information Science, 42(6): 464, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garcia:1992:PLI

[Gar92] D. Linda Garcia. Perspectives on … information technology standards: Standard setting in the United States: Public and private sector roles. Journal of the American Society for Information Science, 43(8):531–537, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1997:LEV

[Gar97] Eugene Garfield. Letters to the Editor: Validation of citation analysis. Journal of the American Society for Information Science, 48(10):962, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1998:LEM

[Gar98] Eugene Garfield. Letters to the Editor: The multiple meanings of impact factors. Journal of the American Society for Information Science, 49(8):768, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gasaway:1998:PLI

[Gas98] Laura N. Gasaway. Perspectives on … Internet issues/publishing issues: Copyright, the Internet, and other legal issues. Journal of the American Society for Information Science, 49(11):1003–1009, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gasaway:1999:PLC

[Gas99] Laura N. Gasaway. Perspective on … copyright and fair-use guidelines for education and libraries: Guidelines for distance learning and interlibrary loan: Doomed and more doomed. Journal of the American Society for Information Science, 50(14): 1337–1341, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gathegi:1992:RSS

[Gat92] John Ng’ang’a Gathegi. Research: The state and society: Intervention in the creation of scientific information in developing countries. Journal of the American Society for Information Science, 43(4):323–333, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gauch:1992:PLH

[Gau92] Susan Gauch. Perspectives on … human-computer interface: Intelligent information retrieval: an introduction. Journal of the American Society for Information Science, 43(2):175–182, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Green:1995:RTRb

[GB95] Rebecca Green and Carol A. Bean. Research: Topical relevance relationships. II. An exploratory study and preliminary typology. Journal of the American Society for Information Science, 46(9):654–662, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goossens:1976:IPF

[GBDH76] Paula C. Goossens, Jaklien M. Bourgois-Derison, and Andre B. Hubert. The internal processing format for bibliographic data at the Royal Library of Belgium. Journal of the American Society for Information Science, 27(3):147–152, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1989:MMK

[GBL89] Michael Gordon, David Blair, and Robert Lindsay. In memoriam: Manfred Kochen 1928–1989: Remembrances of a scholar and a gentle man. Journal of the American Society for Information Science, 40(4):223–225, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Galal:1999:REP

[GCH99] Gehad M. Galal, Diane J. Cook, and Lawrence B. Holder. Research: Exploiting parallelism in a structural scientific discovery system to improve scalability. Journal of the American Society for Information Science, 50(1):65–73, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gaines:1997:MHF

[GCS97] Brian R. Gaines, Lee Li-Jen Chen, and Mildred L. G. Shaw. Modeling the human factors of scholarly communities supported through the Internet and World Wide Web. Journal of the American Society for Information Science, 48(11):987–1003, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gey:1990:LEF

[GD90] Fredric Gey and Daniel P. Dabney. Letters to the Editor: Fulltext against intellectual indexing controversy. Journal of the American Society for Information Science, 41(8):613–614, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1998:RUL

[GD98] Michael D. Gordon and Susan Dumais. Research: Using latent semantic indexing for literature based discovery. Journal of the American Society for Information Science, 49(8):674–685, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Geller:1981:LCR

[GdCD81] Nancy L. Geller, John S. de Cani, and Robert E. Davies. Lifetimecitation rates: a mathematical model to compare scientists’ work. Journal of the American Society for Information Science, 32(1): 1–15, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Geller:1979:BCT

[Gel79] Nancy L. Geller. Brief communication: A test of significance for the Whitworth distribution. Journal of the American Society for Information Science, 30(4):229–231, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gelfand:1996:BRB

[Gel96] Julia Gelfand. Book review: Contextual media: Multimedia and interpretation. Journal of the American Society for Information Science, 47(10):796–797A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gerson:1978:CTP

[Ger78] Gordon M. Gerson. Cliqueing — a technique for producing maximally connected clusters. Journal of the American Society for Information Science, 29(3):125–129, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gerson:1979:BRB

[Ger79] Gordon M. Gerson. Book review: Generalized Data Management Systems and Scientific Information. Report of the Specialist Study on Computer Software. The Use of Generalized Data Management Systems for Handling Scientific Information. OECD Nuclear Energy Agency; 1978: Paris, France: 346 pp. $12.50 hard copy, $3.00 microfiche (CONF 771030). Journal of the American Society for Information Science, 30(2):119, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grossman:1997:RIS

[GFHR97] David A. Grossman, Ophir Frieder, David O. Holmes, and David C. Roberts. Research: Integrating structured data and text: a relational approach. Journal of the American Society for Information Science, 48(2):122–132, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Girill:1991:RESa

[GGJ91] T. R. Girill, Thomas Griffin, and Robert B. Jones. Research: Extended subject access to hypertext online documentation. Parts I and II: the search-support and maintenance problems. Journal of the American Society for Information Science, 42(6):414–426, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Geldenhuys:1996:RCT

[GH96] Gerhard Geldenhuys and Lezánne Human. Research: Comparison of two indices for the investigation of cooperation graphs. Journal of the American Society for Information Science, 47(5): 380–384, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gibson:1986:PLT

[Gib86] Charles E. Gibson. Perspectives on … telecommunications: Principles, developments, prospects: The teleprocessing environment. Journal of the American Society for Information Science, 37(6):429–435, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gilchrist:1978:CCI

[Gil78] Bruce Gilchrist. Coping with catastrophe: Implications to information systems design. Journal of the American Society for Information Science, 29(6):271–277, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Girill:1991:RESb

[Gir91] T. R. Girill. Research: Extended subject access to hypertext online documentation. Part III: the document-boundaries problem. Journal of the American Society for Information Science, 42(6): 427–437, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1989:RRP

[GK89] Michael Gordon and Manfred Kochen. Research: Recall-precision trade-off: a derivation. Journal of the American Society for Information Science, 40(3):145–151, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1991:RUT

[GL91] Michael D. Gordon and Peter Lenk. Research: A utility theoretic examination of the probability ranking principle in information retrieval. Journal of the American Society for Information Science, 42(10):703–714, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1992:RWP

[GL92] Michael D. Gordon and Peter Lenk. Research: When is the probability ranking principle suboptimal? Journal of the American Society for Information Science, 43(1):1–14, January 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garrett:1993:POD

[GL93] John R. Garrett and Patrice A. Lyons. Perspectives on … digital libraries: Toward an Electronic Copyright Management System. Journal of the American Society for Information Science, 44(8): 468–473, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1996:RTD

[GL96] Michael D. Gordon and Robert K. Lindsay. Research: Toward discovery support systems: a replication, re-examination, and extension of Swanson’s work on literature-based discovery of a connection between Raynaud’s and fish oil. Journal of the American Society for Information Science, 47(2):116–128, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gleason:1993:BCQ

[Gle93] John M. Gleason. Brief communications: Questionable validity of Poisson assumptions in a combined loglinear/MDS mapping model. Journal of the American Society for Information Science, 44(1):57–59, January 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Glickert:1972:BCP

[Gli72] Peter Glickert. Brief communications: Psychology and/of computer oriented linguistics. Journal of the American Society for Information Science, 23(2):131–133, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Glickert:1974:POC

[Gli74] Peter Glickert. Patent office classification: Its whats and whys. Journal of the American Society for Information Science, 25(5): 308–311, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Glickert:1978:LEC

[Gli78] Peter Glickert. Letters to the Editor: Cliqueing and patent classification. Journal of the American Society for Information Science, 29(6):314, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gomez:1990:RAR

[GLL90] Louis M. Gomez, Carol C. Lochbaum, and Thomas K. Landauer. Research: All the right words: Finding what you want as a function of richness of indexing vocabulary. Journal of the American Society for Information Science, 41(8):547–559, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Glogoff:1988:RRG

[Glo88] Stuart Glogoff. Research: Reviewing the gatekeepers: a survey of referees of library journals. Journal of the American Society for Information Science, 39(6):400–407, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goodhue:1997:PLIb

[GLS97] Dale Goodhue, Russell Littlefield, and Detmar W. Straub. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: The measurement of the impacts of the IIC on the end-users: the survey. Journal of the American Society for Information Science, 48(5): 454–465, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gluck:1990:RRJ

[Glu90] Myke Gluck. Research: A review of journal coverage overlap with an extension to the definition of overlap. Journal of the American Society for Information Science, 41(1):43–60, January 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gluck:1994:SII

[Glu94] Myke Gluck. Spatial information and information science: Introduction to JASIS’ special topic issue on spatial information. Journal of the American Society for Information Science, 45(9):640– 644, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gluck:1995:BRB

[Glu95a] Myke Gluck. Book review: Hypertext and the technology of conversation: Orderly situational choice. Journal of the American Society for Information Science, 46(5):390–392, June 1995. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gluck:1995:RUP

[Glu95b] Myke Gluck. Research: Understanding performance in information systems: Blending relevance and competence. Journal of the American Society for Information Science, 46(6):446–460, July 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Griffiths:1986:RUI

[GLW86] Alan Griffiths, H. Claire Luckhurst, and Peter Willett. Research: Using interdocument similarity information in document retrieval systems. Journal of the American Society for Information Science, 37(1):3–11, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grant:1986:RCM

[GM86] Frances L. Grant and Robert G. Main. Research: Curriculum 1984: Meeting the needs of the information age. Journal of the American Society for Information Science, 37(1):12–19, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Greene:2000:POD

[GMPS00] Stephan Greene, Gary Marchionini, Catherine Plaisant, and Ben Shneiderman. Previews and overviews in digital libraries: Designing surrogates to support visual information seeking. Journal of the American Society for Information Science, 51(4):380–393, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gleeson:1993:RUD

[GO93] Michael E. Gleeson and John R. Ottensmann. Research: Using data from computerized circulation and cataloging systems for management decision making in public libraries. Journal of the American Society for Information Science, 44(2):94–100, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goetschel:1987:RCA

[Goe87] Roy Goetschel, Jr. Research: Clustering analysis for graphs with multiweighted edges: a unified approach to the threshold problem. Journal of the American Society for Information Science, 38(3):156–161, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Golashvili:1973:PCE

[Gol73] T. V. Golashvili. Problems of compilation and evaluation of property data of substances and materials. Journal of the American Society for Information Science, 24(1):57–64, January 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goldstain:1974:LER

[Gol74a] Charles M. Goldstain. Letters to the Editors: Reference question analysis and search strategy development by man and machine. Journal of the American Society for Information Science, 25(6): 384, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goldstein:1974:BRB

[Gol74b] Charles Goldstein. Book review: Structured Information Files. Richard A. Kaimann: Melville Publishing Company, Los Angeles (1973). Journal of the American Society for Information Science, 25(5):336, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goldstein:1979:LES

[Gol79] Isai Goldstein. Letter to the Editors: Science, Shcharansky, and the Soviets. Journal of the American Society for Information Science, 30(3):173–175, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goldmann:1987:LEI

[Gol87] Nahum Goldmann. Letters to the Editor: Information science research: Where is the meat? Journal of the American Society for Information Science, 38(3):215–216, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gold:1994:PLE

[Gol94] Jon D. Gold. Perspectives on … electronic publishing: An electronic publishing model for academic publishers. Journal of the American Society for Information Science, 45(10):760–764, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goldfarb:1997:SRW

[Gol97] Charles F. Goldfarb. SGML: the reason why and the first published hint. Journal of the American Society for Information Science, 48(7):656–661, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goodhue:1997:PLIa

[Goo97] Dale Goodhue. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: The model underlying the measurement of the impacts of the IIC on the end-users. Journal of the American Society for Information Science, 48(5):449–453, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goodrum:1999:BRB

[Goo99] Abby A. Goodrum. Book review: Principles of multimedia database systems. Journal of the American Society for Information Science, 50(4):382–383, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1979:DSI

[Gor79] Michael D. Gordon. Deficiencies of scientific information access and output in less developed countries. Journal of the American Society for Information Science, 30(6):340–342, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1982:CRV

[Gor82] Michael D. Gordon. Citation ranking versus subjective evaluation in the determination of journal hierarchies in the social sciences. Journal of the American Society for Information Science, 33(1): 55–57, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1990:REE

[Gor90] Michael D. Gordon. Research: Evaluating the effectiveness of information retrieval systems using simulated queries. Journal of the American Society for Information Science, 41(5):313–323, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gordon:1991:RUB

[Gor91] Michael D. Gordon. Research: User-based document clustering by redescribing subject descriptions with a genetic algorithm. Journal of the American Society for Information Science, 42(5): 311–322, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gorbunov:1992:LEM

[Gor92] Audrey L. Gorbunov. Letters to the Editor: Multimedia: What else? Journal of the American Society for Information Science, 43(9):644, October 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gorman:1995:POM

[Gor95] Paul N. Gorman. Perspectives on … medical informactics: Information technology in health care: Information needs of physicians. Journal of the American Society for Information Science, 46(10):729–736, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gould:1974:UPP

[Gou74] Angela M. Gould. User preference in published indexes. Journal of the American Society for Information Science, 25(5):279–286, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goyal:1984:IDS

[Goy84] Pankaj Goyal. An investigation of different string coding methods. Journal of the American Society for Information Science, 35(4):248–252, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goyal:1987:BCI

[Goy87] Pankaj Goyal. Brief communications: Implications of program properties for efficient storage and compilation. Journal of the American Society for Information Science, 38(2):138–141, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Galloway:1983:IPV

[GP83] Emily Galloway and Judith Paris. Information providers and videodisc/optical disk technology. Journal of the American Society for Information Science, 34(6):414–416, November 1983. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gomez:1998:UCC

[GQSdP98] Enrique J. Gómez, José A. Quiles, Marcos F. Sanz, and Francisco del Pozo. A user-centered cooperative information system for medical imaging diagnosis. Journal of the American Society for Information Science, 49(9):810–816, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gupta:1974:OSB

[GR74] Surendra Mohan Gupta and Arunachalam Ravindran. Optimal storage of books by size: an operations research approach. Journal of the American Society for Information Science, 25(6):354– 357, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See notes [Buc75].

Granick:1982:ERC

[Gra82] Lois W. Granick. The emergence and role of common interest groups in secondary information. Journal of the American Society for Information Science, 33(3):175–182, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Greenberg:1990:ROB

[Gre90] Irwin Greenberg. Research: Optimal binary prefix-condition codes with ordinal probabilities. Journal of the American Society for Information Science, 41(5):330–333, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Greenstein:1992:PLI

[Gre92] Shane M. Greenstein. Perspectives on … information technology standards: Invisible hands and visible advisors: an economic interpretation of standardization. Journal of the American Society for Information Science, 43(8):538–549, September 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Green:1995:RTRa

[Gre95] Rebecca Green. Research: Topical relevance relationships. I. Why topic matching fails. Journal of the American Society for Information Science, 46(9):646–653, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Green:1996:BRB

[Gre96] Rebecca Green. Book review: Typologies and taxonomies: an introduction to classification techniques. Journal of the American Society for Information Science, 47(4):328–329, April 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Griffiths:1980:LEV

[Gri80] John B. Griffiths. Letters to the Editor: On the value of information. Journal of the American Society for Information Science, 31(4):303–304, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Ozo80].

Griffith:1981:II

[Gri81] Belver C. Griffith. Introduction: Introduction. Journal of the American Society for Information Science, 32(5):344–346, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grishman:1984:NLP

[Gri84] Ralph Grishman. Natural language processing. Journal of the American Society for Information Science, 35(5):291–296, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Griffith:1988:REF

[Gri88] Belver C. Griffith. Research: Exact fits to large ranked, bibliometric distributions. Journal of the American Society for Information Science, 39(6):423–427, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1975:BRBa

[Gro75a] Audrey N. Grosch. Book review: Australian MARC specification. National Library of Australia. Canberra: National Library of Australia 83 pp. (1973). Journal of the American Society for Information Science, 26(1):66, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1975:BRBb

[Gro75b] Audrey N. Grosch. Book review: Women in librarianship — Melvil’s rib symposium. Edited by Margaret Myers and Mayra Scarborough, New Brunswick, NJ: Bureau of Library and Information Science Research, Rutgers University Graduate School of Library Science, 112 pp. (1975) (ISBN 0-8135-0807-X). Journal of the American Society for Information Science, 26(6):353, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1976:BRBa

[Gro76a] Audrey N. Grosch. Book review: INTERMARC (M): Format bibliographique d’échange pour les monographies: manuel. Group INTERMARC. Paris: diffusion Bibliotheque nationale; Brussels: diffusion Bibliotheque royale Albert Ier. 212 pp. (1975). (ISBN 2-7177-1206-2). Journal of the American Society for Information Science, 27(3):185–186, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1976:BRBb

[Gro76b] Audrey N. Grosch. Book review: Terminology of documentation = terminologie de la documentation = terminologie der dokumentation = terminologie der dokumentation = terminologia de la documentacion. Gernot Versig and Ulrich Neveling, compilers. Paris: The Unesco Press. 274 pp. (1976) (ISBN 92-3-001232-7). Journal of the American Society for Information Science, 27(5): 348–349, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1976:BRBc

[Gro76c] Audrey N. Grosch. Book review: What happens in library filing? Herbert H. Hoffman. Hamden, CT: Linnet Books, 176 pp. (1976). Journal of the American Society for Information Science, 27(5): 350–351, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1977:BRB

[Gro77] Audrey N. Grosch. Book review: Landmarks of Library Literature 1876–1976. Dianne J. Ellsworth and Norman D. Stevens, Eds. Metuchen, NJ: The Scarecrow Press, Inc., 520 pp. (1976) (LC 754.5139; ISBN 0-8108-0899-4). Journal of the American Society for Information Science, 28(5):305–306, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1978:LEW

[Gro78] Audrey N. Grosch. Letter to the Editor: A work study of the review production process. Journal of the American Society for Information Science, 29(1):48, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1987:BRB

[Gro87] Audrey N. Grosch. Book review: Micrographics, 2nd ed. Journal of the American Society for Information Science, 38(3):220, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1988:BRB

[Gro88] Audrey N. Grosch. Book review: Libraries and information science in the electronic age. Journal of the American Society for Information Science, 39(1):24–25, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grosch:1990:BRB

[Gro90] Audrey N. Grosch. Book review: The complete cataloging reference set — Collected manuals of the Minnesota AACR@ trainers. Journal of the American Society for Information Science, 41(2): 147–148, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grose:1996:RRB

[Gro96] Teresa Olwick Grose. Research: Reading the bones: Information content, value, and ownership issues raised by the Native American Graves Protection and Repatriation Act. Journal of the American Society for Information Science, 47(8):624–631, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gebhardt:1978:OPD

[GS78] Friedrich Gebhardt and Imant Stellmacher. Opinion paper: Design criteria for documentation retrieval languages. Journal of the American Society for Information Science, 29(4):191–199, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gelman:1987:RLB

[GS87] E. Gelman and H. S. Sichel. Research: Library book circulation and the beta-binomial distribution. Journal of the American Society for Information Science, 38(1):4–12, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1993:BCK

[GS93a] Eugene Garfield and Irving H. Sher. Brief communications: Key-Words Plus^{T M} — algorithmic derivative indexing. Journal of the American Society for Information Science, 44(5):298–299, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gauch:1993:RES

[GS93b] Susan Gauch and John B. Smith. Research: An expert system for automatic query reformation. Journal of the American Society for Information Science, 44(3):124–136, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gilyarevskii:1993:RRE

[GS93c] R. S. Gilyarevskii and M. M. Subbotin. Research: Russian experience in hypertext: Automatic compiling of coherent texts. Journal of the American Society for Information Science, 44(4): 185–193, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Garfield:1997:LEC

[GS97] Eugene Garfield and Henry Small. Letters to the Editor: Citation format. Journal of the American Society for Information Science, 48(10):963–963b, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gupta:1974:TMN

[GSD74] Anand B. Gupta, Donald L. Shirey, and Anthony Debons. Trends in manpower needs in information science from 1967 to 1982. Journal of the American Society for Information Science, 25(1): 33–43, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Grupp:1998:STI

[GSK98] Hariolf Grupp, Ulrich Schmoch, and Knut Koschatzky. Science and technology infrastructure in Baden-Wuerttemberg and its orientation towards future regional development. Journal of the American Society for Information Science, 49(1):18–29, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gilliland-Swetland:2000:RAD

[GSKL00] Anne J. Gilliland-Swetland, Yasmin B. Kafai, and William E. Landis. Research: Application of Dublin Core metadata in the description of digital primary sources in elementary school classrooms. Journal of the American Society for Information Science, 51(2):193–201, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Glickert:1970:LE

[GT70] Peter Glickert and H. B. Thompson. Letters to the Editor. Journal of the American Society for Information Science, 21(2):167, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Guerra:1978:BCB

[Gue78] Roberto Guerra. Brief communication: Current Contents’ “weekly subject index” as a post-coordinate index. Journal of the American Society for Information Science, 29(1):47, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gull:1985:BRB

[Gul85] C. D. Gull. Book review: Development of resource sharing networks: Networks study no. 22. ABN Conference (1983: Melbourne, Victoria). First ABN Conference, July 12–14, 1983, Melbourne, Papers and Proceedings, Australian Bibliographic Network. Canberra: National Library of Australia; 1983: 214 pp. Journal of the American Society for Information Science, 36(2): 134–135, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gull:1987:HNI

[Gul87] Cloyd Dake Gull. Historical note: Information science and technology: From coordinate indexing to the global brain. Journal of the American Society for Information Science, 38(5):338–366, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Gupta:1990:ROP

[Gup90] Usha Gupta. Research: Obsolescence of physics literature: Exponential decrease of the density of citations to Physical Review articles with age. Journal of the American Society for Information Science, 41(4):282–287, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Goetschel:1987:ROC

[GV87] Roy Goetschel, Jr. and William Voxman. Research: Optimal clustering in graphs with weighted edges: a unified approach to the threshold problem. Journal of the American Society for Information Science, 38(1):13–20, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Griffith:1986:RTM

[GWDS86] Belver C. Griffith, Howard D. White, M. Carl Drott, and Jerry D. Saye. Research: Tests of methods for evaluating bibliographic databases: an analysis of the National Library of Medicine’s handling of literatures in the medical behavioral sciences. Journal of the American Society for Information Science, 37(4):261–270, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hjorland:1995:RTN

[HA95] Birger Hjørland and Hanne Albrechtsen. Research: Toward a new horizon in information science: Domain-analysis. Journal of the American Society for Information Science, 46(6):400–425, July 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hegazi:1987:BCI

[HAA87] Nadia Hegazi, Nabil Ali, and Ehsan Abed. Brief communications: Information content in textual data: Revisited for Arabic text. Journal of the American Society for Information Science, 38(2): 133–137, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haas:1993:RIS

[Haa93] Stephanie W. Haas. Research: Incomplete sentence quotations in books and journals. Journal of the American Society for Information Science, 44(7):398–405, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haas:1996:BRB

[Haa96] Stephanie W. Haas. Book review: The creative process: a computer model of storytelling and creativity. Journal of the American Society for Information Science, 47(3):252–254, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haas:1997:RDV

[Haa97] Stephanie W. Haas. Research: Disciplinary variation in automatic sublanguage term identification. Journal of the American Society for Information Science, 48(1):67–79, January 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haiqi:1997:BCM

[Hai97] Zhang Haiqi. Brief communications: More authors, more institutions, and more funding sources: Hot papers in biology from 1991 to 1993. Journal of the American Society for Information Science, 48(7):662–666, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hales:1980:EII

[Hal80] Celia Hales. The evolution of the international information system on research in documentation (ISORID). Journal of the American Society for Information Science, 31(4):293–297, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hall:1981:PUI

[Hal81] Homer J. Hall. Patterns in the use of information: the right to be different. Journal of the American Society for Information Science, 32(2):103–112, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Halpin:1985:BCT

[Hal85] Peter Halpin. Brief communications: Transferring a national information system from the public sector to the private sector — How the administration on aging did it. Journal of the American Society for Information Science, 36(1):53–55, January 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hall:1995:LEC

[Hal95] Homer J. Hall. Letters to the Editor: Caesar’s horse. Journal of the American Society for Information Science, 46(5):397, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hamilton:2000:RIM

[Ham00] Anne Hamilton. Research: Interface metaphors and logical analogues: a question of terminology. Journal of the American Society for Information Science, 51(2):111–122, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hankins:1975:BPQ

[Han75] Robert B. Hankins. Assets Protection, (a quarterly international periodical). Madison, Wisconsin: Assets Protection Journal Publishing, Vol. 1, Number 1, 78 pp. (Spring 1975). Journal of the American Society for Information Science, 26(6):354, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hansen:1986:REP

[Han86] Kathleen A. Hansen. Research: The effect of presearch experience on the success of naive (end-user) searches. Journal of the American Society for Information Science, 37(5):315–318, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Handley:1991:BCS

[Han91] John C. Handley. Brief communications: Sampling by length. Journal of the American Society for Information Science, 42(3): 229–232, April 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hanson:1992:LEB

[Han92] Dan Hanson. Letters to the Editor: Brush up your Shakespeare. Journal of the American Society for Information Science, 43(2): 151, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hannah:1996:BRB

[Han96] Stan Hannah. Book review: Building IBM: Shaping an industry and its technology. Journal of the American Society for Information Science, 47(3):256–257, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hapke:1999:RHS

[Hap99] Thomas Hapke. Research: History of scholarly information and communication: a review of selected German literature. Journal of the American Society for Information Science, 50(3):229–232, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haris:1970:BRB

[Har70a] Jessica L. Haris. Book review: 5/70-2R. Breve introducción a la computacion Lingüistica; Con Dos Apendices por el Misnio Autor. 1969. Paul L. Garvin. Departmento de Lingüistica, Universidad Nacional Mayor de San Marcos, Lima, Peru. 104 pp. Journal of the American Society for Information Science, 21(5): 376, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harris:1970:BRBc

[Har70b] Jessica L. Harris. Book review: Computerized library catalogs: Their growth, cost, and utility. 1969. J. L. Dolby, V. J. Forsyth, and H. L. Resnikoff. M.I.T. Press, Cambridge, Mass. 164 pp. $10.00. Journal of the American Society for Information Science, 21(3):230–231, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harris:1970:BRBb

[Har70c] Jessica L. Harris. Book review: On research libraries; statement and recommendations of the Committee on Research Libraries of the American Council of Learned Societies. Submitted to the National Advisory Commission on Libraries, November 1967. c1969. M.I.T. Press, Cambridge, Mass. 104 pp. $5.95. Journal of the American Society for Information Science, 21(3):230, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harris:1970:BRBa

[Har70d] Jessica L. Harris. Book review: Scientific and technological communication. 1969. Sidney Passman. Pergamon, Oxford, New York. 151 pp. Journal of the American Society for Information Science, 21(3):229–230, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harmon:1971:OPE

[Har71a] Glynn Harmon. Opinion papers: On the evolution of information science. Journal of the American Society for Information Science, 22(4):235–241, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harris:1971:BRB

[Har71b] Jessica L. Harris. Book review: Current Index to Journals in Education. New York, CCM Information Corporation, 1969-Monthly, with semi-annual and annual cumulations. Journal of the American Society for Information Science, 22(2):143–145, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harte:1971:GEF

[Har71c] Robert A. Harte. Guest editorial: FID looks at the 70’s. Journal of the American Society for Information Science, 22(1):1–4, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1971:BCO

[Har71d] Stephen P. Harter. Brief communications: Optimum number of cards per file guide assuming binary searching. Journal of the American Society for Information Science, 22(2):140, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hartson:1974:RAI

[Har74] H. Rex Hartson. The role of automatic indexing in access control: a modular view. Journal of the American Society for Information Science, 25(3):162–171, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1975:PAAa

[Har75a] Stephen P. Harter. A probabilistic approach to automatic keyword indexing. Part I. On the distribution of specialty words in a technical literature. Journal of the American Society for Information Science, 26(4):197–206, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1975:PAAb

[Har75b] Stephen P. Harter. A probabilistic approach to automatic keyword indexing. Part II. An algorithm for probabilistic indexing. Journal of the American Society for Information Science, 26(5): 280–289, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1975:BRBb

[Har75c] John F. Harvey. Book review: Maurice Falcolm Tauber, a biobibliography, 1934–1973. Marion C. Szigethy, Published for Beta Phi Mu, Nu Chapter, Columbia University School of Library Service by The Scarecrow Press, Inc., 121 pp. (1974). Journal of the American Society for Information Science, 26(6):354–355, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1975:BRBa

[Har75d] John F. Harvey. Book review: The Academic Library: In Honor of Guy R. Lyle. Edited by Ivan Ira Farber and Ruth Walling.

Metuchen, N.J.: Scarecrow Press, 171 pp. (1974). Journal of the American Society for Information Science, 26(2):131, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1976:BRBc

[Har76a] John F. Harvey. Book review: Collective bargaining in libraries. Frederick A. Schlipf, ed. Urbana-Champaign, IL: University of Illinois Graduate School of Library Science, (Allerton Park Institute No. 20) 179 pp. (1975) (ISBN: 0-87845-0424). Journal of the American Society for Information Science, 27(5):350, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1976:BRBb

[Har76b] John F. Harvey. Book review: National library buildings; proceedings of the colloquium held in Rome, 3–6 September 1973. Anthony Thompson, ed. Pollach Munchen: Verlag Dokumentation, IFLA Publications 2; 144 pp. (1975). (ISBN 3-7940-4422-3). Journal of the American Society for Information Science, 27(3): 190–191, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1976:BRBa

[Har76c] John F. Harvey. Book review: Reader on the library building. Hal B. Schell. Englewood. CO: Microcard Editions Books. 360 pp. (August 1975) (ISBN 0-910972-43-5). Journal of the American Society for Information Science, 27(3):188–189, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1977:BRBa

[Har77a] John F. Harvey. Book review: History of Libraries in the Western World. Third edition. Elmer D. Johnson and Michael H. Harris. Metuchen, NJ: Scarecrow Press, 354 pp. (1976). (ISBN 0-8108-0949-4; LC 76-25422). $1 0.00. Journal of the American Society for Information Science, 28(2):124–125, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1977:BRBb

[Har77b] John F. Harvey. Book review: National and International Library Planning, Key Papere Resented at the 40th Session of the IFLA General Council, Washington, D.C., 1974. Robert Vosper and Leone I. Newkirk, Eds. Munich: Verlag Dokumentation, 162 pp.

(1976). (IFLA Publication No. 4; ISBN 3-7940-44 24-X). Journal of the American Society for Information Science, 28(2):125, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harwey:1977:BRB

[Har77c] John F. Harwey. Book review: Farewell to Alexandria: Solutions to Space, Growth and Performance Problems of Libraries. Gore, Daniel, ed. Westport, CT: Greenwood Press, 184 pp., illus. (1976). (LC: 75-5345 ISBN: 0-837-1-8587-4). Journal of the American Society for Information Science, 28(1):62–64, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1978:BRB

[Har78a] Stephen P. Harter. Book review: SPSS as a library research tool. Marchant, Maurice P.; Smith, Nathan M.; Stirling, Keigh H. (Occasional Research Paper, No. 1.) Provo, UT: School of Library and Information Sciences; 9177: 44 pp. Price: $1.50. Journal of the American Society for Information Science, 29(5): 263, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harvey:1978:BRB

[Har78b] John F. Harvey. Book review: The abstract journal, 1790–1920: Origin, Development and Diffusion. Manzer, Bruce M. Metuchen, NJ, Scarecrow Press; 1977: 321 pp. Price $12.50. Journal of the American Society for Information Science, 29(4):213, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1979:BRB

[Har79] Stephen P. Harter. Book review: Evaluating Information: A Guide for Users of Social Science Research. Jeffry Katzer, Kenneth H. Cook, and Wayne W. Crouch. Reading, MA: Addison-Wesley; 1978: 191 pp. Journal of the American Society for Information Science, 30(2):119, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harrington:1983:BRB

[Har83] Jan L. Harrington. Book review: Microprocessors and microcomputers: Hardware and software, 2nd ed. Ronald J. Tocci and Lester P. Laskowski. Englewood Cliffs, NJ; Prentice-Hall; 1982:

404 pp. Journal of the American Society for Information Science, 34(2):158, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1984:SIM

[Har84] Stephen P. Harter. Scientific inquiry: a model for online searching. Journal of the American Society for Information Science, 35(2):110–117, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harmon:1987:BRB

[Har87] Glynn Harmon. Book review: Information specialist as team player in the research process. Journal of the American Society for Information Science, 38(4):316–317, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harmon:1988:BRB

[Har88] Glynn Harmon. Book review: Progress in communication sciences. Journal of the American Society for Information Science, 39(2):99–100, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1990:RST

[Har90] Stephen P. Harter. Research: Search term combinations and retrieval overlap: a proposed methodology and case study. Journal of the American Society for Information Science, 41(2):132–146, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harman:1991:RHE

[Har91] Donna Harman. Research: How effective is suffixing? Journal of the American Society for Information Science, 42(1):7–15, January 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harman:1992:PLH

[Har92a] Donna Harman. Perspectives on … human-computer interface: User-friendly systems instead of user-friendly front-ends. Journal of the American Society for Information Science, 43(2):164–174, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1992:RPR

[Har92b] Stephen P. Harter. Research: Psychological relevance and information science. Journal of the American Society for Information Science, 43(9):602–615, October 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harmon:1993:BRB

[Har93a] Glynn Harmon. Book review: Libraries and expert systems. Journal of the American Society for Information Science, 44(5):302– 303, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1993:LEM

[Har93b] Stephen P. Harter. Letters to the Editor: A model for word clustering. Journal of the American Society for Information Science, 44(4):213, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harris:1994:PLR

[Har94] Donald E. Harris. Perspectives on … the redesign/reengineering of an information services division in a major health science institution: Reassessing user needs. Journal of the American Society for Information Science, 45(5):331–334, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1996:VRA

[Har96] Stephen P. Harter. Variations in relevance assessments and the measurement of retrieval effectiveness. Journal of the American Society for Information Science, 47(1):37–49, January 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1998:SCE

[Har98] Stephen P. Harter. Scholarly communication and electronic journals: an impact study. Journal of the American Society for Information Science, 49(6):507–516, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haspers:1976:YFB

[Has76] Jan H. Haspers. The yield formula and Bradford’s Law. Journal of the American Society for Information Science, 27(5):281–287, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hashim:1985:NCL

[Has85] Elinor M. Hashim. National commission on libraries and information science: a brief overview. Journal of the American Society for Information Science, 36(6):360–363, November 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hastings:1995:BRBa

[Has95a] Samantha Kelly Hastings. Book review: Everybody’s guide to the Internet. Journal of the American Society for Information Science, 46(5):394, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hastings:1995:BRBb

[Has95b] Samantha Kelly Hastings. Book review: Global networks: Computers and international communication. Journal of the American Society for Information Science, 46(9):708–710, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haselkorn:1998:BRB

[Has98] Mark P. Haselkorn. Book review: Ergonomics and safety of intelligent driver interfaces. Journal of the American Society for Information Science, 49(4):384–385, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hawkins:1974:BDB

[Haw74] Donald T. Hawkins. Bibliographic data base usage in a large technical community. Journal of the American Society for Information Science, 25(2):105–108, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hawkins:1977:UUL

[Haw77] Donald T. Hawkins. Unconventional uses of on-line information retrieval systems: On-line bibliometric studies. Journal of the American Society for Information Science, 28(1):13–18, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hawkins:1981:MRO

[Haw81] Donald T. Hawkins. Machine-readable output from online searches. Journal of the American Society for Information Science, 32(4):253–256, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hawkins:1986:PLO

[Haw86] Donald T. Hawkins. Perspectives on … online searching: Online information retrieval reading list. Journal of the American Society for Information Science, 37(4):220–221, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hawkins:1987:LEO

[Haw87] Donald T. Hawkins. Letters to the Editor: Online retrieval. Journal of the American Society for Information Science, 38(4):307, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hawthorne:1990:BRB

[Haw90] Pat Hawthorne. Book review: Public access to government information: Issues, trends and strategies. Journal of the American Society for Information Science, 41(2):153, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hayes:1988:PLE

[Hay88] Robert M. Hayes. Perspectives on … education of the information professional: New dimensions, new directions: Part I. Curricula and accreditation. Education of the information professional: a library school perspective. Journal of the American Society for Information Science, 39(5):312–317, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hayes:1995:BRB

[Hay95] Robert M. Hayes. Book review: Systems analysis for librarians and information professionals. Journal of the American Society for Information Science, 46(5):393, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hayes:1996:MJB

[Hay96a] Robert M. Hayes. In memoriam: Joseph Becker, April 15, 1923– July 23, 1995. Journal of the American Society for Information Science, 47(6):412–414, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hayes:1996:PLC

[Hay96b] Robert M. Hayes. Perspectives on … cost and pricing of library and information services in transition: Cost of electronic reference resources and LCM: the library costing model. Journal of the American Society for Information Science, 47(3):228–234, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hazelton:1998:BRB

[Haz98] Penny A. Hazelton. Book review: Guide to finding legal and regulatory information on the Internet. Journal of the American Society for Information Science, 49(14):1332–1333, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hibner:1996:RIM

[HBG96] Dale Hibner, Richard Bellaver, and Frank M. Groom. Research: Interlinking E-mail systems. Journal of the American Society for Information Science, 47(8):618–623, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haythornthwaite:1999:MDD

[HBJR99] Caroline Haythornthwaite, Geoffrey Bowker, Christine Jenkins, and W. Boyd Rayward. Mapping the dimensions of a dynamic field. Journal of the American Society for Information Science, 50(12):1092–1094, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Horowitz:1984:ROF

[HC84] Irving Louis Horowitz and Mary E. Curtis. Review and opinion: Fair use versus fair return: Copyright legislation and its consequences. Journal of the American Society for Information Science, 35(2):67–74, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harman:1990:RRR

[HC90] Donna Harman and Gerald Candela. Research: Retrieving records from a gigabyte of text on a minicomputer using statistical ranking. Journal of the American Society for Information Science, 41(8):581–589, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hilton-Chalfen:1994:BRB

[HC94] Daniel Hilton-Chalfen. Book review: Adaptive technologies for learning and work environments. Journal of the American Society for Information Science, 45(8):565–566, September 1994. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1996:RCD

[HC96] Stephen P. Harter and Yung-Rang Cheng. Research: Colinked descriptors: Improving vocabulary selection for end-user searching. Journal of the American Society for Information Science, 47(4):311–325, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hill:2000:RAD

[HCL^{+}00] Linda L. Hill, Larry Carver, Mary Larsgaard, Ron Dolin, Terence R. Smith, James Frew, and Mary-Anna Rae. Research: Alexandria digital library: user evaluation studies and system design. Journal of the American Society for Information Science, 51(3):246–259, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hill:1972:LEB

[HD72] Barbara M. Hill and Anthony Debons. Letters to the Editor: Bibliographic coupling. Journal of the American Society for Information Science, 23(4):286, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hart:1988:RII

[HDRD88] Paul Hart, Ph. D., Ronald E. Rice, and Ph. D. Research: Interindustry relations in electronic news services. Journal of the American Society for Information Science, 39(4):252–261, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

He:1998:RCS

[He98] Shaoyi He. Research: Concept similarity and conceptual information alteration via English-to-Chinese and Chinese-to-English translation of medical article titles. Journal of the American Society for Information Science, 49(2):169–175, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

He:2000:TAC

[He00] Shaoyi He. Translingual alteration of conceptual information in medical translation: a crosslanguage analysis between English and Chinese. Journal of the American Society for Information Science, 51(11):1047–1060, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heine:1974:DER

[Hei74] M. H. Heine. Design equations for retrieval systems based on the Swets model. Journal of the American Society for Information Science, 25(3):183–198, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heilprin:1978:BRB

[Hei78] Laurence B. Heilprin. Book review: Perspectives in information science. Debons, Anthony; Cameron, William J., Eds. Leyden, The Netherlands: Noordhoff International Publishers; 1975: 797 pp., illustrations, index. Price: $65. (NATO Advanced Study Institute Series. Series E: Applied Sceince-No. 10.). Journal of the American Society for Information Science, 29(5):264–266, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heilprin:1980:LCT

[Hei80] Laurence B. Heilprin. The library community at a technological and philosophical crossroads: Necessary and sufficient conditions for survival. Journal of the American Society for Information Science, 31(6):389–395, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heim:1986:BRBb

[Hei86a] Kathleen M. Heim. Book review: Atypical careers and innovative services in library and information science. Journal of the American Society for Information Science, 37(1):49–50, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heim:1986:BRBa

[Hei86b] Kathleen M. Heim. Book review: Professional competencies — technology and the librarian. Journal of the American Society for Information Science, 37(1):49, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heim:1987:BRB

[Hei87] Kathleen M. Heim. Book review: Continuing education for the library information professions. Journal of the American Society for Information Science, 38(3):220–221, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heilprin:1988:HBR

[Hei88] Laurence B. Heilprin. Historical: Annual Review of Information Science and Technology (ARIST). Early historical perspective. Journal of the American Society for Information Science, 39(4): 273–280, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heim:1990:BRB

[Hei90] Kathleen M. Heim. Book review: If you want to evaluate your library. Journal of the American Society for Information Science, 41(1):72, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heilprin:1991:RLC

[Hei91] Laurence B. Heilprin. Research: The library community at a technological and philosophical crossroads: Necessary and sufficient conditions for survival. Journal of the American Society for Information Science, 42(8):566–573, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heinan:1994:BRB

[Hei94] Mary Ellen Heinan. Book review: Access versus assets: a comprehensive guide to resource sharing for academic librarians. Journal of the American Society for Information Science, 45(7):539, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heilprin:1995:RP

[Hei95a] John M. Heilprin. Research: Preface. Journal of the American Society for Information Science, 46(8):573, September 1995. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heilprin:1995:RIT

[Hei95b] Laurence B. Heilprin. Research: Information and thermodynamics: Toward a closer unification of information science with other sciences. Journal of the American Society for Information Science, 46(8):579–580, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heilprin:1995:RST

[Hei95c] Laurence B. Heilprin. Research: Science and technology: From prescientific times to the present. Journal of the American Society for Information Science, 46(8):574–578, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heim:1995:BRB

[Hei95d] Michael Heim. Book review: From writing to computers. Journal of the American Society for Information Science, 46(9):707– 708A, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Helmuth:1971:UEI

[Hel71] Nancy A. Helmuth. The use of extracts in information services. Journal of the American Society for Information Science, 22(6): 382–389, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henderson:1988:HTN

[Hen88] Madeline M. Henderson. Historical: Tracking nontraditional information systems. Journal of the American Society for Information Science, 39(4):270–272, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henry:1995:BRB

[Hen95] Charles Henry. Book review: Future libraries: Dreams, madness, and reality. Journal of the American Society for Information Science, 46(10):796–797, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henderson:1996:MCM

[Hen96] Madeline M. Henderson. In memoriam: Calvin N. Mooers, October 24, 1919–December 1, 1994. Journal of the American Society for Information Science, 47(9):659–661, September 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henderson:1997:LE

[Hen97] Albert Henderson. Letter to the Editor. Journal of the American Society for Information Science, 48(8):770, August 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henderson:1999:RIS

[Hen99] Albert Henderson. Research: Information science and information policy: the use of constant dollars and other indicators to manage research investments. Journal of the American Society for Information Science, 50(4):366–379, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henderson:2000:LEA

[Hen00a] Albert Henderson. Letter to the Editors: ARIST. Journal of the American Society for Information Science, 51(12):1154, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henderson:2000:RLL

[Hen00b] Albert Henderson. Research letter: Letter to the Editor. Journal of the American Society for Information Science, 51(5):495, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Herlach:1978:CRI

[Her78] Gertrud Herlach. Can retrieval of information from citation indexes be simplified? Multiple mention of a reference as a characteristic of the link between cited and citing article. Journal of the American Society for Information Science, 29(6):308–310, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Herner:1982:BRB

[Her82] Saul Herner. Book review: Information Retrieval Systems: Characteristics, Testing and Evaluation, 2nd Ed. F. Wilfrid Lancaster. New York: Wiley; 1979. Price: $19.75. Journal of the American Society for Information Science, 33(2):109–110, March 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Herner:1984:BHI

[Her84] Saul Herner. Brief history of information science. Journal of the American Society for Information Science, 35(3):157–163, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hersh:1994:RRR

[Her94] William Hersh. Research: Relevance and retrieval evaluation: Perspectives from medicine. Journal of the American Society for Information Science, 45(3):201–206, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hersh:1995:PLMb

[Her95] William R. Hersh. Perspectives on … medical informatics: Information technology in health care: The electronic medical record: Promises and problems. Journal of the American Society for Information Science, 46(10):772–776, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hersh:1996:LER

[Her96a] William Hersh. Letters to the Editor: Response to Dr. Humphrey’s letter. Journal of the American Society for Information Science, 47(5):408–409, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hert:1996:UGO

[Her96b] Carol A. Hert. User goals on an online public access catalog. Journal of the American Society for Information Science, 47(7): 504–518, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hert:1997:BRB

[Her97] Carol A. Hert. Book review: Design of library automation systems: File structure, data structures, and tools. Journal of the American Society for Information Science, 48(3):281–282, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Herring:1999:RVI

[Her99] Susan Davis Herring. Research: The value of interdisciplinarity: a study based on the design of Internet search engines. Journal of the American Society for Information Science, 50(4):358–365, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:2000:WBA

[HF00] Stephen P. Harter and Charlotte E. Ford. Web-based analyses of E-journal impact: Approaches, problems, and issues. Journal of the American Society for Information Science, 51(13):1159–1176, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haverkamp:1998:RII

[HG98] Donna S. Haverkamp and Susan Gauch. Research: Intelligent information agents: Review and challenges for distributed information sources. Journal of the American Society for Information Science, 49(4):304–311, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haas:2000:RRA

[HG00] Stephanie W. Haas and Erika S. Grams. Research: Readers, authors, and page structure: a discussion of four questions arising from a content analysis of web pages. Journal of the American Society for Information Science, 51(2):181–192, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hlnes:1970:BCP

[HH70] Theodore C. Hlnes and Jessica L. Harris. Brief communications: Permuted title indexes: Neglected considerations. Journal of the American Society for Information Science, 21(5):369–370, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1992:RIS

[HH92] Stephen P. Harter and Patricia A. Hooten. Research: Information science and scientists: JASIS, 1972–1990. Journal of the American Society for Information Science, 43(9):583–593, October 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hersh:1995:PLMa

[HH95a] William R. Hersh and David Hickam. Perspectives on … medical informatics: Information technology in health care: Information retrieval in medicine: The SAPHIRE experience. Journal of the American Society for Information Science, 46(10):743–747, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hersh:1995:EIB

[HH95b] William R. Hersh and David H. Hickam. An evaluation of interactive Boolean and natural language searching with an online medical textbook. Journal of the American Society for Information Science, 46(9):719, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hersh:1995:REI

[HH95c] William R. Hersh and David H. Hickam. Research: An evaluation of interactive Boolean and natural language searching with an online medical textbook. Journal of the American Society for Information Science, 46(7):478–489, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hendry:1997:IIS

[HH97] David G. Hendry and David J. Harper. An informal informationseeking environment. Journal of the American Society for Information Science, 48(11):1036–1048, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hines:1970:ECA

[HHC70] Theodore C. Hines, Jessica L. Harris, and Martin Colverd. Experimentation with computer-assisted indexing: American documentation, volume 20. Journal of the American Society for Information Science, 21(6):402–405, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Higgins:1999:RMI

[Hig99] Margaret Higgins. Research: Meta-information, and time: Factors in human decision making. Journal of the American Society for Information Science, 50(2):132–139, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hill:1996:BRBa

[Hil96a] Janette R. Hill. Book review: Information seeking in electronic environments. Journal of the American Society for Information Science, 47(5):405–406, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hill:1996:BRBb

[Hil96b] Janette R. Hill. Book review: Information seeking in electronic environments. Journal of the American Society for Information Science, 47(5):405–406, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hines:1970:BRB

[Hin70] Theodore C. Hines. Book review: 4/70-2R Theses and Dissertations; a Bibliographical Listing, Keyword Index, and Author Index: Cumulation, 1961–1967. 1970. Prepared by The Catholic University of America Libraries, Fred Blum, Editor. Catholic University of America Press, Washington, D.C. 548 pp. Journal of the American Society for Information Science, 21(4):305–306, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hirst:1978:DIF

[Hir78] Graeme Hirst. Discipline impact factors: a method for determining core journal lists. Journal of the American Society for Information Science, 29(4):171–172, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hirtle:1989:HHN

[Hir89] Peter B. Hirtle. Historical: Historical note: Atherton Seidell and the photoduplication of library material. Journal of the American Society for Information Science, 40(6):424–431, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hirtle:1990:LER

[Hir90] Peter B. Hirtle. Letters to the Editor: Regarding Allen Veaner’s letter. Journal of the American Society for Information Science, 41(6):427, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hirsh:1999:RCR

[Hir99] Sandra G. Hirsh. Research: Children’s relevance criteria and information seeking on electronic resources. Journal of the American Society for Information Science, 50(14):1265–1283, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hiltz:1989:RMA

[HJ89] Starr Roxanne Hiltz and Kenneth Johnson. Research: Measuring acceptance of computer-mediated communication systems. Journal of the American Society for Information Science, 40(6):386– 397, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hill:1999:RCM

[HJD^{+}99] Linda L. Hill, Greg Janée, Ron Dolin, James Frew, and Mary Larsgaard. Research: Collection metadata solutions for digital library applications. Journal of the American Society for Information Science, 50(13):1169–1181, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hert:2000:UAO

[HJD00] Carol A. Hert, Elin K. Jacob, and Patrick Dawson. A usability assessment of online indexing structures in the networked environment. Journal of the American Society for Information Science, 51(11):971–988, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hjorland:1996:LER

[Hjø96] Birger Hjørland. Letters to the Editor: Rejoinder: a new horizon for information science. Journal of the American Society for Information Science, 47(4):333–335, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hjorland:1999:BRB

[Hjø99] Birger Hjørland. Book review: The DDC, the universe of knowledge, and the post-modern library. Journal of the American Society for Information Science, 50(5):475–477, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hjorland:2000:BR

[Hjø00a] Birger Hjørland. Book reviews. Journal of the American Society for Information Science, 51(7):683–685, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hjorland:2000:LER

[Hjø00b] Birger Hjørland. Letter to the Editor: Relevance research: The missing perspective(s): “non-relevance” and “epistemological relevance”. Journal of the American Society for Information Science, 51(2):209–211, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Halpin:1979:APT

[HK79] Peter Halpin and John E. Knight. Air Pollution Technical Information Network: A Revised Approach. Journal of the American Society for Information Science, 30(5):315–316, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hentschke:1980:SAC

[HK80] Guilbert C. Hentschke and Ellen Kehoe. Serial acquisition as a capital budgeting problem. Journal of the American Society for Information Science, 31(5):357–362, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hmeidi:1997:RDI

[HKE97] Ismail Hmeidi, Ghassan Kanaan, and Martha Evens. Research: Design and implementation of automatic indexing for information retrieval with Arabic documents. Journal of the American Society for Information Science, 48(10):867–881, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Higgins:1970:SAF

[HM70] Timothy Higgins and Robert Mays. A solution to the ASIS file management system exercise using PL/ I and three general purpose systems. Journal of the American Society for Information Science, 21(3):209–213, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Humphrey:1987:RKB

[HM87] Susanne M. Humphrey and Nancy E. Miller. Research: Knowledge-based indexing of the medical literature: The Indexing Aid Project. Journal of the American Society for Information Science, 38(3):184–196, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Henrion:1986:REI

[HMNW86] Max Henrion, M. Granger Morgan, Indira Nair, and Charles Wiecha. Research: Evaluating an information system for policy modeling and uncertainty analysis. Journal of the American Society for Information Science, 37(5):319–330, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1997:BCI

[HN97] Stephen P. Harter and Thomas E. Nisonger. Brief communications: ISI’s impact factor as misnomer: a proposed new measure to assess journal impact. Journal of the American Society for Information Science, 48(12):1146–1148, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1998:LERa

[HN98a] Stephen P. Harter and Thomas E. Nisonger. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 49(6):578, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1998:LERb

[HN98b] Stephen P. Harter and Thomas E. Nisonger. Letters to the Editor: Rejoiner — The multiple meanings of impact factors. Journal of the American Society for Information Science, 49(9):864, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:1993:RSR

[HNW93] Stephen P. Harter, Thomas E. Nisonger, and Aiwei Weng. Research: Semantic relationships between cited and citing articles in library and information science journals. Journal of the American Society for Information Science, 44(9):543–552, October 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hjerppe:1989:RCE

[HO89] Roland Hjerppe and Birgitta Olander. Research: Cataloging and expert systems: AACR2 as a knowledge base. Journal of the American Society for Information Science, 40(1):27–44, January 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hockings:1974:BCS

[Hoc74] E. F. Hockings. Brief communications: Selection of scientific periodicals in an industrial research library. Journal of the American Society for Information Science, 25(2):131–132, March 1974. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hocker:1989:BRB

[Hoc89] Susan E. Hocker. Book review: Online database search services directory: a reference and referral guide to more than 1,700 libraries, information firms, and other sources providing computerized information retrieval and associated services using publicly available online databases. Journal of the American Society for Information Science, 40(6):436, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hoey:1972:SUH

[Hoe72] P. O’N. Hoey. Systematic utilization of human resources as an integral part of information science work. Journal of the American Society for Information Science, 23(6):384–391, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Holbrook:1970:TFR

[Hol70] Iola Belle Holbrook. A threaded-file retrieval system. Journal of the American Society for Information Science, 21(1):40–48, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Holm:1971:BCF

[Hol71] B. E. Holm. Brief communications: FID committee information for industry (FID/II). Journal of the American Society for Information Science, 22(6):409–410, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hollenberg:1986:RIP

[Hol86] Dennis Hollenberg. Research: Information processes. I. Hierarchical information systems. Journal of the American Society for Information Science, 37(5):300–306, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hollenberg:1987:LER

[Hol87] Dennis Hollenberg. Letters to the Editor: Rejoinder to Ralph Dumain’s letter. Journal of the American Society for Information Science, 38(4):310, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Holland:1996:ECU

[Hol96] Maurita Peterson Holland. Examples of classes using technological advances for media distribution and collaboration: Collaborative technologies in inter-university instruction. Journal of the American Society for Information Science, 47(11):857–862, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Holmes:1999:BRB

[Hol99] Boyd P. Holmes. Book review: Foundations of library and information science, by Richard E. Rubin. Journal of the American Society for Information Science, 50(9):838–839, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Home:1983:QGF

[Hom83] Esther E. Home. Question generation and formulation: an indication of information need. Journal of the American Society for Information Science, 34(1):5–15, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hooten:1991:RFF

[Hoo91] Patricia A. Hooten. Research: Frequency and functional use of cited documents in information science. Journal of the American Society for Information Science, 42(6):397–404, July 1991. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Horrocks:1978:LEL

[Hor78] Norman Horrocks. Letters to the Editor: Library Association membership. Journal of the American Society for Information Science, 29(3):165–166, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Horrocks:1993:LEG

[Hor93] Norman Horrocks. Letters to the Editor: Getty’s Syndrome^{T M} and its cousins. Journal of the American Society for Information Science, 44(4):213, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hoffman:1993:POD

[HOS^{+}93] Melia M. Hoffman, Lawrence O’Gorman, Guy A. Story, James Q. Arnold, and Nina H. Macdonald. Perspectives on … digital libraries: The RightPages^{T M} Service: an image-based electronic library. Journal of the American Society for Information Science, 44(8):446–452, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Howden:1987:BRB

[How87] Norman Howden. Book review: A micro handbook for small libraries and media centers, 2nd ed. Journal of the American Society for Information Science, 38(6):446, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Howden:1989:BRBb

[How89a] Norman Howden. Book review: Computer industry almanac. Journal of the American Society for Information Science, 40(6): 440, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Howden:1989:BRBc

[How89b] Norman Howden. Book review: Selecting and installing software packages. Journal of the American Society for Information Science, 40(6):440, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Howden:1989:BRBa

[How89c] Norman Howden. Book review: The politics of an emerging profession: The American Library Association 1876–1917. Journal of the American Society for Information Science, 40(6):439–440, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Howard:1994:RPR

[How94] Dara Lee Howard. Research: Pertinence as reflected in personal constructs. Journal of the American Society for Information Science, 45(3):172–185, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harter:2000:RIP

[HP00] Stephen P. Harter and Taemin Kim Park. Research: Impact of prior electronic publication on manuscript consideration policies of scholarly journals. Journal of the American Society for Information Science, 51(10):940–948, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hersh:1996:TOA

[HPH96] William Hersh, Jeffrey Pentecost, and David Hickam. A taskoriented approach to information retrieval evaluation. Journal of the American Society for Information Science, 47(1):50–56, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Herner:1978:AMC

[HS78] Saul Herner and Kurt J. Snapper. The application of multiplecriteria utility theory to the evaluation of information systems. Journal of the American Society for Information Science, 29(6): 289–296, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hunt:1975:MAI

[HSP75] Bernard L. Hunt, Martin Snyderman, and William Payne. Machine-assisted indexing of scientific research summaries. Journal of the American Society for Information Science, 26(4):230– 236, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haas:1996:BCT

[HST96] Stephanie W. Haas, Jeremy Sugarman, and Helen R. Tibbo. Brief communications: A text filter for the automatic identification of empirical articles. Journal of the American Society for Information Science, 47(2):167–169, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Heine:1991:RIO

[HT91] M. H. Heine and J. M. Tague. Research: An investigation of the optimization of search logic for the MEDLINE database. Journal of the American Society for Information Science, 42(4):267–278, May 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hubert:1976:NIB

[Hub76] John J. Hubert. On the Naranan interpretation of Bradford’s Law. Journal of the American Society for Information Science, 27(5):339–341, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hubert:1977:LEL

[Hub77] John J. Hubert. Letters to the Editor: Lotka’s Law in the humanities. Journal of the American Society for Information Science, 28(1):66, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hubert:1978:BCR

[Hub78] John J. Hubert. Brief communications: A relationship between two forms of Bradford’s Law. Journal of the American Society for Information Science, 29(3):159–161, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Huber:1998:RCA

[Hub98] John C. Huber. Research: Cumulative advantage and successbreeds-success: the value of time pattern analysis. Journal of the American Society for Information Science, 49(5):471–476, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hull:1996:SAC

[Hul96] David A. Hull. Stemming algorithms: a case study for detailed evaluation. Journal of the American Society for Information Science, 47(1):70–84, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Humphrey:1970:IPR

[Hum70] Hubert H. Humphrey. Information for peace remarks to the American Society for Information Science. Journal of the American Society for Information Science, 21(1):4–7, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Humphrey:1984:FMM

[Hum84] Susanne M. Humphrey. File maintenance of MeSH headings in MEDLINE. Journal of the American Society for Information Science, 35(1):34–44, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Humphrey:1996:LEI

[Hum96] Susanne M. Humphrey. Letters to the Editor: Information retrieval in medicine. Journal of the American Society for Information Science, 47(5):407–408, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Humphrey:1999:RAI

[Hum99] Susanne M. Humphrey. Research: Automatic indexing of documents from journal descriptors: a preliminary investigation. Journal of the American Society for Information Science, 50(8): 661–674, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Humphrey:2000:LEI

[Hum00] Susanne M. Humphrey. Letter to the Editor: Incremental benefit of human indexing. Journal of the American Society for Information Science, 51(6):582, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hunter:1972:LEP

[Hun72] Paul Hunter. Letters to the Editor: Patent literature project. Journal of the American Society for Information Science, 23(5): 335, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hunt:1996:BRBa

[Hun96a] Deborah Hunt. Book review: Finding government information on the Internet. Journal of the American Society for Information Science, 47(10):792–794A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hunt:1996:BRBb

[Hun96b] Deborah Hunt. Book review: Finding government information on the Internet. Journal of the American Society for Information Science, 47(10):792–794A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hurt:1980:BCE

[Hur80] C. D. Hurt. Brief communication: An examination of differences between two studies of highly cited old papers. Journal of the American Society for Information Science, 31(3):215–216, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hurd:1988:PLE

[Hur88] Julie M. Hurd. Perspectives on … education of the information professional: New dimensions, new direction: Technology: an agent for change in education for information science. Journal of the American Society for Information Science, 39(5):323–326, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hurd:2000:EIO

[Hur00a] Julie M. Hurd. Editorial: Introduction and overview. Journal of the American Society for Information Science, 51(14):1276–1278, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hurd:2000:TSC

[Hur00b] Julie M. Hurd. The transformation of scientific communication: a model for 2020. Journal of the American Society for Information Science, 51(14):1279–1283, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hogge:1971:BCI

[HW71] James H. Hogge and William E. Wright. Brief communications: Involving the User in Applications Software Development. Journal of the American Society for Information Science, 22(5):353– 354, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Harding:1980:BCI

[HW80] Alan F. Harding and Peter Willett. Brief communication: Indexing exhaustivity and the computation of similarity matrices. Journal of the American Society for Information Science, 31(4): 298–300, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haythornthwaite:1998:WFM

[HW98] Caroline Haythornthwaite and Barry Wellman. Work, friendship, and media use for information exchange in a networked organization. Journal of the American Society for Information Science, 49(12):1101–1114, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hedstrom:1999:RLS

[HW99] Margaret Hedstrom and David Wallace. Research: And the last shall be first: Recordkeeping policies and the NII. Journal of the American Society for Information Science, 50(4):331–339, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hsieh-Yee:1993:RES

[HY93] Ingrid Hsieh-Yee. Research: Effects of search experience and subject knowledge on the search tactics of novice and experienced searchers. Journal of the American Society for Information Science, 44(3):161–174, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Haiqi:1998:BCC

[HY98] Zhang Haiqi and Shigeaki Yamazaki. Brief communications: Citation indicators of Japanese journals. Journal of the American Society for Information Science, 49(4):375–379, ???? 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Hamill:1980:UTA

[HZ80] Karen A. Hamill and Antonio Zamora. The use of titles for automatic document classification. Journal of the American Society for Information Science, 31(6):396–402, November 1980. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ingwersen:1997:RDS

[IC97] Peter Ingwersen and Finn Hjortgaard Christensen. Research: Data set isolation for bibliometric online analyses of research publications: Fundamental methodological issues. Journal of the American Society for Information Science, 48(3):205–217, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ingwersen:1997:ERL

[Ing97] Peter Ingwersen. European research letters: Europe and information science. Journal of the American Society for Information Science, 48(12):1139–1141, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Inhaber:1975:CSJa

[Inh75a] H. Inhaber. Canadian scientific journals: Part 1, coverage. Journal of the American Society for Information Science, 26(4):253– 258, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Inhaber:1975:CSJb

[Inh75b] H. Inhaber. Canadian scientific journals: Part II, interaction. Journal of the American Society for Information Science, 26(5):

290–293, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Inhaber:1976:BRB

[Inh76] Herbert Inhaber. Book review: Proceedings of the second open conference on information science in Canada. Ottawa: Canadian Association for Information Science. 294 pp. (1974). Journal of the American Society for Information Science, 27(3):190, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Iivonen:1998:RTN

[IS98] Mirja Iivonen and Diane H. Sonnenwald. Research: From translation to navigation of different discourses: a model of search term selection during the pre-online stage of the search process. Journal of the American Society for Information Science, 49(4): 312–326, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Irish:1989:PLH

[IT89] Peggy M. Irish and Randall H. Trigg. Perspectives on … hypertext: Supporting collaboration in hypermedia: Issues and experience. Journal of the American Society for Information Science, 40(3):192–199, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1971:BCF

[Jac71] E. B. Jackson. Brief communications: FID at Buenos Aires, September 14 to 24, 1970. Journal of the American Society for Information Science, 22(1):64, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1972:BCR

[Jac72] Ecckne B. Jackson. Brief communications: Roman showcase for training in information work. Journal of the American Society for Information Science, 23(1):66–68, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1981:LEP

[Jac81] E. B. Jackson. Letters to the Editor: Perspectives on …. Journal of the American Society for Information Science, 32(5):336, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1991:HHN

[Jac91] Eugene B. Jackson. Historical: Historical note: Out of the cloudy crystal ball. Journal of the American Society for Information Science, 42(10):747–753, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jacobson:1994:PLE

[Jac94] Thomas L. Jacobson. Perspectives on … electronic publishing: The electronic publishing revolution is not “Global”. Journal of the American Society for Information Science, 45(10):745–752, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1996:BRB

[Jac96] J. R. Jackson. Book review: The Internet compendium: Subject guides to health and science resources. Journal of the American Society for Information Science, 47(12):953, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jaffe:1998:BRB

[Jaf98] Valerie Jaffe. Book review: The Internet searcher’s handbook: Locating information, people, and software. Journal of the American Society for Information Science, 49(6):567, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jahoda:1974:RQA

[Jah74] Gerald Jahoda. Reference question analysis and search strategy development by man and machine. Journal of the American Society for Information Science, 25(3):139–144, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jain:1972:SLB

[Jai72] A. K. Jain. Sampling in-library book use. Journal of the American Society for Information Science, 23(3):150–155, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jakobsson:1988:ROP

[Jak88] Matti Jakobsson. Research: One-pass text compression with a subword dictionary. Journal of the American Society for Information Science, 39(4):262–269, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Janes:1991:BCB

[Jan91] Joseph W. Janes. Brief communications: The binary nature of continuous relevance judgments: a study of users’ perceptions. Journal of the American Society for Information Science, 42(10): 754–756, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Janes:1994:ROP

[Jan94] Joseph W. Janes. Research: Other people’s judgments: a comparison of users’ and others’ judgments of document relevance, topicality, and utility. Journal of the American Society for Information Science, 45(3):160–171, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jarvelin:1989:RAQ

[Jär89] Kalervo Järvelin. Research: An approach to query cost modelling in numeric databases. Journal of the American Society for Information Science, 40(4):236–245, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jasperse:1994:PLE

[Jas94] Jaap A. Jasperse. Perspectives on … electronic publishing: Primary science on CD-ROM: The New Zealand experiment. Journal of the American Society for Information Science, 45(10):777– 784, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jasperse:1997:BRB

[Jas97] Jaap A. Jasperse. Book review: Scholarly publishing: the electronic frontier. Journal of the American Society for Information Science, 48(1):89–90, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Johnson:1971:HAI

[JB71a] Claire Johnson and Eleanor Briggs. Holography as applied to information storage and retrieval systems. Journal of the American Society for Information Science, 22(3):187–192, May 1971. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1971:WMA

[JB71b] K. Sparck Jones and E. O. Barber. What makes an automatic keyword classification effective? Journal of the American Society for Information Science, 22(3):166–175, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Juergensmeyer:1985:AID

[JB85] John E. Juergensmeyer and Sarah G. Bishop. Access to information: the dream and reality. Journal of the American Society for Information Science, 36(6):383–388, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jordan:1977:CRT

[JBH77] Sara R. Jordan, Antony F. R. Brown, and Fred C. Hutton. Computerized Russian translation at ORNL. Journal of the American Society for Information Science, 28(1):26–33, January 1977. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Julius:1977:VEW

[JBS^{+}77] Marc Julius, Charles E. Berkoff, Alvin E. Strack, Frank Krasovec, and A. Douglas Bender. A very early warning system for the rapid identification and transfer of new technology. Journal of the American Society for Information Science, 28(3):170–174, May 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jeroski:1978:BCA

[JD78] Sharon Jeroski and Donna Dartnell. Brief communications: The anti-verb list. Journal of the American Society for Information Science, 29(3):158, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jeng:1991:RKR

[Jen91] Ling Hwey Jeng. Research: Knowledge representation of the visual image of a title page. Journal of the American Society for Information Science, 42(2):99–109, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jeng:1995:BRBa

[Jen95a] Ling Hwey Jeng. Book review: Organizing knowledge. (2nd ed.). Journal of the American Society for Information Science, 46(5): 394–395, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jeng:1995:BRBb

[Jen95b] Ling Hwey Jeng. Book review: Using subject headings for online retrieval: Theory, practice and potential. Journal of the American Society for Information Science, 46(5):396, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jerman:1970:PDC

[Jer70] Max Jerman. Promising Developments in Computer-Assisted Instruction. Journal of the American Society for Information Science, 21(4):285–292, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1987:RPR

[JF87] William P. Jones and George W. Furnas. Research: Pictures of relevance: a geometric analysis of similarity measures. Journal of the American Society for Information Science, 38(6):420–442, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1990:RIS

[JGR90] Leslie P. Jones, Edward W. Gassie, Jr., and Sridhar Radhakrishnan. Research: INDEX: the statistical basis for an automatic conceptual phrase-indexing system. Journal of the American Society for Information Science, 41(2):87–97, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1995:RIT

[JGR^{+}95] Susan Jones, Mike Gatford, Steve Robertson, Micheline Hancock-Beaulieu, Judith Secker, and Steve Walker. Research: Interactive thesaurus navigation: Intelligence rules OK? Journal of the American Society for Information Science, 46(1):52–59, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1975:SGR

[JJ75] Eugene B. Jackson and Ruth L. Jackson. Speakers at the 1972 Gordon Research Conference: Their backgrounds and literature records as selection aids for “out-of-specialty” scientists and others. Journal of the American Society for Information Science, 26(3):166–173, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1977:ISL

[JJ77] Eugene B. Jackson and Ruth L. Jackson. The industrial special library universe — a “base line” study of its extent and characteristics. Journal of the American Society for Information Science, 28(3):135–152, May 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1980:CIS

[JJ80] Eugene B. Jackson and Ruth L. Jackson. Characterizing the industrial special library universe. Journal of the American Society for Information Science, 31(3):208–214, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jackson:1991:MSF

[JK91] Eugene B. Jackson and Donald H. Kraft. In memoriam: Stephen E. Furth, 1909–1991. Journal of the American Society for Information Science, 42(10):702, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

John:1997:EMA

[JM97] Bonnie E. John and Matthew M. Mashyna. Evaluating a multimedia authoring tool. Journal of the American Society for Information Science, 48(11):1004–1022, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jarvelin:1993:RDI

[JN93] Kalervo Järvelin and Timo Niemi. Research: Deductive information retrieval based on classifications. Journal of the American Society for Information Science, 44(10):557–578, December 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Johns:1984:ADM

[Joh84] Gerald Johns. Application: Dictionary making by conference and committee: NACA and the American Aeronautical Language, 1916–1934. Journal of the American Society for Information Science, 35(2):75–81, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Johnson:1986:BRB

[Joh86] Donald Clay Johnson. Book review: The library in the information revolution: Proceedings of the Sixth Congress of Southeast Asian Librarians, Singapore 30 May-3 June 1983. Journal of the American Society for Information Science, 37(5):349, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1971:CWP

[Jon71] Kevin P. Jones. Compound words: a problem in post-coordinate retrieval systems. Journal of the American Society for Information Science, 22(4):242–250, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1973:DIE

[Jon73a] Karen Sparck Jones. Does indexing exhaustivity matter? Journal of the American Society for Information Science, 24(5):313–316, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1973:LE

[Jon73b] Karen Sparck Jones. Letters to the Editors. Journal of the American Society for Information Science, 24(2):166–167, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1973:ECC

[Jon73c] Kevin P. Jones. The environment of classification: the concept of mutual exclusivity. Journal of the American Society for Information Science, 24(2):157–163, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1974:ECP

[Jon74] Kevin P. Jones. The environment of classification Part II: How we classify. Journal of the American Society for Information Science, 25(1):44–51, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1976:FSC

[Jon76] Warren T. Jones. A fuzzy set characterization of interaction in scientific research. Journal of the American Society for Information Science, 27(5):307–310, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1986:BRB

[Jon86] Bush Jones. Book review: Architecture of systems problem solving. Journal of the American Society for Information Science, 37(5):352, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1989:ERL

[Jon89] Richard M. Jones. European research letters: Online catalogue research in Europe. Journal of the American Society for Information Science, 40(3):153–157, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jones:1991:RRA

[Jon91] Karen Sparck Jones. Research: The role of artificial intelligence in information retrieval. Journal of the American Society for Information Science, 42(8):558–565, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jordan:1989:LEA

[Jor89] John S. Jordan. Letters to the Editor: Articles by Jung Soon Ro and Don Swanson. Journal of the American Society for Information Science, 40(5):362–363, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ji-ren:1988:RKT

[Jr88] Bao Ji-ren. Research: K, θ, M model: a bridge and a language. Journal of the American Society for Information Science, 39(1): 17–21, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Janes:1996:NIR

[JR96] Joseph W. Janes and Louis B. Rosenfeld. Networked information retrieval and organization: Issues and questions. Journal of the American Society for Information Science, 47(9):711–715, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Jimison:1995:PLM

[JS95] Holly Brügge Jimison and Paul Phillip Sher. Perspectives on … medical informatics: Information technology in health care: Consumer health informatics: Health information technology for consumers. Journal of the American Society for Information Science, 46(10):783–790, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Julien:1999:RBA

[Jul99] Heidi E. Julien. Research: Barriers to adolescents’ information seeking for career decision making. Journal of the American Society for Information Science, 50(1):38–48, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kantor:1976:LIU

[Kan76a] P. B. Kantor. The library as an information utility in the university context: Evolution and measurement of service. Journal of the American Society for Information Science, 27(2):100–112, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kantor:1976:AA

[Kan76b] Paul B. Kantor. Availability analysis. Journal of the American Society for Information Science, 27(5):311–319, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kantor:1978:BCN

[Kan78] Paul B. Kantor. Brief communications: A note on cumulative advantage distributions. Journal of the American Society for Information Science, 29(4):202–204, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See reply [de 78].

Kantor:1987:BCM

[Kan87] Paul B. Kantor. Brief communications: A model for the stopping behavior of users of online systems. Journal of the American Society for Information Science, 38(3):211–214, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kantor:1994:BRB

[Kan94] Paul B. Kantor. Book review: Operations research for libraries and information agencies. Journal of the American Society for Information Science, 45(2):118–119, March 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Karamuftuoglu:1998:CIR

[Kar98] Murat Karamuftuoglu. Collaborative information retrieval: Toward a social informatics view of IR interaction. Journal of the American Society for Information Science, 49(12):1070–1080, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kasabov:1996:BRBa

[Kas96a] Nikola Kasabov. Book review: Fril — fuzzy and evidential reasoning in artificial intelligence. Journal of the American Society for Information Science, 47(10):790–791, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kasabov:1996:BRBb

[Kas96b] Nikola Kasabov. Book review: Fril — fuzzy and evidential reasoning in artificial intelligence. Journal of the American Society for Information Science, 47(10):790–791A, October 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Katzer:1972:DSD

[Kat72] Jeffrey Katzer. The development of a semantic differential to assess users’ attitudes towards an on-line interactive reference retrieval system. Journal of the American Society for Information Science, 23(2):122–128, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Katopol:2000:BRB

[Kat00] Patricia F. Katopol. Book review: New organizational designs: Information aspects, by Bob Travica. Journal of the American Society for Information Science, 51(11):1061–1062, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kayany:1998:CUO

[Kay98] Joseph M. Kayany. Contexts of uninhibited online behavior: Flaming in social newsgroups on Usenet. Journal of the American Society for Information Science, 49(12):1135–1141, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kazlauskas:1975:BRBb

[Kaz75a] Edward John Kazlauskas. Book review: Managing education for results. Richard W. Hostrop, Homewood, Illinois: ETC Publications, 248 pp. (1975). Journal of the American Society for Information Science, 26(6):355–356, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kazlauskas:1975:BRBa

[Kaz75b] Edward John Kazlauskas. Book review: Staff development and continuing education programs for library personnel: Guidelines and Criteria. Barbara Conroy, Boulder, Colorado: Western Interstate Commission for Higher Education, 23 pp. (1974). Journal of the American Society for Information Science, 26(2):130, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kazlauskas:1977:BRB

[Kaz77] Edward John Kazlauskas. Book review: Effecting Change in Library Education. Allan F:. Hershfield. Curriculum Design for Library and Information Science. Robert S. Taylor, Ed. Syracuse, NY: Syracuse University School of Library Science, 87 pp. (1973). (Education and Curriculum Series No. 1). Journal of the American Society for Information Science, 28(2):124, March 1977. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kazlauskas:1979:BRB

[Kaz79] Edward John Kazlauskas. Book review: Education and Large Information Systems. R. A. Buckingham, Amsterdam: North-Holland; 1977: 197 pp. Price: $26.75. Journal of the American Society for Information Science, 30(1):61, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1978:EIR

[KB78] Donald H. Kraft and Abraham Bookstein. Evaluation of information retrieval systems: a decision theory approach. Journal of the American Society for Information Science, 29(1):31–40, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochan:1992:RPF

[KB92] Carol Ann Kochan and John M. Budd. Research: The persistence of fraud in the literature: The Darsee case. Journal of the American Society for Information Science, 43(7):488–493, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kambil:1998:PCI

[KB98a] Ajit Kambil and David Bodoff. Partial coordination. II. A preliminary evaluation and failure analysis. Journal of the American Society for Information Science, 49(14):1270–1282, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kaminer:1998:RBA

[KB98b] Noam Kaminer and Yale M. Braunstein. Research: Bibliometric analysis of the impact of Internet use on scholarly productivity. Journal of the American Society for Information Science, 49(8): 720–730, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1990:LEV

[KBN90] Manfred Kochen and Victor Janvier Bikai-Nyunai. Letters to the Editor: Vision and design in education for information professionals. Journal of the American Society for Information Science, 41 (5):390, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Katzer:1971:LE

[KBR^{+}71] Jeffrey Katzer, Charles L. Bernier, Dana L. Roth, T. Mark Hodges, Eloise C. Foster, and A. W. Calhoun. Letters to the Editors. Journal of the American Society for Information Science, 22(6):411–414, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kling:1999:RCD

[KC99] Rob Kling and Holly Crawford. From retrieval to communication: the development, use, and consequences of digital documentary systems. Journal of the American Society for Information Science, 50(12):1121–1122, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

King:1994:PSO

[KDJ94] D. King, J. Daroussin, and M. Jamagne. Proposal for a spatial organization model in soil science (the example of the European Communities Soil Map). Journal of the American Society for Information Science, 45(9):705–717, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kleyle:1990:RBF

[KdK90] R. Kleyle and A. de Korvin. Research: A belief function approach to information utilization in decision making. Journal of the American Society for Information Science, 41(8):569–580, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kazlauskas:1989:RDA

[KDS89] Edward John Kazlauskas, Michele F. DeYoe, and Katherine R. Smith. Research: A descriptive analysis of the characteristics of the microcomputer periodical literature. Journal of the American Society for Information Science, 40(4):262–268, July 1989. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kim:1999:REN

[KE99] Su Hee Kim and Caroline M. Eastman. Research: An experiment on node size in a hypermedia system. Journal of the American Society for Information Science, 50(6):530–536, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kearsley:1983:IV

[Kea83] Greg Kearsley. Instructional videodisc. Journal of the American Society for Information Science, 34(6):417–423, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kegan:1970:MUW

[Keg70] Daniel L. Kegan. Measures of the usefulness of written technical information to chemical researchers. Journal of the American Society for Information Science, 21(3):179–186, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Keith:1970:GEM

[Kei70] Nathan R. Keith, Jr. A General Evaluation Model for An Information Storage and Retrieval System. Journal of the American Society for Information Science, 21(4):237–239, July 1970. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kennickell:1987:PLF

[Ken87a] Ralph E. Kennickell, Jr. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: The U.S. Government Printing Office — Marketing and publishing. Journal of the American Society for Information Science, 38(1):68–70, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kent:1987:HNC

[Ken87b] Allen Kent. Historical notes: Cyril Cleverdon. Journal of the American Society for Information Science, 38(3):152–155, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kent:1987:HNP

[Ken87c] Allen Kent. Historical notes: Perspectives. Journal of the American Society for Information Science, 38(3):147–151, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kertesz:1974:HCF

[Ker74] François Kertesz. How to cope with the foreign-language problem: Experience gained at a multidisciplinary laboratory. Journal of the American Society for Information Science, 25(2):86–104, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Keren:1977:PSI

[Ker77] Carl Keren. A policy for science information activities in small countries: the case of Israel. Journal of the American Society for Information Science, 28(1):44–57, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Keren:1984:LEI

[Ker84] Carl Keren. Letters to the Editors: On information science. Journal of the American Society for Information Science, 35(2):137, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kostoff:1999:RHS

[KET99] R. N. Kostoff, Henry J. Eberhart, and Darrell Ray Toothman. Research: Hypersonic and supersonic flow roadmaps using bibliometrics and database tomography. Journal of the American Society for Information Science, 50(5):427–447, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Keyes:1996:BRB

[Key96] Alison M. Keyes. Book review: Beyond the information systems outsourcing bandwagon: the insourcing response. Journal of the American Society for Information Science, 47(9):720–722A, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Keyes:1997:BRB

[Key97] Alison M. Keyes. Book review: Information and business performance: a study of information systems and services in high performing companies. Journal of the American Society for Information Science, 48(8):763–764, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1992:BCQ

[KF92] Frederick G. Kilgour and Nancy L. Feder. Brief communications: Quotations referenced in scholarly monographs. Journal of the American Society for Information Science, 43(3):268–270, April 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krevitt:1972:BCC

[KG72] Beth Krevitt and Belver C. Griffith. Brief communications: A comparison of several Zipf-type distributions in their goodness of fit to language data. Journal of the American Society for Information Science, 23(3):220–221, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krauze:1971:CRG

[KH71] Tadeusz K. Krauze and Claude Hillinger. Citations, references and the growth of scientific literature: a model of dynamic interaction. Journal of the American Society for Information Science, 22(5):333–336, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kazlauskas:1980:AMT

[KH80] Edward John Kazlauskas and Thomas D. Holt. The application of a minicomputer to thesaurus construction. Journal of the American Society for Information Science, 31(5):363–368, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1995:BCL

[KH95] Michael Koenig and Toni Harrell. Brief communications: Lotka’s Law, Price’s urn, and electronic publishing. Journal of the American Society for Information Science, 46(5):386–388, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Khawam:1991:REG

[Kha91] Yves J. Khawam. Research: Epistemological grounds for cybernetic models. Journal of the American Society for Information Science, 42(5):372–377, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kidd:1990:RMR

[Kid90] J. S. Kidd. Research: Measuring referencing practices. Journal of the American Society for Information Science, 41(3):157–163, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kies:1998:BRB

[Kie98] Jonathan Kies. Book review: User and task analysis for interface design. Journal of the American Society for Information Science, 49(14):1334, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1987:HNP

[Kil87] Frederick G. Kilgour. Historical note: A personalized prehistory of OCLC. Journal of the American Society for Information Science, 38(5):381–384, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1993:BCLb

[Kil93a] Frederick G. Kilgour. Brief communications: Lack of indexes in works on information science. Journal of the American Society for Information Science, 44(6):364, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1993:BCLa

[Kil93b] Frederick G. Kilgour. Brief communications: Locating information in an Egyptian text of the 17th century B.C. Journal of the American Society for Information Science, 44(5):292–297, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1995:BCC

[Kil95a] Frederick G. Kilgour. Brief communications: Cataloging for a specific miniature catalog. Journal of the American Society for Information Science, 46(9):704–706, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1995:RES

[Kil95b] Frederick G. Kilgour. Research: Effectiveness of surname-title-words searches by scholars. Journal of the American Society for Information Science, 46(2):146–151, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1997:DPT

[Kil97] Frederick G. Kilgour. Development of particular techniques: Origins of coordinate searching. Journal of the American Society for Information Science, 48(4):340–348, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kim:1973:TFT

[Kim73] Chai Kim. Theoretical foundations of thesaurus-construction and some methodological considerations for thesaurus-updating. Journal of the American Society for Information Science, 24(2): 148–156, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kim:1982:RLS

[Kim82] Chai Kim. Retrieval language of social sciences and natural sciences: a statistical investigation. Journal of the American Society for Information Science, 33(1):1–7, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kim:2000:RMH

[Kim00] Hak Joon Kim. Research: Motivations for hyperlinking in scholarly electronic articles: a qualitative study. Journal of the American Society for Information Science, 51(10):887–899, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

King:1976:CRA

[Kin76] Rosemary King. A comparison of the readability of abstracts with their source documents. Journal of the American Society for Information Science, 27(2):118–121, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

King:1982:MSI

[Kin82] Donald W. King. Marketing secondary information products and services. Journal of the American Society for Information Science, 33(3):168–174, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

King:1983:OPF

[Kin83] Donald W. King. Obituary: Professor Fritz Machlup (1902– 1983). Journal of the American Society for Information Science, 34(4):243, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kinnucan:1992:RSR

[Kin92] Mark T. Kinnucan. Research: The size of retrieval sets. Journal of the American Society for Information Science, 43(1):72–79, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

King:1998:PLI

[Kin98a] Donald W. King. Perspective on … Internet issues/publishing issues: Some economic aspects of the Internet. Journal of the American Society for Information Science, 49(11):990–1002, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kingma:1998:BRB

[Kin98b] Bruce R. Kingma. Book review: Costing and pricing in the digital age: a practical guide for information services. Journal of the American Society for Information Science, 49(7):668, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kirson:1973:BCS

[Kir73] Benjamin L. Kirson. Brief communications: Scientific and technological information systems in the Soviet Union. Journal of the American Society for Information Science, 24(4):306–307, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kirtland:1981:MMC

[Kir81] Monika Kirtland. Macro- and microthesauri: Changes occurring in MeSH-derived thesauri and a solution to some related search handicaps. Journal of the American Society for Information Science, 32(4):249–252, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kiresuk:1993:POK

[Kir93] Thomas J. Kiresuk. Perspectives on … knowledge utilization: The evaluation of knowledge utilization: Placebo and nonspecific effects, dynamical systems, and chaos theory. Journal of the American Society for Information Science, 44(4):235–241, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kjell:1974:BCP

[Kje74] Bradley Kjell. Brief communications: Performance of Kilgour’s truncation algorithm in files of different subjects. Journal of the American Society for Information Science, 25(1):70–71, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kappelman:1999:RGR

[KJR99] Leon A. Kappelman, Jerry L. Johnson, and Kathy Rosmond. Research: Government’s role in reducing “Year-2000” risks. Journal of the American Society for Information Science, 50(4):354–357, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kokol:1996:BCL

[KK96] Peter Kokol and Tatjana Kokol. Brief communications: Linguistic laws and computer programs. Journal of the American Society for Information Science, 47(10):781–785, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1979:PBS

[KL79] Donald H. Kraft and James W. Liesener. Planning and budgeting for school media programs at the building, district, and regional levels: O.R. in the Little Red Schoolhouse. Journal of the American Society for Information Science, 30(1):41–50, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kantor:1998:RTM

[KL98a] Paul B. Kantor and Jung Jin Lee. Research: Testing the maximum entropy principle for information retrieval. Journal of the American Society for Information Science, 49(6):557–566, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Khan:1998:RST

[KL98b] Kushal Khan and Craig Locatis. Research: Searching through cyberspace: the effects of link display and link density on information retrieval from hypertext on the World Wide Web. Journal of the American Society for Information Science, 49(2):176–182, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Khan:1998:STC

[KL98c] Kushal Khan and Craig Locatis. Searching through cyberspace: the effects of link cues and correspondence on information retrieval from hypertext on the World Wide Web. Journal of the American Society for Information Science, 49(14):1248–1253, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Klopfenstein:1989:RFC

[Klo89] Bruce C. Klopfenstein. Research: Forecasting consumer adoption of information technology and services — Lessons from home video forecasting. Journal of the American Society for Information Science, 40(1):17–26, January 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Katzer:1973:BCD

[KM73] Jeffrey Katzer and Patricia Moell. Brief communications: On the dimensionality of users’ attitudes toward on-line retrieval systems — a replication? Journal of the American Society for Information Science, 24(4):307–308, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1978:LEW

[KM78] Manfred Kochen and Charles T. Meadow. Letters to the Editor: Writing quality and peer review. Journal of the American Society for Information Science, 29(5):260, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Klobas:1995:RIT

[KM95] Jane E. Klobas and Tanya McGill. Research: Identification of technological gatekeepers in the information technology profession. Journal of the American Society for Information Science, 46(8):581–589, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kling:1999:RSC

[KM99] Rob Kling and Geoffrey McKim. Research: Scholarly communication and the continuum of electronic publishing. Journal of the American Society for Information Science, 50(10):890–906, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:2000:RSP

[KM00a] Frederick G. Kilgour and Barbara B. Moran. Research: Surname plus recallable title word searches for known items by scholars. Journal of the American Society for Information Science, 51(1): 83–89, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kling:2000:JMT

[KM00b] Rob Kling and Geoffrey McKim. Not just a matter of time: Field differences and the shaping of electronic media in supporting scientific communication. Journal of the American Society for Information Science, 51(14):1306–1320, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kilgour:1999:RRE

[KMB99] Frederick G. Kilgour, Barbara B. Moran, and John R. Barden. Research: Retrieval effectiveness of surname-title-word searches for known items by academic library users. Journal of the American Society for Information Science, 50(3):265–270, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kahle:1993:POD

[KMG^{+}93] Brewster Kahle, Harry Morris, Jonathan Goldman, Thomas Erickson, and John Curran. Perspectives on … digital libraries: Interfaces for distributed systems of information servers. Journal of the American Society for Information Science, 44(8):453– 467, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kaplan:2000:DPI

[KN00] Nancy R. Kaplan and Michael L. Nelson. Determining the publication impact of a digital library. Journal of the American Society for Information Science, 51(4):324–339, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Knox:1977:LEI

[Kno77] William T. Knox. Letters to the Editor: Information science and structure. Journal of the American Society for Information Science, 28(2):129–132, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1972:BCE

[Koc72] M. Kochen. Brief communications: On the economics of information. Journal of the American Society for Information Science, 23(4):281–283, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1979:BRB

[Koc79] Manfred Kochen. Book review: Information Systems and Networks. K. Samuelson, H. Borko, and G. X. Amey. New York: Elsevier North-Holland; 1977. Journal of the American Society for Information Science, 30(4):238, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1980:LEC

[Koc80] Manfred Kochen. Letters to the Editor: Cybics. Journal of the American Society for Information Science, 31(4):302, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1981:TCF

[Koc81] Manfred Kochen. Technology and communication in the future. Journal of the American Society for Information Science, 32(2): 148–157, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochtanek:1983:BCD

[Koc83] Thomas R. Kochtanek. Brief communication: Document clustering, using macro retrieval techniques. Journal of the American Society for Information Science, 34(5):356–359, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1984:ISR

[Koc84a] Manfred Kochen. Information science research: the search for the nature of information. Journal of the American Society for Information Science, 35(3):194–199, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1984:LEI

[Koc84b] Manfred Kochen. Letter to the Editor: Ithiel de Sola Pool: In memoriam. Journal of the American Society for Information Science, 35(4):253, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1984:TPI

[Koc84c] Manfred Kochen. Toward a paradigm for information science: the influence of Derek de Solla Price. Journal of the American Society for Information Science, 35(3):147–148, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1987:OPE

[Koc87] Manfred Kochen. Opinion papers: Ethics and information science. Journal of the American Society for Information Science, 38(3):206–210, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1975:BRB

[Koe75] Mike Koenig. Book review: Review of information retrieval and documentation in chemistry. Charles H. Davis and James E. Rush. Westport, Connecticut: Greenwood Press. Journal of the American Society for Information Science, 26(2):129–130, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1979:LSC

[Koe79] Michael E. D. Koenig. On-line serials collection analysis. Journal of the American Society for Information Science, 30(3):148–153, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1981:LEK

[Koe81] M. E. D. Koenig. Letters to the Editor: Kroeber’s “superorganics”. Journal of the American Society for Information Science, 32 (5):335–336, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1983:BIV

[Koe83] Michael E. D. Koenig. Bibliometric indicators versus expert opinion in assessing research performance. Journal of the American Society for Information Science, 34(2):136–145, March 1983. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1988:LEE

[Koe88] Michael E. Koenig. Letters to the Editor: Electronics, the cost disease, and the operation of libraries. Journal of the American Society for Information Science, 39(1):28, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1992:LEB

[Koe92a] Michael E. D. Koenig. Letters to the Editor: Bush’s contribution to the vision of electronic information systems. Journal of the American Society for Information Science, 43(10):676–677, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1992:LEE

[Koe92b] Michael E. D. Koenig. Letters to the Editor: Erratum in “Entering Stage III. The Convergence of the Stage Hypotheses”. Journal of the American Society for Information Science, 43(7):518, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Koe92c].

Koenig:1992:RES

[Koe92c] Michael E. D. Koenig. Research: Entering state III. The convergence of the stage hypotheses. Journal of the American Society for Information Science, 43(3):204–209, April 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See erratum [Koe92b].

Koenig:1993:BCE

[Koe93] Michael E. D. Koenig. Brief communications: On the existence of Mikhailov. Journal of the American Society for Information Science, 44(10):588–589, December 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See response [Koe94].

Koenig:1994:LEM

[Koe94] Michael E. D. Koenig. Letters to the Editor: A. I. Mikhailov — a response. Journal of the American Society for Information Science, 45(7):540, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Koe93].

Koehler:1999:RAW

[Koe99] Wallace Koehler. Research: An analysis of web page and web site constancy and permanence. Journal of the American Society for Information Science, 50(2):162–180, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kohut:1974:CAO

[Koh74] Joseph J. Kohut. A comparative analysis of obsolescence patterns of the U.S. Geoscience literature. Journal of the American Society for Information Science, 25(4):242–251, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koller:1974:LES

[Kol74] Herbert R. Koller. Letters to the Editors: September–October, 1973 JASIS contents. Journal of the American Society for Information Science, 25(2):134, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koller:1975:BRB

[Kol75] Herbert R. Koller. Book review: Computers and early books, anonymous. London: Mansell, 131 pp. (1974). Journal of the American Society for Information Science, 26(5):308–309, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Korfhage:1974:ICS

[Kor74] Robert R. Korfhage. Informal communication of scientific information. Journal of the American Society for Information Science, 25(1):25–32, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kostoff:1994:RFR

[Kos94] Ronald N. Kostoff. Research: Federal research impact assessment: State-of-the-art. Journal of the American Society for Information Science, 45(6):428–440, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kokol:1999:RCN

[KPZ^{+}99] Peter Kokol, Vili Podgorelec, Milan Zorman, Tatjana Kokol, and Tatjana Njivar. Research: Computer and natural language texts — a comparison based on long-range correlations. Journal of the American Society for Information Science, 50(14):1295–1301, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kang:1980:AFD

[KR80] Jong H. Kang and William B. Rouse. Approaches to forecasting demands for library network services. Journal of the American Society for Information Science, 31(4):256–263, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1973:DTV

[Kra73] Donald H. Kraft. A decision theory view of the information retrieval situation: an operations research approach. Journal of the American Society for Information Science, 24(5):368–376, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1974:LED

[Kra74a] Donald H. Kraft. Letters to the Editors: A Decision Theory View of the Information Retrieval Situation. Journal of the American Society for Information Science, 25(3):208, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kramer:1974:LEN

[Kra74b] Arthur D. Kramer. Letters to the Editors: NAPS Procedure Change. Journal of the American Society for Information Science, 25(2):134, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1978:TRA

[Kra78] Donald H. Kraft. A threshold rule applied to the retrieval decision model. Journal of the American Society for Information Science, 29(2):77–80, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1979:BRB

[Kra79] Donald H. Kraft. Book review: The CODASYL Approach to Data Base Management. T. William Olle. Chichester, England: Wiley-Interscience; 1978: 287 pp. Price: $21.50. Journal of the American Society for Information Science, 30(4):239–240, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1985:LEE

[Kra85] Donald H. Kraft. Letters to the Editors: Editor’s reply. Journal of the American Society for Information Science, 36(5):354, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1986:EN

[Kra86a] Donald Kraft. Editor’s notes. Journal of the American Society for Information Science, 37(2):96, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1986:EES

[Kra86b] Donald H. Kraft. Editorial: An editorial statement: In search of good papers and scholarship. Journal of the American Society for Information Science, 37(2):55–56, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1987:EHA

[Kra87a] Donald H. Kraft. Editorial: Happy anniversary to the American Society for Information Science (ASIS). Journal of the American Society for Information Science, 38(5):319–320, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1987:EHN

[Kra87b] Donald H. Kraft. Editorial: Historical notes. Journal of the American Society for Information Science, 38(3):145–146, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1987:EPR

[Kra87c] Donald H. Kraft. Editorial: The peer review process for the Journal of the American Society for Information Science (JASIS). Journal of the American Society for Information Science, 38(2): 81–82, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1988:ENC

[Kra88a] Donald H. Kraft. Editorial: A new citation format for JASIS. Journal of the American Society for Information Science, 39(1): 1, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1988:EHB

[Kra88b] Donald H. Kraft. Editorial: Honoring the JASIS referees over the last few years. Journal of the American Society for Information Science, 39(4):223, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1988:E

[Kra88c] Donald H. Kraft. Errata. Journal of the American Society for Information Science, 39(4):287, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1989:EHA

[Kra89a] Donald H. Kraft. Editorial: Honoring additional JASIS referees. Journal of the American Society for Information Science, 40(2): 67, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1989:ERB

[Kra89b] Donald H. Kraft. Editorial: Reducing the backlog in JASIS. Journal of the American Society for Information Science, 40(5): 301–303, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1990:ERB

[Kra90] Donald H. Kraft. Editorial: Reducing the backlog in JASIS even further. Journal of the American Society for Information Science, 41(4):235–237, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1991:EAI

[Kra91a] Donald H. Kraft. Editorial: An alert to the information and library science community: Arise, mes enfants! Aux armes, citoyens! (to the barricades, comrades). Journal of the American Society for Information Science, 42(8):545, September 1991. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1991:EHD

[Kra91b] Donald H. Kraft. Editorial: Happy Days Are Here Again. Journal of the American Society for Information Science, 42(10):701, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1991:ENC

[Kra91c] Donald H. Kraft. Editorial notes: A clarification of the policy on printing letters to the Editor. Journal of the American Society for Information Science, 42(2):77, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1991:E

[Kra91d] Donald H. Kraft. Erratum. Journal of the American Society for Information Science, 42(2):115, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1992:ECA

[Kra92a] Donald H. Kraft. Editorial: A call for action in response to happy days. Journal of the American Society for Information Science, 43(3):203, April 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1992:EHB

[Kra92b] Donald H. Kraft. Editorial: Honoring the JASIS referees over the last seven years. Journal of the American Society for Information Science, 43(5):335–336, June 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1992:LER

[Kra92c] Donald H. Kraft. Letters to the Editor: Rejoinder to Dan Hanson’s letter. Journal of the American Society for Information Science, 43(2):151, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1993:END

[Kra93] Donald H. Kraft. Editorial: New designs. Journal of the American Society for Information Science, 44(3):123, April 1993. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1994:EES

[Kra94a] Donald H. Kraft. Editorial: Electronic submissions to JASIS are now acceptable. Journal of the American Society for Information Science, 45(10):725, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1994:EST

[Kra94b] Donald H. Kraft. Editorial: Special topics issues of JASIS. Journal of the American Society for Information Science, 45(3):123, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1997:EBW

[Kra97] Donald H. Kraft. Editorial: JASIS will be fourteen in 1998 (issues per annual volume, that is)! Journal of the American Society for Information Science, 48(12):1083–1084, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1999:MBC

[Kra99a] Donald Kraft. In memoriam: A brief commentary on Bob Korfhage. Journal of the American Society for Information Science, 50(4):290, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1999:EBS

[Kra99b] Donald H. Kraft. Editorial: JASIS standards. Journal of the American Society for Information Science, 50(6):483, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kraft:1999:ECM

[Kra99c] Donald H. Kraft. Editorials: The coming of the millennium and the Journal of the American Society for Information Science (JA-SIS). Journal of the American Society for Information Science, 50(1):1–3, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1981:IOB

[KRC81] Manfred Kochen, Victoria Reich, and Lee Cohen. Influence on online bibliographic services on student behavior. Journal of the American Society for Information Science, 32(6):412–420, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kovacs:1995:RSC

[KRD95] Diane K. Kovacs, Kara L. Robinson, and Jeanne Dixon. Research: Scholarly E-conferences on the academic networks: How library and information science professionals use them. Journal of the American Society for Information Science, 46(4):244–253, May 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kling:1998:SII

[KRH98] Rob Kling, Howard Rosenbaum, and Carol Hert. Social informatics in information science: an introduction. Journal of the American Society for Information Science, 49(12):1047–1052, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krisciunas:1977:LEL

[Kri77] Kevin Krisciunas. Letters to the Editor: Lotka’s Law — year by year. Journal of the American Society for Information Science, 28(1):65–66, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krieger:1986:BRBa

[Kri86a] Tillie Krieger. Book review: An introduction to online searching. Journal of the American Society for Information Science, 37(1): 47, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krieger:1986:BRBb

[Kri86b] Tillie Krieger. Book review: Computer-readable databases: a directory and data sourcebook. Journal of the American Society for Information Science, 37(3):174, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krikelas:1986:BRB

[Kri86c] James Krikelas. Book review: The user friendliness of the library catalog. Journal of the American Society for Information Science, 37(1):47–48, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krieger:1987:BRB

[Kri87] Tillie Krieger. Book review: Encyclopedia of information systems and services: 1985–86. 6th ed. International volume. Journal of the American Society for Information Science, 38(3):218, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Krikelas:1989:BRB

[Kri89] James Krikelas. Book review: Information seeking: Basing services on users’ behaviors. Journal of the American Society for Information Science, 40(6):439, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kochen:1970:ECV

[KS70] Manfred Kochen and A. Bertrand Segur. Effects of cataloging volume at the library of congress on the total cataloging costs of American research libraries. Journal of the American Society for Information Science, 21(2):133–139, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kim:1977:SCI

[KS77] Chai Kim and Eui Hang Shin. Sociodemographic correlates of intercounty variations in the public library output. Journal of the American Society for Information Science, 28(6):359–365, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koll:1990:RFV

[KS90] Mathew Koll and Padmini Srinivasan. Research: Fuzzy versus probabilistic models for user relevance judgments. Journal of the American Society for Information Science, 41(4):264–271, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Koenig:1997:RWS

[KS97] Michael Koenig and Patrizia Sione. Research: The world at stage III, but the Net at stage II. Journal of the American Society for Information Science, 48(9):853–859, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kopcsa:1998:STM

[KS98] Alexander Kopcsa and Edgar Schiebel. Science and technology mapping: a new iteration model for representing multidimensional relationships. Journal of the American Society for Information Science, 49(1):7–17, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kent:1990:RSF

[KSDR90] A. Kent, R. Sacks-Davis, and K. Ramamohanarao. Research: A signature file scheme based on multiple organizations for indexing very large text databases. Journal of the American Society for Information Science, 41(7):508–534, October 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kotamarti:1990:RAT

[KT90] Usha Kotamarti and Alan L. Tharp. Research: Accelerating text searching through signature trees. Journal of the American Society for Information Science, 41(2):79–86, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuch:1975:BRB

[Kuc75] T. D. C. Kuch. Book review: Improving access to library resources: The Influence of Organization of Library Collections, and of User Attitudes Toward Innovative Services. Richard M. Dougherty and Laura L. Blomquist. Metuchen, NJ: The Scarecrow Press xii + 180 pp. (1974). Journal of the American Society for Information Science, 26(1):66–67, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuch:1976:BRBb

[Kuc76a] T. D. C. Kuch. Book review: The sourcebook of library technology, a cumulative edition of library technology reports, 1965– 1975. 1976 ed. Chicago, IL: American Library Association. 17 pages + thirty 105MM microfiche, NMA format. (1976). Journal of the American Society for Information Science, 27(5):351, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuch:1976:BRBa

[Kuc76b] T. D. C. Kuch. Book review: Universal bibliographic control — a long term policy, a plan for action. Dorothy Anderson, Pullach bei Muchen: Verlag Dokumentation, 87 pp. (1974). Journal of the American Society for Information Science, 27(3):186–187, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuch:1978:BCR

[Kuc78] T. D. C. Kuch. Brief communications: Relation of title length to number of authors in journal articles. Journal of the American Society for Information Science, 29(4):200–202, July 1978. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuch:1980:LEB

[Kuc80] T. D. Kuch. Letters to the Editor: JASIS publication statistics. Journal of the American Society for Information Science, 31(4): 302–303, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [MZ79] and reply [MZ80].

Kuch:1982:TAI

[Kuc82] T. D. C. Kuch. Thematic analysis in information science: the example of “literature obsolescence”. Journal of the American Society for Information Science, 33(2):69–75, March 1982. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuehl:1972:MPE

[Kue72] Philip G. Kuehl. Marketing perspectives for “ERIC-like” information systems. Journal of the American Society for Information Science, 23(6):359–364, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kugel:1971:BCD

[Kug71] Peter Kugel. Brief communications: Dirty Boole? Journal of the American Society for Information Science, 22(4):293–294, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuhlthau:1991:RIS

[Kuh91] Carol C. Kuhlthau. Research: Inside the search process: Information seeking from the user’s perspective. Journal of the American Society for Information Science, 42(5):361–371, June 1991. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuhlthau:1999:RRE

[Kuh99] Carol Collier Kuhlthau. Resarch: The role of experience in the information search process of an early career information worker: Perceptions of uncertainty, complexity, construction, and sources. Journal of the American Society for Information Science, 50(5): 399–412, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuney:1982:BRB

[Kun82] Joseph H. Kuney. Book review: The Future of the Printed Word: The Impact and Implications of the New Communications Technology. Philip Hills, Ed. Westport, CT: Greenwood Press, 1981: 172 pp. Price $25.00. Journal of the American Society for Information Science, 33(6):371, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kuvshinoff:1973:LES

[Kuv73] Boris W. Kuvshinoff. Letters to the Editor: On self as subject. Journal of the American Society for Information Science, 24(4): 309, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kelly:1977:BCS

[KW77] Patrick Kelly and Francis W. Wolek. Brief communications: STI in the land of RaD. Journal of the American Society for Information Science, 28(2):121–123, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Klempner:1981:LEC

[KW81] Irving M. Klempner and Allan Wittman. Letters to the Editor: Copyright and information transfer. Journal of the American Society for Information Science, 32(2):159–160, March 1981. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kimber:1997:AHH

[KW97] W. Eliot Kimber and Julia A. Woods. Application of HyTime hyperlinks and finite coordinate spaces to historical writing, analysis, and presentation. Journal of the American Society for Information Science, 48(7):603–613, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kwok:1985:PTI

[Kwo85] K. L. Kwok. A probabilistic theory of indexing and similarity measure based on cited and citing documents. Journal of the American Society for Information Science, 36(5):342–351, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kwok:1999:REM

[Kwo99] K. L. Kwok. Research: Employing multiple representations for Chinese information retrieval. Journal of the American Society for Information Science, 50(8):709–723, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kies:1998:CCS

[KWR98] Jonathan K. Kies, Robert C. Williges, and Mary Beth Rosson. Coordinating computer-supported cooperative work: a review of research issues and strategies. Journal of the American Society for Information Science, 49(9):776–791, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Kerre:1986:BCU

[KZC86] Etienne E. Kerre, Rembrand B. R. C. Zenner, and Rita M. M. De Caluwe. Brief communications: The use of fuzzy set theory in information retrieval and databases: a survey. Journal of the American Society for Information Science, 37(5):341–345, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liebscher:1997:RFI

[LAD97] Peter Liebscher, Eileen G. Abels, and Daniel W. Denman. Research: Factors that influence the use of electronic networks by science and engineering faculty at small institutions. Part II. Preliminary use indicators. Journal of the American Society for Information Science, 48(6):496–507, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

LaFollette:2000:OFS

[LaF00] Marcel C. LaFollette. Observations on fraud and scientific integrity in a digital environment. Journal of the American Society for Information Science, 51(14):1334–1337, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lampson:1999:BRB

[Lam99] Marc Lampson. Book review: Technology and privacy: the new landscape, edited by Philip E. Agre and Marc Rotenberg. Journal of the American Society for Information Science, 50(7):631–633, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lampert:2000:BRB

[Lam00a] Lynn D. Lampert. Book review: Online retrieval: a dialogue of theory and practice, by Geraldine Walker and Joseph Janes. Journal of the American Society for Information Science, 51(12): 1152–1153, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lampson:2000:BR

[Lam00b] Marc Lampson. Book review. Journal of the American Society for Information Science, 51(2):206–208, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1971:CEA

[Lan71] F. W. Lancaster. The cost-effectiveness analysis of information retrieval and dissemination systems. Journal of the American Society for Information Science, 22(1):12–27, January 1971. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Landau:1973:BRB

[Lan73] Herbert B. Landau. Book review: The Evaluation of Information Services and Products. Donald W. King and Edward C. Bryant. Information Resources Press. Washington, D.C. 308 pp. Index. $15.00. Journal of the American Society for Information Science, 24(1):74–75, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Landau:1976:BRB

[Lan76] Herbert B. Landau. Book review: Guide to Reference Sources in the Computer Sciences, Ciel Carter. New York, Macmillan Information, 237 pp. (1974). Journal of the American Society for Information Science, 27(1):75, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1977:BRB

[Lan77a] F. W. Lancaster. Book review: Applications of Operations Research Models to Libraries: A Case Study of the Use of Monographs in the Francis A. Countway Library of Medicine, Harvard University. Chen, Ching-chih Cambridge, MA: The MIT Press (1976) (ISBN 0-262-03056-X). Journal of the American Society for Information Science, 28(1):61–62, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Landau:1977:BRB

[Lan77b] Herbert B. Landau. Book review: The Social Use of Information: Ownership and Access. Andrew E. Wessel. New York: Wiley-Interscience, 244 pp. (1976). Journal of the American Society for Information Science, 28(5):303–304, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1978:BRB

[Lan78] F. W. Lancaster. Book review: The subject approach to information. 3rd ed. A. C. Foskett. Hamden, CT: Linnet Books; 1977: 476 pp. ISBN: 0-208-01546-9. Price: $12.00. Journal of the American Society for Information Science, 29(6):319–320, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1980:BRB

[Lan80] F. W. Lancaster. Book review: Online Searching: The Impact on User Charges of the Extended Use of Online Information Services. T. P. Barwise. Paris: International Council of Scientific Unions Abstracting Board; 1979. Journal of the American Society for Information Science, 31(2):127–128, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1981:LEO

[Lan81a] F. W. Lancaster. Letters to the Editor: Online searching and paperless publication — Author’s reply. Journal of the American Society for Information Science, 32(3):236, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Rot81].

Landau:1981:BRB

[Lan81b] Herbert B. Landau. Book review: Thesauri and thesauri construction. Maxine Maccafferty London: ASLIB; 1977: 191 pp. (ASLIB Bibliography No. 7.). Journal of the American Society for Information Science, 32(3):237, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1982:BRB

[Lan82a] F. W. Lancaster. Book review: The Cashless Society: EFTs at the Crossroads. A. Bequai. New York: Wiley; 1981: ISBN 0–471– 05654–5. 298 pp. Price: $21.95. Journal of the American Society for Information Science, 33(3):142, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Langlois:1982:STM

[Lan82b] Richard N. Langlois. Systems theory and the meaning of information. Journal of the American Society for Information Science, 33 (6):395–399, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1987:BRBa

[Lan87a] F. W. Lancaster. Book review: Electronic publishing plus: Media for a technological future. Journal of the American Society for Information Science, 38(3):219, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1987:BRBb

[Lan87b] F. W. Lancaster. Book review: New directions in library and information science education. Journal of the American Society for Information Science, 38(4):314, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1989:BRB

[Lan89a] F. W. Lancaster. Book review: The electronic era of publishing: an overview of concepts, technologies and methods. Journal of the American Society for Information Science, 40(6):438, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1989:LEA

[Lan89b] F. W. Lancaster. Letters to the Editor: Article by Jung Soon Ro. Journal of the American Society for Information Science, 40(5): 362, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1990:BRB

[Lan90] F. W. Lancaster. Book review: India’s national library: Systematization and modernization. Journal of the American Society for Information Science, 41(2):149, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1993:BRB

[Lan93] F. W. Lancaster. Book review: Vocabulary control and search strategies in online searching. Journal of the American Society for Information Science, 44(2):121, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Langenberg:1994:PLR

[Lan94] Donald N. Langenberg. Perspectives on … the redesign/ reengineering of an information services division in a major health science institution: Information technology and the university: Integration strategies for the 21st century. Journal of the American Society for Information Science, 45(5):323–325, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1997:LED

[Lan97] F. W. Lancaster. Letters to the Editor: Dictionary of bibliometrics. Journal of the American Society for Information Science, 48(5):480, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1998:LERa

[Lan98a] F. W. Lancaster. Letters to the Editor: Retrieval experiments: Full text versus human indexing versus automatic indexing. Journal of the American Society for Information Science, 49(5):483– 484, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1998:LERb

[Lan98b] F. W. Lancaster. Letters to the Editor: Retrieval experiments: Full text versus human indexing versus automatic indexing. Journal of the American Society for Information Science, 49(5):484, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lapinski:1999:BRB

[Lap99] P. Scott Lapinski. Book review: Towards the Digital Library: The British Library’s Initiatives for Access Program. Journal of the American Society for Information Science, 50(10):956–957, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lapinski:2000:BRB

[Lap00] P. Scott Lapinski. Book review: Books, bytes, and bridges: Libraries and computer centers in academic institutions, edited by Larry Hardesty. Journal of the American Society for Information Science, 51(13):1248–1249, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Larson:1974:BRB

[Lar74] Signe Larson. Book review: A Guide to Sources of Consumer Information. Sarah M. Thomas and Bernadine Weddington. Washington: Information Resources Press, 1973. 177 pp. ISBN 0-87815-010-2 LC 73-77342. Journal of the American Society for Information Science, 25(4):273, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Larson:1990:BRB

[Lar90] Ray R. Larson. Book review: Hypertext hands-on!: an introduction to a new way of organizing and accessing information. Journal of the American Society for Information Science, 41(8): 612, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Larson:1991:RDS

[Lar91] Ray R. Larson. Research: The decline of subject searching: Long-term trends and patterns of index use in an online catalog. Journal of the American Society for Information Science, 42(3):197–215, April 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Larson:1992:LEE

[Lar92a] Ray R. Larson. Letters to the Editor: Erratum in “Experiments in Automatic Library of Congress Classification”. Journal of the American Society for Information Science, 43(7):518, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Lar92c].

Larson:1992:REAa

[Lar92b] Ray R. Larson. Research: Evaluation of advanced retrieval techniques in an experimental online catalog. Journal of the American Society for Information Science, 43(1):34–53, January 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Larson:1992:REAb

[Lar92c] Ray R. Larson. Research: Experiments in automatic Library of Congress classification. Journal of the American Society for Information Science, 43(2):130–148, March 1992. CODEN AISJB6.

ISSN 0002-8231 (print), 1097-4571 (electronic). See erratum [Lar92a].

Larsen:1999:RBB

[Lar99] Poul Steen Larsen. Research: Books and bytes: Preserving documents for posterity. Journal of the American Society for Information Science, 50(11):1020–1027, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lawton:1979:SEE

[LAT79] Stephen B. Lawton, Ethel Auster, and David To. A Systems Evaluation of the Educational Information System for Ontario. Journal of the American Society for Information Science, 30(1): 33–40, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Law:1974:CTC

[Law74] Gordon T. Law, Jr. Cold type composition: Its impact on library and information science. Journal of the American Society for Information Science, 25(5):319–326, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lawrence:1976:BCP

[Law76] Barbara Lawrence. Brief communication: Preparation of desktop literature indexes for ongoing research projects. Journal of the American Society for Information Science, 27(4):260–261, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lawani:1982:HCA

[Law82] Stephen M. Lawani. On the heterogeneity and classification of author self-citations. Journal of the American Society for Information Science, 33(5):281–284, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Layne:1994:PLI

[Lay94] Sara Shatford Layne. Perspectives on … indexing: Some issues in the indexing of images. Journal of the American Society for Information Science, 45(8):583–588, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lazonder:2000:LER

[Laz00] Ard W. Lazonder. Letter to the Editors: Rejoinder: Differences between novice and experienced users in searching information on the World Wide Web. Journal of the American Society for Information Science, 51(12):1155, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lawani:1983:VCC

[LB83] Stephen M. Lawani and Alan E. Bayer. Validity of citation criteria for assessing the influence of scientific publications: New evidence with peer assessment. Journal of the American Society for Information Science, 34(1):59–66, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1988:PLI

[LB88] Lois F. Lunin and Marion J. Ball. Perspectives on … integrated academic information management systems (IAIMS): Introduction and overview. Journal of the American Society for Information Science, 39(2):101–104, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1989:PLI

[LB89] Lois F. Lunin and Marion J. Ball. Perspectives on … information science and health informatics education: Introduction and overview. Journal of the American Society for Information Science, 40(5):364–366, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1994:PLR

[LB94] Lois F. Lunin and Marion J. Ball. Perspectives on … the redesign/reengineering of an information services division in a major health science institution: Introduction and overview. Journal of the American Society for Information Science, 45(5):319–322, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Large:2000:WCR

[LB00] Andrew Large and Jamshid Beheshti. The web as a classroom resource: Reactions from the users. Journal of the American Society for Information Science, 51(12):1069–1080, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Large:1994:RMC

[LBBR94] Andrew Large, Jamshid Beheshti, Alain Breuleux, and André Renaud. Research: Multimedia and comprehension: a cognitive study. Journal of the American Society for Information Science, 45(7):515–528, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Large:1995:RMC

[LBBR95] Andrew Large, Jamshid Beheshti, Alain Breuleux, and Andre Renaud. Research: Multimedia and comprehension: the relationship among text, animation, and captions. Journal of the American Society for Information Science, 46(5):340–347, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Large:1996:REA

[LBBR96] Andrew Large, Jamshid Beheshti, Alain Breuleux, and Andre Renaud. Research: Effect of animation in enhancing descriptive and procedural texts in a multimedia learning environment. Journal of the American Society for Information Science, 47(6):437–448, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lazinger:1997:RIU

[LBIP97] Susan S. Lazinger, Judit Bar-Ilan, and Bluma C. Peritz. Research: Internet use by faculty members in various disciplines: a comparative case study. Journal of the American Society for Information Science, 48(6):508–518, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liddy:1987:RSD

[LBKO87] Elizabeth Liddy, Susan Bonzi, Jeffrey Katzer, and Elizabeth Oddy. Research: A study of discourse anaphora in scientific abstracts. Journal of the American Society for Information Science, 38(4):255–261, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Larson:1976:CPB

[LBP76] Julian R. Larson, Gail A. Bernhard, and Mary Ellen Padin. Comparison of printed bibliographic descriptions distributed by BIO-SIS, CAS, and Ei. Journal of the American Society for Information Science, 27(1):46–52, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lazonder:2000:RDB

[LBW00] Ard W. Lazonder, Harm J. A. Biemans, and Iwan G. J. H. Wopereis. Research: Differences between novice and experienced users in searching information on the World Wide Web. Journal of the American Society for Information Science, 51(6):576–581, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leimkuhler:1971:AML

[LC71] Ferdinand F. Leimkuhler and Michael D. Cooper. Analytical models for library planning. Journal of the American Society for Information Science, 22(6):390–398, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1983:POO

[LC83] Lois F. Lunin and Elliot Cole. Perspectives on…: Office automation: Impact on organization, user, planner. Journal of the American Society for Information Science, 34(3):208–209, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Langley:1984:AML

[LC84] Pat Langley and Jaime G. Carbonell. Approaches to machine learning. Journal of the American Society for Information Science, 35(5):306–316, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1987:PLF

[LC87] Lois F. Lunin and Joseph P. Caponio. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Introduction and overview. Journal of the American Society for Information Science, 38(1):25–29, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1988:PLE

[LC88] Lois F. Lunin and Marianne Cooper. Perspectives on … education of the information professional: New dimensions, new directions: Introduction and overview. Journal of the American Society for Information Science, 39(5):307–311, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1990:PLE

[LC90] Lois F. Lunin and Susan Crawford. Perspectives on … evaluation of scientific information: Peer review and the impact of new information technology: Introduction and overview. Journal of the American Society for Information Science, 41(3):213–215, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1999:PLC

[LCB99] Lois F. Lunin, Kenneth D. Crews, and Dwayne K. Buttler. Perspective on … copyright and fair-use guidelines for education and libraries. Journal of the American Society for Information Science, 50(14):1303, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liew:1985:ICM

[LCL85] Chong K. Liew, Uinam J. Choi, and Chung J. Liew. Inference control mechanism for statistical database: Frequency-imposed data distortions. Journal of the American Society for Information Science, 36(5):322–329, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lucier:1985:OPC

[LD85] Richard E. Lucier and James F. Dooley. Opinion papers: Cosmology and the changing role of libraries: an analogy and reflections. Journal of the American Society for Information Science, 36(1): 44–47, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lindberg:1988:PLI

[LD88] Donald A. B. Lindberg and M. D. Perspectives on … integrated academic information management systems (IAIMS): Part I. The IAIMS initiatives of NLM. Institutional planning for advanced information services. Journal of the American Society for Information Science, 39(2):105–106, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1991:PLIa

[LD91] Lois F. Lunin and George D’Elia. Perspectives on … integrated information centers within academic environments: Introduction and overview. Journal of the American Society for Information Science, 42(2):116–119, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1997:LER

[LD97] Lois F. Lunin and George D’Elia. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 48(12):1153, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:2000:RVD

[LDS^{+}00] Yew-Huey Liu, Paul Dantzig, Martin Sachs, James T. Corey, Mark T. Hinnebusch, Marc Damashek, and Jonathan Cohen. Research: Visualizing document classification: a search aid for the digital library. Journal of the American Society for Information Science, 51(3):216–227, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1985:MPL

[LE85] Lois F. Lunin and Beth Krevitt Eres. Miscellaneous: Perspectives on … international information issues. Journal of the American Society for Information Science, 36(3):143, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leatherdale:1975:BRB

[Lea75] Donald Leatherdale. Book review: Indexing languages and thesauri: Construction and Maintenance. Dagobert Soergel, Melville Publishing Company, Los Angeles, California: X1iii 632 pp. (1974). Journal of the American Society for Information Science, 26(6):353, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leatherdale:1976:BRB

[Lea76] Donald Leatherdale. Book review: Population and family planning classification. Jacqueline P. Forget. London: International Planned Parenthood Federation. 106 pp. (1975). Journal of the American Society for Information Science, 27(3):187, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leach:1998:BRB

[Lea98] Michael R. Leach. Book review: Java in a nutshell: a desktop quick reference. Journal of the American Society for Information Science, 49(14):1329–1330, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lee:1996:BRBa

[Lee96a] Alfred T. Lee. Book review: Global perspectives on the ecology of human-machine systems. Journal of the American Society for Information Science, 47(6):479–480, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lee:1996:BRBb

[Lee96b] Alfred T. Lee. Book review: Local applications of the ecological approach to human-machine systems. Journal of the American Society for Information Science, 47(6):480, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lee:2000:WC

[Lee00a] Hur-Li Lee. What is a collection? Journal of the American Society for Information Science, 51(12):1106–1113, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lee:2000:RPT

[Lee00b] Wade M. Lee. Research: Publication trends of doctoral students in three fields from 1965–1995. Journal of the American Society for Information Science, 51(2):139–144, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lehnus:1972:WCW

[Leh72] Donald J. Lehnus. Who cited what? A citation analysis of the four basic cataloging texts. Journal of the American Society for Information Science, 23(2):100–108, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lehr:1992:PLI

[Leh92] William Lehr. Perspectives on … information technology standards: Standardization: Understanding the process. Journal of the American Society for Information Science, 43(8):550–555, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leimkuhler:1973:ADL

[Lei73] Ferdinand F. Leimkuhler. ASIS Distinguished Lecture — 1972: Operations research and information science — a common cause. Journal of the American Society for Information Science, 24(1): 3–8, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lenk:1983:MFB

[Len83] Peter Lenk. Mappings of fields based on nominations. Journal of the American Society for Information Science, 34(2):115–122, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lerner:1979:BCP

[Ler79] Alexander Lerner. Brief communications: To protest or to ignore? A scientist’s alternatives in reacting to the persecution of his colleagues. Journal of the American Society for Information Science, 30(6):353–355, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lesk:1989:BRB

[Les89] Michael Lesk. Book review: Text processing and document manipulation: Proceedings of the international conference, Nottingham, April 1986. Journal of the American Society for Information Science, 40(2):137–139, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lesk:1996:BRB

[Les96] Michael Lesk. Book review: The future of software. Journal of the American Society for Information Science, 47(4):326–327, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1973:BCM

[Lev73a] Emil H. Levine. Brief communications: MIS spinoff of a conventional STINFO retrieval system. Journal of the American Society for Information Science, 24(4):305, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1973:LEN

[Lev73b] Emil H. Levine. Letters to the Editor: National Security Council Directive Establishes Data Index System. Journal of the American Society for Information Science, 24(1):76, January 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1974:BCE

[Lev74] Emil H. Levine. Brief communications: Effect of instantaneous retrieval on indexing criteria. Journal of the American Society for Information Science, 25(3):199–207, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1976:BRBa

[Lev76a] Marilyn Levine. Book review: The measurement of verbal information in psychology and education. Klaus Weltner. New York, Heidelberg: Springer-Verlag. 185 pp. (1973) (Translated from the German by Barbara M. Crook) (ISBN 0-387-06335-8). Journal of the American Society for Information Science, 27(3):186, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1976:BRBb

[Lev76b] Marilyn M. Levine. Book review: Language-operational-gestalt awareness: a radically empirical and pragmatical phenomenology of the processes and systems of library experience. Eugene Edward Craziano. Tempe, AZ: The Association for Library Automation Research Communications, xviii +457 pp. (1975). Journal of the American Society for Information Science, 27(5):351– 352, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1977:IAA

[Lev77] Marilyn M. Levine. The informative act and its aftermath: Toward a predictive science of information. Journal of the American Society for Information Science, 28(2):101–106, March 1977. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levitan:1982:IRG

[Lev82] Karen B. Levitan. Information resources as “Goods” in the life cycle of information production. Journal of the American Society for Information Science, 33(1):44–54, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1998:LET

[Lev98] Emil Levine. Letters to the Editor: Total recall on relevance article? Journal of the American Society for Information Science, 49(5):482, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levering:1999:PLC

[Lev99] Mary Levering. Perspective on … copyright and fair-use guidelines for education and libraries: What’s right about fair-use guidelines for the academic community? Journal of the American Society for Information Science, 50(14):1313–1319, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lewis:2000:BR

[Lew00] Bryan Lewis. Book reviews. Journal of the American Society for Information Science, 51(5):490–491, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leydesdorff:1997:RWW

[Ley97] Loet Leydesdorff. Research: Why words and co-words cannot map the development of the sciences. Journal of the American Society for Information Science, 48(5):418–427, May 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leydesdorff:1998:LER

[Ley98] Loet Leydesdorff. Letters to the Editor: Reply about using cowords. Journal of the American Society for Information Science, 49(1):98–99, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1993:POD

[LF93] Lois F. Lunin and Edward A. Fox. Perspectives on … digital libraries: Introduction and overview. Journal of the American Society for Information Science, 44(8):440–445, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1994:PLI

[LF94] Lois F. Lunin and Raya Fidel. Perspectives on … indexing: Introduction and overview. Journal of the American Society for Information Science, 45(8):569–571, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lundquist:1999:RPR

[LFHG99] Carol Lundquist, Ophir Frieder, David O. Holmes, and David Grossman. Research: A parallel relational database management system approach to relevance feedback in information retrieval. Journal of the American Society for Information Science, 50(5): 413–426, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lefever:1972:MUV

[LFS72] Maureen Lefever, Barbara Freedman, and Louise Schultz. Managing an uncontrolled vocabulary ex post facto. Journal of the American Society for Information Science, 23(6):339–342, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ludwig:1975:FAD

[LG75] B. M. Ludwig and H. P. Glockmann. The formal analysis of document retrieval systems. Journal of the American Society for Information Science, 26(1):51–55, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lang:1996:RSM

[LG96] William Steve Lang and Mary Grigsby. Research: Statistics for measuring the efficiency of electronic information retrieval. Journal of the American Society for Information Science, 47(2):159– 166, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lindsay:1999:RLB

[LG99] Robert K. Lindsay and Michael D. Gordon. Research: Literaturebased discovery by lexical statistics. Journal of the American Society for Information Science, 50(7):574–587, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1976:MMD

[LH76] Lois F. Lunin and Miriam T. Heath. A multifaceted medical data information system and one product: the index-handbook of ototoxic agents. Journal of the American Society for Information Science, 27(3):153–161, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1984:IOI

[LH84] Lois F. Lunin and Robert M. Hayes. Introduction and overview: Introduction and overview. Journal of the American Society for Information Science, 35(3):156, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1986:PLO

[LH86] Lois F. Lunin and Donald T. Hawkins. Perspectives on … online searching: Introduction and overview. Journal of the American Society for Information Science, 37(4):179–182, July 1986. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1992:PLH

[LH92] Lois F. Lunin and Donna Harman. Perspectives on … humancomputer interface: Introduction and overview. Journal of the American Society for Information Science, 43(2):152–155, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1995:RST

[LH95a] Robert M. Losee and Stephanie W. Haas. Research: Sublanguage terms: Dictionaries, usage, and automatic classification. Journal of the American Society for Information Science, 46(7):519–529, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1995:PLM

[LH95b] Lois F. Lunin and William R. Hersh. Perspectives on … medical informatics: Information technology in health care: Introduction and overview. Journal of the American Society for Information Science, 46(10):725–728, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lund:1997:PLI

[LH97] Douglas Lund and Nancy Herther. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: The technological structure and services of the IIC. Journal of the American Society for Information Science, 48(5):446–448, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1986:LE

[LHM^{+}86] Marilyn M. Levine, C. D. Hurt, William E. McGrath, Jean E. Crampon, and Martha Cornog. Letters to the Editors. Journal of the American Society for Information Science, 37(5):354–356, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lidstone:1996:BRB

[Lid96] Herbert J. Lidstone. Book review: Computational organization theory. Journal of the American Society for Information Science, 47(4):329–330, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lievrouw:1994:STI

[Lie94] Leah A. Lievrouw. Special topic issue: Information resources and democracies: Information resources and democracy: Understanding the paradox. Journal of the American Society for Information Science, 45(6):350–357, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liebscher:1996:BRB

[Lie96a] Peter Liebscher. Book review: The myth of the electronic library. Journal of the American Society for Information Science, 47(2):179–181, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lievrouw:1996:BRB

[Lie96b] Leah A. Lievrouw. Book review: The information society: a study of continuity and change. Journal of the American Society for Information Science, 47(3):250–251, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lifshin:1993:RCA

[Lif93] Arthur Lifshin. Research: Citation analysis of Geochimica et Cosmochimica Acta, 1951–1960. Journal of the American Society for Information Science, 44(6):322–326, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lilley:1978:BRB

[Lil78] Dorothy B. Lilley. Book review: Trends in governance of libraries (library trends, Vol. 26, No. 2, Fall, 1977). Summers, William F. (Issue Editor). Urbana, IL: University of Illinois Library School; 1977: 298 pp. LC 54-62638; ISSN 0024-2394. Quarterly subscription price $15.00/year. Individual issues $4.00. Journal of the American Society for Information Science, 29(4):213–214, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lindsley:1970:AIT

[Lin70] Thomas F. Lindsley. Application of IBM TEXT-PAC to the ASIS file management exercise. Journal of the American Society for Information Science, 21(3):224–227, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Line:1977:LELa

[Lin77a] M. B. Line. Letters to the Editor: a library network model. Journal of the American Society for Information Science, 28(1): 65, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Line:1977:LELb

[Lin77b] M. B. Line. Letters to the Editor: A library network model. Journal of the American Society for Information Science, 28(6): 377, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lindquist:1978:GDI

[Lin78] Mats G. Lindquist. Growth dynamics of information search services. Journal of the American Society for Information Science, 29(2):67–76, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lindberg:1987:PLF

[Lin87a] Donald A. B. Lindberg. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Section II. Health sciences libraries. National Library of Medicine: the view at 150 years. Journal of the American Society for Information Science, 38(1):34–39, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Line:1987:LER

[Lin87b] Maurice B. Line. Letters to the Editor: Relationship between productivity & obsolescence. Journal of the American Society for Information Science, 38(4):307–308, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Linn:1996:IDL

[Lin96] Marcia C. Linn. Instances of distance learning: Cognition and distance learning. Journal of the American Society for Information Science, 47(11):826–842, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lin:1997:RMD

[Lin97] Xia Lin. Research: Map displays for information retrieval. Journal of the American Society for Information Science, 48(1):40–54, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lipetz:1983:LEC

[Lip83] Bemami Lipetz. Letters to the Editor: Citation relation evaluation clarification. Journal of the American Society for Information Science, 34(4):295–296, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lipetz:1999:RAB

[Lip99] Ben-Ami Lipetz. Research: Aspects of JASIS authorship through five decades. Journal of the American Society for Information Science, 50(11):994–1003, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Litzinger:1996:BRB

[Lit96] Mary Ellen Litzinger. Book review: Designing and writing online documentation: Hypermedia for self-supporting products. Journal of the American Society for Information Science, 47(6):478– 478B, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:1996:RDS

[Liu96] Ziming Liu. Research: Dissipative structure theory, synergetics, and their implications for the management of information systems. Journal of the American Society for Information Science, 47(2):129–135, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:1997:BRB

[Liu97a] Geoffrey Z. Liu. Book review: Symbolic projection for image information retrieval and spatial reasoning. Journal of the American Society for Information Science, 48(10):960–961, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:1997:RSV

[Liu97b] Geoffrey Z. Liu. Research: Semantic Vector Space Model: Implementation and evaluation. Journal of the American Society for Information Science, 48(5):395–417, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:1997:RCT

[Liu97c] Ziming Liu. Research: Citation theories in the framework of international flow of information: New evidence with translation analysis. Journal of the American Society for Information Science, 48(1):80–87, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:1998:BRB

[Liu98] Geoffrey Z. Liu. Book review: Automated information retrieval: Theory and methods. Journal of the American Society for Information Science, 49(10):953–955, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1970:BCI

[LJ70a] F. W. Lancaster and Grace T. Jenkins. Brief communications: On the importance of root-stem truncation in word-frequency analysis. Journal of the American Society for Information Science, 21(5):368–369, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1970:BCQ

[LJ70b] F. W. Lancaster and Grace T. Jenkins. Brief communications: “quality control” applied to the operations of a large-information system. Journal of the American Society for Information Science, 21(5):370–371, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lee:1991:RSP

[LK91] Jung Jin Lee and Paul B. Kantor. Research: A study of probabilistic information retrieval systems in the case of inconsistent expert judgments. Journal of the American Society for Information Science, 42(3):166–172, April 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leung:1997:RSL

[LK97] Chi-Hong Leung and Wing-Kay Kan. Research: A statistical learning approach to automatic indexing of controlled index terms. Journal of the American Society for Information Science, 48(1):55–66, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1985:BTA

[LL85] F. W. Lancaster and Ja-Lih Lee. Bibliometric techniques applied to issues management: a case study. Journal of the American Society for Information Science, 36(6):389–397, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1986:PLT

[LL86] Lois F. Lunin and Larry L. Learn. Perspectives on … telecommunications: Principles, developments, prospects: Introduction and overview. Journal of the American Society for Information Science, 37(6):401–404, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1991:PLIb

[LL91] Lois F. Lunin and Clifford A. Lynch. Perspectives on … imaging. Advanced applications: Introduction and overview. Journal of the American Society for Information Science, 42(8):575–577, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lin:1991:RGR

[LLM91] Xia Lin, Peter Liebscher, and Gary Marchionini. Research: Graphical representations of electronic search patterns. Journal of the American Society for Information Science, 42(7):469–478, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1978:ABP

[LM78] F. W. Lancaster and John Martyn. Assessing the benefits and promise of an international information program (AGRIS). Journal of the American Society for Information Science, 29(6):283– 288, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Longo:1981:CDA

[LM81] Rose Mary Juliano Longo and Ubaldino Dantas MacHado. Characterization of databases in the agricultural sciences. Journal of the American Society for Information Science, 32(2):83–91, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Landry:1973:OPD

[LMM^{+}73] B. C. Landry, B. A. Mathis, N. M. Meara, J. E. Rush, and C. E. Young. Opinion papers: Definition of some basic terms in computer and information science. Journal of the American Society for Information Science, 24(5):328–342, September 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Larson:1996:CID

[LMO^{+}96] Ray R. Larson, Jerome McDonough, Paul O’Leary, Lucy Kuntz, and Ralph Moon. Cheshire II: Designing a next-generation online catalog. Journal of the American Society for Information Science, 47(7):555–567, July 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Landau:1982:ITM

[LMSC82] Herbert B. Landau, Jerome T. Maddock, F. Floyd Shoemaker, and Joseph G. Costello. An information transfer model to define information users and outputs with specific application to environmental technology. Journal of the American Society for Information Science, 33(2):82–91, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1982:FIA

[LN82] F. W. Lancaster and Julie M. Neway. The future of indexing and abstracting services. Journal of the American Society for Information Science, 33(3):183–189, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lee:1999:RTS

[LNL99] Kin Hong Lee, Mau Kit Michael Ng, and Qin Lu. Research: Text segmentation for Chinese spell checking. Journal of the American Society for Information Science, 50(9):751–759, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Longo:1990:RIT

[Lon90] Rose Mary Juliano Longo. Research: Information transfer and the adoption of agricultural innovations. Journal of the American Society for Information Science, 41(1):1–9, January 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Longfellow:1992:LER

[Lon92] Charles A. Longfellow, Jr. Letters to the Editor: Regarding “Lotka’s Law and the Kolmogorov–Smirnov Test: An Error in Calculation”. Journal of the American Society for Information Science, 43(7):518, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Lou92a].

Losee:1987:RPR

[Los87] Robert Losee. Research: Probabilistic retrieval and coordination level matching. Journal of the American Society for Information Science, 38(4):239–244, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1988:RPE

[Los88] Robert M. Losee. Research: Parameter estimation for probabilistic document-retrieval models. Journal of the American Society for Information Science, 39(1):8–16, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1990:RID

[Los90] Robert M. Losee, Jr. Research: Information in a data collection: Models of database and library quality. Journal of the American Society for Information Science, 41(5):359–367, July 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1992:RGC

[Los92] Robert M. Losee, Jr. Research: A Gray code based ordering for documents on shelves: Classification for browsing and retrieval. Journal of the American Society for Information Science, 43(4): 312–322, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1996:BRBa

[Los96a] Robert Losee. Book review: Measuring information: an information services perspective. Journal of the American Society for Information Science, 47(10):794–795A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1996:BRBb

[Los96b] Robert Losee. Book review: Measuring information: an information services perspective. Journal of the American Society for Information Science, 47(10):794–795A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1996:ERP

[Los96c] Robert M. Losee. Evaluating retrieval performance given database and query characteristics: Analytic determination of performance surfaces. Journal of the American Society for Information Science, 47(1):95–105, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1997:RDI

[Los97a] Robert M. Losee. Research: A discipline independent definition of information. Journal of the American Society for Information Science, 48(3):254–269, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1997:RCB

[Los97b] Robert M. Losee. Research: Comparing Boolean and probabilistic information retrieval systems across queries and disciplines. Journal of the American Society for Information Science, 48(2): 143–156, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:2000:WIR

[Los00] Robert M. Losee. When information retrieval measures agree about the relative quality of document rankings. Journal of the American Society for Information Science, 51(9):834–840, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Loughner:1991:LEC

[Lou91] William Loughner. Letters to the Editor: Concerning paper by Kevin Cook. Journal of the American Society for Information Science, 42(2):113, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Loughner:1992:BCL

[Lou92a] William Loughner. Brief communications: Lotka’s Law and the Kolmogorov–Smirnov test: an error in calculation. Journal of the American Society for Information Science, 43(2):149–150, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See comment [Lon92].

Loughner:1992:LER

[Lou92b] William Loughner. Letters to the Editor: Rejoinder to Charles A. Longfellow, Jr.’s letter. Journal of the American Society for Information Science, 43(7):518, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lovins:1971:EES

[Lov71] Julie B. Lovins. Error evaluation for stemming algorithms as clustering algorithms. Journal of the American Society for Information Science, 22(1):28–40, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lowe:1971:ERK

[Low71] Thomas C. Lowe. Effectiveness of retrieval Key Abbreviation Schemes. Journal of the American Society for Information Science, 22(6):374–381, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Levine:1983:LE

[LP83] Marilyn Levine and Timothy Perper. Letters to the Editor. Journal of the American Society for Information Science, 34(2):156– 157, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1994:PLE

[LP94] Lois F. Lunin and Robin P. Peek. Perspectives on … electronic publishing: Introduction and overview. Journal of the American Society for Information Science, 45(10):726–729, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Logan:1997:I

[LP97] Elisabeth Logan and Marvin Pollard. Introduction. Journal of the American Society for Information Science, 48(7):581–582, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Losee:1999:RMS

[LP99] Robert M. Losee and Lee Anne H. Paris. Research: Measuring search-engine quality and query difficulty: Ranking with target and freestyle. Journal of the American Society for Information Science, 50(10):882–889, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Landry:1970:OPT

[LR70] Bertrand C. Landry and James E. Rush. Opinion paper: Toward a theory of indexing — II. Journal of the American Society for Information Science, 21(5):358–367, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lowe:1972:ADS

[LR72] Thomas C. Lowe and David C. Roberts. Analysis of directory searching. Journal of the American Society for Information Science, 23(3):143–149, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Long:1979:RLA

[LR79] Susan Long and David Rawnsley. The role of a linking agent in an information center network. Journal of the American Society for Information Science, 30(6):330–333, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1989:PLH

[LR89] Lois F. Lunin and Roy Rada. Perspectives on … hypertext: Introduction and overview. Journal of the American Society for Information Science, 40(3):158–163, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lipetz:1970:HMC

[LS70] Ben-Ami Lipetz and Czetong T. Song. How many cards per file guide? Optimizing the two-level file. Journal of the American Society for Information Science, 21(2):140–141, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liston:1971:SAD

[LS71] David M. Liston, Jr. and Mary L. Schoene. A systems approach to the design of information systems. Journal of the American Society for Information Science, 22(2):115–122, March 1971. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1980:LSC

[LS80a] F. W. Lancaster and Linda C. Smith. On-line systems in the communication process: Projections. Journal of the American Society for Information Science, 31(3):193–200, May 1980. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Line:1980:LEJ

[LS80b] Maurice B. Line and A. Sandison. Letters to the Editor: Journal evaluation. Journal of the American Society for Information Science, 31(6):453, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1984:IO

[LS84a] Lois F. Lunin and Linda C. Smith. Introduction and overview. Journal of the American Society for Information Science, 35(5): 278–279, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1984:PLA

[LS84b] Lois F. Lunin and Linda C. Smith. Perspectives on …: Artificial intelligence: Concepts, techniques, applications, promise. Journal of the American Society for Information Science, 35(5): 277, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Logan:1987:RIC

[LS87] Elisabeth L. Logan and W. M. Shaw, Jr. Research: An investigation of the coauthor graph. Journal of the American Society for Information Science, 38(4):262–268, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1988:PLC

[LS88] Lois F. Lunin and Peter B. Schipma. Perspectives on … CD-ROM for information storage and retrieval: Introduction and overview. Journal of the American Society for Information Science, 39(1):30–33, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Latta:1992:RVR

[LS92a] Gail F. Latta and Keith Swigger. Research: Validation of the repertory grid for use in modeling knowledge. Journal of the American Society for Information Science, 43(2):115–129, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1992:PLI

[LS92b] Lois F. Lunin and Michael B. Spring. Perspectives on … information technology standards: Introduction and overview. Journal of the American Society for Information Science, 43(8):521– 525, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1993:POK

[LS93] Lois F. Lunin and Susan Salasin. Perspectives on … knowledge utilization: Introduction and overview. Journal of the American Society for Information Science, 44(4):214–216, May 1993. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lending:1997:PLI

[LS97] Diane Lending and Detmar W. Straub. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: Impacts of an Integrated Information Center on faculty end-users: a qualitative assessment. Journal of the American Society for Information Science, 48(5):466–471, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Leighton:1999:RFP

[LS99] H. Vernon Leighton and Jaideep Srivastava. Research: First 20 precision among World Wide Web search services (search engines). Journal of the American Society for Information Science, 50(10):870–881, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lied:1974:GSM

[LT74] Terry R. Lied and Don L. Tolliver. A general statistical model for increasing efficiency and confidence in manual data collection systems through sampling. Journal of the American Society for Information Science, 25(5):327–331, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lucey:1993:LEM

[Luc93] John Lucey. Letters to the Editor: Machine aided indexing. Journal of the American Society for Information Science, 44(7):430, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lancaster:1996:EIK

[LUH^{+}96] F. W. Lancaster, Jacob W. Ulvila, Susanne M. Humphrey, Linda C. Smith, Bryce Allen, and Saul Herner. Evaluation of interactive knowledge-based systems: Overview and design for empirical testing. Journal of the American Society for Information Science, 47(1):57–69, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Luke:1996:LEI

[Luk96] Michael Luke. Letters to the Editor: Information retrieval in the interspace. Journal of the American Society for Information Science, 47(8):657, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1976:BPT

[Lun76] Lois F. Lunin. The biblio-profile — a two-in-one package of information: Its preparation, production, marketing, uses. Journal of the American Society for Information Science, 27(2):113–117, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lundeen:1980:BRB

[Lun80] Gerald W. Lundeen. Book review: The Dissemination of Information. J. E. Rowley and C. M. D. Turner. Boulder, CO: Westview Press: 1978: 356 pp. Price: $19.25. Journal of the American Society for Information Science, 31(5):382–383, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1982:MI

[Lun82] Lois F. Lunin. Miscellaneous: Introduction. Journal of the American Society for Information Science, 33(6):374, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1983:PIO

[Lun83] Lois F. Lunin. Perspectives: Introduction and overview. Journal of the American Society for Information Science, 34(6):405–407, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lunin:1990:POA

[LW90] Lois F. Lunin and Howard D. White. Perspectives on … author cocitation analysis: Introduction and overview. Journal of the American Society for Information Science, 41(6):429–432, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lawrence:1974:MLS

[LWG74] Barbara Lawrence, Ben H. Weil, and Margaret H. Graham. Making on-line search available in an industrial research environment. Journal of the American Society for Information Science, 25(6): 364–369, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:1999:RHM

[LWZ99] James Liu, Yan Wu, and Lina Zhou. Research: A hybrid method for abstracting newspaper articles. Journal of the American Society for Information Science, 50(13):1234–1245, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Liu:1991:RDC

[LY91] Chengwen Liu and Clement Yu. Research: Data compression using word encoding with Huffman code. Journal of the American Society for Information Science, 42(9):685–698, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lingras:1998:DMU

[LY98] P. J. Lingras and Y. Y. Yao. Data mining using extensions of the rough set model. Journal of the American Society for Information Science, 49(5):415–422, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lynch:1977:VGR

[Lyn77] Michael F. Lynch. Variety generation — a reinterpretation of Shannon’s mathematical theory of communication, and its implications for information science. Journal of the American Society for Information Science, 28(1):19–25, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lynch:1988:RSR

[Lyn88] Clifford A. Lynch. Research: Some reliability issues in very large databases. Journal of the American Society for Information Science, 39(6):408–420, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lynch:1990:BRB

[Lyn90] Clifford Lynch. Book review: The linked systems project: a networking tool for libraries. Journal of the American Society for Information Science, 41(4):305–306, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lynch:1991:PLI

[Lyn91a] Clifford A. Lynch. Perspectives on … imaging. Advanced applications: The technologies of electronic imaging. Journal of the American Society for Information Science, 42(8):578–585, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lynch:1991:RNR

[Lyn91b] Clifford A. Lynch. Research: Nonmaterialized relations and the support of information retrieval applications by relational database systems. Journal of the American Society for Information Science, 42(6):389–396, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lynch:1994:PLE

[Lyn94] Clifford A. Lynch. Perspectives on … electronic publishing: The integrity of digital information: Mechanics and definitional issues. Journal of the American Society for Information Science, 45(10): 737–744, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lynch:1998:PLI

[Lyn98] Clifford Lynch. Perspective on … Internet issues/technological issues: The evolving Internet: Applications and network service infrastructure. Journal of the American Society for Information Science, 49(11):961–972, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Lytle:1988:PLE

[Lyt88] Richard H. Lytle. Perspectives on … education of the information professional: New dimensions, new directions: Part III. The fellow travelers. Information resource management: Research, education, and practice. Journal of the American Society for Information Science, 39(5):337–339, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meincke:1976:KSC

[MA76] Peter P. M. Meincke and Pauline Atherton. Knowledge space: A conceptual basis for the organization of knowledge. Journal of the American Society for Information Science, 27(1):18–24, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacKenzie:1972:LEA

[Mac72] A. Graham MacKenzie. Letters to the Editor: Analytical models for library planning. Journal of the American Society for Information Science, 23(3):222, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Machlup:1979:MRE

[Mac79a] Fritz Machlup. Meeting reviews: An economist’s reflections on an Institute for the Advanced Study of Information Science. Journal of the American Society for Information Science, 30(2):111–113, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Macleod:1979:SLD

[Mac79b] Ian A. Macleod. SEQUEL as a Language for Document Retrieval. Journal of the American Society for Information Science, 30(5): 243–249, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mackesy:1982:PSA

[Mac82] Eileen M. Mackesy. A perspective on secondary access services in the humanities. Journal of the American Society for Information Science, 33(3):146–151, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Macleod:1991:RTR

[Mac91] Ian A. Macleod. Research: Text retrieval and the relational model. Journal of the American Society for Information Science, 42(3):155–165, April 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacLennan:1996:BRB

[Mac96] Alan MacLennan. Book review: The artificial life route to artificial intelligence: Building embodied, situated agents. Journal of the American Society for Information Science, 47(6):482–483, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacRoberts:1997:LER

[Mac97] Michael MacRoberts. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 48(10):963–963a, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Madden:1996:BRBa

[Mad96a] Kenneth G. Madden. Book review: Information management for the intelligent organization: the art of scanning the environment. Journal of the American Society for Information Science, 47(10): 795–796A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Madden:1996:BRBb

[Mad96b] Kenneth G. Madden. Book review: Information management for the intelligent organization: the art of scanning the environment. Journal of the American Society for Information Science, 47(10): 795–796A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Main:1990:BRB

[Mai90] Linda Main. Book review: Effective information centers: Guidelines for MIS and IC managers. Journal of the American Society for Information Science, 41(2):149–150, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mai:1999:BRB

[Mai99] Jens-Erik Mai. Book review: Indexing and abstracting in theory and practice, by F. W. Lancaster. Journal of the American Society for Information Science, 50(8):728–730, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mai:2000:BR

[Mai00] Jens-Erik Mai. Book reviews. Journal of the American Society for Information Science, 51(5):491–492, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Maloney:1974:TVT

[Mal74] Ruth Kay Maloney. Title versus title/abstract text searching in SDI systems. Journal of the American Society for Information Science, 25(6):370–373, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Malone:1999:BRB

[Mal99] Cheryl Knott Malone. Book review: Information literacy: Essential skills for the information age. Journal of the American Society for Information Science, 50(13):1257–1258, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Malone:2000:BRB

[Mal00] Cheryl Knott Malone. Book review: Inventing the Internet, by Janet Abbate. Journal of the American Society for Information Science, 51(8):787–788, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McAllister:1980:CPC

[MAN80] Paul R. McAllister, Richard C. Anderson, and Francis Narin. Comparison of peer and citation assessment of the influence of scientific journals. Journal of the American Society for Information Science, 31(3):147–152, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mansfield:1982:UIS

[Man82a] Jerry W. Mansfield. Using An INS Simulation to Forecast Manpower Needs. Journal of the American Society for Information Science, 33(4):217–222, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mansfield:1982:SMO

[Man82b] Una Mansfield. The systems movement: an overview for information scientists. Journal of the American Society for Information Science, 33(6):375–382, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Maruyama:1971:FRR

[Mar71] Lenore S. Maruyama. Format recognition: a report of a project at the library of congress. Journal of the American Society for Information Science, 22(4):283–287, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcus:1972:BCR

[Mar72a] Richard S. Marcus. Brief communications: Retrieval parameters in growing data bases. Journal of the American Society for Information Science, 23(5):333–334, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marron:1972:LEE

[Mar72b] Harvey Marron. Letters to the Editor: ERIC diagnosis. Journal of the American Society for Information Science, 23(1):71, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marson:1972:LED

[Mar72c] H. Marson. Letters to the Editor: Dissemination and utilization by practitioners. Journal of the American Society for Information Science, 23(2):135–136, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin:1975:BRB

[Mar75] Jean K. Martin. Book review: Essays on bibliography. Compiled and edited by Vito J. Brenni. Metuchen, NJ: Scarecrow Press, 552 pp. (1975). Journal of the American Society for Information Science, 26(6):354–355, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Maron:1977:IRM

[Mar77a] M. E. Maron. On indexing, retrieval and the meaning of about. Journal of the American Society for Information Science, 28(1): 38–43, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin:1977:BRB

[Mar77b] Jean K. Martin. Book review: Advances in Librarianship. Volume 6. Melvin J. Voigt and Michael H. Harris, eds. New York: Academic Press, 291 pp. (1976). (ISBN 0-12-785006-6). Journal of the American Society for Information Science, 28(2):127–128, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcus:1978:LEI

[Mar78] Richard S. Marcus. Letters to the Editor: The importance of recall. Journal of the American Society for Information Science, 29(6):314–315, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Maron:1979:DI

[Mar79] M. E. Maron. Depth of indexing. Journal of the American Society for Information Science, 30(4):224–228, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marchant:1980:BRB

[Mar80] Maurice P. Marchant. Book review: Introduction to the Use of Computer Packages for Statistical Analyses. Moore, Richard W., Englewood Cliffs, NJ: Prentice-Hall; 1978: 115 pp. Price: $9.95. Journal of the American Society for Information Science, 31(1): 64, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Maron:1982:OPA

[Mar82a] M. E. Maron. Opinion paper: Associative Search Techniques versus Probabilistic Retrieval Models. Journal of the American Society for Information Science, 33(5):308–310, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marsh:1982:VT

[Mar82b] Fred E. Marsh, Jr. Videodisc Technology. Journal of the American Society for Information Science, 33(4):237–244, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcus:1983:ECE

[Mar83a] Richard S. Marcus. An experimental comparison of the effectiveness of computers and humans as search intermediaries. Journal of the American Society for Information Science, 34(6):381–404, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin:1983:POO

[Mar83b] Thomas H. Martin. Perspectives On…: Office automation technology and functions: an overview. Journal of the American Society for Information Science, 34(3):210–214, May 1983. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marton:1983:CLH

[Mar83c] János Marton. Causes of low and high citation potentials in science: Citation analysis of biochemistry and plant physiology journals. Journal of the American Society for Information Science, 34(4):244–246, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcus:1985:BRB

[Mar85] Richard S. Marcus. Book review: Intelligent systems, the unprecedented opportunity. Jean E. Hayes and Donald Michie., Eds. Chichester: Ellis Horwood Ltd.; 1983: Isbn 0470-27501-4. Price: $39.95. Journal of the American Society for Information Science, 36(2):135–137, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martinez:1988:LEO

[Mar88] Clara Martinez. Letters to the Editor: The NOT operator. Journal of the American Society for Information Science, 39(3):220, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marchionini:1989:RIS

[Mar89] Gary Marchionini. Research: Information-seeking strategies of novices using a full-text electronic encyclopedia. Journal of the American Society for Information Science, 40(1):54–66, January 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marchionini:1992:PLH

[Mar92a] Gary Marchionini. Perspectives on … human-computer interface: Interfaces for end-user information seeking. Journal of the American Society for Information Science, 43(2):156–163, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin:1992:RII

[Mar92b] F. Douglas Martin. Research: Information interactions between members of science-profession dyads as reflected by journal use: Ichthyology and fisheries biology. Journal of the American Society for Information Science, 43(4):276–283, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martinez:1994:AIT

[Mar94] Michael E. Martinez. Access to information technologies among school-age children: Implications for a democratic society. Journal of the American Society for Information Science, 45(6):395– 400, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marchionini:1995:BRB

[Mar95] Gary Marchionini. Book review: A small matter of programming perspectives on end user computing. Journal of the American Society for Information Science, 46(1):78–79, January 1995. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marvin:1996:BRB

[Mar96] Stephen Marvin. Book review: Library patrons and the law. Journal of the American Society for Information Science, 47(9):722, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin:1998:ITE

[Mar98] Stana B. Martin. Information technology, employment, and the information sector: Trends in information employment 1970– 1995. Journal of the American Society for Information Science, 49(12):1053–1069, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marty:1999:MIC

[Mar99] Paul F. Marty. Museum informatics and collaborative technologies: the emerging socio-technological dimension of information science in museum environments. Journal of the American Society for Information Science, 50(12):1083–1091, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marty:2000:LED

[Mar00] Paul F. Marty. On-line exhibit design: the sociotechnological impact of building a museum over the World Wide Web. Journal of the American Society for Information Science, 51(1):24–32, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mason:1997:SRS

[Mas97] James David Mason. SGML and related standards: New directions as the second decade begins. Journal of the American Society for Information Science, 48(7):593–596, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mathews:1970:UTS

[Mat70] William D. Mathews. Using the TIP system in the ASIS file management exercise. Journal of the American Society for Information Science, 21(3):204–208, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mattessich:1982:SAV

[Mat82] Richard Mattessich. The systems approach: Its variety of aspects. Journal of the American Society for Information Science, 33(6): 383–394, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Matthews:1985:BRB

[Mat85] Joseph R. Matthews. Book review: Cable for information delivery: a guide for librarians, educators and cable professionals. Brigitte L. Kenney, Ed. White Plains, NY: Knowledge Industry Publications; 1984: 172 pp. Price: $34.50 hard cover; $27.50 soft cover. Journal of the American Society for Information Science, 36(2):139, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mattison:1997:BRBb

[Mat97a] David Mattison. Book review: Technology and copyright law: a guidebook for the library, research, and teaching professions. Journal of the American Society for Information Science, 48(12): 1150–1151, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mattison:1997:BRBa

[Mat97b] David Mattison. Book review: The war of desire and technology at the close of the mechanical age. Journal of the American Society for Information Science, 48(6):574–575, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mather:2000:RLA

[Mat00] Laura A. Mather. Research: A linear algebra measure of cluster quality. Journal of the American Society for Information Science, 51(7):602–613, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mauerhoff:1971:BCJ

[Mau71] Georg Mauerhoff. Brief communications: Journal coverage overlap in Current Contents(R) editions. Journal of the American Society for Information Science, 22(5):352–353, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mayes:1978:BCC

[May78] Paul B. Mayes. Brief communications: A comparison of the readability of synopses and original articles for engineering synopses. Journal of the American Society for Information Science, 29(6): 312–313, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

May:1997:RAC

[May97] Andrew D. May. Research: Automatic classification of E-mail messages by message type. Journal of the American Society for Information Science, 48(1):32–39, January 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McAllister:1971:HFD

[MB71] Caryl McAllister and John M. Bell. Human factors in the design of an interactive library system. Journal of the American Society for Information Science, 22(2):96–104, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mankin:1981:ADB

[MB81a] Carole J. Mankin and Jacqueline D. Bastille. An analysis of the differences between density-of-use ranking and raw-use ranking of library journal use. Journal of the American Society for Information Science, 32(3):224–228, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1981:PJU

[MB81b] Katherine W. McCain and James E. Bobick. Patterns of journal use in a departmental library: a citation analysis. Journal of the American Society for Information Science, 32(4):257–267, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mahapatra:1986:RIP

[MB86] Manoranjan Mahapatra and Subal Chandra Biswas. Research: Interdependence of PRECIS role operators: a quantitative analysis of their associations. Journal of the American Society for Information Science, 37(1):20–25, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meghabghab:1991:RAI

[MB91] G. Meghabghab and D. Bilal. Research: Application of information theory to query negotiation: Toward an optimal questioning strategy. Journal of the American Society for Information Science, 42(6):457–462, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moffat:1995:RSG

[MB95] Alistair Moffat and Timothy A. H. Bell. Research: In situ generation of compressed inverted files. Journal of the American Society for Information Science, 46(7):537–550, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Montgomery:1976:CBM

[MBFK76] K. Leon Montgomery, Stephen Bulick, John Fetterman, and Allen Kent. Cost-benefit model of library acquisitions in terms of use: Progress report. Journal of the American Society for Information Science, 27(1):73–74, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marchionini:1994:BCE

[MBH94] Gary Marchionini, Diane Barlow, and Linda Hill. Brief communications: Extending retrieval strategies to networked environments: Old ways, new ways, and a critical look at WAIS. Journal of the American Society for Information Science, 45(8): 561–564, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mokhtarian:1998:IGO

[MBS98] Patricia L. Mokhtarian, Michael N. Bagley, and Ilan Salomon. The impact of gender, occupation, and presence of children on telecommuting motivations and constraints. Journal of the American Society for Information Science, 49(12):1115–1134, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin-Bautista:1999:RFG

[MBVL99] María J. Martín-Bautista, María-Amparo Vila, and Henrik Legind Larsen. Research: A fuzzy genetic algorithm approach to an adaptive information retrieval agent. Journal of the American Society for Information Science, 50(9):760–771, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1989:ROA

[MCBC89] Charles T. Meadow, Barbara A. Cerny, Christine L. Borgman, and Donald O. Case. Research: Online access to knowledge: System design. Journal of the American Society for Information Science, 40(2):86–98, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McClure:1978:BCA

[McC78] Charles R. McClure. Brief communications: Academic librarians’ jobtype and contact with information sources. Journal of the American Society for Information Science, 29(6):311–312, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCarn:1980:MIL

[McC80] Davis B. McCarn. MEDLINE: an introduction to on-line searching. Journal of the American Society for Information Science, 31(3):181–192, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1984:RLA

[McC84] Katherine W. McCain. Research: Longitudinal author cocitation mapping: the changing structure of macroeconomics. Journal of the American Society for Information Science, 35(6):351–359, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1986:RCA

[McC86] Katherine W. McCain. Research: Cocited author mapping as a valid representation of intellectual structure. Journal of the American Society for Information Science, 37(3):111–122, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1988:BCE

[McC88] Katherine W. McCain. Brief communications: Evaluating cocited author search performance in a collaborative specialty. Journal of the American Society for Information Science, 39(6):428–431, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1989:RDC

[McC89] Katherine W. McCain. Research: Descriptor and citation retrieval in the Medical Behavioral Sciences literature: Retrieval overlaps and novelty distribution. Journal of the American Society for Information Science, 40(2):110–114, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1990:POA

[McC90] Katherine W. McCain. Perspectives on … author cocitation analysis: Mapping authors in intellectual space: a technical overview. Journal of the American Society for Information Science, 41(6):433–443, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1991:RME

[McC91] Katherine W. McCain. Research: Mapping economics through the journal literature: an experiment in journal cocitation analysis. Journal of the American Society for Information Science, 42(4):290–296, May 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1995:RBC

[McC95] Katherine W. McCain. Research: Biotechnology in context: a database-filtering approach to identifying core and productive non-core journals supporting multidisciplinary R&D. Journal of the American Society for Information Science, 46(4):306–317, May 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1996:BRB

[McC96] Katherine W. McCain. Book review: Dictionary of bibliometrics. Journal of the American Society for Information Science, 47(9): 716–717A, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:1997:LED

[McC97] Katherine W. McCain. Letters to the Editor: Dictionary of bibliometrics — response to Lancaster. Journal of the American Society for Information Science, 48(5):480–481, May 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCain:2000:SDR

[McC00] Katherine W. McCain. Sharing digitized research-related information on the World Wide Web. Journal of the American Society for Information Science, 51(14):1321–1327, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mittman:1973:MDS

[MCD73] B. Mittman, R. Chalice, and D. Dillaman. Mixed data structures in a multi-purpose retrieval system. Journal of the American Society for Information Science, 24(2):135–141, March 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mcdonough:1982:MPD

[Mcd82] Carol C. Mcdonough. Measurement of the potential demand for academic and professional journals: A methodology. Journal of the American Society for Information Science, 33(5):321– 324, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McDonough:1999:RDS

[McD99] Jerome P. McDonough. Research: Designer selves: Construction of technologically mediated identity within graphical, multiuser virtual environments. Journal of the American Society for Information Science, 50(10):855–869, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McGee:1973:AMC

[McG73] Rob McGee. An analysis of manual circulation systems for academic libraries. Journal of the American Society for Information Science, 24(3):210–217, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McGill:1976:KIS

[McG76] Michael J. McGill. Knowledge and information spaces: Implications for retrieval systems. Journal of the American Society for Information Science, 27(4):205–210, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mcgrath:1979:TMP

[Mcg79] William E. Mcgrath. Two models for predicting subject circulation: A contribution to the allocation problem. Journal of the American Society for Information Science, 30(5):264–268, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McGrath:1984:AMS

[McG84] William E. McGrath. Applications of multidimensional scaling: Comment on “A method for studying intercorrelated circulation patterns in library systems”. Journal of the American Society for Information Science, 35(2):122–123, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Smi84a].

McGlamery:1995:BRB

[McG95] Susan McGlamery. Book review: Cases in online search strategy. Journal of the American Society for Information Science, 46(1): 77–78, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McGrath:1996:BRB

[McG96a] William E. McGrath. Book review: Chaos and catastrophe theories. Journal of the American Society for Information Science, 47(4):330–332, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McGrath:1996:RPA

[McG96b] William E. McGrath. Research: Periodicity in academic library circulation: a spectral analysis. Journal of the American Society for Information Science, 47(2):136–145, February 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McIlwaine:1997:DPT

[McI97] I. C. McIlwaine. Development of particular techniques: The Universal Decimal Classification: Some factors concerning its origins, development, and influence. Journal of the American Society for Information Science, 48(4):331–339, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McKechnie:1996:BRBa

[McK96a] Lynne (E. F.) McKechnie. Book review: Genreflecting: a guide to reading interests in genre fiction. Journal of the American Society for Information Science, 47(5):404–405A, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McKechnie:1996:BRBb

[McK96b] Lynne (E. F.) McKechnie. Book review: Genreflecting: a guide to reading interests in genre fiction. Journal of the American Society for Information Science, 47(5):404–405A, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McKnight:2000:PCI

[McK00] Cliff McKnight. The personal construction of information space. Journal of the American Society for Information Science, 51(8): 730–733, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McLaughlin:1979:SSA

[McL79] Patricia McLaughlin. Symposium: Symposium on alternative methods of disseminating information. Journal of the American Society for Information Science, 30(4):214, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McLain:1990:BRBa

[McL90a] John P. McLain. Book review: Bibliometrics toolbox. Journal of the American Society for Information Science, 41(1):70–71, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McLain:1990:BRBb

[McL90b] John P. McLain. Book review: User-oriented computer languages: Analysis & design. Journal of the American Society for Information Science, 41(1):72, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McMullen:1973:LEI

[McM73] R. M. McMullen. Letters to the Editor: Information-value and cost. Journal of the American Society for Information Science, 24(5):404, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McMurry:1976:BRB

[McM76] Glenn McMurry. Book review: Film Literature Index. Volume 1/Number 4/1973: Annual Cumulation. Vincent J. Aceto, Jane Graves and Fred Silva, eds. New York: Filmdex, Inc., 430 pp. (1975). Journal of the American Society for Information Science, 27(4):262–263, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McNeese:1998:I

[McN98] Michael D. McNeese. Introduction. Journal of the American Society for Information Science, 49(9):773–775, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin:1973:UII

[MCT73] Thomas H. Martin, James Carlisle, and Siegfried Treu. The user interface for interactive bibliographic searching: an analysis of the attitudes of nineteen information scientists. Journal of the American Society for Information Science, 24(2):142–147, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mandersloot:1970:TCS

[MDS70] Wim G. B. Mandersloot, Eleanor M. B. Douglas, and Neville Spicer. Thesaurus control — The selection, grouping, and crossreferencing of terms for inclusion in a coordinate index word list. Journal of the American Society for Information Science, 21(1): 49–57, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McQuire:1998:RAN

[ME98] April R. McQuire and Caroline M. Eastman. Research: The ambiguity of negation in natural language queries to information retrieval systems. Journal of the American Society for Information Science, 49(8):686–692, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1972:OPI

[Mea72] Charles T. Meadow. Opinion papers: Issues in information technology in the United States. Journal of the American Society for Information Science, 23(5):292–298, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1978:ENC

[Mea78a] Charles T. Meadow. Editor’s note: Call for papers. Journal of the American Society for Information Science, 29(4):170, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1978:ENI

[Mea78b] Charles T. Meadow. Editor’s notes: Information science and world politics. Journal of the American Society for Information Science, 29(5):218, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1979:ENS

[Mea79a] Charles T. Meadow. Editor’s note: Scope and administration of JASIS. Journal of the American Society for Information Science, 30(4):178–179, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1979:EN

[Mea79b] Charles T. Meadow. Editor’s notes. Journal of the American Society for Information Science, 30(3):122, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1979:ENN

[Mea79c] Charles T. Meadow. Editor’s notes: A new JASIS department. Journal of the American Society for Information Science, 30(6): 323, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1979:ENH

[Mea79d] Charles T. Meadow. Editor’s notes: Homage to JASIS referees. Journal of the American Society for Information Science, 30(6): 324, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1980:ENR

[Mea80] Charles T. Meadow. Editor’s note: On refereeing of JASIS papers. Journal of the American Society for Information Science, 31(5):307–308, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1981:ENC

[Mea81] Charles T. Meadow. Editor’s note: Changes in the editorial staff. Journal of the American Society for Information Science, 32(5): 324, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1982:BRBb

[Mea82a] Charles T. Meadow. Book review: The Soul of a New Machine. Tracy Kidder. Boston: Little Brown; 1981. Journal of the American Society for Information Science, 33(5):349–350, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1982:BRBa

[Mea82b] Charles T. Meadow. Book review: Word frequencies of spoken American English. Hartvig, Dahl, Essex, CT, USA, and Aylesbury, Buckinghamshire, England: a verbatim book, distributed by Gale Research Company; 1979. Journal of the American Society for Information Science, 33(1):61, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1982:ENNb

[Mea82c] Charles T. Meadow. Editor’s note: A New Feature: The European Research Letter. Journal of the American Society for Information Science, 33(5):261, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1982:ENNa

[Mea82d] Charles T. Meadow. Editor’s notes: Notice concerning notices. Journal of the American Society for Information Science, 33(2): 68, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1983:ENS

[Mea83a] Charles I. Meadow. Editor’s note: Scope and administration of JASIS. Journal of the American Society for Information Science, 34(6):372–373, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1983:BRB

[Mea83b] Charles T. Meadow. Book review: Dictionary of computing, data communications, hardware and software basics, digital electronics. Frank J. Galland. Chichester, England: John Wiley & Sons, Inc.; 1982: 330 pp. Journal of the American Society for Information Science, 34(4):302, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1983:BCU

[Mea83c] Charles T. Meadow. Brief communication: User adaptation in interactive information retrieval. Journal of the American Society for Information Science, 34(4):289–291, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1983:LEA

[Mea83d] Charles T. Meadow. Letters to the Editor: Admissibility of delegates to international meetings. Journal of the American Society for Information Science, 34(4):300, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Dav84, Mea84c].

Meadow:1984:ENNb

[Mea84a] Charles T. Meadow. Editor’s note: A new editor for JASIS. Journal of the American Society for Information Science, 35(6): 323, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1984:ENNa

[Mea84b] Charles T. Meadow. Editor’s note: A note on double blind refereeing. Journal of the American Society for Information Science, 35(4):203, July 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1984:LEA

[Mea84c] Charles T. Meadow. Letters to the Editors: Admissibility of delegates — author’s reply. Journal of the American Society for Information Science, 35(2):135, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Mea83d, Dav84].

Meadow:1985:LER

[Mea85] Charles T. Meadow. Letters to the Editors: Relevance? Journal of the American Society for Information Science, 36(5):354–355, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1986:PLT

[Mea86] Charles T. Meadow. Perspectives on … telecommunications: Principles, developments, prospects: Networks and distributed information services. Journal of the American Society for Information Science, 37(6):405–408, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1987:LEH

[Mea87] Charles T. Meadow. Letters to the Editor: History of databases. Journal of the American Society for Information Science, 38(4): 309, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1994:EBR

[Mea94] Charles T. Meadow. Editorial: JASIS: a reflection ten years after. Journal of the American Society for Information Science, 45(8): 545–546, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1995:RLI

[Mea95] Charles T. Meadow. Research: A life in the information trade. Journal of the American Society for Information Science, 46(3): 217–224, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1997:LER

[Mea97] Charles T. Meadow. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 48(10):964–964a, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1998:LEW

[Mea98] Charles T. Meadow. Letters to the Editor: The whole history of relevance. Journal of the American Society for Information Science, 49(4):392, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:2000:LED

[Mea00] Charles T. Meadow. Letter to the Editors: Differences between novice and experienced users in searching information on the World Wide Web. Journal of the American Society for Information Science, 51(12):1154, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mechling:1999:RBF

[Mec99] Jerry Mechling. Research: Better funding for government IT: Views from the front line. Journal of the American Society for Information Science, 50(4):305–313, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Melnyk:1972:SPA

[Mel72] Vera Melnyk. 1972 student paper award man-machine interface: Frustration. Journal of the American Society for Information Science, 23(6):392–401, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Menou:1985:OSM

[Men85] Michel J. Menou. An overview of social measures of information. Journal of the American Society for Information Science, 36(3): 169–177, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mendelsohn:1992:RTT

[Men92] Loren D. Mendelsohn. Research: Technology transfer policy: Its role as a scientific and technical information policy and its impact on technological growth. Journal of the American Society for Information Science, 43(1):80–88, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mendelsohn:1995:POC

[Men95a] Loren D. Mendelsohn. Perspectives on … the Chemist’s workstation: Introduction and overview. Journal of the American Society for Information Science, 46(8):609–610, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Menou:1995:LEM

[Men95b] Michel J. Menou. Letters to the Editor: Modernizing book reviews. Journal of the American Society for Information Science, 46(9):717, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meredith:1971:MAA

[Mer71] J. C. Meredith. Machine-assisted approach to general reference materials. Journal of the American Society for Information Science, 22(3):176–186, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mess:1995:POC

[Mes95] John A. Mess. Perspectives on … the Chemist’s workstation: Design considerations for creating a chemical information workstation. Journal of the American Society for Information Science, 46(8):632–637, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moll:1988:PLE

[MF88] Joy K. Moll and Barbara J. Flood. Perspectives on … education of the information professional: New dimensions, new directions: Information science and information-related educational programs: Their diversity and accreditation. Journal of the American Society for Information Science, 39(5):318–322, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McMurtray:1972:CPC

[MG72] Frances McMurtray and John M. Ginski. Citation patterns of the cardiovascular serial literature. Journal of the American Society for Information Science, 23(3):172–175, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Masys:1987:PLF

[MH87] Daniel R. Masys and Susan M. Hubbard. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Technical information programs of the National Cancer Institute. Journal of the American Society for Information Science, 38(1):60–64, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1982:CIIa

[MHA82a] Charles T. Meadow, Thomas T. Hewett, and Elizabeth S. Aversa. A computer intermediary for interactive database searching. I. Design. Journal of the American Society for Information Science, 33(5):325–332, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1982:CIIb

[MHA82b] Charles T. Meadow, Thomas T. Hewett, and Elizabeth S. Aversa. A computer intermediary for interactive database searching. II. Evaluation. Journal of the American Society for Information Science, 33(6):357–364, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Metzler:1989:RCO

[MHCW89] Douglas P. Metzler, Stephanie W. Haas, Cynthia L. Cosic, and Leslie H. Wheeler. Research: Constituent object parsing for information retrieval and similar text processing problems. Journal of the American Society for Information Science, 40(6):398–423, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mick:1983:POP

[Mic83] Colin K. Mick. Perspectives On…: planning for office automation. Journal of the American Society for Information Science, 34(3):229–232, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Michel:1994:RWU

[Mic94] Dee Andy Michel. Research: What is used during cognitive processing in information retrieval and library searching? Eleven sources of search information. Journal of the American Society for Information Science, 45(7):498–514, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mignon:1979:BRB

[Mig79] Edmond Mignon. Book review: On-Line Information Retrieval Sourcebook. J. L. Hall. London: ASLIB; 1977: 267 pp. Price $43.50. ISBN: 0-85142-106-7. Journal of the American Society for Information Science, 30(1):63, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mignon:1980:BRB

[Mig80a] Edmond Mignon. Book review: Progress in Communication Sciences, Volume 1. Melvin J. Voigt and Gerhard J. Hanneman. Norwood, NJ: ABLEX Publishing Corp.; 1979: 198 pp. ISBN 0-89391-010-4; ISSN 0163-5689. Journal of the American Society for Information Science, 31(5):381–382, September 1980. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mignon:1980:LEI

[Mig80b] Edmond Mignon. Letters to the Editor: Indexing and behavior. Journal of the American Society for Information Science, 31(4):

302–303, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Miller:1971:EMT

[Mil71] William L. Miller. The efficiency of MEDLARS titles for retrieval. Journal of the American Society for Information Science, 22(5): 318–321, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Miller:1987:LEO

[Mil87] Barbara Miller. Letters to the Editor: Online information retrieval reading list. Journal of the American Society for Information Science, 38(4):307, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Millican:1990:BRB

[Mil90] Rita Millican. Book review: The twenty-first century; technology’s impact on academic research and law libraries. Journal of the American Society for Information Science, 41(4):306–307, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Miller:1993:BRB

[Mil93] Wendy Miller. Book review: Envisioning information. Journal of the American Society for Information Science, 44(4):212, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Milstead:1994:PLI

[Mil94] Jessica L. Milstead. Perspectives on … indexing: Needs for research in indexing. Journal of the American Society for Information Science, 45(8):577–582, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Milstead:1995:LEK

[Mil95] Jessica L. Milstead. Letters to the Editor: Kudos for Delphi study. Journal of the American Society for Information Science, 46(10):800, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Milekic:2000:DDE

[Mil00] Slavko Milekic. Designing digital environments for art education/ exploration. Journal of the American Society for Information Science, 51(1):49–56, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mims:1985:BRB

[Mim85] Ted Mims. Book review: MacMillan easy home computer series. (1) The Apple IIe User’s Guide. Mark Andrews. 121 pp.

(2) The Commodore 64 User’s Guide. Jonathan Sacks and Mark Andrews. 121 pp. (3) The IBM PC User’s Guide. Michael Bane and Mark Andrews. 121 pp. (4) The TI 99/4A User’s Guide. Bill Brewer and Mark Andrews. 121 pp. Roger Sharpe, Ed. New York: MacMillan; 1984. Price: $5.95 each. Journal of the American Society for Information Science, 36(4):273, July 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Minker:1978:OPS

[Min78] Jack Minker. Opinion paper: Science, Shcharansky, and the Soviets. Journal of the American Society for Information Science, 29 (5):219–224, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Miyamoto:1998:ARS

[Miy98] Sadaaki Miyamoto. Application of rough sets to information retrieval. Journal of the American Society for Information Science, 49(3):195–205, March 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mizzaro:1997:RWH

[Miz97] Stefano Mizzaro. Relevance: the whole history. Journal of the American Society for Information Science, 48(9):810–832, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mizrachi:1998:BRB

[Miz98] Yonathan Mizrachi. Book review: Modern information systems for managers. Journal of the American Society for Information Science, 49(6):571–573, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moody:1972:DUG

[MK72] D. W. Moody and Olaf Kays. Development of the U.S. Geological Survey bibliographic system using Gipsy. Journal of the American Society for Information Science, 23(1):39–49, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marulli:1979:BCB

[MK79] Luciana Marulli and Michael E. D. Koenig. Brief communications: Bradford distribution of data elements. Journal of the American Society for Information Science, 30(2):107–108, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcus:1978:CIT

[MKB78] Richard S. Marcus, Peter Kugel, and Alan R. Benenfeld. Catalog information and text as indicators of relevance. Journal of the American Society for Information Science, 29(1):15–30, January 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1983:ENB

[MKDL83] Charles T. Meadow, Donald H. Kraft, M. Carl Droit, and Lois F. Lunin. Editor’s note: JASIS editorial changes. Journal of the American Society for Information Science, 34(6):371, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McCarn:1990:RMM

[ML90] Davis B. McCarn and Craig M. Lewis. Research: A mathematical model of retrieval system performance. Journal of the American Society for Information Science, 41(7):495–500, October 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mick:1980:TUU

[MLC80] Colin K. Mick, Georg N. Lindsey, and Daniel Callahan. Toward usable user studies. Journal of the American Society for Information Science, 31(5):347–356, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Murugesan:1978:VNC

[MM78] Poovanalingam Murugesan and Michael J. Moravcsik. Variation of the nature of citation measures with journals and scientific specialties. Journal of the American Society for Information Science, 29(3):141–147, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacRoberts:1987:BCA

[MM87] M. H. MacRoberts and B. R. MacRoberts. Brief communications: Another test of the normative theory of citing. Journal of the American Society for Information Science, 38(4):305–306, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacRoberts:1988:BCA

[MM88] M. H. MacRoberts and B. R. MacRoberts. Brief communications: Author motivation for not citing influences: a methodological note. Journal of the American Society for Information Science, 39(6):432–433, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacRoberts:1989:RPC

[MM89] Michael H. MacRoberts and Barbara R. MacRoberts. Research: Problems of citation analysis: a critical review. Journal of the American Society for Information Science, 40(5):342–349, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacRoberts:1997:BCC

[MM97] M. H. MacRoberts and B. R. MacRoberts. Brief communications: Citation content analysis of a botany journal. Journal of the American Society for Information Science, 48(3):274–275, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ma:1998:TTB

[MM98] Wei-Ying Ma and B. S. Manjunath. A texture thesaurus for browsing large aerial photographs. Journal of the American Society for Information Science, 49(7):633–648, May 15, 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McClure:1999:RDA

[MMB99] Charles R. McClure, William E. Moen, and John Carlo Bertot. Research: Descriptive assessment of information policy initiatives: the government information locator service (GILS) as an example. Journal of the American Society for Information Science, 50(4):314–330, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Miyamoto:1981:DCT

[MN81] Sadaaki Miyamoto and Kazuhiko Nakayama. Determination of the conservation time of periodicals for optimal shelf maintenance of a library. Journal of the American Society for Information Science, 32(4):268–274, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McAllister:1983:CRP

[MN83a] Paul R. McAllister and Francis Narin. Characterization of the research papers of U.S. medical schools. Journal of the American Society for Information Science, 34(2):123–131, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Miyamoto:1983:TTS

[MN83b] S. Miyamoto and K. Nakayama. A technique of two-stage clustering applied to environmental and civil engineering and related methods of citation analysis. Journal of the American Society for Information Science, 34(3):192–201, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moghdam:1975:UTL

[Mog75] Dineh Moghdam. User training for on-line information retrieval systems. Journal of the American Society for Information Science, 26(3):184–188, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Molka:1992:PLI

[Mol92] Judith A. Molka. Perspectives on … information technology standards: Surrounded by standards, there is a simpler view. Journal of the American Society for Information Science, 43(8): 526–530, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Molina:1995:RDA

[Mol95] Maria Pinto Molina. Research: Documentary abstracting: Toward a methodological model. Journal of the American Society for Information Science, 46(3):225–234, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Momenee:1982:NDC

[Mom82] Karen Momenee. A notation for describing the components of library information systems. Journal of the American Society for Information Science, 33(5):270–280, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Montgomery:1972:LIS

[Mon72] Christine A. Montgomery. Linguistics and information science. Journal of the American Society for Information Science, 23(3): 195–219, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Montgomery:1973:BCP

[Mon73] Ruth Reinke Montgomery. Brief communications: Productivity in toxicity papers. Journal of the American Society for Information Science, 24(5):402–403, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Monsell:1981:RPM

[Mon81] Stephen Monsell. Representations, processes, memory mechanisms: the basic components of cognition. Journal of the American Society for Information Science, 32(5):378–390, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mondschein:1994:RSI

[Mon94] Lawrence G. Mondschein. The role of spatial information systems in environmental emergency management. Journal of the American Society for Information Science, 45(9):678–685, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moor:1972:DPT

[Moo72] William C. Moor. The development and preliminary test of behavior related dimensions of information systems. Journal of the American Society for Information Science, 23(1):50–57, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moore:1973:IST

[Moo73] J. R. Moore. On interrelationships of the sciences and technology as expressed by a categorized list of journals and modified by a classification system. Journal of the American Society for Information Science, 24(5):359–367, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moore:1986:BRB

[Moo86] Douglas C. Moore. Book review: The Wordtree. Journal of the American Society for Information Science, 37(2):92–94, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Morrow:1976:GFU

[Mor76] Deanna I. Morrow. A generalized flowchart for the use of OR-BIT(R) and other on-line interactive bibliographic search systems. Journal of the American Society for Information Science, 27(1):57–62, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Morton:1978:DMT

[Mor78] Sara E. Morton. Designing a multilingual terminology bank for united states translators. Journal of the American Society for Information Science, 29(6):297–303, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Morse:1981:IEB

[Mor81] Philip M. Morse. Implications of the exact Bradford distribution. Journal of the American Society for Information Science, 32(1): 43–50, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Morris:1994:RTU

[Mor94] Ruth C. T. Morris. Research: Toward a user-centered information service. Journal of the American Society for Information Science, 45(1):20–30, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moss:1970:BCA

[Mos70] R. Moss. Brief communications: An application of managerial cost accounting to a science information center. Journal of the American Society for Information Science, 21(2):163–164, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mosher:1972:LEA

[Mos72] Edward J. Mosher. Letters to the Editor: 1972 ASIS meeting. Journal of the American Society for Information Science, 23(4):

286, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mount:1977:NSB

[Mou77] Ellis Mount. A national standard for bibliographic references. Journal of the American Society for Information Science, 28(1): 1–12, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moss:1983:LEO

[MOV83] R. Moss, T. Olafsen, and L. Vokac. Letters to the Editor: Optimal values of recall and precision. Journal of the American Society for Information Science, 34(4):294, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mowshowitz:1992:RMVa

[Mow92a] Abbe Mowshowitz. Research: On the market value of information commodities. I. The nature of information and information commodities. Journal of the American Society for Information Science, 43(3):225–232, April 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mowshowitz:1992:RMVb

[Mow92b] Abbe Mowshowitz. Research: On the market value of information commodities. II. Supply price. Journal of the American Society for Information Science, 43(3):233–241, April 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mowshowitz:1992:RMVc

[Mow92c] Abbe Mowshowitz. Research: On the market value of information commodities. III. Demand price. Journal of the American Society for Information Science, 43(3):242–248, April 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martin:1974:ECA

[MP74] Jean K. Martin and Ronald G. Parsons. Evaluation of current awareness service for physics and astronomy literature. Journal of the American Society for Information Science, 25(3):156–160, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Melin:1998:HCB

[MP98] Göran Melin and Olle Persson. Hotel cosmopolitan: a bibliometric study of collaboration at some European universities. Journal of the American Society for Information Science, 49(1):43–48, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Minker:1973:DRE

[MPW73] Jack Minker, Eero Peltola, and Gerald A. Wilson. Document retrieval experiments using cluster analysis. Journal of the American Society for Information Science, 24(4):246–260, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mostafa:1998:FMD

[MQP98] J. Mostafa, L. M. Quiroga, and M. Palakal. Filtering medical documents using automated and human classification methods. Journal of the American Society for Information Science, 49(14): 1304–1318, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcus:1981:TCIa

[MR81a] Richard S. Marcus and J. Francis Reintjes. A translating computer interface for end-user operation of heterogeneous retrieval systems. I. Design. Journal of the American Society for Information Science, 32(4):287–303, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcus:1981:TCIb

[MR81b] Richard S. Marcus and J. Francis Reintjes. A translating computer interface for end-user operation of heterogeneous retrieval systems. II. Evaluations. Journal of the American Society for Information Science, 32(4):304–317, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Macleod:1987:RAM

[MR87] Ian A. Macleod and A. R. Reuber. Research: The Array Model: a conceptual modeling approach to document retrieval. Journal of the American Society for Information Science, 38(3):162–170, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Min:1994:RME

[MR94] Zheng Min and Roy Rada. Research: MUCH electronic publishing environment: Principles and practices. Journal of the American Society for Information Science, 45(5):300–309, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mitchell:1973:SSL

[MRW73] P. C. Mitchell, J. T. Rickman, and W. E. Walden. SOLAR: a storage and on-line automatic retrieval system. Journal of the American Society for Information Science, 24(5):347–358, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mathis:1973:IAA

[MRY73] Betty A. Mathis, James E. Rush, and Carol E. Young. Improvement of automatic abstracts by the use of structural analysis. Journal of the American Society for Information Science, 24(2): 101–109, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

MacIntyre:1976:BCC

[MS76] Walter M. MacIntyre and John M. Sowden. Brief communications: Current awareness with an augmented chemical titles file. Journal of the American Society for Information Science, 27(3): 171–174, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Marcoux:1997:WSW

[MS97] Yves Marcoux and Martin Sévigny. Why SGML? Why now? Journal of the American Society for Information Science, 48(7): 584–592, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meho:2000:RCR

[MS00] Lokman I. Meho and Diane H. Sonnenwald. Research: Citation ranking versus peer evaluation of senior faculty research performance: a case study of Kurdish scholarship. Journal of the American Society for Information Science, 51(2):123–138, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mukhopadhyay:1986:RIS

[MSHB86] Uttam Mukhopadhyay, Larry M. Stephens, Michael N. Huhns, and Ronald D. Bonnell. Research: An intelligent system for document retrieval in distributed office environments. Journal of the American Society for Information Science, 37(3):123–135, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McKinin:1991:RMF

[MSJM91] Emma Jean McKinin, MaryEllen Sievert, E. Diane Johnson, and Joyce A. Mitchell. Research: The MEDLINE/full-text research project. Journal of the American Society for Information Science, 42(4):297–307, May 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Maricic:1998:CCV

[MSPPM98] Siniša Maričić, Jagoda Spaventi, Leo Pavičić, and Greta Pifat-Mrzljak. Citation context versus the frequency counts of citation histories. Journal of the American Society for Information Science, 49(6):530–540, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McFarland:1996:BRB

[MT96] V. Marielle McFarland and Hermann L. Totten. Book review: Out in the cold — Academic boycotts and the isolation of South Africa. Journal of the American Society for Information Science, 47(5):399–401, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Muddamalle:1998:RNL

[Mud98] Manikya Rao Muddamalle. Research: Natural language versus controlled vocabulary in information retrieval: a case study in soil mechanics. Journal of the American Society for Information Science, 49(10):881–887, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moens:1999:RAL

[MUD99] Marie-Francine Moens, Caroline Uyttendaele, and Jos Dumortier. Research: Abstracting of legal cases: the potential of clustering based on the selection of representative objects. Journal of the American Society for Information Science, 50(2):151– 161, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Murphy:1973:BCL

[Mur73] Larry J. Murphy. Brief communications: “Lotka’s Law in the humanities?”. Journal of the American Society for Information Science, 24(6):461–462, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Murphy:1974:LEO

[Mur74] Larry J. Murphy. Letters to the Editors: Other Peoples Mistakes. Journal of the American Society for Information Science, 25(2): 134, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Murphy:1993:BRB

[Mur93] Cynni Murphy. Book review: The reconfigured eye: Visual truth in the post-photographic era. Journal of the American Society for Information Science, 44(9):553, October 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Musa:1986:RSP

[Mus86] F. A. Musa. Research: A system for processing bilingual Arabic/ English text. Journal of the American Society for Information Science, 37(5):288–293, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moed:1995:RIA

[MV95] H. F. Moed and Th. N. Van Leeuwen. Research: Improving the accuracy of Institute for Scientific Information’s Journal Impact Factors. Journal of the American Society for Information Science, 46(6):461–467, July 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McLean:1991:RDD

[MW91] Stuart McLean and Charlotte Weise. Research: Digress: a deductive interface to a relational database. Journal of the American Society for Information Science, 42(1):49–63, January 1991. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

McKibbon:1995:PLM

[MWH^{+}95] K. Ann McKibbon, Nancy Wilczynski, Robert S. Hayward, Cynthia J. Walker-Dilks, and R. Brian Haynes. Perspectives on … medical informatics: Information technology in health care: The medical literature as a resource for health care practice. Journal of the American Society for Information Science, 46(10):737–742, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mitroff:1972:DIS

[MWR72] Ian I. Mitroff, James Williams, and Eugene Rathswohl. Dialectical inquiring systems: a new methodology for information science. Journal of the American Society for Information Science, 23(6):365–378, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1995:RSU

[MWY95] Charles T. Meadow, Jiabin Wang, and Weijing Yuan. Research: A study of user performance and attitudes with information retrieval interfaces. Journal of the American Society for Information Science, 46(7):490–505, August 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Moore:1985:LIS

[MY85] Bessie Boehm Moore and Christina Carr Young. Library/ information services and the nation’s elderly. Journal of the American Society for Information Science, 36(6):364–368, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Myburgh:1998:BRBa

[Myb98a] Sue Myburgh. Book review: Entertainment technology and tomorrow’s information services. Journal of the American Society for Information Science, 49(2):184, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Myburgh:1998:BRBb

[Myb98b] Sue Myburgh. Book review: Netizens: On the history and impact of Usenet and the Internet. Journal of the American Society for Information Science, 49(11):1037–1038, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Myburgh:1999:BRB

[Myb99] Sue Myburgh. Book review: Understanding information retrieval interactions: Theoretical and practical implications. Journal of the American Society for Information Science, 50(13):1257, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Mylonas:1992:PLH

[Myl92] Elli Mylonas. Perspectives on … human-computer interface: An interface to classical Greek civilization. Journal of the American Society for Information Science, 43(2):192–201, March 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Martiner:1978:BCU

[MZ78] Clara Martiner and Irving Zarember. Brief communications: OR NOT: the unused operator. Journal of the American Society for Information Science, 29(4):207–208, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Meadow:1979:SSA

[MZ79] Charles T. Meadow and Mary Ann Zaborowski. Some statistical aspects of JASIS publications. Journal of the American Society for Information Science, 30(6):368–371, November 1979. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [MZ80, Kuc80].

Meadow:1980:LEB

[MZ80] Charles T. Meadow and Mary Ann Zaborowski. Letters to the Editor: JASIS publication statistics — authors’ reply. Journal of the American Society for Information Science, 31(4):302–303, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Kuc80].

Nadziejka:1985:LEF

[Nad85] David E. Nadziejka. Letters to the Editors: Federal information policy. Journal of the American Society for Information Science, 36(3):220, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nahl:1998:PLI

[Nah98] Diane Nahl. Perspectives on … Internet issues/information use issues: Learning the Internet and the structure of information behavior. Journal of the American Society for Information Science, 49(11):1017–1023, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nance:1970:AML

[Nan70] Richard E. Nance. An analytical model of a library network. Journal of the American Society for Information Science, 21(1): 58–66, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nance:1995:BRBb

[Nan95a] William Nance. Book review: Collective intelligence in computerbased collaboration. Journal of the American Society for Information Science, 46(10):793–795, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nance:1995:BRBa

[Nan95b] William Nance. Book review: Groupware in the 21st century: Computer supported cooperative working toward the millennium. Journal of the American Society for Information Science, 46(9): 710–712A, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Narin:1975:NPC

[NC75] Francis Narin and Mark P. Carpenter. National publication and citation comparisons. Journal of the American Society for Information Science, 26(2):80–93, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Neufeld:1986:PLO

[NC86] M. Lynne Neufeld and Martha Cornog. Perspectives on … online searching: Database history: From dinosaurs to compact discs. Journal of the American Society for Information Science, 37(4): 183–190, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Narin:1972:ISJ

[NCB72] Francis Narin, Mark Carpenter, and Nancy C. Berlt. Interrelationships of scientific journals. Journal of the American Society for Information Science, 23(5):323–331, September 1972. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Newman:1971:LE

[NCH^{+}71] Simon M. Newman, John E. Creps, Brigitta Hilliker, Dana L. Roth, E. H. Brenner, and Bill Dantzic. Letters to the Editors. Journal of the American Society for Information Science, 22(5):

356–357, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nakkouzi:1990:RQF

[NE90] Ziad S. Nakkouzi and Caroline M. Eastman. Research: Query formulation for handling negation in information retrieval systems. Journal of the American Society for Information Science, 41(3):171–182, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Neeley:1981:MSS

[Nee81] James D. Neeley, Jr. The management and social science literatures: an interdisciplinary cross-citation analysis. Journal of the American Society for Information Science, 32(3):217–223, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Neill:1980:BCM

[Nei80] S. D. Neill. Brief communication: Mysteries of the deep: Models of the universe of knowledge. Journal of the American Society for Information Science, 31(5):375–377, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nelson:1997:BRB

[Nel97a] Jerold Nelson. Book review: Introduction to United States government information sources (5th ed.). Journal of the American Society for Information Science, 48(5):477–478, May 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nelson:1997:MJT

[Nel97b] Mike Nelson. In memoriam: Jean Tague-Sutcliffe, 1931–1996. Journal of the American Society for Information Science, 48(10): 866, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Neuman:1995:RHS

[Neu95] Delia Neuman. Research: High school students’ use of databases: Results of a national Delphi study. Journal of the American Society for Information Science, 46(4):284–298, May 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Newman:1973:BRB

[New73] Simon M. Newman. Book review: Information Systems, Services, and Centers. Herman M. Weisman. New York: Becker and Hayes, Inc. (A Subsidiary of John Wiley and Sons, Inc.). (1972). 265 + xvi pages, including index, selected bibliography, appendices, and discussion problems and exercises. $10.50. Journal of the American Society for Information Science, 24(1):75, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Newman:1976:BRB

[New76] Simon M. Newman. Book review: Your computer and the law. Robert P. Bigelow and Susan H. Nycum. Englewood Cliffs, N.J.: Prentice-Hall, Inc., 283 pp. (1975). Journal of the American Society for Information Science, 27(2):124–125, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Newhagen:1994:MUP

[New94] John E. Newhagen. Media use and political efficacy: the suburbanization of race and class. Journal of the American Society for Information Science, 45(6):386–394, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nicholls:1987:BCE

[Nic87] Paul Travis Nicholls. Brief communications: Estimation of Zipf parameters. Journal of the American Society for Information Science, 38(6):443–445, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nicholls:1989:RBM

[Nic89] Paul Travis Nicholls. Research: Bibliometric modeling processes and the empirical validity of Lotka’s Law. Journal of the American Society for Information Science, 40(6):379–385, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nicholson:2000:RRW

[Nic00] Scott Nicholson. Raising reliability of web search tool research through replication and chaos theory. Journal of the American Society for Information Science, 51(8):724–729, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Niehoff:1976:DIE

[Nie76] R. T. Niehoff. Development of an integrated energy vocabulary and the possibilities for on-line subject switching. Journal of the American Society for Information Science, 27(1):3–17, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nisonger:1999:RBL

[Nis99] Thomas E. Nisonger. Research: JASIS and library and information science journal rankings: a review and analysis of the last half-century. Journal of the American Society for Information Science, 50(11):1004–1019, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nitecki:1979:NNI

[Nit79] Joseph Z. Nitecki. National network of information in Poland. Journal of the American Society for Information Science, 30(5): 274–279, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Newman:1997:REQ

[NJWC97] D. R. Newman, Chris Johnson, Brian Webb, and Clive Cochrane. Research: Evaluating the quality of learning in computer supported co-operative learning. Journal of the American Society for Information Science, 48(6):484–495, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Norcio:1983:HMO

[NK83] A. F. Norcio and Stephen M. Kerst. Human memory organization for computer programs. Journal of the American Society for Information Science, 34(2):109–114, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nyce:1989:EIP

[NK89] James M. Nyce and Paul Kahn. Epilogue: Innovation, pragmaticism, and technological continuity: Vannevar Bush’s memex. Journal of the American Society for Information Science, 40(3): 214–220, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ng:2000:PEN

[NK00] Kwong Bor Ng and Paul B. Kantor. Predicting the effectiveness of naïve data fusion on the basis of system characteristics. Journal of the American Society for Information Science, 51(13):1177– 1189, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nance:1972:IND

[NKB72] Richard E. Nance, Robert R. Korfhage, and U. Narayan Bhat. Information networks: Definitions and message transfer models. Journal of the American Society for Information Science, 23(4): 237–247, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Noreault:1977:ARO

[NKM77] Terry Noreault, Matthew Koll, and Michael J. McGill. Automatic ranked output from boolean searches in SIRE. Journal of the American Society for Information Science, 28(6):333–339, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Neway:1983:BCC

[NL83] Julie M. Neway and F. W. Lancaster. Brief communication: The correlation between pertinence and rate of citation duplication in multidatabase searches. Journal of the American Society for Information Science, 34(4):292–293, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nelson:1974:EDI

[NM74] John A. Nelson and Ian I. Mitroff. An experiment in dialectic information systems. Journal of the American Society for Information Science, 25(4):252–262, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Noyons:1999:RCM

[NML99] E. C. M. Noyons, H. F. Moed, and M. Luwel. Research: Combining mapping and citation analysis for evaluative bibliometric purposes: a bibliometric study. Journal of the American Society for Information Science, 50(2):115–131, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nederhof:1992:RIC

[NN92] A. J. Nederhof and E. C. M. Noyons. Research: International comparison of departments’ research performance in the humanities. Journal of the American Society for Information Science, 43(3):249–256, April 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Noma:1985:TEP

[NO85] Elliot Noma and Dominic Olivastro. Are there enduring patents? Journal of the American Society for Information Science, 36(5): 297–301, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Noma:1984:CCA

[Nom84] Elliot Noma. Co-citation analysis and the invisible college. Journal of the American Society for Information Science, 35(1):29–33, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Norton:1996:BRB

[Nor96] Melanie J. Norton. Book review: Expertise and technology, cognition and human-computer cooperation. Journal of the American Society for Information Science, 47(9):722–725, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Norton:1998:BRB

[Nor98] Melanie J. Norton. Book review: Web security & commerce. Journal of the American Society for Information Science, 49(10): 955–956, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Novotny:1982:BRB

[Nov82] Eric J. Novotny. Book review: Multinational Computer Nets: The Case of International Banking. Richard H. Veith. Lexington, MA: D. C. Health and Company (Lexington Books); 1981: 133 pp., xvii, Bibliography, Index. Price: $18.95. ISBN 0–669– 04092–4. Journal of the American Society for Information Science, 33(5):349, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nowak:1981:LET

[Now81] Ildiko D. Nowak. Letters to the Editor: Translations of scientific and technical articles. Journal of the American Society for Information Science, 32(1):77, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Narin:1976:SBL

[NPG76] Francis Narin, Gabriel Pinski, and Helen Hofer Gee. Structure of the Biomedical Literature. Journal of the American Society for Information Science, 27(1):25–45, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Neelameghan:1985:ICI

[NT85a] A. Neelameghan and J. Tocatlian. International cooperation in information systems and services. Journal of the American Society for Information Science, 36(3):153–163, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nelson:1985:SSR

[NT85b] Michael J. Nelson and Jean M. Tague. Split size-rank models for the distribution of index terms. Journal of the American Society for Information Science, 36(5):283–296, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nahl:1996:ACS

[NT96] Diane Nahl and Carol Tenopir. Affective and cognitive searching behavior of novice end-users of a full-text database. Journal of the American Society for Information Science, 47(4):276–286, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nugent:1991:PLI

[Nug91] William R. Nugent. Perspectives on … imaging. Advanced applications: Virtual reality: Advanced imaging special effects let you roam in cyberspace. Journal of the American Society for Information Science, 42(8):609–617, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Noyons:1998:MSD

[NvR98] E. C. M. Noyons and A. F. J. van Raan. Monitoring scientific developments from a dynamic perspective: Self-organized structuring to map neural network research. Journal of the American Society for Information Science, 49(1):68–81, ???? 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Nederhof:1991:RQJ

[NZ91] A. J. Nederhof and R. A. Zwaan. Research: Quality judgments of journals as indicators of research performance in the humanities and the social and behavioral sciences. Journal of the American Society for Information Science, 42(5):332–340, June 1991. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OConnor:1970:APD

[O’C70] John O’Connor. Answer-providing documents: Some inference descriptions and text-searching retrieval results. Journal of the American Society for Information Science, 21(6):406–414, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OConnor:1973:TSR

[O’C73] John O’Connor. Text searching retrieval of answer-sentences and other answer-passages. Journal of the American Society for Information Science, 24(6):445–460, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OConnor:1980:APR

[O’C80] John O’Connor. Answer-passage retrieval by text searching. Journal of the American Society for Information Science, 31(4): 227–239, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oen:1988:PLE

[OC88] Carol Oen and Marianne Cooper. Perspectives on … education of the information professional: New dimensions, new directions: Professional identity and the information professional. Journal of the American Society for Information Science, 39(5):355–357, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OConnor:1993:BRB

[O’C93] Brian O’Connor. Book review: Language and representation in information retrieval. Journal of the American Society for Information Science, 44(2):116–117, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OConnor:1995:BRB

[O’C95] Brian C. O’Connor. Book review: Introduction to automation for librarians (3rd ed.). Journal of the American Society for Information Science, 46(1):79–80, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Orwig:1997:RGS

[OCN97] Richard E. Orwig, Hsinchun Chen, and Jay F. Nunamaker, Jr. Research: A graphical, self-organizing approach to classifying electronic meeting output. Journal of the American Society for Information Science, 48(2):157–170, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Otten:1970:OPT

[OD70] Klaus Otten and Anthony Debons. Opinion paper: Towards a metascience of information: Informatology. Journal of the American Society for Information Science, 21(1):89–94, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

ONeill:1987:RCD

[ODVG87] Edward T. O’Neill, Martin Dillon, and Diane Vizine-Goetz. Research: Class dispersion between the Library of Congress Classification and the Dewey Decimal Classification. Journal of the American Society for Information Science, 38(3):197–205, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olsgaard:1981:BCI

[OE81] John N. Olsgaard and John Edward Evans. Brief communications: Improving keyword indexing. Journal of the American Society for Information Science, 32(1):71–72, January 1981. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olle:1970:SAF

[OG70] T. William Olle and Andre M. Gagnoud. A solution to the ASIS file management exercise using RCA’s UL/1. Journal of the American Society for Information Science, 21(3):214–218, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oppenheim:1980:LEH

[OG80] C. Oppenheim and Eugene Garfield. Letters to the Editor: Highly cited old papers. Journal of the American Society for Information Science, 31(3):219, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ottensmann:1993:RIT

[OG93] John R. Ottensmann and Michael E. Gleeson. Research: Implementation and testing of a decision support system for public library materials acquisition budgeting. Journal of the American Society for Information Science, 44(2):83–93, March 1993. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OHara:1970:BCC

[O’H70] F. M. O’Hara, Jr. Brief communications: The corners and edges of the precision-recall square. Journal of the American Society for Information Science, 21(2):166, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OHaver:1995:POC

[O’H95] Thomas C. O’Haver. Perspectives on … the Chemist’s workstation: CHEMCONF: an experiment in international online conferencing. Journal of the American Society for Information Science, 46(8):611–613, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oh:1998:RDR

[Oh98] Sam G. Oh. Research: Document representation and retrieval using empirical facts: Evaluation of a pilot system. Journal of the American Society for Information Science, 49(10):920–931, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ojala:1986:PLO

[Oja86] Marydee Ojala. Perspectives on … online searching: Views on end-user searching. Journal of the American Society for Information Science, 37(4):197–203, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olden:1987:OPS

[Old87] Anthony Olden. Opinion papers: Sub-Saharan Africa and the paperless society. Journal of the American Society for Information Science, 38(4):298–304, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See comments [Aiy88, Tia89].

Olden:1988:LER

[Old88] Anthony Olden. Letters to the Editor: Rejoinder to W.O. Aiyepeku’s comments. Journal of the American Society for Information Science, 39(3):220, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Aiy88].

Olden:1991:BRB

[Old91] Anthony Olden. Book review: Translating an international education to a national environment: Papers presented at the international doctoral student conference sponsored by the doctoral guild at the University of Pittsburgh school of library and information science, September 23–25, 1988. Journal of the American Society for Information Science, 42(7):538–539, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olsen:1974:BRB

[Ols74] James L. Olsen, Jr. Book review: Biographic Guide for Editors & Authors, BIOSIS, CAS, Ei., Washington, DC: American Chemical Society, 362 pp. (1974). [$9.50 (1–5). Quantity discount.]. Journal of the American Society for Information Science, 25(6): 381, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olsen:1975:BRB

[Ols75] James L. Olsen, Jr. Book review: How to find out, G. Chandler. Oxford: Pergamon Press, 194 pp. (1974). Journal of the American Society for Information Science, 26(5):307–308, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olsen:1976:BRB

[Ols76] James L. Olsen, Jr. Book review: University science and engineering libraries. Ellis Mount. Westport, CT: Greenwood Press, 214 pp., (1975). (ISBN-0-8371-7955-6). Journal of the American Society for Information Science, 27(4):262, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olsgaard:1989:BRB

[Ols89] John N. Olsgaard. Book review: Education for professional librarians. Journal of the American Society for Information Science, 40(2):134–135, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olsen:1999:MIP

[Ols99] Kai A. Olsen. In memoriam: An international perspective. Journal of the American Society for Information Science, 50(4):290– 291, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ogawa:1989:BCH

[OMY^{+}89] Haruyuki Ogawa, Nobuyuki Midorikawa, Chie Yoshikawa, Ken ichiro Saito, Hiroshi Itsumura, Masatsugo Kaneko, and Emiko Niki. Brief communications: How much of cited conference materials can be found using bibliographic tools? Journal of the American Society for Information Science, 40(5):350–355, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Onuigbo:1982:BCP

[Onu82] Wilson I. B. Onuigbo. Brief communications: Printer’s Devil and reprint requests. Journal of the American Society for Information Science, 33(1):58–59, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oliver:1997:RUC

[OO97] Ron Oliver and Helen Oliver. Research: Using context to promote learning from information-seeking tasks. Journal of the American Society for Information Science, 48(6):519–526, June 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OConnor:1999:RUR

[OOA99] Brian C. O’Connor, Mary K. O’Connor, and June M. Abbas. Research: User reactions as access mechanism: an exploration based on captions for images. Journal of the American Society for Information Science, 50(8):681–697, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oppenheim:1974:LEM

[Opp74] C. Oppenheim. Letters to the Editors: More on Windsor & Windsor. Journal of the American Society for Information Science, 25 (2):134, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oppenheim:1998:LEF

[Opp98] Charles Oppenheim. Letters to the Editor: Fractional counting of multiauthored publications. Journal of the American Society for Information Science, 49(5):482, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oppenheim:1978:HCO

[OR78] Charles Oppenheim and Susan P. Renn. Highly cited old papers and the reasons why they continue to be cited. Journal of the American Society for Information Science, 29(5):225–231, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oravec:2000:BRB

[Ora00] Jo Ann Oravec. Book review: The clock of the long now: Time and responsibility, by Stewart Brand. Journal of the American Society for Information Science, 51(10):966–967, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oromaner:1977:DCP

[Oro77] Mark Oromaner. The diffusion of core publications in American sociology. Journal of the American Society for Information Science, 28(1):34–37, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Osborne:1998:PLI

[Osb98] Larry N. Osborne. Perspectives on … Internet issues/human communications issues: Topic development in USENET newsgroups. Journal of the American Society for Information Science, 49(11):1010–1016, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ottaviani:1994:RFN

[Ott94] J. S. Ottaviani. Research: The fractal nature of relevance: a hypothesis. Journal of the American Society for Information Science, 45(4):263–272, May 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Olafsen:1982:OVR

[OV82] Tore Olafsen and Libena Vokac. Optimal values of recall and precision. Journal of the American Society for Information Science, 33(2):92–96, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oluic-Vukovic:1992:RJP

[OV92] Vesna Oluić-Vuković. Research: Journal productivity distribution: Quantitative study of dynamic behavior. Journal of the American Society for Information Science, 43(6):412–421, July 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oluic-Vukovic:1993:LEW

[OV93] Vesna Oluić-Vuković. Letters to the Editor: Why has Bradford’s Law been an elusive phenomenon so far? Journal of the American Society for Information Science, 44(3):182–183, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oluic-Vukovic:1995:LED

[OV95] Vesna Oluić-Vuković. Letters to the Editor: On the dynamic behavior of Bradford’s Law. Journal of the American Society for Information Science, 46(10):798–799, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oluic-Vukovic:1997:BDC

[OV97] Vesna Oluić-Vuković. Bradford’s distribution: From the classical bibliometric “law” to the more general stochastic models. Journal of the American Society for Information Science, 48(9):833–842, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Oluic-Vukovic:1998:RSG

[OV98] Vesna Oluić-Vuković. Research: Simon’s generating mechanism: Consequences and their correspondence to empirical facts. Journal of the American Society for Information Science, 49(10):867– 880, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

OKeefe:1999:RMS

[OWF99] Karen M. O’Keefe, Barbara M. Wildemuth, and Charles P. Friedman. Research: Medical students’ confidence judgments using a factual database and personal memory: a comparison. Journal of the American Society for Information Science, 50(8):698–708, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ozog:1979:BCV

[Ozo79a] Stanley Ozog. Brief communications: On the value of information. Journal of the American Society for Information Science, 30 (5):310–315, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ozog:1979:LBC

[Ozo79b] Stanley Ozog. The logics and boundary concepts of information acquisition. Journal of the American Society for Information Science, 30(6):364–367, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ozog:1980:LEV

[Ozo80] Stanley Ozog. Letters to the Editor: On the value of information — author’s reply. Journal of the American Society for Information Science, 31(4):304, August 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Ozo79a].

Pachevsky:1982:PIS

[Pac82] Talmon Pachevsky. Problems of information services with respect to integration of the sciences. Journal of the American Society for Information Science, 33(3):115–123, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paisley:1971:OPI

[Pai71] William Paisley. Opinion papers: Improving a field-based “Ericlike” information system. Journal of the American Society for Information Science, 22(6):399–408, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paisley:1990:POA

[Pai90] William Paisley. Perspectives on … author cocitation analysis: An oasis where many trails cross: the improbable cocitation networks of a multidiscipline. Journal of the American Society for Information Science, 41(6):459–468, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paijmans:1993:RCD

[Pai93a] Hans Paijmans. Research: Comparing the document representations of two IR-systems: CLARIT and TOPIC. Journal of the American Society for Information Science, 44(7):383–392, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paisley:1993:POK

[Pai93b] William Paisley. Perspectives on … knowledge utilization: Knowledge utilization: the role of new communication technologies. Journal of the American Society for Information Science, 44(4):222–234, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paice:1996:RME

[Pai96] Chris D. Paice. Research: Method for evaluation of stemming algorithms based on error counting. Journal of the American Society for Information Science, 47(8):632–649, August 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Palmer:1987:PLF

[Pal87] Raymond A. Palmer. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Effect of federal programs on health sciences libraries. Journal of the American Society for Information Science, 38(1):40–47, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Palmer:1999:ASI

[Pal99a] Carole L. Palmer. Aligning studies of information seeking and use with domain analysis. Journal of the American Society for Information Science, 50(12):1139–1140, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Palmer:1999:RSS

[Pal99b] Carole L. Palmer. Research: Structures and strategies of interdisciplinary science. Journal of the American Society for Information Science, 50(3):242–253, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pao:1978:ATA

[Pao78] Miranda Lee Pao. Automatic text analysis based on transition phenomena of word occurrences. Journal of the American Society for Information Science, 29(3):121–124, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pao:1982:CCM

[Pao82] Miranda Lee Pao. Collaboration in computational musicology. Journal of the American Society for Information Science, 33(1): 38–43, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pao:1986:EEL

[Pao86a] Miranda Lee Pao. An empirical examination of Lotka’s Law. Journal of the American Society for Information Science, 37(1): 26–33, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pao:1986:REE

[Pao86b] Miranda Lee Pao. Research: An empirical examination of Lotka’s Law. Journal of the American Society for Information Science, 37(1):26–33, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Papadakis:1984:LEI

[Pap84] Emmanuel P. Papadakis. Letters to the Editors: Information retrieval and the loss of information. Journal of the American Society for Information Science, 35(1):56, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Papadakis:1988:LEC

[Pap88] Emmanuel P. Papadakis. Letters to the Editor: Color mismatching to invigorate audiences at slide presentations. Journal of the American Society for Information Science, 39(3):220–221, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Papadakis:1993:LEI

[Pap93] Emmanuel P. Papadakis. Letters to the Editor: The impact of proceedings upon authors and journals. Journal of the American Society for Information Science, 44(3):183, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Papadakis:1996:LED

[Pap96] Emmanuel P. Papadakis. Letters to the Editor: Disinformation from research. Journal of the American Society for Information Science, 47(9):726–727, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Papadakis:1998:LEP

[Pap98] Emmanuel P. Papadakis. Letters to the Editor: The price of information not acted upon. Journal of the American Society for Information Science, 49(11):1045, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Parker:1982:BMMa

[Par82a] Ralph H. Parker. Bibliometric models for management of an information store. I. Differential utility among items. Journal of the American Society for Information Science, 33(3):124–128, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Parker:1982:BMMb

[Par82b] Ralph H. Parker. Bibliometric models for management of an information store. II. Use as a function of age of material. Journal of the American Society for Information Science, 33(3):129–133, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Parker:1982:BMMc

[Par82c] Ralph H. Parker. Bibliometric models for management of an information store. III. Developing an empirical model. Journal of the American Society for Information Science, 33(3):134–138, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paris:1983:BVO

[Par83a] Judith Paris. Basics of Videodisc and Optical Disk Technology. Journal of the American Society for Information Science, 34(6): 408–413, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Parker:1983:TTD

[Par83b] Lorraine M. Purgailis Parker. Towards a theory of document learning. Journal of the American Society for Information Science, 34(1):16–21, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Park:1994:RTT

[Par94] Taemin Kim Park. Research: Toward a theory of user-based relevance: a call for a new paradigm of inquiry. Journal of the American Society for Information Science, 45(3):135–141, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Parker:1997:BRB

[Par97] Bret H. Parker. Book review: Computer networks: a systems approach. Journal of the American Society for Information Science, 48(9):861–862, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Park:2000:UUP

[Par00] Soyeon Park. Usability, user preferences, effectiveness, and user behaviors when searching individual and integrated full-text databases: implications for digital libraries. Journal of the American Society for Information Science, 51(5):456–468, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paskoff:1990:BRB

[Pas90] Beth M. Paskoff. Book review: The library microcomputer environment: Management issues. Journal of the American Society for Information Science, 41(2):150, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pathan:1981:LEM

[Pat81] A. Majid Pathan. Letters to the Editor: MEDLINE: an introduction to online searching. Journal of the American Society for Information Science, 32(3):235, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Patterson:1985:BRB

[Pat85] Charles D. Patterson. Book review: Evaluation of reference services. Bill Katz and Ruth A. Fraley, Eds. New York: Haworth Press; 1984: 334 pp. $29.95 (hard). Journal of the American Society for Information Science, 36(6):418–419, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Patterson:1990:BRB

[Pat90] Charles D. Patterson. Book review: Education of library and information professionals: Present and future prospects. Journal of the American Society for Information Science, 41(2):148, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Proper:1999:RWI

[PB99] H. A. Proper and P. D. Bruza. Research: What is information discovery about? Journal of the American Society for Information Science, 50(9):737–750, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paolini:2000:VMT

[PBL^{+}00] Paolo Paolini, Thimoty Barbieri, Paolo Loiudice, Francesca Alonzo, Marco Zanti, and Giuliano Gaia. Visiting a museum together: How to share a visit to a virtual world. Journal of the American Society for Information Science, 51(1):33–38, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1995:RCR

[PBvR95] H. P. F. Peters, R. R. Braam, and A. F. J. van Raan. Research: Cognitive resemblance and citation relations in chemical engineering publications. Journal of the American Society for Information Science, 46(1):9–21, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peck:1985:BRB

[Pec85] Gail Peck. Book review: Accounting for librarians and other not-for-profit managers, G. Stevenson Smith. Chicago: American Library Association; 1984: 470 pp. Journal of the American Society for Information Science, 36(1):57, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peek:1994:PLE

[Pee94] Robin P. Peek. Perspectives on … electronic publishing: Where is publishing going? A perspective on change. Journal of the American Society for Information Science, 45(10):730–736, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peek:1996:BRB

[Pee96a] Robin Peek. Book review: Out in the cold — Academic boycotts and the isolation of South Africa. Journal of the American Society for Information Science, 47(5):399, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peek:1996:EPG

[Pee96b] Robin Peek. Electronic publishing grows up. Journal of the American Society for Information Science, 47(9):665–668, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pelton:1993:BRB

[Pel93] Joseph Pelton. Book review: The emerging worldwide electronic university. Journal of the American Society for Information Science, 44(5):300–301, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Penner:1970:PCU

[Pen70] Rudolf J. Penner. The practice of charging users for information services: a state of the art report. Journal of the American Society for Information Science, 21(1):67–74, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Penniman:1993:BRB

[Pen93] W. David Penniman. Book review: The measurement and evaluation of library service (2nd ed.). Journal of the American Society for Information Science, 44(5):300, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pepinsky:1974:MSR

[Pep74] Harold B. Pepinsky. A metalanguage for systematic research on human communication via natural language. Journal of the American Society for Information Science, 25(1):59–69, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peritz:1982:MCC

[Per82] Bluma C. Peritz. Matched Case-Control Studies in Citation Analysis. Journal of the American Society for Information Science, 33 (5):333–337, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peritz:1983:BCN

[Per83] Bluma C. Peritz. Brief communication: A note on “scholarliness” and “impact”. Journal of the American Society for Information Science, 34(5):360–362, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Perry:1985:DHT

[Per85] J. A. Perry. The Dillion Hypothesis of Titular Colonicity: an empirical test from the ecological sciences. Journal of the American Society for Information Science, 36(4):251–258, July 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peritz:1992:OPO

[Per92] B. C. Peritz. Opinion papers: On the objectives of citation analysis: Problems of theory and method. Journal of the American Society for Information Science, 43(6):448–451, July 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Persson:1994:RIB

[Per94] Olle Persson. Research: The intellectual base and research fronts of JASIS 1986–1990. Journal of the American Society for Information Science, 45(1):31–38, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Petroski:1994:PLR

[Pet94] S. Paul Petroski. Perspectives on … the redesign/reengineering of an information services division of a major health science institution: Redeploying resources. Journal of the American Society for Information Science, 45(5):335–338, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1996:BRBc

[Pet96a] Thomas A. Peters. Book review: Computerization and controversy: Value conflicts and social choices (2nd ed.). Journal of the American Society for Information Science, 47(12):953–955, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1996:BRBa

[Pet96b] Thomas A. Peters. Book review: Human factors in information systems: Emerging theoretical bases. Journal of the American Society for Information Science, 47(8):655–656, August 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1996:BRBb

[Pet96c] Thomas A. Peters. Book review: Human factors in information systems: Emerging theoretical bases. Journal of the American Society for Information Science, 47(8):655–656, August 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1998:BRBa

[Pet98a] Thomas A. Peters. Book review: Culture of the Internet. Journal of the American Society for Information Science, 49(2):188–190, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1998:BRBb

[Pet98b] Thomas A. Peters. Book review: Privacy on the line: the politics of wiretapping and encryption. Journal of the American Society for Information Science, 49(14):1339, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1999:BRB

[Pet99] Thomas A. Peters. Book review: Ethics, information and technology: Readings, edited by Richard N. Stichler and Robert Hauptman. Journal of the American Society for Information Science, 50(8):726–728, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Philip:1992:ERL

[Phi92] G. Philip. European research letters: Speech technology developments in Europe. Journal of the American Society for Information Science, 43(1):92–95, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Picard:1986:BRB

[Pic86] Robert G. Picard. Book review: Reporters under fire: U.S. media coverage of conflicts in Lebanon and Central America. Journal of the American Society for Information Science, 37(4):276, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pichappan:1996:BCL

[Pic96] P. Pichappan. Brief communications: Levels of citation relations between papers. Journal of the American Society for Information Science, 47(8):650–652, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pierce:1992:ROM

[Pie92] Sydney J. Pierce. Research: On the origin and meaning of bibliometric indicators: Journals in the social sciences, 1886–1985. Journal of the American Society for Information Science, 43(7): 477–487, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pierce:1999:RBC

[Pie99] Sydney J. Pierce. Research: Boundary crossing in research literatures as a means of interdisciplinary information transfer. Journal of the American Society for Information Science, 50(3):271–279, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Piternick:1975:DSJ

[Pit75] Anne B. Piternick. Derivation of a sample of journal issues for tests of availability and use. Journal of the American Society for Information Science, 26(5):269–270, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Piternick:1982:LES

[Pit82] Anne B. Piternick. Letter to the Editors: Standardization of journal titles in databases. Journal of the American Society for Information Science, 33(2):105, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Phenix:1987:BRB

[PJ87] Katharine Phenix and Jorg R. Jemelka. Book review: Computer graphics glossary. Journal of the American Society for Information Science, 38(3):218–219, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Parker:1990:RDO

[PJ90] Lorraine M. Purgailis Parker and Robert E. Johnson. Research: Does order of presentation affect users’ judgment of documents? Journal of the American Society for Information Science, 41 (7):493–494, October 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Praunlich:1978:BDN

[PK78] Peter Praunlich and Michael Kroll. Bradford’s distribution: a new formulation. Journal of the American Society for Information Science, 29(2):51–55, March 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Palmquist:2000:RCS

[PK00] Ruth A. Palmquist and Kyung-Sun Kim. Research: Cognitive style and on-line database search experience as predictors of Web search performance. Journal of the American Society for Information Science, 51(6):558–566, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Penniman:1974:BCF

[PKK74] W. David Penniman, Richard E. Krohn, and Gabor J. Kovacs. Brief communications: A framework for the study of emerging network technology. Journal of the American Society for Information Science, 25(6):378–380, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pratt:1979:LE

[PLM^{+}79] Allan D. Pratt, Donald Leatherdale, Charles T. Meadow, Charles H. Davis, and Debora Shaw. Letters to the Editor. Journal of the American Society for Information Science, 30(5):317– 319, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Poincot:2000:MIS

[PLM00] Philippe Poinçot, Soizick Lesteven, and Fionn Murtagh. Maps of information spaces: Assessments from astronomy. Journal of the American Society for Information Science, 51(12):1081–1089, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Paterno:2000:ELA

[PM00] F. Paternò and C. Mancini. Effective levels of adaptation to different types of users in interactive museum systems. Journal of the American Society for Information Science, 51(1):5–13, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pinski:1979:SPL

[PN79] Gabriel Pinski and Francis Narin. Structure of the psychological literature. Journal of the American Society for Information Science, 30(3):161–168, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pearce:1996:TED

[PN96] Claudia Pearce and Charles Nicholas. TELLTALE: Experiments in a dynamic hypertext environment for degraded and multilingual data. Journal of the American Society for Information Science, 47(4):263–275, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Plaunt:1998:RAB

[PN98] Christian Plaunt and Barbara A. Norgard. Research: An association-based method for automatic indexing with a controlled vocabulary. Journal of the American Society for Information Science, 49(10):888–902, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pommer:1974:LEP

[Pom74] Alfred M. Pommer. Letter to the Editors: Protection against photocopying by printing black on red. Journal of the American Society for Information Science, 25(4):274, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pommer:1975:LE

[Pom75] Alfred M. Pommer. Letter to the Editors. Journal of the American Society for Information Science, 26(6):356, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pope:1975:BLP

[Pop75] Andrew Pope. Bradford’s Law and the periodical literature of information science. Journal of the American Society for Information Science, 26(4):207–213, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pouris:1989:REA

[Pou89] Anastassios Pouris. Research: Evaluating academic science institutions in South Africa. Journal of the American Society for Information Science, 40(4):269–272, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Poyer:1982:SCI

[Poy82] Robert K. Poyer. Science Citation Index’s coverage of the preclinical science literature. Journal of the American Society for Information Science, 33(5):317–320, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peek:1998:PLI

[PPP98] Robin Peek, Jeffrey Pomerantz, and Stephen Paling. Perspective on … Internet issues/publishing issues: The traditional scholarly journal publishers legitimize the Web. Journal of the American Society for Information Science, 49(11):983–989, ???? 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Perry:1998:RSC

[PR98] Claudia A. Perry and Ronald E. Rice. Research: Scholarly communication in developmental dyslexia: Influence of network structure on change in a hybrid problem area. Journal of the American Society for Information Science, 49(2):151–168, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pratt:1977:MCC

[Pra77] Allan D. Pratt. A measure of class concentration in bibliometrics. Journal of the American Society for Information Science, 28(5): 285–292, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Prabha:1983:BCS

[Pra83] Chandra G. Prabha. Brief communication: Some aspects of citation behavior: a pilot study in business administration. Journal of the American Society for Information Science, 34(3):202–206, May 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pratt:1990:BRB

[Pra90] Gordon Pratt. Book review: Communicating ideas — the crisis of publishing in a post-industrial society. Journal of the American Society for Information Science, 41(1):67–69, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Prescrel:1970:BRB

[Pre70] Barbara M. Prescrel. Book review: 4/70-4R Advances in Librarianship. Volume 1. 1970. Melvin J. Voight, Editor. Academic Press, New York and London. 294 pp. Journal of the American Society for Information Science, 21(4):306–307, July 1970. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Preschel:1971:BRN

[Pre71a] Barbara H. Preschel. Book review: Nelson, Carnot E., and Pollock, Donald K., editors. Communication Among Scientists and Engineers. Lexington, Mass., Heath, 1970. 346 pp. Journal of the American Society for Information Science, 22(3):231, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Preschel:1971:BRB

[Pre71b] Barbara M. Preschel. Book review: The Role of the Library in Relation to Other Information Activities; a State-Of-The-Art Review. Painter, Ann F. Washington, D.C., U.S. Army, Office of the Chief of Engineers, 1968. 85 pp. (TISA Project Report, No. 23). Journal of the American Society for Information Science, 22(3):231–232, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Preschel:1971:BRS

[Pre71c] Barbara M. Preschel. Book review: Smith, Gloria L., and Meyer, Robert S., editors. Library Use of Computers, an Introduction. New York, Special Libraries Association, 1969. (SLA Monograph, No. 3) 114 pp. Journal of the American Society for Information Science, 22(3):231, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Preschel:1988:PLE

[Pre88] Barbara M. Preschel. Perspectives on … education of the information professional: New dimensions, new directions: Education of the information professional: What employers want. Journal of the American Society for Information Science, 39(5):358–361, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Price:1972:LEC

[Pri72] Douglas S. Price. Letters to the Editor: Caveat Presentor. Journal of the American Society for Information Science, 23(2):137– 139, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Price:1983:OPP

[Pri83] Douglas S. Price. Opinion paper: Possible impact of electronic publishing on abstracting and indexing. Journal of the American Society for Information Science, 34(4):288, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Packer:1979:ISC

[PS79] Katherine H. Packer and Dagobert Soergel. The importance of SDI for current awareness in fields with severe scatter of information. Journal of the American Society for Information Science, 30(3):125–135, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Prade:1987:RFR

[PT87] Henri Prade and Claudette Testemale. Research: Fuzzy relational databases: Representational issues and reduction using similarity measures. Journal of the American Society for Information Science, 38(2):118–126, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Piternick:1983:LEE

[PTH83] Anne B. Piternick, Murray Turoff, and Starr Roxanne Hiltz. Letters to the Editor: The electronic journal. Journal of the American Society for Information Science, 34(4):297–298, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pullen:1974:LED

[Pul74] Keats A. Pullen, Jr. Letters to the Editors: A decision theory view of the information retrieval situation. Journal of the American Society for Information Science, 25(1):73–74, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peters:1994:RDC

[PvR94] H. P. F. Peters and A. F. J. van Raan. Research: On determinants of citation scores: a case study in chemical engineering. Journal of the American Society for Information Science, 45(1):39–49, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Penrod:1988:PLI

[PW88] James I. Penrod and Ronald Witte. Perspectives on … integrated academic information management systems (IAIMS): IAIMS infrastructure: Technological base. Journal of the American Society for Information Science, 39(2):118–125, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pao:1989:RRE

[PW89] Miranda Lee Pao and Dennis B. Worthen. Research: Retrieval effectiveness by semantic and citation searching. Journal of the American Society for Information Science, 40(4):226–235, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Peat:1991:RLT

[PW91] Helen J. Peat and Peter Willett. Research: The limitations of term co-occurrence data for query expansion in document retrieval systems. Journal of the American Society for Information Science, 42(5):378–383, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

oopovic:1992:RES

[PW92] Mirko Popovič and Peter Willett. Research: The effectiveness of stemming for natural-language access to Slovene textual data. Journal of the American Society for Information Science, 43(5): 384–390, June 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pollock:1983:CCS

[PZ83] J. J. Pollock and A. Zamora. Collection and characterization of spelling errors in scientific and scholarly text. Journal of the American Society for Information Science, 34(1):51–58, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Pollock:1984:SDD

[PZ84] Joseph J. Pollock and Antonio Zamora. System design for detection and correction of spelling errors in scientific and scholarly text. Journal of the American Society for Information Science, 35(2):104–109, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Persin:1996:RFD

[PZSD96] Michael Persin, Justin Zobel, and Ron Sacks-Davis. Research: Filtered document retrieval with frequency-sorted indexes. Journal of the American Society for Information Science, 47(10):749– 764, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Qin:2000:LER

[Qin00a] Jian Qin. Letter to the Editor (Reply): Incremental benefit of human indexing. Journal of the American Society for Information Science, 51(10):968, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Qin:2000:RSS

[Qin00b] Jian Qin. Research: Semantic similarities between a keyword database and a controlled vocabulary database: an investigation in the antibiotic resistance literature. Journal of the American Society for Information Science, 51(2):166–180, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Qiu:1993:RMM

[Qiu93] Liwen Qiu. Research: Markov models of search state patterns in a hypertext information retrieval system. Journal of the American Society for Information Science, 44(7):413–427, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Qin:1997:RTL

[QLA97] Jian Qin, F. W. Lancaster, and Bryce Allen. Research: Types and levels of collaboration in interdisciplinary research in the sciences. Journal of the American Society for Information Science, 48(10):893–916, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Quandt:1996:RSM

[Qua96] Richard E. Quandt. Research: Simulation model for journal subscription by libraries. Journal of the American Society for Information Science, 47(8):610–617, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Radecki:1982:PAI

[Rad82a] Tadeusz Radecki. A probabilistic approach to information retrieval in systems with boolean search request formulations. Journal of the American Society for Information Science, 33(6):365– 370, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Radecki:1982:SMB

[Rad82b] Tadeusz Radecki. Similarity measures for boolean search request formulations. Journal of the American Society for Information Science, 33(1):8–17, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rada:1989:PLH

[Rad89] Roy Rada. Perspectives on … hypertext: Writing and reading hypertext: an overview. Journal of the American Society for Information Science, 40(3):164–171, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rafsnider:1975:CBU

[Raf75] Jean Rafsnider. Complementarity between a use study and a citation study in considering publication practices in the physical and social sciences. Journal of the American Society for Information Science, 26(6):344–345, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Raghavan:1996:BRB

[Rag96] Vijay Raghavan. Book review: Measurement in information science. Journal of the American Society for Information Science, 47(3):257–258, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rao:1980:DSP

[Rao80] I. K. Ravichandra Rao. The distribution of scientific productivity and social change. Journal of the American Society for Information Science, 31(2):111–122, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Raphael:2000:BRB

[Rap00] Randy Raphael. Book review: Data on the Web: From relations to semistructured data and XML, by Serge Abiteboul, Peter Buneman, and Dan Suciu. Journal of the American Society for Information Science, 51(12):1050–1052, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rastogi:1980:RBD

[Ras80] Kunj B. Rastogi. Retrieval behavior of derived truncated search keys for a large on-line bibliographic file. Journal of the American Society for Information Science, 31(2):84–88, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rashid:1990:RBA

[Ras90] Haseeb F. Rashid. Research: Book availability as a performance measure of a library: an analysis of the effectiveness of a health sciences library. Journal of the American Society for Information Science, 41(7):501–507, October 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rasmussen:1999:MMR

[Ras99] Edie Rasmussen. In memoriam: In memoriam Robert R. Korfhage. Journal of the American Society for Information Science, 50(4):288, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rawlins:1993:POD

[Raw93] Gregory J. E. Rawlins. Perspectives on … digital libraries: Publishing over the next decade. Journal of the American Society for Information Science, 44(8):474–479, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rayward:1994:RVX

[Ray94] W. Boyd Rayward. Research: Visions of Xanadu: Paul Otlet (1868–1944) and hypertext. Journal of the American Society for Information Science, 45(4):235–250, May 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rayward:1997:POH

[Ray97] W. Boyd Rayward. Paul Otlet and his successors: The origins of information science and the International Institute of Bibliography/International Federation for Information and Documentation (FID). Journal of the American Society for Information Science, 48(4):289–300, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rayward:1999:RHG

[Ray99] W. Boyd Rayward. Research: H. G. Wells’s idea of a World Brain: a critical reassessment. Journal of the American Society for Information Science, 50(7):557–573, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rzasa:1972:MEU

[RB72] Philip V. Rzasa and Norman R. Baker. Measures of effectiveness for a university library. Journal of the American Society for Information Science, 23(4):248–253, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rice:1983:UCM

[RB83] Ronald E. Rice and Christine L. Borgman. The use of computermonitored data in information science and communication research. Journal of the American Society for Information Science, 34(4):247–256, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rada:1989:RRD

[RB89] Roy Rada and Ellen Bicknell. Research: Ranking documents with a thesaurus. Journal of the American Society for Information Science, 40(5):304–310, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Regazzi:1980:LSD

[RBR80] John J. Regazzi, Bruce Bennion, and Susan Roberts. On-line systems of disciplines and specialty areas in science and technology. Journal of the American Society for Information Science, 31(3): 161–170, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robertson:1974:LED

[RC74] S. E. Robertson and William S. Cooper. Letters to the Editors: In Defence of Relevance. Journal of the American Society for Information Science, 25(3):208–209, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rajashekar:1995:RCA

[RC95] T. B. Rajashekar and W. Bruce Croft. Research: Combining automatic and manual index representations in probabilistic retrieval. Journal of the American Society for Information Science, 46(4):272–283, May 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rogers:1994:PPS

[RCJS94] Everett M. Rogers, Lori Collins-Jarvis, and Joseph Schmitz. The PEN project in Santa Monica: Interactive communication, equality, and political action. Journal of the American Society for Information Science, 45(6):401–410, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ramsey:1999:POI

[RCZS99] Marshall C. Ramsey, Hsinchun Chen, Bin Zhu, and Bruce R. Schatz. Perspectives on … visual information retrieval interfaces: A collection of visual thesauri for browsing large collections of geographic images. Journal of the American Society for Information Science, 50(9):826–834, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Raghavan:1998:I

[RDS98] Vijay V. Raghavan, Jitender S. Deogun, and Hayri Sever. Introduction. Journal of the American Society for Information Science, 49(5):397–402, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Regazzi:1988:RPM

[Reg88] John J. Regazzi. Research: Performance measures for information retrieval systems — an experimental approach. Journal of the American Society for Information Science, 39(4):235–251, July 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Reilly:1975:BRB

[Rei75] Kevin D. Reilly. Book review: Library planning and decisionmaking systems. M. Hamburg, R. Clelland, M. Bommer, L. Ramist and R. Whitfield. Cambridge, MA: MIT Press. 274 pp. (1974). Journal of the American Society for Information Science, 26(1):65, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Reilly:1977:BRB

[Rei77] Kevin D. Reilly. Book review: Cannons’ Bibliography of Library Economy, 1876–1920: An Author Index with Citations. Anne H. and Melbourne Jordan, Eds. Metuchen, NJ: Scarecrow Press, 473 pp. (1976). Journal of the American Society for Information Science, 28(6):375, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Reintjes:1984:AMT

[Rei84] J. Francis Reintjes. Application of modern technologies to interlibrary resource-sharing networks. Journal of the American Society for Information Science, 35(1):45–52, January 1984. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Reid:1985:BRB

[Rei85] Marion T. Reid. Book review: Resource sharing and information networks, vol. 1, nos. 1/2-(Fall/Winter 1983–). New York: The Haworth Press, Inc. Annual subscription rates: $33 for individuals; $60 for institutions and libraries. Journal of the American Society for Information Science, 36(4):273–274, July 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Reno:1970:CC

[Ren70] Charles Reno. Computers Are for Concepts. Journal of the American Society for Information Science, 21(4):293–298, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Renfro:1980:OIS

[Ren80] Charles G. Renfro. An online information system for aggregate state and local area economic data. Journal of the American Society for Information Science, 31(5):319–333, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Repo:1989:RVI

[Rep89] Aatto J. Repo. Research: The value of information: Approaches in economics, accounting, and management science. Journal of the American Society for Information Science, 40(2):68–85, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ruocco:1997:RCC

[RF97] Anthony S. Ruocco and Ophir Frieder. Research: Clustering and classification of large document bases in a parallel environment. Journal of the American Society for Information Science, 48(10):932–943, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:2000:SRV

[RF00] Mark Rorvig and Steven Fitzpatrick. Shape recovery: a visual method for evaluation of information retrieval experiments. Journal of the American Society for Information Science, 51(13): 1205–1210, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robbin:1997:RET

[RFK97] Alice Robbin and Lee Frost-Kumpf. Research: Extending theory for user-centered information services: Diagnosing and learning from error in complex statistical data. Journal of the American Society for Information Science, 48(2):96–121, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Richmond:1970:ANE

[RG70] Phyllis A. Richmond and Marcia K. Gill. Accommodation of nonstandard entries in a serial list made by computer. Journal of the American Society for Information Science, 21(4):240–246, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rohde:1997:PLI

[RH97] Nancy Rohde and Nancy Herther. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: Implications of the IIC project for the education of information professionals. Journal of the American Society for Information Science, 48(5):474–475, May 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:1999:POIc

[RH99] Mark Rorvig and Matthias Hemmje. Perspectives on … visual information retrieval interfaces: Conference notes — 1996: Foundations of advanced information visualization for visual information (retrieval) systems. Journal of the American Society for Information Science, 50(9):835–837, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Richmond:1970:LE

[RHH^{+}70] Phyllis A. Richmond, C. R. G. Holmes, Iola Belle Holbrook, C. Secebarth, John V. Howard, Dana L. Roth, and F. M. O’Hara, Jr. Letters to the Editor. Journal of the American Society for Information Science, 21(6):429–430, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Richmond:1970:OPF

[Ric70] Phyllis A. Richmond. Opinion paper: The final report of the comparative systems laboratory: a review. Journal of the American Society for Information Science, 21(2):160–162, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Richmond:1976:CPS

[Ric76] Phyllis A. Richmond. Classification from PRECIS: Some possibilities. Journal of the American Society for Information Science, 27(4):240–247, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Richmond:1977:BRB

[Ric77] Phyllis A. Richmond. Book review: Major Classification Systems: The Dewey Centennial. Papers presented at the Allerton Park Institute, 9–12 November 1972. Kathryn Luther Henderson, Ed. Urbana-Champaign, IL: University of Illinois, Graduate Center of Library Science, 182 pp. 1976. (Allerton Park Institute No. 21; ISBN 0-87845-044-0; LC 76-236331). Journal of the American Society for Information Science, 28(2):126–127, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rice:1987:BRB

[Ric87] James Rice. Book review: Universities, information technology, and academic libraries: the next twenty years. Journal of the American Society for Information Science, 38(6):446–447, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Richards:1988:HNI

[Ric88] Pamela Spence Richards. Historical note: Information science in wartime: Pioneer documentation activities in World War II. Journal of the American Society for Information Science, 39(5): 301–306, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Richardson:1989:BRB

[Ric89] John Richardson, Jr. Book review: Federal information policies in the 1980’s: Conflicts and issues. Journal of the American Society for Information Science, 40(2):140–141, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rice:1995:EWW

[Ric95] Ronald E. Rice. Editorial: Writing wrongs. Journal of the American Society for Information Science, 46(5):322–324, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rice:1997:LER

[Ric97] Ronald E. Rice. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 48(10):964–964b, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rietdyk:1988:PLC

[Rie88] Ron J. Rietdyk. Perspectives on … CD-ROM for information storage and retrieval: Creation and distribution of CD-ROM databases for the library reference desk. Journal of the American Society for Information Science, 39(1):58–62, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robertson:1976:RWS

[RJ76] S. E. Robertson and K. Sparck Jones. Relevance weighting of search terms. Journal of the American Society for Information Science, 27(3):129–146, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Radhakrishnan:1979:LLC

[RK79] T. Radhakrishnan and R. Kernizan. Lotka’s Law and Computer Science Literature. Journal of the American Society for Information Science, 30(1):51–54, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robson:1976:AAP

[RL76] Alan Robson and Janet S. Longman. Automatic aids to profile construction. Journal of the American Society for Information Science, 27(4):213–223, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rieusset-Lemarie:1997:POH

[RL97] Isabelle Rieusset-Lemarié. Paul otlet and his successors: P. Otlet’s Mundaneum and the international perspective in the history of documentation and information science. Journal of the American Society for Information Science, 48(4):301–309, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:1999:POIa

[RL99] Mark Rorvig and Lois F. Lunin. Perspectives on … visual information retrieval interfaces: Introduction and overview: Visualization, retrieval, and knowledge. Journal of the American Society for Information Science, 50(9):790–793, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robb:1997:BCD

[RM97] David J. Robb and Angela McCormick. Brief communications: Decision support for serials deselection and acquisition: a case study. Journal of the American Society for Information Science, 48(3):270–273, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rivera–Morales:1999:PLC

[RM99] Noemí A. Rivera-Morales. Perspective on … copyright and fairuse guidelines for education and libraries: Fair-use guidelines: a selected bibliography. Journal of the American Society for Information Science, 50(14):1353–1357, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rowland:1996:EDE

[RMMS96] Fytton Rowland, Cliff McKnight, Jack Meadows, and Peter Such. ELVYN: the delivery of an electronic version of a journal from the publisher to libraries. Journal of the American Society for Information Science, 47(9):690–700, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Romano:1998:ADD

[RNBV98] Nicholas C. Romano, Jr., Jay F. Nunamaker, Jr., Robert O. Briggs, and Douglas R. Vogel. Architecture, design, and development of an HTML/ JavaScript Web-based Group Support System. Journal of the American Society for Information Science, 49(7):649–667, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ro:1988:REAa

[Ro88a] Jung Soon Ro. Research: An evaluation of the applicability of ranking algorithms to improve the effectiveness of full-text retrieval. I. On the effectiveness of full-text retrieval. Journal of the American Society for Information Science, 39(2):73–78, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ro:1988:REAb

[Ro88b] Jung Soon Ro. Research: An evaluation of the applicability of ranking algorithms to improve the effectiveness of full-text retrieval. II. On the effectiveness of ranking algorithms on full-text retrieval. Journal of the American Society for Information Science, 39(3):147–160, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robertson:1975:EIV

[Rob75] Stephen E. Robertson. Explicit and implicit variables in information retrieval (IR) systems. Journal of the American Society for Information Science, 26(4):214–222, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robertson:1978:NFD

[Rob78] Stephen E. Robertson. On the nature of fuzz: a diatribe. Journal of the American Society for Information Science, 29(6):304–307, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See reply [Cer79].

Robertson:1979:FSR

[Rob79] Stephen E. Robertson. On Fuzzy sets: Reply to Cerny. Journal of the American Society for Information Science, 30(6):357–358, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Cer79].

Robinson:1980:CCA

[Rob80a] Barbara M. Robinson. Cooperation and competition among library networks. Journal of the American Society for Information Science, 31(6):413–424, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robinson:1980:ETC

[Rob80b] Jo Robinson. Education and training for computer-based reference services: a case study. Journal of the American Society for Information Science, 31(2):97–104, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robinson:1991:BCA

[Rob91] Michael D. Robinson. Brief communications: Applied bibliometrics: Using citation analysis in the journal submission process. Journal of the American Society for Information Science, 42(4): 308–310, May 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robbin:1996:BRB

[Rob96a] Alice Robbin. Book review: Telecommunications politics: Ownership and control of the information highway in developing countries. Journal of the American Society for Information Science, 47(8):653–654A, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Roberts:1996:HPS

[Rob96b] Judith M. Roberts. Historical perspectives: The story of distance education: a practitioner’s perspective. Journal of the American Society for Information Science, 47(11):811–816, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robertson:1998:MCW

[Rob98] Stephen Robertson. In memoriam: Cyril W. Cleverdon. Journal of the American Society for Information Science, 49(10):866, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robins:2000:RSF

[Rob00] David Robins. Research: Shifts of focus on various aspects of user information problems during interactive information retrieval. Journal of the American Society for Information Science, 51(10): 913–928, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Roderer:1993:MLB

[Rod93] Nancy Roderer. In memoriam: Laurence B. Heilprin, 1906–1993. Journal of the American Society for Information Science, 44(10): 555–556, December 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rodgers:1995:PLM

[Rod95] R. P. Channing Rodgers. Perspectives on … medical informatics: Information technology in health care: Automated retrieval from multiple disparate information sources: the World Wide Web and the NLM’s Sourcerer Project. Journal of the American Society for Information Science, 46(10):755–764, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rollins:1987:BRB

[Rol87] Debra C. Rollins. Book review: The story behind the word. Journal of the American Society for Information Science, 38(4):313, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Romkey:1991:PLI

[Rom91] John Romkey. Perspectives on … imaging. Advanced applications: Whither cyberspace? Journal of the American Society for Information Science, 42(8):618–620, September 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:1990:RSS

[Ror90] Mark E. Rorvig. Research: The simple scalability of documents. Journal of the American Society for Information Science, 41(8): 590–598, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:1993:RMA

[Ror93] Mark E. Rorvig. Research: A method for automatically abstracting visual documents. Journal of the American Society for Information Science, 44(1):40–56, January 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:1999:RVE

[Ror99a] Mark Rorvig. Research: A visual exploration of the orderliness of TREC relevance judgments. Journal of the American Society for Information Science, 50(8):652–660, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:1999:RIS

[Ror99b] Mark Rorvig. Research: Images of similarity: a visual exploration of optimal similarity metrics and scaling properties of TREC topic-document sets. Journal of the American Society for Information Science, 50(8):639–651, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rosenberg:1971:CET

[Ros71a] Victor Rosenberg. Comparative evaluation of two indexing methods using judges. Journal of the American Society for Information Science, 22(4):251–259, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rosenberg:1971:SSM

[Ros71b] Victor Rosenberg. A study of statistical measures for predicting terms used to index documents. Journal of the American Society for Information Science, 22(1):41–50, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rosenberq:1973:TMU

[Ros73] Victor Rosenberq. A Technique for Monitoring User Behavior at the Computer Terminal Interface. Journal of the American Society for Information Science, 24(1):71, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rosenberg:1974:OPS

[Ros74] Victor Rosenberg. Opinion paper. The scientific premises of information science. Journal of the American Society for Information Science, 25(4):263–269, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ross:1976:CAL

[Ros76] Ian C. Ross. A cataloger’s authority list maintenance system. Journal of the American Society for Information Science, 27(4): 224–229, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ross:1979:BCS

[Ros79] Johanna C. Ross. Brief communications: Searching the chemical literature via three on-line vendors: a comparison. Journal of the American Society for Information Science, 30(2):103–106, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rosenberg:1982:IPD

[Ros82a] Victor Rosenberg. Information policies of developing countries: The case of Brazil. Journal of the American Society for Information Science, 33(4):203–207, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ross:1982:LES

[Ros82b] Johanna Ross. Letter to the Editors: Searching the chemical literature. Journal of the American Society for Information Science, 33(2):105–106, March 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rose:1983:ODD

[Ros83] Denis A. Rose. Optical disk for digital storage and retrieval systems. Journal of the American Society for Information Science, 34(6):434–438, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Roth:1979:LES

[Rot79] Dana Lincoln Roth. Letter to the Editors: Searching the chemical literature. Journal of the American Society for Information Science, 30(6):372, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rothman:1981:LEO

[Rot81] John Rothman. Letters to the Editor: Online searching and paperless publication. Journal of the American Society for Information Science, 32(1):77, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See reply [Lan81a].

Rouse:1974:CDP

[Rou74a] William B. Rouse. Circulation dynamics: a planning model. Journal of the American Society for Information Science, 25(6): 358–363, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1974:OSA

[Rou74b] William B. Rouse. Optimal selection of acquisition sources. Journal of the American Society for Information Science, 25(4):227– 231, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1975:ORA

[Rou75] William B. Rouse. Optimal resource allocation in library systems. Journal of the American Society for Information Science, 26(3): 157–165, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1976:BRB

[Rou76a] William B. Rouse. Book review: Applications of operations research models to libraries: A Case Study of the Use of Monographs in the Francis A. Countway Library of Medicine, Harvard University. Ching-chih Chen. Cambridge, MA: MIT Press (1976). Journal of the American Society for Information Science, 27(5): 349, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1976:LNM

[Rou76b] William B. Rouse. A library network model. Journal of the American Society for Information Science, 27(2):88–99, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1979:TMM

[Rou79] William B. Rouse. Tutorial: Mathematical modeling of library systems. Journal of the American Society for Information Science, 30(4):181–192, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rousseau:1988:RFS

[Rou88] Ronald Rousseau. Research: The Fussler sampling technique. Journal of the American Society for Information Science, 39(6): 421–422, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Roundtree:1989:BRBb

[Rou89a] Lynn Roundtree. Book review: MARC for archives and manuscripts: a compendium of practice. Journal of the American Society for Information Science, 40(2):139, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Roundtree:1989:BRBa

[Rou89b] Lynn Roundtree. Book review: MARC for archives and manuscripts: The AMC format. Journal of the American Society for Information Science, 40(2):139, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rousseau:1990:RRB

[Rou90] Ronald Rousseau. Research: Relations between continuous versions of bibliometric laws. Journal of the American Society for Information Science, 41(3):197–203, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rousseau:1992:RBR

[Rou92a] Ronald Rousseau. Research: Breakdown of the robustness property of Lotka’s Law: the case of adjusted counts for multiauthorship attribution. Journal of the American Society for Information Science, 43(10):645–647, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rousseau:1992:RCD

[Rou92b] Ronald Rousseau. Research: Concentration and diversity of availability and use in information systems: a positive reinforcement model. Journal of the American Society for Information Science, 43(5):391–395, June 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1998:CSC

[Rou98] William B. Rouse. Computer support of collaborative planning: an applications report. Journal of the American Society for Information Science, 49(9):832–839, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rowley:1984:LEP

[Row84] J. Rowley. Letters to the Editors: Practice and research in JASIS. Journal of the American Society for Information Science, 35(2): 135, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Royce:1989:BRB

[Roy89] Bert Royce. Book review: Database design for information retrieval. Journal of the American Society for Information Science, 40(5):361, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Roy:1990:BRB

[Roy90a] Loriene Roy. Book review: Success in answering reference questions: Two studies. Journal of the American Society for Information Science, 41(1):69–70, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Royce:1990:BRB

[Roy90b] Bert Royce. Book review: The conceptual foundations of descriptive cataloging. Journal of the American Society for Information Science, 41(3):210–211, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1971:NMB

[RR71] James E. Rush and Phillip M. Russo. A note on the mathematical basis of the SLIC index. Journal of the American Society for Information Science, 22(2):123–125, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1977:AIC

[RR77] William B. Rouse and Sandra H. Rouse. Assessing the impact of computer technology on the performance of interlibrary loan networks. Journal of the American Society for Information Science, 28(2):79–88, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rouse:1978:EPU

[RR78] William B. Rouse and Sandra H. Rouse. The effect of parameter uncertainties on the predictions of a library network model. Journal of the American Society for Information Science, 29(4): 180–186, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rees:1991:PLI

[RRB91] Warren Rees, Nancy Freeman Rohde, and Richard Bolan. Perspectives on … integrated information centers within academic environments: Legal issues for an Integrated Information Center. Journal of the American Society for Information Science, 42(2): 132–136, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1974:JDD

[RSK74] Barbara Rush, Sam Steinberg, and Donald H. Kraft. Journal disposition decision policies. Journal of the American Society for Information Science, 25(4):213–217, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ruge:1991:REE

[RSW91] Gerda Ruge, Christoph Schwarz, and Amy J. Warner. Research: Effectiveness and efficiency in natural language processing for large amounts of text. Journal of the American Society for Information Science, 42(6):450–456, July 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1971:AAI

[RSZ71] J. E. Rush, R. Salvador, and A. Zamora. Automatic abstracting and indexing. II. Production of indicative abstracts by application of contextual inference and syntactic coherence criteria. Journal of the American Society for Information Science, 22(4): 260–274, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Robinson:1981:ILE

[RT81] Earl J. Robinson and Stephen J. Turner. Improving library effectiveness: a proposal for applying fuzzy set concepts in the management of large collections. Journal of the American Society for Information Science, 32(6):458–462, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1971:LEI

[RTH71] James E. Rush, J. Tocatlian, and John G. Helmkamp. Letters to the Editors: Information science terminology. Journal of the American Society for Information Science, 22(2):141–142, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rorvig:1999:POIb

[RTM99] M. E. Rorvig, C. H. Turner, and J. Moncada. Perspectives on … visual information retrieval interfaces: The NASA Image Collection Visual Thesaurus. Journal of the American Society for Information Science, 50(9):794–798, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ruben:1992:RCI

[Rub92] Brent D. Ruben. Research: The communication-information relationship in system-theoretic perspective. Journal of the American Society for Information Science, 43(1):15–27, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rumble:1990:BRB

[Rum90] John Rumble, Jr. Book review: Data base: Structured techniques for design, performance, and management. Journal of the American Society for Information Science, 41(2):154, March 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1970:AFA

[Rus70a] James E. Rush. Administrative and financial aspects of computers in education. Journal of the American Society for Information Science, 21(4):299–302, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1970:EIS

[Rus70b] James E. Rush. Education and information science — a symposium. Journal of the American Society for Information Science, 21(4):265, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1970:F

[Rus70c] James E. Rush. Foreword. Journal of the American Society for Information Science, 21(4):266, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1977:BRB

[Rus77] James E. Rush. Book review: Serials Automation in the United States: A Bibliographic History. Gary M. Pitkin. Metuchen, NJ: Scarecrow Press, pp. 148 (1976). Journal of the American Society for Information Science, 28(6):375–376, November 1977. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1985:BRB

[Rus85] James E. Rush. Book review: Document delivery in the United States. Washington, DC: Council on Library Resources, Inc.; 1983: 78 pp. (A report prepared by Information Systems Consultants, Inc.) available from the council for $5.00 (prepaid only); 1785 Massachusetts Avenue N.W., Washington, DC 20036. Journal of the American Society for Information Science, 36(1):56–57, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rush:1987:BRB

[Rus87] James E. Rush. Book review: Technical standards: an introduction for librarians. Journal of the American Society for Information Science, 38(3):221–222, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Radhakrishnan:1978:OSB

[RV78] T. Radhakrishnan and K. Venkatesh. On optimal storage of books by size. Journal of the American Society for Information Science, 29(5):253–254, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rousseau:1996:RJP

[RV96] Ronald Rousseau and Guido Van Hooydonk. Research: Journal production and Journal Impact Factors. Journal of the American Society for Information Science, 47(10):775–780, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rousseau:1998:LEH

[RV98] Ronald Rousseau and Guido Van Hooydonk. Letters to the Editor: The Harter–Nisonger proposal concerning impact factors. Journal of the American Society for Information Science, 49(6): 577–578, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Rowbottom:1982:ESM

[RW82] Mary E. Rowbottom and Peter Willett. The effect of subject matter on the automatic indexing of full text. Journal of the American Society for Information Science, 33(3):139–141, May 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Raghavan:1986:RCA

[RW86] Vijay V. Raghavan and S. K. M. Wong. Research: A critical analysis of vector space model for information retrieval. Journal of the American Society for Information Science, 37(5):279–287, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ross:2000:REU

[RW00] Nancy C. M. Ross and Dietmar Wolfram. Research: End user searching on the Internet: an analysis of term pair topics submitted to the Excite search engine. Journal of the American Society for Information Science, 51(10):949–958, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ryan:1980:BRB

[Rya80] Frederick W. Ryan. Book review: The Social Sciences and Information (in Russian) (Obshchestvennye nauki i informatiia). V. A. Vinogradov. Moscow: Nauka; 1978; 262 pp. Journal of the American Society for Information Science, 31(3):222–223, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ryan:1993:BRB

[Rya93] Susan M. Ryan. Book review: Depository library use of technology: a practitioner’s perspective. Journal of the American Society for Information Science, 44(10):595, December 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sievert:1991:RIC

[SA91] MaryEllen C. Sievert and Mark J. Andrews. Research: Indexing consistency in Information Science Abstracts. Journal of the American Society for Information Science, 42(1):1–6, January 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaughnessy:1997:PLI

[SA97] Thomas Shaughnessy and Carl Adams. Perspectives on … implementation and evaluation of an integrated information center in an academic environment: The design of the Integrated Information Center project. Journal of the American Society for Information Science, 48(5):434–437, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sachs:1976:AAR

[Sac76] Wladimir M. Sachs. An approach to associative retrieval through the theory of fuzzy sets. Journal of the American Society for Information Science, 27(2):85–87, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sager:1975:SGS

[Sag75] Naomi Sager. Sublanguage grammers in science information processing. Journal of the American Society for Information Science, 26(1):10–16, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1970:APF

[Sal70] Gerard Salton. Automatic processing of foreign language documents. Journal of the American Society for Information Science, 21(3):187–194, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1972:BRF

[Sal72a] G. Salton. Book review: F. W. Lancaster, Vocabulary control for information retrieval. Information Resources Press, Washington, 1972, 233 pages. Journal of the American Society for Information Science, 23(6):407, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1972:GER

[Sal72b] G. Salton. The “generality” effect and the retrieval evaluation for large collections. Journal of the American Society for Information Science, 23(1):11–22, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1972:NCB

[Sal72c] G. Salton. A new comparison between conventional indexing (MEDLARS) and automatic text processing (SMART). Journal of the American Society for Information Science, 23(2):75–84, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1972:BCT

[Sal72d] Gerard Salton. Brief communications: Thoughts on the unisist feasibility study. Journal of the American Society for Information Science, 23(1):68–70, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1973:DIS

[Sal73] G. Salton. On the development of information science. Journal of the American Society for Information Science, 24(3):218–220, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1979:LEA

[Sal79a] Gerard Salton. Letter to the Editors: Automatic text analysis. Journal of the American Society for Information Science, 30(2):

116, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1979:LES

[Sal79b] Gerard Salton. Letter to the Editors: Science, Shcharansky, and the Soviets. Journal of the American Society for Information Science, 30(2):114, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1980:PRI

[Sal80] Gerard Salton. A progress report on information privacy and data security. Journal of the American Society for Information Science, 31(2):75–83, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1984:LES

[Sal84] Gerard Salton. Letters to the Editors: Syntactically based indexing. Journal of the American Society for Information Science, 35 (2):136, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1985:BCN

[Sal85] Gerard Salton. Brief communications: A note about information science research. Journal of the American Society for Information Science, 36(4):268–271, July 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salvendy:1986:BRB

[Sal86] Gavriel Salvendy. Book review: Measuring productivity: Trends and comparisons from the First International Productivity Symposium. Journal of the American Society for Information Science, 37(1):49, January 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1987:LER

[Sal87a] Gerald Salton. Letters to the Editor: Rejoinder to Nahum Goldmann’s letter. Journal of the American Society for Information Science, 38(3):216, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1987:HNP

[Sal87b] Gerard Salton. Historical note: The past thirty years in information retrieval. Journal of the American Society for Information Science, 38(5):375–380, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1996:LEN

[Sal96] Gerard Salton. Letters to the Editor: A new horizon for information science. Journal of the American Society for Information Science, 47(4):333, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Samarajiva:1995:BRB

[Sam95] Rohan Samarajiva. Book review: Measuring the impact of information on development. Journal of the American Society for Information Science, 46(1):75–77, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandison:1974:DUA

[San74] Alexander Sandison. Densities of use, and absence of obsolescence, in physics journals at M I T. Journal of the American Society for Information Science, 25(3):172–182, May 1974. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandison:1975:LE

[San75a] A. Sandison. Letters to the Editors. Journal of the American Society for Information Science, 26(2):132, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandison:1975:BCP

[San75b] Alexander Sandison. Brief communications: Patterns of citation densities by date of publication in physical review. Journal of the American Society for Information Science, 26(6):349–352, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandison:1977:BCM

[San77] A. Sandison. Brief communications: Models for librarians. Journal of the American Society for Information Science, 28(5):300– 302, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandison:1979:LEU

[San79a] A. Sandison. Letter to the Editors: Using the collection control model. Journal of the American Society for Information Science, 30(2):115, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandison:1979:LEF

[San79b] A. Sandison. Letters to the Editor: Fluctuations in ranked orders of periodicals. Journal of the American Society for Information Science, 30(4):234, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandison:1983:LET

[San83] A. Sandison. Letters to the Editor: Thematic analysis and “obsolescence”. Journal of the American Society for Information Science, 34(4):295, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandusky:1997:BRB

[San97] Robert J. Sandusky. Book review: Bringing design to software. Journal of the American Society for Information Science, 48(12): 1149–1150, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sandusky:1998:BRB

[San98a] Robert J. Sandusky. Book review: Practical digital libraries: Books, bytes, and bucks. Journal of the American Society for Information Science, 49(11):1038–1039, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sanidas:1998:BRB

[San98b] Amy E. Sanidas. Book review: Beyond the library of the future: More alternative futures for the public library. Journal of the American Society for Information Science, 49(2):187–188, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sanidas:1999:BRB

[San99] Amy E. Sanidas. Book review: Newspapers of record in a digital age: From hot type to hot link, by Shannon E. Martin and Kathleen A. Hansen. Journal of the American Society for Information Science, 50(9):840–841, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sappington:1994:PLR

[Sap94] Thomas E. Sappington. Perspectives on … the redesign/ reengineering of an information services division of a major health science institution: Realigning and reeducating staff. Journal of the American Society for Information Science, 45(5):339–341, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1971:SRI

[Sar71] Tefko Saracevic. Selected results from an inquiry into testing of information retrieval systems. Journal of the American Society for Information Science, 22(2):126–139, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1975:RRF

[Sar75] Tefko Saracevic. Relevance: a review of and a framework for the thinking on the notion in information science. Journal of the American Society for Information Science, 26(6):321–343, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1999:IS

[Sar99] Tefko Saracevic. Information science. Journal of the American Society for Information Science, 50(12):1051–1063, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Savoy:1993:RSF

[Sav93] Jacques Savoy. Research: Stemming of French words based on grammatical categories. Journal of the American Society for Information Science, 44(1):1–9, January 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Savard:1997:BRB

[Sav97a] Réjean Savard. Book review: Strategic management for public libraries: a handbook. Journal of the American Society for Information Science, 48(9):860–861, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Savoy:1997:RRS

[Sav97b] Jacques Savoy. Research: Ranking schemes in hybrid Boolean systems: a new approach. Journal of the American Society for Information Science, 48(3):235–253, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Savoy:1999:RSP

[Sav99] Jacques Savoy. Research: A stemming procedure and stopword list for general French corpora. Journal of the American Society for Information Science, 50(10):944–952, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schultz:1973:LE

[SB73] Louise Schultz and Robert W. Burns, Jr. Letters to the Editors. Journal of the American Society for Information Science, 24(3): 234–235, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sirkin:1980:LEP

[SB80] Arlene Farber Sirkin and Charles L. Blische. Letters to the Editor: Photographic or pictorial collectors. Journal of the American Society for Information Science, 31(5):380, September 1980. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sustik:1983:RII

[SB83] Joan M. Sustik and Terrence A. Brooks. Retrieving information with interactive videodiscs. Journal of the American Society for Information Science, 34(6):424–432, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1990:RIR

[SB90] Gerard Salton and Chris Buckley. Research: Improving retrieval performance by relevance feedback. Journal of the American Society for Information Science, 41(4):288–297, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Straub:1991:PLI

[SB91] Detmar W. Straub and Cynthia Mathis Beath. Perspectives on … integrated information centers within academic environments: The Integrated Information Center concept. Journal of the American Society for Information Science, 42(2):128–131, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sun:1994:ITN

[SB94] Su-Lien Sun and George A. Barnett. The international telephone network and democratization. Journal of the American Society for Information Science, 45(6):411–421, July 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1985:IIS

[SBA85] Tefko Saracevic, Gilda M. Braga, and Matthew A. Afolayan. Issues in information science education in developing countries. Journal of the American Society for Information Science, 36(3): 192–199, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1983:AQF

[SBF83] G. Salton, C. Buckley, and E. A. Fox. Automatic query formulations in information retrieval. Journal of the American Society for Information Science, 34(4):262–280, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shneiderman:1989:PLH

[SBPP89] Ben Shneiderman, Dorothy Brethauer, Catherine Plaisant, and Richard Potter. Perspectives on … hypertext: Evaluating three museum installations of a hypertext system. Journal of the American Society for Information Science, 40(3):172–182, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sullivan:1990:REU

[SBW90] Michael V. Sullivan, Christine L. Borgman, and Dorothy Wippern. Research: End-users, mediated searches, and front-end assistance programs on Dialog: a comparison of learning, performance, and satisfaction. Journal of the American Society for Information Science, 41(1):27–42, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Siegfried:1993:RPE

[SBW93] Susan Siegfried, Marcia J. Bates, and Deborah N. Wilde. Research: A profile of end-user searching behavior by humanities scholars: The Getty Online Searching Project Report No. 2. Journal of the American Society for Information Science, 44(5): 273–291, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sinha:1976:ACC

[SC76] Bani K. Sinha and Richard C. Clelland. Application of a collection-control model for scientific libraries. Journal of the American Society for Information Science, 27(5):320–328, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sinha:1978:BCU

[SC78] Bani K. Sinha and Richard C. Clelland. Brief communications: Using the collection-control model. Journal of the American Society for Information Science, 29(3):161–163, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1980:LEB

[SC80] Gerard Salton and Ana D. Cleveland. Letters to the Editor: Buck’s prime number coding scheme. Journal of the American Society for Information Science, 31(3):219–220, May 1980. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salasin:1985:ISB

[SC85a] J. Salasin and T. Cedar. Information-seeking behavior in an applied research/service delivery setting. Journal of the American Society for Information Science, 36(2):94–102, March 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salasin:1985:PPC

[SC85b] John Salasin and Toby Cedar. Person-to-person communication in an applied research/service delivery setting. Journal of the American Society for Information Science, 36(2):103–115, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Scacchi:1989:PLH

[Sca89] Walt Scacchi. Perspectives on … hypertext: On the power of domain-specific hypertext environments. Journal of the American Society for Information Science, 40(3):183–191, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schacter:1998:RCI

[SCD98] John Schacter, Gregory K. W. K. Chung, and Aimée Dorr. Research: Children’s Internet searching on complex problems: Performance and process analyses. Journal of the American Society for Information Science, 49(9):840–849, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Scheffler:1973:NPD

[Sch73a] Frederic L. Scheffler. A novel philosophy for the design of information storage and retrieval systems appropriate for the ‘70’s. Journal of the American Society for Information Science, 24(3): 205–209, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwarzlander:1973:BRB

[Sch73b] Harry Schwarzlander. Book review: Problems of Information Science, FID 478, A. I. Chernyi, ed. All-Union Institute for Scientific and Technical Information, Moscow 1972. 240 pages. Journal of the American Society for Information Science, 24(6):463–464, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schultz:1974:BCB

[Sch74] Louise Schultz. Breaking the communication barrier between searcher and literature file: an interactive guide. Journal of the American Society for Information Science, 25(1):1–9, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schorr:1975:BCL

[Sch75] Alan Edward Schorr. Brief communications: Lotka’s Law and Map Librarianship. Journal of the American Society for Information Science, 26(3):189–190, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Scholz:1976:BCM

[Sch76] William H. Scholz. Brief communications: A mathematical model for comparison of bibliographic data bases. Journal of the American Society for Information Science, 27(3):183–184, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schultheisz:1981:TEO

[Sch81] Robert J. Schultheisz. TOXLINE: Evolution of an online interactive bibliographic database. Journal of the American Society for Information Science, 32(6):421–429, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schabas:1982:PRC

[Sch82] Ann H. Schabas. Postcoordinate retrieval: a comparison of two indexing languages. Journal of the American Society for Information Science, 33(1):32–37, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwartz:1987:BRB

[Sch87] Candy Schwartz. Book review: Subject and information analysis. Journal of the American Society for Information Science, 38(4): 315–316, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schipma:1988:PLC

[Sch88a] Peter B. Schipma. Perspectives on … CD-ROM for information storage and retrieval: A CD-ROM database product for oncology. Journal of the American Society for Information Science, 39(1): 63–66, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwerin:1988:PLC

[Sch88b] Julie B. Schwerin. Perspectives on … CD-ROM for information storage and retrieval: CD-ROM: Potential markets for information. Journal of the American Society for Information Science, 39(1):54–57, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwarz:1990:RAS

[Sch90] Christoph Schwarz. Research: Automatic syntactic analysis of free text. Journal of the American Society for Information Science, 41(6):408–417, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schultz:1991:MLM

[Sch91a] Claire K. Schultz. In memoriam: Lea M. Bohnert 1919–1990. Journal of the American Society for Information Science, 42(7): 467–468, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwartz:1991:LEI

[Sch91b] Candy Schwartz. Letters to the Editor: Indexer consistency. Journal of the American Society for Information Science, 42(10):

759, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schauder:1994:REP

[Sch94] Don Schauder. Research: Electronic publishing of professional articles: Attitudes of academics and implications for the scholarly communication industry. Journal of the American Society for Information Science, 45(2):73–100, March 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schamber:1996:WDR

[Sch96a] Linda Schamber. What is a document? Rethinking the concept in uneasy times. Journal of the American Society for Information Science, 47(9):669–671, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schroeder:1996:BRB

[Sch96b] Alan T. Schroeder, Jr. Book review: Innovation and the library: the adoption of new ideas in public libraries. Journal of the American Society for Information Science, 47(6):479, June 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schmidt:1997:SDP

[Sch97a] Colin T. Schmidt. The systemics of dialogism: On the prevalence of the self in HCI design. Journal of the American Society for Information Science, 48(11):1073–1081, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schroeder:1997:BRB

[Sch97b] Alan T. Schroeder, Jr. Book review: Data mining with neural networks: Solving business problems from application development to decision support. Journal of the American Society for Information Science, 48(9):862–863, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwartz:1998:BRB

[Sch98a] Candy Schwartz. Book review: Authoritative guide to Web search engines. Journal of the American Society for Information Science, 49(7):672, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwartz:1998:PLI

[Sch98b] Candy Schwartz. Perspective on … Internet issues/technological issues: Web search engines. Journal of the American Society for Information Science, 49(11):973–982, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schwartz:1999:RRS

[Sch99] Candy Schwartz. Research: The role of SIG/ CON in the advancement of information science. Journal of the American Society for Information Science, 50(11):1037–1039, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schamber:2000:TLI

[Sch00a] Linda Schamber. Time-line interviews and inductive content analysis: their effectiveness for exploring cognitive behaviors. Journal of the American Society for Information Science, 51(8): 734–744, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schroeder:2000:BR

[Sch00b] Alan T. Schroeder, Jr. Book reviews. Journal of the American Society for Information Science, 51(5):493–494, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stern:1973:PIS

[SCLL73] Louis W. Stern, C. Samuel Craig, Anthony J. La Greca, and Gerald J. Lazorick. Promotion of information services: an evaluation of alternative approaches. Journal of the American Society for Information Science, 24(3):171–179, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1983:EIM

[SD83] Debora Shaw and Charles H. Davis. Entropy and information: a multidisciplinary overview. Journal of the American Society for Information Science, 34(1):67–74, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1996:RML

[SD96] Debora Shaw and Charles H. Davis. Research: The Modern Language Association: Electronic and paper surveys of computerbased tool use. Journal of the American Society for Information Science, 47(12):932–940, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sengupta:1997:ESA

[SD97] Arijit Sengupta and Andrew Dillon. Extending SGML to accommodate database functions: a methodological overview. Journal of the American Society for Information Science, 48(7):629–637, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Seglen:1992:RSS

[Seg92] Per O. Seglen. Research: The skewness of science. Journal of the American Society for Information Science, 43(9):628–638, October 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Seglen:1994:RCR

[Seg94] Per O. Seglen. Research: Causal relationship between article citedness and journal impact. Journal of the American Society for Information Science, 45(1):1–11, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sellmer:1971:BRB

[Sel71] E. C. Sellmer. Book review: ATA Professional Services Directory — 2nd Edition. 1969. American Translators Association, Thorofare, N.J. Boris R. Anzlowar, Editor. 132 pp. $15.00 ($12.00 if ordered prepaid). Journal of the American Society for Information Science, 22(3):232, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sempsey:1998:BRBa

[Sem98a] James J. Sempsey. Book review: The death of distance: How the communications revolution will change our lives. Journal of the American Society for Information Science, 49(11):1041–1042, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sempsey:1998:BRBb

[Sem98b] James J. Sempsey. Book review: The virtual workplace. Journal of the American Society for Information Science, 49(14):1338, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sempsey:2000:BRa

[Sem00a] James J. Sempsey. Book review. Journal of the American Society for Information Science, 51(2):206, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sempsey:2000:BRb

[Sem00b] James J. Sempsey. Book reviews. Journal of the American Society for Information Science, 51(7):686–687, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sewell:1982:LEC

[Sew82] Winifred Sewell. Letter to the Editors: Computer aids to learning online retrieval. Journal of the American Society for Information Science, 33(6):372, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sutcliffe:2000:ESE

[SEW00] A. G. Sutcliffe, M. Ennis, and S. J. Watkinson. Empirical studies of end-user information searching. Journal of the American Society for Information Science, 51(13):1211–1231, ???? 2000. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sprague:1978:BBS

[SF78] Robert J. Sprague and L. Ben Freudenreich. Building better SDI profiles for users of large, multidisciplinary data bases. Journal of the American Society for Information Science, 29(6):278–282, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shipman:2000:RGP

[SFB^{+}00] Frank M. Shipman III, Richard Furuta, Donald Brenner, Chung-Chi Chung, and Hao wei Hsieh. Research: Guided paths through Web-based collections: Design, experiences, and adaptations. Journal of the American Society for Information Science, 51(3): 260–272, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1985:AFM

[SFV85] G. Salton, E. A. Fox, and E. Voorhees. Advanced feedback methods in information retrieval. Journal of the American Society for Information Science, 36(3):200–210, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Servi:1980:MPJ

[SG80] Patricia N. Servi and Belver C. Griffith. A method for partitioning the journal literature. Journal of the American Society for Information Science, 31(1):36–40, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Small:1989:LEA

[SG89] Henry Small and Eugene Garfield. Letters to the Editor: Analysis of scientific literature to assist in problem solving. Journal of the American Society for Information Science, 40(3):152, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Srivastava:1970:EDV

[SGS70] R. K. Srivastava, B. Garudadwajan, and P. K. Sinha. Evolution & development of a versatile machine-based information system. Journal of the American Society for Information Science, 21(1): 75–85, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sievert:1989:REI

[SH89] MaryEllen Sievert and Mary Haughawout. Research: An editor’s influence on citation patterns: a case study of Elementary School Journal. Journal of the American Society for Information Science, 40(5):334–341, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stephen:1994:PLE

[SH94] Timothy Stephen and Teresa M. Harrison. Perspectives on … electronic publishing: Comserve: Moving the communication discipline online. Journal of the American Society for Information Science, 45(10):765–770, December 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swigger:1996:RCO

[SH96] Kathleen M. Swigger and Ken Hartness. Research: Cooperation and online searching via a computer-supported cooperative problem solving environment. Journal of the American Society for Information Science, 47(5):370–379, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1975:CSC

[Sha75] W. M. Shaw, Jr. Computer Simulation of the Circulation Subsystem of a Library. Journal of the American Society for Information Science, 26(5):271–279, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaughnessy:1978:RLJ

[Sha78] Thomas W. Shaughnessy. Redesigning library jobs. Journal of the American Society for Information Science, 29(4):187–190, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shackel:1983:IPB

[Sha83a] B. Shackel. Invited papers: The BLEND system: Programme for the study of some “electronic journals”. Journal of the American Society for Information Science, 34(1):22–30, January 1983. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1983:SDA

[Sha83b] W. M. Shaw, Jr. Statistical disorder and the analysis of a communication-graph. Journal of the American Society for Information Science, 34(2):146–149, March 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shatzer:1984:LEU

[Sha84] Linda Shatzer. Letters to the Editor: Universities, research, and policy. Journal of the American Society for Information Science, 35(3):153, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1985:CTC

[Sha85] W. M. Shaw, Jr. Critical thresholds in co-citation graphs. Journal of the American Society for Information Science, 36(1):38–43, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sharma:1986:BRBb

[Sha86a] G. Sharma. Book review: Elements of digital satellite communication. Journal of the American Society for Information Science, 37(4):274, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sharma:1986:BRBa

[Sha86b] G. Sharma. Book review: Spread spectrum communications. Journal of the American Society for Information Science, 37(4): 273–274, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1986:BCF

[Sha86c] W. M. Shaw, Jr. Brief communications: On the foundation of evaluation. Journal of the American Society for Information Science, 37(5):346–348, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1990:BRB

[Sha90] Debora Shaw. Book review: Online searching for end users: an information sourcebook. Journal of the American Society for Information Science, 41(3):211, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1991:RSCa

[Sha91a] W. M. Shaw, Jr. Research: Subject and citation indexing. Part I: the clustering structure of composite representations in the Cystic Fibrosis Document Collection. Journal of the American Society for Information Science, 42(9):669–675, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1991:RSCb

[Sha91b] W. M. Shaw, Jr. Research: Subject and citation indexing. Part II: the optimal, cluster-based retrieval performance of composite representations. Journal of the American Society for Information Science, 42(9):676–684, October 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shapiro:1992:ROB

[Sha92] Fred R. Shapiro. Research: Origins of bibliometrics, citation indexing, and citation analysis: the neglected legal literature. Journal of the American Society for Information Science, 43(5): 337–339, June 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shapiro:1995:BCC

[Sha95] Fred R. Shapiro. Brief communications: Coinage of the term information science. Journal of the American Society for Information Science, 46(5):384–385, June 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shaw:1996:BRB

[Sha96] Debora Shaw. Book review: Handbook of usability testing: How to plan, design, and conduct effective tests. Journal of the American Society for Information Science, 47(3):258–259, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shackel:1997:HCI

[Sha97a] Brian Shackel. Human-computer Interaction — Whence and whither? Journal of the American Society for Information Science, 48(11):970–986, November 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shade:1997:BRB

[Sha97b] Leslie Regan Shade. Book review: High noon on the electronic frontier: Conceptual issues in cyberspace. Journal of the American Society for Information Science, 48(8):764–767, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shade:1998:BRB

[Sha98] Leslie Regan Shade. Book review: Virtual individuals, virtual groups: Human dimensions of groupware and computer networking. Journal of the American Society for Information Science, 49 (2):183–184, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shcharansky:1979:LES

[Shc79] Avital Shcharansky. Letters to the Editor: Science, Shcharansky, and the Soviets. Journal of the American Society for Information Science, 30(4):233, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shera:1971:SRI

[She71] Jesse H. Shera. The sociological relationships of information science. Journal of the American Society for Information Science, 22(2):76–80, March 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sherrod:1972:GEO

[She72] John Sherrod. Guest editorial: October is the fairest month …. Journal of the American Society for Information Science, 23(5):

291, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sherman:1986:PLT

[She86] Arthur E. Sherman. Perspectives on … telecommunications: Perspectives, developments, prospects: Trends in telecommunications technology. Journal of the American Society for Information Science, 37(6):414–417, November 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shearer:1991:BRB

[She91] Kenneth D. Shearer. Book review: Language distribution in databases: an analysis and evaluation. Journal of the American Society for Information Science, 42(7):537–538, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shifiett:1985:BRB

[Shi85] Lee Shifiett. Book reviews: Beyond “1984”: the future of library technical services. Peter Gellatly, Ed. New York: Haworth Press; 1984: 265 pp. Price: $24.95 hard cover. Journal of the American Society for Information Science, 36(2):137–139, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shiflett:1986:BRBb

[Shi86a] Lee Shiflett. Book review: Antiquarian bookselling in the United States: a history from the origins to the 1940s. Journal of the American Society for Information Science, 37(5):352–353, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shiflett:1986:BRBa

[Shi86b] Lee Shiflett. Book review: University science and engineering libraries. Journal of the American Society for Information Science, 37(2):90, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shiflett:1988:BRB

[Shi88] Lee Shiflett. Book review: A difficult balance: Editorial peer review in medicine. Journal of the American Society for Information Science, 39(1):22–23, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shneiderman:1986:RDM

[Shn86] Ben Shneiderman. Research: Designing menu selection systems. Journal of the American Society for Information Science, 37(2): 57–70, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shoffner:1971:SIA

[Sho71] Ralph M. Shoffner. Some implications of automatic recognition of bibliographic elements. Journal of the American Society for Information Science, 22(4):275–282, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shpackov:1992:LEN

[Shp92] A. A. Shpackov. Letters to the Editor: The nature and boundaries of information science(s). Journal of the American Society for Information Science, 43(10):678, December 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See rejoinder [Deb92].

Starks:1972:TMC

[SHS72] David D. Starks, Barbara J. Horn, and Thomas P. Slavens. Two modes of computer assisted instruction in a library reference course. Journal of the American Society for Information Science, 23(4):271–277, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shuman:1978:BRB

[Shu78] Bruce A. Shuman. Book review: Public knowledge, private ignorance: Toward a library and information policy. Wilson, Patrick. (Contributions in Librarianship and Information Science. No. 10.) Westport, CT: Greenwood Press; 1977. Price: $13.50. ISBN: 0-8371-9485-7. Journal of the American Society for Information Science, 29(3):167–168, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shuman:1980:BRB

[Shu80] Bruce A. Shuman. Book review: The Search for a Scientific Profession; Library Science Education in the U.S. and Canada. L. Houser, Alvin M. Schrader. Metuchen, NJ: Scarecrow Press; 1978: 192 pp. Price; $8.00. Journal of the American Society for Information Science, 31(6):455, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sichel:1985:BDW

[Sic85] H. S. Sichel. A bibliometric distribution which really works. Journal of the American Society for Information Science, 36(5):314– 321, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sichel:1992:RNS

[Sic92] H. S. Sichel. Research: Note on a strongly unimodal bibliometric size frequency distribution. Journal of the American Society for Information Science, 43(4):299–303, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Siecel:1970:BRB

[Sie70] Suzanne Siecel. Book review: 5/70-3R. Information Storage and Retrieval Systems for Individual Researchers. 1970. Gerald Jahodn. Wiley-Interscience. New York. 135 pp. Journal of the American Society for Information Science, 21(5):376–377, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Siegel:1986:BRB

[Sie86] Leah H. Siegel. Book review: The microcomputer user’s guide to information online. Journal of the American Society for Information Science, 37(2):92, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sievert:1989:BRBa

[Sie89a] MaryEllen C. Sievert. Book review: Online search analyst, version 1.0. Journal of the American Society for Information Science, 40(2):133, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sievert:1989:BRBb

[Sie89b] MaryEllen C. Sievert. Book review: Search strategies in mass communication. Journal of the American Society for Information Science, 40(2):133–134, March 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sievert:1990:RTY

[Sie90] MaryEllen C. Sievert. Research: Ten years of the literature of online searching: an analysis of Online and Online Review. Journal of the American Society for Information Science, 41(8):560–568, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sievert:1996:FTI

[Sie96] MaryEllen C. Sievert. Full-text information retrieval: Introduction. Journal of the American Society for Information Science, 47(4):261–262, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Simon:1981:IPM

[Sim81] Herbert A. Simon. Information-processing models of cognition. Journal of the American Society for Information Science, 32(5): 364–377, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Simkins:1983:E

[Sim83] Alan Simkins. Editorial. Journal of the American Society for Information Science, 34(1):1–4, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1972:CAT

[SJ72] Kingsley R. Smith II and Theodore M. Judy. Correcting author, title and multiple-term index entries in mechanized bibliography preparation. Journal of the American Society for Information Science, 23(6):379–383, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sligo:2000:KBG

[SJ00] F. X. Sligo and Anna M. Jameson. The knowledge — behavior gap in use of health information. Journal of the American Society for Information Science, 51(9):858–869, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1988:RSIb

[SK88a] Tefko Saracevic and Paul Kantor. Research: A study of information seeking and retrieving. II. Users, questions, and effectiveness. Journal of the American Society for Information Science, 39(3): 177–196, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1988:RSIc

[SK88b] Tefko Saracevic and Paul Kantor. Research: A study of information seeking and retrieving. III. Searchers, searches, and overlap.

Journal of the American Society for Information Science, 39(3): 197–216, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1997:RSVa

[SK97a] Tefko Saracevic and Paul B. Kantor. Research: Studying the value of library and information services. Part I. Establishing a theoretical framework. Journal of the American Society for Information Science, 48(6):527–542, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1997:RSVb

[SK97b] Tefko Saracevic and Paul B. Kantor. Research: Studying the value of library and information services. Part II. Methodology and taxonomy. Journal of the American Society for Information Science, 48(6):543–563, June 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Szava-Kovats:1998:BCN

[SK98] Endre Száva-Kováts. Brief communications: Non-indexed indirect-collective citedness (NIICC). Journal of the American Society for Information Science, 49(5):477–481, ???? 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Szava-Kovats:1999:RIC

[SK99] Endre Száva-Kováts. Research: Indirect-collective referencing (ICR): life course, nature, and importance of a special kind of scientific referencing. Journal of the American Society for Information Science, 50(14):1284–1294, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Savage-Knepshield:1999:IIR

[SKB99] Pamela A. Savage-Knepshield and Nicholas J. Belkin. Interaction in information retrieval: Trends over time. Journal of the American Society for Information Science, 50(12):1067–1082, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1988:RSIa

[SKCT88] Tefko Saracevic, Paul Kantor, Alice Y. Chamis, and Donna Trivison. Research: A study of information seeking and retrieving.

I. Background and methodology. Journal of the American Society for Information Science, 39(3):161–176, May 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salminen:1997:TMD

[SKL97] Airi Salminen, Katri Kauppinen, and Merja Lehtovaara. Towards a methodology for document analysis. Journal of the American Society for Information Science, 48(7):644–655, July 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Skovira:1998:BRB

[Sko98] Robert J. Skovira. Book review: Visual explanations: Images and quantities, evidence and narrative. Journal of the American Society for Information Science, 49(6):574–575, ???? 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schuegraf:1992:RAP

[SKvB92] Ernst J. Schuegraf, Liam T. Keliher, and Martin F. van Bommel. Research: An analysis of personal journal subscriptions of university faculty. Part I. Science. Journal of the American Society for Information Science, 43(1):28–33, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Strong:1983:VRM

[SL83] Harry Strong and David H. Lehman. Videodisc research at the MITRE corporation. Journal of the American Society for Information Science, 34(6):433, November 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smiraglia:1999:RDB

[SL99] Richard P. Smiraglia and Gregory H. Leazer. Research: Derivative bibliographic relationships: the work relationship in a global bibliographic database. Journal of the American Society for Information Science, 50(6):493–504, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Slamecka:1973:ASS

[Sla73] Vladimir Slamecka. An audiographic system for self-instruction. Journal of the American Society for Information Science, 24(6): 405–412, November 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Slamecka:1975:OPP

[Sla75] Vladimir Slamecka. Opinion Paper. Pragmatic observations on theoretical research in information science. Journal of the American Society for Information Science, 26(6):318–320, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Slamecka:1985:ITT

[Sla85] Vladimir Slamecka. Information technology and the third world. Journal of the American Society for Information Science, 36(3): 178–183, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Scheuermann:1998:MQP

[SLC98] Peter Scheuermann, Wen-Syan Li, and Chris Clifton. Multidatabase query processing with uncertainty in global keys and attribute values. Journal of the American Society for Information Science, 49(3):283–301, March 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sleeter:1994:BRB

[Sle94] Ellen Sleeter. Book review: Managing the economics of owning, leasing and contracting out information services. Journal of the American Society for Information Science, 45(3):218, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Slone:2000:EOL

[Slo00] Debra J. Slone. Encounters with the OPAC: On-line searching in public libraries. Journal of the American Society for Information Science, 51(8):757–773, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Segebarth:1970:LE

[SLW70] C. Segebarth, F. W. Lancaster, and Eugene Wall. Letters to the Editor. Journal of the American Society for Information Science, 21(5):372–374, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sun:1992:RIO

[SLZC92] Wei Sun, Lon-Mu Liu, Weining Zhang, and John Craig Comfort. Research: Intelligent OCR processing. Journal of the American Society for Information Science, 43(6):422–431, July 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1975:PES

[SM75] Rowena Weiss Swanson and Joseph Mayer. Performing evaluation studies in information science. Journal of the American Society for Information Science, 26(3):140–156, May 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1978:LEB

[SM78] Tefko Saracevic and Charles T. Meadow. Letters to the Editor: On JASIS refereeing. Journal of the American Society for Information Science, 29(4):209–210, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Settel:1988:PLE

[SM88] Barbara Settel and Donald A. Marchand. Perspectives on … education of the information professional: New dimensions, new directions: Syracuse University School of Information Studies: a tradition of innovation. Journal of the American Society for Information Science, 39(5):331–333, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Svenonius:1990:RAD

[SM90] Elaine Svenonius and Mavis Molto. Research: Automatic derivation of name access points in cataloging. Journal of the American Society for Information Science, 41(4):254–263, June 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salager-Meyer:1991:BCM

[SM91] Françoise Salager-Meyer. Brief communications: Medical English abstracts: How well are they structured? Journal of the American Society for Information Science, 42(7):528–531, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Small:1973:CCS

[Sma73] Henry Small. Co-citation in the scientific literature: a new measure of the relationship between two documents. Journal of the American Society for Information Science, 24(4):265–269, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smalley:1980:CPA

[Sma80] Topsy N. Smalley. Comparing Psychological Abstracts and Index Medicus for coverage of the journal literature in a subject area in psychology. Journal of the American Society for Information Science, 31(3):143–146, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Small:1986:RSS

[Sma86] Henry Small. Research: The synthesis of specialty narratives from co-citation clusters. Journal of the American Society for Information Science, 37(3):97–110, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Small:1999:POI

[Sma99] Henry Small. Perspectives on … visual information retrieval interfaces: Visualizing science by citation mapping. Journal of the American Society for Information Science, 50(9):799–813, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Scheffler:1972:ESU

[SMB72] Frederic Scheffler, Jacqueline March, and John Bernados. An experiment to study the use of boolean not logic to improve the precision of selective dissemination of information. Journal of the American Society for Information Science, 23(1):58–65, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Summers:1983:EPP

[SMC83] Edward G. Summers, Joyce Matheson, and Robert Conry. The effect of personal, professional, and psychological attributes, and information seeking behavior on the use of information sources by educators. Journal of the American Society for Information Science, 34(1):75–85, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smeaton:1987:ERL

[Sme87] Alan F. Smeaton. European research letters: Information retrieval research and ESPRIT. Journal of the American Society for Information Science, 38(1):21–22, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smeaton:1992:ERL

[Sme92] Alan F. Smeaton. European research letters: Europe, ESPRIT II, and information retrieval. Journal of the American Society for Information Science, 43(1):89–91, January 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shuldberg:1993:RDI

[SMHC93] H. Kelly Shuldberg, Melissa Macpherson, Pete Humphrey, and Jamii Corley. Research: Distilling information from text: The EDS TemplateFiller System. Journal of the American Society for Information Science, 44(9):493–507, October 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1974:PPC

[Smi74a] Joshua I. Smith. Position paper. Cutting the NSF-OSIS budget: Potential disaster for information science and technology. Journal of the American Society for Information Science, 25(2):77–85, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1974:SPA

[Smi74b] Linda C. Smith. Student paper award 1974. Systematic searching of abstracts and indexes in interdisciplinary areas. Journal of the American Society for Information Science, 25(6):343–353, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1981:UQN

[Smi81] J. Macgregor Smith. The use of queuing networks and mixed integer programming to allocate resources optimally within a library layout. Journal of the American Society for Information Science, 32(1):33–42, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1984:BCM

[Smi84a] Charles H. Smith. Brief communications: A method for studying intercorrelated circulation patterns in library systems. Journal of the American Society for Information Science, 35(2):118–121, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See comment [McG84].

Smith:1984:AIS

[Smi84b] Linda C. Smith. Artificial intelligence: a selected bibliography. Journal of the American Society for Information Science, 35(5): 317–319, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1996:BRBa

[Smi96a] Donald R. Smith. Book review: The impact of emerging technologies on reference service and bibliographic instruction. Journal of the American Society for Information Science, 47(8):654–655A, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1996:BRBb

[Smi96b] Donald R. Smith. Book review: The impact of emerging technologies on reference service and bibliographic instruction. Journal of the American Society for Information Science, 47(8):654–655A, August 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1997:BRB

[Smi97] Martha Montague Smith. Book review: Internet dreams: Archetypes, myths, and metaphors for inventing the net. Journal of the American Society for Information Science, 48(8):767–769, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1998:BRBc

[Smi98a] Derek G. Smith. Book review: Complete Internet companion for librarians. Journal of the American Society for Information Science, 49(9):861–863, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1998:BRBb

[Smi98b] Derek G. Smith. Book review: Internet besieged: Countering cyberspace scofflaws. Journal of the American Society for Information Science, 49(8):764–766, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1998:BRBa

[Smi98c] Martha M. Smith. Book review: Information science: Integration in perspective. Proceedings of the Second International Conference on Conceptions of Library and Information Science. Journal of the American Society for Information Science, 49(1):96–97, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1999:RBA

[Smi99] Linda C. Smith. Research: Journal of the American Society for Information Science (JASIS): Past, present and future. Journal of the American Society for Information Science, 50(11):965–969, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:2000:BRB

[Smi00a] Donald R. Smith. Book review: Computer-based library information systems designing techniques, by Madan Mohan Kashyap. Journal of the American Society for Information Science, 51(12): 1149, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:2000:BR

[Smi00b] Donald R. Smith. Book reviews. Journal of the American Society for Information Science, 51(7):687, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sievert:1995:PLM

[SMJ95] MaryEllen C. Sievert, Emma Jean McKinin, and E. Diane Johnson. Perspectives on … medical informatics: Information technology in health care: Full-text databases in medicine. Journal of the American Society for Information Science, 46(10):748–754, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Srinivasan:2000:MHI

[SNG00] Uma Srinivasan, Anne H. H. Ngu, and Tom Gedeon. Managing heterogeneous information systems through discovery and retrieval of generic concepts. Journal of the American Society for Information Science, 51(8):707–723, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Snyder:1998:BRB

[Sny98] Herbert Snyder. Book review: The economics of information: a guide to economic and cost-benefit analysis for information professionals. Journal of the American Society for Information Science, 49(4):382–383, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Soergel:1972:UST

[Soe72] Dagobert Soergel. A universal source thesaurus as a classification generator. Journal of the American Society for Information Science, 23(5):299–305, September 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Soergel:1976:USH

[Soe76] Dagobert Soergel. Is user satisfaction a hobgoblin? Journal of the American Society for Information Science, 27(4):256–259, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See reply [Coo76].

Soergel:1994:PLI

[Soe94] Dagobert Soergel. Perspectives on … indexing: Indexing and retrieval performance: the logical evidence. Journal of the American Society for Information Science, 45(8):589–599, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Soergel:1999:ROR

[Soe99] Dagobert Soergel. The rise of ontologies or the reinvention of classification. Journal of the American Society for Information Science, 50(12):1119–1120, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Solomon:1974:DSS

[Sol74] Susan L. Solomon. Determination of satisfactory scale for data processing facilities. Journal of the American Society for Information Science, 25(3):151–155, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Solomon:1993:RCI

[Sol93] Paul Solomon. Research: Children’s information retrieval behavior: a case analysis of an OPAC. Journal of the American Society for Information Science, 44(5):245–264, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Solomon:1997:RDIa

[Sol97a] Paul Solomon. Research: Discovering information behavior in sense making. I. Time and timing. Journal of the American Society for Information Science, 48(12):1097–1108, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Solomon:1997:RDIb

[Sol97b] Paul Solomon. Research: Discovering information behavior in sense making. II. The social. Journal of the American Society for Information Science, 48(12):1109–1126, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Solomon:1997:RDIc

[Sol97c] Paul Solomon. Research: Discovering information behavior in sense making. III. The person. Journal of the American Society for Information Science, 48(12):1127–1138, December 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Summers:1999:ISL

[SOM^{+}99] Ron Summers, Charles Oppenheim, Jack Meadows, Cliff McKnight, and Margaret Kinnell. Information science in 2010: A Loughborough University view. Journal of the American Society for Information Science, 50(12):1153–1162, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Soper:1996:BRB

[Sop96] Mary Ellen Soper. Book review: A library manager’s guide to the physical processing of nonbook materials. Journal of the American Society for Information Science, 47(4):327–328, April 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sorkin:1996:LEL

[Sor96] David E. Sorkin. Letters to the Editor: Legal informatics. Journal of the American Society for Information Science, 47(5):407, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Soultoukis:1996:BRB

[Sou96] Donna Zoccola Soultoukis. Book review: The Internet access cookbook: a librarian’s commonsense guide to low-cost connections. Journal of the American Society for Information Science, 47(9):725–725B, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Saracevic:1973:AAS

[SP73] Tefko Saracevic and Lawrence J. Perk. Ascertaining activities in a subject area through bibliometric analysis. Application to library literature. Journal of the American Society for Information Science, 24(2):120–134, March 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Shepherd:1986:RPQ

[SP86] Michael A. Shepherd and W. J. Phillips. Research: The profilequery relationship. Journal of the American Society for Information Science, 37(3):146–152, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spasser:1997:REF

[Spa97] Mark A. Spasser. Research: The enacted fate of undiscovered public knowledge. Journal of the American Society for Information Science, 48(8):707–717, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spasser:1999:IIS

[Spa99] Mark A. Spasser. Informing information science: the case for activity theory. Journal of the American Society for Information Science, 50(12):1136–1138, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spencer:1971:RTS

[Spe71] Carol C. Spencer. Random time sampling with self-observation for library cost studies: Unit costs of interlibrary loans and photocopies at a regional medical library. Journal of the American Society for Information Science, 22(3):153–160, May 1971. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Speert:1976:LE

[Spe76] Kathryn H. Speert. Letter to the Editor. Journal of the American Society for Information Science, 27(5):353, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Specht:1980:PUO

[Spe80] Jerry Specht. Patron use of an online circulation system in known-item searching. Journal of the American Society for Information Science, 31(5):335–346, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spink:1994:BRB

[Spi94] Amanda Spink. Book review: Online retrieval: a dialogue of theory and practice. Journal of the American Society for Information Science, 45(8):565, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spink:1996:RMS

[Spi96] Amanda Spink. Research: Multiple search sessions model of enduser behavior: an exploratory study. Journal of the American Society for Information Science, 47(8):603–609, August 1996. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spink:1997:RIS

[Spi97a] Amanda Spink. Research: Information science: a third feedback framework. Journal of the American Society for Information Science, 48(8):728–740, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spink:1997:RSI

[Spi97b] Amanda Spink. Research: Study of interactive feedback during mediated information retrieval. Journal of the American Society for Information Science, 48(5):382–394, May 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sprehe:1987:PLF

[Spr87] J. Timothy Sprehe. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Section I. National policy for health information services. Policy on management of Federal Information Resources. Journal of the American Society for Information Science, 38(1):30–33, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sprehe:1999:RGI

[Spr99] J. Timothy Sprehe. Research: Government information: From inaccessibility to your desktop and back again. Journal of the American Society for Information Science, 50(4):340–345, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Speier:1999:RFP

[SPWH99] Cheri Speier, Jonathan Palmer, Daniel Wren, and Susan Hahn. Research: Faculty perceptions of electronic journals as scholarly communication: a question of prestige and legitimacy. Journal of the American Society for Information Science, 50(6):537–543, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1977:RBA

[SR77] Joseph D. Smith and James E. Rush. The relationship between author names and author entries in a large on-line union catalog as retrieved using truncated keys. Journal of the American Society for Information Science, 28(2):115–120, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Smith:1979:AQN

[SR79] J. Macgregor Smith and Nilliam B. Rouse. Application of queueing network models to optimization of resource allocation within libraries. Journal of the American Society for Information Science, 30(5):250–263, September 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steynberg:1995:BCM

[SR95] Susan Steynberg and Steve F. Rossouw. Brief communications: Multiple authorship in biomedical papers: A South African case study. Journal of the American Society for Information Science, 46(6):468–472, July 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Srinivasan:1990:RGT

[Sri90] Padmini Srinivasan. Research: On generalizing the Two-Poisson Model. Journal of the American Society for Information Science, 41(1):61–66, January 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spink:1998:RUS

[SRS98] Amanda Spink, David Robins, and Linda Schamber. Research: Use of scholarly book reviews: Implications for electronic publishing and scholarly communication. Journal of the American Society for Information Science, 49(4):364–374, ???? 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schultz:1970:JDB

[SS70] Wallace L. Schultz and Claire K. Shultz. John Desmond Bernal — Evangelist of science (1901–). Journal of the American Society for Information Science, 21(2):142–144, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spaulding:1976:CAS

[SS76] F. H. Spaulding and R. O. Stanton. Computer-aided selection in a library network. Journal of the American Society for Information Science, 27(5):269–280, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Summers:1989:RCC

[SS89] John E. Summers and Edward G. Summers. Research: The computerized cataloguing of historic watercraft: a case study in information retrieval in museology. Journal of the American Society for Information Science, 40(4):253–261, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sillince:1993:RMP

[SS93] J. A. A. Sillince and M. Sillince. Research: Market pressure and government intervention in the administration and development of molecular databases. Journal of the American Society for Information Science, 44(1):28–39, January 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stuart-Stubbs:1995:BRB

[SS95] Basil Stuart-Stubbs. Book review: Critical approaches to information technology in librarianship, foundations and applications. Journal of the American Society for Information Science, 46(1): 73–75, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Spink:1997:RII

[SS97] Amanda Spink and Tefko Saracevic. Research: Interaction in information retrieval: Selection and effectiveness of search terms. Journal of the American Society for Information Science, 48(8): 741–761, August 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steele:2000:IIR

[SS00] Thomas W. Steele and Jeffrey C. Stier. The impact of interdisciplinary research in the environmental sciences: a forestry case study. Journal of the American Society for Information Science, 51(5):476–484, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sun:1999:RME

[SSD99] Qinglan Sun, Debora Shaw, and Charles H. Davis. Research: A model for estimating the occurrence of same-frequency words and the boundary between high- and low-frequency words in texts. Journal of the American Society for Information Science, 50(3): 280–286, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Samstag-Schnock:1993:RPE

[SSM93] Uwe Samstag-Schnock and Charles T. Meadow. Research: PBS: an economical natural language query interpreter. Journal of the American Society for Information Science, 44(5):265–272, June 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sewell:1986:ROE

[ST86] Winifred Sewell and Sandra Teitelbaum. Research: Observations of end-user online searching behavior over eleven years. Journal of the American Society for Information Science, 37(4):234–245, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schubert:1989:REP

[ST89] András Schubert and András Telcs. Research: Estimation of the publication potential in 50 U.S. states and in the District of Columbia based on the frequency distribution of scientific productivity. Journal of the American Society for Information Science, 40(4):291–297, July 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stanfield:1975:BRB

[Sta75a] Jonathan Stanfield. Book review: Management of information technology case studies. Elizabeth B. Adams, New York, NY, Petrocelli/Charter (1975). Journal of the American Society for Information Science, 26(6):354, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stanton:1975:AMO

[Sta75b] Robert O. Stanton. Applying the management-by-objectives technique in an industrial library. Journal of the American Society for Information Science, 26(6):313–317, November 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stanfield:1977:BRB

[Sta77a] Jonathan Stanfield. Book review: System Analysis for Managerial Decisions, A Computer Approach. P. Ramahgam. New York: John Wiley and Sons, 607 pp. (1976). Journal of the American Society for Information Science, 28(6):376, November 1977. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stanford:1977:BRB

[Sta77b] Jonathan Stanford. Book review: Information Retrieval, British and American, 1876–1976. Metcalfe, John. Metuchen, NJ: The Scarecrow Press (1976). 243 pp. Journal of the American Society for Information Science, 28(1):62, January 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stanoulov:1982:EAI

[Sta82] Nicolay Stanoulov. An Evolutionary Approach in Information Systems Science. Journal of the American Society for Information Science, 33(5):311–316, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Standera:1985:EPP

[Sta85] O. L. Standera. Electronic publishing: the predicament of occasional users in the editorial process. Journal of the American Society for Information Science, 36(4):230–240, July 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stavri:1998:BRB

[Sta98] P. Zoë Stavri. Book review: Automating the lexicon: Research and practice in a multilingual environment. Journal of the American Society for Information Science, 49(4):380–381, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steinacker:1974:IAS

[Ste74a] Ivo Steinacker. Indexing and automatic significance analysis. Journal of the American Society for Information Science, 25(4): 237–241, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stevens:1974:CAL

[Ste74b] Norman D. Stevens. A computer analysis of library postcards (CALP). Journal of the American Society for Information Science, 25(5):332–335, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steele:1977:BRB

[Ste77] Richard A. Steele. Book review: American Overseas Library Technical Assistance. 1940–1970. Beverly J. Brewster, Metuchen, NJ: Scarecrow Press, Inc., 445 pp. (1976) $17.50. Journal of the American Society for Information Science, 28(6):372–373, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stevens:1980:LNR

[Ste80] Norman D. Stevens. Library networks and resource sharing in the united states: an historical and philosophical overview. Journal of the American Society for Information Science, 31(6):405–412, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steinacker:1984:LED

[Ste84] Ivo Steinacker. Letters to the Editor: Derek de Solla Price. Journal of the American Society for Information Science, 35(6):372, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stephenson:1985:BCR

[Ste85] Mary Sue Stephenson. Brief communications: The research method used in subfields and the growth of published literature in those subfields: Vertebrate paleontology and geochemistry. Journal of the American Society for Information Science, 36(2):130– 133, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stern:1990:RUB

[Ste90] Richard E. Stern. Research: Uncitedness in the biomedical literature. Journal of the American Society for Information Science, 41(3):193–196, April 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steckel:1999:BRB

[Ste99a] Mike Steckel. Book review: Deep information: the role of information policy in environmental sustainability, by John Felleman. Journal of the American Society for Information Science, 50(8): 724–725, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stevens:1999:PLC

[Ste99b] Joann Stevens. Perspective on … copyright and fair-use guidelines for education and libraries: The multimedia guidelines. Journal of the American Society for Information Science, 50(14): 1324–1327, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steckel:2000:BRB

[Ste00a] Mike Steckel. Book review: U.S. Government on the Web: Getting the information you need, by Peter Hernon, John A. Shuler, and Robert E. Dugan. Journal of the American Society for Information Science, 51(10):963–964, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Steckel:2000:BR

[Ste00b] Mike Steckel. Book reviews. Journal of the American Society for Information Science, 51(5):488–490, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stirling:2000:BR

[Sti00] Dale A. Stirling. Book reviews. Journal of the American Society for Information Science, 51(5):492–493, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Strong:1974:SPA

[Str74] Suzanne M. Strong. 1973 student paper award. An algorithm for generating structural surrogates of English text. Journal of the American Society for Information Science, 25(1):10–24, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Struminger:1976:BRB

[Str76] Leny Struminger. Book review: Introduction to computer science. Harry Katzan, Jr., New York: Petrocelli Books, 495 pp. (1975). Journal of the American Society for Information Science, 27(2): 125, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Struminger:1978:BRB

[Str78] L. Struminger. Book review: Data structures and programming techniques. Maurer, Herman A. Englewood Cliffs, NJ: Prentice-Hall; 1977: 228 pp. Price: $13.50. Journal of the American Society for Information Science, 29(5):264, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Strong:1982:ASS

[Str82] Gary W. Strong. Adaptive system: the study of information, pattern, and behavior. Journal of the American Society for Information Science, 33(6):400–406, November 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Stroumtsos:1984:LEV

[Str84] Lorraine Y. Stroumtsos. Letters to the Editors: The value of information services. Journal of the American Society for Information Science, 35(2):136, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Studer:1972:MIR

[Stu72] Paul A. Studer. From multidisciplinary to interdisciplinary research: Effects of information systems on formal organizations. Journal of the American Society for Information Science, 23(6): 343–352, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Studer:1978:LEI

[Stu78] Paul A. Studer. Letters to the Editor: Information science and the phenomenon of information. Journal of the American Society for Information Science, 29(4):210–211, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Su:1994:RRR

[Su94] Louise T. Su. Research: The relevance of recall and precision in user evaluation. Journal of the American Society for Information Science, 45(3):207–217, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Subramanyam:1984:BCR

[Sub84] K. Subramanyam. Brief communication: Research productivity and breadth of interest of computer scientists. Journal of the American Society for Information Science, 35(6):369–371, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Summers:1984:ARA

[Sum84] Edward G. Summers. Application: A review and application of citation analysis methodology to reading research journal literature. Journal of the American Society for Information Science, 35 (6):332–343, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sundt:1999:PLC

[Sun99] Christine L. Sundt. Perspective on … copyright and fair-use guidelines for education and libraries: Testing the limits: The CONFU digital-images and multimedia guidelines and their consequences for libraries and educators. Journal of the American Society for Information Science, 50(14):1328–1336, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sutton:1994:RRA

[Sut94] Stuart A. Sutton. Research: The role of attorney mental models of law in case relevance determinations: an exploratory analysis. Journal of the American Society for Information Science, 45(3): 186–200, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sutton:1996:IDL

[Sut96] Stuart A. Sutton. Instances of distance learning: Planning for the Twenty-First Century: The California State University. Journal of the American Society for Information Science, 47(11):821–825, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sutton:1999:RCD

[Sut99] Stuart A. Sutton. Research: Conceptual design and deployment of a metadata framework for educational resources on the internet. Journal of the American Society for Information Science, 50(13):1182–1192, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schuegraf:1994:RAP

[SvB94] Ernst J. Schuegraf and Martin F. van Bommel. Research: An analysis of personal journal subscriptions of university faculty. Part II. Arts and professional programs. Journal of the American Society for Information Science, 45(7):477–482, August 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Svenonius:1972:EIT

[Sve72] Elaine Svenonius. An experiment in index term frequency. Journal of the American Society for Information Science, 23(2):109– 121, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Svenonius:1986:RUQ

[Sve86] Elaine Svenonius. Research: Unanswered questions in the design of controlled vocabularies. Journal of the American Society for Information Science, 37(5):331–340, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Svenonius:1994:PLI

[Sve94] Elaine Svenonius. Perspectives on … indexing: Access to nonbook materials: the limits of subject indexing for visual and aural languages. Journal of the American Society for Information Science, 45(8):600–606, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Santhanam:1991:RMI

[SW91] Radhika Santhanam and Susan Wiedenbeck. Research: Modeling the intermittent user of word processing technology. Journal of the American Society for Information Science, 42(3):185–196, April 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salminen:1992:RTL

[SW92] Airi Salminen and Carolyn Watters. Research: A two-level structure for textual databases to support hypertext access. Journal of the American Society for Information Science, 43(6):432–447, July 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1972:BCT

[Swa72] Rowena W. Swanson. Brief communications: On terminology for search strategies. Journal of the American Society for Information Science, 23(2):133–134, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1973:CTI

[Swa73] Rowena W. Swanson. On Clustering Techniques in Information Science. Journal of the American Society for Information Science, 24(1):72–73, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1975:BRBa

[Swa75a] Rowena Weiss Swanson. Book review: Communication science and technology. Patrick R. Penland. New York: Marcel Dekker, Inc., 205 pp. (1974). Journal of the American Society for Information Science, 26(2):131–132, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1975:BRBb

[Swa75b] Rowena Weiss Swanson. Book review: Proceedings of the 1973 clinic on library applications of data processing: Networking and Other Forms of Cooperation. Journal of the American Society for Information Science, 26(4):263–264, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1976:BRBc

[Swa76a] Rowena Weiss Swanson. Book review: Library, documentation and archives serials, 4th ed., Crazyna Janzing, compiler. K. R. Brown, ed. The Hague, The Netherlands: International Federation for Documentation, (FID Publication No. 532) 203 pp. (1975). Journal of the American Society for Information Science, 27(5):348, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1976:BRBb

[Swa76b] Rowena Weiss Swanson. Book review: Proceedings of the 1974 clinic on library applications of data processing: Applications of minicomputers to library and related problems. F. Wilfrid Lancaster. (Ed.) Urbana-Champaign, Illinois: University of Illinois, Graduate School of Library Science, 195 pp. (1974). Journal of the American Society for Information Science, 27(2):123–124, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1976:BRBa

[Swa76c] Rowena Weiss Swanson. Book review: The Organization of the Library Profession, A Symposium Based on Contributions to the 37th Session of the IFIA General Council, Liverpool, 1971. A. H. Chaplin, ed. The Hague, The Netherlands: International Federation of Library Associations. 132 pp. (1973). Journal of the American Society for Information Science, 27(1):76–77, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1976:BCW

[Swa76d] Rowena Weiss Swanson. Brief communications: A Work Study of the Review Production Process. Journal of the American Society for Information Science, 27(1):70–72, January 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1977:AFD

[Swa77a] E. Burton Swanson. The application of function descriptors to the development of an information system topology. Journal of the American Society for Information Science, 28(5):259–267, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See errata [Swa78a].

Swanson:1977:BRBa

[Swa77b] Rowena Weis Swanson. Book review: Proceedings of the 1975 Clinic on Library Applications of Data Processing: The Use of Computers in Literature Searching and Related Reference Activities in Libraries. F. Wilfrid Lancaster, Ed. Urbana-Champaign, IL: University of Illinois, 159 pp. (1976). (ISBN 0-87845-043-2; LC 76-1790). Journal of the American Society for Information Science, 28(2):125–126, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1977:BRBb

[Swa77c] Rowena Weiss Swanson. Book review: Advances in Librarianship. Volume 7. Melvin J. Voight and Michael H. Harris, Eds. New York: Academic Press, 348 pp. (1977). Journal of the American Society for Information Science, 28(5):304–305, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1978:EAF

[Swa78a] E. Burton Swanson. Errata: The application of function descriptors to the development of an information system typology. Journal of the American Society for Information Science, 29(5): 256, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Swa77a].

Swanson:1978:BRB

[Swa78b] Rowena Weiss Swanson. Book review: Evaluation and scientific management of libraries and information centers. Lancaster, F. W.; Cleverdon, C. W., Ed. Leyden, Netherlands: Noordhoff International Publishing; 1977: 184 pp. (NATO Advanced Study Institutes Series, Series E: Applied Science-No. 18). Journal of the American Society for Information Science, 29(5):263–264, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1978:EIS

[Swa78c] Rowena Weiss Swanson. Education for information science as a profession. Journal of the American Society for Information Science, 29(3):148–155, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1987:RTM

[Swa87] Don R. Swanson. Research: Two medical literatures that are logically but not bibliographically connected. Journal of the American Society for Information Science, 38(4):228–233, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1988:HHN

[Swa88] Don R. Swanson. Historical: Historical note: Information retrieval and the future of an illusion. Journal of the American Society for Information Science, 39(2):92–98, March 1988. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1989:BCS

[Swa89a] Don R. Swanson. Brief communications: A second example of mutually isolated medical literatures related by implicit, unnoticed connections. Journal of the American Society for Information Science, 40(6):432–435, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1989:BCO

[Swa89b] Don R. Swanson. Brief communications: Online search for logically-related noninteractive medical literatures: a systematic trial-and-error strategy. Journal of the American Society for Information Science, 40(5):356–358, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1989:LER

[Swa89c] Don R. Swanson. Letters to the Editor: Response: Absence of citations can be valuable clue. Journal of the American Society for Information Science, 40(3):152, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swanson:1989:LEV

[Swa89d] Don R. Swanson. Letters to the Editor: Verification of results that logically related noninteractive literatures are potential sources of new knowledge. Journal of the American Society for Information Science, 40(3):152, May 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Swift:1979:SAD

[SWB79] Donald F. Swift, Viola A. Winn, and Dawn A. Bramer. A sociological approach to the design of information systems. Journal of the American Society for Information Science, 30(4):215–223, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schultheisz:1978:DIL

[SWK78] Robert J. Schultheisz, Donald F. Walker, and Kay L. Kannan. Design and implementation of an on-line chemical dictionary (CHEMLINE). Journal of the American Society for Information Science, 29(4):173–179, July 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Schipma:1971:CDD

[SWS71] Peter B. Schipma, Martha E. Williams, and Allan L. Shafton. Comparison of document data bases. Journal of the American Society for Information Science, 22(5):326–332, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Sullo:1979:RTC

[SWT^{+}79] Pasquale Sullo, William A. Wallace, Thomas Triscari, Cathy A. Chazen, and James F. Davis. A Reliability Theoretic Construct for Assessing Information Flow in Networks. Journal of the American Society for Information Science, 30(1):25–32, January 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1981:MTI

[SWY81] G. Salton, H. Wu, and C. T. Yu. The measurement of term importance in automatic indexing. Journal of the American Society for Information Science, 32(3):175–186, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Salton:1975:TTI

[SYY75] G. Salton, C. S. Yang, and C. T. Yu. A theory of term importance in automatic text analysis. Journal of the American Society for Information Science, 26(1):33–44, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tagliacozzo:1972:BCS

[Tag72] Renata Tagliacozzo. Brief communications: Some relations between queries and search-terms generated by catalog users. Journal of the American Society for Information Science, 23(4):278– 280, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tagliacozzo:1975:CNI

[Tag75] Renata Tagliacozzo. The consumers of new information technology: a survey of the utilization of MEDLINE. Journal of the American Society for Information Science, 26(5):294–304, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tagliacozzo:1978:SSV

[Tag78] Renata Tagliacozzo. Some stylistic variations in scientific writing. Journal of the American Society for Information Science, 29(3): 136–140, May 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tague:1981:SBS

[Tag81] Jean Tague. The success-breeds-success phenomenon and bibliometric processes. Journal of the American Society for Information Science, 32(4):280–286, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tallman:1974:OPA

[Tal74] Johanna E. Tallman. Opinion paper. An affirmative statement on copyright debate. Journal of the American Society for Information Science, 25(3):145–150, May 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Talavage:1979:FAI

[Tal79] Joseph Talavage. Financial aspects of industrial information centers: a survey. Journal of the American Society for Information Science, 30(3):154–160, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Taniguchi:1996:RSA

[Tan96] Shoichi Taniguchi. Research: A system for analyzing cataloging rules: a feasibility study. Journal of the American Society for Information Science, 47(5):338–356, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Taniguchi:1999:RAO

[Tan99] Shoichi Taniguchi. Research: An analysis of orientedness in cataloging rules. Journal of the American Society for Information Science, 50(5):448–460, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tauber:1974:BRB

[Tau74] Stephen J. Tauber. Book review: The scientific and technical information contained in patent specifications — the extent and tie factors of its publication in other forms of literature. Felix Liebesny, James W. Hewitt, Paul S. Hunter and Maureen Hannah. The Polytechnic of North London School of Librarianship, 57 pp. (July 1973). Journal of the American Society for Information Science, 25(5):339, September 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tauber:1984:LES

[Tau84] Stephen J. Tauber. Letters to the Editor: The structure of Chinese characters. Journal of the American Society for Information Science, 35(6):372, November 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Taylor:1982:VAP

[Tay82] Robert S. Taylor. Value-added Processes in the Information Life Cycle. Journal of the American Society for Information Science, 33(5):341–346, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Taylor:1994:BRB

[Tay94] Robert S. Taylor. Book review: Organizational decision making and information. Journal of the American Society for Information Science, 45(4):284, May 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Taylor:1995:BRB

[Tay95] Robert S. Taylor. Book review: Scientific literature in wartime: the allied-German rivalry, 1939–1945. Journal of the American Society for Information Science, 46(3):241, April 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Triolo:1993:RDM

[TB93] Victor A. Triolo and Dachun Bao. Research: A decision model for technical journal deselection with an experiment in biomedical communications. Journal of the American Society for Information Science, 44(3):148–160, April 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Taghva:1994:REN

[TBCE94] Kazem Taghva, Julie Borsack, Allen Condit, and Srinivas Erva. Research: The effects of noisy data on text retrieval. Journal of the American Society for Information Science, 45(1):50–58, January 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tague:1973:OPI

[TC73] Jean Tague and John Carroll. Opinion Paper. Information in an informationless world or the making of an information scientist. Journal of the American Society for Information Science, 24(1): 40–44, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tolle:2000:CNP

[TC00a] Kristin M. Tolle and Hsinchun Chen. Comparing noun phrasing techniques for use with medical digital library tools. Journal of the American Society for Information Science, 51(4):352–370, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tombros:2000:UPR

[TC00b] Tassos Tombros and Fabio Crestani. Users’ perception of relevance of spoken documents. Journal of the American Society for Information Science, 51(10):929–939, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tebbutt:1970:OPB

[Teb70] Arthur V. Tebbutt. Opinion paper: Barter system for researchers: a proposal. Journal of the American Society for Information Science, 21(6):422–426, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tinker:1999:RCH

[TEC99] William Tinker, Benjamin Evers, and Paul N. Chance. Research: A case history in selective elimination as a solution to the information crises^{1}. Journal of the American Society for Information Science, 50(11):1040–1041, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Templeton:1994:LEN

[Tem94] Maxie Templeton. Letters to the Editor: The need for theoretical studies in information science. Journal of the American Society for Information Science, 45(8):567, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tenopir:1988:BRB

[Ten88] Carol Tenopir. Book review: Information online 86: Proceedings of the first Australian Online Information Conference. Journal of the American Society for Information Science, 39(1):25, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tenopir:1990:BRB

[Ten90] Carol Tenopir. Book review: Online reference and information retrieval. Journal of the American Society for Information Science, 41(2):152–153, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tenopir:1993:BRB

[Ten93] Carol Tenopir. Book review: Online information hunting. Journal of the American Society for Information Science, 44(6):365– 367, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tenopir:1994:BRB

[Ten94] Carol Tenopir. Book review: Online information retrieval: an introductory manual to principles and practice (4th edition). Journal of the American Society for Information Science, 45(3):219– 220, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tessier:1979:BRB

[Tes79] Judith A. Tessier. Book review: On-Line Information Retrieval Systems: An Introductory-Manual to Principles and Practice. Bernard Houghton and John Convey. London: Clive Bingley; 1977. Journal of the American Society for Information Science, 30(4):240, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Testemale:1986:BRB

[Tes86] Claudia Testemale. Book review: Fuzzy relational databases — a key to expert systems. Journal of the American Society for Information Science, 37(4):272–273, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Todeschini:1989:RES

[TF89] Claudio Todeschini and Michael P. Farrell. Research: An expert system for quality control in bibliographic databases. Journal of the American Society for Information Science, 40(1):1–11, January 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tversky:1994:SMM

[TFTB94] Barbara Tversky, Nancy Franklin, Holly A. Taylor, and David J. Bryant. Spatial mental models from descriptions. Journal of the American Society for Information Science, 45(9):656–668, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Turoff:1982:EJP

[TH82] Murray Turoff and Starr Roxanne Hiltz. The electronic journal: A progress report. Journal of the American Society for Information Science, 33(4):195–202, July 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tolle:1985:OSP

[TH85] John E. Tolle and Sehchang Hah. Online search patterns: NLM CATLINE database. Journal of the American Society for Information Science, 36(2):82–93, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thompson:1971:IDI

[Tho71] David A. Thompson. Interface design for an interactive information retrieval system: a literature survey and a research system description. Journal of the American Society for Information Science, 22(6):361–373, November 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thorpe:1972:BCE

[Tho72] Peter Thorpe. Brief communications: An evaluation of the rheumatology coverage of index medicus. A preliminary report. Journal of the American Society for Information Science, 23(6): 406, November 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thompson:1973:FAI

[Tho73a] Charles W. N. Thompson. The functions of abstracts in the initial screening of technical documents by the user. Journal of the American Society for Information Science, 24(4):270–276, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thompson:1973:BCC

[Tho73b] Harold B. Thompson. Brief communications: Check List for Information Environment. Journal of the American Society for Information Science, 24(1):70–71, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thompson:1979:LEI

[Tho79] Harold B. Thompson. Letter to the Editors: Indexing. Journal of the American Society for Information Science, 30(2):115, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thoma:1981:TIO

[Tho81] George R. Thoma. Transmission of information: an overview. Journal of the American Society for Information Science, 32(2): 131–140, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thomas:1982:MRS

[Tho82] Robert J. Thomas. Marketing research in the scientific and technical information services industry: Development and future directions. Journal of the American Society for Information Science, 33(5):265–269, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thompson:1990:RSA

[Tho90] Paul Thompson. Research: A sensitivity analysis of a probabilistic information retrieval system. Journal of the American Society for Information Science, 41(5):348–358, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thompson:2000:BR

[Tho00] Sara R. Thompson. Book reviews. Journal of the American Society for Information Science, 51(5):487–488, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tiamiyu:1989:RSS

[Tia89] Mutawakilu A. Tiamiyu. Research: Sub-Saharan Africa and the paperless society: a comment and a counterpoint. Journal of the American Society for Information Science, 40(5):325–328, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See [Old87].

Tibbo:1994:PLI

[Tib94] Helen R. Tibbo. Perspectives on … indexing: Indexing for the humanities. Journal of the American Society for Information Science, 45(8):607–619, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tilley:1970:BRB

[Til70] Glenda Tilley. Book review: 4/70-3R Conversion of Retrospective Catalog Records to Machine-Readable Form. 1969. U.S. Library of Congress. RECON Working Task Force. Library of Congress, Washington, D.C. 230 pp. Journal of the American Society for Information Science, 21(4):306, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tinker:1970:MFK

[Tin70] John F. Tinker. Microfilm finding-keys: Dual dictionary. Journal of the American Society for Information Science, 21(1):86–88, January 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Toms:1996:REE

[TK96] Elaine G. Toms and Mark T. Kinnucan. Research: The effectiveness of the electronic city metaphor for organizing the menus of Free-Nets. Journal of the American Society for Information Science, 47(12):919–931, December 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Terrant:1974:LEC

[TL74] S. W. Terrant and Ben-Ami Lipetz. Letters to the Editors: Citation of the literature by information scientists in their own publications. Journal of the American Society for Information Science, 25(1):72–73, January 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tenopir:1998:PLI

[TL98] Carol Tenopir and Lois F. Lunin. Perspective on … Internet issues/technological issues: Introduction and overview. Journal of the American Society for Information Science, 49(11):957–960, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Turner:1984:FST

[TO84] Stephen J. Turner and Gregory O’Brien. A fuzzy set theory approach to periodical binding decisions. Journal of the American Society for Information Science, 35(4):228–234, July 1984. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tocatlian:1970:TCP

[Toc70] Jacques J. Tocatlian. Are titles of chemical papers becoming more informative? Journal of the American Society for Information Science, 21(5):345–350, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Todd:1999:RUH

[Tod99] Ross J. Todd. Research: Utilization of heroin information by adolescent girls in Australia: a cognitive analysis. Journal of the American Society for Information Science, 50(1):10–23, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tomer:1992:PLI

[Tom92] Christinger Tomer. Perspectives on … information technology standards: Information technology standards for libraries. Journal of the American Society for Information Science, 43(8):566– 570, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tompson:1996:BRB

[Tom96] Sara Tompson. Book review: Managing Internet information services. Journal of the American Society for Information Science, 47(3):254–255, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tomer:1998:BRBa

[Tom98a] Christinger Tomer. Book review: Internet economics. Journal of the American Society for Information Science, 49(6):569–571, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tomer:1998:BRBb

[Tom98b] Christinger Tomer. Book review: Local area networking for the small library. 2nd ed. Journal of the American Society for Information Science, 49(11):1043–1044, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tompson:1999:BRB

[Tom99] Sara R. Tompson. Book review: Special libraries: a cataloging guide. Journal of the American Society for Information Science, 50(4):383–384, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Torok:1989:BRB

[Tor89] Andrew G. Torok. Book review: Human aspects of library automation: Helping staff and patrons cope. Journal of the American Society for Information Science, 40(2):135, March 1989. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Travica:1998:RIA

[Tra98] Bob Travica. Research: Information aspects of new organizational designs: Exploring the non-traditional organization. Journal of the American Society for Information Science, 49(13): 1224–1244, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tague-Sutcliffe:1992:MBC

[TS92] Jean Tague-Sutcliffe. In memoriam: B. C. (Bertie) Brookes, 1910–1991. Journal of the American Society for Information Science, 43(4):273–275, May 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tague-Sutcliffe:1996:SPE

[TS96] Jean M. Tague-Sutcliffe. Some perspectives on the evaluation of information retrieval systems. Journal of the American Society for Information Science, 47(1):1–3, January 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tsay:1998:LJU

[Tsa98] Ming-Yueh Tsay. Library journal use and citation half-life in medical science. Journal of the American Society for Information Science, 49(14):1283–1292, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tagliacozzo:1971:WRT

[TSK71] Renata Tagliacozzo, Dorothy Semmel, and Manfred Kochen. Written representation of topics and the production of query terms. Journal of the American Society for Information Science, 22(5):337–347, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tsuffis:1979:BRB

[Tsu79] M. L. Tsuffis. Book review: Current Research on Scientific and Technical Information Transfer. New York: Jeffrey Norton Publishers, Inc.; 32 pp., Microfiche 7 pp. Price: $12.95. Journal of the American Society for Information Science, 30(6):373, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tang:1999:RTI

[TSV99] Rong Tang, William M. Shaw, Jr., and Jack L. Vevea. Research: Towards the identification of the optimal number of relevance categories. Journal of the American Society for Information Science, 50(3):254–264, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Trueswell:1979:SCU

[TT79] Richard W. Trueswell and Stephen J. Turner. Simulating circulation-use characteristic curves using circulation data. Journal of the American Society for Information Science, 30(2):83–87, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Turner:1977:IMM

[Tur77] Stephen J. Turner. The identifier method of measuring use as applied to modeling the circulation use of books from a university library. Journal of the American Society for Information Science, 28(2):96–100, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Turner:1998:BRB

[Tur98] James M. Turner. Book review: Digital image access & retrieval. Journal of the American Society for Information Science, 49(8): 766–767, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tijssen:1990:RNC

[TvR90] R. J. W. Tijssen and A. F. J. van Raan. Research: Net citation balances: a measure of influence between scientific journals.

Journal of the American Society for Information Science, 41(4): 298–304, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tyno:1975:BCS

[Tyn75] Siarl Tyno. Brief communications: The Schlichtmeier gap: a report based on recent archeological findings. Journal of the American Society for Information Science, 26(4):259–262, July 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thom:1992:RMW

[TZ92] James A. Thom and Justin Zobel. Research: A model for word clustering. Journal of the American Society for Information Science, 43(9):616–627, October 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Thom:1993:LER

[TZ93] James Thom and Justin Zobel. Letters to the Editor: Rejoinder to letter from Stephen P. Harter. Journal of the American Society for Information Science, 44(4):213, May 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Tzvieli:1990:BRB

[Tzv90] Arie Tzvieli. Book review: Possibility theory: an approach to computerized processing of uncertainty. Journal of the American Society for Information Science, 41(2):153–154, March 1990. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Usdin:1982:LE

[UAW82] B. Tommie Usdin, Robert J. Allan, and Ben H. Weil. Letters to the Editor. Journal of the American Society for Information Science, 33(5):347–348, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Underwood:1996:BRB

[Und96] Peter G. Underwood. Book review: Critical issues in systems theory and practice. Journal of the American Society for Information Science, 47(9):717–719A, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Underwood:1998:BRBa

[Und98a] Peter G. Underwood. Book review: Implementation of organizational innovation: Studies of academic and research libraries.

Journal of the American Society for Information Science, 49(2): 186–187, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Underwood:1998:BRBb

[Und98b] Peter G. Underwood. Book review: Issues and applications of case-based reasoning in design. Journal of the American Society for Information Science, 49(8):762–764, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Usdin:1985:BRB

[Usd85] B. T. Usdin. Book review: ALA glossary of library and information science. Chicago, IL: American Library Association; 1983: 245 pp., $50. Journal of the American Society for Information Science, 36(4):272–273, July 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vallee:1973:IOI

[VA73] Jacques F. Vallee and Gerald L. Askevold. Information organization for interactive use: Design implications in data-base systems. Journal of the American Society for Information Science, 24(4): 287–299, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vagianos:1980:BRB

[Vag80] Louis G. Vagianos. Book review: Electronic Delivery of Documents and Graphics. Daniel M. Costigan. New York: Van Nostrand Reinhold; 1978: xvii, 344 pp. (illustrated index; appendix). ISBN 0-442-80036-3. LC 78-8166. Price: $21.95. Journal of the American Society for Information Science, 31(5):381, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanRyzin:1980:LEE

[Van80] Elizabeth Van Ryzin. Letters to the Editor: Energy technology — a correction. Journal of the American Society for Information Science, 31(6):454, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanDeBogart:1990:BRB

[Van90] Williard Van De Bogart. Book review: Information science: an integrated view. Journal of the American Society for Information Science, 41(2):151–152, March 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanderVeerMartens:1997:BRB

[Van97a] Betsy Van der Veer Martens. Book review: The scholar’s courtesy: the role of acknowledgement in the primary communication process. Journal of the American Society for Information Science, 48(3):279–280, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanHooydonk:1997:BCF

[Van97b] G. Van Hooydonk. Brief communications: Fractional counting of multiauthored publications: Consequences for the impact of authors. Journal of the American Society for Information Science, 48(10):944–945, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanHooydonk:1998:LER

[Van98a] G. Van Hooydonk. Letters to the Editor: Rejoinder. Journal of the American Society for Information Science, 49(5):482–483, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanHooydonk:1998:RSR

[Van98b] G. Van Hooydonk. Research: Standardizing relative impacts: Estimating the quality of research from citation counts. Journal of the American Society for Information Science, 49(10):932–941, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanderVeerMartens:1999:BNR

[Van99] Betsy Van der Veer Martens. Biographical note — Robert S. Taylor. Journal of the American Society for Information Science, 50(12):1109–1110, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vara:1975:BRBa

[Var75a] Albert C. Vara. Book review: Information and the small manufacturer: Report of a Survey of the Information Needs of Small Manufacturers in New South Wales. Carmel Maguire and Robin Kench., the University of New South Wales, Australia, 96 pp.

(1974). Journal of the American Society for Information Science, 26(1):66, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vara:1975:BRBb

[Var75b] Albert C. Vara. Book review: Selected readings on information for industry, FID (international federation for documentation) study committee, “information for industry” (FID/II), FID Publ. 518, Copenhagen: 60 pp. (1974). Journal of the American Society for Information Science, 26(5):308, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vara:1978:BRB

[Var78] Albert C. Vara. Book review: Directory Information Service. Detroit, MI: Gale Research Co.; 1977: over 300 pages per year. Three issues per year. Price: $36.00/yr. Journal of the American Society for Information Science, 29(6):318–319, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Varian:1997:BRB

[Var97] Hal R. Varian. Book review: The economics of communication and information. Journal of the American Society for Information Science, 48(12):1151–1152, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Varlejs:1999:RCP

[Var99] Jana Varlejs. Research: The continuing professional education role of ASIS: Fifty years of learning together, reaching out, seeking identity. Journal of the American Society for Information Science, 50(11):1032–1036, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

vanBlokland:1981:LEE

[vB81] G. G. van Blokland. Letters to the Editor: Efforts to alleviate information inequity. Journal of the American Society for Information Science, 32(3):236, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Verity:1970:CBS

[VC70] John B. Verity and Elizabeth L. Crocker. A Computer-Based System for Serial Records at Lawrence Radiation Laboratory. Journal of the American Society for Information Science, 21(4):

247–253, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VanGyseghem:1998:IUU

[VD98] Nancy Van Gyseghem and Rita De Caluwe. Imprecision and uncertainty in the UFO database model. Journal of the American Society for Information Science, 49(3):236–252, March 1998. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

vandenBesselaar:1996:RMC

[vdBL96] Peter van den Besselaar and Loet Leydesdorff. Research: Mapping change in scientific specialties: a scientometric reconstruction of the development of artificial intelligence. Journal of the American Society for Information Science, 47(6):415–436, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

vanderVet:1999:RCQ

[vdVM99] Paul E. van der Vet and Nicolaas J. I. Mars. Research: Condorcet query engine: a query engine for coordinated index terms. Journal of the American Society for Information Science, 50(6): 485–492, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Veaner:1990:LER

[Vea90] Allen B. Veaner. Letters to the Editor: Re: Peter Hirtle’s note on Atherton Seidell. Journal of the American Society for Information Science, 41(6):427, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Veith:1980:ITD

[Vei80] Richard H. Veith. Informatics and transborder data flow: the question of social impact. Journal of the American Society for Information Science, 31(2):105–110, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Veith:1988:BRB

[Vei88] Richard H. Veith. Book review: Review of “Multinational Computer Nets”. Journal of the American Society for Information Science, 39(1):29, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vizine-Goetz:1993:BRB

[VG93] Diane Vizine-Goetz. Book review: Subject analysis in online catalogs. Journal of the American Society for Information Science, 44(10):593, December 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

VandenBerghe:1998:BIU

[VHdB^{+}98] Herman Van den Berghe, Josee A. Houben, Renger E. de Bruin, Henk F. Moed, André Kint, Marc Luwel, and Eric H. J. Spruyt. Bibliometric indicators of university research performance in Flanders. Journal of the American Society for Information Science, 49(1):59–67, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vigil:1983:POS

[Vig83] Peter J. Vigil. The psychology of online searching. Journal of the American Society for Information Science, 34(4):281–287, July 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vinken:1974:OPD

[Vin74] Pierre J. Vinken. Opinion paper. Developments in scientific documentation in the long term. Journal of the American Society for Information Science, 25(2):109–112, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vinkler:1999:LE

[Vin99] P. Vinkler. Letter to the Editor. Journal of the American Society for Information Science, 50(10):958, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Virgo:1970:EIM

[Vir70] Julie A. Virgo. An evaluation of Index Medicus and MEDLARS in the field of ophthalmology. Journal of the American Society for Information Science, 21(4):254–263, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Voiskunskii:1997:RES

[Voi97] Vladimir G. Voiskunskii. Research: Evaluation of search results: a new approach. Journal of the American Society for Information Science, 48(2):133–142, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Voos:1974:BCL

[Voo74] Henry Voos. Brief communications: Lotka and information science. Journal of the American Society for Information Science, 25(4):270–272, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Voos:1980:BCI

[Voo80] Henry Voos. Brief communication: Implications of holography for information systems. Journal of the American Society for Information Science, 31(6):449–451, November 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Voorbij:1999:RSS

[Voo99] Henk J. Voorbij. Research: Searching scientific information on the Internet: A Dutch academic user survey. Journal of the American Society for Information Science, 50(7):598–615, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

vanRaan:1998:I

[vR98] Anthony F. J. van Raan. Introduction. Journal of the American Society for Information Science, 49(1):5–6, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

vanRijsbergen:1996:RIC

[vRL96] C. J. van Rijsbergen and M. Lalmas. Research: Information calculus for information retrieval. Journal of the American Society for Information Science, 47(5):385–398, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vleduts-Stokolov:1987:RCR

[VS87] Natasha Vleduts-Stokolov. Research: Concept recognition in an automatic text-processing system for the life sciences. Journal of the American Society for Information Science, 38(4):269–287, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vaughan:1999:RJI

[VS99] Misha W. Vaughan and Nancy Schwartz. Research: Jumpstarting the information design for a community network. Journal of the American Society for Information Science, 50(7):588–597, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Vaughan:1997:RMI

[VTS97] Liwen Qiu Vaughan and Jean Tague-Sutcliffe. Research: Measuring the impact of information on development: A LISREL-based study of small businesses in Shanghai. Journal of the American Society for Information Science, 48(10):917–931, October 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1979:ULC

[WA79] Edward S. Warner and Anita L. Anker. Utilizing library constituents’ perceived needs in allocating journal costs. Journal of the American Society for Information Science, 30(6):325–329, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wood:1996:DMC

[WA96] William H. Wood III and Alice M. Agogino. Development of modular curricular materials: Engineering courseware content and delivery: The NEEDS infrastructure for distance independent education. Journal of the American Society for Information Science, 47(11):863–869, November 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wood:1997:LEI

[WA97] Elizabeth H. Wood and Joan S. Ash. Letters to the Editor: IAIMS perspectives. Journal of the American Society for Information Science, 48(12):1153, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Walter:1972:LEI

[Wal72] Pat. L. Walter. Letters to the Editor: Information Service Affects Communication. Journal of the American Society for Information Science, 23(2):136, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wall:1975:SDT

[Wal75] Eugene Wall. Symbiotic development of thesauri and information systems: a case history. Journal of the American Society for Information Science, 26(2):71–79, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Walker:1981:OUI

[Wal81] Donald E. Walker. The organization and use of information: Contributions of information science, computational linguistics and artificial intelligence. Journal of the American Society for Information Science, 32(5):347–363, September 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wallace:1985:USM

[Wal85] Danny P. Wallace. The use of statistical methods in library and information science. Journal of the American Society for Information Science, 36(6):402–410, November 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wall:1986:ROM

[Wal86a] Eugene Wall. Research: An objective method of “weeding” bibliographic databases. Journal of the American Society for Information Science, 37(4):256–260, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wallace:1986:BRBb

[Wal86b] Danny P. Wallace. Book review: Integrated online library systems: Principles, planning and implementation. Journal of the American Society for Information Science, 37(4):275–276, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wallace:1986:BRBa

[Wal86c] Danny P. Wallace. Book review: The library manager’s guide to automation. Journal of the American Society for Information Science, 37(4):271–272, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wallace:1986:RRB

[Wal86d] Danny P. Wallace. Research: The relationship between journal productivity and obsolescence. Journal of the American Society for Information Science, 37(3):136–145, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wallace:1987:EST

[Wal87] Danny P. Wallace. Editorial: Some thoughts on book reviewing. Journal of the American Society for Information Science, 38(4):

227, July 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Walker:1997:LBD

[Wal97] Thomas D. Walker. Literature: Journal of Documentary Reproduction, 1938–1942: Domain as reflected in characteristics of authorship and citation. Journal of the American Society for Information Science, 48(4):361–368, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Walsham:1998:ICP

[Wal98] Geoff Walsham. IT and changing professional identity: Microstudies and macro-theory. Journal of the American Society for Information Science, 49(12):1081–1089, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Walker:2000:BRB

[Wal00a] Billie E. Walker. Book review: The Internet public library handbook, by Joseph Janes et al. Journal of the American Society for Information Science, 51(11):1064–1065, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Walsh:2000:NPS

[Wal00b] Peter Walsh. The neon paintbrush: Seeing, technology, and the museum as metaphor. Journal of the American Society for Information Science, 51(1):39–48, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wanner:1981:CIP

[Wan81] Eric Wanner. Current issues in psycholinguistics. Journal of the American Society for Information Science, 32(6):405–411, November 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wang:1992:RMI

[Wan92] Chih Wang. Research: A Micronesian information system: an application of EATPUT model. Journal of the American Society for Information Science, 43(9):594–601, October 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ware:1973:GSM

[War73] Glenn O. Ware. A general statistical model for estimating future demand levels of data-base utilization within an information retrieval organization. Journal of the American Society for Information Science, 24(4):261–264, July 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1979:BRB

[War79] Edward S. Warner. Book review: Development of a Responsive Library Acquisitions Formula. Glyn T. Evans, Mary H. Beilby, and Roger Gifford. Albany, New York: Suny Central Administration, Office of Library Services; December 1978: 95 pp. (Final Report on Usoe Grant No. g007702305). Price: Not Given. Journal of the American Society for Information Science, 30(6):373– 374, November 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1981:IIA

[War81] Edward S. Warner. The impact of interlibrary access to periodicals on subscription continuation/cancellation decision making. Journal of the American Society for Information Science, 32(2): 93–95, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1988:PLI

[War88] Homer R. Warner. Perspectives on … Integrated Academic Information Management Systems (IAIMS): Clinical linkages: IAIMS at the University of Utah. Journal of the American Society for Information Science, 39(2):135–137, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1991:RQQ

[War91] Amy J. Warner. Research: Quantitative and qualitative assessments of the impact of linguistic theory on information science. Journal of the American Society for Information Science, 42(1): 64–71, January 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1993:RWL

[War93] Julian Warner. Research: Writing and literary work in copyright: a binational and historical analysis. Journal of the American Society for Information Science, 44(6):307–321, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1996:BRBa

[War96a] Julian Warner. Book review: Electric words: Dictionaries, computers, and meanings. Journal of the American Society for Information Science, 47(10):791–792, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1996:BRBb

[War96b] Julian Warner. Book review: Electric words: Dictionaries, computers, and meanings. Journal of the American Society for Information Science, 47(10):791–792A, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1998:BRBa

[War98a] Julian Warner. Book review: Borders in cyberspace: Information policy and the global information infrastructure. Journal of the American Society for Information Science, 49(4):387–388, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1998:BRBb

[War98b] Julian Warner. Book review: The soft edge: a natural history and future of the information revolution. Journal of the American Society for Information Science, 49(11):1041, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1999:IVH

[War99a] Julian Warner. An information view of history. Journal of the American Society for Information Science, 50(12):1125–1126, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Warner:1999:RIS

[War99b] Julian Warner. Research: Information society or cash nexus? A study of the United States as a copyright haven. Journal of the American Society for Information Science, 50(5):461–470, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wasser:1974:ANI

[Was74] Nigel Wasser. An alpha-numeric information storage and retrieval system. Journal of the American Society for Information Science, 25(6):374–377, November 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wasserman:1980:ISD

[Was80] Anthony I. Wasserman. Information system design methodology. Journal of the American Society for Information Science, 31(1): 1–24, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Watters:1989:RLF

[Wat89] C. R. Watters. Research: Logic framework for information retrieval. Journal of the American Society for Information Science, 40(5):311–324, September 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Watson:1998:PLI

[Wat98] Jinx Stapleton Watson. Perspectives on … Internet issues/ information use issues: “If you don’t have it, you can’t find it.” A close look at students’ perceptions of using technology. Journal of the American Society for Information Science, 49(11):1024–1036, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Watters:1999:RIR

[Wat99] Carolyn Watters. Research: Information retrieval and the virtual document. Journal of the American Society for Information Science, 50(11):1028–1029, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wong:1998:ADS

[WBX98] S. K. M. Wong, C. J. Butz, and Y. Xiang. Automated database schema design using mined data dependencies. Journal of the American Society for Information Science, 49(5):455–470, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilde:1975:RDG

[WC75] Daniel U. Wilde and H. David Chafe. Reduction of data generation costs. Journal of the American Society for Information Science, 26(2):103–107, March 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wenger:1977:JEL

[WC77] Charles B. Wenger and Judith Childress. Journal evaluation in a large research library. Journal of the American Society for Information Science, 28(5):293–299, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1984:MCC

[WC84] Howard D. White and Karen Calhoun. Mapping a curriculum by computer. Journal of the American Society for Information Science, 35(2):82–89, March 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1985:LEA

[WC85] Howard D. White and Karen Calhoun. Letters to the Editors: Authors’ Reply. Journal of the American Society for Information Science, 36(5):353–354, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weiss:1992:PLI

[WC92] Martin Weiss and Carl Cargill. Perspectives on … information technology standards: Consortia in the standards development process. Journal of the American Society for Information Science, 43(8):559–565, September 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weissberg:1987:PLF

[WCL87] Alfred Weissberg, Joseph F. Caponio, and Lois F. Lunin. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: The Neurological Information Network. Journal of the American Society for Information Science, 38(1):52–59, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wagner-Dobler:1997:RST

[WD97a] Roland Wagner-Döbler. Research: Science-technology coupling: the case of mathematical logic and computer science. Journal of the American Society for Information Science, 48(2):171–183, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wolfram:1997:BCP

[WD97b] Dietmar Wolfram and Alexandra Dimitroff. Brief communications: Preliminary findings on searcher performance and perceptions of performance in a hypertext bibliographic retrieval system. Journal of the American Society for Information Science, 48(12):1142–1145, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wildemuth:1995:RMS

[WdBFF95] Barbara M. Wildemuth, Ruth de Bliek, Charles P. Friedman, and Dean D. File. Research: Medical students’ personal knowledge, searching proficiency, and database use in problem solving. Journal of the American Society for Information Science, 46(8): 590–607, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weedman:1992:RIF

[Wee92] Judith Weedman. Research: Informal and formal channels in boundary-spanning communication. Journal of the American Society for Information Science, 43(3):257–267, April 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weedman:1999:RCC

[Wee99] Judith Weedman. Research: Conversation and community: the potential of electronic conferences for creating intellectual proximity in distributed learning environments. Journal of the American Society for Information Science, 50(10):907–928, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weil:1970:SWA

[Wei70] Ben H. Weil. Standards for writing abstracts. Journal of the American Society for Information Science, 21(5):351–357, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinert:1973:KSS

[Wei73] Hartmut H. Weinert. A keyword system for small information centers. Journal of the American Society for Information Science, 24(3):180–192, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weiner:1976:BRB

[Wei76] Betty Weiner. Book review: World guide to technical information and documentation services, 2d ed. Paris, France: UNESCO, 515 pp. (1975). Journal of the American Society for Information Science, 27(3):189–190, May 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weilisch:1979:BRB

[Wei79] Hans H. Weilisch. Book review: Library Lit. 8 — The Best of 1977. Bill Katz, Ed. Metuchen, NJ: Scarecrow Press; 1978: 333 pp. ISBN: 0–8108–1125–1. Price: $12.00. Journal of the American Society for Information Science, 30(4):238–239, July 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1988:BRB

[Wei88] Bella Hass Weinberg. Book review: Macrothesaurus for information processing in the field of economic and social development. Journal of the American Society for Information Science, 39(1): 23–24, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1990:LER

[Wei90] Bella Hass Weinberg. Letters to the Editor: Regarding Michael Buckland’s article. Journal of the American Society for Information Science, 41(6):428, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1991:BRB

[Wei91] Bella Hass Weinberg. Book review: Introduction to indexing and abstracting. Journal of the American Society for Information Science, 42(7):532–536, August 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1993:BRBb

[Wei93a] Bella Hass Weinberg. Book review: Automated support to indexing. Journal of the American Society for Information Science, 44(2):119–121, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1993:BRBa

[Wei93b] Bella Hass Weinberg. Book review: Subject indexing: a introductory guide. Journal of the American Society for Information Science, 44(2):117–119, March 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1994:BRBb

[Wei94a] Bella Hass Weinberg. Book review: Challenges in indexing electronic text and images. Journal of the American Society for Information Science, 45(9):718–723, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1994:BRBa

[Wei94b] Bella Hass Weinberg. Book review: File management and information retrieval systems: a manual for managers and technicians (3rd ed.). Journal of the American Society for Information Science, 45(5):310–314, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1995:BRB

[Wei95a] Bella Hass Weinberg. Book review: Indexing books. Journal of the American Society for Information Science, 46(1):65–73, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1995:DBR

[Wei95b] Bella Hass Weinberg. In-depth book review: Art & Architecture Thesaurus (2nd ed.); and Guide to Indexing and Cataloging with the Art & Architecture Thesaurus. Journal of the American Society for Information Science, 46(2):152–160, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1996:BRB

[Wei96a] Bella Hass Weinberg. Book review: Computational linguistics in information science: Information retrieval (full-text or conceptual), automatic indexing, text abstraction, content analysis, information extraction, query languages: Bibliography. Journal of the American Society for Information Science, 47(3):247–249, March 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weingand:1996:BRB

[Wei96b] Darlene E. Weingand. Book review: Software and intellectual property protection: Copyright and patent issues for computer and legal professionals. Journal of the American Society for Information Science, 47(6):480–482A, June 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weibel:1997:MTY

[Wei97a] Stuart Weibel. In memoriam: A tribute to Yuri Rubinsky, August 2, 1952–January 21, 1996. Journal of the American Society for Information Science, 48(7):583, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weinberg:1997:DPT

[Wei97b] Bella Hass Weinberg. Development of particular techniques: The earliest Hebrew citation indexes. Journal of the American Society for Information Science, 48(4):318–330, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weisgerber:1997:DPT

[Wei97c] David W. Weisgerber. Development of particular techniques: Chemical Abstracts Service Chemical Registry System: History, scope, and impacts. Journal of the American Society for Information Science, 48(4):349–360, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1970:BRB

[Wel70] H. Wellisch. Book review: 5/70-1R An introduction to UDC. 1969. Jean M. Perreault. Archon Books & Clive Bingley, Hamden, Conn. Journal of the American Society for Information Science, 21(5):375–376, September 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1971:BRB

[Wel71] H. Wellisch. Book review: 1/71-1R Sacherschliessung an Bibliotheken bei Einsatz elektronischer Datenverarbeitung. (Subject Retrieval in Libraries with the Aid of EDP) 1970. Dietrich Roessler. Köln, Greven Verlag. 117 pp. (Arbeiten aus dem Bibliothekar-Lehrinstitut des Landes Rheinland-Westfalen, Heft 34.). Journal of the American Society for Information Science, 22(1):66–67, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1972:FCI

[Wel72] Hans Wellisch. A flow chart for indexing with a thesaurus. Journal of the American Society for Information Science, 23(3):185– 194, May 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Welsh:1975:BRB

[Wel75] Harry E. Welsh. Book review: Policies and practices in the bibliographic control of United States Government Publications, C. P. Crowers, issue editor, Drexel Library Quarterly 10 (Nos. 1 and 2): 153 pp. (1974). Journal of the American Society for Information Science, 26(3):191, May 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1978:BRB

[Wel78] Hans H. Wellisch. Book review: Bibliographic services throughout the world 1970–74. Marcelle Beaudiquez, ed. Paris: Unesco; 1977. 419 p. Journal of the American Society for Information Science, 29(6):317, November 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1980:BCB

[Wel80a] Hans Wellisch. Brief communication: Journal of the ASIS vs. Nachrichten für Dokumentation: A comparative analysis of citations. Journal of the American Society for Information Science, 31(3):216–218, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1980:BRB

[Wel80b] Hans H. Wellisch. Book review: Indexing Concepts and Methods. Harold Borko and Charles L. Bernier. New York: Academic; 1978: 256 pp. ISBN 0-12-118660-1. Price: $14.00. Journal of the American Society for Information Science, 31(2):126–127, March 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1980:CBC

[Wel80c] Hans H. Wellisch. The cybernetics of bibliographic control: Toward a theory of document retrieval systems. Journal of the American Society for Information Science, 31(1):41–50, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1981:LEI

[Wel81] Hans H. Wellisch. Letter to the Editor: Informatics or cybics? Journal of the American Society for Information Science, 32(4):

318, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1983:AFR

[Wel83] Hans H. Wellisch. The ALA filing rules: Flowcharts illustrating their application, with a critique and suggestions for improvement. Journal of the American Society for Information Science, 34(5):313–330, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1985:E

[Wel85] Hans H. Wellisch. Erratum. Journal of the American Society for Information Science, 36(5):354, September 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1987:LET

[Wel87] Hans H. Wellisch. Letters to the Editor: Toward expert systems. Journal of the American Society for Information Science, 38(3): 216–217, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1994:PLI

[Wel94] Hans H. Wellisch. Perspectives on … indexing: Book and periodical indexing. Journal of the American Society for Information Science, 45(8):620–627, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weller:2000:EPR

[Wel00] Ann C. Weller. Editorial peer review for electronic journals: Current issues and emerging models. Journal of the American Society for Information Science, 51(14):1328–1333, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Westler:1987:PLF

[Wes87] Jean A. Westler. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Section III. Health information analysis centers, clearinghouses, and other special projects. Federal health information clearinghouses. Journal of the American Society for Information Science, 38(1): 48–51, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

West:1988:PLI

[Wes88] Richard T. West. Perspectives on … Integrated Academic Information Management Systems (IAIMS): Programmatic intent of IAIMS — past and future. The National Library of Medicine’s IAIMS Grant Program: Experiences and futures. Journal of the American Society for Information Science, 39(2):142–145, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Westbrook:1989:BRB

[Wes89] J. H. Westbrook. Book review: Data bases and data models. Journal of the American Society for Information Science, 40(6): 436–437, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Westland:1991:REC

[Wes91] J. Christopher Westland. Research: Economic constraints in hypertext. Journal of the American Society for Information Science, 42(3):178–184, April 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Westbrook:2000:CAC

[Wes00] Lynn Westbrook. Conference announcement: Conference announcement: Library research seminar II Partners and connections: Research and practice University of Maryland; November 2-3, 2001. Journal of the American Society for Information Science, 51(11):1066, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wan:1997:REA

[WEWP97] Tian-Long Wan, Martha Evens, Yeun-Wen Wan, and Yuen-Yuan Pao. Research: Experiments with automatic indexing and a relational thesaurus in a Chinese information retrieval system. Journal of the American Society for Information Science, 48(12): 1086–1096, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wood:1972:OLJ

[WFK72] James L. Wood, Carolyn Flanagan, and H. E. Kennedy. Overlap in the lists of journals monitored by Biosis, CAS, and Ei. Journal of the American Society for Information Science, 23(1):36–38, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wood:1973:OAJ

[WFK73] James L. Wood, Carolyn Flanagan, and H. Edward Kennedy. Overlap among the journal articles selected for coverage by BIO-SIS, CAS, and Ei. Journal of the American Society for Information Science, 24(1):25–28, January 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1981:ACL

[WG81] Howard D. White and Belver C. Griffith. Author cocitation: a literature measure of intellectual structure. Journal of the American Society for Information Science, 32(3):163–171, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wright:1970:BRB

[WH70] Keith Wright and Theodore C. Hines. Book review: 4/70-1R Isotopes Information Center Keyword Thesaurus. July, 1969. Isotopes Information Center Staff, Robert H. Lafferty, Jr., Editor. Oak Ridge National Laboratory, Oak Ridge, Tenn. 98 pp. Journal of the American Society for Information Science, 21(4):305, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1991:RIF

[WH91] Arden White and Nelda Rae Hernandez. Research: Increasing field complexity revealed through article title analyses. Journal of the American Society for Information Science, 42(10):731–734, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Witbrock:1998:SRD

[WH98] Michael J. Witbrock and Alexander G. Hauptmann. Speech recognition for a digital video library. Journal of the American Society for Information Science, 49(7):619–632, May 15, 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Whatley:1983:IPH

[Wha83] Allan Whatley. Invited papers: 1984 — a hoax? Journal of the American Society for Information Science, 34(1):86–90, January 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Whitehead:1976:BCR

[Whi76] Daniel L. Whitehead, Jr. Brief communication: Representation of the change of state in an individual resulting from data acquired and processed. Journal of the American Society for Information Science, 27(2):122, March 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1977:TNI

[Whi77] Howard D. White. Toward a national information system for social science data files. Journal of the American Society for Information Science, 28(6):313–322, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1981:CAR

[Whi81] Howard D. White. Cocited author retrieval online: an experiment with the social indicators literature. Journal of the American Society for Information Science, 32(1):16–21, January 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1983:CMS

[Whi83] Howard D. White. A cocitation map of the social indicators movement. Journal of the American Society for Information Science, 34(5):307–312, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1987:HNR

[Whi87] Herbert S. White. Historical note: Random ruminations from the groves of academe. Journal of the American Society for Information Science, 38(5):336–337, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Whitbeck:1989:BRB

[Whi89a] George Whitbeck. Book review: Internationalizing library and information science education: a handbook of policies and procedure in administration and curriculum. Journal of the American Society for Information Science, 40(6):441, November 1989. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1989:BRB

[Whi89b] Herbert S. White. Book review: Education of library and information professionals. Journal of the American Society for Information Science, 40(6):437–438, November 1989. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1991:BCF

[Whi91] Arden White. Brief communications: A further exploration of title size and author number. Journal of the American Society for Information Science, 42(5):384–385, June 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1994:LEC

[Whi94] Herbert S. White. Letters to the Editor: Citation-based auditing of academic performance. Journal of the American Society for Information Science, 45(8):567, September 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1996:LEP

[Whi96] Herbert S. White. Letters to the Editor: Perception studies and the assessment of graduate LIS programs. Journal of the American Society for Information Science, 47(9):726, September 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1998:BRBa

[Whi98a] Jeff White. Book review: Coordinating technology: Studies in the international standardization of telecommunications. Journal of the American Society for Information Science, 49(11):1039–1041, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1998:BRBb

[Whi98b] Jeff White. Book review: Readings in agents. Journal of the American Society for Information Science, 49(14):1330–1332, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1998:BRBc

[Whi98c] Jeff White. Book review: Research misconduct: Issues, implications, and strategies. Journal of the American Society for Information Science, 49(14):1334–1336, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Whitney:1998:BRB

[Whi98d] Gretchen Whitney. Book review: Computer-mediated communication: Linguistic, social, and cross-cultural perspectives. Journal of the American Society for Information Science, 49(9):859–860, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1999:SPW

[Whi99a] Howard D. White. Scientist-poets wanted. Journal of the American Society for Information Science, 50(12):1052–1053, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1999:BRB

[Whi99b] Jeff White. Book review: Ink into bits: a web of converging media, by Charles T. Meadow. Journal of the American Society for Information Science, 50(7):629–630, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Whitmire:2000:BRB

[Whi00] Ethelene Whitmire. Book review: Information seeking in the online age: Principles and practice, by Andrew Large, Lucy A. Tedd, and R. J. Hartley. Journal of the American Society for Information Science, 51(10):964, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wiedenbeck:1986:ROP

[Wie86] Susan Wiedenbeck. Research: Organization of programming knowledge of novices and experts. Journal of the American Society for Information Science, 37(5):294–299, September 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wiggins:1995:POC

[Wig95] Gary Wiggins. Perspectives on … the Chemist’s workstation: CHMINF-L: the chemical information sources discussion list. Journal of the American Society for Information Science, 46(8): 614–617, September 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1970:LE

[Wil70] John H. Wilson, Jr. Letters to the Editor. Journal of the American Society for Information Science, 21(4):303–304, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1971:FMM

[Wil71] J. H. Williams, Jr. Functions of a man-machine interactive information retrieval system. Journal of the American Society for Information Science, 22(5):311–317, September 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1975:BCP

[Wil75] Martha E. Williams. Brief communications: Progress and Problems of the Data Base Community. Journal of the American Society for Information Science, 26(5):305–306, September 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1977:NLD

[Wil77a] Martha E. Williams. Networks for on-line data base access. Journal of the American Society for Information Science, 28(5):247– 253, September 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1977:OPD

[Wil77b] Martha E. Williams. Opinion paper: Data bases — a history of developments and trends from 1966 through 1975. Journal of the American Society for Information Science, 28(2):71–78, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1979:BCU

[Wil79] Patrick Wilson. Brief communications: Utility-theoretic indexing. Journal of the American Society for Information Science, 30(3):169–170, May 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Willett:1982:ERL

[Wil82] Peter Willett. European research letter: Some Current Information Retrieval Research in the United Kingdom. Journal of the American Society for Information Science, 33(5):262–264, September 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Willett:1984:BCN

[Wil84] Peter Willett. Brief communication: A note on the use of nearest neighbors for implementing single linkage document classifications. Journal of the American Society for Information Science, 35(3):149–152, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Willett:1985:E

[Wil85a] P. Willett. Erratum. Journal of the American Society for Information Science, 36(3):210, May 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1985:BRB

[Wil85b] T. D. Wilson. Book review: The competitive strength of the information and communication industry in Europe. H. J. van Houten, Ed. The Hague: Martinus Nijhoff; 1983: 173 pp. (ISBN 90-247-2860-6). Journal of the American Society for Information Science, 36(1):56, January 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1986:PLO

[Wil86] Martha E. Williams. Perspectives on … online searching: Transparent information systems through gateways, front ends, intermediaries, and interfaces. Journal of the American Society for Information Science, 37(4):204–214, July 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilbur:1992:RRT

[Wil92a] W. John Wilbur. Research: Retrieval testing by the comparison of statistically independent retrieval methods. Journal of the American Society for Information Science, 43(5):358–370, June 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wildemuth:1992:REG

[Wil92b] Barbara M. Wildemuth. Research: An empirically grounded model of the adoption of intellectual technologies. Journal of the American Society for Information Science, 43(3):210–224, April 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilbur:1993:RRT

[Wil93a] W. John Wilbur. Research: Retrieval testing with hypergeometric document models. Journal of the American Society for Information Science, 44(6):340–351, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1993:RCE

[Wil93b] Patrick Wilson. Research: Communication efficiency in research and development. Journal of the American Society for Information Science, 44(7):376–382, August 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1994:BRB

[Wil94] Patrick Wilson. Book review: The barefoot expert: the interface of computerized knowledge systems and indigenous knowledge systems. Journal of the American Society for Information Science, 45(3):220–221, April 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1995:BRB

[Wil95a] Patrick Wilson. Book review: Knowledge-based systems for general reference work: Applications, problems, and progress. Journal of the American Society for Information Science, 46(10):792– 793, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1995:RUR

[Wil95b] Patrick Wilson. Research: Unused relevant information in research and development. Journal of the American Society for Information Science, 46(1):45–51, January 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1997:DSL

[Wil97] Robert V. Williams. The documentation and special libraries movements in the united states, 1910–1960. Journal of the American Society for Information Science, 48(9):775–781, September 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilbur:1998:CGI

[Wil98a] W. John Wilbur. A comparison of group and individual performance among subject experts and untrained workers at the document retrieval task. Journal of the American Society for Information Science, 49(6):517–529, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilkins:1998:BRB

[Wil98b] Janie L. Hassard Wilkins. Book review: Books, bricks & bytes: Libraries in the Twenty-First Century. Journal of the American Society for Information Science, 49(14):1336–1338, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1998:BRB

[Wil98c] Robert D. Wilson. Book review: Information graphics: a comprehensive illustrated reference. Journal of the American Society for Information Science, 49(4):383–384, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilensky:2000:RDL

[Wil00a] Robert Wilensky. Research: Digital library resources as a basis for collaborative work. Journal of the American Society for Information Science, 51(3):228–245, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See erratum [Ano08].

Wilkins:2000:BRB

[Wil00b] Janie L. Hassard Wilkins. Book review: Internet policy handbook for libraries, by Mark Smith. Journal of the American Society for Information Science, 51(8):788–789, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:2000:BRB

[Wil00c] Hugh E. Williams. Book review: Information retrieval: Algorithms and heuristics, by David A. Grossman and Ophir Frieder. Journal of the American Society for Information Science, 51(11): 1063–1064, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Windsor:1970:FSC

[Win70a] Donald A. Windsor. A format for the standardized citation of both published and unpublished documents. Journal of the American Society for Information Science, 21(3):195–197, May 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Winings:1970:DSI

[Win70b] J. W. Winings. A data structure for information retrieval. Journal of the American Society for Information Science, 21(2):145– 148, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Windsor:1995:LEA

[Win95] Donald A. Windsor. Letters to the Editor: Abstract concerns. Journal of the American Society for Information Science, 46 (9):717–718, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Windsor:1999:IRI

[Win99] Donald A. Windsor. Industrial roots of information science. Journal of the American Society for Information Science, 50(12): 1064–1065, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wise:1999:REA

[Wis99] James A. Wise. Research: The ecological approach to text visualization. Journal of the American Society for Information Science, 50(13):1224–1233, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wise:2000:RPA

[Wis00] Michael J. Wise. Research: Protein annotators’ assistant: a novel application of information retrieval techniques. Journal of the American Society for Information Science, 51(12):1131–1136, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wittman:1980:LEC

[Wit80] Allan Wittman. Letters to the Editor: Copyright and information transfer. Journal of the American Society for Information Science, 31(5):379, September 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Withington:1981:CTS

[Wit81] Frederic G. Withington. Computer technology: State of the art. Journal of the American Society for Information Science, 32(2): 124–130, March 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wittman:1987:LER

[Wit87] Allan Wittman. Letters to the Editor: Regarding erratum. Journal of the American Society for Information Science, 38(2):142, March 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wittorf:1996:BRBa

[Wit96a] Robert Wittorf. Book review: Learning networks: a field guide to teaching and learning online. Journal of the American Society for Information Science, 47(10):797–798, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wittorf:1996:BRBb

[Wit96b] Robert Wittorf. Book review: Learning networks: a field guide to teaching and learning online. Journal of the American Society for Information Science, 47(10):797–798, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1975:OPT

[WK75] James G. Williams and Chai Kim. Opinion paper. On theory development in information science. Journal of the American Society for Information Science, 26(1):1–9, January 1975. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic). See reply [Art75a].

Walsh:2000:CMC

[WKMG00] John P. Walsh, Stephanie Kucker, Nancy G. Maloney, and Shaul Gabbay. Connecting minds: Computer-mediated communication and scientific work. Journal of the American Society for Information Science, 51(14):1295–1305, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wanger:1980:NLD

[WL80] Judith Wanger and Ruth N. Landau. Nonbibliographic on-line data base services. Journal of the American Society for Information Science, 31(3):171–180, May 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1981:BCL

[WL81] Martha E. Williams and Laurence Lannom. Brief communications: Lack of standardization of the journal title data element in databases. Journal of the American Society for Information Science, 32(3):229–233, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wong:2000:DKN

[WLC00] Man Leung Wong, Kwong Sak Leung, and Jack C. Y. Cheng. Discovering knowledge from noisy databases using genetic programming. Journal of the American Society for Information Science, 51(9):870–881, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wellisch:1978:LEB

[WM78] Hans H. Wellisch and Gharles T. Meadow. Letters to the Editor: The JASIS KWOC index. Journal of the American Society for Information Science, 29(5):260–261, September 1978. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:1998:RVD

[WM98] Howard D. White and Katherine W. McCain. Research: Visualizing a discipline: an author co-citation analysis of information science, 1972–1995. Journal of the American Society for Information Science, 49(4):327–355, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:2000:MBC

[WM00a] Howard D. White and Katherine W. McCain. In memory of Belver C. Griffith. Journal of the American Society for Information Science, 51(10):959–962, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

White:2000:LER

[WM00b] Howard D. White and Katherine W. McCain. Letters to the Editor: Rejoinder: Authors of information science. Journal of the American Society for Information Science, 51(9):882–883, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wilson:1988:PLI

[WMME88] Marjorie P. Wilson, Robert J. McCarter, Jr., Alan B. McKay, and Rosnec Estime. Perspectives on … Integrated Academic Information Management Systems (IAIMS): The management of change: Lessons learned from the IAIMS experience. Journal of the American Society for Information Science, 39(2):113–117, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wright:1999:RHC

[WNAR99] Lawrence W. Wright, Holly K. Grossetta Nardini, Alan R. Aronson, and Thomas C. Rindflesch. Research: Hierarchical concept indexing of full-text documents in the Unified Medical Language System(R) Information Sources Map. Journal of the American Society for Information Science, 50(6):514–523, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wolfe:1971:SEP

[Wol71] Theodore Wolfe. Suggestions for exploiting the potential of online remote access information retrieval and display systems. Journal of the American Society for Information Science, 22(3): 149–152, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wolek:1972:PIC

[Wol72] Francis W. Wolek. Preparation for interpersonal communication. Journal of the American Society for Information Science, 23(1): 3–10, January 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wolek:1986:RMS

[Wol86] Francis W. Wolek. Research: Managerial support and the use of information services. Journal of the American Society for Information Science, 37(3):153–157, May 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wolfe:1987:PLF

[Wol87] Molly A. Wolfe. Perspectives on … the Federal Government and health information: Patterns, impact, expectations: Section V. Public/ private information providers. Evolution of contractorsupported health information activities. Journal of the American Society for Information Science, 38(1):71–75, January 1987. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wolfram:1996:RIR

[Wol96] Dietmar Wolfram. Research: Inter-record linkage structure in a hypertext bibliographic retrieval system. Journal of the American Society for Information Science, 47(10):765–774, October 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wooster:1970:IAC

[Woo70] Harold Wooster. An information analysis center effectiveness chrestomathy. Journal of the American Society for Information Science, 21(2):149–159, March 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wooster:1976:ENL

[Woo76a] Harold Wooster. An experiment in networking: The LHNCBC experimental CAI network, 1971–1975. Journal of the American Society for Information Science, 27(5):329–338, September 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wooster:1976:LEC

[Woo76b] Harold Wooster. Letters to the Editor: The case rests. Journal of the American Society for Information Science, 27(4):264, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Woodward:1977:RRI

[Woo77a] Anthony M. Woodward. The roles of reviews in information transfer. Journal of the American Society for Information Science, 28(3):175–180, May 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wooster:1977:BRB

[Woo77b] Harold Wooster. Book review: The SCATT Report: Designing a National Scientific and Technical Communication System. Russell L. Ackoff, Thomas A. Cowan, Wladimir M. Sachs, Maybeth L. Meditz, Peter Davis, James C. Emergy. and Martin C. J. Elton. Philadelphia, PA: University of Pennsylvania Press, 173 pp. (1976). $12.00 (cloth). Journal of the American Society for Information Science, 28(2):127, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wood:1984:ERL

[Woo84] Frances E. Wood. European research letter: Teaching online information retrieval in united kingdom library schools. Journal of the American Society for Information Science, 35(1):53–55, January 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Woodward:1987:OPC

[Woo87a] Diana Woodward. Opinion papers: Commentary on F. W. Horton’s “An Information Bill of Rights”. Journal of the American Society for Information Science, 38(2):130–132, March 1987. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wooster:1987:HNS

[Woo87b] Harold Wooster. Historical note: Shining Palaces, Shifting Sands: National information systems. Journal of the American Society for Information Science, 38(5):321–335, September 1987. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Woodward:1988:BRB

[Woo88a] Diana Woodward. Book review: Protecting privacy in two-way electronic services. Journal of the American Society for Information Science, 39(6):434–435, November 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Woodward:1988:PLE

[Woo88b] Diana Woodward. Perspectives on … education of the information professional: New dimensions, new directions: Drexel University College of Information Studies: Evolving programs, new connections. Journal of the American Society for Information Science, 39(5):334–336, September 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wooster:1990:BRB

[Woo90] Harold Wooster. Book review: A history of information science 1945–1985. Journal of the American Society for Information Science, 41(8):611–612, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Worthen:1973:EPC

[Wor73] Dennis B. Worthen. The epidemic process and the contagion model. Journal of the American Society for Information Science, 24(5):343–346, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Woodruff:1994:GAG

[WP94] Allison Gyle Woodruff and Christian Plaunt. GIPSY: Automated geographic indexing of text documents. Journal of the American Society for Information Science, 45(9):645–655, October 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weaver:1970:ASR

[WRE^{+}70] Barbara N. Weaver, Albert B. Ruffner, Claire Ellis, Helen Anderson, and Samuel X. Delorenzo. An automated storage and retrieval system for clinical pharmaceutical research documents. Journal of the American Society for Information Science, 21(6): 415–421, November 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wright:1971:BRB

[Wri71a] Keith Wright. Book review: 1/71-2R Library Systems Analysis Guidelines. 1970 Edward A. Chapman, Paul L. St. Pierre, and John Lubans, Wiley-Interscience, New York. 226 pp. Journal of the American Society for Information Science, 22(1):67, January 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wright:1971:BCA

[Wri71b] Keith C. Wright. Brief communications: Assured Failure in Information Center Operations. Journal of the American Society for Information Science, 22(4):294–295, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wright:1972:BCA

[Wri72] William R. Wright. Brief communications: An algorithm for computing correlation matrices. Journal of the American Society for Information Science, 23(2):130–131, March 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wright:1973:SSI

[Wri73] Kieth Wright. Social science information characteristics with particular reference to the educational resources information centers (ERIC). Journal of the American Society for Information Science, 24(3):193–204, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wright:1979:WWG

[Wri79] H. Curtis Wright. The wrong way to go. Journal of the American Society for Information Science, 30(2):67–76, March 1979. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wallace:1971:RFB

[WS71] L. Wallace and Claire K. Schultz. Robert A. Fairthorne — a biographical sketch. Journal of the American Society for Information Science, 22(3):161–165, May 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Whitlock:1975:LE

[WS75] Leigh S. Whitlock and Charles W. Sargent. Letters to the Editors. Journal of the American Society for Information Science, 26(1): 68, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wang:1998:RCM

[WS98] Peiling Wang and Dagobert Soergel. Research: A cognitive model of document use during a research project. Study I. Document selection. Journal of the American Society for Information Science, 49(2):115–133, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Watters:1998:REN

[WSB98] C. R. Watters, M. A. Shepherd, and F. J. Burkowski. Research: Electronic news delivery project. Journal of the American Society for Information Science, 49(2):134–150, February 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Watters:1994:RTO

[WSQ94] Carolyn Watters, Michael A. Shepherd, and Liwen Qiu. Research: Task-oriented access to data files: an evaluation. Journal of the American Society for Information Science, 45(4):251–262, May 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wenger:1979:MEA

[WSS79] Charles B. Wenger, Christine B. Sweet, and Helen J. Stiles. Monograph evaluation for acquisitions in a large research library. Journal of the American Society for Information Science, 30(2): 88–92, March 1979. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wu:1993:RCT

[WT93] Zimin Wu and Gwyneth Tseng. Research: Chinese text segmentation for text retrieval: Achievements and problems. Journal of the American Society for Information Science, 44(9):532–542, October 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Weise:1994:PLR

[WT94] Frieda O. Weise and M. J. Tooey. Perspectives on … the redesign/reengineering of an information services division of a major health science institution: Designing a new facility for information services. Journal of the American Society for Information Science, 45(5):342–344, June 1994. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wu:1995:RAA

[WT95] Zimin Wu and Gwyneth Tseng. Research: ACTS: an automatic Chinese text segmentation system for full text retrieval. Journal of the American Society for Information Science, 46(2):83–96, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wu:1998:RIE

[Wu98] Xindong Wu. Rule induction with extension matrices. Journal of the American Society for Information Science, 49(5):435–454, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wuest:1997:BRB

[Wue97] Ruth Wuest. Book review: Digital image and audio communications: Toward a global information infrastructure. Journal of the American Society for Information Science, 48(3):280–281, March 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wallace:1997:BCD

[WV97] Danny P. Wallace and Connie Van Fleet. Brief communications: Describing technological paradigm transitions: a methodological exploration. Journal of the American Society for Information Science, 48(2):184–187, February 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wang:1985:RTI

[WVE85] Yih-Chen Wang, James Vandendorpe, and Martha Evens. Relational thesauri in information retrieval. Journal of the American Society for Information Science, 36(1):15–27, January 1985. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Windsor:1973:CLI

[WW73] Donald A. Windsor and Diane M. Windsor. Citation of the literature by information scientists in their own publications. Journal of the American Society for Information Science, 24(5):377–381, September 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wang:1999:RCM

[WW99] Peiling Wang and Marilyn Domas White. Research: A cognitive model of document use during a research project. Study II. Decisions at the reading and citing stages. Journal of the American Society for Information Science, 50(2):98–114, ???? 1999. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Watters:2000:RRN

[WW00] Carolyn Watters and Hong Wang. Research: Rating news documents for similarity. Journal of the American Society for Information Science, 51(9):793–804, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Williams:1997:LBH

[WWB97] Robert V. Williams, Laird Whitmire, and Colleen Bradley. Literature: Bibliography of the History of Information Science in North America, 1900–1995. Journal of the American Society for Information Science, 48(4):373–379, April 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ward:2000:RNO

[WWSR00] Robert Ward, Gary Wamsley, Aaron Schroeder, and David B. Robins. Research: Network organizational development in the public sector: a case study of the federal emergency management administration (FEMA). Journal of the American Society for Information Science, 51(11):1018–1032, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Whittemore:1973:GCD

[WY73] Bruce J. Whittemore and M. C. Yovits. A generalized conceptual development for the analysis and flow of information. Journal of the American Society for Information Science, 24(3):221–231, May 1973. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wong:1990:RGB

[WY90a] S. K. M. Wong and Y. Y. Yao. Research: A generalized binary probabilistic independence model. Journal of the American Society for Information Science, 41(5):324–329, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wong:1990:RQF

[WY90b] S. K. M. Wong and Y. Y. Yao. Research: Query formulation in linear retrieval models. Journal of the American Society for Information Science, 41(5):334–341, July 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wong:1992:RIT

[WY92] S. K. M. Wong and Y. Y. Yao. Research: An informationtheoretic measure of term specificity. Journal of the American Society for Information Science, 43(1):54–61, January 1992. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wong:1993:RPM

[WY93] S. K. M. Wong and Y. Y. Yao. Research: A probabilistic method for computing term-by-term relationships. Journal of the American Society for Information Science, 44(8):431–439, September 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wyllys:1987:LEH

[Wyl87] Ronald E. Wyllys. Letters to the Editor: Historical perspective articles. Journal of the American Society for Information Science, 38(5):387, September 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Wong:1991:REA

[WYSB91] S. K. M. Wong, Y. Y. Yao, G. Salton, and C. Buckley. Research: Evaluation of an adaptive linear model. Journal of the American Society for Information Science, 42(10):723–730, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Xiaorong:1992:RSP

[Xia92] Fang Xiaorong. Research: A study of the problem of the aging of books in university libraries: Strategies for countering its effects.

Journal of the American Society for Information Science, 43(7): 501–505, August 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Xie:2000:SII

[Xie00] Hong (Iris) Xie. Shifts of interactive intentions and informationseeking strategies in interactive information retrieval. Journal of the American Society for Information Science, 51(9):841–857, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yager:1987:BCN

[Yag87] Ronald R. Yager. Brief communications: A note on weighted queries in information retrieval systems. Journal of the American Society for Information Science, 38(1):23–24, January 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yannakoudakis:1990:RMC

[YAH90] E. J. Yannakoudakis, F. H. Ayres, and J. A. W. Huggill. Research: Matching of citations between non-standardized databases. Journal of the American Society for Information Science, 41(8):599–610, December 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yang:1990:RLI

[Yan90] Qi Quan Yang. Research: The law of information inconservation. Journal of the American Society for Information Science, 41(6): 418, September 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yang:1997:BCQ

[Yan97] Shu Ching Yang. Brief communications: Qualitative exploration of learners’ information-seeking processes using Perseus Hypermedia System. Journal of the American Society for Information Science, 48(7):667–669, July 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yao:1995:RMR

[Yao95] Y. Y. Yao. Research: Measuring retrieval effectiveness based on user preference of documents. Journal of the American Society for Information Science, 46(2):133–145, March 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yaru:1997:BCS

[Yar97] Dang Yaru. Brief communications: Structural modeling of network systems in citation analysis. Journal of the American Society for Information Science, 48(10):946–952, October 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yeh:2000:CKM

[YCO00] Jian-Hua Yeh, Jia-Yang Chang, and Yen-Jen Oyang. Content and knowledge management in a digital library and museum. Journal of the American Society for Information Science, 51(4):371–379, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yovits:1987:RED

[YdKKM87] M. C. Yovits, A. de Korvin, R. Kleyle, and M. Mascarenhas. Research: External documentation and its quantitative relationship to the internal information state of a decision maker: the information profile. Journal of the American Society for Information Science, 38(6):405–419, November 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yee:1991:RSD

[Yee91] Martha M. Yee. Research: System design and cataloging meet the user: User interfaces to online public access catalogs. Journal of the American Society for Information Science, 42(2):78–98, March 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yerkey:1971:BCO

[Yer71a] A. Neil Yerkey. Brief communications: Occurrence of letters in engineering periodical titles. Journal of the American Society for Information Science, 22(4):290–292, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yerkey:1971:BCU

[Yer71b] A. Neil Yerkey. Brief communications: UNISIST: Growing interest in a worldwide science information system. Journal of the American Society for Information Science, 22(4):288–289, July 1971. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yerkey:1973:MIS

[Yer73] A. Neil Yerkey. Models of index searching and retrieval effectiveness of keyword-in-context indexes. Journal of the American Society for Information Science, 24(4):282–286, July 1973. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yerkey:1983:CAR

[Yer83] A. Neil Yerkey. A cluster analysis of retrieval patterns among bibliogrpahic databases. Journal of the American Society for Information Science, 34(5):350–355, September 1983. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yerkey:1991:LEI

[Yer91] A. Neil Yerkey. Letters to the Editor: Information as thing. Journal of the American Society for Information Science, 42(10): 758, December 1991. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yovits:1985:EAI

[YF85] M. C. Yovits and C. R. Foulk. Experiments and analysis of information use and value in a decision-making context. Journal of the American Society for Information Science, 36(2):63–81, March 1985. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yovits:1981:IFAa

[YFR81a] M. C. Yovits, C. R. Foulk, and L. L. Rose. Information flow and analysis: Theory, simulation, and experiments. I. Basic theoretical and conceptual development. Journal of the American Society for Information Science, 32(3):187–202, May 1981. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yovits:1981:IFAb

[YFR81b] M. C. Yovits, C. R. Foulk, and L. L. Rose. Information flow and analysis: Theory, simulation, and experiments. II. Simulation, examples, and results. Journal of the American Society for Information Science, 32(3):203–210, May 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yovits:1981:IFAc

[YFR81c] M. C. Yovits, C. R. Foulk, and L. L. Rose. Information flow and analysis: Theory, simulation, and experiments. III. Preliminary experiments and analysis. Journal of the American Society for Information Science, 32(4):243–248, July 1981. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yerkey:1990:RSL

[YG90] Neil Yerkey and Maryruth Glogowski. Research: Scatter of library and information science topics among bibliographic databases. Journal of the American Society for Information Science, 41(4):245–253, June 1990. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Ying:1995:BRB

[Yin95] Hao Ying. Book review: Essentials of fuzzy modeling and control. Journal of the American Society for Information Science, 46(10): 791–792, December 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yovits:1993:RAD

[YK93] Marshall C. Yovits and Robert M. Kleyle. Research: The average decision maker and its properties utilizing the generalized information system model. Journal of the American Society for Information Science, 44(6):352–363, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yang:2000:CBD

[YLYY00] Christopher C. Yang, Johnny W. K. Luk, Stanley K. Yung, and Jerome Yen. Combination and boundary detection approaches on Chinese indexing. Journal of the American Society for Information Science, 51(4):340–351, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yaghmai:1984:EST

[YM84] N. Shahla Yaghmai and Jacqueline A. Maxin. Expert systems: a tutorial. Journal of the American Society for Information Science, 35(5):297–305, September 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yuan:1999:RSU

[YM99] Weijing Yuan and Charles T. Meadow. Research: A study of the use of variables in information retrieval user studies. Journal of the American Society for Information Science, 50(2):140–150, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yoon:1994:RPC

[Yoo94] Lanju Lee Yoon. Research: The performance of cited references as an approach to information retrieval. Journal of the American Society for Information Science, 45(5):287–299, June 1994. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Young:1975:BRB

[You75] Arthur P. Young. Book review: Research Libraries and Technology. Herman H. Fussler. Chicago: University of Chicago Press. 91 pp. (1973). Journal of the American Society for Information Science, 26(1):67, January 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yovits:1970:IST

[Yov70] M. C. Yovits. Information science: Toward the development of a true scientific discipline. Journal of the American Society for Information Science, 21(4):268, July 1970. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yu:1977:SPM

[YR77] Clement T. Yu and Vijay V. Raghavan. Single-pass method for determining the semantic relationships between terms. Journal of the American Society for Information Science, 28(6):345–354, November 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yager:1998:RFD

[YR98] Ronald R. Yager and Alexander Rybalov. Research: On the fusion of documents from multiple collection information retrieval systems. Journal of the American Society for Information Science, 49(13):1177–1184, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Young:1993:RGF

[YS93] Degi Young and Ben Shneiderman. Research: A graphical filter/ flow representation of Boolean queries: a prototype implementation and evaluation. Journal of the American Society for Information Science, 44(6):327–339, July 1993. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yu:1974:CAB

[Yu74] Clement T. Yu. A clustering algorithm based on user queries. Journal of the American Society for Information Science, 25(4): 218–226, July 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yu:1976:STC

[Yu76] C. T. Yu. The stability of two common matching functions in classification with respect to a proposed measure. Journal of the American Society for Information Science, 27(4):248–255, July 1976. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yuan:1997:REU

[Yua97] Weijing Yuan. Research: End-user searching behavior in information retrieval: a longitudinal study. Journal of the American Society for Information Science, 48(3):218–234, March 1997. CO-DEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Yang:1996:RUC

[YW96] Yiming Yang and John Wilbur. Research: Using corpus statistics to remove redundant words in text categorization. Journal of the American Society for Information Science, 47(5):357–369, May 1996. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zais:1977:EMA

[Zai77] Harriet W. Zais. Economic modeling: an aid to the pricing of information services. Journal of the American Society for Information Science, 28(2):89–95, March 1977. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zaltman:1974:NII

[Zal74] Gerald Zaltman. A note on an international invisible college for information exchange. Journal of the American Society for Information Science, 25(2):113–117, March 1974. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zamora:1980:ADC

[Zam80] Antonio Zamora. Automatic detection and correction of spelling errors in a large data base. Journal of the American Society for Information Science, 31(1):51–57, January 1980. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zarember:1982:LED

[Zar82] Irving Zarember. Letter to the Editors: Database comparison. Journal of the American Society for Information Science, 33(1): 60, January 1982. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zahir:1992:ROE

[ZC92] Sajjad Zahir and Chew Lik Chang. Research: Online-expert: an expert system for online database selection. Journal of the American Society for Information Science, 43(5):340–357, June 1992. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhu:2000:RVG

[ZC00] Bin Zhu and Hsinchun Chen. Research: Validating a geographical image retrieval system. Journal of the American Society for Information Science, 51(7):625–634, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zeng:1999:RME

[Zen99] Marcia Lei Zeng. Research: Metadata elements for object description and representation: a case report from a digitized historical fashion collection project. Journal of the American Society for Information Science, 50(13):1193–1208, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhang:1998:DRA

[Zha98] Jiajie Zhang. A distributed representation approach to group problem solving. Journal of the American Society for Information Science, 49(9):801–809, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhang:2000:RUI

[Zha00] Yin Zhang. Research: Using the Internet for survey research: a case study. Journal of the American Society for Information Science, 51(1):57–68, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhiyou:1997:LET

[Zhi97] Zhou Zhiyou. Letters to the Editor: The term “information science”. Journal of the American Society for Information Science, 48(12):1153, December 1997. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zins:2000:SSS

[Zin00] Chaim Zins. Success, a structured search strategy: Rationale, principles, and implications. Journal of the American Society for Information Science, 51(13):1232–1247, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zaltman:1972:DTS

[ZK72] Gerald Zaltman and Barbara Marie Köhler. The dissemination of task and socioemotional information in an international community of scientists. Journal of the American Society for Information Science, 23(4):225–236, July 1972. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhang:1999:RDA

[ZK99a] Jin Zhang and Robert R. Korfhage. Research: A distance and angle similarity measure method. Journal of the American Society for Information Science, 50(9):772–778, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhang:1999:RDD

[ZK99b] Jin Zhang and Robert R. Korfhage. Research: DARE: Distance and angle retrieval environment: a tale of the two measures. Journal of the American Society for Information Science, 50(9): 779–787, ???? 1999. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zmud:1975:SCR

[Zmu75] Robert W. Zmud. Selecting computer resources for inclusion within a pricing system. Journal of the American Society for Information Science, 26(6):346–348, November 1975. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zoellick:1988:PLC

[Zoe88] Bill Zoellick. Perspectives on … CD-ROM for information storage and retrieval: CD-ROM software architecture to promote interchangeability. Journal of the American Society for Information Science, 39(1):47–53, January 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zitt:1998:TNT

[ZPB98] Michel Zitt, François Perrot, and Rémi Barré. The transition from “national” to “transnational” model and related measures of countries’ performance. Journal of the American Society for Information Science, 49(1):30–42, ???? 1998. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zurkowski:1984:IAI

[Zur84] Paul G. Zurkowski. Integrating America’s infostructure. Journal of the American Society for Information Science, 35(3):170–178, May 1984. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zimmerman:1988:PLI

[ZvBR88] John L. Zimmerman, Jan H. van Bemmel, and Otto Rienhoff. Perspectives on … integrated academic information management systems (IAIMS): Medical informatics education. Journal of the American Society for Information Science, 39(2):138–141, March 1988. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zhang:2000:SDT

[ZvD00] Ping Zhang and Gisela M. von Dran. Satisfiers and dissatisfiers: a two-factor model for website design and evaluation. Journal of the American Society for Information Science, 51(14):1253–1268, ???? 2000. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zvieli:1986:BRB

[Zvi86] Arie Zvieli. Book review: Query processing in database systems. Journal of the American Society for Information Science, 37(2): 89, March 1986. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zvieli:1987:BRB

[Zvi87] Arie Zvieli. Book review: Languages for automation. Journal of the American Society for Information Science, 38(3):219–220, May 1987. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).

Zweizig:1995:BRB

[Zwe95] Douglas L. Zweizig. Book review: If you want to evaluate your library. (2nd ed.). Journal of the American Society for Information Science, 46(9):710, October 1995. CODEN AISJB6. ISSN 0002-8231 (print), 1097-4571 (electronic).